A Complete Bibliography of Publications in the International Journal of Computer Applications

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

24 January 2025 Version 1.00

Title word cross-reference

−ϕ(ξ) [KZ16b]. {0; 1}^{n} [Sin13]. {2^{(2n+1)} + 1, 2^{(2n+1)}, 2^{(2n+1)} − 1} [MB17]. 2k [MS10a]. 2n [KAR10]. 2 × 1 [IEG14]. 6 × 20 [CSN15]. 8 × 8 [RJS^{+}13]. ^{2} [Hos12]. ^{[}x] [AS13d]. C [DH17, Hab13a, PRP15, SP11a]. F [HAA16, MHE13]. G [ACK12, Ven14b]. gλ [MP15]. K [AlS24a, AlS24b, AA11a, BS21a, BM16a, Bef18, BRS16, BT14b, DAB12, DUP13, DS12c, GNN16, IEME12, JH18, KB16a, MJ15, NDC10, OAR16, OKS18, PA24, PK15b, PZN21, SS14a, SP12, SYR13, ZRP16, AS15a, KAP15, KH20, MHE13, RPV10, SK13b]. L [Ble18, Mus13]. M [DM14, RGCGCMFJ13, MHE13, PS12b]. μ [PBK14]. N [MAt12, RGCGCMFJ13, XL21, GHBN12, MHE13]. N × N [MFM12]. P [PA24]. Q [Nta14, BM16b, GG12, KBA10]. R [DL19, YK12]. S [Wu21, Wu23a]. T [DR15b, EJC16, Wu23b, OKS14]. τ∗ [MP15]. V_{n} [RAA15].

-Anonymity [BM16a, PA24]. -Anti-Fuzzy [AS13b]. -Arithmetic [RAA15]. -ary [DM14, XL21]. -Bernstein [KBA10]. -bit [GHBN12]. -Closed [MP15]. -consecutive-at [MHE13]. -CSRmesh [ZYW^{+}18]. -factor [Nta14]. -fuzzy [Mus13]. -Grams [MAt12]. -Jobs [RGCGCMFJ13]. -LeaSel [Ven14b]. -Machines [RGCGCMFJ13]. -matrices [BM16b]. -Mean [DS12c]. -Means [SK13b, AS15a, RPV10, AlS24b, BS21a, BT14b, DUP13, GNN16, NDC10, OAR16, PK15b, SP11a, SP12, ZRP16, AA11a, DH17, DAB12, Hab13a, MJ15, PRP15, SYR13]. -Metric [ACK12]. -Modes [PZN21]. -Nearest [KH20, AlS24a, Bef18, IEME12, KB16a, OKS18]. -NN [JH18]. -Norm [DR15b]. -Order [OKS14]. -out-of- [MHE13]. -peak [DL19]. -QPSK [SSU10]. -Quadratic [Ble18]. -S [EJC16, Wu23b]. -Sensitive [PA24]. -Systems [Wu23a]. -Trees [YK12]. -Tuples [BM16a]. -type [Wu21]. -Uniform [KAP15]. -Unit [MAA13].

.jpeg [NSS17]. .png [NSS17]. .png/.tif/.jpeg [NSS17]. .tif [NSS17].

/G/1 [AS13d].

0/1 [YMAS17].

1 [ASG10, AS13d, HSK14, NM14, RBB11]. 1-Bit [SSK13a]. 1-Safe [Sin13]. 13 [NYS15]. 130nm [KKS14]. 140 [SVB23]. 140-4G [SVB23]. 16-bit [SLK^{+}13]. 19 [KSA20].

2 [EMF14, HJBJ11, HSK14, LK15, RBB11, SPC13, SK18, SS12e]. 2-Axis [SM14a]. 2-D [SPC13]. 2-tier [DM14]. 2.0 [ESB13, HMMN13, YMA13, ZKN10]. 2.5D [Sha13]. 200-Gb [MSS13]. 200-Gb/s [MSS13]. 2009 [Kar11]. 2D [BA11, BI13, DP12, KEV14, MBKH16, Rab14, RHU12, SAKS24, VPGF10]. 2D-3D [Rab14]. 2DOFLQG [TArM14].

3 [RSS^{+}11, Sal15, SGKS12]. 3-Axis [SZ15]. 3-Connected [KIH15]. 3-D [Sal15]. 3-Dimensional [AAEK18, NAH11]. 3-Disjoint [RSS^{+}11]. 3-Level [SS15a]. 3-Regular [KIH15]. 31A-M-Frozen-in-Idle-State [AKK11]. 32-Bit [MFM12]. 328P [LSM16]. 3D [AJB13, BSAb12, BS12b, GYP^{+}15, HF16, HO15b, HH13, JM15, KS15a, KMT^{+}24, MBKH16, Rab14, RVGP20, RGS15, SA15a, SM13a, Sha13, SRK10, SM19, Sun19, TSRM10, VPGF10, ZEF17]. 3D-Handwritten [KS15a]. 3D-Rebuild [Sun19]. 3G [ATA14, EBO13, EB13].

4 [DV14]. 4A [VB13]. 4B [PB23]. 4EGSOR [SS12c]. 4EGSOR9L [SS12c]. 4G [AAA^{+}17b, Sul17, SVB23, Tha14]. 4T4R [MAM19]. 4th [PT15]. 4U [BA12b].

5 [PUW13].

64-bit [DV14]. 65nm [SAR16b]. 6lo [RKBB14]. 6T [RSK15a].

7-Bit [FPS16].

8 [NM14]. 8-Bit [AKK11, ARS13]. 802.11b [DV16, GK12a, KG15b]. 802.11p [AM15]. 802.15.4 [DKS14, RS13a]. 802.15.4/ZigBee [DKS14]. 802.16 [Kum14a, RKG10]. 802.16e [GDD13, KDTD11, SSG11]. 802.1X [HR21, KA21]. 802.22 [JM14, MJM16]. 8GB [PB23].

9 [MR13a, SOME16]. 90nm [RJS^{+}13].

A-CBF [PA13]. A-r-Star [OHT13]. A3MR [AB20]. A3MR-NEMO [AB20]. AAA [AB20]. AADL [BB10, BBC12b]. Abacus [MB16]. Abadan [TSS12]. Abandoned [PP17b]. ABC [AD24, AAA17a, KC14b, MBL15a, SCE14]. Ability [DUP13, LcCm13]. Abled [PR15]. Abnormality [AS12e, Las10, TV13]. Absolute [DSJ11b, II16]. Absorbing [DGJ12]. Abstractive [SS17a]. AC [Ena12, Iss13, MAI15]. AC/DC [Ena12]. AC4LV [AMS14]. Academic [AAEA16, DUP13, DTO23, Far14a, GC24, JFJN20, KM13b, USC12]. Accelerated [FR14, SDJ13]. Accelerating [KEBF13]. Acceleration [Gha16]. Accelerator [Say22]. Accelerometer [GK15c]. Acceptance [ASS17, AAO22, OAA24, Pra14]. Accepting [KA14]. Access [Abu14, ABF14, AkAA14, AmM12, Aq14a, ASN12, MKOS14, CS16a, CS18, DT14, DV16, EAa14, ETG17, ES12b, GHD17, HSP10, IHH14, sNIALY13, JS18a, KIA15, KSlAM12, KN12c, LK15, MMG16, MK13d, MM14b, NAD19, NMAB14, PA14, Pou13, RK21, SNP15, SOO15, TGL13, TAM14a]. Access-Aware [ETG17]. Accessibility [AOK18, SP15e]. Accessing [SMM11]. ACCFLA [TAM14a]. Accident [EAM14, MH13b, NKM12, PZN21, Rah12]. Accidental [SAAM17]. Accidents [AZY18a]. according [AS18c, EBO13, JM14]. Accounting [OAW12]. Accra [GAA17, OAG17]. Accumulate [GSN14]. Accumulation [RMD10]. Accuracy [AA16b, Ask14, KR17, REA12, SY15, Zon17]. Accurate [Abb12a, AWA12, AHBAR^{+}11, AMK12, Dur14, NM20, TRR16]. Acetate [PN10]. Achieve [DKS14, KMS12, NJB16, PB14]. Achievement [AJ18, AAEA16]. Achieving [EEKG17]. Acknowledgement [SN15c]. Acknowledgment [NKR14]. ACO [JSY14, NS13b, SAKS24, VST17]. Acoustic [DK18, SET15, TS17]. Acquisition [Ba13, DDN14, Mun14, SS10b, ZCZ24, RPMH24]. Acreage [TK^{+}17]. Across [MB15c, VMA23, GKG16]. Action [DL21, EEM12]. Actionable [KS12a]. Activation [TRM23]. Active [AMA15, AAA17a, ASM13, AMK12, BK17, CHG12, HB13, JPP12, KVRS14, KKR^{+}14, NEA17, PMP12, PMH10, SKM16, SDS20, UR12, VRS12]. Activeness [GL16]. Activities [AIG12, SS15b]. Activity [FYM11, GEAE18, Gho15, GK15c, JH18, KBEA18, KDC12, KH13, MMHA11, PM15, RKGK15]. Activity-based [RKGK15]. Actual [HE15, TA10]. Actuating [SPXT17]. Actuator [Cha17, FAK14, KTN12, Nad14]. ACWMF [ST14]. Ad [AMG13, ASNS10, AA12f, AV10, ASD16, AJC14, AS15d, BD10a, BGS11, BD10b, BS12c, DB13b, ES13, GE17, KM16b, KA12e, KBT^{+}21, Mah19, Meg10, NV12, PP13a, PSS11, RD17a, SWP19, SN15c, SSMP13, SF15b, TAM14b, VMT13, YRV17]. Ad-Hoc [DB13b, GE17, NV12, ASNS10, BS12c, ES13, KA12e, PSS11, SSMP13, SF15b, SWP19]. Adaboost [BA15b, Abb12a]. Adaboost-based [BA15b]. Adaptability [OAR16, RA12a]. Adaptation [AK14, FAZ13, KT10, RO11]. Adapted [Kor23]. Adaption [CS12]. Adaptive [Abd18, AAY21, ATM12, ASY17, ASS10, AhAT13, AsS13, AAoS14, AERS11, ASD16, AHK12, AVS13, Bad13b, BECM11, BBD13, BSP11, BWH15, CTG10, Cha14, CWC15, CZW15a, DM10, ElZA13, El 20, EBR12, EJC16, FRR11a, FAK14, GEAE18, GE17, GKM^{+}18, GSA11, HJT14, HKY15, Ibr18, ITK21, JHP12, JT15, JTA19, JTI19, KM17, KM15a, KJK12, KSL^{+}18, KPB12, KA13a, KVKR16, LBA16, MBMB12, Nad14, NEA17, NM13, NH13, Ngu24a, OAA16, ONJ13, Ori14, OSA14, OJ11, PCW10, PPP11, QCERM15, QJY11, RT13, Rad13, RAJ^{+}12, RS12, RS14b, SPP13, SEb13, SDZ15, SGM^{+}11, SAR18, SHD13, SM13c, SDS15, SP18c, TBR15, TOMGBSS13, TNC10, UR16, VK11, WDD16, ZRZ11, ZBI14, ZZ11, UA14]. Adaptive-Transmission-Power [ASD16]. Adaptivity [RA12a]. Add [EAa14]. Add/Drop [EAa14]. Addaptive [Anh14a]. Adder [KKS14, MKWK15, SSK13a, SLK^{+}13]. Addiction [Far24]. Additive [BA23, DD13]. Adenoids [KN12a]. Adhoc [HS15b, KP10a, VG16]. Adiabatic [SN14a, SSS13c]. Adjacency [SSV11]. Adjacent [TKC^{+}13]. Adjectives [TT11]. Adjustable [MMA13]. Adjusted [Jas13]. Adjustment [GEB12, KT14]. Adleman [IPR15]. Administration [GMRK12, LOAO23, SAM14a]. Admission [LG13b, BG17, VMT13]. Adomian [SMAM15]. Adopting [AAI13, ADKsO14, ES15b, FCW14, KSS18, KMG20, NF17]. Adoption [AS18a, ATAAAF19, AnAA14, AO16, EEM12, NOO16, OK14, UAMN17]. ADPS [LHKR16]. Adsorber [IM15]. ADU [Far24]. Adult [CRBJ12]. Advance [ADFAJ15, JM16]. Advanced [AM13a, AGZ15, AAOA23, DS12b, ETG17, JBT23, KB16b, NKH12, PSM15, SG16b, YSH12]. advancements [VJ24]. Advances [HK19]. Advancing [Has24]. Advantages [MK17, SEA16]. Adversarial [Ala16, AK19, ZYL^{+}21]. Adversary [EOO22]. Adverse [ZK12]. Advertisement [DMS13]. Advertising [OSS14]. Adviser [ZAE18]. Advisory [AAA16, IGS17]. AEAR [AS24]. AECS [TVY17]. AeLF [QCERM15]. Aerial [ABL17, MOM17, Rez12, SA15a]. Aerodynamic [HA16]. Aerogel [AHZ15b]. Aerogel/Glass [AHZ15b]. Aeronautical [MC13]. Aerosols [PR12]. AES [BC22, GGM23, MZEE13]. Affect [MSM17]. Affective [BFD14]. Affects [SAA24]. affine [CZW15b]. Affluent [LCP21]. Affordable [ALO17]. Africa [IND13, SP11b]. after [NKD14, ST11]. Against [GJM16, Bat18, BH17b, EOO22, IS21, SN15b, SSS16, SS16d, YS11]. Agarwal [CKD12]. Age [AYAB18]. Aged [PR15]. Agencies [EAE19]. Agent [AAYM13, AH17a, ASS15b, BKB11, CJC19, FH13, JK14b, KP16, KS16a, KGAM10, OOE18, Ojh11, RC12, Sal18, TO13, FTLK14, KAS12]. Agents [AGL17, HA17, Hus12, MAMS12]. Agglomerative [MAMA16]. Aggregate [AC18, HKS12]. Aggregated [RSK15b]. Aggregating [SS23b]. Aggregation [FRR11b, KP16, KMMM12a, KS13a, KPAB14, NS13b, SPP11, SL11]. Aggregator [ABYB22]. Agile [AIG12, BA12a, DR15a, FCW14, MR13b, SC12, GT11]. Aging [KR11b]. Agreement [BSI12]. Agricultural [MJ14, PSS12, RÇ16, SB18a]. Agriculture [OAA24, WKG^{+}18, YS11]. Agriculturists [YS11]. Agro [MM15a]. Agro-ecosystem [MM15a]. Ahead [AJT17, JM18, RSK15b]. AHLs [ASM13]. Ahmad [SU20]. Ahmadu [UDA13]. AHP [BSS12]. AHP-QFD [BSS12]. AI [Amb24, RVGP20, Upa23]. Aid [AO13b, BML22, Bou12a, NAH11, SC19b, CM12]. Aided [HO15b, LP16, MRG19, PBVD22, SBC12]. Aided-Decision [SBC12]. Aids [Gha13, SC14a, BI12, ZRE16]. Air [ASD19, AM17b, BBC12a, EAMD14, NAhA13, NAD19, RMP17, Sal15, SM17b, SJK21, ZAE18]. Air-Gap [BBC12a]. Air-to-Ground [NAhA13]. Air2Day [ZAE18]. Airborne [MM16]. Aircraft [NVR12, dSEdZ14]. Airflow [MB15c, MBR16]. Airline [LPP12]. Airtarget [TS17]. Aklim [ESH^{+}14]. al [CKD12]. Alamouti [IEG14, PT15]. Alanine [KJA15]. Alarming [VSS16]. Alarms [FKT12, SF15a]. Aleatory [DA12b]. Alert [AsRA^{+}12, ARDS18, BKK13, DDC^{+}24, MO17, MIS12]. Alexander [HS18]. Algebra [BD13, SMSM10]. Algeria [BC12, RSB12]. Algorism [Sun19]. Algorithm [AOB12, AD24, Abb12a, AA14a, AS17, AM11, AJ13, AAS14, AA13a, AJ14, MNP11, Ae14a, Abu14, AMS13, ASN11, AS15b, AMA14a, AA12b, AA12c, AM13a, AU23, AMAHH19, AhAT13, AmAdS12, AA18b, Alh12, AA20, AYYB15, AA12e, AAA14b, AI18, AA11a, AOK14, AGL17, AA12f, AS11, AAZN21, ASCR11, ALH^{+}24, AA12g, AAD13, AAY15, ASG11, CK13, AMB11, AA17b, AAMA20, ANA11, AB13c, ur13, AMTM11, AAY17, AMH11, L13, BKK13, BWS^{+}12, BSI12, BS12b, BS21a, BM16a, BRD12, BBD13, BBA12, BS11a, BP15, BO16, BMRA12, BSM^{+}12, CW23, CD16, CM13, Cha14, Cha15, CS15b, CS15c, Cur20, CA16, DMM14a, DT14, DL19, DK13, DCH16, DH19, DN12, DSJ11b, DF15, DD13, DPT14, DBB12, DH12, DRS15, DUP13, DS12b, Dha12, DA12a, DB13b, DPB14, Dur14]. Algorithm [Eja12, ESZED15, EHSA23, ENEBES15, El 20, EBB17, EEKG17, EME13, Ela12, Ela11, EEGHS15, EMF14, ETI14, Far17, FR14, FR15, GNN16, Gha16, GMAR14, GV15, GG11, GJ14, HB12a, Hal13, HMC12a, HMY12, HG11, HF20, HAS17, Hem22, HST19, HKY15, HS10, wHsYwT12, mH14b, Hus16, Ibr18, IE12, IGS17, Isu16, JAG15a, JM18, JPJ18, JPDB14, JSK10, JM15, KK10a, KN12a, KH20, KTS10, KG15a, KRSB12, KAR10, KBTB12, KK20, KBN18, KJ12a, KS15c, KK16a, KB16a, KS16a, KDB17, KSS12, Key12, KKI16, KH12, KIH15, KS16d, KEV14, KKM18, KR10, KAAM11, KA12f, KGK13, KK14, KSK14, KVKR16, KS16e, KNS20, LGC11, LSB17, LBB11, LNM24, LGZX17, MM14a, MMN18, MM12a, Mal13a, MK17, MEAT11, Maz21, MSN12, MRR21, ME15, MM23, MMR14c, MESNH17]. Algorithm [MAM16, MBL15a, MG17b, MB15b, MTM12, MHV^{+}14, MB17, NKM11, NAK^{+}14, NBF14, NW13, NM20, NBKM13, NMB13, NK14a, Ngu24a, NS13c, NO18, NU18, ONK24, Ode14, OP18, OEA18, OON17, OAE16, OC13, OSA14, OG16, OHI15, PR10a, PD14, PDK14, PC11a, PS15b, PPM12, PP10, Pat14, RD13, RB14a, RS13b, RMS12, RAA14, RB11, Ran13, REA12, RPV10, RKEdK14, RK17, RAHH16, RKG10, SH23, SMVK11, SCE14, ST16, SD16, SCD10, SS13b, SSP13a, Sar16a, SE15, SD17, SHR14, SK16, SP12, Sha18, SAS13, She11, SSSS12, SM12a, SQ18, ST11, SGS15, SAk13b, SSS12, SM12b, SM13c, SSV14, SRS14b, SM14b, Sin14, Sin16, SE10, SV14b, SRA15, Sub10, SAP16, SC19b, SB18c, Sul17, VDS12, SV10, SSZ16, SB12c, SS13f, SSY11, SKZ22]. Algorithm [TFT11, Tah17, TK14, TL16, TS14, TS17, TB12c, TNC10, TY15a, TOHS14, TAY15, TY16, UA14, UJ13, VKB14, Ven14a, VM14a, VJD22, VST17, Vya14, WDS13, WKAL24, YAT15a, YEE14, ZTM11, ZZW24, ZZ11, BAA24, GSR21, HLZ^{+}24, SG10, TA12, USK17]. Algorithm-Based [ZZW24]. Algorithmic [KSMM15, NK15]. Algorithms [AII13, ALO17, Abu17, AO13a, AHAT16, AF15, ANA11, BJ12, BAS12a, Bef18, BMU14, BV14, BOB17, BS12c, CJ10, CS11, DXT23a, DL19, DRFF15, DAB12, DL13, DB13a, EBAS11, EEH14, EHA18, EMS19, FS11, GMM21, GEG23, GMS12, Hos12, HEN13, JA16a, JI10b, JS18b, KS14a, KVK19, KH15, KS10, KSS13, KSL^{+}18, KC11, KJ12b, KB16c, KK10c, KG15c, LSS20, MSM15, MAM16, MG19, NP13, NmBV13, Neb13, Ngu16, NMK15, OJ11, Oua12, PG15a, PKDS17, PB15, PRT^{+}16, Ps14b, MP17, Red19, Rok20, RJA23, SE12a, SE14, SSAK10, SP15b, SPC13, SAS21, SG16a, SAK18, SABA12, SAKS24, SAM14a, SSG11, SA12e, Sri12, SJV^{+}15, SA14, TR14b, TN14b, WG14]. Aliasing [PPP11]. ALIDS [Bad13b]. Alignment [ADH17, JJA14, KEBF13, LBB11, PK12a, RD13, SAK18, TERCM15]. Alignments [GK10]. Alleviation [TD21]. Allocation [ASP17, MKOS14, BJ12, BB14a, BA11, DL23, ED19, FHM16, GSB12, HYA12, HEN13, KAA13, RBF12, RSSV14, RRR11, SSO14, TR12, VT15, VMT13]. Alpha [KGK13, RA18]. Alpha-Power [RA18]. Alphabet [AS12a]. Alphabets [BO21, NS13a]. Altera [MFM12]. Alternate [PT13]. alternating [TA12]. Alternative [GGX12, IR16]. Altitude [CSN15]. ALU [ARS13, ACK14, KKS14]. Alumina [EH14]. Alzheimer [ASH16b, RMD10, VMM14]. Amalgamation [Kum14a]. Amazigh [BO21, EREM11]. Ambiguation [Eg13]. Ambiguity [SR13b]. Ambulance [BJPP15]. AMDFLT [DPB14]. AMHCD [EREM11]. AMI [NWS15]. among [Add12, Ash19, KAL15, MMH12, MAI15, SS14d]. Amplifier [AM14, BKM12, KVRS14, LGM12b, MDGS13, PBK14, VRS12, LGM12a]. Amplifiers [SMT19]. Anacardium [AGG16]. Analog [KN12b, GBM10]. Analogical [HMAA13, MRG19]. analogue [LGM12a]. Analogy [ABB20]. Analysis [AN16a, AC18, ARP17, am12a, AA16a, APU16, AR15, AMN12, Ade13a, Afr14, AJ10, AKK11, Agg17, AJSG11, AHZ14, Ahm20, ATA22, AMBM11, Al 13, ARK13, ARL16, AB12b, AV13, Ans14, ASA14, Ara15, AK11, AS13d, AZMP17, AMS14, MKOS14, BEE13, BS17, BAO16, BSS15, BMU14, Bej16, BBB16, BRR16, BKM15, BSP11, BOB17, BBL13, BOS15, BB13b, BJS14, BAK20, BG20, BT14b, ÇGDT18, Cha12, CM15a, CR12, CN10, CT10, CB16, CKD12, DFGG15, DXT23a, DS10, DDC12, DP15b, DDM^{+}23, Deb20, DB13a, DDDD12, DDN14, DM14, EH16, EK14b, El 13, EMS19, Eln14, ELDM13, EE13b, FJM16, FM13, Fen18, FLLR20, GJ15a, GA15, GK12a, Gho15, GGI14, GMS12, GS11, GGM23, GK15c, HF16, HKA15, HAAA15, HZS14, HS12, HAT16, HSP10, HM18, HS24, IS21]. Analysis [II16, IMIA11, IRT11, IAA13, JZS13, JPAK11, JI10a, JC14, JD15, JKS^{+}12, JK14a, KRR11, KHAE19, KVN10, KJ12a, KV16a, KVRS14, Ker13, KWS^{+}11, Kha13, KE14, KVK19, KNI14, KH15, KRH13, KJJ15, KGAM10, KG11, KAAM11, KKK13, Kum14a, KRR15, KK16b, KK22, KRR13, KNV10, KC15c, KG15c, LR11, LR12, Lic20, LCP21, Liu24, MKN12, MO17, MM13, MLRU12, MHE15, MRH15, MAI12, MS17, MG12, MKA17, Mal13c, MS11, MGA20, MMTS13, MRR21, MSP12, MRJN14, MRM13, MMR14c, MJM16, MSS13, MEE16, MH12b, MSR16, Nad17, NAZ17, NAhA13, NCL11, NC14, NM20, NM19, NYS15, ND11, NMK15, OMH11, OA15, OSS14, PUW13, PN12, PVR20, PPR17, PPU11, PB16, PCC12, PIB11, PDG16, PBVD22, PSS11, PMA13, PPP11, PRT^{+}16, PA19, PAH17, PMP14]. Analysis [RS13a, RSB12, RKR15, RKA13, RSSV14, RM12, RP10, RN24, RH15, RRR11, Ras16, RD17a, RMG13, RK17, RHK15, RT20, Rok20, RJA23, RJC17, SDB15a, SZ15, SZOK24, SAA24, SPBP17, SSS13a, SS12b, SOT18, SHR11, SBR18, SG14, SCW16, SS10e, SSV11, SG12a, SS13d, SG16a, SG16b, SRK10, SSG14, SM14b, SZUSNK19, SK14b, SA12d, SMT17, Sub20, SSRD12, SGJ12, SV12b, SW17, SSZ16, SB11c, TT14, TN14a, TSE17, TAM14b, TCS14, TC10, Tha11, TD16b, TOHS14, TAH13, Vad20, VR14, VB10, VK11, VM12a, VSN13, VDM22, WKC11, WAG11, WH21, WRJ23, Wu23b, YAE11, ZD15]. Analytic [Agg13, LAM14, MBR16]. Analytical [AZY18a, BGS11, BNS15, CCG12, DKP21, DDDD12, DKS14, Job17a, KT14, KMMM12a, SSO14, SNJ16, SKGP12, TJ12, HAb14]. Analytics [Amb24, PK19, SMGS17, TVS19]. Analyze [BA23, EAEE14, KA11, LcCm13, MD14a]. Analyzer [BK10b, RK14, SFAS19, TT11]. Analyzers [AG13]. Analyzing [DD20, GJP13, KAR12, RD13, SSMP13]. Anatomy [EAEE15]. Anchor [SV14b]. Ancient [SS12g]. Android [KDB17, AZ16, AYS19, AIP18, AA17b, CS15b, DBKN20, HM18, IOO20, JKTM15, KMT^{+}24, MP12b, NA18, PA23a, PA23b, RK16, RA23, USC12, Ver16, WT13, WES17]. Android-based [KDB17, PA23a, RA23, WES17]. ANFIS [BWH15, AJM^{+}15, KP15a, LLSL12, SND13, SSA16]. Angle [ESH^{+}14, EBO13, WSW^{+}13]. Angles [Abd14b, AN16b, MZ18, Rag15]. Angular [PB14, RRVP12]. Animal [OC20, OFI17]. Animation [PZ10]. Anisotropic [Sam13]. ANLRED [S13]. Anlysis [Jov15]. ANN [AO13a, AkC13, BHC12, BAA24, FA13, KS13b, MK13c, Mos11, SND13, SSG14, SMT17]. ANN-based [FA13, MK13c]. Annealing [AKS^{+}13, FD14, GMAR14, GEG23, KJ12b, TB12c, Zon17]. Annotations [Far14b, SC15a]. Annual [Ras16, Ras17]. ANODR [SS10d]. Anomalies [Cur20]. Anomalous [SF15b]. Anomaly [AAC16, AW20, AH12c, BS21a, BKSB15, DDC^{+}24, EA11, HJB^{+}14, ISM^{+}17, PM21, Vad20]. Anonymity [BM16a, MC17a, PA24]. Anonymous [AE14b]. Answer [GF14, She11]. Answering [CSM21]. Ant [AA18b, AF15, BH17b, FK12, HSMb12, HMC12a, JSY14, JK14b, KS15c, PCR12, SF15a, SK13a, SMS16, SHR14, TAH13]. Antenatal [SP11b]. Antenna [AS14a, AVRN12, AK18b, AI13a, AAA^{+}17b, AGZ15, EMA11, FR14, FSG17, JTI19, KCD13, KNGT10, LK16b, MT13b, PVR20, PA17, PB16, PK12b, PS16a, RC14, RK15, RT20, SKS13, SS13a, SA16, SSG14, SKJ15, SSK13b, TRDB24]. Antennas [AH12a, Ade13b, DJA13, JD13, JBK13, TR15]. Anti [AS13b, HHS14, Mis14, SMSM10, SZEM16, ZCZ24]. Anti-counterfeit [ZCZ24]. Anti-Money [HHS14]. Anti-Spyware [SZEM16]. Anti-Windup [Mis14]. Antialiasing [RM16]. Anticancer [FMS13]. Anticipated [RGR15]. Anticipating [Abb12b]. Antirobustness [BO17]. Antiviral [ZRE15]. AntMeans [HMC12a]. Anyigba [ASD19]. AO [JJA14]. AODV [AJSG11, BGS11, DS10, DA10, DGPP15, HS15d, KJJ15, KK12b, KC15c, MAM15, MS10b, PIB11, SS10d, SG12a, SA15c, SRC12, TY15b, TKR15, VV12]. AOMDV [ES13, KVKR16, KC15c, TKR15]. APC [GDS16]. APD [BOS15]. Aperiodic [ARHERZ10]. Aperture [Has14]. API [HS24, AAS24, CW23, MKAR22, PSG12, PA23a, RA23, SRK10]. APIs [AS24]. App [AHAPT15, SDW21, Ver16]. Appetite [TD16b]. Appliances [Joh14, Kad16, MA15, SGA13]. Applicability [MR13a, RMS14a]. Application [Ade11, AAT16, Ali19, ARL16, AAR13, AIP18, Ara15, AER19, BK14, BD13, Bar14, BS16, BK17, BML22, BYKO17, CN12, CK12, DV16, DV14, Eja12, EKI16, EK14a, EME15, Far24, FLM11, GEAE15, GVRP10, GK12b, GSCT12, GPB22, GKK13, Has16, HBN12, IUN10, JD13, JSSA20, JFJN20, JI10a, KNSS11, KA12d, KHSE12, KBN18, KMG21, KM16d, LR12, LBNP12, LAAI23, LSP17, LCP21, LAM14, MG17a, Mal13b, MK18a, MM20b, MM14b, MNB13a, MNB13b, MP12b, MK18b, MTM12, ML13, NI14, Nad17, NM14, NA18, NKM12, OFE15, PAK16, PWA12, PA17, PP16, PA23a, Pra14, PP12b, PW23, RT11, RK16, RKGK16, RA12b, RHT19, RDD^{+}12, SMSC14, SS10a, Sal14, SN15b, SOT18, SYR17, SP11b, SPP16, SS23a, TSS12, TY13a, TG18, TO13, WES17, Yak18]. Application [YMA13, Zhu21, AB17]. Applications [AKK11, ACK12, AK18b, AMK13, ATB12, AYS19, AS13a, AAaNJ13, AG14, AAR14, Bon15, CS16a, CHG12, CKA18, DFGG15, DBKN20, EADEA17, FEE13, FCW14, FSG17, GC24, GE12, GENERM16, HSK14, KB14a, Ker13, KD22, KD15, KK10c, LSK17, LK16b, MKWK15, MDM17, MMK16, NSMK17, PCS18, PK19, PK12b, PS16a, PF20, Ras17, RGK17, RK15, SZM18, SS14c, SSS15a, SM10, SAR18, MNB^{+}22, SSP13b, SRS14b, SKJ15, SET15, SA12e, TRDB24, TD16b, TAM13, TS16, VRS12, YAE11, ZEE15]. Applied [ASS10, ABD22, CAL^{+}14, DBM14, DGH16, DXT23b, HES14, İHŞ17, JTA19, KAFS15, Lic20, MGA20, MRG19, SAk13b, SP15e, TN14a, TERCM15, TOHS14, YMAS17, DDN14]. Applying [BM15a, Ghe16, JAG15a, MK13b, MJ14, Rad13, SsTM12, ZRZ11]. Appointment [AB16a, EKI16]. Appraisal [RGK17]. Approach [ASH16a, AEk13, ADH17, AHH13, ADT17, AEEO24, AD14a, AAA20, Agg13, AHZ15a, AAB22, AAA11, AHBG19, AS18a, AzSKM12, AS14c, AS18b, Alb15, ARHERZ10, AJI15, AGRR10, AOK18, AH12c, ASK15, AYAI17b, AK15, AS18d, ASH13, AAC13, AA11b, aAO14, AGH14, AAS13, BCO13, BB14a, Bad13b, BM13b, BN14, BS12b, BP16, BRD15, BS14a, BMM13, BRR16, BYT14, BK10b, BP15, BBL13, BKB11, BA15b, BMRA12, IU18, Cai18, CA11, CR12, CJK17, CGKP13, DAaGZ13, DDB11, DPJB12, DP11, DD20, DKP10, DRBR24, DBB24, Eis14, ESEMD21, EBR12, EMA11, ESB13, ESE16, EnFn14, FML20, FRR11b, FZM12, FY12, FCW14, FCS11, FTLK14, Gaa16, GEAE18, GPWS18, GSP18, GMRK12, GP12, GSA11, GFMS14, GSB12, GCS14, GENERM16, GJM16, GSA12b, GJ14, HM13, HZS14, HF20, HK10, HSH17]. Approach [HPR17, HAb14, IPB14, IM12, Ira12, sNIALY13, IND13, JSSA20, JS18a, Jah18, JS12b, JPJ18, JS12c, JA11, JP15, JVG15, JLCV16, Jay12, JP16, JGZ13, JS15, KA12c, KSA20, KK18, KB12a, KS12a, KRSB12, KV15, KK16a, KT12a, KST16, KP15b, KJJ15, KBK12, KSC11, KJKK12, KRM12, KA13a, KS13b, KAR15, KSMM15, KRR13, Kum10, KR16, LR11, LCP21, MM13, MC17a, MM11a, MD13, Mal13b, Mal13c, MJ15, MB13a, MP12a, MM16, MA10, MNB13a, MNB13b, ME15, MMPH15, MR12b, MB13b, MIS12, MD14b, MAMS12, MVC10, NP19, NKR14, NTY19, NK15, NUK16, NM20, Neg13, NSM14, NJB16, OOE18, OP18, OO21, OE14, PDK14, PB13, PSA18, PRP15, PR16, PG11, PS13, PS10, PMM18, PMD13, RR16, RUB11, RTD10, RRudS12, RSA11, RAR24, RGK17, RML11, RGCGCMFJ13]. Approach [RL14, SSG13, SB21, SE17, SB13a, SF15a, SS10b, SN10a, SMS16, SG13a, ŞAG17, SP15c, SSS15a, SA12b, SHR11, STG12, SP15d, SA12c, SD12, SC15b, SSC15a, SGT^{+}11, SNJ16, SSC11, SZUSNK19, ST14, SA12d, SRM12, Sta22, SV12b, SB18b, SMR16, TGcO14, TB14, TERCM15, TSY18, TV13, TK^{+}17, TS11, UNM18a, UNM18b, US15, VG12, VMM14, VP10, Win14, Wu21, YRA23, YAA14, YMR^{+}12, ZCDR13, AR16, DS12a, FD14, GN14, GAA17, NM24, RP15, RH14]. Approaches [AA14d, AB13c, ARR20, CP15b, DFGG15, EAEE14, GJS17, GV15, HEBR11, HFSM12, MSM17, Ori14, OOC18, RKR11, SRTB11, BR11a, KAS12]. Approachment [SDW21]. Appropriate [AA18b, EG14]. Approximate [AS18b, BHK13, NW13, ZYW^{+}18]. Approximation [KBK10, Kha20, KC15b, LA17, MMA13, SLU14]. Apps [AIP18]. Apriori [ur13, STS19]. AQM [OMH11, S13, SN10a]. AR-HMM [Isl18]. Arabia [AS19]. Arabian [Alt11]. Arabic [Ae14a, AbAAw12, AAG13, AHM17, AAS13, DAB^{+}16, DB12a, Eg13, ETKE20, EEE14, EA18, EEL11, Far12, GF14, ME14, MBE10, MAt12, She11, TB13, Yac15]. ARABIRS [Yac15]. Arbitrarily [JT15]. Arcade [DSS18]. Archetype [ADS16]. Architectural [RM11, USC12]. Architecture [AM11, AHZAH16, AD18, ABAO24, ABBG14, BECM11, BBB16, BK10a, BM15c, BK11, CS16a, CDC18, DLBM12, DH19, DBB24, EzHEgT13, Far13, FSNR18, GSA12a, KUM13a, KPV12, KB16b, Kum19, LHKR16, MM12a, MM15c, OAW12, OV13, PA23a, RSJT14, RBN^{+}16, RD17b, SE16, Sm13b, She12, SM15b, SK10c, SR14, TY15a, TO13, TAM13, UAA16, VMA23, VV18, JSD14, MN24, Sri12, Liu24]. Architectures [AB17, CLBA^{+}12, HSMb12, LB11, MKCV13, RS16, RRB16, SB11b, SSM12]. Archiving [OJO11]. ARDL [Mal13c]. Arduino [AIC19, RR18]. Area [APU16, AICN14, Ala16, AKG20, CPSP13, GL16, GMRK12, GK12a, Jay15, JMM21, KAAA13, KK10c, LSB17, Mal10, PA13, Par14a, PP16, SK15a, SSK13a, SM13c, TAMA13, TMS15, UDA13, YK13]. Arhythmia [KKM18]. Aricha [SAAM17]. ARIMA [HJBJ11, HLZ^{+}24]. Arithmetic [GRP13, PR10b, RAA15, RS11, SR14, AS16]. Arm [SPXT17]. Aroko [Olu14]. Arranged [SM17b]. Arrangements [GM14b, rPGM14, GM15b]. Arranging [SCL14]. Array [BSM17, GA15, Kum19, LR12, SM14a, SM17b, SC19b, Tah14]. Arrays [AA16a, Bou12b, CR15, DRS15, FR14, ZYW^{+}18]. Arrhythmia [IHB^{+}15]. Arrival [Abd14b, AOO20, AS15b, KS18, KC15a, KK17b, MG12]. Arrivals [KRR15, RRR11, SRSV14, RSSV14]. Art [AMG^{+}12, HO15c, Ish14, NP13, PS15e, RRB16, SAK18, NUK16, HaB13b]. Artefacts [MN22]. Artery [HS14]. Article [MD13]. Articles [CPG15, KVK19]. Articulated [SS12d]. Articulations [MBKC15]. Artifacts [AR15, HSH17]. Artificial [ASD19, AD24, AA14a, AAA14b, AHH14, AI18, AGG16, ASG11, ANA11, BSAb12, BI12, BS11a, BWH15, BML22, CP15a, DK11, DXT23b, DTO23, EBAS11, GRS22, GGK14, GYP^{+}15, JIF12, Jab14, JPDB14, JAB22, KSKS23, KCD13, KSK14, LDB^{+}22, MK18a, Mes16, MK15a, MB15b, MRG19, MG19, NO18, PC11a, QKRB11, RBF12, RM15, SHSA12, SO14a, SGS15, SC13, SSA16, SSK13b, SBC12, SSJ12, Tah17, TBN^{+}15, TOMGBSS13, YWT12, ZEB17]. Artisanal [SKZ22, SVB23]. ARX [OJR11]. ary [DM14, XL21]. AS-Cycle [ZYL^{+}21]. Asantewaa [AP16]. ASP.NET [GENERM16]. Aspect [ASB12, GAZ13, MMHA11]. Aspect-Oriented [ASB12, GAZ13, MMHA11]. AspectJ [SMM11]. Aspects [DP16, IAQ11, MB16]. ASR [KBEA17]. Assembly [SS10a, STG13]. Assessing [AK18a, ATM18, HHHK21, RKGK16]. Assessment [AHA14, AAO22, AAA11, AAO11, AMY23, AS19, ur13, BC12, CB16, DA12b, ENE13, GSP18, GFMS14, HB16, KA12a, KRH13, KB15, LS13, MM14b, MAM19, MJP18, MRG19, OAA24, PFS12, Poo22, RÇ16, Ras17, RR22, Sab16, She11, Sub24]. Assign [CS18]. Assignment [CK13, WDS13, GHD17]. Assist [IA18, LSA^{+}17]. Assisted [PG15b, SZS15, RV12b, Tha14]. Assisting [EAE19]. Assistive [AMS14, Sal19, SS23a]. Associated [BFD15, Mes23, SSV11, VV14]. Association [AMBS16, DBM14, KS12a, LHM15, Mah13, OP18, PP12a, PP10, STG12, VMM14, YRV14]. Associations [EsGK14]. Associative [El 13, NmEM13, ZYM13]. Assorted [KTDG12]. Assurance [Kas17, MKSB22]. Assuring [BNS15]. Asymmetric [KMS15]. Asymmetrical [PB16]. Asymptotic [EEB16]. Asynchronous [MS14b, SSK^{+}10, SM12a]. AT89C5130A [AKK11]. AT89C5130A/31A [AKK11]. AT89C5130A/31A-M-Frozen-in-Idle-State [AKK11]. ATAM [GSA12a]. ATC [Pri13]. Athens [PTPS19]. ATM [LEO15, SYP10]. ATMA [MP12b]. ATMega [LSM16]. Atmospheric [AN16b, PR12, SR15a]. Atom [MFM12]. Atomic [BAIs^{+}14]. Atoms [SPB12]. Atrophy [TRR16]. Attachment [Pra14]. Attack [ARAA11, AM17a, AK16, AMY23, AV10, AV13, Bat18, BH17b, BO17, BKSB15, DSB12, JSK17, KBT^{+}21, KJJ15, LKS14, MAM15, PK15a, PC15b, PA19, RSM15, MP17, RC12, SNP15, SU12, SSS16, TY15b, UJ14, VR11, YRA23, ZTL^{+}14]. Attacked [MRH15]. Attacks [AMG^{+}12, AS15d, Cur20, EOO22, KC20, KSS^{+}14b, MA12, NKG12, NBL15, Par11, RMVP14, SG13a, SC16, SS16d, UC15]. Attendance [JJKD14, OAO18]. Attenuation [Kum14a]. Attitude [HZ14, SFA14, SZ15, SDZ15, SKM16]. Attitudes [Add12]. Attribute [GHH15, IA15, KK10a, PA24, PP13c, VG12]. Attributes [ST16, YRV14]. Auction [MK17, TR12]. Audio [DD15, HO15c, IR10, KFEN21, Lah16, PG13a, SS10c, WDN^{+}20, YBA12]. Audio-Visual [IR10]. Audit [RGNA13, SU20]. Auditing [BR15, SS16a]. Auditory [RHK15, SV14a]. Augmented [AJ18, Ari23, CHG12, GVW24]. AUN [UI16]. Authentic [EsGK14]. Authenticated [SYP10, VV12]. Authentication [AN14, AmM12, AAS24, AS24, LEO15, BBM14, BCBS14, BHC12, CM12, DRBR24, DBB24, EAJ^{+}15, EME13, Far13, GC24, GCP13, HMGV12, HR21, JS14, JP16, KA21, KT12b, KK13c, KMG20, LBNP12, MK13a, MIAEFS15, MNYMGG12, Okp17, RW20, RDD^{+}12, STM13, SSS16, ZAHH16, SMJ12, Liu24]. Authentication/Authorization [AS24]. Authentications [SC16]. Authoring [MB13a]. Authorís [Nas11]. Authority [EzHEgT13]. Authorization [AS24, CS18]. Authorized [BR15]. Authorship [EK14b]. Auto [AS24, BO16, IGS17, KM13a, MNHB12, SS10a, ZYM13]. Auto-Associative [ZYM13]. Auto-delegation [KM13a]. Auto-Flow [SS10a]. Auto-Immunity [BO16]. Auto-Scheduling [IGS17]. Autocorrelation [DB10]. AutocorrelationTechniques [NRR12]. Autoimmunity [BO17]. Automata [Al 14, DDDD12, DB12b, ElR16, HJG^{+}13, Hew11, KR15, MM15a, MK15b, NSM13, SZZR11, TAM14a, ZSR14]. Automate [PS21]. Automated [AmAdS12, AS13a, CHW16, CS16b, Cha12, DSDD12, ENC12, EEA19, GS16a, GP15, GEG23, Hik13, KSV10, Kha20, KK15c, LHKR16, NHS11, OFI17, RBDD16, SB11a, SWH16, SNA13, UAA16, UNM18a, WBL^{+}23]. Automatic [AD24, ATH12, Ahm14a, ASB15, AJI15, ASH16b, AERE21, BS12a, BB12a, Cai17, COWK14, DAB^{+}16, DP15b, DMB21, DRS13, DBW19b, Far12, GGK14, GJM16, HMN18, ISH12b, Jay12, KN17b, KNS17, KKR^{+}14, KLS13, LC17a, MO17, MB11a, NR16, NRP12, Obe23, ONK24, OLT14, PFS12, PG13b, PMH10, RNM14, RRB12, Rez12, SPK16, SC15a, Sm13b, TArM14, VPGF10, Win14, WHP14]. Automatically [EAE19, KJ14]. Automating [Tes15, Upa23]. Automation [BM15a, BA16, CMKS15, NA18, PHK^{+}15, PCS18, PSA18, WKG^{+}18]. Automaton [BP11, HKS12]. Automorphism [Ela11]. Automotive [PRK17, TB12a]. Autonomous [AAA14b, CBhM12, HM13, JPAK11, KNI14, MOM17, SA15a, SBHM13, MNB^{+}22, TVY17]. Autonomy [Kor23]. Autoregressive [GCK17, IM15, Mal13c]. Availability [MJM17, MY13, WBK22]. Available [ETD13, RB16]. AVC [AJN18, AE12a, DLBM12, JKS^{+}12, PB19, Sar13, BWS^{+}12]. Average [AB12b, HO15a, Kei14, MRP16, Neg13, OA12, RM15]. Averaging [MP12a, OJR11]. Aversion [SNP15]. Aviation [KAFS15]. AVISPA [VV12]. AVL [Tri10]. Avoid [MAM15]. Avoidance [HM13, HEs12, KAS13, S13, YSH12]. Avoiding [KA12b, MFE13]. AVR [ASS10]. Aware [AS18b, AS18d, BM14b, CSKB11, CTG10, DF15, ETG17, Fla18, KZ16a, KM16b, KDTD11, MB13a, NV12, PSS12, RGS15, SMSC14, Sal18, SAR18, SM13d, SDS15, KVR14, Meg10]. Awareness [Ash19, DSS13, aKNP^{+}15, ZRE16]. AWGN [AEE15, AAaN15, BR11b, IS21, MA17, PAH17, SGD13]. Axes [SPXT17]. Axial [GK15c]. Axially [SK15b]. Axis [AVT13, Rag15, SZ15, SKM16, SM14a]. Azeotropic [PSD13].

b [KB14a, BB13a, SAK13a, SN10b, TOMGBSS13, VR12, VR15b, YPS19]. B-DEC [YPS19]. B-scan [SN10b]. B-Spline [SAK13a, VR12, TOMGBSS13]. B-splines [VR15b]. B2C [Sri12]. Back [AS12a, ASC15, AB13c, LG11, RV16, SO14a, SSP13a, Sud10]. Back-Prop [RV16]. Back-Propagation [AB13c, SSP13a]. Background [KA13a, KY15, MKK13, SB21]. Backhaul [Kum14a]. Backlash [GAA15]. Backlogging [BMF11]. Backoff [JT11]. Backordering [TYSS15]. Backpropogation [IUN10]. Backscattering [ASK17]. Backstepping [IHH11]. Backtracking [EDD19]. Backward [AV15, TY15a]. Bacterial [JA22, KJ12a, KR10, YAT15b]. Bad [BD18, PRT^{+}16]. Baffles [Hus15]. Bag [BP16]. Bag-of-Features [BP16]. BAGA [MHV^{+}14]. Bagging [MV12b, VK13]. Baghdad [ANAK17]. Bagley [ŞAG17]. Baids [RC12]. Balance [aAO14, DBW19b, FSER20]. Balanced [DBB12, EmdHNmAW13, ES15b, KAFS15, RRudS12]. Balancer [ABBG14]. Balancing [CJK^{+}12, sMTEkEdAr12, EMF14, ETG17, GK12a, Kon15, KG15c, Ngu24b, OAA13, PPP16a, SD18, SP13, SH15, STG13, SVJ15, Tri10, YMA16, PPM12]. Band [AK18b, AI13a, DJA13, LW16, MDGS13, MFE13, PVR20, PA17, PK12b, SDD14, SKJ15, VMT13, MT13b, PS16a, PPP11, RC14]. Bands [BI19, CA16, JM14, SHS23]. Bandwidth [AMN12, BBMA13, DD14, ElZA13, JD13, JMM21, KCD13, KA17, LGM12a, MSP12, OMH11, RSSV14, RS11, RRR11, SMT19, SRM13, VD15, VMT13, YRV17]. Bangla [AA17a]. Bangladesh [am12a, APU16, Rah12, RM15, SAAM17]. Bank [AP16, IS21, PJ12, Sab16, CTG10, PPP11]. Banking [DS12a, HMN18, HMB13, HHS14, NF17, OOC18, Sab16, TS15, ZUDS16]. Bankruptcy [ZRZ11]. Banks [KK14, SE14, SCE14]. BANs [TAMA13]. Banten [RHT19]. Bargaining [FHM16]. Barrier [KAA20]. Barrier-Free [KAA20]. barriers [GT11]. Bars [DC22, HBN12]. Base [DSA11, IEME12, PA23b, USC12, Ask14, BSG15, BLA12, DSS13, MEE15, OO14, TK^{+}17]. Base-Fuzzy [PA23b]. Based [AD24, Ade13b, AR10, AG10, AHZ15a, AASAA15, AS24, AVH10, ABAO24, BBC12a, CN10, DS10, DP11, DL23, DKP10, DRBR24, DBB24, Far14b, GC24, GVW24, HK10, HS10, HR14b, IR10, JG11, JA11, KGAM10, KJKK12, MAaM14, Man10, MAt12, Ngu24b, ONK24, PVL11, S24, SS10a, SA10, ST11, SMK17, SP10, TNC10, VN10, ZZW24, ZKN10, AN14, ASH16a, AC18, Abb12b, AII13, AS17, AJ15, AE13, AHAS^{+}17, AAEK18, AAH21, ATH12, ABYB22, AHZAH16, Abd14a, ADT17, AH12a, AJ14, Abd14b, AH12b, ATM12, ANE^{+}13, AYAB18, AZM13, AD14a, ASY17, AAA20, Ade13a, AM19a, ASN11, AKH15, Agn15, AJSG11, AS15b, AS12b, Ahm14a, AKN13, AEA17, AAO11, AH17a, AD11, AHBG19, AMAHH19, AS18a, ATA22, Aa13c, Ak13, AmAr14, AmM12, Am13b, AQ15]. based [Aq14a, ArAn13, AsS13, AIC19, AH14, AH11, Ali12, ARK13, ASH15, AAA16, AHH14, AQ14b, ASM13, AMG^{+}12, Als18, AT17, AOK14, AGL17, AA12f, AA14d, AMS16, AC15, AKA^{+}18, AAZN21, ASCR11, AERS11, AOK18, CK13, AYM16, Anh15b, ASK15, AIP18, AA17b, AER19, AAO13, AVS13, ALA12, AG14, AAMA20, AE17, CT13, ANA11, AKB16, AAaN15, ur13, aAO14, AM14, AHS15, AZY18b, BA15a, BSAb12, BM14a, BAM13, BA11, Bar14, BS12b, BS21a, Bas12b, BECM11, Baz18, BKM12, BMU14, BRD12, BB12a, BFD15, BS14a, BBHL12, BBC12b, BS11b, BH17b, BM13c, BYT14, BSP11, BV21, BNS15, BI13, BC22, BKB11, Bou12a, Bou13, Bou12b, BA15b, BKSB15, CS13, CHW16, MRTB16, COWK14, CS16b, CJK16, CT14, CQM15, CSP14, CSN15, CPSP13, CT10, CS15c]. based [CC17, CHG12, CKA18, DXT23a, DL19, Dak23, DR14, DK11, DSB12, DMS13, DN12, DSDD12, DV16, DSA11, DSJ11b, DDL18, DRS13, DT22, DBM14, DPT14, DK18, DBB12, DR12, DS12a, DR15b, DSM14, DS12b, Dha12, DS12c, DBW19a, DB13b, Dur14, CB12, EAS16, EUT17, EER14, EEH14, EUEI23, ElZA13, EsGK14, EsTeEh^{+}13, El 20, EAE19, ETKE20, EBB17, EGES15, EME13, EAEE15, EEL11, ENE14, Ela12, EAEa12, Elc11, Elg13, ESR13, EIB^{+}11, EJC16, EHHM17, EAF^{+}16, EB13, EAO15, ED19, EnFn14, FA13, Far13, Far11, Far17, FRR11a, FRR11b, FAZ13, eFKS13, FR12, FSG17, FH13, FNH13, FSER20, FPA13, Gaa16, GEAE15, GPWS18, GJ15b, GJP13, GCK17, GHD17, GSA11, GAZ13, GSB12, GP15, GK13, GHH15, GRS22, GT13, GPB22, GGX12, GJ14, GOCA15, HJBJ11, HKA15]. based [HMO17, HAAA15, HMC12a, HZ14, HMGV12, HaB13b, HWT^{+}13, HPC13, HIE14, HAS17, HS12, HHHK21, HE15, HB16, HAAT16, HMB13, HM18, Hew11, HR22, Hik13, HTPB11, HR14a, HST19, HAH15, HZLL15, HLZ^{+}24, mH14b, HK15, Hus12, HEN13, VNM15, Ibr14, IHH14, IAH14, IM12, İHŞ17, IKY22, Isl18, IN15, IA15, Isu16, JHP12, JA16a, JA16b, JA12, JSD14, Jah18, JVJ12, JBK13, JS14, JC14, JJN14, JM18, JS12c, JAsA13, Jas13, JK14b, JM15, KM17, KAA13, KS14a, KM13a, KFEN21, KHAE19, KB12a, KC14a, KSS18, KUM13a, KCD13, KKT15, KRP12, KHSE12, Kas17, KAR12, KS14b, KG15b, KS15c, KP15a, KA16b, KV16a, KA10, KPM16, KJKM21, KDB17, Kei14, KTBK12, KMB14, KMMM12a, KMMM12b, KT12a, KAqC14, KAP11, KJK12, KNS17, KS16d, KB14b, KBM22, KMT^{+}24, KW16]. based [KA21, KBAQ15, KKK12, KN12c, KR14, KS13a, KK15b, KR10, Kum12, KA12f, KAS12, KPV12, KGK13, KVR14, Kum14a, KS15e, KC15b, KVS16, KS16e, KS13c, LJ16, LBNP12, Las10, LHKR16, LcCm13, LC17a, LC17b, LSSD20, LMZ12, LKS14, LWL^{+}23, LGZX17, MWS20, MM13, MBKC15, MN15, Mah19, MAA13, MH12a, MC17a, MK13c, MM11a, Mal13a, MK17, MJ15, MSM17, MMG16, MB13a, MK18a, MA16, Mar12, MRP16, MA12, MP12a, MM18, MS10b, Maz21, MRR21, ME15, MAE16b, MB13b, MK18b, MMR14c, MSA16, MSA17, MAE19, MNR14, MZAE16, Mos11, MAM12, MABE16, MDNA12, MCA14, MEKS15, MNYMGG12, MZ18, MI13, Mug24, MJP18, MGB11, Mun14, MMK16, MEH14, Nad14, NTY19, NUK16, NEA17, NK19, NV12, NP11, NRRM12, NSM13, ND17, ND15, NNR16, Ngu21, Ngu23, NJB16]. based [ND11, MSA18, NF15, NMK14, NMK15, NS23, Nta14, OMM14, Ode14, OC20, OOE18, OE14, OS16, OFE15, ONJ13, OOC18, OCZ13, PAK16, PPM12, PB13, Par14a, Par14b, PRP15, PG11, PP13c, PB19, PS16b, PC15a, PSG12, PM15, PG15b, PRT^{+}16, PMM18, PDR11, PA23a, PP12b, PA12, PWG15, PG13c, PGS13, PS12b, PA23b, QT15, QKRB11, QJ13, RMP11, RRVP12, RS13a, RTK15, RUB11, RR11, RÇ16, RMS12, RSHI12, RBN^{+}16, RSA16, RM16, RAA14, VGS11, RJS^{+}13, RS12, RAE15, RB11, Ran13, REA12, RKR11, RV13, RB14b, RW20, RB15, RMD10, RAHH16, RS14b, RKGK15, RP24, RG21, RH14, RS13d, RAB12, RDE^{+}15, RS16, RGJ11, RGNA13, RA23, RPS12, RL14, RJC17, RD17b, RK21, SV12a, fSG13, SSG13, DBS15, SMN^{+}22, SA15a, SE16, SKS13, SAY16, SN15a, SM17a, SS14b]. based [SP15b, SO14a, SYAS11, SO14b, SIAI14, SN15b, SN15c, SOOA13, SCD10, SMS16, SZZR11, SPC13, SK18, SS12b, SG13b, SP15c, SC15a, SAS21, SS13c, SPP11, SSP15a, STE14, SHD13, SSAZ13, SHS23, SI13, Sha14, SG12a, SSK13a, SAS13, SSS13b, SS13d, SSA14, SN14a, SC15b, She11, SA15c, SAI16, SM12a, SB11a, SGS15, SRS14a, SKT14, SSC15a, SY15, SDB15b, SGS18, SK13b, SBM^{+}15, SMM11, SSG11, SH12, SK12b, SSS13c, SP15e, SC15c, SSA16, SE11, SK14b, SC16, SA12d, SVJ15, SA12e, SB18a, SMT17, SL15, SB12b, SC19b, Suj14, SZS15, Sun19, SN16, SJV^{+}15, SJ19, SR17, SB12c, TBKG13, TGL13, TFT11, TT14, Tah14, TX20, TAMA13, TBN^{+}15, TL16, TAS^{+}16, TAM19, TB14, TGR12, Tha13, TS15, TY15a, Tha11, TMP11, TRSK15, TY13b, TM13, TAH13, TY16, Tua17]. based [USR17, UAA16, UNM18b, UR12, UR16, US15, UO23, VKB14, VRS12, VK13, VM12a, Ven14b, VD15, Ver17, VMT13, Vya14, VDM22, WC15, WDAJ15, WDS13, WES17, XCY11, Yac15, YRV14, YYS^{+}23a, YRMB15, YEE14, YMA13, You13, You18, YPS19, ZEB17, ZEfNF13, ZAER12, ZD15].

Based-on [WDAJ15]. Baseline [AHM14b, BWS^{+}12, CK12]. Basic [AMM21, MC13, IHB^{+}15]. Basis [DHT13, MK18a, MKT17, VR15b]. BASK [SA12e]. Bat [GV15, RAA14, TAY15, UA14]. Batch [AD14a, DH12, GN14, KC15a, RBT10]. Batch-based [AD14a]. Bathymetry [NRL17]. Batteries [SD18]. Battery [BAK14, MA16]. Battling [Far24]. Bayer [MM21]. Bayes [AU23, Bef18, GS16a, HS24, SN15a, TK15]. Bayesian [FRR11a, JA11, MHV^{+}14, NCL11, RJC17, SJ12, SC15b, Sin16, VV14]. BB [VM12b]. BB-BC [VM12b]. BBO [JSK10, SS13a]. BC [VM12b]. BCH [MA13a]. BCI [AkC13, KAR12, ZEB17]. BCO [JSY14]. BDD [SSC11]. be [DBM14]. Beacon [RS13a, RKB19]. Beacons [YK13]. Beam [BBD13, JTI19, VR15a]. Beam-Forming [JTI19]. Beamforming [KK17a, KT15, VK11]. Beams [DRS15, IFT14]. Beans [MBA20]. Becoming [LSM16, SHHL13]. Bed [MMHP17, SFA14, DGPP15]. Bee [AD24, AA14a, AAA14b, ASG11, JSY14, KSK14, MB15b, NO18, PC11a, SGS15, SVJ15, Tah17]. Before [NG10]. Beginning [MG19]. Behavior [AB13b, DL23, EEB16, EN13, FSNR18, FK15, GA15, GSP18, HK15, IFT14, MS14b, NMN15, PP13a, RD13, RKR10, SEE13, SS15b, Vad20, Ali19]. Behavioral [NU12, SS21]. Behaviors [ZRP16]. Behaviour [DJ12, JSD14, KS16c]. Behavioural [KAAM11]. behind [NUK16]. Bello [UDA13]. BEMD [EAMAES16]. BEMD-DWT-DCT-SVD [EAMAES16]. Benchmark [AS13a, KJ12b, PB23, YAT15b]. Benchmarking [PKK18]. Benefit [JH18, MG12, Qte14]. Benefits [Asa13, OD13, TGR12]. Bengali [ASCR11]. Benign [Sha16]. Beowulf [PKK18, PB23]. Ber [SNJ16, AT15, BR11b, Kum14a, SSU10]. Bernoulli [AS13d]. Bernstein [KBA10]. Bertoni [AMM21]. Best [APU16, DH21, SZUSNK19]. betterh [AkC13]. Between [Alg19, DC22, ABA^{+}10, Ahm14a, AAAA17, AYS19, ASB12, BB13a, BNA14, CKA18, Elg13, Far13, FS11, GKG16, GSV20, HITA20, Kum14b, Ms13, RA10b, Sak15, SH15, Uzo13, YMR^{+}12, ZAAS18]. beyond [Rah13]. Beyza [MJ14]. Bézier [ME14]. BG [SMSM10]. BGPS [DH12]. BHnFDIA [RMVP14]. Bi [HSS11, MK19, ES15b]. Bi-Directional [MK19]. Bi-quadratic [HSS11]. Bianchi [AHH20]. Bias [RSK15a]. Biasing [FA19]. Bibliometric [BAK20]. Bidding [JM18, RSK15b]. Bidirectional [KTS17, PSR13]. Big [AAT16, MA22, BBB16, DSM^{+}20, ES15a, GMM21, KPC16, MREd14, NP19, PCS18, SMGS17, SS23b, TOHS14, TVS19, SE17, Nor13]. Bilateral [AOK14, Anh14b, KM17, QJY11, RV12a]. Bilatetal [Anh14a]. Bilingual [MH13a]. Binarization [CAL^{+}14, KKBS12]. Binary [AMS13, AHBH11, AhAT13, Al 14, DP15a, FTA18, GHBN12, rPGM14, HJG^{+}13, KPV13, LKO16, NK15, NJAK15, NMB13, PFS12, PR10b, SK15b, SSC11, Sud10, Tha14]. Binning [GM14b]. Binocular [SM19]. Bio [KS13b, NDC10, Neb13, PR10a, PSM15]. Bio-fuels [KS13b]. Bio-Inspired [Neb13]. Bio-Medical [PSM15]. Bio-Metric [PR10a]. Biofeedback [LSA^{+}17]. Biogeography [SKS13]. Bioinformatics [fSG13]. Biological [FLLR20, Wu21, Wu23a]. Biology [ETI14]. Biomass [BJS14]. Biomedical [AK18b, SET15]. Biometric [AOSM12, AAS24, AnAA14, AS13c, AA11b, Dha12, EAJ^{+}15, El 14, EAEa12, KC14a, KGK10, KS15e, KMG20, OS14, RDD^{+}12, SS10f]. Biometrics [AEKE15, BAK20, MR11, OAO18, Sab16, WES17]. Bioparticle [DK18]. Biosensor [SS10a]. Bioterrorism [YS11]. Bipartite [KIH15, RS13c]. Biped [RA12b]. Biphase [BKM12]. Birim [ATA19]. Bis [PN10]. Bisecting [PK15b]. Bit [AKK11, AMHAh13, ARS13, AmAr14, AGRR10, EAF^{+}16, FPS16, Fla18, GEB12, IS21, KWS^{+}11, KG11, MFM12, NW13, OG16, SD17, SSK13a, SA15d, DV14, GHBN12, SLK^{+}13]. Bit-Stream [AMHAh13, KWS^{+}11]. Bit-Vector [NW13]. Bitcoin [A18]. Bitherapy [ZRE15]. Bitmap [SEA16]. Bitopological [KETESH17]. Bits [AD18, MA13a]. Bivariate [AO13a, EBAS17]. BJWT [AAS24]. Black [AAC16, AV10, KBT^{+}21, KJJ15, MAM15, NBL15, NMB13, RMVP14, TY15b]. Blackhole [RC12]. Blades [RHM15]. Blast [KPG13, L13]. BLDC [AH16, CP15a, VGJ12]. Blend [ADVB10]. Blended [MRG19, Zae13]. Blending [AAS13]. Blight [JA22]. Blind [AMK13, DBB24, Ela11, HO15c, JSC12, KH12, MK14, PKP13, PVL11, SA14, TSY18]. Block [BS10, De 20, DXT23b, EME13, GSA12b, KTDG12, KH15, KJKK12, MZEE13, MRP15, SG16a, SSG11, TR15, ABA^{+}10]. Block-Matching [MRP15]. Blockade [LDB^{+}22]. Blockchain [GJUL23, GC24, KMS19, RD17b, YPS19]. Blockchain-Based [GC24]. Blocking [LK16a]. Blocks [BWS16]. Blog [GK12b]. Blogs [Rok20]. Blood [AAA20, AS18c, EUEI23, HS14, JIB14, KK15c, MABE16]. Bloom [NS23, PA13, urPP13]. Blossom [Zhu21]. Blowfish [ARMS16, SSS12]. Blue [BD10b]. Blueprint [NS23]. Bluetooth [AJ16, JLCV16, Kad16, Ver16]. BMH [KTS10]. Board [HPM10, LcCm13]. Bode [BM13c]. Body [AKG20, EAEE15, ETLE13, GMRK12, TSCD15, TAMA13]. BOEbot [CS15b]. Boids [RBF12]. Bold [YM12]. Bombay [KGAM10]. Bond [NEA17]. Bone [GYP^{+}15]. Bones [DPS21]. Book [ATM18, OEA18]. Booking [PW23]. Books [EAE19]. Boolean [SCD10, SKA13]. Boost [AM17c, HR14b, KM15a]. Boosting [FRR11a]. Boot [HKR17]. Bootstrap [FRR11b, LCP21, SS23b]. Bot [APS18, ENC12, TD21]. Both [ISH12b, AAEEA15]. Botnet [AB13b, AMG^{+}12]. Botnet-based [AMG^{+}12]. Bottling [ATH12]. Bounce [BSZ13]. Boundary [HJ13, SAK13a, VR12, VR15b]. Bounded [CDC17, VMA23]. Bounds [Fen18, Mal13b]. Bovine [DSP^{+}21]. Box [AAC16, DDL16, RSA11]. BP [MWS20]. BPM [urPP13]. BPNN [Bar16, VKJ^{+}15]. BPSK [SA12e, SSU10]. Brahmi [HPM20]. Brain [Aa13c, ASH15, AVH10, BADR16, CA11, ESAA18, GJ15a, HSD13, JARM13, KS14a, Las10, NJ16, Obe23, RV16, RMD10, SDB15b, SCK17, TV13]. Breaches [OOC18]. Break [MTKP15, RGCGCMFJ13]. Break-Down [RGCGCMFJ13]. Breakdown [CCE20]. Breakdowns [AER18]. Breast [DS19, EE13a, FMS13, GS16a, MPR14, MSD12, RR16, SVJ14, SR17].

Bridge [MLRU12]. Bridging [Ahm14a, SB10]. Brief [PCKD17]. Brightness [Kau15]. Brillouin [TIT14]. Broadband [CS16a, DV16, DJA13, OS16, PA17, VM14b]. Broadcast [MCB12]. Broadcasting [GRC13, SA15a, SAS24]. Broadside [AA16a]. Broken [DC22, HBN12]. Broker [GRL20, NJB16]. Broker-based [NJB16]. Brown [JA22]. Brownian [Lic20]. Browser [AR15]. Browsers [GS16b]. Browsing [NBF14]. BRS [BBC12b]. Brushless [AAEEA15, SO14b]. BT [BD10b]. BTC [TMK^{+}10]. BTE [AO13b]. BTZA [PN10]. Buck [KM15a]. Buck-Boost [KM15a]. Buckeling [CM15b]. Budding [NU12]. Budgeting [RK19]. Bufferbloat [SPP13]. Build [Fad18, HR22, SWH16]. Builder [Kar11, KW14]. Building [AA14b, AZY18a, De 20, EAE19, Far13, FLM11, FNH13, GBHI12, IEL18, IAA13, JBP16, Kum13b, MBR16, OO14, RRAFAM19]. Buildings [Rez12]. Bulk [KRR15, MTM12, RKR17, RRR11]. bunched [BB12b]. Bundles [PP17a]. Burners [TSS12]. Burst [AAY17, DDC12]. Bus [CSKB11]. Business [AnAA14, ADKsO14, KAFS15, MBMA11, NSS15, PK12a, Roc17, RML11, RGJ11, SS14e, SNA13, VV18]. Businesses [IA18, OK14]. BW [HKA15]. Byzantine [BSI12].

C [Kar11, MEAT11, PA17]. C# [Kar11]. C-Band [PA17]. C4.5 [BD13, SH23, SAS13]. C5.0 [Bad13b]. CA [CSKB11, JT11]. CA-ESB [CSKB11]. CA1 [TM13]. Cabinet [YPS19]. Cache [CTG10, CLBA^{+}12, DL23, EGES15, RS12, RGR15, SK14a]. Caching [Abd13, GE17]. Caesar [SA15b]. Cage [RUB11]. Calcification [SS11]. Calculate [AP15, HMY12]. Calculating [BN14]. Calculation [Joh14, MKM17]. Calibration [AMM21, DMB21, ND15]. Call [ASA14, BG17, Uzo13, VMT13]. Calls [MKOS14]. CAM [SDS20]. Camera [Am13b, AH13c, GRC13, HR22, JD15, MKAR22, PKDS17]. Camouflage [MG14c]. Campus [ATA14, GAA17, NOO16, OAG17]. can [AkC13, LJ16, TCS19]. Cancelable [KC14a]. Cancellation [SN14b, US12]. Cancer [AEEO24, AHBG19, DS19, EE13a, eFKS13, FMS13, GVRP10, GS16a, HES14, JARM13, KE14, KM15e, ND11, OAA16, PSM15, RR16, RVGP20, SSA11, SVJ14, SSC15b, TVGT18]. Cancers [Abb12a]. Candidate [Upa23]. Canny [ABD22, FR15, HKY15, TH14, Vya14]. Capability [SSY11]. Capacitance [NMK14, NMK15]. Capacities [MOOOOMT18]. Capacitive [BSM17]. Capacitor [BLL13, KR10, KRM12, KK14, NSMK17]. Capacitors [OAA13]. Capacity [AAH21, ESKM20, FM13, JS12d, KG15b, MA12, SSAK10, SB11c, VT15, MDCS22]. Capelin [EBAS11]. Capture [Abd14a]. Car [AMA15, AER18, Cha12, KBN18]. Carbon [AMSS13, SMHS10, SMK17]. Card [HS24, JS14, hKhHL13, SAP16, TB12a, YJ14]. Cardinal [QT15, SS23b]. Cardiovascular [Abb12b, AE13, EUEI23]. Care [FKT12, OFI17, RK16, RV12b, XCY11]. CARES [IGS17]. Caristi [LAZ17]. Carl [FTA18]. Carlo [Lic20, MSEEs13]. Carpet [SSK13b]. carried [Ans14].

Carrier [Suj14, VSN13]. Carry [SLK^{+}13]. Carrying [CHG12, KG15b]. Cars [Nad17]. cart [JKM^{+}17]. CARVER [YS11]. Cascaded [AKA^{+}18, Suj14]. Case [Ahm20, ATA19, ATAAAF19, AR17b, AI13b, AP16, BAO16, BI12, BWH15, BML22, DRFF15, DDBD22, EBB17, GP13, GAA17, GK12b, GAZ13, yGL13, HS24, IPB14, JSD14, KTA17, KDT16, KW16, KS13c, MEAT11, MMTS13, MDCS22, MAM12, MABE16, MTM12, MI13, NKM12, OAG17, OC13, PK19, RSS^{+}11, RKBB14, Sm13b, SS14d, SE11, SNA13, TTA10, TOD13, VS12, YBM16, GPB22, KAFS15, OK14]. Case-based [EBB17, MAM12, MI13]. Cases [ATA19, RHS18]. Cashew [AGG16]. Cast [MRTB16, PMH10]. Castellated [IFT14]. Cat [RS14b, TY16]. Cat2Vec [NM24]. Catalog [SZOK24]. Catalytic [Bou13, TSS12]. Catalyzing [WBK22]. Catastrophic [Poo22]. Catchments [SPBP17]. Categorical [BS21a, REA12, SSS13a]. Categorization [CR12, ISH12b, MJ15]. Cathode [CKS13]. Cauchy [TA12]. Cauer [KNT11]. Causal [HTM14]. Cause [MMPH15, SPSR12]. Caused [NO18]. Causes [PKJ14]. Causing [IAQ11, MD14a]. Cavitation [SSRD12]. Cavity [BWS16, EAMD14]. CBC [DXT23b]. CBF [PA13]. CBIR [IKY22, KMS14]. CBR [EBR12, PIB11, SG12a]. CBRP [KC15c]. CBT [AJ13]. CCII [KN12b]. CCTV [MKAR22]. CDMA [Agn15, BBD13, MN15, TJ12]. CDMCA [KBTB12]. CDR [AP15]. CDT [Mal13a]. Ceiling [NSM14]. Cell [BO16, ETG17, Job17b, KS15d, KC14b, KKK12, MDNA12, MB13c, RSK15a, SP15c, SM10, Ahm20]. Cells [Abb12a, LAA^{+}17, ZAAS18, LK16a]. Cellular [Al 14, AWK19, DDDD12, DB12b, DA18, ElR16, MT13b, MAM19, NSM13, SK15d, ZSR14]. Census [BMD16, VD10]. Center [HAS20, JND15]. centered [AAA11]. Centers [Car18, MC12, MGA20]. CentOS [KP13]. Central [ASD19, ATA19]. Centralized [LG13b, PG11]. Centre [KA17]. Centric [LG13b, Sub24]. Centrifuge [MM20b]. Centrist [TL16]. Centroid [GPR14, SB13b, TSA12]. Centroids [SK13b]. Cepstral [AS12a, MSM15, NRR12]. Ceramic [BLL13, ISH12b]. Cerebrovascular [PNIS16]. Cerenkov [BB12b]. Cerrado [AGG16, FDG17]. Certificate [Abd13, AV10, GC24]. Certificateless [AR17a, HMB13, KBAQ15, TS15]. Cetres [KAS13]. CFBLB [FSER20]. CFD [SSRD12, TSS12]. CFO [Agn15]. CGLCM [RKR17]. Chain [A18, BOS15, DN12, FS11, GM15c, KK22, MSEEs13, MABE16, PG13b, PS15e, UAMN17]. Chaining [AV10]. Challenge [SPXT17]. Challenged [ZGAS^{+}11]. Challenges [AMBS16, ASASHF17, Asa13, ETD13, JS12c, KBT^{+}21, Liu24, MNN10, NF17, SA12a, ZUDS16, ZEE15]. Challenging [IR10]. Chameleon [TS15]. Champ [PMD13]. Chance [PBG15]. Change [EIB^{+}11, Nim15]. Changeable [HB14b]. Changed [KP10b]. Changes [JSS^{+}17b]. Changing [BS18]. Channel [AJ15, AEE15, Ade13a, CK13, BJ12, BSZ13, CP14, DJJ13, DA18, ETG17, EN13, FKFU12, GDJ12, GG12, GDD13, HS15b, IS21, KA12a, KA16b, KSZK15, KDTD11, LDB^{+}22, MN15, MZEE13, MA17, MK14, Nim15, PKJ15, PM13, PAH17, PS12b, RMG13, SE12a, SSG11, SGD13, SNJ16, SSU10, UCNC13, YR12, Dar12].

Channel-Gain [ETG17]. Channels [AH12a, AMN12, AAaN15, BR11b, BFD15, DM14, EE13b, Mal10, MA12, NAhA13, SS10e, TR15]. Chaos [AS12b, ATA22, AHS15, MKB13, RS14b, RJA23, SA12d, TRSK15]. Chaos-based [AS12b, ATA22]. Chaotic [AS12b, AhAT13, AQ15, AH14, ASC14, Baz18, ESE16, HZ14, HS15a, IPR15, MN15, MAE16b, MSA17, MAE19, MBL15a, Raf14, RAE15, TJ12, TC16, ZEfNF13]. Character [BG13, CJK17, Dum20, EREM11, JKC11, KSRS12, KS15a, KA12d, PRT^{+}16, YM12]. Characteristic [EEL11, Ena12, SS17c, AA11b]. Characteristics [ASB12, AM14, BR13, CCG12, HA16, LG11, ND15, RSB12, PVL11]. Characterization [EB21, GM15a, KAP15, MAaM14, PVL10, SM13a, CS12]. Characters [AA17a, ME14, OMF^{+}14, PC11b, SS12g]. Charcoal [LSM16]. Charge [SD18, TAM13]. Charger [MA16]. Charging [KTS17, MA16]. Charging/Discharging [KTS17]. Chat [HJT14]. Chebyshev [JV14]. Check [BD10b]. Checker [ASH16a, MAt12, SBKS16]. Checking [BBC12b]. Checkpointing [ARHERZ10]. Chemical [AEA17, HD14, PD14, PS12a, SK15b, YJ11, VJR11]. Chemotaxis [KJ12a, YAT15b]. Chest [RRB12]. Cheybshev [GC18]. Chi [Ae14a, SS23b]. Chi-Square [SS23b]. Chikungunya [AR16]. Child [PKJ14]. Children [ASH15, DDBD22, EFA16]. Chinese [JKC11, STM13]. Chip [CTG10, Fla18, JM15, KRMB10, SMSC14, Sub10]. Chittagong [APU16]. Chloride [KNV10]. Choice [AZY18a, AA18b, HBA13, HA17]. Cholera [UN13]. Choosing [KJ14, WG14]. Chromatograms [AAR13]. Chromosome [KSS13, KSL^{+}18]. Churn [AST11, JBT23, KHK13]. CICO [BD10b]. CICO-BT [BD10b]. CIECAM02 [PWG15]. Cipher [AS14d, BRD12, EEGHS15, JP15, MZEE13, MC17a, SA15b, ABA^{+}10, ABA^{+}10]. Ciphering [RAE15]. Ciphers [De 20, DXT23b]. Circuit [AB12a, Cha17, GSN12, SOME16, SM12a]. Circuit-level [AB12a]. Circuits [ARS13, AMB11, AC13, Jah18, KKT15, MKWK15, NRRM12, SPK16, YWT12]. Circular [DRS15, GB18, Hus15, KS17, MTJ13, PK12b, PS15c, RT20, RPR12, SAY16, SM14a]. Circular-Linear [RPR12]. Circulation [EAE19]. Circumventing [PKJ15]. CiRM [RMH20]. Cities [RKBB14, TVS19]. Citizen [BBB16, RMH20]. Citizens [SS23a, PWA12]. Citrus [OOO19]. City [ATAAAF19, AAAA21, MBM14, Rah12, RD17a, RMH20, TMC17, APU16, PTPS19]. Civil [KAFS15]. Civilian [ABL17]. CLARK [HA16]. CLARK-Y [HA16]. Clarkdale [MFM12]. Class [AYAB18, Ade11, BBA12, CWC15, CZW15a, CZW15b, DP15a, Fad18, FTBA16, GJM16, JV17, KHK13, NS13a, OLT14, PMA13, TG18]. Classical [GF17, MVC10]. Classification [AAEK18, Ae14a, AD14a, AHBH11, AHBG19, AHBAR^{+}11, AlS24a, AJI15, ASH15, ASH16b, AMG^{+}12, AAG13, AV13, AS18d, AMJ16, Ba13, BK14, BIE^{+}12, BS21b, CS14, CS16b, DSP^{+}21, DP15a, EAE19, El 13, EME15, Elc11, EIB^{+}11, ESAA18, eFKS13, GJ15a, GB11, GMS12, GJM16, GQD^{+}22, Hem22, HS24, HIS15, JV17, JSK10, KSA20, KV14, KBTB12, KB16a, KA10, KTDG12, KSS14a, KST16, KKDC12, KM16d, KM15e, KSSM24, LsAAA13, LR11, LT12, MNB13a, MSD12, MJ12, MV12b, MEKS15, MTM12, NmBV13, NDC10, NM20, NP15, NOO16, Nur11, OOO19, PP13d, PS13, PBVD22, PK16, PPP22, RV16, RAR24, Rez12, RACI24, SDS20, SE17, SH23, SN15a, SSA11, SSS15b, SPL15, SAS13, SKC15, SAM14a, Sud10, SGJ12, TT14, TN14a, TS12, TS17, Tes15, TK15, TK^{+}17, TY13b, UNM18b, US15]. Classification [Upa23, VM12a, YWT12, YRV14, YRMB15, YRA23, ZJR^{+}19, ZRP16, ZPAQP11, Zyo12]. Classificaton [EAB13]. Classifier [AST11, BG16, BSS15, BO21, DS19, FA13, FRR11a, FRR11b, GS16a, Gho15, Hab13a, HAAT16, IEME12, KST16, KS13a, MM21, NP12, ND11, NS13c, PK15a, QJ13, SJA13, SEF17, SJ12, SS23b, TB13, TK15]. Classifiers [GJM16, Hik13, PFS12, Sin16, SB12b]. Classifies [KB15]. Classify [ARAA11, AkC13, GS16a]. Classifying [IU18, SN16]. Classroom [JHP12]. CLBP [AHBH11]. Clean [AMT^{+}17]. Cleaning [AA12g]. Cleansed [RAR24]. Cleansing [BNMM21, SRA15]. Clearness [ANAK17]. CLEFIA [BDC17]. Click [TY13b]. Client [Ade11, ARL16]. Clients [MSR16]. Climate [Amb24]. Climbing [FM20]. Clinical [AA14b, BML22]. Clinics [SP11b]. Clipping [HO15a]. Clique [AMS13]. Clock [HPM10, IOO20]. Clone [AV13, EmErB13]. Closed [BRS16, CSM21, DAG16, JV14, MP15, MB12]. Closed-Domain [CSM21]. Closeness [SU14]. Closing [BWH15]. Closure [TR13]. Cloud [AC18, AHH13, Abd14a, AAC16, ASC15, AM17a, AKN13, ASASHF17, AMA14c, AAZN21, ASS15a, ARL16, AG14, AAMA20, AZ17, BBM14, BB14a, BM13b, BH14, BADR16, Bat18, BL15, BR15, CMKS15, CS15c, CDC18, CKA18, DP16, DH21, EER14, EEH14, EzHEgT13, EMS19, EEEE16, ED19, EHI17, EEM12, GM16, GEAE15, GHD17, GRL20, HAS20, KKR18, KM13b, KAS13, KAR15, KSMM15, KC15b, KG15c, MB15a, MK13d, MC12, MR15, MMH12, MMHP17, MBMA11, MEH14, NJB16, OAA14, OOC18, PPP16a, PP13c, SP16, SSM12, SP13, SY17, Sal19, SOOA13, SMS16, SG13a, SS14e, SVJ15, Sri12, Sub20, SKGP12, TAM14a, THAP19, WDAJ15, YEE14, ZAHH16]. Cloud-based [AAMA20, CKA18, ED19]. Clouds [EEM12, IBS16]. CLR [SHR11]. Clump [UNM18b]. Cluster [Ade13a, BA15a, BT14b, GSD14, Ish12a, KS14b, KMMM12a, KMMM12b, KE14, MM14a, MAMA16, NV12, PKK18, PB23, RD13, SHHL13, SFH17, SSS16, TBKG13, TB14, Ver17, YMA16, CS11]. Clustered [AMG13, ETG17, ETI14, HST19, SHHL13, Zyo12]. Clustering [AHA14, AkAA14, AlS24b, AS12c, AA11a, AK14, Anh15a, BS12b, BS21a, Ber11, DH17, DSD15, DSA11, DS12b, DS12c, FH13, GMM21, GNN16, GMS12, IE12, IM12, JS12a, JG12, JVG15, MPR14, KSS12, KAAM11, KPAB14, KT15, KWM16, LR12, LSS20, LSSD20, LT12, MMP21, MG14b, MEAT11, MMR14b, MNB13b, MIS12, MJS15, MAMA16, NDC10, NP15, OAR16, OO21, OHI15, PPM12, PK15b, PSM15, PZN21, RBN^{+}16, REA12, RB15, RL14, SP18a, SS14a, SF15a, Sal17, SK18, SSS13a, SC15a, SP12, SK13b, VDS12, SSY11, TBN^{+}15, TM13, VN10, YRMB15, YRA23, ZYM13, ZD15, ZRP16, WG14]. Clusters [Cha15, PS15d, TS17, ZPAQP11]. CM [SSA11]. CMN [RSHI12].

CMOS [AC13, Kei14, KN12b, LGM12b, Nah13, PBK14, RJS^{+}13, RSK15a, SMT19, SLK^{+}13, TMS15]. CNC [SM14a]. CNN [RVGP20, SDS20]. Co [DH19, FSER20, FR15, Jah18, MGB11, PS12a, PPP16b]. Co-design [DH19]. Co-evolutionary [Jah18]. Co-occurrence [MGB11, PPP16b]. Co-operative [FSER20]. Co-Simulation [FR15]. Co-Substituted [PS12a]. COA [AA13b]. Coalitional [JM14]. COBIT [RR22]. Cockroach [ED18]. COCOMO [GJP13]. Coconut [LSM16]. Code [AR10, AS16, AHAT16, BM13a, DBKN20, DKP21, EmErB13, FPS16, GC24, GVW24, GSV20, HAH15, KG11, LPP12, NI15, SS16a]. Code-Level [DKP21]. Codebooks [SSAK10]. Codecs [IRT11, MSP12, ARK13]. Coded [AEE15, AB12b, KAA13, PDK12, SR13b, Mah16]. Codes [AAGM15, Ish14, JS12d, KSBM17, KBM22, MA13a, Oud13, PDK12, SS10e, Tha11, VK11]. Coding [AmAr14, ArAn13, FKFU12, Fla18, GRP13, JKS^{+}12, KTDG12, MZEE13, MNB13c, PB19, PAH17, TR15, TAM14b, TR14a, Tha11, THAP19, ADZ11]. Codon [Bar14]. Codon-based [Bar14]. Coefficient [AS18c, MSEEs13, SU20]. Coefficients [AS12a, DB10, MSM15, YT15]. Cogeration [Car18]. Cognitive [AH12a, AK18a, AVRN12, AA14c, CP15b, DT14, DKP21, DCM15, DS13, ESE13, EF15, Elm22, FAZ13, FHM16, GKM^{+}18, GMRK12, VNM15, JBK13, KAA13, KIA15, KT15, LK15, MTD14, MD14a, MFE13, MA17, PPB15, PKJ14, RJC17, SIAI14, SG14, SSSS12, STD14, SRS14a, SZS15, SJV^{+}15, SS17c, TBKG13, WH21]. Cohen [PMA13]. Coherence [KKM18, Sam13]. Coherency [CLBA^{+}12]. Cohesive [DMS13]. Coils [SVB13]. Coinfection [ZRE15]. Cointegration [Mal13b]. Cold [MHE15, MRH15]. Collaboration [BADR16, MAMS12]. Collaborative [HN16, KK10c, NK14b, OEA18, SP15b]. Collapsed [KBBM15, Rez12]. Collect [RBO19]. Collecting [AMT^{+}17]. Collection [AsRA^{+}12, GGI14, KAA20, UR16, VD10]. Collections [OFE15]. Collective [HMMN13]. Collector [IM15]. Collectors [AN16b]. Collision [Ibr16, KA13b]. Collision-Keys [KA13b]. Collocation [VR12]. Colon [ND11]. Colonies [HMC12a]. Colony [AD24, AA14a, AAA14b, ASG11, AF15, BH17b, FK12, HSMb12, JSY14, JK14b, KS15c, KSK14, MB15b, NO18, PCR12, PC11a, SK13a, SHR14, SGS15, SVJ15, Tah17, TAH13]. Color [Anh15b, AP15, CHW16, DMM14a, GP15, Gum19, HK18, HCL15, JS15, KA16b, KTDG12, KH12, KR14, KA13a, KS14c, LP16, MA12, MM21, MSA16, MSA17, MAE19, Moh13b, MMK11, NAZ17, Neb13, Pil13, PWG15, RPR12, SH12, SRS12, SMR16]. Colored [GH14, MS10a]. Colorless [RK19]. Colour [AGRR10, SCH21]. Column [HKE^{+}14]. Combating [AZA23]. Combination [BRD12, BO21, DB17, EAEE13, GJ13a, GNN16, HB14b, IPR15, RL14, SH12, VJR11]. Combinatorial [BAS12a, Cha14]. Combined [ASH16a, CR12, CJK17, GV15, GDJ12, HKF12, ITK21, LSSD20, LLSL12, MIAEFS15, MAH15, SNB18, ST14, YMR^{+}12]. Combining [Bef18, DB12b, GDS16, MPR14, MNR14, PFS12, SB12a, TBKA13, WRJ23, VS10]. Come [SSN18]. Comment [KV16b]. Commerce [Bej16, Sri12]. Commercial [PSG12]. Commit [MG17a]. Commitment [Oud13]. Common [AER18, JPP12, SPSR12]. Common-mode [JPP12]. Communicating [EAM14, Kor23]. Communication [AEE15, Ade13a, AR17a, Ali12, AEKA11, AD14b, AM10, BSZ13, CSP14, CSN15, EEE14, EA18, ETLE13, EAa14, ESE16, Fla18, GSR21, Hir12, IEG14, JA16b, JD13, JTS15, JYP10, KMS15, KNGT10, KG11, KD15, KRR15, MT13b, NAhA13, OJO11, OHI15, PVR20, RSSV14, RRR11, RMG13, RS13d, SRSV14, SHS23, SSC11, Sta22, Sul17, SA14, SS23b, TCS14, TO13, US14, VM14b]. Communications [AAaAD12, FHM16, HS12, JA16a, KSS^{+}14b, MM15c, SIAI14, TSCD15, YYS^{+}23b]. Communities [BOB17]. Community [AEk13, AnAA14, AS18d, MOOOOMT18, NP15]. Community-Aware [AS18d]. Commutation [VGJ12]. Commutative [SSV11]. Compact [FSG17, HPM10, JD13, MDM17, PA17, PS16a, SSAZ13, SKJ15]. Companies [JM18, TUS20]. Company [KAFS15]. Comparative [Agg17, ABA^{+}10, AAA^{+}17b, Ans14, AK11, AC13, BEE13, BA11, CV13, CKD12, DC22, Eba13, EMS19, GBPTEQ^{+}13, Ghe16, HS11, HE15, Hik13, HB12b, Jag15b, JSY14, KBTB12, KS10, MAI12, MAP13, MRG19, MVC10, NJAK15, NM14, OON17, OJ11, PN12, PD14, PB16, PG11, PAH17, RVKB12, RKA13, RH15, RS11, SDB15a, SJA13, SSS13a, SS10e, SAK13a, SG16a, SN10b, SZUSNK19, TCS14, UA14, VSN13, YAE11, YAT15b, ZD15]. Comparatively [AHZ14]. Comparator [GHBN12, SSK13a]. Compared [MAM16]. Comparing [MR11, She12]. Comparision [DGPP15]. Comparison [ALO17, ARZ15, AA12c, AD13, AAAA17, Alg19, BB12b, Bou13, DL19, DSM^{+}20, DUP13, EME15, EE13b, FS11, FkKB12, GDD13, İHŞ17, Isu16, KR17, Kar11, KTBK12, KMS14, KP13, KK10b, KKK13, KWM16, Lah11, LDB^{+}22, MS10b, MRJN14, Ngu16, OMM14, PG15a, PB19, Pet23, RKGK15, SN15a, SPC13, SS10d, SG12a, SGD13, SA15d, SAR16b, SJ14, SA12e, SK15c, SKGP12, SA14, TA10, Wis13, XL21, YT15, YMR^{+}12, BR11a, NS24]. Comparisons [MAI15]. Comparitive [MHV^{+}14]. Compartive [KRH13]. Compatible [EKA13, KEBF13]. Compensation [AY18, MRP15, RMS14b, Agn15]. Compensators [mH14b]. Competencies [MJP15]. Competency [SAH12]. Competitive [ASG11, AMB11, AMTM11, BS11a, DPT14, DB13a, MM14a, MTM12]. Compilation [Ahm14a]. Compiled [RAC12]. Compiler [SFAS19]. Compilers [Ahm14a]. Complement [NBR10]. Complementary [MTJ13]. Complete [RS13c]. Completion [AMTM11]. Complex [BOB17, CKD12, HMC12a, NLC12, PCD14, PA19, SC19a, SS15c, SW17, TO13]. Complexity [DKP21, GC12, MMA13, MB13b, MSS13, PT13, TMP11, TOHS14, WH21]. Compliance [DRBR24, GPB22, HPC13]. Complicated [Mos14]. Complier [BK10b]. Component [AT17, CT10, GJ15a, HS24, KVN10, KJKK12, MKK13, MB13b, SPSR12]. Components [CBhM12, Es12a, HS15c, NP12, RH14, TO13]. Composite [AD15, AHZ15b, ADVB10, EH16, HAAA15, IFT14, RMS14a, RKGK16, SEE13]. Composition [YEE14]. Compound [AHBH11, mH14b, KRR15, RRR17]. Comprehensive [GMS12, HK18, HMGV12, ISM^{+}17, PK15a, SWP19, WRJ23, YT15]. Compressed [ASII11, KPV13, KVKR16, NBF14]. Compression [AAY21, AR10, Ak13, AmAr14, Al 14, AE12c, Bha17, CFB11, DR14, EUT17, GB18, Hos12, JNR19, JP15, JS18b, KFEN21, KR17, KAAA13, KDS10, MB11b, MM21, MBMB12, MH11, NB11, PB14, RFAH14, RS11, SP11a, SGM^{+}11, SCL14, SL14, SD12, SAk13b, SK12b]. Compressor [HB12a, SSS13c]. Compromised [HZS14]. Computation [AK18a, DB10, Muk13, MB16, Ori14, SHHL13, SA13a, SQ18, SPB12]. Computational [AOAG13, AAO22, BT14a, BBB16, DDBD22, FDG17, GJS17, KM15e, MB12, Olu14, PT13, SGKS12, SS12g, TR14a, YAT15b]. Computed [KR11b, RVGP20]. Computer [Abb12b, AAN12, AARA19, Add12, AAI13, AYE13, Akg13, AAR14, BADR16, DSP^{+}21, Eja12, ESB13, GK13, KAL15, KM12, LP16, MBA20, Muk13, MMK16, MRG19, NM13, PBVD22, PPR^{+}12, RRB16, SM14a, SM18, She11, SP15e, SSC11, SR14, VR14, Zhu21]. Computer-Aided [PBVD22]. Computer-based [MMK16]. Computerized [GYP^{+}15]. Computers [KE14]. Computing [AHH13, Abd14a, AA13a, AM17a, ASASHF17, AAEA16, Alh12, AM12b, AAZN21, ASS15a, ALH^{+}24, AZ17, BBM14, BH14, BADR16, Bat18, CS15c, CKA18, DP17, DKP10, EER14, EHSA23, sMTEkEdAr12, EMS19, EHI17, GM16, GEAE15, GB15, HAS20, HSK14, IMM12, KDBT12, KM13a, KT12a, KM13b, KAR15, KSMM15, KG15c, LMZ12, MK13d, MC12, MR15, MT13a, MAM12, MBMA11, MEH14, NJB16, NOO16, OAW12, Pat14, PPP16a, RR16, RS13b, RBO19, fSG13, SSM12, SP13, SY12, Sal19, SMS16, SG13a, SS14e, SS16b, SVJ15, Sri12, Sub20, Tha11, VSA15, You13, ZAHH16, ZKN10, OOC18]. Concealment [FIS16]. Concentration [AK11, HD14, Kal11]. Concept [AAA21, GG11, LSP17, NHMR13, OOJ17, PR15, PRR15, PSG12, PB14, RV13, MNB^{+}22, TTA10, TUS20, UDA13]. Concept/Ontology [NHMR13]. Conception [FAA13]. Concepts [NM19]. Conceptual [AZY18b, AMS14, GS12b, MKM17, OSS14]. Concerns [GAZ13, Sab16, SG12b]. Concise [AKH15, SSP13b]. Concrete [RAC12]. Condition [HIU11, KIH15, KJK12, NM13, PB12, SMVK11, SB11c]. Conditional [HST19, S22, SRS14b]. Conditioner [AKB16]. Conditioning [Fer14, NAD19, RMP17, GBM10]. Conditions [AN16b, HJ13, ML13, PM13, RHM15, YRV17, YB12, ASNS10]. Conducted [JPP12]. Conductivity [BH17a, EH14, HD14, Haz14, HGK15, ILG15, SH17, TH15]. Confidence [Ask14]. Confidentiality [KK13c, MM11a]. Configuration [BS12c, DKS14, HZLL15]. Configurations [SZ15]. Confining [AS14a]. Conflation [PWG15]. Confronting [CHG12]. Confusion [AlS24g]. Congestion [AAN12, ASY17, AQ14b, AAD13, BSG15, HEs12, IAH14, IMIA11, JTS15, MG14a, NHH10, NKM11, RMP11, RSP11, S13, VM14a].

Congruency [ASK15]. Conjecture [MB12]. Connect [PJ12]. Connected [ER12, HE15, JPAK11, JJN14, KETESH17, KIH15, MKK13, RRB12]. Connecting [AM15]. Connection [KSS13]. Connections [PIB11]. Connectivity [AA13a, AAAA21, AGL17, AT15, DNP15, Far17, NS16]. Connectivity-based [Far17]. Consecutive [EEGHS15, MHE13]. Consensus [KJKK12, ZEF17]. Consequences [MBMA11]. Consideration [Ota13]. Considerations [AKK11, RGS15]. Considering [AMN15a, FR12, MHE15, MRH15, MJM17, MB15b, NVR12, RD17a, ACK14]. Consignable [ZJR^{+}19]. Consistency [GE17, PP17b]. Consistent [MMR14b]. Consolidation [KSMM15, MC12, SMS16]. Constant [EN13, QT15, SSMP13, TH15]. Constant-Time [QT15]. Constellation [GB11]. Constrained [BM13c, KAS12, OC13, PDK14, PBG15, Tah17]. Constraint [GJ13a, HES14, KAA13, MJM17, SAM14b, SSY11]. Constraint-based [KAA13]. Constraints [BRS16, CTG12, EIB^{+}11, Ker13, SB12c, VGo12, VM12b]. Construct [Wu23a]. Constructing [AaEeK11]. Construction [AZY18a, AIRR19, BP11, CJC19, FRR11b, IHD12, NK15, ONK24, OS16, RBT10, RS13d]. Constructive [MMR14c, SM12b]. Consultation [AIP18]. Consumers [Far13]. Consumption [AII13, AkAA14, ANAH20, BKK13, DAB12, EE16b, HB12a, HBS12, Kei14, KMMM12a, KMMM12b, MBM14, MFA13, NJB16, SDG18]. Contact [EAMD14]. Contactless [AA11b, MMG16]. Contacts [HB14a]. Container [CD13, RK21]. Containing [PB15]. Content [BAM13, CJK16, EME13, HWT^{+}13, HR14a, JA12, KDBT12, KR14, MB13a, NRV12, PWG15, RMS12, RO11]. Content-based [BAM13, HR14a, PWG15]. Contents [NAK^{+}14]. Context [Anh15b, BM14b, BK10b, BOK15, CSKB11, EnFn14, KVR14, MB13a, MM21, RS11, SAR18, SAM14a, Zar24]. Context-Aware [SAR18, KVR14]. Context-based [EnFn14]. Contextual [Alt11, EIB^{+}11, EE16b, yGL13, IHH14]. Contingency [SMVK11]. Continuation [Sou15]. Continues [TR12]. Continuous [HJG^{+}13, HB16]. Continuously [PCKD17, SH17]. Contour [AMK12, KKR^{+}14, PMP12, PMH10, UR12]. Contourlet [CT10, HAAT16, NP11, ZvAS12]. Contraction [WWSS12]. Contrast [ASK15, BMM12, EEEE16, Hal13, KT14, Kau15, MSS14, LG13a]. Contribution [AMAB15]. Contributions [Sul17]. Control [AAN12, Abd18, AHAS^{+}17, AA12a, ANE^{+}13, ASY17, AAEEA15, AMA15, AhAT13, AmM12, AZS16, ASN12, AQ14b, AAoS14, AR16, Als18, Alz12, AAD13, AYM16, AVS13, ABS^{+}12, AK11, BSG15, BFA13, BG17, BSP11, Bou13, Bou12b, CHW16, CP15a, CQM15, CPSP13, CWC15, CZW15a, CPD13, CA16, DK13, DK11, Deb20, DBW19b, DBW19a, EUEI23, EJC16, EOO22, ES12b, FZM12, GM20, GAA15, HAA16, HES14, HB13, HZ14, HZK12, HTM14, HDMK14, HSP10, HR22, HS10, HHS15, Ibr14, IHH14, IAH14, IMIA11, sNIALY13, JV14, JBK13, JTS15, JTA19, KTN12, Kal11, KM15b, KMB14, KN17a, KCS14, KN12c, KNR15, LEE13, LSM16, LRL17, MBKC15, MNB13a, MNB13b, MM17, MBZA16, MBE16, Mos11, Nad14, NVR12, NEA17, NR12, NAD19, NM16, NMAB14, NU18, OS16, OA12]. Control [PKDS17, PCW10, RMP11, RK21, S22, SHSA12, SZ15, SDZ15, SKM16, SND13, SM14a, SK15a, SO14b, SR15a, SK12a, ST12, SOO15, SSP15a, STE14, SWH16, SK10c, SPXT17, TAM14a, TAM19, TY15b, VMT13, WDD16, YMR^{+}12, ZRE16, VJR11]. Control-SMC [ANE^{+}13]. Controllable [HMAA13]. Controlled [AH16, CS15b, KN12b, KR10, KRM12, MP12a, Ngu24b, OCZ13, PMK12, RRA17, RN24, SPP13, SNB18, TVY17, VM14a, Ver16, DSA11]. Controller [AM17c, AmH13, ATM12, ANE^{+}13, ASS10, AA13b, AAEEA15, AD13, AMA15, AAA17a, AW20, AAA14b, AHK12, BM13c, Eja12, EHHM17, FAK14, HB13, HTPB11, HR14b, IHH11, Iss13, KNI14, LT13, LGZX17, MLRU12, Mis14, MBZA16, MBE16, NVR12, NA18, NK14a, Ngu16, Ngu23, Ngu24a, PCKD17, PC11a, Par14a, Par14b, QA14, SFA14, SM17a, SND13, SM14a, SO14b, SR15a, SGB16, TUS20, TNC10, TArM14, Tua17, USC12, VGJ12, WDD16, YSH12]. Controllers [AS17, AKK11, AK11, HR16, HAA16, PCW10, SK15c, TOMGBSS13, Wu23a]. Controlling [BBMA13, GRS22, Kad16, NK14a, RKE23, SGA13, You18]. Controls [GBPTEQ^{+}13]. Convection [BH17a, BWS16, CP14, GDJ12, HY13, HR13, HJ13, KM16a, Sal15, SM17b, SJK21, WBS16]. Convective [DGJ12, DJJ13, Haz14, ILG15, SDJ13, TH15, TRM23]. Conventional [AK11, MRJN14, SNJ16]. Convergence [AAD13, CKD12, CRA12]. Conversion [HB13, Isu16, JJN14, KMB14, MCaC14, Rab14, SU18, SS12b, Win14]. Converter [AM17c, AKB16, Ena12, HR14b, IAM16, KM15a, KTS17, MLRU12, MDP12, MH12b, RURS13, LGM12a]. Converters [AM14, MAI15]. Converting [Kha13]. Convex [HSS11, PK19]. Conveyor [Mug24]. Conveyors [BKM12]. Convolutional [ACB17, AM19b, BV21, FSNR18, KSSM24, MGL22, Obe23, SS21]. CONWIP [TB12b]. Cooling [AN16b, EN13, Job17b]. Cooperation [FKFU12]. Cooperative [BADS17, DCM15, FM20, GE17, Ish12a, KAA13, KT15, KK16b, LKS14, PPB15, RMG13, SHS23]. Coordinated [AS15d, PC11a]. Coordinates [SM19]. Coordinating [AOAG13]. Coordinator [HS15d]. Copula [Poo22]. Copy [AE12a, DDG16, SB11a]. Copy-Move [DDG16, SB11a]. Copying [MBCL17]. Copying-Intensive [MBCL17]. Copyright [AH17b, AS13c, DBB24, MA12, YBA12]. CORBA [She12]. CORDIC [NBKM13]. Core [Akg13, Am13b, AB13c, CDC18, HTPB11, OA15, RPV10, SK14a, Sar13, LB11, OA15]. Corner [AK18b, PR16, SG10, TL16]. Corona [DNP15, DM18]. Coronavirus [DPS21]. Corpus [GK10]. Correcting [Ish14, MA13a]. Correction [AD18, DKB12, Far11, FPS16, Fla18, JKC11, MS10a, RTK15, RNM14, Zaf11]. Correctness [EnFn14]. Correlated [EE13b]. Correlation [AlS24d, ASB12, AS12e, BNS17, CKA18, MIS12, PR16, Uzo13]. Corruption [ASN11]. Cosecant [CR15]. Cosine [AlS23b, NA20, PKP13, PCD14, REA12, TK15, YT15]. Cosines [JFJN20]. Cosmological [AHH20]. Cost [AG10, AM12b, Ans14, ASK17, AG13, BM13b, DDC12, DS12b, GRL20, KHAE19, KS15b, MHE15, MG12, MKWK15, Nur11, Ori14, PGPM23, Rab14, Rah14, Sar16a, SM12b, TY13a, GBM10, KAS12]. Cost-Benefit [MG12]. Cost-Effective [BM13b, PGPM23]. Cost-Throughput [DDC12]. Costrank [HMY12]. Costs [OHI15]. Cotangent [JFJN20]. Cots [KJKK12]. Cotton [PSA18]. Couette [CDC17, DSJ11a, DGJ12]. Counter [JT11]. Counteractive [Tal11]. counterfeit [ZCZ24]. Countermeasures [ZUDS16]. Counterpart [RBB11]. Counting [PA13]. Countries [MAMA16, SOOA13, UAMN17]. Coupled [AS12b, CS16a, El 20, GGX12, SVB17, WOBK18]. Coupled-Neural [El 20]. Coupling [NVR12]. Coupon [TE16]. Course [DRFF15, ESZED15, IGS17, JHP12, Kas17, PP12a]. Courseware [AMS14, NAH11, Sub24]. Cover [LMZ12, MAE16b, PPP22, SS13c, ZAAS18, AMJ16]. Coverage [DBM14, EBO13, GVW24, HAH15, KNS17, MDCS22]. CoveringSet [AA12f]. CoveringSet-based [AA12f]. Covid [KNS20, KSA20, SDS20]. Covid-19 [KNS20, KSA20]. COVID-CAM [SDS20]. Coworking [PW23]. Cox [VV14]. CP [Hos12]. CPL [SN14a]. CPLD [HPM10]. CPU [AAS14, Hus16, MB15a, MAM16]. CPU-Scheduling [Hus16]. CPW [KCD13]. CPW-Fed [KCD13]. CR [VSA15]. CR-OLAP [VSA15]. Cracking [Bou13]. Cracks [JBP16]. Crafting [KC20]. Crawler [HK10, KP10b, KN11, MMTS13]. Crawling [ETD13, SS13d, SRM12]. Creating [Far14a, Far24, MM15b, MM11b, MG14c]. Creation [BAM13, CHG12, USK17]. Credit [Gho11, IPB14, LKS14]. Cricket [Kha11]. Crime [ABAO24, KC20, MM18]. Crisp [SKA13]. Criteria [APU16, AMAB15, AKG20, RR11, STG13, Sal14]. Criterion [KNT11, MP12a, Nad14]. Critical [am12a, AA14b, AMS14, BKB11, IAQ11, PPP11, Raf15a, RPV10, SG16b, SSZ16]. CRM [HS18, RHT19]. CRN [BNS17]. CRONE [YMR^{+}12]. Crop [AMJ16, RÇ16, SCH21, TK^{+}17, WKG^{+}18, VJ24]. Cross [AAY21, CK13, AHAPT15, BLQ21, HKU19, HKL16, KN12c, KD15, LM20, MBR16, NVR12, SJ12, SSS15b]. Cross- [MBR16]. Cross-Coupling [NVR12]. Cross-Entropy [SSS15b]. Cross-Layering [AAY21]. Cross-Platform [BLQ21, HKL16]. Cross-tier [HKU19]. Cross-Validated [LM20]. Crossbreed [BG16]. Crosscutting [SG12b]. Crossover [HB14b, OJ11, SJA13]. Crosstalk [Fla18]. Crowd [EEM12]. Cryptanalysis [KK15a, SS14b]. Crypto [JP15]. Crypto-Compression [JP15]. Cryptograhic [UC15]. Cryptographic [De 20, aKNP^{+}15, KT12b, RJA23, US14, AAAOG20]. Cryptography [ASH16a, AC18, AQ15, AJ16, CM12, DXT23a, De 21, DBB24, HMB13, JP15, JM16, KV15, KVS16, MNYMGG12, Olu14, Red19, SK16, SSA14, VS10]. Cryptology [GSR21]. Cryptoprocessor [BM15c]. Cryptosystem [AC18, AT17, BM20a, IN15, RMSS13, WDAJ15]. Crysanthemun [MKT17]. Crystal [PN10, RHU12, SAS21]. Crystals [KJA15]. CS [TJ12]. CS-QCSK [TJ12]. CSFs [RRudS12]. CSMA [ASNS10, JT11]. CSMA/CA [JT11]. CSRmesh [ZYW^{+}18]. CSTR [AK11, Kal11]. CSUA [DKP10]. CT [AD24, AHBG19, AKS^{+}13, KKR^{+}14, RRB12, SErEs^{+}12, SSC15b, Zaf11]. Cubes [RTM12]. CubeSat [GB11]. Cuckoo [Aly13, AK14, Ghe16, RS13b, RKEdK14, SV14b]. CUDA [AA21, JI10b, KEBF13]. CUK [AA11a]. Cultural [Ari23, KTA17]. Cup [AP15]. Curbing [OOC18]. Currency [RD17b]. Current [AMBS16, BKM12, GDJ12, KN12b, Liu24, MA16, MH12b, ND17, SWP19, SC15c, SVB13]. Currently [DL19, DSD15]. Curriculum [AARA19]. Cursor [JP16]. Curve [AC18, Abd17, ABB20, AlS24c, AJ16, CM12, DBB24, GSR21, JIA13, KVS16, MRJN14, MNYMGG12, SSY11, De 21]. Curvelet [KP15a, PGS13]. Curves [HO15b, ME14, WDAJ15]. Custody [PS15e]. Custom [Elm22]. Customer [AST11, HS18, JBT23, KB16a, KHK13, LWL^{+}23, PUW13, PS21, RK16, RHT19, SS15b, TSE17, TB12b]. Customers [KC15a, RK16, Sab16, NS24]. Customization [ESB13]. Cut [AK18b, KGK13]. Cuts [HMC12b]. Cutting [KTN13]. cWRR [AOB12]. Cyber [AYYB15, BC22, GK12b, KC20, KN17b, KC14c, MTKP15, NTL18]. Cyber-Physical [AYYB15, BC22, NTL18]. Cybercrime [ARP17, AZA23]. Cybersecurity [Has24]. Cycle [DM18, DKS14, SR23, ZYL^{+}21]. Cycles [SOT18]. Cycling [GKG16]. Cycloconverter [Alg19]. Cyclomatic [WH21]. Cylinder [SK15b]. Cylinders [SM17b]. Cylindrical [HJ13].

D [MFM12, Sal15, SPC13, KM13a]. D2D [FHM16]. D3LS [EMA11]. DAG [NHMR13]. DAGitizer [AM12b]. Dahlan [SU20]. Daily [Gho15, KDC12]. Dam [AJM^{+}15, FJM16]. Damage [KJA15, OE14]. Damping [Ibr14, OS16, RJC15]. Dangerous [SOME16]. Dangling [Jay12]. DAQ [MZAE16]. DAQ-based [MZAE16]. Dark [OAE16]. Dashboard [AA18c, MREd14]. Data [AC18, AAN^{+}19, ARAA11, ATA14, AAT16, AMHAh13, AHK14, AXKT19, ANM17, Ahm20, AS18a, ADFAJ15, AkC13, Al 14, AA13d, AA18a, Alb15, AH11, AA12d, ASH15, AHH14, AEKA11, AP16, AOK18, AA12g, ASA14, AHAT16, ASH13, AG14, AE17, Asi15, AAOA23, AAS13, MKOS14, MA22, Ba13, BNMM21, BM15a, BAM13, BLQ21, BH14, BS12b, Bat18, Bej16, BBB16, BBL15, BS16, BMD16, BR15, BD18, BC22, BS21b, IU18, Car18, CA11, CPD13, CS15c, DD15, DCH16, DSM^{+}20, DH21, DP11, DS15, DDN14, DKP10, DA18, DRBR24, Eba13, ES15a, EA11, Far17, GJ13a, GSR21, GEA16, GVRP10, GSP18, GJUL23, GMM21, GNN16, GK12b, GENERM16, GB18, GGI14, GB14, HMC12a, HMN18, HF20, HHEY19, HAS20, HHHK21, Has24, Hem22, Hir12, HASH17, HSS11, Hus12, VNM15, IBS16, IHH14]. Data [Ish14, ISM^{+}17, JS12a, Jag15b, JS18a, JPJ18, JARM13, JVF11, JGZ13, KPC16, KBK10, KFEN21, KT10, KBTB12, KP16, KV15, KG15b, KMMM12a, KE14, KK11, KP13, KDC12, KM15d, KM16d, KA17, KDS10, KAS13, Kum13b, KPAB14, KM15e, KVKR16, KR16, LT12, MN17, MBAR12, MREd14, MAI12, MM11a, Mal10, MC13, MG14b, MKSB22, MC12, MM20a, MA12, MC16, MM14b, MM16, MMR14b, MBM13, MAk12, MAM19, MM12b, Mun14, MH13b, NHS11, NP19, NJ16, NRL17, NA16, NM14, NM20, NSS17, ND17, NH13, ND11, NTL18, OP18, OOC18, PRR15, Pet23, RMVP14, RJSJ16, RKR15, Raf14, RBN^{+}16, RHS18, REA12, RPV10, Ras16, RRudS12, RAR24, RA10b, RBO19, SsTM12, SE17, dSEdZ14, SN15a, SMGS17, SY17, SZOK24, SN15c, SPBP17, STS19, SSS13a, SAAM17, SA12b, Sel22, SD17, SGM^{+}11]. Data [SPP11, SYR17, SCL14, SD12, SQ18, SP11b, SEA16, SBM^{+}15, SKS16, SL11, SRA15, SASS11, SMT17, Sub20, SB18b, SVJ14, VDS12, SS23b, TSE17, TK15, TOHS14, TS11, TVS19, TM13, THAP19, UI16, UR16, UDA13, VR14, VK13, VG12, VD10, WC15, Win14, Wis13, YK16, ZYM13, MJ14, PK15b, RPMH24, Nor13, PMP14]. DATA* [HKS12]. Data-Driven [DRBR24]. Data-Link [MC13]. Database [ADFAJ15, AA12d, BM16a, BBL15, Bon15, DPJB12, DS12a, DS12c, EzHEgT13, EREM11, Fad18, FLLR20, GS12b, GH14, KSS14a, KPV13, KVS16, LOAO23, Ota13, RAB12, SP16]. Databases [AAP13, Bar16, CV13, Eba13, KW14, KSC11, SB13a, SC15a, SP15d, SA12c, WAM12, Wis13]. Databricks [KSKS23]. Dataset [BRR16, Gho15, RR16, RP15, SS14a]. Datasets [BD13, BS21a, GPC16, PN11, RD13, SP15c, SW17, US15, SP18a]. Daubechies [JNR19, KP15b, MD14b]. Day [AJT17, JM18, RSK15b]. Day-ahead [RSK15b]. Daycare [MGA20]. Days [SKG14]. DB [BBHL12]. DB-SEP [BBHL12]. DB4 [MD14b]. DBA [KRR15]. DBCrypto [DPJB12]. DBMS [RKY12]. DBPTGMF [KS14c]. DBSCAN [AM13a, KJ14]. dc [MLRU12, AAEEA15, CPSP13, DD21, Ena12, IAM16, JTA19, KTN12, KTS17, MAI15, MBZA16, NK14a, OAA13, SO14b]. DC-DC [IAM16, KTS17]. DC-link [OAA13]. dc/dc [MLRU12]. DCA [BOK15]. DCSL [KAP15]. DCT [EAMAES16, AHS15, Bar16, Dha12, IKY22, MAE19, SYAS11, SAk13b]. DCT-based [SYAS11]. DCTM [NA20]. DDOS [SNP15, AW20, AMG^{+}12, BKSB15]. DE-based [Ngu23]. De-Mystifying [BM15a]. De-Noising [AHM14b, KAP11, RVKB12, SM15a]. DEA [PMP14]. Dead [WDD16]. Dead-zone [WDD16]. Deadline [SV10]. Deadlock [GH14]. Deaf [EEE14, EA18]. Deal [DA12b]. Dealing [AARA19]. death [MN24]. Debit [BOS15]. Deblurring [VGP15, ZYL^{+}21]. Debugging [MR12a]. DEC [YPS19]. Decay [PBVD22]. Decentralized [Ala16, FTBA16, GC24, Hus12, Ira12, WBK22]. Deception [OKS18]. Decision [AA18c, AA14b, AAT16, AKG20, AO13b, Bad13b, BWS^{+}12, Bef18, BBC12a, BV14, CJC19, CD16, Eis13, EFA16, EE13a, FAA13, GHH15, GOCA15, HEEG15, HS24, HAA12, IA18, IR16, IR10, IA15, IAA13, JBP16, KM13a, KB14a, LCA12, MH12a, MR12b, Mos14, MCA14, NMS10, NWS15, PSG12, RR11, RÇ16, SN15a, SS16a, ST11, SSC11, SBC12, SB18a, SVJ14, ZAA13, Sal14]. Decision-Making [AA18c, CJC19, Mos14, Sal14]. Decisional [RT11]. Decisions [AAAA17]. Declarative [Dix11]. Decline [MMPH15, MRJN14]. Decode [GDS16]. Decoded [DV14]. Decoder [DLBM12, Rah14, SC19a]. Decoding [AhAT13, BV14, VNM15, MMA13]. Decomposed [DM22]. Decomposing [PF20]. Decomposition [AAEK18, ELDM13, GM15b, GM15d, KVN10, KK12a, KH12, MG14c, PKP13, QJY11, SMAM15, TS14]. Decompositiong [KAP11]. Decontamination [SGD^{+}20]. Deconvolution [JSC12]. Decoupled [AA12a, ABS^{+}12]. Decoupling [FZM12]. Decreased [MA16]. Decreasing [Ara15]. Decryption [VS10]. dedicated [KTN12]. Deduction [ADFAJ15]. Deduplication [BR15]. Deep [AEEO24, AHM17, DL21, ETD13, KBEA18, NPA^{+}24, ONK24, PPP22, RN24, RACI24, SDS20, SS21, SYR17]. Deep-Learning [SYR17]. Deeply [ENC12]. DeepSpeech [PH23]. Defect [DDDD12, RP24]. Defected [SRM13]. Defects [hC13]. Defend [SN15b]. Defense [SS16d]. Defensive [SU12]. Defined [AYYB15, HZBG13, KA14]. Defining [Dix11]. Deformation [RTK15]. Deformation-Map [RTK15]. Deformed [RTK15]. Degree [AJSG11, MBKC15, TArM14, XL21]. Degree-of-Freedom [MBKC15]. Degrees [SS12d]. Dehydration [AGG16]. Delay [AMNA18, BLA12, HITA20, JA16a, KR11b, NKR14, OMH11, PVA13, MP17, SPP13, SC14b, SMK14]. delegation [KM13a]. Delivery [BM13b, CM13, JKM^{+}17, MMTS13, PCS18, SOOA13, Tha13]. Delphi [Kar11]. Delta [BI12, LGM12a]. Demand [ASD16, KJTM15, KP10a, Mal13b, MS11, MK13d, RMP11, RSK15b, RA23, SMK14]. Demilitarized [KBBM15]. Demographic [GVRP10]. Demonstration [TGcO14]. Dempster [AB14a, RMD10]. Dendritic [BO16]. Denial [AMG^{+}12]. Denoising [Anh14a, Gha13, JT15, KM17, KJK12, RT13, RC15, SHD13, SC15b, SS13e, RV12a]. Dense [DN12, JTKC14, SM13a]. Density [BS12b, Guo18, PIB11, RAJ^{+}12, RSRT20, RRR17, SP15c, SR15b, SR17, TOHS14]. Deorbit [AMN15a]. Dep [CV13]. Dep-Miner [CV13]. Department [SNA13, KSL^{+}18]. Dependence [Ker13]. Dependency [CV13, KW14, MB13b]. Dependent [BMF11, KS18, KC12, MJM17, SMK14]. Deployed [ATA14]. Deploying [RR12]. Deployment [ABYB22, FR12, GRL20, Nta14, Sab16]. deposit [NS24]. Depot [FS11]. Depots [FS11]. Depression [CRBJ12]. Depth [PP15, SYAS11]. Deregulated [NHH10, NKM11, RSP11]. Derivation [KA10]. Derivative [Isl18]. Derivatives [KS15d]. Derived [BYT14, KS12b]. Deriving [JT11]. Dermoscopic [NPA^{+}24]. Descent [AlS24e, AlS24f, AlS24h, AlS24i, KPR14, KS16e]. Describe [EK14a]. Described [SK12a]. Describing [MM11b]. Description [NK12]. Descriptor [TL16]. Descriptors [PK16, SB11a, SB13b, TSA12]. Design [AHAS^{+}17, ATH12, AA16a, AMN15b, AMN15a, AB17, Afr14, Afr17, ASS10, AB16a, ADFAJ15, Aq14a, AsJ12, AAA17a, AIRR19, AMAB15, AAA14b, AB12a, ACK14, AMB11, AAGM15, AA17b, AGZ15, AC13, AG13, AMS14, BH16, BSM17, Ber11, BV14, BK10b, BA16, BA12c, CS13, CPSP13, CM15b, CBhM12, CS11, DK18, DS12a, DDDD12, DR15b, DS12c, EmdHNmAW13, ETKE20, ENE14, Far11, Far24, FAK14, GHBN12, GSN12, GD13, GSN14, GC18, GB11, GE12, GMAR14, HE15, HR22, HTPB11, HKY15, Ibr14, IOO20, IHH11, IhZ12, JSS17a, JV17, JPP12, JVF11, Joh14, KSP11, KUM13a, KDT16, KBBM15, KKS14, hKhHL13, KNI14, KK17a, KM16c, KKK12, KN11, Kum13b, KNR15, Kum19, LGM12b, LEE13, LT13, LAA^{+}17, LWL^{+}23, LK16b, MKM17, MLRU12, MDGS13, MP10, MA15, MDP12, MDCS22, Mis14, MT13b]. Design [MKWK15, MBE10, MBCL17, MI13, MBL15b, MNN10, Nad17, NVR12, NRRM12, Ngu22, Nta14, OJR11, OO14, OS14, PR15, PVR20, PKDS17, PC11a, PPB15, PBK14, Pou13, PP16, PA19, PW23, RHU12, RV12b, RW20, RR18, RT20, RKS17, SE12a, SE14, SFA14, SZ15, SDZ15, SS13a, SMSC14, SM14a, SS10a, SCE14, SN15b, Say22, Sel22, Sha14, SSK13a, SS14e, SM12a, SWH16, SRM13, SSS13c, SSG14, SSK13b, SA12e, SS23a, SLK^{+}13, Tah14, TB13, TG18, TMS15, TRDB24, TO13, TAM13, TArM14, UDA13, UO23, VM14b, VD10, WDD16, WNE21, YJ14, YMR^{+}12, ZAAS18, ZYW^{+}18, ZAER12, ZCZ24, DH19, PA13]. Designed [SS12d]. Designing [AH12c, DPB14, HPM10, KA11, Mah16, MBAR12, NRR^{+}10, RRAFAM19, RM11, SSV14]. Designs [AAA^{+}17b, MM15b]. Desktop [VJD22]. destination [CW23]. Destructive [Ela12, MMR14c]. Detail [ASA14]. Detect [ASA14, BSS15, Cur20, DC22, EAEE14, HZS14, MPR14, PB13, SJV^{+}15]. Detecting [AP15, EAEE15, HHEY19, Has15, HHS14, JBP16, MK13c, OKS18, SO14a, SK13b]. Detection [AD24, ACB17, AOM12, ADZ11, AEk13, AEEO24, AAA20, AD18, AAC16, AM13a, AHBAR^{+}11, AB13b, AIC19, ATB12, AE12a, ASH16b, AKA^{+}18, AB14d, AH12c, AV13, ABAO24, AV18, ARDS18, AS12e, BS12a, Bad13b, BAO16, Bar14, BS21a, BNS17, Bef18, BS14a, BBD13, BD18, BK10a, BOB17, BML22, BKS14, BA15b, BIAE24, BKSB15, CB16, DL19, Dak23, DD15, DSB12, DSDD12, DPS21, DDM^{+}23, DRS13, DT22, DS19, DP11, DDC^{+}24, DDG16, DB12b, EUT17, EmErB13, EAEE13, EAEE15, EIB^{+}11, EA11, Far11, FRR11a, FAA13, FPS16, Fla18, FR15, FTLK14, GD13, GB15, GAA17, GH14, GB14, HM18, HR22, HST19, HIS15, HKY15, HSD13, Ibr16, IEL18, IHB^{+}15, ISM^{+}17, IM16, JAG15a, JZS13, JS12b, JK14b, JSK17, KSKS23, KSS18, KR17, KNA12, KT12a, KSV10, KKDC12, KMG21, KKM18]. Detection [KB16c, Las10, LC17a, LDB^{+}22, MK12, MAE16a, MKK13, MK18a, MDD21, MSN12, MM23, MAARIH16, MIAEFS15, MBP16, MHV^{+}14, NR16, NSA14, NC14, NDC10, NSM13, NP15, NKM12, NK12, OOO19, ONJ13, PP13b, PSA18, PP17b, PM21, PC15b, MN22, PB12, MP17, RPN14, RMP17, RV16, RH15, RGNA13, RC12, RACI24, SOME16, SDS20, SF15a, SCH21, SM18, SS21, SBHM13, SS17b, SM15a, SABA12, SM15b, SB11a, STD14, SF15b, SG10, SS12e, SSV14, SS15b, SS16c, SP18c, SRA15, SB18b, SW17, SS17c, SMR16, SSC15b, SKZ22, TSRM10, TBN^{+}15, TBKA13, TS17, TSY18, TY15b, TOO16, UNM18a, Vad20, VKB14, VKJ^{+}15, VST17, VDM22, WDN^{+}20, WHP14, WWSS12, YM12, YRMB15, ZCZ24, ZJR^{+}19, Zon17, ZTL^{+}14, GSCT12]. Detective [SS23b]. Detector [AH12a, AC15, Jas13, MRP16, MA17, MMK11, MMK16, PR16, TH14]. Detectors [PBG12, SYAS11]. Deteriorating [DMM14b]. Deterioration [TYSS15]. Determination [ADS16, BLA12, CC17, IA15]. Determine [KA11, SZUSNK19]. Determining [Has15, PZN21, SYAS11, SGS18]. Deterministic [BSZ13, BP11, JPJ18]. Devanagari [SGT^{+}11, YM12]. Develop [SS10b, TUS20]. Developed [Dar13]. Developer [MR12b]. Developing [AA18c, AARA19, Al 13, AWK19, AKG20, AB11, ASH13, Asa13, EKES20, ESEMD21, EFA16, Far13, KN12c, LSP17, MKM17, MAMA16, NF15, OJO11, SZM18, SJK21, SOOA13, UAMN17, GP13]. Development [AOM12, ASY17, AB16a, AAO11, ABAO24, AHAPT15, AAA19, AZMP17, AZY18b, BA12c, CGKP13, CDC18, DR15a, EUT17, ENE13, EE16a, Far24, FLM11, FCW14, GB11, GAZ13, GK10, GOCA15, HK19, HPR17, IUN10, IAQ11, Ira12, IA15, JTI19, KRR11, KSP11, KDT16, KAZA13, KMT^{+}24, LcCm13, MKSB22, MKT17, MMTS13, MM15c, MBE10, MZAE16, NJAK15, OMM14, Ojh11, PS15d, Rah12, REUHTG18, RK14, RKBB14, SS10a, SOT18, SR23, SP18b, SS12d, SKG17, SC12, SPP16, Suj14, VDS12, SJ19, SK15d, TTA10, ZAA13, ZK12]. Development-Oriented [FLM11]. Deviance [AMA14b]. Deviation [Jas13]. Device [Ali19, Cai17, Deb20, GJUL23, HR22, HZLL15, Joh14, JKTM15, RW20, SSA^{+}23, SS16c]. Devices [AM17b, Alg19, BR11a, CA16, GE12, IhZ12, MNB13c, Mun14, PB13, Pri13, SET15, KC12, RSP11]. Devnagari [NRR12]. DEVS [MBM14]. DF [Kam17, PG15b]. DF-ICF [Kam17]. DFIG [ASP17, AB14b, HB13]. DFRFT [HKF12]. DFT [SNJ16]. DG [LRRR11, YR12]. DG-MOSFET [YR12]. DGD [AJ10]. DGs [MBL15a]. Dhaka [AAAA21, BWH15, Rah12, SAAM17]. DHCP [RJSJ16]. DHT [SGD13]. Diabetes [AGFA16, FAA13, KBTB12, KW16, Mes23, MJP18, ZK12]. Diabetic [BS12a, DSDD12, Hab13a, Hem22, LAAI23, PA12]. Diabolo [AS14a]. Diagnose [AER18]. Diagnosing [AERE21, GVRP10, JBP16, SN16]. Diagnosis [BBC12a, DPB14, EBR12, EFA16, EE13a, HBN12, JIB14, JSSA20, Jag15b, JA22, KST16, Mes23, OO18, PA12, RVGP20, SASS11, SBC12, SVB13, SVJ14, UN13, VMM14]. Diagnostic [AAA20, MH14a, MNR14]. Diagnostics [PBVD22]. Diagonal [RM16, WAG11]. Diagram [Fad18, KH13, Roc17]. Diagrams [KNS17, MMHA11, OLT14, SSC11]. Dialectal [SU18]. DIC [AH11]. Dictionary [GGX12, KDS10]. Dielectric [KJA15, PS12a]. Dielectrophoretic [DK18]. Diesel [ANAH20]. Difference [AOO20, AH13a, DSJ11b, EEB16, LG11]. Different [ARZ15, AN16b, ARK13, AAqA14, AS18c, AASAA15, AAA^{+}17b, AD14b, AB13c, AC13, BA12a, FA19, GBPTEQ^{+}13, HJE14, Hik13, IRT11, KJK12, MK12, Moh13b, NJAK15, RSHI12, RHM15, RD17a, SSS13a, SAAM17, SM15b, SA15d, SNJ16, SA12e, TBN^{+}15, TR14a, VGP15, YB12, AM17b, CS12, CS10, EME15, ESAA18, Job17b, KK22, MAI15, PAH17, SYR13, VM14a]. Differential [DVS13, DRS15, KJ14, KKS14, KK15a, Nah13, SNA15, SMAM15, SSP13a, TMP11, PVA13]. Differentially [PR15]. Differentiation [DRJ14]. Diffie [ABB20]. Diffuse [ANAK17]. Diffusion [EAS16, JM16, KSRS12, KAP11, KP15b, SK15b, SS13e]. Digest [AM13a]. Digit [Zhu21]. Digital [ADZ11, ARS13, AIRR19, AMB11, AS13c, Ban11, Baz18, BMM12, CM15a, ETKE20, ENE13, GM20, GC18, HCL15, HPM10, HKF12, HB12b, IOO20, IUN10, ITK21, JG11, KK12a, aKNP^{+}15, Kum14b, MA12, MDP12, MAE16b, MG14c, NK19, NRRM12, Nta14, OD13, OFE15, PP13b, PVL11, PS15e, PAH17, PS12b, RB14a, RNM14, RD17b, S14, SS13c, SCW16, SSK11, SCV13, SR15b, SVMLb13, TSY18, TV10, US14, WHP14, YPS19, LGM12a]. Digitally [KN12b, Sta22]. Digits [AJI15, RSA11]. Digraph [RS13c]. Dijkstra [KK20]. Dilated [DSDD12, BS12a]. Dimension [PUW13, PPP16b, ZYM13]. Dimensional [AAEK18, DR15a, MMR14b, MAE16b, NAH11, SSK11, TAY15, YAT15b, ZYM13, MRM13]. Dimensionality [CR12, MN17, ST11]. Dimensions [NOO16, RT11, WRJ23]. Diode [PV17]. Dioxide [AMSS13, SMHS10]. Direct [AE12b, AP16, BFA13, CZW15a, KS12b, KRR15, Mal13c, RM16, RRR11]. Directable [KR15]. Directed [AEA17, KBG11]. Direction [AS15b, EBO13, KAR10, KK17b]. Directional [HR22, MK19, NK12, SVMLb13]. Directionlet [RT13]. Directions [KBT^{+}21]. Directive [ASK15]. Dirichlet [VV14]. Dirty [RMP17]. Disabilities [DDBD22]. Disability [SKGV17]. Disable [JKM^{+}17]. Disabled [SS23a]. Disambiguation [SG15]. Disaster [PWA12, SM15a, SGS18]. Disasters [OE14]. Disc [AP15, FSG17, Hus15, TRR16, WHP14]. Discard [IMIA11, OMH11]. Discharge [BR13]. Discharging [KTS17]. Disclosure [CPD13]. Discounted [CD16]. Discoveries [Kon15]. Discovery [AmAdS12, AM12b, DDL18, EBB17, Kha20, KSC11, LKS14, Mee12, NG10, NM19, OAA14, PG11, SZZR11, SPL15, SSP15b]. Discrete [AAEK18, AAoS14, Alz12, AAO13, AA21, CR15, DDG16, EJC16, HJG^{+}13, JNR19, KA16b, MBKH16, NA20, NK19, PKP13, PGS13, SS13c, SSJ12, TIA15, TG18, GSCT12, SGD13]. Discrete-Time [EJC16, Alz12]. Discriminant [MRR21]. Discrimination [KA13a, MHJ15]. Discriminative [CCJ14]. Discs [SR15c]. Disease [AG10, ASH16b, ARR20, BRR16, BIAE24, EUEI23, GVRP10, Jag15b, KH20, KBTB12, KE14, LABV21, LHKR16, NJ16, OOO19, PK16, RP15, RA10b, SCH21, SASS11, SMT17, SGJ12, VMM14, ZK12]. Diseased [RMD10]. Diseases [Abb12b, AHBAR^{+}11, EFA16, JA22, KM16c, Mes23, MBP16, SN22]. Dish [ST12]. Dishonest [SN15b]. Disjoint [GSA12b, RSS^{+}11]. Disjunct [AmAdS12]. Disk [MM20b]. Dislocation [BAIs^{+}14]. Disorder [JIB14]. Disorders [FY12, GB15, KST16]. Disparities [MAMA16]. Dispatch [AMAHH19, DGH16, DDL16, GV15, LSB17, OC13, Ori14, PS15b, VM12b]. Dispersal [AAMA20, THAP19]. Dispersed [HS10]. Dispersion [CCG12, PR12]. Displacement [AM17b]. Display [AIRR19, CKS13]. Displays [KKK12]. Dissemination [DS15]. Dissimilar [RHS18]. Dissimilarity [HR14a, Ms13]. Dissipation [HD14, Haz14, HGK15]. Dissipations [SDJ13]. Distance [AJ19, ASD16, BS21a, BBHL12, Cai18, CWT15, DS12c, EGES15, GHH15, GPC16, HAH15, JBK13, JVG15, LW16, MS11, NS15b, NS16, PG13b, SS14a, SSA^{+}23, SP12, SNB18, SB13b, SYR13, TSA12, TH14, USR17, YWT12, BIE^{+}12]. Distance-based [BS21a, BBHL12, EGES15]. Distant [ALA12]. Distillation [DK11, HKE^{+}14, PSD13]. Distinctive [PN14]. Distinctiveness [BCC13]. Distinguish [RA10b]. Distorted [NR16]. Distortion [Alg19, AGRR10]. Distributed [AMA14a, AMG^{+}12, AA12f, AIG12, AB11, AT14, BK10a, BA15b, EmErB13, ENEBES15, GH14, GSB12, Hem12, IK11, JIA13, Ker13, KN11, KNV10, Kum10, LJ16, LMZ12, MP10, MT13a, MEE16, NI15, RBF12, RHD12, RGR15, RB16, SSM12, Sal18, SP15c, SK10b, SS13f, TAP15]. Distributing [sMTEkEdAr12]. Distribution [AYAB18, ATA19, AM17b, Ali19, Ble18, EBAS17, EDR15, EBAS15, ETG17, GSB12, HWT^{+}13, HS10, JT11, Job17a, KRMB10, KMG21, KK14, LRRR11, MAaM14, Moh14, MBL15a, Poo22, RA18, TYSS15]. Distributions [EADEA17, RPR12]. Distributive [Mal13c]. District [KDT16, KMG20]. Disturbance [OS16, SR15a]. Disturbances [Alz12, SN16]. Dither [Gum19]. Divergence [KBG11]. Diverse [SZOK24]. Diversified [KD15]. Diversity [Ari23, ES13, RK15, SB12a]. Division [HK10]. Divisor [SSV11]. Divulging [SNP15]. DL [GDD13]. DL-PUSC [GDD13]. DNA [AAP13, AS18b, GJ14, Hos12, KBK10, KEBF13, PS10, SS12f, SA12d, TKC^{+}13, TR14a]. DNN [HHHK21]. DNS [KP13]. do [AkC13]. DOA [TY15a]. Docking [FMS13]. Document [AlS23a, GJ13b, KSRS12, KKBS12, OFE15, SS12a, SPL15, SK13b, SK14b, Sud10, Tes15, SO14c]. Documentations [CS15a]. Documents [AK14, FH13, Gum19, IM12, Mah13, SS12g, WG14]. Dodoma [Ana15, SS14d]. DOE [CS11]. Does [GM15b]. Dogri [MML15]. DOI [SYAS11]. Domain [AK18a, AQ15, ASK15, AAaAD12, BA15a, BG16, CSM21, CS14, EAF^{+}16, GK15c, HO15a, HB12b, JT15, KTA17, KA16b, KS16b, KA14, KJ15, Lah16, Man10, NAK^{+}14, NR12, PG13a, PB14, Roc17, SS10b, SS15a, SCW16, SErEs^{+}12, TERCM15, AM17c, ZAER12]. Domains [EME15, VMA23]. Domestic [Deb20, IA15]. Dominated [PS15b]. Dominating [DM18]. Domination [BS10]. Domino [MM17, SSD^{+}13]. Door [MK12, NKH12, TV10]. Doping [YR12]. Dot [ZC24, DDDD12]. Dotty [HKS12]. Douala [MDCS22]. Double [FML20, Hus15, LC17a, MSA16, RD17b, STD14, TR12, SQ18]. Doubly [AB14b, KMB14, TOMGBSS13]. Doubly-Fed [AB14b]. Down [LSA^{+}17, MMPH15, MAI15, MTKP15, RM16, RGCGCMFJ13, MDP12, DOC^{+}19]. Down* [RA10a]. Downlink [NNR16]. Downstream [Ahm20]. DPCM [SGM^{+}11]. DPVM [ST14]. DQPSK [ASK17]. Dravidian [GP13]. Drawings [NK19]. DRC [LJ16]. Drift [AHM14b, PRR15]. Drilling [Naj12]. Drive [CP15a, Ena12, EEM12, GJG16, HR16, Iss13, PCW10, SV12a, Sal19].

DriveCare [WBL^{+}23]. Driven [AAT16, BWS16, DRBR24, KJJ15, MB15c, NM13, OOJ17, SNB18, UI16, VGJ12, BRR16, MBR16, TB12b]. Driver [ARDS18, DDM^{+}23, Kum12]. Drivers [WBL^{+}23]. Driving [SSO14]. DRLSE [Vya14]. DRM [GRP13]. Drop [EAa14, PSS11, Uzo13]. Dropout [LSSD20]. Dropping [ZTL^{+}14]. Drowsiness [BSS15, DDM^{+}23, VDM22]. DRR [DP17]. DRSA [KAR10]. Drug [ZK12]. Drugs [FMS13]. Drunk [EAB13]. DRX [HITA20]. Dryer [Dar13]. DS [MN15]. DS-CDMA [MN15]. DSD [DS12c]. DSDV [SZUSNK19]. DSM [PV17]. DSmT [EIB^{+}11]. DSP [AJ14, GBM10, SSP15a]. DSP-based [AJ14]. dsPIC [MA15]. DSR [ASNS10, DA10, DGPP15, DB13b, KC15c, PIB11, PSS11, RSRT20, SS10d, SZUSNK19, TKR15]. DSS [KK15b]. DST [HS15c, SYAS11]. DSTBC [NAhA13]. DTA [KM13a]. DTA-d [KM13a]. DTC [OA12, SV12a]. DTCWT [MSS14]. DTN [MG17a]. DTNs [MK17]. Dual [ARMS16, AS11, BFA13, DJA13, MAaM14, OA15, PVR20, PK12b, RC14, SSK^{+}10, SSD^{+}13, SKJ15, UJ13, VR15a]. Dual-Band [PVR20]. Dual-Beam [VR15a]. Duct [SJK21]. due [AJ10, GT11, SK15b, Tal11]. Dufour [SH17]. Duo [OA15]. Duplicate [DRS13]. Duplication [AXKT19, SRA15]. During [MC16, Naj12, EAM14, EE16b, EN13, GT11, PWA12, PNIS16, Raf15a]. DUROID [KNGT10]. Dusty [Haz14]. Duty [ASN12, DKS14, GKG16]. DVCC [Nah13]. Dwesa [MMTS13]. DWT [EAMAES16, AAaAD12, BG16, BHC12, Dha12, HKF12, MSS14, MAE19, RRVP12, SS15a, SAk13b, SNJ16, TAS^{+}16]. DWT-OFDM [AAaAD12]. Dyadic [AH13a]. Dye [KJA15]. DYMO [AJSG11, GCP13]. Dynamic [AOM12, Abu17, AMNA18, AS14c, ASN12, AYM16, Ara15, AAY17, BBM14, Bou13, BA15b, CJK^{+}12, COWK14, CLBA^{+}12, DD14, DSA11, DP17, DA18, EBB17, EMF14, FJM16, FYM11, GK12a, HJE14, Hal13, HN16, Hus12, JV12, KPC16, KT10, KY15, LK15, LGZX17, MSM15, MAM16, MCB12, Nad14, NC14, Neg13, Ngu18, Ori14, PN11, PS15b, PPP16a, RSSV14, RRR11, S14, SRTB11, STD14, SL11, SRM12, TBA12, YRV17]. Dynamic-Double-Threshold [STD14]. Dynamical [Eln14, S23]. Dynamically [KA12b]. Dynamics [HG13, HEN13, Ibr17, NVR12, NLC12, PCD14, RP14].

e-Administration [SAM14a]. E-Assessment [AAO11]. E-Banking [Sab16]. e-CiRM [RMH20]. E-Commerce [Bej16]. E-CRM [RHT19]. E-Governance [LG13b]. E-Government [GM13, AS19, FIA11, ZHUR16]. E-health [AM10]. E-Healthcare [SZM18]. E-Human [Qte14]. e-Initiative [VB10]. E-Learner [HMMN13]. E-Learning [ATAAAF19, AERS11, BK11, OAR16, PA23a, SEb13, Add12, AA14c, ATA14, YMA13, VS12, HMMN13]. E-Mail [KG15b, NUK16]. E-payments [AO16]. E-Procurement [UAMN17]. E-Shape [KNGT10]. E-Shaped [AI13a]. E-Shopping [JKM^{+}17]. E-testing [JHP12]. E-Voting [KS15e, AAMA20]. E/R [ATA19]. EA [WC15]. Each [NVR12]. EAP [HR21]. EAP-PEAP [HR21]. Ear [WES17]. Early [AEEO24, AS14b, ADFAJ15, DSDD12, DPS21, IMIA11, JAG15a, NC14, OMH11, RKS16, RKS17, SG12b, SSC15b, ZEB17]. Earnest [SNP15]. Earth [BAK14, NB11, RACI24]. Earthing [YA18]. Earthquake [KDBT12, MA13b]. Easing [Ban11]. East [KDT16]. Easy [AMT^{+}17, EEE14, EA18]. EC [RMS14a]. ECC [CM12, KVS16, MK13a, PR10b, RS11]. Eccentric [DNP15]. ECDH [ABB20]. ECDLP [SS14b]. ECG [AHM14b, AG13, AS12e, AE12c, DL19, EAJ^{+}15, GC18, IHB^{+}15, JVJ12, KK18, KHAE19, ML12, MD14b, SN14b, WWSS12]. Echelon [FS11]. Echo [Lah16]. ECLAT [DM17]. Eclectic [MR12b]. ECN [SN10a]. Eco [SS16b]. Eco-Friendly [SS16b]. Ecoli [IU18]. Economic [AMAHH19, BJS14, DDL16, GV15, Lah11, LSB17, OC13, Ori14, PS15b, TR12, VM12b]. Economyh [RK13]. Ecosystem [EBAS11, HMMN13, MM15a]. EDF [NS13b]. Edge [AC15, ALH^{+}24, DB12b, FR15, HST19, HKY15, JT15, JND15, Jas13, KT12a, MSN12, NS13a, NSM13, NK12, PP13b, PBG12, PSR13, RS14b, SM15b, TH14, VKB14]. Edge-Preserving [JT15]. Edges [AMK12]. EDU [Far14a]. Education [AA18c, AKN13, AA13d, AER18, AERE21, AIO11, BT16, HEN13, LSSD20, MREd14, MKSB22, urPP13, Ana15, NWS15]. Education- [AA13d]. Educational [AARA19, am12a, AHA14, ASASHF17, EAE14, Elg13, HJT14, KMT^{+}24, MM11b, NMAB14, OAG17]. EEG [ALO17, AkC13, BSS15, BI19, BK13, EAB13, HSH17, KAR12, KB15, MEKS15, TT14, TN14a, VDM22]. EEG-Signal [KB15]. EEHC [SA10]. EEHC-Based [SA10]. EES [Sel22]. Effect [AO13a, ARZ15, AH13a, CM15b, DBKN20, DJ12, EMA11, FKFU12, FM13, Gho15, GDJ12, HS14, JA16b, KR17, KJJ15, KJ12b, KR11b, MML15, MRP16, MB15b, MB15c, NVR12, NKD14, PR12, PRK17, RHM14, RSRT20, Rus15, SJA13, SCD14, SS14a, SR15a, SBR18, SPW15, SK12c, TIT14, VG16, WAmMS14, YJ11, SSMP13]. Effective [AZY18a, AG10, AS13a, ASK17, BM13b, CP15b, Cha15, DD14, DAB12, GF14, GRL20, Has15, HAb14, JI10a, Jay12, KSS14a, KP10b, MD13, MEAT11, MRP16, PGPM23, RS11, SS14c, SD17, SP12, SGJ12, TT11, TAM13]. Effectively [Ota13]. Effectiveness [MNP11, AnAA14, NAH11, ZRE15, VS12]. Effects [Ahm20, AMY23, AJ19, BH17a, BR13, CP14, CNP19, CDC17, DSJ11a, DGJ12, EAMD14, Ena12, HD14, HGK15, HA16, HS15d, MFA13, NR12, PKJ15, Pra14, SS13a, SH17, SDJ13, SM13a, STG12, TH15, VV18, Zaf11]. Effectuate [HN16]. Efficiency [BT14a, CC17, DKS14, ENEBES15, ESAA18, EG15, IA18, Iss13, Joh14, KPC16, LC17b, LAA^{+}17, MS17, MG17b, NS13c, Ota13, PS15a, PB19, PG15b, PMP14, RRB16, SKS10, SB12a, US15]. Efficient [AAP13, APK11, Al 13, AE12a, AS11, ASCR11, ALH^{+}24, AA12g, AAY15, AAR14, AB20, AK11, AE14b, MA22, BD10a, Bat18, BRR16, BS14b, BCC12, BA15b, CT10, CC17, DSA11, DS15, ES15a, EBR12, FSNR18, FHM16, GSN14, Gaa16, GEAE15, GEA16, GRP13, GMM21, GGI14, GB14, GR14, Guo18, HKU19, HNW17, HBS12, IE12, JS12a, JBK13, JS14, JA11, JYP10, Joh14, KK18, KS12a, KG15a, KS15b, KRP12, KP16, KS15c, KC14b, KAAA13, KEV14, Kub17, KG11, KA13a, KPAB14, KSMM15, KJ15, LA15, MMP21, MC13, MK13d, MP10, MC12, MP12a, Mee12, MKWK15, MH11, NS15a, NA20, Naj12, NSS15, NMSA12, NRV12, NTL18, Ode14, OC20, PA13, PA14, PPB15, PG13b, PS15d, PR10b, RMVP14, RKA13, Rah14, RKR11, RK17, SSG13, SV18a, S24, SS16a, SK18, Say22, SGM^{+}11, SPP11]. Efficient [SP15d, SK10b, SD12, SSK13a, SC15b, SS16b, SC19a, SK12b, SSS13c, SC14b, SL11, SRA15, SC16, SR15b, SV12b, SLK^{+}13, TK14, Tha14, TMS15, TMP11, UR16, Upa23, VKB14, VJD22, VSA15, WC15, ZAHH16, ZC24, DB13b, YK13]. Effort [GJP13, SM12b]. Efforts [RHT19]. EGIS [LPP12]. Egypt [AAAA17, AER18, AERE21]. Egyptian [KAFS15]. Eigen [MRTB16, DP12, DP15b, KNI14]. Eigenface [PC15a]. Eigenstructure [ASP17]. EIRDA [SPP11]. Ekeland [LAZ17]. Elaboration [JAB22, MM20b]. Elastic [COWK14, CP14, GM16, HO15b, KC15b]. Elasticity [HSEA15, SO13]. Elastogram [BI13]. ELCC [RMP11]. Elderly [KDC12, SS23a]. Election [AA12f, BBHL12, TB14]. Elections [IND13]. ELECTRE [SGS18]. Electric [AS14a, AN16b, Car18, EB21, KTN12, KSP11, MMN18, Mis14, PR12]. Electrical [BAK14, KTS17, NMK14, NMK15, NO18, RMP17, SGA13, RPMH24]. Electrification [BDN13]. Electrocardiograms [MDD21]. Electrocardiography [KAqC14]. Electroencephalogram [MZAE16]. Electromagnetic [BB12b, DL13, JPP12, KTN13, SDD14]. Electromechanical [BAK14]. Electromyography [ELDM13]. Electron [BB12b, BAIs^{+}14, KDBT12]. Electronic [AER19, ADKsO14, Cai17, De 21, DLR19, Dix11, Nas11, NS23, Okp17, PP16, RHT19, RMH20, Sri12, TOD13, ZEE15]. Electronics [NL15, RP10]. Electrostatic [Cha17, Nad14]. Element [DL13, LNM24, SKS13, SS13a, SAK13a]. Elements [AA16a, AWK19, BK17, EMA11]. Elevation [SVMLb13]. Elevator [CS13]. ElGamal [MSR16]. eliciting [GT11]. Eliminate [JPP12, TIT14]. Eliminating [SS13c]. Elimination [LKS14, MK19, NSMK17, RMVP14, RSM15]. Elitist [OC13]. Elliptic [AC18, Abd17, ABB20, AJ16, CM12, De 21, DBB24, GSR21, KVS16, MNYMGG12, WDAJ15]. Elliptic-curve [De 21]. Elman [Ibr17]. ElT [SPC13]. EM-Field [LNM24]. EMACA [ASNS10]. Embedded [Afr14, AKK11, AHZ15a, BM13a, BM15c, DJJ13, DSS18, HKR17, HTPB11, LJ16, MBA20, MPJ18, Mun14, SS14c]. Embedding [AHZ14, JS12d, LcCm13]. embeddings [NM24]. Embodied [MKM17]. EMD [BIE^{+}12]. Emergency [AzSKM12, BYKO17]. Emerging [LABV21]. EMG [KST16, RRA17]. EMG-Controlled [RRA17]. Emission [AMSS13, GV15, PS15b, SMHS10]. Emopad [BFD14]. Emotion [AAG13, BRD15, BI19, CCJ14, CT14, JV17, PG13c, SB18b, TD21]. Emotional [AAS13, JHB17, Mar12]. Emphasis [DSD15]. Empirical [AJ13, ASNS10, AKS16, CKA18, ELDM13, GSV20, HR14a, Kum10, SN15a, Wis13].

Employee [KM16d, UAMN17, WDN^{+}20]. Employees [SAM14a, YBM16]. Employing [AZY18a, GG11, SSD^{+}13, TH14, VJR11]. EMTN [BD10a]. Emulating [ZMB14]. Emulation [IhZ12]. Emulator [BS12c]. Enabled [TVS19, RS13a]. Enabling [SC12]. Enclosed [AH11]. Enclosure [AHZ15b, HJ13]. Encoder [BWS^{+}12, KWS^{+}11]. Encoding [ASH16a, Al 14, AS12d, HK11, JP15, MSA16, SSA14, SA12d, SKGP12, XL21, NM24]. Encounter [TTA10]. Encrypted [AA12d, SP16, SY17]. Encryption [AS12b, AMA14a, ATA22, AH14, Ali16, AAOA23, AHS15, Baz18, BRD12, BM20b, DPJB12, ElR16, GGM23, GJ14, HS15a, HSP10, Hos12, IPR15, KM15c, aKNP^{+}15, KD22, KB16b, MC17a, MIAEFS15, MSA16, MSA17, MAE19, NBR10, PP13c, RAHH16, RJA23, S24, SG16a, SQ18, SA12d, SKGP12, THAP19, VS10, ZEfNF13]. Encryption-Watermarking [MIAEFS15]. End [IAH14, Moh13a, MH12b, RVGP20, SV14a]. End-to-End [IAH14, Moh13a, RVGP20]. Ended [SMC17]. Endless [OD13]. Endoscopy [MN22]. Enegy [KV16a]. Enema [GMRK12]. Energy [AII13, AH12a, MS16, BKK13, BT14a, BCC12, BDN13, BJPP15, BAIs^{+}14, BJS14, Car18, CC17, DBKN20, DSA11, DF15, DAB12, DGH16, DD20, DB13b, DSS13, DKS14, EhMA13, ETLE13, EB21, EE16b, EG15, FHM16, GHD17, GP14, HB13, HITA20, JJN14, KTN12, KP16, KS14b, KS15c, KMB14, KMMM12a, KMMM12b, KJTM15, KPAB14, KSMM15, KC15b, KVKR16, LA15, LC17b, LCP21, MKM17, MBM14, MFA13, MK18a, Meg10, MJS15, MA17, MCaC14, NS15b, NJB16, OAS19, OC20, PS15a, PSS12, PPM12, PPU11, PPB15, PG15b, RMVP14, RKA13, RP14, Ras16, Ras17, RR18, RK17, RB15, RBDD16, RRB16, DBS15, SS12b, SPP11, SK10b, SB15, SS16b, STD14, SM13d, SL11, SC16, SS17c, TK14, TRM23, Tua17, UR16, WBK22, YK13]. Energy-Alert [BKK13]. Energy-Aware [DF15, SM13d, Meg10]. Energy-based [GHD17]. Energy-Efficiency [EG15]. Energy-Efficient [OC20, SS16b, SC16, YK13]. Energy-Power [KC15b]. Energy-saving [KTN12]. Enery [PS15d]. EnEVoS [Okp17]. Enforcementof [GPB22]. Engagement [AA14c]. Engine [AZY18b, Far14a, FAZ13, KN11, Ran13, RV13, RMS14b, RACI24, SA13b, Ven14a]. Engineering [ANM17, AICN14, BS18, CAS14, DS12c, JS12c, JIA13, REUHTG18, RGK17]. Engineers [BA12c, MJP15]. Engines [Rah13, Ven14a]. English [AS12a, DDD12, MH13a]. Enhance [AA14c, AA18b, AS12d, AOK18, IA18, KV15, MOM17, MBL15a, NVV16]. Enhanced [Abb12a, ABYB22, Abd13, AOO20, AH17a, AHH14, AT17, AM12b, ALH^{+}24, LEO15, BRD12, DSA11, DM10, ETLE13, EEGHS15, GSK13, HEs12, ISH12b, JV12, KP15b, KA12f, KS14c, MDP12, NP12, OAW12, Okp17, RMP11, RSM15, RAJ^{+}12, RC15, RS14b, SOOA13, SK13a, SFH17, SP11a, SGM^{+}11, She11, SF15b, SGS13, SSP15b, TOO16, Ven14b, AAAOG20]. Enhanced-Time [AOO20]. Enhancement [AmH13, AAB22, AJ19, ADS16, CT13, BKK15, BS17, BMM12, CS16a, CS12, EUEI23, EEEE16, GCK17, HS11, Hal13, HHS15, JD13, KM17, Kau15, KA12e, KAP11, KC12, KR10, KRM12, MTJ13, MP12a, MSS14, PN14, PSR13, Pri13, RVKB12, SV14a, SAY16, SMT19, SMVK11, SA12c, SaDN13, SC14a, SRM13, SKS16, TY16, Tua17, US15, LG13a, VT15]. Enhancements [LB11]. Enhancing [AS14a, HAPG17, AS14d, Amb24, AAS24, AP16, AHAT16, Asi15, AT14, Bef18, BM20a, EER14, Eis14, EsTeEh^{+}13, HST19, IE12, KK13c, KSSM24, LS17, MZEE13, NS13b, NA16, OG16, PN11, PA24, RBN^{+}16, RHD12, REA12, RS11, RP24, SP15e, SGKS12, Sub24, Tha13, Ver17, VS12, MG17b]. Enriched [BB14b]. Enrichment [Far14a, KB12a, SCV13]. Enrolled [NR16]. Ensemble [ADFAJ15, ELDM13, FRR11b, Gho11, Hem22, Mes23, NS13c, RP15, RP24, SSS13a, SY15, WOBK18]. Ensembled [SS23b]. Entanglement [SPB12]. Enterprise [ASS17, AMNA18, BB13b, CSKB11, GS14, LSP17, NOO16, NMAB14, RRudS12]. Enterprises [BA12a, OK14, PGPM23, PMP14, RHT19]. Entities [DDD12, Zar24]. Entity [AAA19, Far14b, KB14b]. Entropic [MNB13a, PDR11]. Entropies [SAS13]. Entropy [AC15, JP15, KB14a, KK15c, MM18, VGS11, RA10b, SSS15b, YMA16]. Entropy-based [AC15]. Entry [SP15e]. Envelopment [MAI12, PMP14, VR14]. Environment [AAC16, AM17a, AsS13, AD14b, AH13b, AB20, BS11b, BCC12, BJS13, BOK15, Car18, CDC18, DE13, EzHEgT13, HKU19, HKR17, HDMK14, Hem12, IR10, JFJN20, JND15, JKS^{+}12, JSS^{+}17b, KKR18, KS15b, KBAQ15, KAR15, KY15, Mal13a, MB15a, NM11, PM21, RKB19, RTD10, RBO19, SWP19, SSS15a, SDG18, STD14, SJV13, SN10b, SVJ15, Sri12, SS17c, SKGP12, USK17, Ven14b, VM12b]. Environmental [Nur11, SK12c]. Environments [Ala16, DKP10, EGES15, GEAE18, JHB17, MK12, NKM11, RSHI12, RK21, SA15a, SP15b, SAKS24, MNB^{+}22, SKZ22, TA10]. EOG [GRS22]. ePerícia [RBO19]. Epidemic [ARZ11, LRL17]. Epilepsy [ZEB17]. Epileptic [KSV10, NMN15, SMN^{+}22]. Epistemic [DA12b, Fen18]. EPML [KLS13]. EPMP [AAYM13]. EPON [IhZ12]. EPSF [SAW16]. Epsilon [FSG17]. Equal [GM14a]. Equalization [AJ15, AAaAD12, Kau15, SD18, SA14, TS17]. Equalizer [IS21]. Equation [EG14, FTA18, KZ16b, Nim15, ŞAG17, TA12]. Equations [MNP11, AFW14, DVS13, EEB16, PVA13, RS14a, SNA15, SMAM15, WAG11]. Equi [PS15d]. Equi- [PS15d]. Equilibrium [PP12b]. Equilization [DR14]. Equivalent [Cha17]. Erasure [THAP19]. Erl [KUM13a]. Erlang [ASG10]. Erlang-K [ASG10]. ERP [MBMA11]. Erratic [PB12]. Error [AD18, AKS16, DM14, Far11, Fla18, GG12, IS21, ITK21, Ish14, Jah18, JKS^{+}12, JM16, KG11, Mal10, MM14b, MA13a, NS16, NTL18, SS14a, Sar13, SRK10, SA15d, SPW15, SV14b]. Errors [SPSR12]. ERS [NM14]. ERS-1 [NM14]. ESB [CSKB11, KR11a]. ESO [NSA14]. Essay [EEA19]. Essential [DM22, MJP15]. Establishing [BCBS14, EsGK14]. Establishment [Ana15, JSS17a, PSS12]. Estimate [AJM^{+}15, DDS20]. Estimated [MNB13b, Naj12]. Estimating [AOO20, AY18, BI19, GPR14, Mal13b].

Estimation [AJ15, Abd14b, ASN11, AS15b, Aly13, AB13c, AMJ16, BCO13, BFD15, CWT15, DK13, EDR15, EM23, GJP13, GDD13, HSEA15, KCD13, KG15b, KMMM12b, KH15, KNS20, KK17b, MSEEs13, MRP15, NMSA12, NRR12, OMM14, Ps14b, RUB11, Rag15, RB15, RRR17, SO13, Sar13, SJV13, SSG11, SM12b, SM13c, SC15c, SV18b, SPSR12, TBA12, TH14, TY15a, TK^{+}17, TOHS14, TN14b, Dar12]. Estimator [FA19, HBN12, SND13, SJV13]. Ethernet [IhZ12, YJ14]. Ethical [VV18]. ETL [GS14, NVV16]. ETM [NM14]. ETM-8 [NM14]. ETTH [Ans14]. Euclidean [BHK13, PG13b]. Euler [Rag15]. European [ESH^{+}14]. Evacuation [SGS18]. Evaluate [AKH15, Ask14, GM13, KB12a, YMA13]. Evaluating [AMA14b, AYS19, AZY18b, BKK15, CS15a, hC13, ESAA18, MR13b, NS16, RKR11, RMS14a, SMM11, TCS19]. Evaluation [ATA14, AAA11, AB16b, AlS24g, AASAA15, Aly13, AC15, AS13a, AJC14, AAaAD12, AIK11, BR11b, BADS17, BA11, BGS11, BBMA13, BOS15, CS11, CS10, DA10, DGH16, DS12c, GSA12a, GS16b, GK12a, GRC13, GR14, GK15b, HS11, Hik13, HR14a, HS15c, KA12a, KRMB10, KM15c, KBEA18, KSlAM12, KS12b, KKK12, MN15, MH13a, MC17b, Mar12, MBKH16, MMHP17, MFM12, NP13, Ngu22, OON17, OV13, Oua12, OLT14, PR15, PCR12, PK15a, PFS12, PNIS16, PM13, RSHI12, MP17, RHK15, SS11, SS14a, SRSV14, SH15, SE15, SN14b, SZEM16, SGS13, SCK17, TB13, TR14a, TJ12, TOO16, UCNC13, Zae13, BME19]. Evaluations [Akg13, GS12b]. Evaporator [AHAS^{+}17]. Event [AAoS14, FTLK14, KSS18, KSV10, MBKH16]. Events [ZK12, MN24]. Evidence [AWK19, DE13, RTD10, RMD10, YPS19]. Evidential [MH12a]. EVM [HB14a]. Evolution [DRS15, KJ14, MBM13, MNB13a, PVA13, SS13a, dSEdZ14, SSP13a, SBS13, TMP11, TKR15]. Evolutionary [AmAdS12, AMB11, HS10, Ibr18, KS16e, MEAT11, MNB13b, Ori14, PKB13, RB14b, RGCGCMFJ13, SA13a, SM12a, Sin16, Jah18]. Evolutions [KZ16b]. Evolved [TNKS15]. Evolving [PCKD17, TERCM15]. Exact [AHH20, BHK13, FCS11]. Exam [RDD^{+}12]. Examination [AAYM13, GSV20]. Example [ZHUR16]. Excavators [SVB23]. Excel [DD21, PRK17]. Excel-VBA [DD21]. Exchange [AAS24, BWH15, Maz21, UCNC13, YAA14, KGAM10]. Exchanger [Hus15]. Excitation [MML15]. Excited [Mos11, SS12b]. Exclusion [SS13f]. Executable [KLS13]. Execution [GVW24]. Execution-Based [GVW24]. Exercises [LBA16]. Exertion [NKD14]. Exhaustive [GP13]. Exist [GM15b]. Existence [PVA13, SKA13]. Existing [DL19, DSD15, GK10, GK15b, RS16, SB11b]. eXogenous [IM15]. Exon [TR14a]. Exp [KZ16b]. Expansion [BDC17, CSM21, KZ16b, LA17, SSA14]. Expansions [LSK17]. Expectation [SNB18]. Experience [AWK19, Ana15, HG13, HHS15]. Experiences [Kas17]. Experiment [HNW17, RPV10, SM19]. Experimental [ANAK17, AA13d, AM17b, AAA21, AHH14, AVH10, AAC13, ARR20, DDN14, DC22, EME15, Hus15, IMIA11, MK14, PIB11, RHM15, RKGK15, SJK21].

Experiments [CS11, CHG12, MM15c]. Expert [AZY18a, AER18, CP15b, GM13, HFSM12, IGS17, IA18, JIB14, KM16c, MJP18, NF15, Pil13, SN16]. Experts [RHM14]. Expire [AS24]. Explainable [RVGP20]. Explicit [RMS12]. Exploitation [AMK13, SS14d, WBK22]. Exploiting [DDG16]. Exploration [Has24, KAL15, Naj12, RKR15]. Exploring [AA18b]. Explosion [ASII11, GNN16]. Exponential [EDR15, EBAS15, KB14a, MAaM14, MHJ15, Moh14, SVRB15]. Exponential-Type [MAaM14]. Exponentiated [AR17a, Moh14]. Expression [AH12b, BRD15, BP11, DP15b, JS12a, KBK10, Kum12, MAk12, MMR14c, PZ10, TM13]. Expressions [BSS15, Dur14, GM14a, rPGM14, GM15c, GM15a, GM15b, GM15d, KVN10, PG13c]. Expressionsh [GM14b]. Expressway [NCL11]. Extendable [BK11]. Extended [AS13d, BNMM21, BS12b, Ela11, KZ16b, RUB11, RTM12, SP15a, SHR14]. Extendible [BS12c]. Extending [BM16b]. Extensible [AM11, HR21, SRF12, Bor19]. Extension [EBAS17, EDR15, GB14, MS14a]. Extensions [AA12g]. Extensive [GK15a]. External [Alz12, GM15a, MKSB22, PPB15]. Extra [AB12a, MM20b]. Extract [AD11, ES15a, Elm22]. Extracted [SAW16, TMK^{+}10]. Extracting [CD13, Gum19, SKC15]. Extraction [AAP13, AlS23a, AV18, BLQ21, CAL^{+}14, CT10, Dha12, Ela12, FPA13, GPWS18, GJ13b, Ham14, IEL18, JVF11, JS15, KK18, KS12a, KMS14, KC14b, KM15d, KR14, Kum12, KA12g, MB11a, MM20a, MD14b, MMR14c, NJAK15, NSS17, OOO19, Pil13, PG13b, RMAM19, RH15, SB21, SA12b, SYR17, SSAZ13, SBS13, VD15, AA11b, BG16]. Extractive [GEG23, OO21]. Extrema [NMN15]. Extreme [AHZAH16, EBAS17, JS12b, SY15, SS15c, SD21]. Extremely [ESH^{+}14, Obe23]. Exudates [ASB15, SMR16]. Eye [ARDS18, BAK20, JD15, JND15, Nta14, SP15e, TVY17, GRS22]. Eye-Controlled [TVY17]. Eye-Tracker [SP15e]. Eyes [ABL17]. EZW [AMHAh13].

F [SJV13, LA17, RB14a]. F-QTS [RB14a]. Faber [AA21]. Face [ACB17, AOM12, AEKE15, AB14c, AKSEB15, AR17b, AH11, ASM13, APS18, BCBS14, Bar16, BMU14, BV21, CT10, DSJ11b, DP15b, DT22, DHT13, EAEE13, EAEE14, Far13, HaB13b, HAAT16, HH13, JA11, KPV13, KKSS12, MKAR22, MEE15, MRM13, MMK11, PG15a, PB12]. Face-API [MKAR22]. Face94 [Bar16]. Facebook [AH17b, MMTS13]. Faces [A17]. Facial [AH12b, BSS15, BRD15, DDM^{+}23, Dur14, EAEa12, HO15b, Ham14, KVN10, Kum12, MKAR22, PZ10, PG13c, SO13, TD18]. Facilitate [JKM^{+}17, Roc17]. Facilities [PCS18]. Facility [HDMK14, KK22, MAA13, NRR^{+}10]. Fact [SEA16]. Factor [AAH21, Als18, BHK13, EME13, HITA20, KA12a, PS13, SJV13, SBM^{+}15, Nta14, Liu24]. Factor-Type [SJV13]. Factoring [SS14b]. Factorization [AAR13, MK14, RS13c]. Factors [AZY18a, AS18a, AAEA16, CB16, DR15a, Mes23, MD14a, Qte14, WSW^{+}13]. Factory [RRudS12]. Facts [Pri13, AD13, BR11a, KC12, RSP11, SK15c]. Fading [AH12a, AMN12, BR11b, DM14, EE13b, GG12, MN15, PM13, RMG13, SS10e, SSU10, SB11c, TR15, PS12b]. Failure [ARHERZ10, BKM15, KP10a, MC17b, Poo22, RHS18, RK17, YA18]. Failures [Es12a, MAA13, MHE15, SPSR12]. Faint [ZC24]. Fair [FHM16]. Fairly [KA17]. Fake [JP16, SR23]. Falcon [JJA14]. Falcon-AO [JJA14]. Fall [MMPH15]. Falling [AIRR19]. False [BMM13, FKT12, RMVP14]. Familiarity [KB15]. Family [AHM17]. Fan [MM17]. Fan-in [MM17]. FAP [CC17]. FAP-based [CC17]. FaRes [DT22]. Farey [HSP10]. Farmer [OAA24]. Farms [BML22]. Farrow [HIE14]. Farsi [Far14b]. Farthest [DUP13]. Fast [AAP13, AHBAR^{+}11, BWS^{+}12, Cai18, DSJ11b, Ela11, GR14, KS14a, KPS11, Key12, KH15, KRM12, NK12, PMP12, PGS13, RSS^{+}11, RS13d, SM15b, YYS^{+}23b, ZEF17, ZD15, SGD13]. Fast-Multiplying [Key12]. Faster [AzSKM12, MM23]. Fasting [EUEI23]. Fatal [MH13b]. Fatality [KNS20]. Fatigue [ARDS18, BSS15]. Fault [AEA17, ASS15a, AT14, BM14a, BBC12a, DBB12, FR12, HZK12, HBN12, JM15, KRSB12, KKDC12, KMG21, MK18a, MH14a, NSA14, NEA17, NP12, Rah14, RB16, SK18, SYR17, STE14, SBC12, SVB13, YWT12, GSCT12]. Fault-Tolerant [HZK12]. Faults [BSI12, MT13a, NO18, SBM^{+}15]. Faulty [BBL13, YWT12]. FBMC [TCS14]. FBNN [PG15a]. FCFS [Mal13a]. FCM [FkKB12, KS14a]. FCRM [AB14d]. FD [TCS19]. FDE [Agn15]. FDMA [PG15b, RM12]. FEA [BBL13]. FEA-TFR [BBL13]. Feasibility [CGKP13, MJ14]. Feathering [MRP16]. Feature [AN14, Aa13c, AlS24d, ASM13, AV18, BSAb12, BNS17, CAL^{+}14, CT10, Dha12, ESR13, ESAA18, FPA13, GPWS18, Gho15, HM18, Hik13, IR10, JA12, JK14a, JS15, KA10, KMS14, KSS14a, KR14, Kum12, KA12f, KB16c, Ms13, MN16, MD14b, MG17b, MTM12, NmBV13, OOO19, PG15a, RMAM19, RH15, SCH21, Sha13, SYR17, SBS13, SM15a, Sin16, SP10, TBN^{+}15, NS24]. Feature-Vectors [Ms13]. Features [AZM13, AA14b, AD11, AS24, ADS16, APS18, BP16, CCJ14, DM10, DB10, DSM14, DDG16, Far14a, GJ13a, GP15, Ham14, Has15, IEL18, IN15, KB12b, LP16, MGB11, NC14, Oua12, Par11, RRVP12, RKR17, RKR11, Sak15, SU18, SAW16, TS12, TBKA13, Tes15, TMK^{+}10, VKJ^{+}15, WG14, ZPAQP11, BIE^{+}12, BG13, LT12, NM24, VS10]. FEC [AsS13]. Fed [AH16, AB14b, IAM16, KCD13, KMB14, MA15, MA16, PB16, TOMGBSS13]. Federated [KSlAM12, KR11a]. Federation [TAM14a]. Feed [ES15c, Jab14, RC14, RV16, Sud10]. Feed-forward [Jab14]. Feedback [CWC15, CZW15b, FPA13, GF14, KVRS14, MRG19, RMS12, SSG13, SRSV14, Tua17, VGo12]. Feedback/Assessment [MRG19]. Feedforward [BBZG^{+}17, SSP13a]. Fees [TUS20]. Feistel [RAHH16]. Female [Nad17]. Femtocell [AAMG13, BG17, HKU19, VT15]. Ferrites [PS12a]. Ferromagnetic [MMK16]. Fetal [KK18, KAqC14]. FETs [PKJ15]. Few [KRR11]. FFANN [AG10]. FFmpeg [PB19]. FFT [SGD13, AAaAD12, HS12]. FFT-OFDM [AAaAD12]. FHQA [PG15b]. FI [DNR15, ARP17, AZ16, DV16, EBO13, GP14]. FIAM [KSlAM12]. Fiber [AJ10, AHZ15b, Aq14a, ASK17, CCG12, DPT14, LSB12, MM15c, PSA18]. Fibonacci [RSC16, RTM12, SVB17]. Fibre [Nta14]. Fibrosis [PCC12]. Fidelity [ADVB10, GG11]. Field [AS14a, BH17a, BSP11, ETD13, HS14, JH18, LNM24, MG19, PR12, PSS12, SM14a, TMC17, XRL14, YSH12, ZYW^{+}18]. Field-programmable [ZYW^{+}18]. Fields [PR10b]. Fifteen [SS13a]. Fifth [BFD15, VR12]. Fifth-order [BFD15]. Fight [SR15a, YS11]. File [Al 16, AA12g, AT14, Jay12, NA20, RGR15]. Files [DD15]. Filled [CP14]. Filter [Akg13, Anh14b, Anh14a, ABD22, BK10a, CP12, DH17, HIE14, IS21, ITK21, JPP12, KM17, KA16b, KKDC12, KA12f, KS16e, KK15c, MSS14, PA13, PT15, PPP11, QJY11, RUB11, RM16, RMP17, SS11, SE14, SAY16, SCE14, SIAI14, SGB16, SSR14, SKT14, SR15b, TT11, TVGT18, TN14b, UJ14, GC18]. Filtering [BD18, CCG19, HO15a, Ibr18, KBK12, LC17a, LC17b, MRR21, MVC10, NAZ17, NB11, OEA18, PN14, PPP11, RAJ^{+}12, SF15a, SM15a, SS13e, SK14c, RV12a]. Filters [BSS15, HCL15, HaB13b, KB14b, NL15, SE12a, SV18a, SJ12, SCW16, GBM10]. Finance [IPB14]. Financial [Agg17, DH21, DP11, HMN18, HIS15, IA18, KGAM10, KR11a, Pet22, SO14a]. Find [GJG16, Sab16]. Finding [AMT^{+}17, GKK13, KPM16, MA10, MMPH15, QT15, SSC11]. Findings [SU20]. Fine [HSMb12, SRS14a]. Fine-grained [HSMb12]. FinFET [NRRM12]. Finger [AWA12, AOSM12, KBK12, MBKC15]. Finger-Knuckle [AOSM12]. Fingerprint [AEKE15, AnAA14, AA11b, BCC13, BCO13, Bha17, CAL^{+}14, CWT15, EAJ^{+}15, EAEa12, IN15, JJKD14, KJ15, OAO18, TIA15, UAA16, YA14]. Fingerprints [KC14a, NR16, NBKM13]. Finite [BP11, DL13, Hew11, LNM24, LG11, MM15a, SSV11, SAK13a]. Fins [Job17b]. FIR [Ibr18, PT15]. Fire [BS14a, IAA13, MAARIH16, NSM14, NO18, OE14, Raf15a, Raf15b, SS12e]. Firebase [You18]. Firefly [KKM18, Sha18]. Fires [Alb15]. Firewall [BK10a, HJB^{+}14, KT12b]. Fireworks [BAA24]. First [AJ13, Add12, AZY18a, AAA19, ARZ11, DUP13, KS14b, SSN18, SRSV14, SSR14]. Fish [AI18, SB18c]. Fisher [DP15b, MRR21]. Fisherface [AH12b]. Fisheye [DKB12]. FISODATA [FkKB12]. Fitness [KSS13]. Fitting [AlS24c, NMK14, NMK15]. Five [HKA15, NSM13]. Fix [AER18]. ’Fixed [Ali16, ACK12, Eba13, LAZ17]. Fixed-Data [Eba13]. Fixing [AERE21]. FL [Far14b, GJP13]. FL-COCOMO [GJP13]. Flags [LP16]. Flame [NSM14]. Flare [Tal11]. Flat [CKS13, CS16a, ISH12b]. Flatness [STE14]. Flaw [ISH12b]. FLD [BMU14]. Flexible [AMTM11, FIS16, GN14, GF17, GKG16, Jov15, TG18]. FLFC [Ode14]. Flight [BR13, KNR15, Sha18]. Flip [Jay15]. Flip-Flop [Jay15]. Floating [DDS20]. Flood [AM17a, SRC12, SSS16]. Flop [Jay15]. Florianopolis [DDS20]. Flow [AM17b, AS18c, AMTM11, BMF11, BAF15, BS11a, BB13a, CP14, CDC17, DSJ11a, DGJ12, DJJ13, DB17, ED18, GN14, GSB12, GENERM16, GDJ12, GSA12b, HIU11, HAT16, HS14, HD14, HGK15, wHsYwT12, ILG15, Kub17, KC15b, MJM17, MEE16, MB15c, NMK14, RGCGCMFJ13, SS10a, Sal15, SJK21, SDJ13, SK15b, SK10c, SSRD12, SSY11, TH15, TRM23, TAH13]. Flow-Shop [wHsYwT12, RGCGCMFJ13]. Flowcharts [MGL22]. Flower [PS15b]. Flowers [MKT17]. Flows [BBMA13, KAS12]. Floyd [SM14b]. Fluctuations [RK13]. Fluid [BH17a, BB13a, Bou13, DGJ12, EH14, Haz14, SH17, SK15b, TH15]. Flux [EN13, TH15]. Fly [EEM12, RMS14a]. Flying [Ade13a]. FMI [IGS17]. fMRI [VNM15]. Focus [KS16b]. Focused [BK17, HK10, SS13d]. Focusing [PG11]. Fog [FSER20]. Fog-based [FSER20]. Foil [BAIs^{+}14]. Fold [MK13c]. Folding [GM15c]. Foliage [KNSS11]. Followed [KS16e]. Follower [HZ15]. Following [AZS16]. Font [AHM17]. FOO [SaDN13]. Food [STS19]. Footprint [ND15, BDC17]. FOPID [Abd18]. Foraging [Fer14, KJ12a, KR10, YAT15b]. Force [AWA12, DK18]. Forced [DJJ13, HY13, MKB13]. Forcing [AAaAD12]. Forecast [AMSS13, Sta22]. Forecasting [AJT17, AMAY20, AE17, BNMM21, HJBJ11, HMN18, HEEG15, KRR13, MAP13, PKB13, PRK17, QKRB11, Raf14, SAA24, SVRB15, SB18a, YK16]. Forecasts [LLSL12]. Foreground [SB21]. Foreign [Mal13c, Maz21, PSA18]. Forensic [AR15, DE13, DP16, DRJ14, MM23, RBO19, SR23, SB11a]. Forensics [AN14, ARL16, GS16b, HPR17, WK16]. Forest [Alb15, AR17b, BS14a, KNS20, MM16, SS12e, NS24]. Forex [Maz21, MAP13]. Forgery [DDG16, KSS^{+}14b, SB11a, TSY18]. Form [AaEeK11, JV14, KNT11, Pet23]. Formal [AH13b, HK11, NBR10, RML11, RGNA13, SH15, SA12c, TB15, VV12]. Formalism [HK11, MBM14]. Formalization [AAR11, BKB11]. Format [HKS12]. Formation [Ade13a, BAIs^{+}14, KN17a, MB11a, Naj12, PS15d]. Formatting [TGcO14]. Forming [JTI19]. Forms [FA19]. Formulae [Kha20]. Formulating [TS11]. Formulation [AOK14, CCG12, GFMS14, HN16]. Forums [TGL13]. Forward [AV15, ES15c, GDS16, KK10a, RV16, SU12, Sud10, TY15a, Jab14]. Forwarding [ASY17, GJ15b, SE12b, YYS^{+}23a, ARZ11]. found [GT11]. Foundations [BB13b]. Four [Ara15, DMM14a, HKA15, MBKC15, RK15, SKM16]. Four-Color [DMM14a]. Fourier [SGD13, Gaa16, RAE15, RDE^{+}15, SB13b, TSA12, YT15]. Fourth [HII15]. FP [HF20, MM18, RBT10]. FP-Growth [HF20, MM18]. FP-tree [RBT10]. FPAB [JSK10]. FPAB/BBO [JSK10]. FPG [RRudS12]. FPG-Data-Mining [RRudS12]. FPGA [ZYW^{+}18, BDC17, DV14, EAF^{+}16, ER12, Gha13, HAS17, Jah18, LcCm13, MC13, MA13a, MFM12, NBKM13, NK12, RSA16, RMSS13, SM14a, SFH17, SM15b, SA12e, SK14c, ZAER12].

FPGA-based [SA12e, ZAER12]. Fractal [LK16b, NGA14, SSK13b, SGKS12]. Fractals [DLN12, NLC12]. Fractional [DD13, KS15d, KKI16, KNR15, MBE16, PMM18, RAE15, SO14b, SNA15, SMAM15, ŞAG17]. Fractions [HSP10]. Fragile [EME13, MIAEFS15]. Fragments [Far14a]. Frame [AJN18, HAS17, KKGG12, RM12]. Framework [ASS17, AJ13, AAYM13, AMG13, AA14c, AHZ15a, AH17a, AH17b, ANAH20, AS19, AKA^{+}18, AS24, AIO11, AS13a, AHAPT15, BEE13, BAM13, BNS17, BECM11, BBL15, Bor19, DL21, DP15a, DP16, DBM14, DGH16, EAEa12, EEM12, GJP13, GVW24, GK15b, HKR17, HPR17, HKL16, HMMN13, JYP10, Jay12, JBT23, KZ16a, KS15d, KSlAM12, LOAO23, MR15, Maz21, MMH12, MIS12, MR13b, RT11, RR22, RGNA13, RRP16, SSO14, SZM18, STM13, SAR18, SYR17, SGB14, SU14, TM18, THAP19, UN13, VN10, ZMB14, VS12]. Frameworks [ES15a, TAP15, VB13]. FRANSAC [ZEF17]. Fraud [Dak23, DP11, HIS15, SS15b]. Frauds [Pet22]. Fraudulent [SO14a]. FRBF [MEE15]. Free [AR16, BB12b, BR13, DGJ12, DJJ13, EH16, JIA13, KM16a, KAA20, LRL17, PPP11, SDJ13, SAM14b, TH15, ZZ11]. Free-Parameter [ZZ11]. Freedom [MBKC15, SS12d, TArM14]. Freeway [MKN12]. Freon [RMP17]. Frequency [AS12a, AlS23a, AlS23b, BCO13, BBC12a, BK17, BI19, HO15a, HB12b, Ibr14, JM14, KTN13, KJ15, MM13, MSM15, MDM17, PVR20, PMA13, SK15a, Sak15, SCW16, SA10, SRS14a, TAM19, Wu23b, PC11a]. Frequency-Inverse [AlS23a]. Frequent [DTS16, JS18a, KPM16, KP10a, MM18, Mee12, MA10, SP15d, SV12b, TKC^{+}13]. Friedrich [FTA18]. Friend [PG13b]. Friendly [SS16b, VM12b]. Friendship [Bad13a]. Frog [Ngu24a]. Front [GF17, MH12b]. Frontal [MMK11]. Fronts [AE12b]. Frozen [AKK11]. FRR [KA12b]. Fruits [FDG17, TM18]. FSO [SA15d]. FSR [MS10b]. FT [JM15, SRC12]. FT-AODV [SRC12]. FT-Z-OE [JM15]. FTM [RMR14]. FTP [OCZ13, SG12a]. FTTH [Aq14a, Ans14]. Fuel [ANAH20, HB12a, Ori14, SM10]. fuels [KS13b]. Full [AAA17a, BDC17, KKS14, MGA20, Raf15a, Raf15b, SSK13a]. Full-key [BDC17]. Full-Time [MGA20]. Fully [KBBM15, KEV14, KN11]. FUN [CV13]. Function [AA14a, AJ18, ASN11, Ara15, Ba13, BM13c, BYT14, DHT13, EM23, EHHM17, Fer14, KRMB10, LA17, Mal13b, MK18a, NMK14, NMK15, Ori14, PN12, PCD14, SB13b, SD15, TS15, THT11, TK15, Tua17, UJ13, AAAOG20]. Functional [AYS19, CV13, MKT17, MB13b, PVA13, STG13, SG12b]. Functioned [AAO13]. Functions [De 20, DM18, EEGHS15, GJP13, KJ12b, MC13, MM15b, SS14a, VR15b, KAS12, LAZ17]. Fundamental [BOS15]. Fundamentals [DD21]. Funds [Agg17]. Fundus [ASB15, BS12a, SMR16]. Fungal [SGD^{+}20]. Furnace [KS13b, LSM16]. Fused [SErEs^{+}12]. Fusion [AEKE15, AAA16, ASK15, CA11, DJG12, DM10, El 20, EAEa12, EIB^{+}11, IR10, KP15a, KK16a, KA16b, KS16b, KDC12, KS10, KR16, MR11, PRP15, PGS13, RG21, MS12, SGB14, SL11, SYP10, SR17, YRMB15].

Fusion/Contextual [EIB^{+}11]. Future [BADR16, Kor23, Lic20, Liu24]. Futuristic [JI10a]. Fuzzifiers [Eg13]. Fuzziness [BSS12, NRP12]. Fuzzy [AST11, Abd18, AA12a, AVT13, AAY21, HAPG17, Agg13, AAEEA15, AMAY20, AMA15, Aa13c, AQ14b, Als18, AERS11, AVH10, AS13b, CT13, AK11, ur13, AGH14, Ba13, BK14, BS14a, BBL15, Ber11, BSP11, BWH15, BP15, BIAE24, CT14, CC17, DD13, DB13a, DMM14b, DR15b, DAG16, DPB14, EKES20, EKI16, EJC16, EOO22, EG15, FIA11, FH13, GPWS18, GSD14, GJ15b, GJP13, GPS14, GT13, GM13, HJBJ11, Hab13a, HR16, HCL15, HSK14, HR14b, HAA12, HMAA13, IPB14, IAH14, IAA13, JAG15a, JTA19, JAsA13, KV14, KN12a, KR15, KBN18, KA16b, KB16a, KB14a, Kei14, KN17a, KDC12, KM16c, KKDC12, KH13, KS13a, KBG11, KRR13, KR16, LK15, LT13, LA15, LC17a, MD14a, MHJ15, MH14a, MBE16, MCA14, MJP18, MGB11, NVR12, NK14a, Ngu16, Ngu23, Ngu24a]. Fuzzy [NMN15, ND11, NF15, NM16, OAA16, OE14, OOJ17, OESKF16, Oud13, PAK16, PRP15, PCW10, PS10, PKJ14, PMM18, PA23b, QA14, RB14a, RR11, RÇ16, RV13, RB14b, RMD10, RG21, RBB11, RKS16, RKS17, SV12a, SS16a, SRTB11, SK18, SP11a, SMSM10, SKG14, SKC15, SS12e, SSA16, SP18c, SPW15, SRS12, SZS15, SN16, SSY11, TK14, TB14, TV13, TK^{+}17, TS11, UR12, UN13, Wu21, Wu23a, Wu23b, YAA14, ZRZ11, AHK12, Mus13, Sal14, VJR11, Kei14]. Fuzzy-Dempster [RMD10]. Fuzzy-GA [Kei14]. Fuzzy-MSFLA [Kei14]. FVSI [KRM12].

G [AS13d, Ngu24a, SRM13]. G-Shaped [SRM13]. G1 [Sri12]. G2 [Tha14]. GA [AMN15b, BR11a, CQM15, DRS13, DH12, Kei14, PD14, SJA13, SAKS24, UJ13]. Gabor [BSS15, CP12, HaB13b, KBK12, KS16e, KK15c, Las10]. Gain [ETG17, IHH11, LEE13, ND11, SKS13, SMT19, LGM12a]. Gait [AN16a, DRBR24, Hik13, RA12b, SB12b]. Galerkin [SAK13a, VR15b]. Gallium [CN12]. Game [Bor19, DDBD22, JM14, KMT^{+}24, LAM14, MABE16, RKGK16, SZOK24, SAP16, HLZ^{+}24]. Gamepad [BFD14]. Games [BMRA12, CHG12, DDBD22, EK14a, EG15, NM13]. Gamma [KRMB10, RRR17, SSU10]. Ganesha [NWS15]. Gap [Ahm14a, BBC12a, OSS14, SDD14, SB10]. Gaps [BKK15]. Garbage [AMT^{+}17]. Garment [Ngu24b]. Gas [BKS14, HB12a, PA19, SOME16, SOT18, SGD^{+}20, TSS12, FEbO13]. Gastroenteritis [BIAE24]. Gate [SM14a, SN14a, Tah14, VSS16, ZYW^{+}18]. Gated [SN14a]. Gates [AARA19, HMAA13, ND17, PMU17]. Gateway [GCP13, SGS13]. Gateways [MMP21]. Gathering [RTD10]. Gauss [FTA18]. Gaussian [AU23, GS16a, Ibr14, TD16a, TArM14]. Gb/s [MSS13]. GBM [JA22]. Gbps [CSN15]. GCOA [RKEdK14]. GDI [SSK13a]. GDPS [VRS10]. Gen [ST16, TC10]. Gender [GMS12, Raj19, SB12b, TD18]. Gene [AM19a, eFKS13, JS12a, KBK10, MAk12, MMR14c, SSA11, TM13]. General [Es12a, MFM12, S23, SKC15, SDS15, SNA13]. Generalisation [LM20]. Generalised [BB13a]. Generalization [OKS14, PN11, PPK13].

Generalized [ASR15, DAG16, EBAS17, FA19, KR15, KBG11, MAaM14, Moh14, Mus13, PA14, RSC16, SVB17, SSU10, TR13, Zar24]. Generate [BRS16, Fad18, NRL17, SSC11]. Generating [Dur14, PH23, SCD10, Sin13]. Generation [Afr17, AMAB15, Bas12b, BC12, BK10b, BI13, BJS14, DGJ12, DRS15, DP11, DDC^{+}24, GF17, GSB12, HS10, IK11, JM18, JSS^{+}17b, Job17a, KC14a, KSS18, KV15, KSS14a, KN17b, LCP21, LBA16, MK18a, MEE16, NM13, RSJT14, RA12b, S14, SN15a, SR13a, SCD10, SSP15a, Sm13b, SVRB15, SS15c, VG12, VR15a]. Generation-based [Bas12b]. Generation/Absorbing [DGJ12]. Generative [ZYL^{+}21]. Generator [AB14b, ANAH20, ASG11, BRD12, DD21, JPAK11, KJKM21, KMB14, MBE10, Mos11, SSY11, TOMGBSS13]. Generators [SS12b]. Generic [AE13, BMRA12, DKB12, Gho15, PS14a, SBS13]. Genetic [MNP11, ASP17, Abu17, AAA14b, AAY15, ANA11, AB13c, AMH11, BJ12, BP15, BMRA12, CW23, CM13, Cha14, Cha15, DF15, DRFF15, DAB12, DL13, DA12a, EBAS11, Eja12, ESZED15, GMAR14, GEG23, GG11, HB12a, HEN13, KA16a, KRSB12, KK16a, KB16a, KS16a, Kha20, KSS13, KSL^{+}18, KJ12b, KGK13, KSMM15, LBB11, LNM24, Maz21, MRR21, MHV^{+}14, NKM11, NK14a, NU18, OP18, OON17, OC13, OJ11, PD14, PDK14, Pat14, PB15, Ps14b, RMS12, RKEdK14, Red19, Rok20, RL14, ST16, SCD10, Sar16a, SK16, SAM14a, SM12b, SRS14b, Sin14, SV10, SB12c, TB12c, UJ13, VM14a, WDS13, WG14, ZZW24]. GeneticAlgorithm [SS12b]. Genetically [TNKS15]. Genomes [NBF14]. Genomic [TR14a]. Genre [CS16b, LAM14, SSS15b]. Genres [ETKE20]. Gensets [DDL18]. Geo [KTA17, LOAO23]. Geo-location [LOAO23]. Geo-spatial [KTA17]. Geocasting [Par11]. Geographic [SE12b, SE16, SJ19, YYS^{+}23a]. Geographical [ATA19, AMA14b, El 15, KAR10, YK13]. Geological [NM14]. Geometric [AS13c, BRD15, DS12b, EAEE14, FR12, GQD^{+}22, JND15, MA12, SB13b, TX20]. Geometric-inspired [GQD^{+}22]. Geometrical [AV18, BSZ13, Ham14, Mos14]. Geometries [AM17b]. Geometry [BCBS14, DJG12, NGA14, SM13a, SGKS12]. GeoProcessing [TAP15]. GES [SSA14]. Gesture [GK13, GSK13, JKTM15, KS15a, NS13a]. Gestures [GK15a, PB13]. get [HNW17]. Gezira [ZAA13]. GF [BBA12]. Ghana [ATA19, AP16, ADKsO14, DPB14, IPB14, KAA20, KMG20, OAG17, SNA13, SVB23, TOD13, YA14]. Ghanaian [Add12, AnAA14, Dak23]. GhanaToll [KAA20]. GHz [PS16a, VRS12]. GI [Sri12]. GI/G1 [Sri12]. gigahertz [SDG18]. Gimbal [AVT13, SZ15]. GIS [PS15c, PSG12]. Giving [KKI16]. Glass [AHZ15b, ZAAS18]. Glaucoma [AV18, AP15, WSW^{+}13]. GLCM [AA16b, BG13, RKR17, ZPAQP11]. Global [AMSS13, AOK14, ESR13, Eln14, İHŞ17, KB12a, KTBK12, PB13, PK16, RGR15, RV12b, REUHTG18, UR12]. Glottal [MML15, VKJ^{+}15]. Glow [BR13]. GLSC [PDR11]. GMDH [AkC13]. GME [SDS15]. GMM [KA13a]. Go [Tal11]. Goal [BP15, Hus16, KJKK12, Roc17]. Going [Rah13]. Gold [SKZ22]. Golden [AB12b]. Golf [Nad17]. Gompertz [NMK14, NMK15]. Gondar [YBM16]. Good [PRT^{+}16]. Goods [RA23]. Google [CW23, CDC18, PSG12, RACI24]. Gossip [DGP14, SA15c]. Governance [LG13b, TOD13]. Government [AOK18, AO16, GM13, MMTS13, ZHUR16, AS19, FIA11, am12a]. Governor [TArM14]. GPGPU [TD16b]. GPON [Aq14a]. GPR [CCG19]. GPRS [EBO13, EB13]. GPRS/3G [EBO13, EB13]. GPS [BNS15, EBO13, EB13, GJG16, NKM12, SA15a, SMK17, SK15d]. GPS-Impaired [SA15a]. GPU [AA21, EHSA23, JI10b, KEBF13, SGT^{+}11]. Graceful [Bad13a]. Grade [JD15]. Gradient [AlS24e, AlS24f, AlS24h, AlS24i, KTBK12, KPR14, KS16e, KS14c, MB11a, PR16, SS23b, VKB14]. Gradients [JND15]. Grading [PCC12]. Graduates [urPP13]. Grain [HST19]. grained [HSMb12]. Grammatical [SDB15b]. Grams [MAt12, RKR17]. Granular [BECM11, KV14]. Graph [AA20, AS18d, BM15b, BRS16, BH17b, DM18, DS12b, EAS16, ESR13, ETI14, FM20, HTM14, HMC12b, JI10b, KPM16, KIH15, KAAM11, MB12, NJ16, NEA17, NS15b, OLT14, RAA15, SS21, SSV11, SK13b, SK14b, ZZ11, IM12, ND15]. Graph-based [EAS16]. Graph-theoretic [MB12]. Graphic [sMTEkEdAr12]. Graphical [CKS13, JP16, SAS21]. Graphs [AD15, Bad13a, BS10, DNP15, DS12b, KAP15]. Gravitational [AS17, AE12b]. Gravity [FJM16]. Gray [ADZ11, LKS14, PPP16b, RAA14]. Gray-Level [RAA14]. Grayscale [AAB22, ElR16]. GRDP [IA15]. Greedy [DT14, Meg10, OSA14]. Green [BLL13, GP14, KT15, NJB16, RMS14b, SMS16, SS16b]. Grenade [GNN16]. Grey [ALH^{+}24, HCL15, II16, NMK14, NMK15, YMAS17]. GRGA [HB14a]. Grid [Alh12, AA14d, AS11, AM12b, AH13b, AG14, CJ10, DBB12, DS19, DP17, sMTEkEdAr12, GSA11, HE15, Hem12, JPAK11, JJN14, KRSB12, MMN18, MM11a, Mal13a, OAW12, Pat14, PS16b, RS13b, RHD12, SK15a, SZZR11, SH15, SS10d, SE10, Tha13, TAH13, TY16, Ven14b, WBK22, ZKN10, OAS19]. Grids [AOAG13, KM16b, TR12]. Grimace [Bar16]. Grooves [MZ18]. Gross [IA15]. Ground [HA16, NAhA13, RÇ16, SRM13, TMC17]. Group [Eis13, EE16a, LGZX17, NmBV13, S14, SSA14, SS13f, TE16]. Grouping [MEH14, PT13]. Groups [GGX12]. Growing [MM16, TVGT18]. Growth [AA18b, ADS16, BB12b, HF20, KA16a, MM18, MSA18, PVL10, RRR13, ZSR14]. GSM [MKOS14, MDP12, MNR14, SGA13, UCNC13]. GSOFT [PS14a]. GSPN [KH13]. GTO [Alg19]. Guava [PDG16]. GUI [AYS19]. Guidance [EAE14, TY13a]. Guide [Cai17, PCKD17]. Guided [CJ10, DL13, EAMD14, ESB13, JTS15, MBM13]. Guidelines [ETD13]. Gujarati [PY19]. Gumbel [EM23]. Gummel [SSAZ13]. GVF [AMK12]. GWEs [Zar24]. Gylph [WG14]. Gylph-clustering [WG14].

h [AS13b]. H. [RDD^{+}12]. H.264 [AJN18, AE12a, AJC14, DLBM12, JKS^{+}12, KWS^{+}11, PB19, Sar13]. H.264/AVC [AJN18, AE12a, DLBM12, JKS^{+}12, PB19, Sar13]. H.264/SVC [AJC14, KWS^{+}11]. H.265 [PB19]. H.265/HEVC [PB19].

H264 [AB17, BWS^{+}12]. H264/AVC [BWS^{+}12]. H8 [YMR^{+}12]. HAAR [TAS^{+}16, LSK17, NK19, YT15]. Habits [CAS14]. HACBA [ZCDR13]. Hacking [KA21]. Hadamard [VGS11]. Hadoop [AXKT19, AT14, GD13, GMM21, SBR18]. Hadoop.TS [KK13a]. Half [MLRU12]. Halftone [JM16]. Hall [DSJ11a, GDJ12, SDJ13, CDC17]. Hamiltonian [KIH15]. Hammerstein [BFD15, SK12a]. Hamming [FPS16, HAH15]. Hand [DJG12, GK13, JKTM15, NS13a, PC11b, SG13b, SS16c, SM19]. Handling [AHH13, KHK13, RP15, RAR24, CCE20, NM24]. Handoff [BG17, KA12c]. Handover [FEE13, HKA15, UCNC13, VjS12]. Handovers [THEE11]. Handshakes [KK15b]. Handwriting [AbAAw12, BB12a]. Handwritten [AA17a, AD11, BG13, BHC12, EREM11, JAsA13, KS15a, KK13b, MGL22, RMAM19, RSA11]. HandwrittenText [Far12]. Hanger [Ngu24b]. Haralick [BIE^{+}12]. Harbor [PCS18]. Hard [ASB15, BV14, SB12c]. Hardware [AHZAH16, DLBM12, FR15, HKY15, KFEN21, MHE15, MM12a, NK12, SFH17, SK14c]. Hardware-based [KFEN21]. Harmonic [Alg19, NL15, NSMK17, SGB16]. Harmonics [MK19]. Harmony [AMS13, DMM14a, Gha16, SE15]. Harris [JH18, PR16, SB11a]. Hartley [SGD13]. Harvest [BS14b, VJ24]. Harvested [BJPP15]. Harvesting [MJS15, SB15]. Hash [AA12d, BBA12, De 20, SS13b, TFT11, TS15, AAAOG20]. Hashing [AAP13, BHK13]. Hassle [JIA13]. Hate [WK16]. Haulage [PPR^{+}12]. Hausdorff [PPP16b]. Having [SRSV14, KRR15, RRR11]. Hazard [EM23]. HBV [ZRE15]. HBV-HDV [ZRE15]. HCFA [CS16a]. HCI [BA12c]. HCV [El 13]. HD [AD18, BWS^{+}12]. HDV [ZRE15]. Head [GSD14, Ish12a, KS14b, SHHL13, SSS16, TB14]. Header [EUT17, KAAA13, Moh13a]. Heads [MM14a]. Headset [ALO17]. Health [AB16a, AIP18, BS16, CNP19, DRBR24, EUEI23, KMG20, MOOOOMT18, PUW13, UO23, ZEE15, AM10]. Healthcare [AAT16, Asi15, BM14b, ESEMD21, JH18, KMS19, MBAR12, SZM18, SOOA13]. Healthy [AE13, PNIS16, RA10b]. Hearing [AO13b, Gha13, SC14a, SC19b]. Heart [Abb12b, AU23, ARR20, BRR16, Jag15b, KH20, KM16c, KKM18, KKK13, KR11b, PP16, RA10b, Sak15, SASS11, VDS12]. Heat [Afr17, AHZ15b, BWS16, DGJ12, EN13, FDG17, HIU11, HAT16, Haz14, HY13, HR13, HJ13, Hus15, Job17a, KM16a, SM17b, SJK21, SK15b, TH15, TRM23]. Heating [AN16b, BWS16, MA15, Nim15, WBS16]. Hebbian [FYM11]. Hedge [Agg17]. hedonic [SAS24]. Held [SS16c]. Helicopter [IEG14]. Hellman [ABB20]. Hello [SSS16]. Help [GJG16]. Helpdesk [PS21]. Hemorrhages [Obe23]. Henon [AHS15, ZEfNF13]. Hepatitis [MEAT11, RP15]. hERG [LDB^{+}22]. Hermite [RTM12]. Hertz [SHS23]. Hetergeneous [PP12b]. Heterogeneous [Abu14, AB17, AD14b, BA15a, BBHL12, FEE13, GHD17, GPS14, Hem12, KA12c, KMS15, KC15a, KA13b, LCA12, LABV21, LMZ12, MC12, NV12, RR11, Sal17, SK10b, SV18b, TSRM10]. HetNets [ETG17]. Heuristic [AS12c, AA14d, AK15, CJ10, FAZ13, HS15c, PB13, RS13d, SS14c, STG12, WC15, ZPAQP11, Alh12, GF17, MMP21, SE12a]. Heuristic-based [PB13]. Heuristics [EHA18]. HEVC [PB19]. Hexaband [SSG14]. Hh [LT13]. Hidden [AW20, DI10, GCK17, MP12a, MZ18, SRM12]. Hide [STG12]. Hiding [AMBS16, AMHAh13, Ish14, Lah16, MM12a, MAE16b, MAk12, NA16, SSAK10, SS10c, SD17, SK16, STG12, Sou15, WC15]. Hierarchal [Ish12a]. Hierarchical [ETI14, HSP10, JV17, KA13a, LSS20, LAM14, MK13a, MAMA16, OO21, PPU11, SHHL13, STM13, SS14a, SV12b, ZYW^{+}18, ZCDR13]. Hierarchy [AZY18a, Agg13]. High [AM11, AS15b, ARS13, Ali12, AT15, ADVB10, BBMA13, BIE^{+}12, BK17, BAIs^{+}14, BOS15, CSN15, DV14, GGI14, Guo18, IHH11, IAM16, Joh14, KKT15, KKS14, KDS10, LEE13, LAA^{+}17, Mal10, MA12, MMR14b, MM21, NRRM12, ND17, OMH11, PB14, Raf15a, RAJ^{+}12, RJS^{+}13, RB16, SSD^{+}13, SR15b, SLK^{+}13, TAS^{+}16, Tha14, TAY15, TAH13, TY16, CCE20, XL21, YBA12, YAT15b, ZYM13, Zon17, MH12b, PB23]. High-Accuracy [Zon17]. High-Bandwidth-Delay [OMH11]. High-Degree [XL21]. High-Density [Guo18]. High-performance [NRRM12]. High-speed [Ali12]. Higher [AA18c, AKN13, ASASHF17, AERS11, AIO11, CAS14, DB10, GSP18, HKE^{+}14, HEN13, LSSD20, MKSB22]. Highest [KS14b]. Highly [SS15a, SI13, MMA13]. Highway [SAAM17]. Highways [VSS16]. Hilbert [RG21, SN16]. Hill [FM20, ABA^{+}10]. Himalayan [SKS10]. Hindi [DDD12, GK10, KSRS12, MML15, RK14]. Hip [SSK11]. Hippocampus [TM13]. Hirota [RS14a]. Histogram [DR14, İHŞ17, Kau15, NS13a, PSR13, PDR11, PWG15, RAA14, SH12]. Histogram-based [İHŞ17]. Histograms [AH13a]. Histopathological [SNB18, UNM18a, UNM18b]. Historical [Far12, RHS18, WG14]. Hit [MB13c]. HIV [BI12, SP11b, ZRE16]. HIV/AIDS [BI12, ZRE16]. HMM [AH11, DI10, Isl18, MS14b, MZ18]. Hoc [AMG13, ASD16, AJC14, AS15d, BD10a, BGS11, BD10b, DB13b, GE17, KM16b, KBT^{+}21, NV12, PP13a, RD17a, TAM14b, VMT13, ASNS10, AA12f, AV10, BS12c, ES13, KA12e, Mah19, Meg10, PSS11, SN15c, SSMP13, SF15b, YRV17, SWP19]. HOG [JH18]. Hole [AV10, KBT^{+}21, KJJ15, LKS14, MAM15, NBL15, RMVP14, TY15b]. Holes [NMB13]. Holistic [Has24, MR12b]. Holomorphic [SS15c]. Holonomic [TX20]. Home [Aq14a, APS18, CMKS15, DLR19, GRC13, JKM^{+}17, Kad16, Mah16, PHK^{+}15, RRAFAM19, MN24]. Homeostatic [FYM11]. HOMER [BDN13]. Homes [VM14b]. Homogeneous [HTPB11, KRR15, RSSV14, SRSV14, YRV17]. Homomorphic [De 21]. Homotopy [SNA15]. Honeybee [FK15]. Honeybees [Fer14]. Hop [ARZ11, SK18]. Horizon [AYAI17b, HBN12, YRMB15, VJ24]. Horizontal [ASM13, DJJ13, MA10, Nim15, RHM15, SJK21, SM17b]. Horse [PMD13]. Hose [FK12]. Hospital [KAqC14, KLS13, KSL^{+}18, SSN18]. Hospitality [ADKsO14]. Host [WT13]. Hosted [TA10]. Hosting [PB19]. Hough [JKC11, TBKA13]. Hourly [ANAK17]. Hours [KG15c]. House [MH12a]. Household [Joh14]. Hoyt [GG12]. HPDDRR [JMM21]. HPL [PB23]. HPSO [MG14b]. HRV [RA10b]. HSI [PSA18]. HTTP [KG15b, SG13a]. HTTP-DOS [SG13a]. Hue [RG21]. Huffman [ASH16a, AR10, AS16, AHAT16, Hos12, RS11, XL21]. Huge [ZMB14, SP18a]. Human [AKSEB15, AS18c, ASM13, AC15, ALA12, BRD15, CS15b, DP15b, EAEE15, ESB13, GK13, GRS22, GK15c, HASH17, JH18, KBEA18, KKSS12, MBKC15, MSM17, MS14b, MAMS12, MG19, PK12a, PFS12, PM15, Qte14, SP15e, SPSR12, KBA10]. Human-Agents [MAMS12]. Human-Computer [ESB13, SP15e]. Humanoid [SM19]. Hunting [MH12a]. HVCFPMINETREE [MA10]. HVS [PVL11]. HWPDE [SA12b]. HWSN [KPAB14]. Hybrid [AS17, ADH17, ASP17, Abu17, AA17a, Agg13, AMAY20, AQ15, AMAB15, AMK20, AB14d, ANA11, BR15, BDN13, BAA24, CS16a, CLBA^{+}12, Cur20, DF15, DP11, DB12a, DHM15, DA12b, EBAS11, ESKM20, ESEMD21, EAEE13, EAEE15, EIB^{+}11, EEEE16, EB21, ETG17, EDD19, FSNR18, GD13, GE17, Gho11, HM13, HMC12a, HZK12, HSH17, wHsYwT12, IR10, JV12, JA12, Jag15b, KKR18, KM17, KM15b, KBN18, KA15, KV16a, KMS15, KTBK12, aKNP^{+}15, KKDC12, KPB12, LGC11, LBB11, MAE16a, MG14b, MJM17, MH11, MAP13, MHV^{+}14, NKR14, NTY19, NMS10, NmBV13, Okp17, OON17, PP13b, PP10, PR16, RV12a, RJS^{+}13, RAE15, RSK15b, RG21, SS17a, SSN18, SSP13a, SM10, SGB16, SS17b, SK10b, SAI16, SSSS12, SU14, SLK^{+}13, SS13f, TJ12, UAA16, UC15, UR12, VGo12, WKC11]. Hybrid [WBK22, YRA23, JSK10]. Hybridization [GEG23]. Hybridized [MABE16, SSK13a]. Hybridizing [KB16a]. Hydraulic [EN13]. Hydrocephalus [HSD13]. Hydrogen [AMN15b]. Hydromagnetic [CDC17, TRM23]. Hydrometers [DMB21]. Hydrothermal [JPDB14]. Hyper [KA10]. Hyperbolic [LSB12, PCD14]. Hyperbolic-Secant [LSB12]. Hypercube [AHZ14, AHZ15a]. Hypergraph [AEA17]. Hypergraph-based [AEA17]. Hyperline [BK14]. Hyperspectral [AAEK18, ANM17, LW16, MN17, SGM^{+}11]. Hypertension [WSW^{+}13]. Hypotheses [EADEA17]. Hysteresis [SGB16].

I-FOO [SaDN13]. I-V [LG11]. i.e [DSD15]. i3 [OA15]. i5 [MFM12]. IARMMD [Mah13]. IBA [PSD13]. IBAc [PSD13]. IBBB [AJN18]. ICF [Kam17]. ICI [IEES16]. ICM [EIB^{+}11]. ICT [AZMP17, BT16, NS14, SKS10, Tal11]. ID [NUK16]. ID3 [SBS13]. IDEaaS [CDC18]. Ideal [BM13c]. Ideals [SMSM10]. Identical [CCE20]. Identification [AN16a, AZY18a, ASB15, ASM13, AAO13, ALA12, BM16a, BFD15, BS14b, Bha10, Bou12a, DJG12, DK11, Dar13, DUP13, DSM14, Dha12, DXT23b, EKA13, EAEa12, GJ13a, GJS17, Hab13a, HKE^{+}14, HAAT16, Ibr17, Ibr18, IR10, JHB17, KNSS11, KKGG12, KBM22, KKR^{+}14, LP16, MZ18, MAMA16, NJ16, NU12, Ngu18, OAA16, Obe23, Ojh11, OS14, PS10, PG13b, RKR17, RUB11, RSM15, SS11, SN22, SAS21, SYR17, SSS13b, SBM^{+}15, SG12b, SS10f, SB12b, TSCD15, TIA15, TR14a, TMP11, WES17, WWSS12, YWT12, YB15, YJ11, Roc17]. Identified [OS16, GT11]. Identify [AI13b, ABD22, DDBD22, ETKE20, JSS^{+}17b, RW20, ZRP16, Sal14]. Identifying [DAB^{+}16, FY12, HK10, HASH17, MG14c, NUK16, Pet22]. Identities [RSC16]. Identity [AmM12, GJUL23, KSlAM12, KBAQ15]. Identity-based [AmM12, KBAQ15]. IDEs [ZGAS^{+}11]. IDF [AlS23a, Kam17, KSS18]. Idle [AKK11]. IDM [RMR14]. IDM-IM [RMR14]. IDM-LC [RMR14]. IDS [AsRA^{+}12, Ish12a]. IDSS [AA18c]. IEEE [AM15, DV16, DKS14, EK14a, GK12a, GDD13, HR21, JM14, KG15b, KA21, KDTD11, Kum14a, MJM16, RS13a, RKG10, SSG11]. IEM [LG13a]. IFDMA [MAH15]. iFUM [KPS11]. IGBT [Alg19]. IGP [YAE11]. iHHO [RAR24]. iHHO-SMOTe [RAR24]. II [GJP13, Isu16, SOT18]. IIR [Ibr18, Kum19]. IIRS [ASS15b]. IK [RA12b]. IK-PSO [RA12b]. Ikegami [AMM21]. ILabs [MM15c]. ILC [AYAI17b, AYAI17a]. Ill [KA14]. Ill-Defined [KA14]. Illegal [ASA14, VV18]. Illuminated [LG11]. Illumination [DP15b, PB12, SVMLb13]. ILSB [SD17]. IM [RMR14]. Image [AN14, AAEK18, AAY21, AR10, Afr14, ASN11, AS12b, AMA14a, AB14c, AAB22, Akg13, ATA22, Ak13, AmAr14, AS12d, AA12e, AAA16, AS16, AMS16, AGRR10, Anh14a, AVH10, ASK15, AA17b, CT13, AHS15, BKK15, BS17, BAM13, Baz18, BRD12, BS14b, BHC12, BT14b, CFB11, CJK16, CM15a, DD15, DSDD12, DT22, DB13a, DBB24, Eis14, El 20, ElR16, EAEa12, Ela11, EAF^{+}16, EAO15, FSA12, FTA18, FPA13, Gaa16, GB11, GEB12, GGX12, GGM23, GJ14, HK18, HS11, HAAA15, Has14, HR14a, HKF12, IUN10, ITK21, IPR15, ISH12b, JG11, JA12, JZS13, JT15, JNR19, JSC12, JSK10, JS15, JS18b, KT14, KKBS12, KP15a, KK16a, KA16b, KS16b, KA10, KK12a, KTDG12, KTBK12, KSS14a, KT12a, KAP11, KJK12, KD22, KR14, KS10, KSC11, KPR14, KS16e, KSSM24, KB12b, KKSS12]. Image [KS14c, LSK17, LKO16, LZW15, MC17a, MB11b, MKAR22, MNB13a, MNB13b, MAE16b, MM23, MSD12, MM21, MIAEFS15, MSA17, MESNH17, MAE19, MH11, MGB11, NAZ17, NTY19, NRP12, NAK^{+}14, NSS17, Neb13, MSA18, NMK15, PKP13, PRP15, PR16, PSR13, PP13d, PK19, PVL11, PDG16, PBVD22, PB14, PG13b, PDK12, PWG15, PGS13, QJY11, RVKB12, RT13, RMS12, RFAH14, RV12a, RC15, RM16, RN24, RB11, RKR11, RS14b, RG21, RDE^{+}15, RJA23, RPR12, SSG13, S24, SO13, SEF17, SA13a, SAW16, SPC13, SP11a, SS13c, SS15a, SK16, SCW16, SHD13, Sha18, SC15b, SM15a, SAk13b, SH12, SK12b, SS13e, SA12d, SR15b, SRS12, SP10, Sud10, SMR16, TAS^{+}16, TD18, TSY18, TMK^{+}10, TRSK15, TC16, TM18, US14, UR12, LG13a, VKB14, VGP15, WDN^{+}20, ZEfNF13, ZvAS12, ZAER12]. Image [ZCZ24, ZYL^{+}21, ZD15, AKS^{+}13, IEL18]. ImageFARMER [BAM13]. Imagery [BS17, HBS15, JVG15, PPP22, RACI24]. Images [AD24, AAH21, HAPG17, ASB15, Aa13c, Al 14, AMA14b, ALA12, AS13c, BS12a, MS16, BG16, BP16, BIE^{+}12, CA11, CM12, CD13, DSP^{+}21, DPT14, EAMAES16, Elm22, EEEE16, ESAA18, GJ13b, GGK14, GJM16, GK15b, HCL15, İHŞ17, IKY22, KSRS12, MPR14, KS14a, KG15a, KE16, KS15d, aKNP^{+}15, KH12, KKR^{+}14, KP15b, KSC11, KPV13, KJ15, Las10, LW16, MS10a, MRR21, MSD12, MSS14, MSA16, Moh13b, MBP16, MG14c, NPA^{+}24, Obe23, PN12, PB15, MN22, RB14a, RTK15, RAE15, RMD10, RRB12, Rez12, SS12a, SS10c, SSS15a, SABA12, SNB18, SAk13b, SN10b, SErEs^{+}12, SVMLb13, TRR16, TV13, UNM18a, UNM18b, VMM14, WHP14, ZC24, HSD13]. Images- [GJ13b]. Imaging [ASH15, HSEA15, NMK14, PP10, SSS15a, SSK11]. Imbalance [AA18a, KHK13, MD14a]. Imbalanced [RAR24]. Immigrants [Abu17]. Immigration [LRL17]. Immune [JPDB14, Mes16, ZEB17]. Immunity [BO16]. Immunization [CGKP13]. Immunotherapy [HES14]. Impact [AR17b, AYS19, AnAA14, AZMP17, AZY18b, BS18, BKK13, BT16, HHS15, KS12b, KEA19, LB11, Mah19, Mes23, NS16, NBL15, OAS19, Qte14, RMR14, RSK15a, SS16b, SVRB15, Tes15, YR12, YBM16, ZRE16]. Impacts [IAQ11, MD14a, SAS24, TRM23]. Impaired [SA15a]. impedance [RPMH24]. Impending [Tal11]. Imperceptibility [TAS^{+}16]. Imperfect [MR12a, MAA13, NYU14]. Imperial [DPT14]. Imperialist [AMB11, AMTM11, BS11a, MM14a, MTM12]. Implantable [ETLE13]. Implementation [AOM12, AJ14, AB17, APK11, ARS13, Aq14a, AlS23a, AlS23b, AlS24c, AlS24a, AlS24b, AlS24d, AlS24e, AlS24f, AlS24g, AlS24h, AlS24i, Ali12, AVS13, Ari23, Asa13, AA21, BA23, BAO16, BA16, CS13, CS15c, DFGG15, DB12a, DNR15, DV14, EAF^{+}16, ER12, GM20, GC18, Gha13, GE12, GGI14, GPB22, GM13, HZBG13, HS18, HKY15, HR14b, IGS17, IhZ12, JSS17a, JV17, JVF11, KSKS23, Kam17, KKT15, KKS14, hKhHL13, KS10, KN11, KAAM11, KK14, KD15, MC13, MA15, Mes16, Mis14, MA13a, MAM12, MBL15b, Nad17, NW13, NBKM13, NSMK17, OJR11, OO14, OS14, PZN21, PW23, PR10b, PA23b, RSA16, RJS^{+}13, RKE23, RR18, RMSS13, RA23, SFA14, SDZ15, SE17, SH23, SM14a, SS13b, SFH17, STS19, SR23, SP18b, SC19a, SGS18]. Implementation [SC13, SGT^{+}11, SU20, Sri12, SK14c, TUS20, TOD13, TC10, UI16, USC12, UO23, VD10, Yak18, ZTM11]. Implementations [BDC17, PTPS19, SAS21]. Implemented [Bar16, HPM10, Jah18, SFAS19]. Implementing [GMRK12, KBEA17, NK12]. Implication [UAMN17]. Implications [RK13]. Implicit [RMS12]. Importance [Ash19, YMA13]. Important [CRBJ12]. Imprecise [HDMK14]. Improve [AARA19, Afr17, L13, EEAE17, EMF14, JAG15a, KPC16, MMN18, MEH14, NmBV13, NJB16, RAR24, SGB16, VjS12]. Improved [AAS14, Afr14, AM13a, AAY17, AMH11, BSI12, BP16, CS14, DCH16, DH19, DP12, DP15a, DA18, ES13, EUT17, EAMAES16, EEL11, GSK13, HAS17, IK11, JJA14, KPS11, KKS14, KS14b, KS17, KDS10, KSK14, MWS20, MK13a, NMSA12, Neg13, Ngu24a, OMM14, OC20, OEA18, PDR11, QJY11, RC15, SS17b, SM15a, SSC15a, Tah17, ZTM11, KS16d]. Improvement [AEE15, AICN14, AH13b, AKB16, DB17, DGP14, ETG17, GCP13, GSB12, GS11, HB16, IKY22, JI10b, KBG11, LSS20, LC17a, LC17b, Moh13a, PT15, QA14, SS14b, SA16, SQ18, SE10, Sun19, TB12a, VR14, WKAL24].

Improvements [WKAL24]. Improving [AM15, AAD13, AG14, CC17, Eis13, EEH14, EHSA23, FEE13, HFSM12, IAH14, Job17b, JS12d, KTS10, KNS17, KA13b, MOOOOMT18, MJS15, MY13, Neg13, PPP16a, RGR15, RB15, RM11, SK14a, Sha18, TBKA13, VKB14]. Impulse [MSN12, OSS14, SSR14, SR15b]. Impulsive [KY15]. Impulsively [HD14]. in-Home [Mah16]. In-Memory [AG14]. Incentives [DRBR24]. Incidence [II16]. Incident [Abb12a, TRR16]. Incidents [Raf15a]. Including [SSAZ13, ZAAS18, AKK11, SSY11]. Inclusion [RML11]. Incompressible [BB13a]. Incorporating [ADS16, ATM18, SV14a, SU18, SS17b, SS14e]. Increase [HS18, RHT19, SSAK10, ZAAS18]. Increasing [AAH21, AA16b, ESKM20, NS13c]. Incremental [AD14a, Anh15a, KSS18, KP10b, MN16, MN17, Mah13, OMM14, RBT10, SP15d, SRK10]. Incubator [FZM12, SsTM12]. independent [JHB17, RMAM19]. Indeterministic [PKJ14]. Index [AWA12, ANAK17, CCG12, DNP15, KRM12, MBKC15, PUW13, SEA16, SA12d, Ver17, VDM22]. Indexing [AS12c, FAR16, GB18, GR14, Jay12, MAE16b, SD17, SRM12, TD16a, Yac15]. India [AKN13, GT11, SKS10, SSC15b]. Indian [AS12e, BA12a]. Indicator [ADVB10, IA15]. Indicators [AZMP17, AZY18b, EmdHNmAW13, FEbO13, Maz21, RML11]. Indirect [EJC16, SND13]. Individuals [JKM^{+}17]. Indonesia [Ari23, HN16, KDT16, MA13b, RHT19, SM23]. Indonesian [ASI17, HS24, SM23]. Indoor [Cai18, IHD12, JLCV16, JSS^{+}17b]. Induced [MB15c, SMK14, WBS16]. Induction [AB14b, BFA13, BBC12a, BSP11, DK13, DC22, GAA15, HAA16, HTM14, HBN12, IHH11, KMB14, MA15, Mis14, Mos11, NYU14, RUB11, SV12a, SM17a, SND13, SS12b, SSP15a, SMC17, TOMGBSS13]. Inductor [VRS12]. Industrial [AER18, AERE21, Dar13, Mug24, PSA18]. Industries [MMPH15]. Industry [ADKsO14, LLSL12, MBCL17, PRK17, TB12a]. Inequalities [EADEA17]. Infant [AG10]. Infants [HSD13]. Infarction [MDD21]. Infective [LRL17]. Inference [BWH15, Kam17, OAA16, PAK16, RKS16, SSA16, SP18c, VV14]. Inferred [KA14]. Infinite [BB13a, PVA13]. Inflation [DMM14b, SMK14]. Inflows [Mal13c]. Influence [AAAA17, KM16a, KJA15, KVRS14, MOOOOMT18, MK15b, OS16, PP15, REUHTG18, RM11, SAS24]. Influencing [AAEA16]. Informal [yGL13]. Information [AB14a, ASS17, ATA19, AAA11, AAO11, ASB15, AIO11, ASS15b, AAMA20, ASI17, AGH14, BAF15, ESEMD21, EAEE14, Elm22, EE16b, GFMS14, HBS15, HBA13, KDT16, KK11, KRH13, KBG11, KMG20, Lah11, LHM15, MS10a, MBAR12, MOOOOMT18, MHJ15, NM11, NmEM13, NYU14, ND11, NWS15, PWA12, PK12a, PP12a, PPR17, PRP15, Qte14, RGJ11, RGNA13, RA23, RRP16, SS12a, SSAK10, SSS15a, SK16, SLU14, SAH12, Sou15, SB18a, SJ19, TKC^{+}13, TUS20, TS11, USC12, VK13, VS10, VJD22, Yac15]. Informational [HTM14]. Informative [HHS15]. Infrared [JAB22, KMG21]. Infrastructure [AAAA21, GRL20, KG15b, MK13d, PGPM23, Yak18]. Infrastructures [KTA17, PTPS19, THAP19]. Inheritance [GC12, MR13a]. Inhibitors [NYS15]. Inhomogeneous [KS15d]. Initial [AYAI17a, SK13b]. Initializing [KS14a]. Initiated [KP16]. Initiative [GP14, VB10]. Injection [AK16, DSB12, NKG12, ND17, RMVP14]. Injuries [KAL15]. Inner [EAMD14]. Innovative [APK11, Cai17, CA16, Sha14]. Inpainting [A17]. Input [AAaNJ13, HZ14, IM15, LHKR16, RURS13, RK15]. Input-based [LHKR16]. Inrush [SC15c]. Inscriptions [HPM20]. INSENS [SC16]. Inserting [RBDD16]. Insertion [AHAT16, AA17b]. Insider [Bat18]. Insight [NL15, SG15, TSE17]. Insights [GT11]. Inspect [KN12a]. Inspection [ATH12, ISH12b, MBA20, MHE15, MRH15, SS12d]. Inspired [AII13, Cha14, CQM15, Ghe16, Hew11, Neb13, Pri13, SP15b, SHR14, TT14, GQD^{+}22, Ven14a]. Instability [OMH11, RSK15a]. Instance [AD14a]. Instance-based [AD14a]. Instances [KKR18]. Institutes [am12a, AIO11]. Institution [LSSD20]. Institutional [DSM14]. Institutions [ASASHF17, ED19, KM13b]. Insulation [SK12c]. Insurance [DRBR24]. Integer [AMHAh13, LK16a, VRS10]. Integrated [AB12a, BSS12, BG17, CDC18, DK18, DA18, HN16, JP15, KR11a, KMG20, MKSB22, MKCV13, NRV12, NAH11, PB16, RSJT14, RSM15, RJC15, RMS14b, RR22, SOME16, TM13, TVS19]. Integrating [HFSM12, NAD19, SP16, TB12c]. Integration [AIO11, APS18, BBL15, DVS13, Elg13, Hus12, IAA13, MMP21, NM14, NKM12, OAS19, Pil13, PGS13, SK15d, MN24]. Integrity [DE13, GJUL23, IND13, KPC16, NTY19]. Integro [PVA13, SNA15, SMAM15]. Integro-Differential [SNA15, PVA13]. Intel [Ask14, LB11, MFM12, OA15]. Intelligence [AMSS13, AHH14, BML22, CP15a, DDBD22, DXT23b, Has24, JGZ13, JS15, KSKS23, KAFS15, KM15e, Man10, Mar12, NSM14, OO18, RBF12, SP18a, SM18, SS14e, SBC12, SS12g]. Intelligent [AHAS^{+}17, Abd14a, AYYB15, AERE21, AERS11, ASS15b, AIP18, BM14b, DP16, DR12, DL23, ENC12, EAE19, EEE14, EA18, ENE14, EFA16, GB15, Has16, HHS14, HMMN13, JHP12, KT10, KS15e, MLRU12, MAARIH16, ML12, MHV^{+}14, Nur11, OOE18, OAA24, PKDS17, fSG13, SS17a, Sal18, SMS16, SM18, She11, Sta22, UO23, Ven14a]. Intensity [ASK17, PB15, Rad13, RG21]. Intensive [ETD13, FKT12, MBCL17]. Intent [Ahm14a]. Intent-based [Ahm14a]. Inter [AJN18, BA15a, CSP14, CSN15, MM21, PPP11]. Inter-band [PPP11]. Inter-Color [MM21]. Inter-Domain [BA15a]. Inter-Satellite [CSP14, CSN15]. Interaction [CS15b, GK13, KKK12, MTJ13, OOJ17, SYAS11]. Interactive [AJB13, BMD16, HDMK14, HPR17, Hus16, KK11, KS13b, MM15c, MTKP15, RMS12, Rus15, SAW16, SCV13, SC16, Sub24, SJ19, TGcO14]. Interchange [MML15]. Interconnected [FTBA16, PMK12, Par14a, Par14b, PC11a]. Interconnection [AHZ15a].

Interconnections [RSS^{+}11]. Interconnects [KRMB10]. Interdisciplinary [MRG19]. Interest [AKSEB15, CJK16, FML20, SPP11, SB11a]. Interested [ES15a]. Interests [SSC15a]. Interface [BADR16, DOC^{+}19, ESB13, GRS22, IOO20, JD15, JVF11, KKK12, KW14, MB13b, PJ12, RA12a, SP15e, SPW15, YJ14]. Interfaces [BS11b, GE12]. Interference [Ade13b, Agn15, EF15, EE13b, HKU19, JPP12, KA12a, KAA13, MFE13, SB12a, VG16]. Interior [SGS13]. Interleaving [CSP14]. Intermediate [AWA12]. Internal [Alz12, SC16, SU20, WBS16]. Internet [ASII11, SZM18, AM15, ASII11, AAAA21, AAAA17, ANAH20, JMM21, KP13, MMN18, MNR14, MSA18, OOC18, PAK16, PMBP20, PC15b, PA23b, ZRP16]. Internetworks [AAA14a]. Interoperability [SB10]. Interoperable [MMH12]. Interpolation [BI13, HIE14, Kub17, RTM12, SM14a, SErEs^{+}12]. Interpolative [Abd14b]. Interpretive [STG13]. Interruption [AS13d]. Interruptions [ASG10]. Intersection [DTS16]. Intersections [Rah12]. Interuptor [KTN13]. Interval [Alb15, MRTB16, DTS16, HJBJ11, KB14a, KNT11, LK15, OESKF16, PDK14]. Intervals [Ask14]. Intervention [GAZ13]. Intervertebral [SR15c]. Intra [BWS^{+}12, SM13a]. Intra-organelle [SM13a]. Introducing [Ari23, BAM13]. Introduction [CCG12, KSS13]. Intrusion [AB14d, Bad13b, BAO16, Bef18, BK10a, BA15b, CB16, EUT17, FRR11a, GD13, GAA17, JS12b, JK14b, KB16c, MAE16a, MHV^{+}14, ONJ13, PC15b, RS13d, RGNA13, SF15a, SM18, SS17b, TSRM10, TBN^{+}15, VST17]. Intrusion-Tolerant [BK10a]. Intrusive [ARDS18]. Intuitionistic [DAG16, EKI16, HAA12, KB14a]. Intuitive [SFVA14]. Invariance [MGB11]. Invariant [AHBH11, DP15b, MA12, PB12, S22]. Invasive [AA12b, AA12c, NAD19]. Inventory [DMM14b, FS11, HF20, HDMK14, NU12, RA23, SMK14, TYSS15]. Inverse [ASN11, AlS23a, BS10, DL13, Gha16, KP15b, MAaM14, NLC12, Ngu21, RA18, RRR13, RA12b, SS12d]. Inversion [GSR21, SC19a]. Inverted [CQM15, HA16, TRDB24]. Inverter [AH16, MK18a, MK19, NSMK17, OA12, OAA13, SM10, SMC17, Suj14, TAM19, VSN13]. Invertible [BBA12]. Investigate [AM14]. Investigating [BS12c, Mes23, OAE16]. Investigation [AJT17, AHZ15b, AM17b, ASC14, AKS16, BWS16, EH14, HK18, HY13, Hus15, KAR10, LMZ12, MRJN14, MVC10, Nim15, PP15, PSM15, RRA17, RHM14, RHM15, SC14a, SKG17, SA15c]. Investigations [MBR16]. Investment [DP15a, DA12a, Mal13c, PGPM23]. Involvement [BA12c]. IoMT [GGM23]. IOT [MSA18, WKG^{+}18, AIC19, ANAH20, AYM16, BC22, DSS18, DDL18, GSR21, GJUL23, KBAQ15, Kor23, LWL^{+}23, MDCS22, NAD19, RKE23, SGD^{+}20, You18]. IoT-based [AIC19, DDL18]. IoTs [MMN18]. IP [KAR10, MV12a, RSJT14]. IP-MPLS [RSJT14]. IPFC [AD13]. IPS [HKL16, PPM12]. IPTV [AMBM11]. IPv4 [AIK11]. IPv4/IPv6 [AIK11]. IPV6 [KAAA13, AIK11, GCS14, Moh13a, SGS13]. Iran [TSS12]. Iraq [ANAK17, AAqA14, Ras16, Ras17]. Iraqi [AJT17, TAH13, TY16]. IRBM [TM13]. IRC [AB13b]. Iris [AEKE15, AOSM12, AA16b, ALA12, DP12, Dha12, GK15b, KR17, KA15, Oud13, PPP16b, SS14a, Sar14, ZvAS12]. Irises [OAE16]. Iron [PMH10]. Irradiation [BAIs^{+}14]. Irregularities [AJ10]. Irregularity [AD15]. Irrevocable [KC14a]. Irrigation [KDM17, MCA14, RBDD16, ZAA13]. IRSs [SHS23]. ISA [BM13a, KT10]. Ishikawa [CRN10, NLC12, PCD14]. ISM [CA16]. Isodata [MNB13b]. Isogeometric [FJM16]. Isolated [Bha10, ME14, OMF^{+}14, SSJ12]. Isolation [GJ15b, MTKP15, NSA14]. Isomorphism [AA20, BM15b, IM12]. Isoperimetric [HES14]. Isopistonic [ESH^{+}14]. Isotropic [BSZ13]. ISPWM [SM10]. Issue [AA18a, ES12b]. Issues [Al 13, AJ16, DDD12, GP13, GT11, PS15a, SA12a, WAM12, MNN10]. Italic [YM12]. ITE [AO13b]. Items [DMM14b, HHEY19, STG12]. Itemset [DTS16]. Itemsets [MA10]. Iterated [PN12]. Iterates [CRN10, PCD14]. Iteration [CD16, KKI16, NLC12]. Iterative [Anh14b, KSZK15, OSA14, SD16, SS12c, WKAL24, CKD12]. iTV [OSS14]. IWAK [AH11]. IWAK-means [AH11].

J [NK14b]. J-Wiki [NK14b]. Jaccard [SU20]. Jacobi [NK15]. Jacobian [AFW14]. Jalan [TUS20]. Java [Akg13, JVF11, MA13b]. Jaw [AYE13]. JDRA [PG15b]. Jeddah [ATAAAF19]. Jelinski [MR12a]. Jellyfish [AS15d]. JEP [MA13b]. Job [AJ13, AA14d, AF15, AZ17, CS11, Dak23, GM15d, GF17, GSA12b, LK16a, Pat14, RS13b, RB14b, Sri12, TB12c]. Jobs [AOAG13, GSA12b, RGCGCMFJ13]. Join [GM14a]. Joint [AS18d, RUB11, TG18, VT15, ZBI14]. Jointly [OHI15]. Joke [KJKM21]. Jones [EAEE13, EAEE14]. Jordanian [OAS19]. Joule [HGK15, SDJ13]. JPEG [LBNP12]. JPEG2000 [KR17, MB11b]. JSON [AAS24, BB14b, Pet23]. Julia [CRN10, DLN12, SGKS12]. Jumps [AJC14]. Jungck [PCD14]. Juridical [NK14b]. Jute [MMPH15]. JVM [SHR11]. JWT [AAS24].

Kali [KA21]. Kalman [KKDC12, PN14, RUB11, SAY16, TN14b]. Kanban [TB12a]. Kaput [Tal11]. Karbala [TMC17]. Kasner [AHH20]. KDD [KSC11]. Kekre [AP15]. Kenya [AO16, MOOOOMT18, NKM12, OK14]. Kenyan [OK14]. Kepler [EG14]. Kernel [AZ16, BFD15, EM23, MSS14, SO13, WOBK18, ZD15]. Kernelized [JS12b]. Kernels [NP15]. Kerner [AYM16]. Kernighan [OON17]. Key [AC18, Ali16, AAS24, AV15, Baz18, Bha10, DXT23a, EmdHNmAW13, HSP10, KA13b, MMR14a, PP13c, RMSS13, RML11, SG16a, BDC17]. Key-Aggregate [AC18]. Key-based [Baz18]. Keying [TJ12]. Keys [AT17, KA13b]. Keystroke [BBM14]. Keyword [AlS23a, Ali16, Bha10, KA12g, MM23, Rah13, SB13a, TS10, VJD22]. KFCG [KK10b, SSAK10]. KFDA [VD15]. KFDA-based [VD15]. KFP [Mee12]. KFP-Tree [Mee12]. KGAWSN [DAB12]. Kidney [AAN^{+}19, HS11]. Kinematics [Ngu21, RA12b, SDW21, SS12d]. Kinetics [AEA17]. Kite [DPS21]. Kkn [IR16]. KKT [AJ18]. Klassen [MAMA16]. KMF [TA12].

KMGEM [GNN16]. Knapsack [Ghe16, HBA13, HA17, LGC11, RMSS13, YMAS17]. KNN [AlS24a, EAE14, PG15a, SS17b, VST17, ZPAQP11]. KNN-ACO [VST17]. Knot [MV12b]. Knowledge [AK18a, AH13c, Ba13, BRR16, Hew11, HFSM12, Ibr17, KS14a, KSC11, MBAR12, MBM13, MR13b, NM19, NHMR13, NK14b, Pil13, RS12, SS10b, SSS13b, SA13b, TK^{+}17]. Knowledgebase [Sel22]. Knuckle [AOSM12]. Kolkata [JIF12]. KOMOS [TS10]. KPE [SSAK10]. KSA [Far13]. KSBC [NS13c]. Ku [AI13a]. Kuala [Yak18]. Kumaraswamy [Moh14]. Kumaraswamy-Generalized [Moh14]. Kutta [HII15]. Kuwait [AS18a, AAAA17]. kV [TY16]. KVM [CJK^{+}12]. KWH [LWL^{+}23]. KWIQPLUS [IPB14].

L [KJA15]. L-Alanine [KJA15]. L0 [KS14c]. L2 [CTG10]. Lab [DD21]. Label [AICN14, AS18d, SR15c, EME15]. Labeled [KNA12]. Labeling [Bad13a, MKK13]. Laboratory [Ali12, RHM15, SN10b]. LabVIEW [KHAE19, DBW19a]. Lack [Oua12]. Lacunar [PNIS16]. Lag [Mal13c]. Laid [DSD15]. Laminar [SM17b]. Laminated [EH16]. Laminates [SEE13]. Land [AMJ16, MJ14, PPP22]. Landmark [TX20]. Landmark-based [TX20]. Landscape [LABV21]. Landuse [GJM16]. Language [MS16, Eg13, EEE14, Far14b, FLM11, HPM20, KM15c, KW14, NS13a, RAC12, RHK15, SK10a, She11, TSA12, THT11, You13, EA18]. Languages [BK10b, GP13, Kar11, MML15, Sm13b, SM23]. Lankan [MJP15]. Laplace [AYAB18, SS12f, TA12]. Laplacian [SS12c]. LAR [Mah19]. Large [AM19a, ANM17, BAM13, BS12b, ÇGDT18, ESH^{+}14, FTBA16, Hus12, JTKC14, JI10b, KK13a, KKK12, MMP21, NOO16, PS14a, PM13, RD13, RP14, RPV10, SE17, SP15c, SP15d, SL15, TB13, WAG11, Zhu21]. Large-Scale [AM19a, ÇGDT18, KK13a, TB13, WAG11, BAM13, PS14a]. Laser [BB12b, KJA15, PBVD22]. Lash [PRK17]. last [CTG10]. Latency [Ker13]. Latent [AS12c, BEE13]. Lateral [HSD13, NVR12]. Latitude [KDBT12]. Lattice [DXT23a, KE16]. Lattice-based [DXT23a]. Lattices [AS12b]. LATU [KJA15]. Laundering [HHS14]. Laws [MC17b]. Layer [ARMS16, AZ16, CK13, FM13, HZBG13, IhZ12, KP13, KSSM24, LsAAA13, MJM16, PCC12, RC14, SP11b, ZAAS18, EAMD14, FKT12, HAAT16]. Layered [Bad13b, MS14b, AHK12]. Layering [AAY21]. Layout [ETI14, MA10]. Layout- [MA10]. Lazy [TK15]. LBG [KK10b, SSAK10]. LBP [DDG16]. LC [RMR14]. LCV [HS15b]. LCV-MMAC [HS15b]. LDA [CS16b, HH13]. LDA-based [CS16b]. LDPC [BV14]. Leach [KS14b, KV16a, PPU11]. Lead [DMM14b, Kor23]. Lead-Time [DMM14b]. Leader [AA12f]. Leadership [ZHUR16]. Leaf [CT13, CP12, JA22, MBP16, SN22]. Leak [BKS14, RMP17]. Leakage [ARS13, AC13, DD15, PV17, SAR16b]. Lean [NSS15, PW23]. Leap [NG10]. Leaping [Ngu24a]. Learnable [SBS13]. learned [NM24]. Learner [HMMN13, MTKP15, YMA13]. Learners [AMS14]. Learning [AHZAH16, ARAA11, AEEO24, AS14b, AD14a, AAO22, Agg17, ATA14, ATAAAF19, ANAH20, AAG13, Alt11, AERS11, AYAI17a, AER19, AS18d, Asa13, AB13c, ARR20, Ba13, BNS17, BECM11, BYT14, BS21b, BK11, CHG12, DL21, Dak23, DDS20, DB13a, DC22, DDBD22, CB12, EUEI23, ENE13, Far17, FNH13, GDS21, yGL13, GGX12, GK15c, HEBR11, Hem22, HN16, HAb14, JS12b, JS12c, JBT23, KSA20, KNA12, KDB17, KVK19, KB16c, LR11, LAAI23, MAK13, MB13a, MM20b, MDD21, Mes23, MMR14c, MK15b, MEKS15, Mug24, MRG19, NUK16, Nas11, NPA^{+}24, OAR16, ONK24, ONJ13, PS14a, PRR15, PPP22, PA23a, QCERM15, RRA17, RN24, RP24, RA12a, RACI24, Rus15, SEb13, SN22, SDS20, SP15b, SZZR11, SC15a, SP18b, Sha16, SYR17, SHD13, SAS13, SY15, SS15c, Sin16, SD21, Sub24]. Learning [SKGV17, SJ19, TAM14a, TGcO14, Tit16, Vad20, Ver17, YRA23, YMA13, Zae13, ZZ11, Add12, AA14c, HLZ^{+}24, MTKP15, VJ24, VS12, HMMN13]. Learning-based [Dak23, NUK16]. LeaSel [Ven14b]. Least [AlS24e, AlS24h, AlS24i, AM12b, DK11, EAF^{+}16, GEB12, HS15b, OG16, SD17, SErEs^{+}12, Tah17, Zyo12, MHE13]. least- [MHE13]. Least-Squares [SErEs^{+}12, Zyo12]. LECIM [Al 13]. Lecture [Far14a]. Lecturer [AAO11]. Legacy [RT11, RS16, RGJ11]. Legal [VV18]. Legendre [DVS13, LA17, SNA15]. LEM [SBS13]. Lending [IPB14]. Length [AS18b, Ali16, NSM14]. Lengths [AAGM15, Moh13b]. Lesion [NPA^{+}24, SDB15b]. Lesions [AD24]. Less [SCE14, HIE14, PMP12, SE12a, SE14]. Lesson [NS23]. Lethal [SCH21]. Level [ARP17, ADZ11, AAY21, AEKE15, AAO22, AI13b, AE12c, LEO15, MS16, BCO13, DKP21, DPJB12, EUEI23, EOO22, HAS17, KSRS12, KKBS12, KS15d, KMS12, KKSS12, MFA13, NMK14, NMK15, NSMK17, PPP16b, PBG15, RAA14, RM11, RRP16, MS12, SS15a, SNB18, TYSS15, VSN13, AB12a, BL15, CTG10, DM10, KR11a, MMP21, RS13d]. Levels [AK18a, ASM13, PB15, SE12b, SM15b]. Levenberg [JS12b]. Leveraging [Amb24, PS21, PK12a, Sub24, Upa23]. Levitated [DL13]. Levy [Sha18]. Levy-Flight [Sha18]. Lexical [BK10b, SS13d]. Lexicon [GK15a]. LFC [PMK12, Par14a, Par14b]. Li [OAE16, GP14, PS12a]. Li-Fi [GP14]. Library [Ade11, CBhM12, CS18, JA12, OFE15, RAC12, RR22]. License [Cha12, DDB11, MWS20, MPJ18, SSO14]. Lid [BWS16]. LiDAR [MM16]. Life [AYAB18, Gho15, MD14a, MJS15, SR23, SK12c]. Lifetime [LA15, LS17, NM16, RR12, ZAAS18, SKS16]. Lifting [KK12a, SE15]. Light [JA22, NU18]. Lights [AYM16, KCS14]. Lightweight [HSP10, HKL16, KB16b, SFVA14]. Like [AFW14]. Limit [EF15]. Limitations [FM13]. Limited [Fla18, GJ13a]. Lin [OON17]. Line [ATH12, AH13c, FSG17, HZ15, Kha13, Ngu24b, PMK12, STG13, SGT^{+}11, VM14b, YRMB15, MAK13]. Linear [AA16a, ARZ15, Agg13, AlS24e, Ara15, AVS13, BSS15, BP15, Dar13, DD13, FR14, HIE14, KK15a, KP15b, KKK13, Kum19, LK16a, MGA20, MVC10, OA12, PBG15, PMM18, RAJ^{+}12, RPR12, SV18a, SNA15, SMAM15, Say22, SS13e, Tah17, TArM14, Zaf11, DKB12, KSRS12, TNC10]. linearity [JS12d].

Linearly [MAH15, SRF12]. Lines [SBM^{+}15, SS12g, SVB13]. Linguistic [NHMR13]. Link [HK10, KP10a, MC13, Nta14, RMP11, SA15d, USK17, OAA13, TG18]. Linked [Far17]. Linking [KB15, PP17a]. Links [JM15, Rok20]. Linpack [PB23]. Linux [AZ16, BAO16, DSS18, DDN14, HKR17, KA21, LcCm13, LB11, MPJ18, Mun14, NRR^{+}10]. Lion [AA18b, Raj19]. Lion/Lioness [Raj19]. Lioness [Raj19]. Lip [MZ18]. List [KS17]. Lists [Abd13]. Literacy [Add12]. Literature [Bey14, JG11, RS16]. Live [Alt13, AAR11, KS15d, MB15a, SA15a]. Liver [OAA16]. LJF [AJ13]. LLC [SK14a]. LLWNN [MAP13]. LMMSE [SC19a]. LMS [CS18, Ibr18]. Load [AJT17, AMG13, AMAHH19, AY18, ABBG14, aAO14, BNMM21, BHK13, CJK^{+}12, DBW19b, sMTEkEdAr12, EMF14, ETG17, FSER20, GK12a, GSB12, JIA13, KA12a, Kon15, KRR13, KG15c, MMN18, OC13, Ori14, PC11a, PPP16a, QKRB11, Raf15a, SP13, SK15a, SH15, SVJ15, SSY11, Uzo13, YMA16]. Load-frequency [PC11a]. Loaded [SSG14]. Loads [IFT14, KC12]. Local [AHBH11, Anh14b, Anh14a, AZY18b, COWK14, ESR13, FPA13, GPWS18, GL16, GK12a, JT15, KSBM17, KM17, KTBK12, KB14b, KPV13, KK15c, LKO16, MRP16, NMN15, PRP15, PK16, SBM^{+}15, SRS12]. Locality [BHK13]. Localization [Cha12, DDB11, DSJ11b, EB13, JND15, MIAEFS15, MN22, RKB19, SV14b, TL16, JD15]. Localized [Nim15]. Locating [NHH10, RSP11]. Location [APU16, Ade13b, AY18, BKS14, DSS13, GJG16, HS15d, KRP12, KDC12, KRM12, MBL15a, SMVK11, SSA^{+}23, SGS18, SM13c, TKC^{+}13, YWT12, AGRR10, LOAO23]. Locations [HMO17, SAAM17, Ras16]. Loci [EEL11]. Lock [NKH12, TV10]. Locked [ElZA13, PS16b, PS12b, TAM13]. Log [AA12g, KV15, LZW15, PK15b, ZRP16]. Log-transform [LZW15]. Logic [AARA19, AAY21, HAPG17, Als18, AK11, BS14a, BSP11, BKB11, BIAE24, CHW16, CC17, DPB14, EBR12, EOO22, EG15, GPWS18, GSD14, GPS14, GT13, IPB14, IAA13, JAG15a, JTA19, JAsA13, KBN18, KKS14, KDC12, LK15, LA15, MGB11, NVR12, NRRM12, NK14a, Ngu16, Ngu24a, NMN15, NM16, OE14, OOJ17, Olu14, PS10, PA23b, RÇ16, Rah14, RJS^{+}13, RBB11, RKS17, SV12a, SSK^{+}10, SK18, SSD^{+}13, SSK13a, SKG14, SN14a, SSS13c, SLK^{+}13, Tah14, TK14, ND17, VJR11]. Logical [MK13b, SE12b]. Logistic [AlS24f, Baz18, CK12, CRA12, TB12a, NS24]. Logistics [KUM13a, MABE16]. Lognormal [SSMP13]. Logo [DSM14, GP15]. Logs [MM14b, Mee12, SL15]. LOKS [AV15]. LOM [EK14a, EK14a]. Lomax [EBAS15]. Long [CSN15, NK15, SMN^{+}22, SAA24, VRS10, FMS13]. Longer [AJ13]. Longitudinal [SR15a]. Look [NG10]. Looking [RKR17]. Loop [AVT13, ElZA13, PS16b, PS12b, SFA14, Sal15]. Loops [BAIs^{+}14, ElZA13, TAM13]. Lorenz [HS15a]. LOSEPs [AZY18b]. Loss [BR11a, CS10, GSB12, mH14b, IK11, KK14, LPP12, LRRR11, SM13d, RHU12]. Losses [KR10, KC14c]. Lossless [AMHAh13, AmAr14, CM12, KDS10, MM21, SGM^{+}11]. Lost [DMM14b]. Lot [BMF11]. Lotka [LEE13]. Low [AKK11, ARS13, AB12a, ACK14, AV15, AC13, AG13, AMS14, BDC17, CKT10, ETLE13, HIE14, Ibr14, JD15, Jay15, JM15, KDBT12, KFEN21, KHAE19, KKT15, KVRS14, KN12b, KAAA13, LGM12a, MM17, MKWK15, MSS13, Nur11, PBK14, Rab14, RHU12, SMT19, Sar13, SM13d, TY13a, TMP11, TOHS14, VRS12, YK13, MH12b]. Low-Cost [Rab14, TY13a]. Low-footprint [BDC17]. Low-loss [RHU12]. Low-noise [AB12a]. Low-Overhead [AV15]. Low-Power [SMT19]. Low-Resources [KFEN21]. Low-Voltage [SMT19]. Loyalty [HS18]. LP3 [VR11]. LPC [Isl18, RSHI12]. LPDA [PS16a]. LQR [ASP17]. LR [CS10]. LR-WPAN [CS10]. LSB [ARMS16, AS16, AHAT16, AA17b, EAF^{+}16, EAO15, MAk12, OG16, Rad13, SS10c, TC16]. LSH [KSS18]. LSO [SYAS11]. LSTM [GK21, SMN^{+}22]. Ltd. [AP16]. LTE [AAMG13, AGZ15, ETG17, HKU19, HZBG13, JA16a, JA16b, Kum14a, MAM19, NNR16, RM12]. LTE-Advanced [AGZ15]. LTE-based [JA16b]. LTSA [ZZ11]. Luczak [AE13]. Luczak-based [AE13]. Luggage [LPP12]. Lumber [SR15c]. Lumpur [Yak18]. Lung [AD24, AEEO24, PK16, RVGP20, SSC15b, TVGT18]. LUT [SFH17]. LUV [AP15]. LuYAP [SYAS11]. LVCSR [GP13]. LWM [CM12]. LWP [GPWS18]. Lyapunov [EHHM17, Nad14]. Lysine [AA18a]. LZ [LAZ17]. LZ-functions [LAZ17].

M [ASG10, KMMM12a, KMMM12b, AS13b, PBK14, AKK11, ASG10, AS13d, KSBM17, SB11c]. M-QAM [SB11c]. M-Ticketing [KSBM17]. M/M/1 [KMMM12a, KMMM12b, ASG10]. M2M [JA16b]. MAC [BADS17, DKS14, FM13, GSN14, GB14, HS15b, IhZ12, JV12, RKG10, ZBI14]. MACA [ASNS10]. Machine [AHZAH16, ARAA11, AS14b, AA17a, Agg17, AA12b, AA18a, ANAH20, AERE21, Als18, AAG13, ARR20, BAK14, BNS17, BBC12a, BYT14, BK13, BS21b, Dak23, DK13, DDS20, DMB21, DR12, DB17, DDD12, DC22, EUEI23, EEA19, EE16a, GDS21, Gho11, GRS22, GK10, GK15c, Has24, Hem22, JA16a, JS12b, JBT23, KVK19, KJTM15, KC15b, KB16c, LR11, LAAI23, MH13a, MJ15, MK13d, MM20b, MDD21, Mes23, MJM17, MEKS15, Mug24, NUK16, OOO19, ONJ13, PFS12, PC11b, PNIS16, PG13c, RRA17, SP18a, SN22, Sha16, SAS13, SY15, SS15c, SD21, SM23, TGcO14, CCE20, Vad20, Ver17, YRA23, Zyo12, AA12c, HLZ^{+}24, VJ24]. Machine-Based [BBC12a]. Machine-Learning [KB16c]. Machine-to-Machine [JA16a]. Machinery [US15]. Machines [CJK^{+}12, FKT12, GN14, IEME12, KAS13, MB15a, OA12, RGCGCMFJ13, SM14a, SYR17, CCE20]. Macroblock [FIS16]. Macrocell [BG17]. Macrocell-Femtocell [BG17]. MAEDA [KP16]. Magnet [CPSP13, SHSA12]. Magnetic [ASH15, BH17a, HS14, KS14a, NJ16]. Magnetohydrodynamic [HAT16]. Magnitude [DB10, Isu16]. MAHEFT [GSA11]. MAHEFT-based [GSA11]. Mahogany [RACI24]. Mail [KG15b, MSR16, RKGK15, NUK16]. Main [DSD15]. Maintain [EMF14, HBS15, KPC16]. Maintainability [SZEM16]. Maintenance [AGL17, ASG11, BKM15, JSS17a, KMG21, NI14, PSS12, SE11]. MAIZ [PWA12]. Maize [SCH21]. Major [AG10, IAQ11]. Make [BB10, AAT16].

Makeover [SATS13]. Makespan [SV18b]. Making [AA18c, AKG20, CJC19, IA18, IA15, KB14a, Mos14, RR11, SH15, HAA12, KKS14, Sal14]. Malaria [AAA20, DPB14, OO18]. Malaysia [Yak18]. Males [NKD14]. Malfunctions [AERE21, HKE^{+}14]. Malicious [HK15, SJV^{+}15, TGL13, TOO16, UC15]. Malignant [Sha16]. Malware [HM18, RH15, SS21]. Mammogram [ADZ11, KP15b]. Mammograms [JZS13, Zyo12]. Mammographic [BMM13]. Mammography [BMM12]. Manage [AER19, NK14b]. Managed [Kar11, Kum14b]. Management [ADH17, Abd13, ADT17, AGFA16, AG10, AAA21, ANAH20, AMK20, CS18, CGKP13, DGH16, DA12a, EEAE17, GJUL23, HJB^{+}14, HS18, IAQ11, JMM21, KB16a, KSlAM12, KW16, KAA20, KA13b, Kum10, KMG20, MBM14, MMN18, MBAR12, MG14a, MC12, MJ14, MR13b, MCA14, NHH10, NKM11, NI14, NCL11, OE14, PPM12, PS15d, QCERM15, RSP11, REUHTG18, RBM19, RBO19, RHT19, RMH20, SM15a, SC14b, TUS20, USK17, USC12, YRV17, YPS19, ZHUR16, ZAA13]. Manager [ARMS16]. Managing [AIG12, TS11]. Mandelbrot [SGKS12]. MANET [AASAA15, AD14b, AE14b, BA15a, BKK13, BSG15, CB16, DF15, DSS13, GJ15b, HBS12, Ish12a, KJJ15, KC15c, LKS14, MKN12, MS10b, PPM12, RKA13, RSM15, RSRT20, SG12a, SSP15b, SJ14, TY15b, TKR15, VR11, ZTL^{+}14, JC14, KRP12, RC12]. MANETS [MS11, AMY23, JV12, JSS17a, Kum10, MAM15, SFVA14, WAmMS14]. Maneuver [AMN15a]. Manhattan [FD14]. Manifold [VNM15]. Manipulation [BA12b, WBL^{+}23]. Manipulator [MBKC15, MBE16, Ngu18, Ngu21, Ngu23]. Manipulators [Gha16, Ngu24a]. Manipuri [KK13b]. Manual [Ahm14a, SE15]. Manufacturing [ASH13, GYP^{+}15, Jov15, LSP17, Ngu22, VR14]. Manuscript [ASI17]. Manuscripts [Far12]. MAODV [SS16d]. MAP [PSG12, AS12b, AQ15, AA12d, AHS15, CRA12, Dum20, FY12, GM14a, GM15a, IHD12, JSS^{+}17b, KPC16, MP12a, OO14, RTK15, RR16, SDS20, SP15d, AKS^{+}13, SV14b]. MAP-Model [AKS^{+}13]. Map-Reduce [RR16]. Mapper [GM15c]. Mapping [AN14, AB14c, ATA19, AAqA14, DMM14a, DBB12, GM15d, GH14, KH13, Kum19, LAZ17, MBM13, RPV10, RS14b, SMSC14, SMK17, SS15c]. Mapreduce [MM12b, GM14b, GM14a, rPGM14, GM15b, GM15d, SBR18, TR14b]. Maps [AH14, CW23, KTBK12, MAE16b, MD14a, PKJ14, RAE15]. Marathi [DDD12]. Margin [KA11]. Mark [TVGT18]. Marked [ArAn13]. Marker [SNB18]. Market [AS18a, ASG11, DH17, FEbO13, GK21, Jab14, JM18, KGAM10, Lah11, Lic20, NMS10, OE14, RK13, Raf14, RSK15b, SAA24]. Marketing [AP16]. Marketplace [PSG12]. Markets [Agg17, NHH10]. Marking [JJKD14, PBG12]. Markov [A18, AW20, AAMG13, CD16, DI10, GCK17, KHSE12, KK22, MSEEs13, MP12a, MZ18]. Markovian [KS18, SDB15a]. Marquardt [JS12b]. Mars [PMD13]. Mashup [PWA12]. Mask [KTBK12]. Masked [Oud13]. Masking [SV14a]. Mass [BMM13, FDG17, MSD12, SK15b, TH15, YAA14]. MaSSEETL [GS14].

Masses [ADZ11, JZS13]. Master [L13, CM13, Ngu24b]. Master-Slave [CM13, Ngu24b]. Mastitis [DSP^{+}21]. Match [Kha11]. Matched [SIAI14]. Matched-Filter-based [SIAI14]. Matchers [TERCM15]. Matching [AS18b, AE12a, AA12e, AOK14, BCC13, BC12, CJK16, CJK17, DSJ11b, EBB17, HMO17, KH15, MRP15, NW13, ND15, Ode14, Oua12, PRT^{+}16, Ps14b, RB14a, Rad13, WKAL24]. Mate [BA23, MP12b]. Material [AHZ15b, DS12c, HAAA15, PMH10, PVL10]. Materials [CM13, OAR16]. Maternal [NS14]. Mathematical [ABAO24, Apo19, BI13, CWT15, DSDD12, GFMS14, GSA12b, HSK14, KKI16, LBA16, MM15b, PR10a, PMM18, SK12a]. Mathematics [AARA19, RJS^{+}13, Sha14]. MATLAB [AM17c, AS15b, Ans14, GPR14, KSP11, Kum14a, AB13a, CBhM12, IMM12, JNR19]. MATLAB/Simulink [KSP11, AB13a, CBhM12]. Matrices [BBA12, DN12, JSSA20, OMM14, BM16b]. Matrix [AlS24d, AlS24g, AAR13, Bou13, DVS13, Far11, IMM12, LBNP12, MFM12, MK14, PPP16b, PMU17, SSV11, SC19a, SAk13b, SN10b, TFT11, TY15a, WKC11, ZC24]. Mattress [Ngu22]. Maturity [RRP16]. Max [ARZ11, KV14, SMS16, SSS15b, SKC15, DSS13]. Max-Min [SMS16]. Max30102 [AU23]. Maximal [TKC^{+}13]. Maximization [JM18, KK14, MK15b, NHH10, SKS13, SNB18]. Maximizing [AOB12, ENEBES15, ES15b, LA15]. Maximum [AMS13, AMTM11, KM12, LW16, MHE15, MAK13, MG12, MCaC14, SC13, SC15c]. MAYA [SRK10]. MB [MFE13]. MB-OFDM [MFE13]. MC [BBD13, Agn15]. MC-CDMA [Agn15]. MCDM [AKG20]. MCS [SV14b]. MCS-Map [SV14b]. MCUs [AKK11]. MD5 [SS13b]. MDA [DS12a, SS14e]. MDBUTMF [ITK21]. MDC [ES13]. MDFT [SE14, SCE14]. Mead [CCG12]. Mean [Anh14b, DS12c, EME13, HITA20, MJM17, SSR14, SB13b]. Means [AlS24b, BS21a, BT14b, DUP13, GNN16, KM17, NDC10, OAR16, PK15b, SS14a, SP11a, SP12, SK13b, ZRP16, AS15a, AH11, AA11a, CCG19, DH17, DAB12, Hab13a, MJ15, PRP15, RPV10, SYR13, WG14]. Measure [AJ18, AYS19, CS12, GC12, KB12a, Ms13, PDR11, RT11, RB11, SP12, SU14, YJ11]. Measurement [AWA12, BSM17, NMK14, NMK15, RRP16, SBHM13, TTA10]. Measurements [Ae14a, AA16b, ESH^{+}14, MDNA12, LG13a]. Measures [AAA11, DKP21, DRJ14, HR14a, JFJN20, JBP16, KBG11, MHJ15, RHD12, MS12, SPSR12, Tal11, YB12]. Measuring [AYE13, SM19]. MEC [PPM12]. Mechanical [KJA15, OS16, SEE13]. Mechanism [AsRA^{+}12, AMK20, AB14d, BNMM21, DL21, DKS14, EHI17, FYM11, HEs12, HSP10, KN12c, LPP12, NMK14, NMK15, RGR15, RAB12, S13, SPP13, SD18, SS10b, SN15c, SK18, VSS16, VjS12, YYS^{+}23b]. Mechanisms [KK11, PG11, SU12]. Mechanismsfor [FIA11]. Media [AAAA17, Ari23, Far14a, HY13, HR13, HJ13, ILG15, JTS15, OD13, PH23, Rok20, SAA24, WK16, SAS24]. Median [SKT14, SR15b, SK14c, TVGT18]. Mediated [Muk13]. Medical [Add12, AAI13, ANM17, ASK15, BS21b, CPG15, El 20, FLM11, GGM23, JSSA20, KMS19, KN17b, KKR^{+}14, Mah13, Man10, MH11, PN12, PSM15, RV12a, RKR11, S24, SSS15a, SASS11, US14, UR12, Wis13]. Medicinal [AAO13]. Medium [BA12a, BB13a, CP14, CA16, DJJ13, GS14, HIU11, sNIALY13, KIA15, OK14, PGPM23, PMP14, RHT19, SOO15, TH15, WBS16]. Medium- [PGPM23]. Medium-sized [PMP14]. MEECDA [KPAB14]. Meeting [AIC19]. Mel [AS12a, Isl18, MSM15, RSHI12]. Mel-Frequency [MSM15]. Mel-LPC [Isl18, RSHI12]. Mellitus [MJP18]. Membership [Ba13, GJP13]. Membrane [SM13a]. Memory [AS17, AAA21, AG14, CLBA^{+}12, DN12, Far11, HSMb12, Mes16, NmEM13, Pou13, RPV10, RB16, RKY12, SMN^{+}22, SAA24]. Memristor [Apo19, ZZW24]. MEMS [BSM17, CM15b, Cha17]. Mental [CNP19, FY12]. Menus [MGA20]. Merging [GM15d, KAFS15]. MES [BM20b]. Mesh [AJB13, CK13, BA11, EsGK14, GMAR14, HS15d, LNM24, MNHB12, TBKG13]. Mesh-based [BA11]. Message [AR17a, AS12d, MM12a, MC17a, SS13f]. Message-Peer-to-Peer [MC17a]. Messages [TCS19]. Meta [AK15, EHA18, FAZ13, GM15c, GM15d, GF17, MMP21, SE12a]. Meta-Heuristic [AK15, FAZ13, GF17, MMP21, SE12a]. Meta-Heuristics [EHA18]. Metabolites [NJ16]. Metadata [MM11b, QJ13, SB13a]. Metaheuristic [BAS12a, GN14, MPR14, LM20, SE14, SB18c]. Metaheuristics [BG20, Ghe16]. Metal [MMK16]. Metastatic [PS10]. Meteorological [Alb15]. Meter [LWL^{+}23]. Metering [MDCS22, SDG18]. Method [ADZ11, AHZAH16, AH12a, AZY18a, AM19b, AFW14, AA12d, AS12d, AHH14, AAY15, AMH11, L13, BNMM21, BA23, BD10a, Bar14, Bas12b, BLA12, BL15, BS11a, BB14b, BS14b, CT10, CCG12, CC17, CPG15, DPS21, DSJ11b, DD13, DL13, DKB12, DB12a, DA12b, EAE14, EAEE13, EEAE17, EAO15, eFKS13, GEAE15, GS11, Guo18, GPR14, HII15, HSEA15, HDMK14, HHS14, IA15, JVJ12, KA16a, KM17, KS15d, KSZK15, KZ16b, Kub17, KP10b, KPR14, KB15, KR11b, LGM12b, Las10, Lic20, LMZ12, LAM14, MS14a, MK13b, MK13c, MA16, MKT17, MBMB12, Muk13, NA20, NBR10, NA16, NmBV13, ND15, OJR11, OV13, OAA13, PUW13, PP13b, PT13, PS13, PB14, PGS13, Rad13, RMP17, SDS20, SA15b, SMAM15, SSN18, Sam13, STS19, SP11a, SS12c, SR23, SHD13, SS17b, SB15, SB11a]. Method [SGS18, SVRB15, SC16, SU20, SR15c, TS10, TR12, TIT14, TRSK15, TM18, US12, VR12, VR15b, WAG11, ZZW24, Zhu21, ZC24, RPMH24, Sal14, VS10]. Method-based [GEAE15]. Methodological [TERCM15]. Methodologies [BA12a, CKS13]. Methodology [DDC^{+}24, GBHI12, GRC13, JSC12, KKBS12, KSS^{+}14b, KKSS12, KWM16, LRRR11, NRL17, Nta14, She11, SDB15b]. Methods [AGN14, AR17b, AKG20, AC15, BA15b, EME15, GE17, GSP18, GF17, GMAR14, HJE14, HK19, HS24, İHŞ17, JZS13, JD15, KSRS12, KK11, KJK12, KK17b, LG11, NVV16, NAZ17, PN14, PP13d, PSM15, RFAH14, RH15, RS14a, RMS14a, SN15a, SA13a, SNA15, SSS13a, SAK13a, SO14c, SGJ12, TAS^{+}16, TR14a, TB15]. Metric [ACK12, AHH20, GL16, KBM22, MR13a, MB13b, PR10a, RM11, LG13a, WH21, WRJ23, Pou13]. Metrics [BIE^{+}12, hC13, GC12, KS16c, KS12b, NP12, PFS12, RKS17, Sab16, SHR11, SYR13]. Metro [EAa14]. MFA [Liu24]. MFCC [DB10]. MFSPFA [KA12f]. MGS [SSA11]. MGS-CM [SSA11]. MHD [DSJ11a, DGJ12, HIU11, HD14, HGK15, ILG15, SH17, SDJ13, SK15b, TH15]. MHDmixed [GDJ12]. mHealth [HPC13]. MHz [KIA15, LGM12a]. Micro [ASS17, IPB14, Nad14, Rok20, RHT19, SS11, MNB^{+}22, SKJ15, TAY15, VMA23, YAT15b]. Micro-Actuator [Nad14]. Micro-Mouse [MNB^{+}22]. Micro-services [VMA23]. Microarray [KBK10, RKR15]. Microarrays [KE14, SSA11]. Microbes [LT12]. Microbial [AHH14]. Microcalcification [ZPAQP11]. Microcalcifications [Zyo12]. Microchannel [HAT16]. Microcontroller [Ibr16, JVJ12, LSM16, SS10a, SMK17]. Microcontrollers [KNV10]. Microelectrode [SM13a]. Microfluidic [SET15]. Microgrid [DGH16, JPAK11]. Microgrids [DDL16]. Microparticles [SET15]. Microphone [SC19b]. Micropolar [BH17a, SH17, TH15]. Microprocessor [Afr17]. Microservice [SP18b]. Microservices [PF20]. Microstrip [AK18b, AI13a, DJA13, KNGT10, PVR20, PB16, PK12b, RC14, SA16, SSG14, TRDB24]. Mid [Alg19]. Mid-Point [Alg19]. Migration [DBB12, MB15a, NHS11, NRR^{+}10, RS16, RGJ11, SS13a, SMS16]. Mihailov [KNT11]. Military [ATB12, JSD14]. Mimic [MBKC15]. MIMO [AJ15, AVRN12, AMN12, Ade13a, ARZ15, AB12b, AGZ15, AAaN15, BSZ13, BBD13, CWC15, CZW15b, Dar12, EE13b, KA12a, MAM19, PDK12, RPN14, SR13b, SB11c, VK11]. MIMO-OFDM [AJ15, Dar12, PDK12, VK11]. Min [KV14, MMA13, SMS16, SKC15, RSS^{+}11]. Min-Max [KV14, SKC15]. Min-Sum [MMA13]. Minahasa [KDT16]. Minahasa-Indonesia [KDT16]. Mind [PF20]. Miner [CV13, OFE15]. Miniature [SWH16]. Miniaturized [CHW16]. Minima [MP12a]. Minimal [BM14a, BP11, DM18, GM14b, LMZ12, RA10a]. Minimization [CCG12, DSS13, GSB12, mH14b, Kei14, KR10, KK14, KS14c, OE14, SV12a, SI13, SV14b, SZS15, UR12, BR11a]. Minimize [AMTM11, KC14c, MJM17]. Minimized [SAk13b]. Minimizing [ARP17, OHI15]. Minimum [AT15, Hus16, LRRR11, Mal13a, MM18, NS15b, SKA13, GM16, FD14]. Mining [AAN^{+}19, AHK14, ABF14, AS18a, ADFAJ15, AA13d, ASH15, AP16, ASS15b, ASH13, AE17, BAM13, Bej16, BS16, CTG12, CRBJ12, CV13, DP11, DD20, DM17, DKP10, DTS16, EmErB13, EA11, GSP18, GK12b, HMC12a, HF20, HFSM12, HMMN13, Jag15b, JARM13, KPS11, KBTB12, KV15, KM16d, KSC11, Mah13, Man10, MM20a, MM18, MBM13, MJ12, MM12b, MH13b, NP13, NJ16, OFE15, PS21, PRR15, PP12a, PP10, RJSJ16, Raf14, RPV10, RKR10, RRudS12, SsTM12, SBKS16, STS19, SSS13a, SAAM17, SA12a, SP15d, SPL15, SBM^{+}15, SSP13b, SASS11, SMT17, SL15, SV12b, SVJ14, SKZ22, SVB23, TSE17, TKC^{+}13, TK15, Ver17, WK16, Wis13, YK16, LT12]. minislot [DCM15]. Minitab [PRK17]. MINPHIS [GSA12a]. Minutiae [KBK12, UAA16]. Minutiae-based [UAA16]. MIPv4 [THEE11]. MIPv6 [THEE11]. MIRFCM [MD14a]. Misbehavior [AJSG11]. Misbehaviour [GJ15b]. Mishaps [SY12]. MISING [RHK15]. MISO [PT15]. Missing [EMA11, KBK10, RP15, VG12]. Mission [OHI15]. Mitigate [AJ13, SG13a]. Mitigating [AK16, AK19, LT13, PP13a, PS15a]. Mitigation [Agn15, HKU19, MAARIH16, PV17, SPP13]. Mitosis [UNM18a]. Mixcolumn [KB16b]. Mixed [BSZ13, BH17a, BWS16, Isu16, LK16a, SJK21, ST14]. Mixed-Bounce [BSZ13]. Mixed-Radix [Isu16]. Mixing [QCERM15, SSAK10]. Mixture [PSD13, RHM14, RRR17, RPR12, SB21, SK15b, TD16a]. MLP [SP11b]. MMAC [HS15b]. MMP [NYS15]. MMP-13 [NYS15]. MMSE [Agn15, BBD13]. MMSE-FDE [Agn15]. MNEWQUE [SN10a]. Mobile [Abd18, AARA19, Abd14b, Abu17, AOAG13, AOO20, ASC15, AH17a, AzSKM12, AR17a, ARHERZ10, AAA14b, Alt11, AA12f, ALH^{+}24, AB11, AV10, AV13, ASD16, AHAPT15, AER19, AM10, AK15, Asa13, AB20, AS15d, BBM14, BJ12, BSZ13, BD10a, BGS11, BSI12, Bey14, BS12c, CS16a, CGKP13, DD14, DB13b, DKP10, ES13, EGES15, El 14, EHHM17, EHI17, Far24, FM13, FCW14, FTLK14, GEAE15, GE17, GJG16, GBPTEQ^{+}13, GE12, GCS14, GPB22, GDD13, HM13, HKA15, HPC13, HMB13, HS15b, HAb14, IE12, KS15b, KP16, KS15c, KS16a, KMS15, aKNP^{+}15, KA12e, KNI14, KHK13, KN12c, KBT^{+}21, KDTD11, KR11a, MV12a, Meg10, NF17, NV12, NM16, Nur11, OK14, PP13a, QA14, RK16, RBM19, RBO19, RC12, SEb13, Sal19, SN15c, ST12, SS12c, SF14, SP18b, SDW21, SAR18]. Mobile [SBHM13, Sha14, SB15, SSG11, SM13c, SE10, SV14b, SZUSNK19, SPP16, SYP10, Sul17, TX20, TTA10, TS15, TE16, UR16, VjS12, VG16, WT13, Yak18, YSH12]. Mobile-Phone [AzSKM12]. Mobile-to-Mobile [BSZ13]. Mobiles [Hus12]. Mobility [AJSG11, AASAA15, AAMG13, AB20, GB14, KS12b, KA13b, Kum10, MKN12, Mah19, MFA13, RS13a, RMR14, RMS14b, VM14a, WAmMS14]. Mod [AAH21]. Mod.B [PCKD17]. modal [SP15e]. Modalities [KGK10, MS12]. Mode [AA12a, ANE^{+}13, ABS^{+}12, BWS^{+}12, DXT23b, EKA13, ELDM13, GAA15, JPAK11, KN12b, KSlAM12, LSB12, SFA14, SKM16, JPP12]. Model [AN16a, AAN^{+}19, Abb12b, AE13, ASS17, ASP17, AB14b, Ade13a, AOAG13, AM19a, AA17a, AKH15, AHH20, AMAY20, AZS16, AB13b, AlS24g, AA18b, AAAA21, AW20, ASC14, AI13b, ASN12, AJM^{+}15, AAoS14, AR16, Aly13, Alz12, AAMG13, AH13b, AB11, ABAO24, AKS^{+}13, ASH13, AE17, Ask14, AAR11, AGH14, AZY18b, AMS14, BSS12, BSZ13, BLA12, BBZG^{+}17, BBC12b, Ber11, BM13c, BSP11, CJC19, COWK14, Cha17, CJK17, Dak23, DI10, DV16, DF15, DB17, DDDD12, DP11, DS13, DKS14, EKES20, EBAS11, EUEI23, El 15, EIB^{+}11, EJC16, FJM16, FPS16, FH13, GCK17, Gho11, GK21, HJBJ11, HF16, HJG^{+}13, HKA15, HZK12, HHEY19, HHHK21, HMB13, HS14, HIS15, HMC12b, IHH14, IUN10, IM15, JIF12, JND15, JH18, JV17, KUM13a, KRMB10, KHSE12, KMMM12a, KMMM12b, KSZK15]. Model [KKI16, KNS17, KLS13, KGAM10, KNT11, KAS13, KY15, KRR15, KK22, KA12g, KMG20, LABV21, LRL17, LSSD20, MKN12, MR12a, MG17a, MMG16, MP12a, ME15, MAP13, MABE16, MDNA12, MCA14, MZ18, NSM13, NR12, Ngu22, Nim15, NAH11, OAA14, OAR16, OC20, OAA24, OE14, OJR11, OS16, OON17, ONJ13, OSS14, OO14, OA12, OKS18, PS14a, PP12a, PPR17, PP13c, PH23, PMH10, PA23a, Pra14, Ps14b, PS12b, RBF12, RSSV14, Rah12, RRR13, RMD10, RSK15b, RRR17, RPR12, RD17b, RK21, fSG13, SEb13, SA15a, SDZ15, SO14a, SK12a, SN15b, SOOA13, SPBP17, SRSV14, SS10d, SSAZ13, SSMP13, SKG14, SS14e, SBS13, SP11b, SBM^{+}15, SRK10, SM12b, SMK14, SV18b, SGJ12, TSRM10, TT14, TN14a, TUS20, TAM19, TYSS15, TD16a, TM13, UR12, USC12, CCE20, VN10, YK16, YJ11, ZRE15]. Model [ZRE16, ZRZ11, ZJR^{+}19, HLZ^{+}24, Ojh11, OS14, PPM12, Sal14, Sri12]. Model-based [SA15a]. Model-View-Controller [TUS20, USC12]. Model-View-ViewModel [PA23a]. Modeling [AEk13, ATM12, AB13a, AAO22, AEKA11, AERS11, ADS16, AVS13, BAF15, BB12a, BJS14, CS16b, CS11, CWT15, CS15c, EAS16, Eln14, EB21, Fer14, FK15, GA15, GH14, GYP^{+}15, HJT14, HWT^{+}13, HTM14, HW18, JHP12, KM12, LNM24, MBM14, MM20b, MBKH16, MAMS12, MAM12, MG17b, NCL11, OAR16, PP17b, RH14, RKS17, Roc17, RKY12, RML11, RMH20, Sak15, SEE13, SCH21, SRTB11, Sel22, Sha13, SSAZ13, SHS23, STG13, SKJ15, SC15c, SAH12, UDA13, YB12, VS12, YB15]. Modelling [Apo19, JPAK11, OFE15, RAC12, SET15]. Modelling-based [OFE15]. Models [Agg17, AEA17, ASR15, AAqA14, AASAA15, AMM21, BM13b, CS10, CDC18, DSM^{+}20, EER14, EE13a, ETI14, GS12b, HK18, HSK14, HK11, KA16a, KRH13, KKR^{+}14, KS12b, Mah19, MM15a, Mes23, MRJN14, MRG19, RP24, SDB15a, SH15, SBS13, SD21, SVMLb13, TAP15, TB15, WAmMS14, ZSR14, GSCT12]. Moderation [AAYM13]. Modern [ESE16, BM20b]. Modes [GGM23, MS17, PZN21]. Modification [DBM14, TSS12, XRL14]. Modified [AmH13, ASY17, Agn15, ABD22, BK14, CD16, DMM14a, DDC12, DNP15, DJA13, EDR15, EAO15, FM20, JS18a, KA16a, Kam17, KK20, KC14b, LGC11, LGZX17, MR12a, MAM15, MG14a, MP12a, MK19, MAMA16, NK15, Ngu23, SE14, SP11a, SOT18, SHD13, SP12, SM12b, SV14b, SK14c, SKZ22, UNM18b, ZTL^{+}14, Hos12, USK17]. Modular [Ira12]. Modulated [KHSE12]. Modulation [ARZ15, AE12b, AAaNJ13, BR11b, DM14, EME13, MS17, PAH17, SA15d, SNJ16, Suj14, SB11c, ZBI14, ASK17, GG12]. Modulator [SA12e]. Modulators [SA12e]. Module [AB13a, RR18, RSS^{+}11, SSK13a, ZAAS18]. Modules [KN12b, PPK13, Tah14, ZAAS18]. Moduli [MB17]. Moisture [JYP10, SO13]. Molecular [RP10, SW17]. Molecules [SPB12]. MoMath [Sha14]. Moment [SH12]. Moments [EADEA17, MEE15]. Monetary [CTG12]. Money [DMM14b, HHS14, IPB14, Mal13b]. Monitor [AH17a, AsJ12, Deb20, PGPM23, PP16, SKGV17]. Monitored [AB16a].

Monitoring [ATH12, AZA23, AEA17, AU23, AKSEB15, DDL18, DBW19a, HZLL15, JJN14, JYP10, KAqC14, KDM17, LSM16, MO17, MMTS13, MNR14, MMK21, MSA18, ONK24, OAO18, RKE23, RV12b, RRB16, SSA^{+}23, SMK17, SS15b, SSA16, SGD^{+}20, TAMA13, TY13a, UN13, UO23, WKG^{+}18, WNE21, ZAE18]. Monitoring/Diagnostic [MNR14]. Monodomain [NR12]. Monolithic [PF20]. Monopole [FSG17, JD13]. Monopoly [NS15b]. Monopropellant [AMN15b]. Monoxide [SMK17]. Monsoon [RM15]. Monte [Lic20, MSEEs13]. Moodle [SF14]. Moranda [MR12a]. Morocco [ZAE18]. Morphological [AAB22, DSDD12, LT12, NmEM13, RK14, TT11, TMK^{+}10]. Morphology [MPR14, Kum12]. Mortality [NS14]. Mosaicing [AMS16, RB11]. MOSFET [Alg19, YR12]. Most [Has15, NM20]. Motif [Bar14, Kon15]. Motion [AH13c, COWK14, DT22, HMC12b, KDC12, KH15, LBNP12, MBKC15, MRP15, SM17a, Sar13, Sha13, YSH12, Zaf11, ZYL^{+}21]. Motivation [WDN^{+}20, SAS24]. Motor [AA12a, AH16, ANE^{+}13, AAEEA15, AHK12, BBC12a, BSP11, CPSP13, DD21, Ena12, GAA15, HR16, HAA16, HTM14, HBN12, IHH11, JTA19, Mis14, MBZA16, NK14a, NKM12, RUB11, SV12a, SHSA12, SO14b, SSP15a, SMC17, VGJ12, SND13]. Motor-Generator [DD21]. Motor-pump [ANE^{+}13]. Motors [BFA13, DC22, SM17a]. Mouse [MNB^{+}22]. Move [DDG16, SB11a]. Movement [HR22, Lic20, MC16, NM16]. Movements [AYE13, BAK20, GRS22]. Movie [CS16b]. Moving [EBO13, FM20, HGK15, HBN12, HMC12b, KY15, SH17, SDJ13, SK15b, SSV14, VD15, YK12]. Mozilla [PH23]. MPI [WKC11, DHM15, EMF14]. MPI-2 [EMF14]. MPK [AS12d]. MPLS [AICN14, Moh13a, RSJT14]. MPLS-TE [AICN14]. MPPT [ANE^{+}13, KM15a]. MPSoC [CLBA^{+}12, SS14c]. MPSOCs [SB11b]. MQ [SOME16]. MQ-9 [SOME16]. MQS [NNR16]. MQTT [DDL18, SDG18]. MR [CA11, DB13a, ESAA18, Las10, SErEs^{+}12]. MRC [AD18]. MRI [Aa13c, ASH16b, HSD13, KM17, KB12b, RMD10, SV18a, SDB15b, SCK17, TV13, ZD15]. MSA [SDD14]. MSAA [AAGM15]. MSAProbs [EHSA23]. MSFLA [Kei14]. MSLDIP [AS12d]. MSMEs [RHT19]. MST [AGL17]. MST-based [AGL17]. MTSF [MC17b]. Muller [Far11]. Multi [AM11, APU16, Abu17, AMN12, AAC16, AA12b, AA12c, Akg13, Am13b, AB13b, AMA14b, Ala16, AS18b, AMAB15, Alh12, AI18, AKG20, ASS15b, CK13, AHK12, ASH13, AF15, ABBG14, BD13, BAF15, BL15, BBD13, BKB11, BMRA12, CW23, CJC19, CTG10, CCG19, DR15a, DM10, DVS13, DMM14b, DR15b, DAG16, DCM15, DA12a, DL23, EME15, FKT12, FS11, FEbO13, FK12, Fla18, FTLK14, GJM16, GGX12, HAAT16, HS15b, HTPB11, HBA13, HA17, IA15, JS14, JV17, KK20, KRP12, KS16b, Kha13, KH12, KST16, KVK19, KC15a, KSS13, KSL^{+}18, KR11a, KAS12, LsAAA13, LSB17, LB11, MS14a, MMP21, MR15, MD14a, MT13b, MAM16, MHE13, NS13a, NK12, OOE18, Ojh11, OS14, OHI15, PA14, Par14a, Pat14, PA24, Poo22, MN22, PBG15, PMM18, RPV10, SK14a, SM17a]. Multi [ST16, Sal18, SCD10, SK18, SP11b, SM13c, SP15e, SD15, SVJ15, SO14c, SV10, TO13, USR17, VRS12, YK13, ZCZ24, ZJR^{+}19, Sal14, Liu24]. Multi- [MD14a]. Multi-Agent [BKB11, CJC19, OOE18, Sal18, TO13, FTLK14]. Multi-Algorithm [AM11]. Multi-Area [LSB17, Par14a]. Multi-Attribute [PA24]. Multi-bank [CTG10]. Multi-Beacons [YK13]. Multi-Beam-MIMO-MC-CDMA [BBD13]. Multi-Bit [Fla18]. Multi-Channel [HS15b]. Multi-choice [HBA13, HA17]. Multi-Chromosome [KSS13, KSL^{+}18]. Multi-Class [GJM16, JV17]. Multi-Classifier [KST16]. Multi-Cloud [AAC16, MR15]. Multi-Colony [FK12]. Multi-Core [Akg13, HTPB11, RPV10, SK14a, LB11]. Multi-Criteria [APU16, AMAB15, AKG20, Sal14]. Multi-department [KSL^{+}18]. Multi-Depot [FS11]. Multi-destination [CW23]. Multi-Dimensional [DR15a]. Multi-Directional [NK12]. Multi-document [SO14c]. Multi-factor [Liu24]. Multi-Focus [KS16b]. Multi-Gen [ST16]. Multi-Groups [GGX12]. Multi-hop [SK18]. Multi-Hose [FK12]. Multi-label [EME15]. Multi-Layer [LsAAA13, SP11b, FKT12, HAAT16]. Multi-layered [AHK12]. Multi-Length [AS18b]. Multi-Level [PBG15, BL15, DM10, KR11a, MMP21]. Multi-Machine [AA12b, AA12c]. Multi-minislot [DCM15]. Multi-Mobile [SM13c]. Multi-modal [SP15e]. Multi-Model [ZJR^{+}19, OS14]. Multi-Object [MN22]. Multi-Objective [DR15b, DA12a, FTLK14, MS14a, OHI15, Pat14, SVJ15, AI18, AF15, FEbO13, SCD10, SD15]. Multi-Objectives [Abu17]. Multi-ontology [BAF15]. Multi-Path [AMN12, KRP12]. Multi-Perspective [KVK19]. Multi-phase [AB13b]. Multi-player [BMRA12]. Multi-queue [MAM16]. Multi-Relational [BD13]. Multi-resolution [KH12]. Multi-Robot [Ala16]. Multi-server [JS14]. Multi-Slotted [MT13b]. Multi-Social [USR17]. Multi-spectral [ZCZ24]. Multi-Stage [ASH13, ABBG14]. Multi-State [MHE13, Poo22]. Multi-temporal [AMA14b]. Multi-Threading [DL23]. Multi-Types [KC15a]. Multi-UAV [KK20]. Multi-User [BBD13]. Multi-Warehouses [DMM14b]. Multi-Wavelet [CCG19]. Multi-wavelets [DVS13]. MultiAuthority [PP13c]. Multiband [KNGT10, LK16b, TRDB24]. Multibeam [Ade13b]. Multiblocks [SK10c]. Multicarrier [AAaAD12, AAaN15, IEES16, IS21, TCS14, VM14b]. Multicast [AV10, KK16b, STM13, SFVA14, SI13, VR11]. Multicasting [AMBM11]. Multichannel [DT14, LCP21, SCL14]. Multicomputers [BA11]. Multicore [SSS15a, SAK18]. Multicorns [DLN12]. Multidimensional [DSM^{+}20, HBA13, HA17, JNR19, LGC11]. Multidisciplinary [AMN15b, AMN15a]. Multifont [ME14]. Multigrid [PMH10]. Multihop [ESE13, KV16a, KPAB14, KK16b]. Multihop-Multicast [KK16b]. Multilabel [AS18d]. Multilayer [AM19a, SM13a]. Multilevel [AH16, MKOS14, BP15, ESE16, JVJ12, MK18a, MK19, RAA14, SM10, SKT14, SMC17, Suj14, TAM19]. Multilingual [GJ13a]. Multimedia [AAY21, AAR11, CS16a, GRP13, HKL16, HAb14, IM12, KDS10, Kum13b, NAH11, OJO11, Rus15, SJ19, WAM12, XRL14]. Multimodal [AOSM12, ASK15, AS13c, DM10, EAJ^{+}15, El 14, KGK10, MR11, RG21, SS10f, TD21, UJ13]. Multiobjective [DD13, OESKF16]. Multipath [AAaN15, BSG15, IS21, SCD14]. ’Multiple [FYM11, AH12a, AD18, AkAA14, AlS24h, AR17a, AAaNJ13, AS15d, EAEE15, EE13b, GM15d, GF17, GRL20, VNM15, JBK13, Key12, MM18, MBL15a, MDNA12, PB15, PJ12, RR11, RK15, SDB15a, SSA11, SSK^{+}10, STG12, SAK18, SS15c, SPXT17, TR15, TS12, THAP19, WRJ23]. Multiple-Server [SDB15a]. Multiplexer [SSS13c, TMS15]. Multiplexers [EAa14]. Multiplexor [MM17]. Multiplication [DN12, MFM12, RMS14a, TFT11, VRS10, WKC11, Zhu21]. Multiplicative [SVB17]. Multiplier [BM15c, GSN12, HIE14, RJS^{+}13, SE12a, SE14, SCE14, ZYW^{+}18]. Multiplier-Less [SCE14, HIE14, SE12a, SE14]. Multiply [GSN14]. Multiplying [Key12]. Multipoint [AMNA18]. Multiprocessor [AAC13, SRF12, SJV13]. Multiprocessors [CTG10]. Multiprotocol [AICN14]. Multirate [GBM10, ITK21]. Multiresolution [AJB13]. Multiservice [MMG16]. Multisets [EKI16]. Multispectral [KE16]. Multistage [GJP13, RMSS13, SM15a]. Multistep [SLU14]. Multitemporal [MRP15]. Multithreaded [SSS15a]. Multithreading [IMM12]. Multivariable [ABS^{+}12, TNC10, YMR^{+}12]. Multivariate [BNMM21, NYS15]. Multiview [MS14b, MNB13c, PB12]. Mumbai [KSBM17]. MUMCUT [BM14a]. Municipality [ATA19]. Museum [yGL13]. Music [Dur14, ETKE20, KB15, NKD14, NM13, SSS15b, TD16a, Ade13b]. Mutated [Sha18]. Mutation [HB14b, KSS13, KSL^{+}18, SGS15, SMM11, UA14]. Mutual [MC17a, SSS15a, SS13f]. MVA [Tua17]. MVC [RBN^{+}16, TUS20, USC12]. MVVM [PA23a]. MW [SVRB15]. Myanmar [THT11]. Myocardial [COWK14, MDD21]. Mystifying [BM15a]. MZI [ND17].

n [MAA13, NYS15, SHR11, VRS12]. Naīve [FRR11a]. NAC [NMAB14]. Nairobi [NKM12]. Naive [AU23, HS24, Bef18, GS16a]. Nakagami [GG12, PS12b, SS10e]. Nakagami- [GG12, PS12b]. Name [CD13, hKhHL13, SNA13]. Named [AAA19, DDD12, Far14b, KB14b]. NameNode [AT14]. NAND [SN14a]. NAND/NOR [SN14a]. Nano [DPT14, EH14, KMS12, Pou13, YR12]. Nanofluid [AHZ15b, BWS16, TRM23, WBS16]. Nanorobot [PS10]. Nanostructured [PS12a]. Nanotube [CN12]. Narrow [DRS15]. NAS [KS16d]. NAS-RIF [KS16d]. NASA [SM12b]. Nash [PP12b]. National [OAS19]. Nations [Asa13]. Natural [AAA16, BP16, Far14b, HB12a, HW18, HPM20, HR13, HJ13, ILG15, KM15c, KW14, LCP21, NGA14, PM21, Raf15b, Sal15, SM17b, SK10a, You13]. Nature [AII13, KR11b, SP15b]. Nature-Inspired [AII13]. Naval [CNP19].

Navigation [HM13, LHM15, NH13, RN24, SJ12]. NAWA2 [VR11]. NB [MDCS22]. NB-IoT [MDCS22]. NBRUrp [EADEA17]. NCCCN [KAA13]. NDVI [SS16c]. Near [JAB22, NVV16, SV14a]. Near-end [SV14a]. Nearest [AGN14, AlS24a, Bef18, BHK13, IEME12, KH20, KB16a, OKS18, TV13, YK12]. Necrosis [SCH21]. Nectar [Fer14, FK15]. Need [AKN13]. Negative [FSG17, KC15a, MAE16a, MK14, RSK15a, AAR13]. Negotiation [AS15c]. Nehalem [Ask14]. Neighbor [AGN14, BHK13, IEME12, KH20, KB16a, QT15]. Neighborhood [GG11, KAAM11, NSM13, DBS15, SSV11, SSR14, ZZ11]. Neighbors [AlS24a, Bef18]. Neighbour [OKS18, TV13, YK12]. Neighbourhood [BCC13]. Nelder [CCG12]. Nematodes [Bou12a]. NEMIDSOFT [Bou12a]. NEMO [AB20]. Neo [NU12]. Neonatal [NS14, SsTM12]. Neonate [FZM12]. Nephropathy [LAAI23]. NER [Far14b]. NER-FL [Far14b]. Nested [ESR13]. Net [GH14, JG12, KK14, MM13, SKA13, Sin13]. NetFlow [ARAA11]. Nets [AERS11, LR12]. Network [AOB12, ASD19, ACB17, AAN12, AHZAH16, ADT17, Abd14b, AH12b, AMNA18, AM19a, ASNS10, ATA14, AHZ15a, ASC15, AM19b, AICN14, AZ16, Aq14a, ArAn13, AsS13, Al 16, AR17a, Ali19, AMY23, AGG16, AA16b, AKA^{+}18, AH12c, ASD16, ABAO24, AE17, AB20, AB13c, BK14, BI12, BADR16, BM16a, Bef18, BS14a, BBHL12, BS11a, BH17b, BG17, BV21, BWH15, BJPP15, BD10b, BA15b, BO17, ÇGDT18, CS16a, CPSP13, CWC15, CT10, DT14, DK11, DCH16, DH19, DMS13, DA10, DHT13, DB13b, Dum20, EsGK14, El 20, Ela12, FD14, FSNR18, FSA12, FYM11, FK15, FTLK14, GJ15a, GL16, GMRK12, GAA17, GSB12, GMAR14, GPS14, GGI14, GRS22, GGK14, GJM16, HKA15, HMY12, HaB13b, HWT^{+}13, Hik13, HPR17, HS15d, IE12, IUN10, ISM^{+}17, sNIALY13, IM16, JIF12, Jab14, JTKC14]. Network [JS18b, JM15, KNSS11, KAA13, KV14, KCD13, KG15b, KS16a, KBEA18, KIA15, KC12, KT12b, KD15, KRR15, KB16c, KSSM24, KNV10, KR16, LsAAA13, LCA12, Las10, LcCm13, LDB^{+}22, MWS20, MAE16a, MS17, MK18a, MK15a, MEE15, MMR14c, MH14a, MKCV13, MJS15, NS15a, NJAK15, NS16, NOO16, NM16, NMAB14, OMH11, Obe23, OAG17, OON17, OHI15, PS15a, PSS12, PKB13, PPU11, PGPM23, PS13, PM15, PA12, QKRB11, RS13a, RR11, RSSV14, RM15, RHD12, RP14, RV16, RV12b, RRR11, RV13, RD17a, RMG13, RR12, SHSA12, SE12b, SMSC14, SO14a, SIAI14, SAA24, SN15c, ST12, SSP13a, SRSV14, SOO15, SMJ12, SR23, SK10b, SC15b, SKC15, SSSS12, SF15b, SC13, SS12e, SM13c, SGS13, SC14b, SC15c, SKS16, SSC11, Sin16, SL11, SZUSNK19, SK15c, Sub10, SW17]. Network [SV10, SK15d, SSY11, TT14, TN14a, TBN^{+}15, TAM14b, TMC17, TB14, TMP11, UCNC13, UR16, USK17, UDA13, UO23, Uzo13, Vad20, VGo12, VD15, YWT12, YRA23, YAA14, YJ14, ZYL^{+}21, ZCDR13]. Network-based [AsS13]. Network-on-Chip [SMSC14]. Networked [NKR14]. Networking [USR17, YBM16]. Networks [AII13, AAN12, AAY21, ATM12, Abu14, Abu17, AS12a, ASY17, AOO20, AMG13, AHZ15a, ABB20, AmM12, AZS16, ARHERZ10, ARK13, AQ14b, AKG20, AEKA11, AAA^{+}17b, AMS16, AT15, CK13, AB11, AV13, AYAI17a, AJC14, ANA11, AM14, AS15d, AIK11, MKOS14, BSAb12, BJ12, BT14a, BD10a, BGS11, BBZG^{+}17, BOB17, BS12c, CC17, DS10, DDC12, DSA11, DAB12, DCM15, DTO23, DA18, DKS14, DGP14, ES13, ESE13, EBAS11, EF15, ETLE13, EAa14, Eln14, EG15, ES15c, FKFU12, FSA12, FEE13, FR12, GEA16, GE17, GKM^{+}18, GSD14, GVRP10, GHD17, GK12a, GDS16, GS11, HZS14, Has16, HS15b, HK15, Ibr17, JI10a, JLCV16, JYP10, JAB22, JSK17, JMM21, KA12c, KA12d, KFEN21, KMMM12a, KMMM12b, KA12e, KAAA13, KSV10, KCS14, KP10a, KBT^{+}21, KC11, KPV12, KA13b]. Networks [KT15, LCA12, LSS20, LA15, MM14a, MLRU12, MK13a, MAK13, MFA13, Mal10, MMP21, MP10, MTD14, Meg10, MK15a, MK15b, MMK21, MAM19, MCB12, MGL22, MNHB12, MRG19, MNN10, NCL11, NV12, NM19, NMAB14, OC20, PP13a, PTPS19, PC11b, PSS11, PS15d, PJ12, QA14, RMVP14, RSJT14, RR11, MP17, RS12, RB15, RKG10, RRB16, RJC17, DBS15, SE16, STM13, dSEdZ14, SWP19, Sal17, SM18, SS21, SPP11, SSS15b, SS14d, SA10, SKS10, SU12, SP11b, SSS12, SM13d, SSS16, SAH12, SV14b, SSK13b, SB12a, Sud10, SZS15, SS16d, SJV^{+}15, SSJ12, TBKG13, TBR15, TK14, TAMA13, TOMGBSS13, TAM14b, TCS14, Tit16, TOO16, VT15, VjS12, VG16, VMT13, XRL14, YRV17, YK13, YYS^{+}23a, ZMB14, ZZW24, ZBI14, ZYM13, AV10, PM13]. Networks- [EBAS11]. Neural [ASD19, ACB17, AAN12, Abd14b, AH12b, ATM12, ASP17, AS12a, AHZ15a, ASC15, AM19b, AsS13, AZS16, AGG16, Alz12, AMS16, AA16b, AYAI17a, AE17, ANA11, AB13c, BSAb12, BK14, BI12, BBZG^{+}17, BS11a, BV21, BWH15, CQM15, CWC15, CT10, DK11, DCH16, DH19, Dum20, DTO23, EBAS11, El 20, Ela12, Eln14, ES15c, FSNR18, FSA12, FK15, GJ15a, GVRP10, GRS22, GGK14, GJM16, Hik13, VNM15, Ibr17, IUN10, JIF12, Jab14, JAB22, JS18b, KNSS11, KV14, KA12d, KCD13, KJKM21, KBEA18, KSV10, KSSM24, KRR13, KNV10, LsAAA13, Las10, LDB^{+}22, MWS20, MM13, MLRU12, MK12, MAK13, MK18a, MK15a, MEE15, MMR14c, MH14a, MGL22, MRG19, NJAK15, Obe23, OON17, PKB13, PS13, PC11b, PM15, PA12, QKRB11, RM15, RV16, fSG13, SHSA12, SO14a, SAA24]. Neural [SRTB11, SSP13a, SS21, SSS15b, SKC15, SP11b, SC13, SC15c, SSA16, SAH12, SSK13b, Sud10, SSJ12, SSY11, TT14, TN14a, TBN^{+}15, TOMGBSS13, VGo12, YWT12, ZZW24, ZYM13]. Neural-Fuzzy [KRR13]. Neuro [AST11, AMAY20, AHK12, BWH15, EKES20, KKDC12, KR16, OAA16, PP10, SK12b, SP18c, TS11, ZRZ11]. Neuro- [TS11]. Neuro-Fuzzy [AST11, EKES20, KKDC12, KR16, OAA16, ZRZ11, AHK12]. Neuro-Fuzzy-Particle [AMAY20]. Neuro-Wavelet [SK12b]. NeuroFuzzy [MAP13]. Neurological [KAR12]. Neuromuscular [KST16]. Neuron [ASC14, FYM11, VGJ12]. Neuronal [BB12a, HJG^{+}13]. Neuroph [IU18]. Neutral [PVA13]. Neutralization [SRTB11]. Neutrosophic [DM22, Eis14, JFJN20, SEF17]. Newly [SS12d]. News [AB16b, GS12a, KVK19, MD13]. Newton [WAG11]. Newtonian [HS14]. Newton’s [AFW14]. Newton’s-Like [AFW14]. Next [FCS11, RSJT14].

NextGen [BK10b]. NFC [MMG16, SRS14b]. Ni [PS12a]. NIDS [PA13]. Niger [BI12]. Nigeria [ASD19, BI12, ED19, KAL15, LOAO23, MBL15b, OE14, SSO14, ST12, UI16, WBK22]. Nigerian [OC13, RK13]. Nine [NSMK17, SKS13, SS12c]. Nine-Element [SKS13]. Nine-Level [NSMK17]. Nine-Point [SS12c]. NIRR [AAS14]. Nitric [PNIS16]. Nitride [CN12]. NLO [PVL10]. NLOS [TSCD15]. NLP [Upa23]. NMF [KPB12]. NN [Dum20, JH18, SSY11]. NN-OPF [SSY11]. NN-SOM [Dum20]. NNAPC [Kal11]. No [Ibr17, LK16a, MJM17]. No-wait [MJM17]. NoCs [MBKH16]. Nodal [MAMS12]. Node [AJSG11, BKK13, BKS14, EE16b, FKFU12, FR12, HKA15, JTKC14, KA13b, KRR15, PIB11, RS13a, RK17, RSRT20, SRSV14, SF15b, SJV^{+}15, Tha13, USK17, SKS16, Tri10]. Nodes [BD10a, GKG16, HZS14, MFA13, SM13c, TOO16, SRSV14]. Nodules [ZJR^{+}19]. Noise [ASK17, GC18, Guo18, HCL15, JPP12, KG15a, KVRS14, KJK12, KY15, MSN12, MRR21, ML12, OJR11, PBK14, RAJ^{+}12, RSA16, RAR24, SV14a, SV18a, Sam13, SN14b, STD14, SSR14, SKT14, SN10b, ST14, SR15b, US12, Uzo13, VRS12, AB12a]. Noises [BOS15]. Noising [AHM14b, KAP11, RVKB12, SM15a]. Noisy [DB10, FAt13, GK15b, Isl18, RSHI12, YRV14]. Non [AhAT13, AAR13, Anh14b, Anh14a, AVS13, ARDS18, BS12a, BSZ13, BBA12, CZW15b, DSDD12, DSM^{+}20, DKB12, EAB13, HS14, JPJ18, JTKC14, KSRS12, KKR18, KM17, KP15b, KKK13, MC17a, MA12, MVC10, MK14, NAD19, OA12, Ori14, PS15b, RS13a, RAJ^{+}12, SE12a, SB21, SRSV14, SG12b, SS13e, Tha14, TNC10, AAAOG20, KH12, KAS12, RMS14a]. Non-affine [CZW15b]. Non-Beacon [RS13a]. Non-Binary [AhAT13, Tha14]. non-Blind [KH12]. non-cryptographic [AAAOG20]. Non-Deterministic [JPJ18]. Non-dilated [BS12a]. Non-Dominated [PS15b]. Non-Drunk [EAB13]. Non-Functional [SG12b]. Non-geometric [MA12]. Non-Intrusive [ARDS18]. Non-invasive [NAD19]. Non-Isotropic [BSZ13]. Non-Linear [KP15b, KKK13, MVC10, RAJ^{+}12, SS13e, DKB12, KSRS12, TNC10]. Non-Local [Anh14b, Anh14a, KM17]. Non-Negative [MK14, AAR13]. Non-Newtonian [HS14]. Non-Preemptive [KKR18]. Non-Relational [DSM^{+}20]. Non-repeating [SB21]. Non-Smooth [Ori14, KAS12]. Non-Stationary [MVC10]. Non-uniform [JTKC14, SE12a]. non-window [RMS14a]. Nonaffine [Alz12]. Nondeterministic [Ala16]. Noninvasive [COWK14]. Nonlinear [AFW14, AZS16, AAA17a, Alz12, ABS^{+}12, BNMM21, BRD12, CWC15, CZW15a, CZW15b, CCG12, Dar13, EEGHS15, EJC16, Eln14, FTBA16, HII15, IM15, JV14, KM15b, KN17a, KZ16b, LLSL12, NSA14, OJR11, SV18a, SK12a, Sam13, TG18, TMP11, TN14b, VGo12, VJR11, WAG11, PN10, Sal14]. Nonlinear-Quadratic [JV14]. Nonlinearity [KR11b, YJ11]. Nonsubsampled [NP11]. Nontraditional [VP10]. NOOR [CRA12, CKD12]. NOR [SN14a]. Norm [DR15b]. Normal [ETG17, RRR17, YA18]. Normality [AKS16]. Normalization [AHK14, DM22, KA16b]. Normalized [AS13a, NRV12, PR16]. North [ASD19, PPR17]. North-Central [ASD19]. NoSQL [FLLR20]. Notched [SEE13]. Notification [DDM^{+}23, You18, MN24]. Notifier [SF14]. Notion [EnFn14]. Novel [Abd17, AHH13, AB13a, AA11a, AAY15, ASK15, AMTM11, LG13b, BNMM21, BSS12, BSZ13, BK10a, Dar13, DDB11, DCM15, DS12c, DKP10, Far14b, Far17, FIS16, HCL15, Jay15, JLCV16, JS15, KV14, KB12a, KC14a, KV15, KA10, Kha11, KM16c, KSS^{+}14b, KSC11, KAS12, KGK13, KK14, KD15, LGM12b, Mah13, MM12a, MJ15, MSN12, Mes16, MT13b, MRJN14, MIS12, MSA17, MG19, NTY19, Ojh11, Ori14, PR10a, PT15, PV17, PP10, PK12b, PC15a, PB14, Pou13, PSM15, RMAM19, RJS^{+}13, RB11, RKR11, RW20, SEb13, SB21, SS10b, ST16, SN15b, SSK^{+}10, SK18, SA12b, SHR11, SSD^{+}13, SS17b, SP11b, SB13b, ST14, SA12d, SU14, SR15c, VR11, HBS12]. Novelty [MA10]. NPC [OA12, OAA13]. NS2 [CN10, KC15c]. NS3 [SJ14]. NSCT [ASK15]. NTRU [BM20a]. Nuclear [DJ12, SNB18]. Numeracy [AI13b]. Numeral [JAsA13]. Numerals [EEL11, KK13b]. Numerical [AA14a, AM17b, AE12b, Ba13, BWS16, CDC17, HF16, HAT16, HII15, Job17a, Nim15, PP15, Raf15a, Raf15b, RAC12, SM14a, ŞAG17, UA14, VR15b, WBS16, YAT15b]. Nutrients [MGA20]. Nutritional [SH23]. Nyquist [MAH15].

Object [AC15, BSAb12, BP16, BECM11, BA12b, CS18, DPS21, DKP21, EBO13, FML20, FPA13, GGK14, HAb14, Ira12, KS16c, Key12, KY15, MKK13, Mar12, NJAK15, NP12, OAR16, PP17b, MN22, RV12b, RKS17, RA12a, SAY16, SJ12, Sel22, SBHM13, SB13b, SM19, SKZ22, TGR12, VD15, YK12]. Object-based [FPA13]. Object-Oriented [DKP21, RKS17, RV12b]. Objective [AS11, DR15b, DA12a, FTLK14, GRC13, MS14a, OHI15, Pat14, PMM18, SVJ15, SV10, AI18, AF15, FEbO13, SCD10, SD15]. Objectives [Abu17, Alh12, GF17, STG13]. Objects [FNH13, HMC12b, Kor23, SSV14, SM19, KZ16a]. Observation [AMBM11, FM20, NB11]. Observer [Bou13, IHH11, LEE13, TG18]. Observers [FTBA16, NEA17]. Obstacle [HM13, SBHM13, YSH12]. Obtained [Kam17, MM14b]. occidentale [AGG16]. Occluded [CM15a]. Occlusion [A17]. Occurrence [AZY18a, LAAI23, TS10, MGB11, PPP16b]. OCR [BO21]. Ocular [WSW^{+}13]. Odd [Bad13a]. ODEs [HII15]. OE [JM15]. OFDM [AMN12, AJ15, AEE15, AJ14, AAaAD12, Dar12, HO15a, HS12, MFE13, PT13, PDK12, SCD14, SSG11, SGD13, SNJ16, SSU10, TCS14, Tha14, VK11]. OFDM-BPSK [SSU10]. OFDM-IEEE [SSG11]. OFDM/WOFDM [AMN12, AMN12]. OFDMA [RM12, VT15]. Off [WBK22, TS14]. Off-Grid [WBK22]. Office [AM17b, BA16, Kad16]. Officers [WBL^{+}23]. Offline [DI10, RRVP12, RMAM19]. Offloading [GEAE15]. Offset [MMA13, SRS14a]. Offshore [PA19]. Oil [ATH12, AHH14, BS11a, EOO22, JAB22, JGZ13, MM20a, MM20b, PA19, SABA12]. OLAP [VSA15]. Oldroyed [BB13a]. Oldroyed-B [BB13a]. Olive [ATH12, JAB22, MM20a, MM20b]. OLSR [RSRT20, SZUSNK19, ZTL^{+}14]. OMR [Am13b]. On-Chip [Fla18, KRMB10]. On-Demand [ASD16, MK13d]. On-line [MAK13]. on-the-fly [RMS14a]. One [AVT13, AJT17, AK18a, BBA12, GJM16, TFT11, TG18, TY13b, AB14c, MB15c]. One-Against-Rest [GJM16]. One-Click [TY13b]. One-link [TG18]. Online [ASS17, AGFA16, AS18a, BA15b, Dak23, DP11, GS12a, HN16, ISM^{+}17, MM15c, NM19, OEA18, RK16, RDD^{+}12, SS12a, SS14d, SJ19, Tit16, YBM16]. Onlooker [KSK14]. onto [BBL15]. Ontological [DP16, ESB13, SS13d]. Ontologies [ENE14, TERCM15]. Ontology [ADH17, AGH14, BBL15, BS11b, El 15, FLM11, GBHI12, HEBR11, JJA14, KTA17, KB12a, NHMR13, OV13, PP17a, RV13, SS10b, SU14, VK13, ZK12, BAF15]. Ontology-based [BS11b]. ONU [IhZ12]. OPAP [ADVB10]. OPC [DGH16, DDL16]. OPC-UA [DGH16, DDL16]. Open [AA18c, ATAAAF19, AZ16, AOK18, BT16, DLR19, GAA17, GS14, HN16, KS16c, NA16, SMC17, WSW^{+}13]. Open-Ended [SMC17]. Open-Source [DLR19]. Opening [Kau15]. Opening-by-Reconstruction [Kau15]. Openings [MB15c, MBR16]. OpenMP [WKC11, BM15b, DHM15, UJ13]. OpenMP-MPI [DHM15]. Opensource [NMAB14]. Operant [Fer14]. Operasional [TUS20]. Operated [MM17]. Operating [ARL16, EhMA13, Mun14, NRR^{+}10, SS17c, YB12]. Operation [DD21, KM12, LS17, MB11a, YA18]. Operational [DVS13, KK10c, LGM12b, MIS12, SMT19]. Operations [Has24, Naj12, PR10b, RGR15, RMS14a]. operative [FSER20]. Operator [AOM12, NSS17, SJA13]. Operators [KAL15, KJ12b, MBL15b, OJ11, TMK^{+}10, TR13]. OPF [MB15b, SSY11]. OPFET [LG11]. Opinion [MMTS13, MJ12, SA12a, WK16]. Opinions [Sab16]. Opportunistic [ASY17]. Opposing [MB15c]. Optic [AP15, Nta14, TRR16, WHP14]. Optical [AS14a, AM14, BOS15, CCG12, DDC12, EAa14, GEB12, GS11, KJA15, LSB12, MM15c, MSA16, MSA17, NDC10, ND17, PN10, RHU12, RK19, Sam13]. Optimal [AR10, AMAB15, AAA14b, AJI15, AR16, ADS16, BHK13, ED18, Es12a, EHI17, FTBA16, GKG16, GK15a, HES14, HS10, mH14b, IK11, Iss13, JV14, JI10a, JM18, Jas13, KSA20, KNI14, KK17a, KBEA17, KAS12, KJKK12, KRM12, KAS13, KK14, KNR15, LRL17, MG14a, MA16, MBL15a, NHH10, NSM13, NR12, RSK15b, SsTM12, MS12, SHHL13, SMVK11, Sal19, SSC11, SSY11, TAH13, TArM14, ZRE15, ZRE16]. Optimation [HR21]. Optimised [PT13]. Optimising [RA10a]. Optimization [AII13, AA14a, AS17, AHAS^{+}17, ADT17, AA13a, AMN15b, AMN15a, ASS10, AA12b, AA12c, AM17a, AMAY20, AA18b, AAZN21, ALH^{+}24, AMB11, AYAI17b, AK15, AF15, aAO14, AZ17, BAS12a, Ber11, BH17b, BM13c, Bon15, BG20, Cha14, Cha15, DCH16, DR15b, DDL16, DB12b, ED18, FEbO13, FR14, FCS11, GHBN12, GSN12, GEB12, GMAR14, GG11, GQD^{+}22, HB12a, HNW17, HHHK21, HKR17, wHsYwT12, Hus12, ITK21, JM18, JTA19, JSY14, JIA13, JK14b, MPR14, KA16a, KTN13, KS15b, KJ12a, KS15c, KS16a, KMS12, KK17a, KSL^{+}18, KK14, KC14c, LGC11, LSB17, MS14a, MM14a, MG14a, MG14b, MM20a, Maz21, ME15, MM12b, MCaC14, MB13c, MNB13c, Mug24, Muk13, MG19, Ngu16, Ngu24a, Ngu24b, PCR12, PD14, PDK14, PSD13, PA23a, RS13b, RP14, RHM14, RB14b, RKEdK14, RK17]. Optimization [RR12, DBS15, SKS13, SAY16, Sar16a, SE15, SMHS10, SSS15b, SR13b, SAP16, SB18c, SV10, Tah17, Tha13, TAY15, TAH13, TY16, UA14, UJ13, VM12a, VR15a, YMAS17, YAT15b, Zyo12, AD24, VJR11]. Optimize [AGG16, ARZ11, EE16b, KA16a, MABE16, NKR14, NK14a, WKG^{+}18]. Optimized [BJ12, BM15c, CR15, DBKN20, DSJ11b, Jay15, JMM21, KS16e, KK15c, LCA12, MK18a, MMK11, NH13, NB11, SFH17, SS15a, SM12a, SC13, MNB^{+}22, SM13d, SRS14b, Sub10, TS14, VM12b, Zon17]. Optimizer [DDL16]. Optimizing [AA13b, AS15a, Al 16, AAOA23, EF15, HZ15, KS13b, MNB13a, MDNA12, OP18, RBN^{+}16, SS13a, SSM12, SAKS24]. Optimum [ASS10, AGRR10, BDN13, GSB12, HE15, HMAA13, KA11, LAA^{+}17, ND17, RPN14, SP15d, YAT15a]. Options [GRL20]. Optisystem [BOS15]. OR-Assisted [PG15b]. Oral [PCC12]. Orange [KJA15]. OrBAC [MMG16]. Orbit [BAK14, CRA12]. Orbital [RP10]. Orchestration [Sm13b, TO13]. Order [AERS11, AMTM11, DB10, DVS13, HKE^{+}14, HII15, KB14a, Kub17, KNT11, KNR15, MAaM14, MBE16, OKS14, SO14b, SSR14, SPW15, VR12, VR15b, YJ11, BFD15, PT15]. Order-Fuzzy-PID [MBE16]. Ordered [Eja12]. Ordering [FIS16, SSC11, TS10]. ordinal [NM24]. organelle [SM13a]. Organic [PVL10]. Organization [HAS20, Sou15]. Organizational [BB13b, Qte14]. Organizations [SOOA13]. Organize [RL14]. Organized [SS10b]. Organizing [Dum20, FY12, MK12, NJAK15]. Orientation [AA11b]. Oriented [AM11, ASB12, BSP11, CTG12, DKP21, FLM11, GAZ13, HAb14, Ira12, KUM13a, KS16c, LS13, MMHA11, Mar12, NS13a, NP12, Ojh11, RKS17, Sm13b, She12, SA13b, TB12a, MN24, RV12b, SZZR11, TB12b]. Original [MBCL17, NRR12]. Ornamental [AAO13]. Orthogonal [MN15]. Orthographic [VPGF10]. Orthopedic [PR15]. OSCI [ZKN10]. Oscillating [GDJ12]. Oscillation [RJC15]. Oscillations [Ibr14, LT13]. Oscillator [Kei14, MKB13]. Oscillators [BK17]. Oscillatory [HIU11]. Osmotic [AGG16]. OSN [NP15]. OSNs [SS14d]. OSTBC [DAaGZ13]. Osteo [SM13a]. Osteoporosis [ADFAJ15]. Other [NVR12]. OTP [AAS24]. OTP-JWT [AAS24]. OTSU’S [TM18]. Out-of-Hospital [KAqC14]. Outage [KK16b]. Outbreak [NO18]. Outcomes [ATA19]. Outdoor [MDNA12]. Outlet [PJ12]. Outlier [SG10, SBM^{+}15]. Outliers [RAR24, SP15c]. Outline [CHG12]. Output [AB12a, AAaNJ13, CZW15b, GM15a, RK15, SSS15b, SPW15]. Outsourced [BH14, JS18a]. Overcome [HK15]. Overcoming [Rag15]. Overflow [PF20].

Overhead [AV15, DSS13, JM15, SI13]. Overlapped [AAR13]. Overlapping [VSN13]. Overlay [MTD14]. Overload [KAS13]. Overview [BA12c, JI10b, RKG10, SWP19, SP13, SE11, SASS11, Sul17, TAMA13, TIT14]. OWL [FLM11, PP17a]. OWL-S [FLM11]. Ownership [AAA21]. Oxidation [ABD22]. Oxide [PNIS16]. Oxygen [WNE21]. Ozone [SGD^{+}20].

P [AK11, LCP21]. P&O [KM15a]. P. [RDD^{+}12]. P2MP [IhZ12]. P2P [RMP11, RS12]. P9 [RS13c]. P9-factorization [RS13c]. Packet [AAaNJ13, BK10a, GJ15b, GDS16, JA16a, KC20, KS16d, PVL11, SV10, TIA15, Tha13, ZTL^{+}14]. Packets [ArAn13, Gha13, TY15b]. PACS [OJO11]. Paddy [MBP16, SN22]. PAF [FA13]. Page [AaEeK11, BN14, CS14, SJA13]. PageRank [HG11, SHR14]. Pages [KP10b, KA12g, SA12b, SHR14]. PAIN [Ban11]. Paired [GKG16]. Pairing [SSS13b]. Pairs [GKG16]. Pairwise [KETESH17, OMM14, KEBF13]. Pakistan [AWK19, Mal13b, Mal13c]. Palm [DJG12, DM10]. Palmprint [AOSM12, HMGV12]. Palmtop [NKH12]. Panacea [BO17]. Pancasila [KMT^{+}24]. Panel [SC13]. Panorama [SR13a]. Pansharpening [KE16]. Paper [AAYM13, EDD19]. Papers [DD20]. PAPR [AMN12, HS12, MAH15]. Papua [IA15]. Parabola [YYS^{+}23a]. Parabolic [CCG12]. Parabolic-Index [CCG12]. Paradigm [SS13d, Zar24]. Parallel [Akg13, AS18b, AAR14, L13, BB13a, CS11, DN12, DTS16, EHSA23, EmErB13, GN14, GK10, HSMb12, HMY12, IMM12, KEV14, Kon15, KP10a, MG12, MC17b, RR16, RGR15, RTM12, SAK18, SQ18, SK10c, SGT^{+}11, Sin14, TB13, CCE20, VSA15, WKAL24, RPMH24]. Parallel-Series [MC17b]. Parallelism [SM15b]. Parallelization [BM15b, DHM15]. Parallelized [JK14b]. Parallelizing [RD13]. Parameter [DK13, FA19, HBN12, HMC12b, KC14b, Ngu18, RUB11, ST16, SSAZ13, TB14, ZZ11, ZPAQP11, SV10]. Parameterization [JV14]. Parameterized [BA15b]. Parameters [HAPG17, AsJ12, Aly13, AM14, DS13, DKS14, HSK14, IHB^{+}15, Isl18, JV12, KRR11, KJ14, KMS14, KMS12, KS13b, MDGS13, Maz21, MNB13b, Ngu16, OMH11, PMM18, Ps14b, SDZ15, SM12b, YR12, PSS12]. Parametric [KBG11, PZ10]. Parasite [AAA20]. Pareto [GF17]. Pareto-Front [GF17]. Parikh [BM16b]. Parity [ASH16a, FPS16, Pou13]. Pariwisata [HS24]. Parkinson [SMT17]. Parser [BK10b]. Parsing [KW14, SP15a]. Part [EAMD14, CDC18, SOT18, SGKS12]. Partial [A17, DK11, DMM14b, NAH15, SM11, TYSS15, YT15]. Partially [CP14, SP15a]. Participants [BNA14]. Participatory [KZ16a]. Particle [AS17, AHAS^{+}17, ASS10, AA12c, AM17a, AHBG19, AMAY20, AZ17, Ber11, BM13c, DCH16, DH17, DB12b, FR14, GEB12, GQD^{+}22, wHsYwT12, LGC11, MG14a, MG14b, PCR12, RHM14, SAY16, SMHS10, VR15a, Zyo12, VJR11]. Particles [KK17a, ME15]. Partitioned [PN12, SE16, SEA16]. Partitioning [BA11, CTG10, HJE14, SS14c]. Parts [EAEE14, PRK17, VR14]. Party [NI14]. Pashto [KAZA13]. Pass [HIE14]. Passenger [AAA17a]. passing [SS13f]. Passive [BLL13, FKFU12, TS17]. Password [ARMS16, JP16]. Past [HGK15, SH17, SDJ13, TRM23, HD14]. Patch [AK18b, AI13a, KCD13, KNGT10, LK16b, MT13b, PVR20, PB16, PK12b, RC14, SA16, SSG14, SKJ15, TRDB24]. patched [DDN14]. Patching [AA20]. Patents [DD20]. Path [Abd18, AMN12, AM12b, CS10, DM18, CB12, ES13, rPGM14, GBPTEQ^{+}13, JSS17a, KK20, KRP12, MC17a, MKT17, PSS12, RK17, RS13d, SAY16, SS12c, SAKS24, MNB^{+}22, YMR^{+}12]. Path-based [MC17a]. Path-Loss [CS10]. Path-Planning [MNB^{+}22]. Pathfinder [OHT13]. Pathology [MK13c, NDC10, VKJ^{+}15]. Paths [Kum14b, RA10a, RSS^{+}11]. Patient [AG13, HPC13, Hem22, RV12b, SSA16, XCY11]. Patients [GOCA15, LAAI23]. Patrolling [Ala16, FTLK14]. Pattern [AO13a, AHBH11, BK14, CTG12, CR12, DDM^{+}23, DM17, EmErB13, GM16, GPWS18, HMN18, JS18a, KV14, LKO16, NP13, NJ16, PK15a, Pet22, PG13b, PA23a, SB21, SCD10, SP15d, SPL15, SL15, SB12b, SV12b, Zon17]. Patterns [ABF14, AW20, BRS16, CR15, DDBD22, FR12, Gum19, KPV13, LR12, LHM15, MM18, Mee12, MI13, PR10a, PZN21, SRS12, TKC^{+}13, VR15a, WAmMS14]. Pause [ASNS10]. Payload [AAH21]. Payment [SMK14]. Payments [OK14, AO16]. PC [AZA23, CS11, SGA13]. PC-cluster [CS11]. PC FFIW [AA12e]. PCA [Bar16, BMU14, KRR11, KK16a, MEE15, MBMB12, PG15a, SHD13]. PCA-based [BMU14]. PCA-BPNN [Bar16]. PD [BBMA13, MBKC15, Ngu23, Ngu24a]. PDA [VD10]. PDITT [HN16]. Peak [AB12b, HO15a, KG15c, DL19]. Peak-to-Average [AB12b, HO15a]. PEAP [HR21]. Pedestrian [OOJ17]. Pediatric [GOCA15]. Pediatrics [BIAE24]. Peer [Ali12, Alt13, HK15, MC17a, NM11, RMP11, RD17b, VV18]. Peer-to-Peer [Alt13, HK15, NM11, RMP11, RD17b, VV18, Ali12]. Pen [GJG16, OD13]. Penalty [KA12b]. Pencil [EDD19, TY15a]. Pendulum [CQM15]. Penetration [AZ16, KA21]. People [DOC^{+}19, EAB13, EAEE15, LSA^{+}17, RDD^{+}12]. Pepper [Guo18, KG15a, RAJ^{+}12]. Perceived [KEA19]. Perception [NNR16]. Perceptions [Sab16]. Perceptron [AM19a, HAAT16, LsAAA13]. Perceptrons [FKT12, SP11b]. Perceptual [SS15a, TS12]. Perfect [Olu14, SCE14]. Performance [AOB12, AEE15, AAA14a, AS14b, Ade13a, AO13a, Afr17, AJ10, AA14c, AJSG11, AA12c, Ahm14a, Akg13, ARS13, AD13, ATA22, Al 16, AS12d, AMY23, AASAA15, AH13b, AV13, Ans14, AF15, AG14, AAaAD12, AAOA23, ARZ11, AIK11, L13, MKOS14, BR11a, BADS17, BGS11, BBMA13, Bef18, BKM15, BSP11, Bon15, ÇGDT18, CN10, CS10, DXT23a, DS10, DA10, DB17, DB13a, DGPP15, DS12c, DTO23, DKS14, ESE13, EmdHNmAW13, ES15b, EAa14, EMA11, FEE13, FLLR20, GSA12a, GCP13, GMS12, GS11, GSK13, GG12, GK15b, IS21, IKY22, IRT11, IM15, JAG15a, JA16a, JA16b, JZS13, JC14, JD15, Job17b, JI10b, KA12a, Kar11, KM15c, KV16a, KTBK12, KWS^{+}11, KBEA18, KP13, KH15, KM16d, KDS10, KK10b, Kum14a, KS14c, KC15c, KG15c, MKN12, MLRU12, MN15, Mah19, MFA13, Mal10].

Performance [MS11, Mar12, MS10b, MBKH16, MSP12, MM21, MK18b, MJM16, MSS13, MMHP17, MFM12, MA17, MB13c, MSR16, NP13, NKR14, NAhA13, NS13b, NmBV13, NM14, NSS15, NMK15, NS24, OMH11, OA15, Oua12, PCR12, PT15, PD14, PB23, PPP16a, Poo22, PRT^{+}16, PPR^{+}12, PM13, RS13a, RM12, RSHI12, RFAH14, RBN^{+}16, MP17, RJS^{+}13, RGR15, RHM14, RSK15a, RPV10, RD17a, RMG13, RKGK15, RGK17, RKY12, RML11, RRP16, RJA23, RJC17, SK14a, SDB15a, SJA13, SA16, SRSV14, SOT18, SHR11, SSD^{+}13, SBR18, SS10d, SKS10, SS10e, SG12a, SN14b, SZEM16, SA15c, SQ18, SGD13, SGS13, SM14b, SA15d, SNJ16, SJ14, SD21, SZUSNK19, SA12d, SA12e, SSU10, SW17, SLK^{+}13, SCK17, SA14, THEE11, TAM14b, TOD13, TR14a, TA10, TKR15, UCNC13, UDA13, CCE20, VKB14, VK11, VM12a]. Performance [VjS12, WKC11, WAmMS14, WAG11, XL21, YT15, YB12, Zae13, CS12, EHSA23, NRRM12, RSRT20, TBKA13]. Performances [BME19, BOS15, KBBM15, LB11, MR11]. Performed [PB13]. Performing [CAS14]. Perimeter [Meg10]. Periodic [AE12b, HJ13, TCS19]. Periodicals [EAE19]. Periodogram [ELDM13]. Peripapillary [TRR16]. Peripheral [RRB12]. Perishable [SM11]. Permanent [CPSP13, SHSA12]. Permeability [Ber11, JGZ13]. Permeable [Haz14]. Permissible [SMK14]. Permission [HM18]. Permission-based [HM18]. Permutation [Ela11, wHsYwT12, RGCGCMFJ13]. Peroxide [AMN15b]. Persian [KB14b, RSA11, TT11]. Persistent [IBS16, SG16b]. Person [AN16a, DJG12, NUK16, SS10f]. Personal [Ana15, CPSP13, Dha12, EAEa12, KAAA13, Mal10, RMS14b, YK13]. Personality [NU12]. Personalization [HMMN13]. Personalize [KN15]. Personalized [AIO11, BK11, Cha15, CB12, GS12a, HASH17, KA12g, Ran13, RV13, SA13b]. Persons [PR15]. Perspective [DRJ14, KVK19, OKS14, RKS16, RKS17, VMA23, VV18]. Perspectives [AAI13, MJP15, NOO16]. pertained [AS13a]. Pertinence [BN14]. Perturbation [Eba13, EsGK14, KKI16, SNA15]. Perturbation-based [EsGK14]. Perturbation-Iteration [KKI16]. Perturbations [Eln14]. Pervasive [KNGT10]. PESH [ZEfNF13]. Pesticides [SS16c]. PET [RMD10, SYAS11, VMM14]. Petri [AERS11, GH14, LR12, SKA13, Sin13]. Petrov [VR15b]. PF [MH12b]. PFIMII [DTS16]. pH [SRTB11]. Phalanges [AWA12]. Phase [Alg19, AYM16, ASK15, BFA13, DDDD12, DBW19b, DBW19a, ElZA13, Ena12, GM14a, HJE14, HIE14, wHsYwT12, KC14a, KMG21, KSK14, MKM17, MG17a, Mis14, MSA16, MAI15, Nim15, NMK14, NSMK17, PT13, PMK12, PS16b, PS12b, SM17a, SSK^{+}10, SR13b, SBM^{+}15, TAM13, TM13, AB13b, SS12b]. Phase-Locked [ElZA13, PS16b]. Phases [PMH10]. Phenomenon [KAR12, MR12a, NGA14]. Phishing [Has15, KSKS23, PAK16]. Phone [AzSKM12, GPB22, HPC13, KKK12, You18]. PhoneGap [SPP16]. Phoneme [HW18]. Phonemes [RHK15]. Phones [AWK19, CGKP13, SB15]. Phonetic [She11]. Phosphate [GMRK12]. Photocopy [DRJ14]. Photodetection [BOS15]. Photodiodes [BOS15]. Photonics [RHU12]. Photos [AH17b, RNM14]. Photovoltaic [ANE^{+}13, AB13a, Bou12b, CBhM12, DDC^{+}24, GA15, IAM16, Job17b, KM15a, MAK13, MA16, MCaC14, RURS13, RBDD16, YAT15a]. Photovoltaic- [ANE^{+}13]. PHR [PP13c]. Phrase [GR14]. Physical [AE13, AYYB15, BC22, FM13, HZBG13, KN17b, MJM16, NKD14, SSA^{+}23, NTL18]. Physically [ZGAS^{+}11]. Physicians [AAI13]. Phytochemicals [FMS13]. PI [PCKD17, PKK18, PB23, AAEEA15, AK11, BSP11, Bou12b, CSN15, HR16, HAA16, Par14b, SGB16, TOMGBSS13, HKR17, TVY17]. PIC18F452 [KNV10]. Pick [CM13]. Picture [OJO11, USR17]. Pictured [Moh13b]. PID [AmH13, AS17, ATM12, ANE^{+}13, AZS16, AAA17a, AAA14b, AK11, BM13c, Eja12, KNR15, LGZX17, MBE16, SO14b, SR15a, ST12]. PIECC [GSR21]. Pieces [AA20]. Piecewise [ASR15]. PII [DH21]. Pilot [AJ15]. PIN [BOS15]. Pineview [MFM12]. Pineview-D [MFM12]. Pinglish [TT11]. Pipe [Hus15]. Pipeline [BKS14, DH12]. Pipelining [SFH17]. Piracy [VV18]. Pitch [NRR12]. Pixel [AGRR10, ADVB10, GEB12, HMO17, MNB13a, PRP15, RM16, SCH21, SSA14]. Pixels [SAW16, SR17]. Place [YBM16]. Placement [mH14b, JTKC14, KAS13, KK14, MK13d, MG14a, MBZA16, OO14, SS10d, SD15]. Plagiarism [SBKS16]. Plan [IE12, IA15, Jag15b, NS23]. Planar [AGZ15, BB12a, FM20, KIH15]. Plane [AmAr14, HR13, SN15c, ZEF17]. Planning [AA18c, BDN13, DF15, FCS11, HO15b, HPR17, IK11, KK20, LSP17, MNHB12, RRudS12, SS12c, SM11, SSK11, SAKS24, MNB^{+}22, MDCS22, TB12b]. Plant [AHBAR^{+}11, AAO13, CP12, JA22, KNSS11, LABV21, MM15a, MBP16, PA23b, RKE23, SVRB15, VJR11]. Plants [SMHS10, HDMK14]. Plastic [HO15b]. Plasticity [FYM11]. Plate [ABD22, BI13, Cha12, DDB11, HD14, HGK15, ILG15, MWS20, SH17, SDJ13]. Plates [BB13a, CDC17, EH16, MPJ18]. Platform [ATA14, AHAPT15, BAO16, BLQ21, CSN15, DLR19, HKR17, HKL16, KN12c, KD15, USC12, ZTM11, ZKN10]. Platforms [AHZAH16, AAO22, Tit16, SAS24]. Player [PH23, BMRA12]. Playlist [Dur14]. PLC [CS13, HZLL15, HR14b, Mah16, PCS18, SM17a]. Plum [Zhu21]. Plum-Blossom [Zhu21]. plus [EBO13]. PM [HR16, KM12]. PMD [AJ10]. PMSM [PCW10]. Pneumatic [SPXT17]. PNN [Elc11, Lah11, MRM13]. PNN-based [Elc11]. Poem [DAB^{+}16]. Poetry [AAG13]. Point [ACK12, Aa13c, Alg19, AYAI17b, AYAI17a, GSR21, LAZ17, MAK13, MB15b, MCaC14, SJ12, SS12c, SAK13a, SC13]. PointBug [CS15b]. Pointers [Jay12]. Pointing [KKK12]. Points [ABD22, FML20, Kum14b, SB11a, WBL^{+}23]. Poisson [KRR15, MRH15, RSSV14, SRSV14]. Pol [MKB13]. Polarization [CSP14]. Polarized [RT20]. Pole [MBZA16]. Pole-Placement [MBZA16]. Police [Nor13, RN24, SG13b, WBL^{+}23]. Policies [SK14a]. Policing [Nor13]. Policy [BG17, CD16, EGES15, LS17, PP13c]. Pollination [PS15b]. Poly [Abb12b, VS10]. Polycarbonate [KMS12]. Polygonal [SRK10]. Polyhedron [BS12b]. Polynomial [Ak13, AlS24c, AlS24i, DNP15, OJR11]. Polynomials [JV14, RSC16, KBA10]. Polytechnic [am12a]. PON [ASK17, RK19]. Poon [SSAZ13]. Poor [AAI13, AA14b]. Popularity [KVR14]. Population [AA14d, DDS20, UJ13]. Porcine [BML22]. Porosity [AB13c, JGZ13, ST16]. Porous [BB13a, CP14, CDC17, DJJ13, HIU11, HGK15, HY13, HR13, HJ13, Nim15, TH15]. Port [KA21]. Port-based [KA21]. Portable [AR15, AB12a, GS16b, hKhHL13, MNR14, MNB13c, OD13, RHM15, MZAE16]. Portal [MREd14, MM12b]. Portals [AS19]. Portfolio [KA11]. Ports [RK15]. POS [PY19]. Pose [Rag15, TX20]. Position [Abd14b, AOO20, AASAA15, BYKO17, EAM14, Ham14, HR22, Mah19, MBZA16, NM24]. Position-Based [AASAA15]. positional [NM24]. Positioning [Cai18, DNR15, JLCV16, SV14b]. Positive [BMM13, HSS11]. Possibilistic [SP11a]. Possible [CHG12, Pet22]. Post [MRH15, Zaf11, SG10]. post-processing [SG10]. Postoperative [XCY11]. Postponement [SM11]. Potential [DUP13, Joh14, RÇ16, Ras16, Ras17, SM13a, YSH12]. Potentiality [BC12]. Power [AmH13, AJT17, AKK11, ASS10, AA12b, AA12c, ARS13, AD13, AkAA14, AMAB15, Alg19, Als18, AB12a, AB12b, ACK14, ASD16, AKB16, AC13, BR11a, BK17, BD18, BJPP15, CKT10, DDN14, DDL16, ED18, FHM16, GS11, HO15a, HBS12, HB13, mH14b, Ibr14, II16, Isl18, JBK13, JJN14, Jay15, Joh14, KKT15, Kei14, KMB14, KAAA13, Kub17, KC12, KAS12, KK14, KC15b, KRR13, LGM12a, MLRU12, MMN18, MAK13, MK18a, MTD14, MM17, MKWK15, MEE16, MCaC14, NL15, NHH10, NKM11, OC13, PV17, PMK12, Par14a, Par14b, PS16b, PPR^{+}12, QKRB11, RSP11, RA18, RKE23, RURS13, RJC15, RK19, Rok20, SMT19, Sar13, SK18, SMHS10, SGB16, SC13, SSS13c, SS15c, SAR16b, SBC12, SK15c, SZS15, SN16, SLK^{+}13, SSY11, TBA12, TMS15, TD16b, TAH13, VRS12, VT15, VM12b]. Power [VM14b, VDM22, YA18, PC11a]. powered [EhMA13]. Powering [SB15]. PPE [ONK24]. ppp [AA20]. PPUDP [NP19]. Practical [BBA12, GFMS14, HZLL15, RTD10, TFT11]. Practice [AA14b]. Practices [Amb24, DH21]. Pre [BB12b, DCH16, GSK13, Nta14]. Pre-bunched [BB12b]. Pre-Deployment [Nta14]. Pre-Processing [GSK13, DCH16]. Precise [BB10]. Precision [BKM12, Nah13]. Precursor [KDBT12]. Predict [Alb15, BS11a, CK12, MA13b, NSM14, RR16, STS19]. Predicted [EGES15]. Predicting [AAN^{+}19, AW20, BWH15, Cur20, DTO23, KS16c, KM16c, LAAI23, LSSD20, LHM15, MB13b, NM20, NO18, PFS12, RRudS12, SO14a, SAAM17, SD21, VDS12, NS24]. Prediction [ASD19, Abb12a, AST11, AS14b, AJN18, ANAK17, AA18a, AHH14, AAMG13, ASH13, Ask14, ARR20, BI12, Ber11, BYT14, DL21, DH17, DB17, DKS14, EKES20, EUEI23, EmdHNmAW13, El 13, EEAE17, FA13, GDS21, GT13, GK21, HHHK21, Hem22, IM15, JIF12, Jab14, JBT23, JARM13, JGZ13, KH20, KBTB12, KV16b, KJTM15, KM16d, KHK13, KNV10, Lah11, LHKR16, MM13, MB15a, NMS10, NP12, NMN15, NPA^{+}24, OC20, PAK16, PMD13, Rah12, RM15, RKGK15, RP24, RKS16, SY12, ST16, SCH21, Sha16, SKG14, SK10c, SS15c, SASS11, SK12c, ZEB17, ZRZ11, BAA24]. Prediction-based [OC20]. Predictions [Amb24]. Predictive [AmAr14, HZK12, JBT23, ND11, SASS11, SGJ12, TAM19]. predSucc [AA18a]. predSucc-Site [AA18a]. Preemption [AzSKM12]. Preemptive [BB14a, KKR18]. Preference [MK13b]. Preferred [NKD14]. Premature [WWSS12]. Preoperative [SSK11]. Preparata [JS12d]. Preparation [AAYM13]. Prepared [PS12a]. Preprocessing [ALO17, AR17b, KP15b, RKR15]. Preprocessor [BNMM21]. Presence [BH17a, Es12a, MT13a, MTKP15, PR12, SDZ15, SK15b, SV14a, SB12a]. Presentation [AAR11, JHP12, Rus15, SHR14]. Preservation [CPD13, GPC16, Kau15, PA24, SRS14b]. Preserving [AMBS16, AHK14, Bat18, DBB24, HSS11, JS18a, JT15, KMS19, NP19, PN11, Pou13, VN10]. PRESS [BLQ21]. Pressure [BSM17, BR13, DJ12, EN13, Naj12, PSD13]. Pretreatment [CFB11]. Prevalence [BI12, KAL15]. Prevent [CJK17, LPP12, TGL13]. Preventing [NKG12, SC16]. Prevention [AM17a, AV10, AS15d, FAA13, JSK17, KSS^{+}14b, ZRE16, ZTL^{+}14]. Preventive [KMG21]. Prewitt [NSS17]. Price [BWH15, GDS21, GT13]. Prices [RK13, SAA24]. Pricing [DD14]. Pride [VR14]. Primary [AI13b, DBB12, EF15, MY13, Sha14, WSW^{+}13]. Prime [SF15b]. Primitive [BO21]. Principal [CR12, CT10, GJ15a, HS24, KVN10]. Principle [LAZ17]. Principles [Ira12, JAG15a]. Print [DRJ14, KBK12, RDE^{+}15]. Print-Scan [RDE^{+}15]. Printed [MT13b, OMF^{+}14, PS16a]. Prior [Ibr17]. Priori [BBL15]. Prioritised [FIS16]. Prioritization [AS15c, BM14a, JV12, MAM16, VB13]. Prioritize [BG17]. Prioritized [MKOS14]. Prioritizing [RHS18, RRudS12]. Priority [AZY18a, Ara15, IA15, KS17, KM12, KAR15, PS15c, SSN18]. Prism [Kar11]. Privacy [AMBS16, AHK14, AMK13, AMA14c, Asi15, Bat18, CPD13, DBB24, GS16b, HBS15, JS18a, KMS19, KZ16a, KSlAM12, KN15, NP19, PN11, PA24, SRS14b, Sub20, VN10, ZUDS16]. Privacy-Preserving [DBB24]. Privacy-related [AMK13]. Private [AMNA18, ATAAAF19, MMHP17, SY17]. Proactive [MKN12, Nor13, SAR18, UR16, VjS12]. Probabilistic [AA16b, DDDD12, GJ15a, GT13, JVG15, LC17a, LC17b, PMD13]. Probability [Fen18, KIA15, KK16b, SHHL13, TSRM10, TOHS14]. Probe [RC14]. Proble [TB12c]. Problem [AMS13, AMAHH19, AS14c, AI18, AF15, ANA11, AMTM11, CW23, CM13, DMM14a, DD13, DMM14b, EBAS11, FS11, GN14, GMAR14, GV15, Has16, HBA13, HA17, wHsYwT12, JV14, JSY14, KJKK12, LGC11, LSB17, LRL17, MJM17, MB15b, NR12, OON17, OESKF16, OJ11, PBG15, PMM18, RGCGCMFJ13, SK13a, STG13, SGS15, SDS15, SS13f, TNKS15, WDS13, YMAS17, TA12, FD14]. Problems [AJ13, Abu17, BP15, Cha14, DL13, EAS16, EDR15, ESZED15, GM15c, GAZ13, Ghe16, JI10a, LK16a, MS14a, Mos14, Ori14, PK19, SAK13a, Tah17, TN14b, TAY15, UA14, VGo12, VR12, VR15b, YAT15b, HAA12]. Procedure [BG16, JA16b, SSAZ13, VPGF10]. Procedures [AV13, EE16b]. Process [AAYM13, AZY18a, AO13a, Agg13, AMK20, AGG16, AVS13, LG13b, Car18, DK11, DSDD12, EKI16, FLM11, FYM11, Ibr17, JAB22, KHSE12, KMS12, KSL^{+}18, KS13b, KJ15, LAM14, MM20a, MM20b, MAM16, MB15c, Ngu22, NI15, PF20, Pil13, PA19, REUHTG18, Roc17, RS16, RML11, RGJ11, SRTB11, SFAS19, SP10, VV14, Ver17, GT11]. Processes [CD16, EMF14, KLS13, YJ11]. Processing [AAA16, AB11, DC22, ETKE20, ESR13, ELDM13, FAR16, GEA16, GMS12, GSK13, HK18, HPM20, IEG14, ITK21, ISH12b, JAB22, KK13a, KSC11, KPV12, LSK17, MSD12, Mun14, MVC10, MSA18, PK19, PDG16, RGR15, RBT10, SP16, SK10a, SPBP17, SAS21, SJV13, SV18b, TAP15, TR14a, UI16, WDN^{+}20, YK12, You13, ZAER12, DCH16, SG10]. Processor [BA11, HYA12, NK12, OA15]. Processors [Am13b, BM13a, BSM^{+}12, DL23, Far14b, MFM12, RPV10, SK14a]. Procurement [UAMN17]. Producing [NS23]. Product [DNP15, DM18, HF20, HEEG15, HS18, HHS15, IA15, KS13a, SF15b, Zhu21]. Production [GSA12b, HDMK14, Ngu24b, SM11, TB12b]. Productivity [VR14, YBM16]. Products [CM13, SM11]. Profession [CNP19]. Professionals [Ash19, NU12]. Proffering [SZZR11]. Proficient [RDD^{+}12]. Profile [BWS^{+}12, EK14a, KZ16a, KR10, NK14a, Ran13, Tua17, USR17]. Profiles [ATM18, KDBT12, RHM15]. Profiling [ur13, Cuf14, MK18b, USR17]. Profit [JM18]. Program [FDG17, MKSB22, MA13b, Say22]. Programmable [CHW16, KN12b, SM14a, Tah14, ZYW^{+}18]. Programmatic [DBKN20]. Programme [HN16]. Programmers [ZGAS^{+}11]. Programming [Agg13, AAR14, BP15, COWK14, DD13, DR15b, Hus16, IMM12, Kha20, KJKK12, MGA20, Mar12, MAk12, MMR14c, PBG15, PMM18, Sub24, VGo12]. Programs [ATAAAF19, JFJN20, LK16a, MMHA11, PS14a, ZRE16]. Progress [AIG12]. Progression [WSW^{+}13]. Project [ADH17, ASP17, AI18, BSS12, CS15a, ED19, HB14a, IAQ11, JS12c, REUHTG18]. Project-based [JS12c]. Projected [SP12]. projection [VJ24]. Projective [PPK13]. Projector [KKK12]. Projects [AZY18b, CK12, DR15a, ML13, SM12b, You18, GT11]. Prolong [NM16]. PROMETHEE [MS14a]. Prominence [AD14b]. Prominent [Mes23]. Promotion [HF20]. Prop [RV16]. Propagation [AS12a, ASC15, AE12b, AB13c, EAMD14, LSB12, ME15, MDNA12, SO14a, SSP13a, Sud10]. Propeller [SSRD12]. Properties [BYT14, DM22, EDR15, HK19, KJA15, Moh13b, Moh14, PN10, SV14a, SVB17, TRM23, PS12a]. Property [GPB22, MR13a]. Property-related [GPB22]. Proportion [EAEE15]. Proposal [Mun14, PF20]. Proposed [AH17b, AS19, AER19, BBB16, CM13, ESEMD21, GC12, GK15b, MNYMGG12, NAH15, NF15, Sha18, Tit16, THAP19, UN13, ZEfNF13]. Proposing [AH13b, AMS14, DN12, MMH12]. Proprietary [ARL16].

Propulsion [AMN15b, AMN15a]. Prosody [MBE10]. Prostheses [RRA17]. Protected [HR21]. Protecting [AH17a, IND13, NTY19]. Protection [AH17b, Ali12, AJ19, AS13c, Far11, KN15, KK12b, MM15a, MA12, PA24, RRAFAM19, RAB12, YBA12, YA18]. Protein [NW13, TR14a, BYT14]. Proteins [AA18a]. Proteus [GM20]. Protocol [ASNS10, AKG20, ASD16, ARZ11, BA15a, BGS11, Bat18, BBHL12, ÇGDT18, CPSP13, CLBA^{+}12, DSD15, DSA11, DDL18, DGP14, ETLE13, EG15, GSD14, GCP13, GB14, HR21, HS15d, JC14, JMM21, KRP12, KV16a, KJJ15, KK12b, KPAB14, KVKR16, KC15c, LJ16, LA15, MK13a, Mah19, MAM15, MG17a, MS11, MP10, Meg10, MY13, NV12, Par11, PPU11, PSS11, PJ12, RHD12, Sal18, SSK^{+}10, SPP11, SFVA14, SaDN13, SI13, SK10b, SG12a, SRC12, SM13d, SGS13, SRS14b, SSP15b, SL11, TBKG13, TCS19, TC10, TY15b, TKR15, UR16, VV12, ZTL^{+}14, CS12]. Protocol-Protected [HR21]. Protocols [ASNS10, AMY23, AASAA15, AD14b, AV10, AM10, AE14b, BADS17, DSD15, DGPP15, IAH14, KP13, KA21, KS12b, MKN12, MV12a, MS10b, NBL15, RKA13, RSRT20, SOO15, SSMP13, SJ14, SZUSNK19, WAmMS14, YAE11, ZBI14]. Prototype [AZA23, GPB22, KM13b, Tit16]. Prototyping [TVY17]. Provably [De 20, RAHH16]. Provenance [PP17a]. Provide [GH14, KVKR16, OV13]. Provider [BS16, GM13]. Providers [MMH12]. Providing [NJB16]. Province [IA15, PPR17]. Provisioning [aAO14, MK15a, WT13]. Proximal [AWA12]. Proximities [yGL13]. Proximity [AJ18]. Proxy [Abd17]. Pruned [BK14]. Pruning [GM15b, SY15]. Pseudo [RSA16, SSK^{+}10, VGJ12, VPGF10]. Pseudo-Wireframe [VPGF10]. Pseudorandom [Lah16]. Pseudotime [Ahm20]. PSK [GG12]. PSM [DBB12]. PSO [AmH13, ASN11, AMA15, BR11a, Bou12b, DH19, DRS13, GQD^{+}22, Key12, KAS12, Par14a, Par14b, Pri13, RA12b, SMVK11, SAKS24, ZPAQP11]. PSO-based [ASN11]. PSO-KNN [ZPAQP11]. PSO-PI [Bou12b]. PSONN [ML12]. PSS [LT13]. Psychological [Pra14]. PTL [RJS^{+}13, SSK13a, SLK^{+}13]. PTL/CMOS [RJS^{+}13, SLK^{+}13]. PTS [PT13]. Public [ATAAAF19, Ali16, BR15, BG20, DXT23a, DH21, EUEI23, IBS16, KDT16, MMR14a, PP13c, RMSS13, RBO19, SK15d]. Publication [HHEY19, PN11]. Publishing [NP19]. Pulmonary [ZJR^{+}19]. Pulse [El 20, LSB12, MM17, SSP15a, SR13b, Suj14]. Pulses [MAH15]. PUM [Tha11]. Pump [NA18, TAM13, ANE^{+}13]. Pump-Phase [TAM13]. Pumping [YAT15a]. Punctured [PDK12]. Punjabi [GK10]. Pupil [AMK12]. Pupils [AI13b]. Purchase [OSS14]. Purchasing [AAAA17]. Pure [CWC15]. Pure-Feedback [CWC15]. Purely [WBS16]. Purging [BO17]. Purging/Antirobustness [BO17]. Purpose [MFM12]. Purposes [MMTS13]. PUSC [GDD13]. Pushing [BRS16]. Puzzle [AA20]. PV [ANE^{+}13, HE15, LAA^{+}17, NSMK17, SVRB15, ZAAS18]. PVD [ADVB10, VP10]. PVT [ACK14]. PWL [MMA13]. PWM [VSN13]. Pyramidal [MS14b]. Pyramids [AAB22, KS16b]. Python [AlS23a, AlS23b, AlS24c, AlS24a, AlS24b, AlS24d, AlS24e, AlS24f, AlS24g, AlS24h, AlS24i].

Q [SMSM10]. Q-Fuzzy [SMSM10]. QAM [DM14, SB11c]. QCSK [TJ12]. QEMU [CJK^{+}12]. QEMU-KVM [CJK^{+}12]. QFD [BSS12]. QFT [YMR^{+}12]. QLMW [DVS13]. QMM [AMM21]. QoS [ASNS10, AICN14, ARK13, AJC14, CS16a, CJ10, DSA11, DKS14, MS17, Moh13a, NV12, SDS15, YEE14]. QoS-based [YEE14]. QoS-RCBRP [NV12]. QoSin [SSM12]. QOSTBC [DAaGZ13]. QPSK [SSU10]. QR [GC24, KSBM17, LPP12]. QTS [RB14a]. Quad [MT13b, PN12, RB14a]. Quad-band [MT13b]. Quad-tree [PN12]. Quadcopter [AS17]. Quadratic [Ble18, Bou13, DT14, DVS13, Ibr14, JV14, RSA16, TArM14, Tua17, HSS11]. Quadrotors [KN17a]. Quadtree [Ak13, QT15]. Quadtree-based [QT15]. Quality [AOB12, HAPG17, AHA14, ATA14, AICN14, AsS13, ANAH20, AS19, ASH13, ASB12, AKB16, BNS15, DA18, EsTeEh^{+}13, ENE13, GSA12a, GRC13, GM13, HST19, JJN14, JIA13, Kas17, KSZK15, KMS12, MBA20, MKSB22, MNB13a, MNB13b, MSP12, MK15a, MMK21, MNN10, NP12, PFS12, PDG16, SS15a, SGB16, SU20, SN16, TTA10, ZAE18]. Quantification [PMH10, RA10b]. Quantitative [BMU14, Jov15, MGA20, MB13b]. Quantitatively [HSD13]. Quantization [AR10, AE12c, KSS12, KPG13, KK10b, LBNP12, SP11a, SK12b]. Quantized [AH13a, LKO16]. Quantum [AA14a, Cha14, CQM15, DDDD12, MAM16, Neg13, PMU17, Pri13, TT14, TN14a, Tha11]. Quarter [AMA15]. Quartic [VR12]. Quasi [AE12b, MCB12]. Quasi-Dynamic [MCB12]. Quasi-Periodic [AE12b]. Quaternion [KR14]. Queries [Bas12b, Dix11, ES15a]. Query [AA12d, AB11, BB14b, Bha10, Cha15, CJK17, DPJB12, Eba13, FAR16, GR14, HNW17, Hus12, KS15b, KW14, KPV12, Ojh11, RS12, SP16, SB13a, TY13b, YK12]. Query-Set [Eba13]. Question [CSM21, She11]. Question-Answering [CSM21]. Queue [AS13d, DA18, IMIA11, KS18, KC15a, SSN18, SJV13, MAM16]. Queueing [ASG10]. Queuing [EER14, IMM12, KMMM12a, KMMM12b, Sri12]. Quick [AB16a, Nta14]. Quintic [VR15b]. Quranic [KBEA17].

R [ATA19, OHT13, Hos12]. R.H.S. [STG12]. RACH [JA16b]. Racing [LM20, PMD13]. Radar [SR13b]. Radial [DHT13, HO15b, KK14, LRRR11, MK18a, MEE16, TN14a]. Radiating [WBS16]. Radiation [ANAK17, DGJ12, Haz14, SK15b, TRDB24]. Radio [AH12a, Abu14, AVRN12, Ade11, CK13, DT14, DCM15, ESE13, EF15, FAZ13, GKM^{+}18, GHD17, HZBG13, JBK13, KAA13, KIA15, KT15, LK15, MTD14, MFE13, MA17, PPB15, RJC17, SIAI14, SG14, SSSS12, STD14, SRS14a, SZS15, TBKG13]. Radix [Isu16, PS15c]. Radon [HH13]. Rail [SSK^{+}10]. Railway [GDS21, GKK13]. Rainfall [AE17, MM13, RM15]. Ramped [HD14]. Random [AR17b, BNMM21, DS12b, EsGK14, HG11, HS15b, IMIA11, JAG15a, JPJ18, KRSB12, KNS20, MKN12, MSA16, NSS17, OMH11, Pou13, NS24, ZEF17]. Random-Walk [BNMM21]. Randomization [TC16]. Randomized [AAEK18, GM15a, Sal17]. Randomizers [MM12a]. Randomly [SS15a].

Randomness [DXT23a]. Range [BSM17, BI19, ES15a, PS15c]. Ranger [BA23]. Rank [KBM22, SHR14]. Ranking [Far17, JK14a, KP10b, Ran13, RV13, RL14, SU14, Ven14a, Ver17]. Rapid [GYP^{+}15, Rab14]. Raspberry [HKR17, PKK18, PB23, PCKD17, TVY17]. Raspberry-Pi [TVY17]. RASSTD [PG13a]. Rat [TM13]. Rate [AU23, AKS16, BSG15, BS11a, BB12b, CKD12, EM23, EE16b, IS21, Jah18, KS18, KSS13, KSL^{+}18, KG11, KR11b, LC17a, Mal10, ND17, NS14, PR12, PB14, PP16, RA10b, Sak15, STS19, SA15d, Uzo13, YK13]. Rates [KNS20]. Rating [DP15a, KRM12, KB15]. Ratio [AB12b, KAR10]. Rational [EEB16, HSS11]. Raw [MMK21]. Ray [CKS13, MB11a, SSK11]. Rayleigh [ASK17, BR11b, KA12a, MN15, MA17, RRR13, RMG13, SS10e]. RBF [VD15]. RBFN [DHT13]. RBFNN [PG15a]. RC [KRMB10]. RC4 [ABA^{+}10]. RC6 [Baz18]. RCBRP [NV12]. RCGA [LRRR11, SD15]. RCPLAG [SN14a]. RDF [AOK18]. RDF-based [AOK18]. RDUP3 [USR17]. Re [ASK17, Elc11, PN11, SMJ12]. Re-authentication [SMJ12]. Re-modulation [ASK17]. Re-publication [PN11]. Re-weighting [Elc11]. Reach [CSN15, SM19]. Reachability [PP12b]. Reachable [YK12]. Reaction [AEA17, AE12b, HD14, PR12, PD14, SKM16]. Reactions [SK15b]. Reactive [BKK13, CS12, HB13, KRP12, KS12b, MKN12, VM12b]. Reactivity [PNIS16]. Reactor [DJ12]. Read [RGR15]. Reading [Far12, IHD12]. Ready [MJM17, SJV13]. Real [AK15, AG13, CM13, CKT10, DRFF15, DGPP15, EhMA13, EMA11, EHHM17, FEE13, GK13, HYA12, IM16, JD15, JND15, JSS^{+}17b, KS15a, KSS18, Ker13, Key12, KDC12, KD22, LB11, Mal13b, MK18b, MBZA16, MNR14, NVV16, NS13b, PC15b, RP14, SPP13, SCH21, SPBP17, SA12c, SSA16, SP18c, SJ14, SGD^{+}20, THEE11, TBA12, TG18, VM12b, VD15, YRMB15, YAE11, BAA24, MN24]. Real-Time [HYA12, Ker13, Key12, RP14, SCH21, SA12c, YRMB15, EhMA13, JSS^{+}17b, LB11, SPP13, SGD^{+}20, THEE11, BAA24, MN24]. Real-World [VD15]. Realistic [RMR14, SWP19, TAM14b]. Reality [Ari23, CHG12]. Realization [KM15b]. Realizing [GCS14]. Really [ABYB22]. Reanalyzing [OAE16]. Reasoning [KW16, KS13a, MH12a, MAM12, MI13, SK10a, SE11]. Reasoning-based [MH12a]. Rebuild [Sun19]. RecB [SL15]. Receding [AYAI17b]. Received [RKB19, SM13c]. Receiver [LGM12a, LGM12b, RT20, SS17c]. Receivers [ARZ15]. Receiving [MC13]. Recency [CTG12]. Recently [JA12]. Recipients [HS24]. Recitation [KBEA17]. Recognition [AH12b, AZM13, AS12a, AO13a, AR17b, AJI15, AA16b, AHM17, APS18, ALA12, AAA19, MS16, BSAb12, Bar16, BP16, BRD15, BMU14, BA12b, BG13, Bha10, BI19, BV21, CP12, CCJ14, CT14, CT10, DI10, DP12, DDB11, DP15b, DM10, DT22, DB10, DHT13, Dha12, Dum20, EEL11, EREM11, FAt13, Far12, FSNR18, GEAE18, GB11, Gho15, GP15, GK13, GGK14, GSK13, GK15c, HaB13b, Hik13, HH13, Isl18, JA11, JAsA13, JH18, JKTM15, KS15a, KA12d, KDBT12, KR17, KBA10, KVN10, KA15, KBEA18, KDC12, KBEA17, KB14b, KK10b, KBK12, Kum12, KPV13, KK13b, MWS20, MSM17, MSM15, ME14, MKAR22, MEE15, MS14b, MRM13, MBP16, MGL22, MPJ18, NKH12, NS13a, NBKM13, OAE16, OMF^{+}14, PG15a, PPP16b, PC15a, PC11b, Pet22, PM15, PG13c, RSHI12, Raj19]. Recognition [RSA11, RDD^{+}12, SMN^{+}22, Sar14, SG13b, SAI16, SP18c, SSJ12, TSA12, TD18, TD21, US12, YYS^{+}23b]. Recognizer [Far14b]. Recognizing [A17, AA17a]. Recommendation [CW23, DBM14, GS12a, SZOK24, SSC15a, TE16, Ven14a]. Recommender [AS15c, KVR14, OEA18, SP15b]. Recommenders [SN15b]. Reconfigurable [MA22, MC13, MDP12, PVR20, SE12a, SN14a]. Reconstruct [PR10a]. Reconstruction [AR15, El 20, GGX12, Has14, ITK21, Kau15, NMK15, OSA14, SCE14, SD16, SPC13]. Record [DRS13, KMS19, SRA15]. Recorded [EAB13, TS11]. Recording [AYE13]. Recordings [HSH17]. Records [ASA14, BB14b, ZEE15]. Recovering [CPG15]. Recovery [AHH14, KP10a, KJKK12, NTL18]. Recruitment [Dak23, Ver17]. Rectangular [DJA13, ES15a, LK16b, MBR16, PG15a, SJK21, SM14b, SKJ15]. Rectifier [MA15, MH12b, Nah13]. Rectifiers [BKM12]. Recurrence [MAaM14, MA13b, RR16]. Recurrent [SAA24]. Recursive [DK13, MP12a, Ode14]. Recursive-based [Ode14]. RED [BBMA13, IMIA11, OMH11]. RED-AQM [OMH11]. RED-PD [BBMA13]. Redesigned [BM13a]. Redeye [RNM14]. Redirection [HWT^{+}13]. Reduce [Afr17, DAB12, GM14a, GM15a, Jah18, KPC16, KVKR16, MC16, RR16, STG12, SP15d, GBM10]. Reduced [CM15b, Eja12, FSA12, MMA13]. Reduced-Complexity [MMA13]. Reducer [GM15c]. Reducing [FKT12, GM15d, KY15, MBM14, NS14, RKR15, RAR24, SR15a]. Reduct [KK10a]. Reduction [ADZ11, AMNA18, AkAA14, AsRA^{+}12, AC13, BM13a, BMM13, BK17, CR12, EN13, GC18, GHH15, HO15a, HJG^{+}13, IEES16, IK11, KK10a, KNT11, KG15c, LJ16, MN17, Moh13a, MAH15, PT13, Sam13, SMHS10, SGB16, ST11, SKS16, SAR16b, SN10b, UJ14, ZYM13]. Redundancy [HAS17]. Redundant [AAO13, PK19, SI13, YB12]. Reed [Far11, Ish14, KG11]. Reengineering [BB13b, RT11, RS16]. Reference [AH17a, BM13c, BSP11, DA10, PKP13, SDZ15, SMR16]. References [BBD13]. Referencing [BN14]. Referentially [NBF14]. Referentially-compressed [NBF14]. Refinement [AKA^{+}18, KN15, LNM24, NTL18]. Refinery [TSS12]. Reflections [Ana15]. Reflector [MS10a]. Reflex [KV14]. Reforming [TSS12]. Refractivity [Ade11]. Refresh [AS24]. Regarding [Ash19]. Regards [BA12c]. Regeneration [KNS17]. Regime [YR12]. Regime-TCAD [YR12]. Region [AKSEB15, CJK16, EGES15, HKS12, IKY22, MM16, RSK15b, SKS10, TVGT18, TM13]. Region-based [IKY22]. Regional [IA15, MAMA16, SM23]. Regions [ES15a, TR14a]. Register [DL23, SK10c]. Registrar [SNA13]. Registration [Aa13c, MV12a, RMD10, SSS15a, SNA13]. Regressio [NYS15]. Regression [ASR15, AlS24c, AlS24e, AlS24f, AlS24h, AlS24i, Ask14, BNMM21, CK12, FA19, HAH15, KV16b, KJTM15, LSSD20, RHS18, SO13, SBR18, WOBK18, NS24]. Regular [BP11, KIH15]. Regularization [PMP12]. Regularized [FSA12, SNB18]. Regulate [Wu21]. Regulation [VS12]. Regulator [TArM14]. Regulatory [AM19a]. Reinforced [IEME12]. Reinforcement [Far17, ONK24, SJV^{+}15]. Rejuvenation [RT11]. Related [ZUDS16, AMK13, GPB22]. Relational [BD13, BLQ21, BBL15, CV13, DSM^{+}20, Fad18, II16, KVS16, Pet23, SB13a, Sel22]. RelationalJSON [BB14b]. Relations [DM22, JSSA20, MAaM14, RP24]. Relationship [AAO11, BI19, GSV20, HS18, KB16a, RHT19, RMH20, Sak15]. Relative [AIC19, DLN12, Ham14, II16, KK10a, USR17]. Relay [AM15, AJ19, GDS16, JTKC14, MK17, RMG13, RR12, SZS15]. Relaying [BADS17, ESE13]. Relays [KT15]. Release [ADS16, FCS11]. Relevance [FPA13, PS13, SSG13]. Reliability [AM11, AA18b, ADS16, KA16a, MR12a, MAA13, MC17b, MM14b, MHE13, PG15b, RRR13, RM11, RKS16, RKS17, MS12, SH15, SPSR12, TB15, YA14, BAA24]. Reliable [Fla18, JTKC14, Kub17, LHKR16, NV12, OAA14, SPP11, TBR15]. Reliance [SA10]. Relieved [RS13d]. Reluctance [AHK12]. Remainder [STM13]. Remaining [KS14b]. Remedial [JBP16]. Remedy [MMPH15]. Remote [AMJ16, BS17, BNA14, DLR19, JVG15, KSSM24, KS14c, MOM17, MSS14, PP16, SDG18, SWH16, UO23]. Remotely [KHAE19, MJ14]. Removal [AHM14b, EEEE16, Guo18, HSH17, KG15a, KA13a, MSN12, ML12, SV18a, SSR14, SF15b, SG10, ST14, SR15b, UC15, YRV14, AVH10]. Removing [Gum19, RAJ^{+}12]. Rename [DL23]. Rendering [sMTEkEdAr12, Par11]. Rendezvous [DBS15]. Renewable [EhMA13, MK18a, OAS19]. Repair [Es12a, KM12, KK22, MAA13, MHE15, MRH15, MG12]. Repairable [MRH15]. Repeated [HO15a]. Repeating [SB21, SB21]. REPET [SB21]. Repetitive [KAL15]. Replacement [EGES15, JPJ18, KM12, SK14a, SSK11]. Replicated [BH14, GM14a]. Replication [MY13, SBR18, TBR15]. Report [LC17b, SU20]. Reporting [BYKO17, FK15, SO14a]. Reports [SN15c]. Repository [FNH13]. Represent [PR10a]. Representation [Bha17, MRTB16, ESR13, FML20, Hew11, VNM15, KSS13, KSL^{+}18, PMU17, ST16, SK14b, ZK12]. Representations [PK19, TS12]. Reproductive [BML22]. Reprogramming [EE16b]. Repulsion [SP11a]. Reputation [HK15]. Request [BA11, HWT^{+}13]. Request-Partitioning-based [BA11]. Requirement [JS12c, REUHTG18, SS13c]. Requirements [AS15c, BS18, KA11, KS13c, RKS17, VB13, GT11]. Rescue [HHHK21]. Research [AMBS16, BKK15, EREM11, MNN10, OAG17, XCY11, HLZ^{+}24]. Resemblance [SPP16]. Reservation [DP17, DA18, SNA13]. Reserved [KC15a]. Reservoir [AA13a, AJM^{+}15, Ber11]. Reservoirs [JGZ13]. Reset [LC17a, SSK^{+}10]. Residual [AR15, AM19b]. Residue [RSA16]. Resilience [JKS^{+}12, VMA23]. Resilient [ASK17, OAG17, RDE^{+}15]. Resistance [PP15]. Resistant [AGRR10]. Resolution [AS15b, BIE^{+}12, DT22, DL13, GH14, GGX12, MESNH17, KH12]. Resolve [Jay12, OJ11]. Resolving [AA18a, MD14a]. Resonance [AS14a, ASH15, DM10, KS14a, NJ16]. Resonant [KTS17]. Resource [ADT17, AK15, aAO14, BB14a, DP17, KAA13, LSP17, MDP12, PK12a, Qte14, SS14c, SZZR11, TR12]. Resources [DH12, Dix11, ENE13, HEN13, KFEN21, MM11b, ASNS10]. Resourcesin [MK13d]. Respect [CS14, RKB19, SWP19, Mar12, RGK17]. Respiration [AKSEB15]. Respiratory [BML22]. Response [AMM21, El 13, FJM16, KG15c, LJ16, Wu23b]. Responses [KB15]. Responsive [BH16]. Rest [GJM16, RA23]. Restaurant [UI16]. RESTfull [PA23a]. Restoration [ASN11, ARHERZ10, FSA12, HO15b, MESNH17, RDE^{+}15, ZC24]. Restricted [KAAM11]. Result [AC18, AS12c, CB16, GGI14, NC14, Rus15]. Resulting [A18]. Results [ACK12, ES15b, IGS17, JG12, Kam17, Ran13, RV13, RL14, TS10]. Resume [Upa23]. Resupplying [TB12a]. Retention [RHT19]. Retinal [ASB15, BS12a, DSDD12, KPR14, KS16e, KK15c, SMR16, TRR16, WHP14]. Retinopathy [BS12a, DSDD12, Hab13a, PA12]. Retrial [ASG10, KC15a]. Retrieval [AB14a, AAA11, ASB15, ASS15b, ASI17, AGH14, BAM13, CJK16, Eis14, Elg13, FPA13, HR14a, JA12, KTBK12, KR14, KB12b, LKO16, MGB11, PWG15, RMS12, RKR11, SSG13, S24, SEF17, SLU14, SH12, TS12, TMK^{+}10, TD16a, VK13, VJD22, Yac15, YT15]. Retrieved [JA12]. Retrieving [GP12]. RETT [TC10]. RETT-Gen [TC10]. Returns [GK21]. Reusability [SC12]. reveal [FMS13]. REVEDERE [NI15]. Revenue [TUS20]. Reversible [AMHAh13, GHBN12, GSN12, GSN14, MKWK15, Pou13, RB14a, Rah14, SPK16, SR14, Tah14]. Review [Alt13, ASS15a, ABL17, Bey14, Cuf14, DS15, GB15, GJ13b, Hem12, JG11, Joh14, NA18, NGA14, NI15, NMAB14, PP13d, PMBP20, PKJ15, RS16, RPS12, SSC15b, FIA11]. Revised [Bad13a, Rad13]. Revocation [Abd13]. Reward [CD16]. Rewriting [BKB11]. RF [CM15b, MDGS13, MDNA12, RR18, SDG18, TAM13]. RFB [PC11a]. RFID [GKK13, KAA20, KR16, NMSA12, NAD19, SS13b, SYP10, TV10]. RGB [AAH21, BS12a]. Rhodopsin [NW13]. Rhythm [DAB^{+}16]. Rice [JA22, MTM12]. Rich [Elm22]. Rician [BR11b, EE13b]. Ridge [AA11b, BCO13, CWT15, FA19, SO13]. RIF [KS16d]. Rigid [Aa13c]. Ring [BSZ13, Kei14, SSP15b, TS16]. Rings [SSV11]. Ripples [SV12a]. RISC [BM13a, DV14]. Risk [AAN^{+}19, DP11, DA12b, GFMS14, JARM13, KRH13, LS13, MJP18, RR22, WSW^{+}13]. Risks [KA11]. Risky [HHS14]. Rivers [BI12]. Rivest [IPR15]. RLS [SM13c]. RLS-Received [SM13c]. RNA [LBB11]. RNS [Isu16, MB17]. RO4003 [KNGT10]. Road [AEKA11, MB11a, MBL15b, SM15a, TBKA13, VSS16]. Roadmap [AO16, KJ15]. Roads [RN24]. Robin [AAS14, AMK20, AAY17, EEKG17, MAM16, Neg13]. Robot [Abd18, Ala16, AAA14b, CS15b, EHHM17, Gha16, HZ15, KNI14, MBE16, NH13, Ngu18, Nur11, Rag15, SS12c, SBHM13, SS12d, SAKS24, MNB^{+}22, SM19, TSA12, TG18, Ver16, YSH12]. Robotic [LK16a, MBKC15, Ngu24a, PKDS17, SWH16]. Robotics [LBNP12, SM14a]. Robots [GBPTEQ^{+}13, HM13, TX20]. Robust [ACB17, Ade13b, AA12e, AS13c, BMM12, BMM13, CM15b, DB10, Dha12, DNR15, DBB24, EUT17, Eja12, EAMAES16, Ela11, Eln14, EAF^{+}16, GSCT12, HMO17, Ira12, JS18a, JPJ18, JND15, KK12a, KN17a, KH12, KAAA13, KEV14, KNR15, KK17b, MK18a, NP11, OJO11, PG13a, PVL11, PB12, RB14a, VGS11, S13, ST12, SS15a, SYR17, SGB14, SP10, TL16, TY15a, TRSK15, TC16, US12, YBA12]. Robustness [BMU14, GG11]. ROF [Kum14a]. Rogowski [SVB13]. ROI [BV21]. Role [ANM17, CSP14, rPGM14, HAA12, HAb14, KRR11, KN12c, NS14, SK10c, Sou15]. Roll [NVR12]. Romantic [NKD14]. Room [AIC19, AM17b, HJT14, PW23]. Root [ABF14, SLK^{+}13, USK17]. Root-set [ABF14]. ‘ROR’ [ST14]. Rossler [HS15a]. Rotary [KS13b, SYR17]. Rotated [GG12]. Rotating [BB13a, CP14, CDC17, DSJ11a, GDJ12, WBS16]. Rotation [AHBH11, GDJ12, PB12, Rag15]. Rotors [DC22]. Rough [AS18a, DB17, Eis13, MMPH15, NMS10, NYU14, OESKF16, PRP15, Pil13, fSG13, ST11, VG12, XCY11]. Rough-Fuzzy [PRP15]. Rough-Neural [fSG13]. Round [AAS14, AMK20, AAY17, EEKG17, MAM16, Neg13, SRF12]. Round-Robin [MAM16]. Route [CW23, CC17, KP10a, KVKR16, LS17, PS12a]. Routed [SA10]. Router [AOB12, AB20, ABBG14, HWT^{+}13]. Router-based [HWT^{+}13]. Routing [AAN12, Abu17, AMNA18, ASNS10, AMY23, AASAA15, AKG20, AD14b, AV10, ASD16, AE14b, ARZ11, BA15a, BT14a, BSG15, CM13, DF15, DGPP15, DSS13, DGP14, EEH14, FS11, FK12, GJ15b, GCP13, GKG16, Has16, HS15d, JC14, JM15, KRP12, KS15c, KA12e, KS12b, KK12b, KC15c, LA15, LS17, MKN12, Mah19, MAM15, MFA13, MP10, MS10b, Meg10, NS15a, NV12, NBL15, PPU11, PSS11, RKA13, RA10a, RSRT20, SE12b, SE16, Sal18, SN15b, SK13a, SG12a, SSMP13, SM13d, SGS13, SJ14, SZUSNK19, TBKG13, TBR15, TC10, TKR15, VR11, WAmMS14, YK13, YAE11, ZCDR13, CS12]. Rover [BA23]. Rover/Ranger [BA23]. Row [Pet22]. RPOA [aAO14]. RPOAWLB [aAO14]. RRH [GHD17]. RRNS [AD18]. RSA [IPR15, AHAT16, AA17b, MSR16]. RSFT [SE10]. RSMA [Ode14]. RSMA-FLFC [Ode14]. RSS [ABYB22, RKB19]. RSSI [JLCV16]. RSVP [MSP12]. RT [DDN14, KNGT10]. RT-patched [DDN14]. RTRL [TNC10]. Rule [AMBS16, ur13, CT14, DSB12, Ela12, Hew11, KS13a, Man10, MMR14c, NYU14, OP18, PAK16, PP12a, PP10, PRT^{+}16, PA23b, RP24, SN15a, STG12, SZS15, TK15, TK^{+}17, TM13, VMM14]. Rule- [CT14]. Rule-based [Hew11]. Rules [AmAdS12, ASN12, Ba13, DBM14, KS12a, LHM15, Mah13, MM20a, SKC15, VG12]. Rummy [SAP16]. Runge [HII15]. Rural [AP16, BDN13]. Rust [AAA21].

S [KM12, MSS13, AM17c, EJC16, FLM11, MMG16, NK14b, PA17, Wu23b].

S-Band [PA17]. S-domain [AM17c]. S-OrBAC [MMG16]. S/W [KM12]. SaaS [SS14e]. SAD [TH14]. Safe [GMRK12, MAARIH16, RN24, Sin13]. Safeguarding [DE13]. Safety [BS18, HBS15, OOJ17]. Sahasralingam [MB16]. Salatiga [PMP14]. Sales [Bej16, DMM14b, HEEG15, HS18, STS19]. Salesman [AS14c, ANA11, CW23, OJ11, SGS15]. Salesperson [OON17]. Saliency [AN14]. Salt [Guo18, KG15a, RAJ^{+}12]. Salt-and-Pepper [Guo18, RAJ^{+}12]. SAM [BIE^{+}12]. Sample [BYKO17, SS17c, ZEF17]. Samples [KNA12]. Sampling [KAR10, OP18, PPP11, RM16]. SAN [UDA13]. Sanchez [JSSA20]. SAR [SABA12]. SAT [SCD10]. Satellite [Ade13a, Ade13b, BAK14, BS17, BL15, CSP14, CSN15, EEEE16, FAK14, GGK14, GJM16, HZ14, IEG14, IEL18, JSK10, KE16, KSSM24, LAA^{+}17, NB11, PPP22, RACI24, SFA14, SZ15, SDZ15, ST12, STE14, VjS12]. Satellites [BIE^{+}12]. Satisfaction [KEA19, PUW13, Qte14]. Satisfying [DM22]. Saturated [WBS16]. Saturation [RG21, WNE21]. Saudi [AS19, Alt11]. Saving [DBKN20, GS11, HITA20, MTD14, RR18, KTN12]. Savings [KK14, RMP17]. SAX [SP15a]. Sc-FDMA [RM12, PG15b]. SCADA [AKH15, Mug24, PTPS19, PA19]. Scala [SE17]. Scalability [AS15a, Ans14, BME19, CKA18, EEH14, Par11]. Scalable [BA15a, KN11, SI13, ZMB14]. Scalar [HAA16, Sud10]. Scalarization [AJ18]. Scale [AM19a, BA12a, BAIs^{+}14, ÇGDT18, GS14, Hus12, JTKC14, KK13a, MGB11, MG14c, NOO16, PGPM23, PM13, RP14, RHK15, SE17, SL15, SKZ22, SVB23, TB13, WAG11, BAM13, PS14a]. Scaled [Raf15a, Raf15b]. Scaling [GM16, II16, MB17, NSS15]. Scan [AHBG19, Jay15, Kha13, RRB12, RDE^{+}15, SSC15b, Zaf11, SN10b]. Scanner [KTS10]. Scanning [Sar14]. Scans [RVKB12, RVGP20, Zaf11]. SCARA [Ngu21]. Scattering [MDGS13, TS17, TIT14]. Scenario [AAAA21, CN10, KN17b, Kum14a, NCL11, RRudS12, SWP19, SJ14]. Scenarios [LCP21, RD17a, TAM14b, VM14a]. SCGNN [AJ15]. Schauder [AA21]. Schedulability [Ker13]. Schedule [AS13d, GKG16, GSA12b, SV10]. Scheduled [KA17]. Scheduler [JMM21, KDTD11, LMZ12, Pat14]. Scheduling [AOB12, AJ13, AAS14, AMK20, AI18, AA14d, AS11, AAZN21, ALH^{+}24, ASG11, AF15, AAC13, AMTM11, AZ17, AAY17, BMF11, BB14a, CJ10, CS11, DBB12, DH12, EER14, ENEBES15, EMF14, EHI17, GN14, GSA11, GF17, HB14a, Hem12, wHsYwT12, Hus16, IGS17, JA16a, JPDB14, KKR18, KRSB12, KM16b, KMS15, KC11, KAR15, LK16a, Mal13a, MT13a, MJM17, MCB12, MEH14, NNR16, OOE18, RS13b, RB14b, RGS15, RKG10, RGCGCMFJ13, SRF12, Sin14, Sri12, SV10, TB12a, Tha13, TB12c, TR14b, TNKS15, TB12b]. Schema [Asi15, SEA16, Win14]. Scheme [AM15, Abd18, Abd17, Abu17, Ade13b, AS12b, Ali12, ASK17, AS13c, MKOS14, BC22, CSP14, DD14, DCM15, DSS13, DA18, ESKM20, EsGK14, ETG17, FIS16, GEAE15, GEA16, GPS14, HBS12, HMB13, HITA20, IEG14, JBK13, JS14, JM16, KFEN21, KBM22, KBAQ15, KJKK12, KA13b, LC17a, LC17b, Mah16, MK12, MKA17, MA12, Mis14, MH11, MNYMGG12, MZ18, NMSA12, NNR16, NTL18, OAW12, Oud13, PG13a, PBG12, RPN14, VGS11, S24, SS14b, Sal17, SRF12, SSS13b, STD14, SF15b, SDS15, TOO16, UCNC13, VT15, ZEfNF13, ZAHH16, ZAA13, GHD17, KM13a]. Schemes [ATA22, AMK13, AAC13, BR11b, JTS15, Liu24, MS17, MIAEFS15, PG15b, PAH17, TMP11, CKD12]. Scholarship [HS24]. School [Add12, AI13b, LSSD20, Sha14, Ana15]. Schools [CJC19]. SCIDR [BA15a]. Science [Amb24, Sta22, Zhu21]. Scientific [Al 13, CHG12, KKR18, SV18b]. Scintillation [SYAS11]. SCIT [GB18]. SCLEA [AA12f]. Scopes [Al 13]. Scorbot [BA12b]. Scorbot-ER [BA12b]. Score [AEKE15, DP11, HK10, MS12]. Scorecard [EmdHNmAW13, Kha11, RRudS12]. Scorecards [ES15b, KAFS15]. Scores [MAt12]. Scoring [Bar14, CS15a, EEA19, Gho11, GF14, IPB14, MK13b, SSA11]. Scout [BA23]. Scrambled [SA15b, SS15a]. Screen [HS15c, SCV13]. Screening [Agg13, LPP12, SSC15b, Upa23]. Script [KK13b, SS12g, YM12]. Scripting [RAC12]. Scrubbing [HAS17]. Scrum [MR13b]. Scrutinizing [ASS15b]. SDLC [Ash19]. SDN [ADT17, AW20]. SDSS [PSG12]. Seafarer [CNP19]. Search [AAP13, AS17, AGN14, AMS13, AS12c, Ali16, Aly13, ATM18, AZY18b, Bha10, BHK13, CS14, Cha15, DMM14a, DS19, Far14a, FM20, FLLR20, Gha16, HHHK21, JG12, JTA19, KS17, KN15, KN11, KAAM11, KS13c, MRP15, PDK14, PS15c, Rah13, RB11, Ran13, RV13, RS13d, RPS12, SS12a, SB13a, SS10b, SE15, SAKS24, SA13b, SV14b, TS10, TB12c, TY13b, Ven14a, AK14]. Search-based [KS13c]. Searchable [S24]. searched [FMS13]. Searching [MI13, PS15c, PSG12, RS12, RL14, SY17, SU20, VJD22]. Seasonality [SVRB15]. Seat [AAA17a]. Seawall [CCG19]. Secant [LSB12]. Second [LEO15, DVS13, Kub17, KNT11, MSS13, SCV13, SPW15]. Second-Level [LEO15]. Secondary [MTD14, SS17c]. Secrecy [AV15, Olu14]. Secret [AS12d, BNA14, JM16, KK15b, PA14, SS12f]. Section [SRS14a, YWT12]. Sector [DDN14, SVB23, ZUDS16]. Secure [AM10, AB20, De 20, Eba13, GSR21, GJ15b, GRP13, GGI14, HMB13, JS14, JPJ18, JM16, KAAA13, KD22, KSS^{+}14b, MC17a, MR15, NmEM13, RB14a, RAHH16, S24, SIAI14, US14, YBA12, ZAHH16, MNYMGG12, Ish12a]. Secured [ARMS16, ABB20, AAMA20, AAOA23, BR15, EUT17, Hir12, KN12c, SS15a, SK16, SGB14, SS23b]. Securing [AS24, BH17b, BC22, DH21, ESE16, EOO22, aKNP^{+}15, NM11, THAP19]. Security [ARP17, AHH13, AKH15, AH17a, ASASHF17, ATA22, AQ15, AR17a, AJ16, AS14d, AT17, AAS24, AHAT16, Asi15, LEO15, MA22, BRD12, Bey14, CS15c, DLR19, ESKM20, El 14, FIA11, GJUL23, GFMS14, GCS14, GGM23, HMB13, HR21, IBS16, KC20, KV15, KM16b, KA12e, KSlAM12, KRH13, KC12, KR11a, KT12b, KK13c, KS15e, KVKR16, MO17, MOM17, MKAR22, MMK16, NHS11, NKH12, NTY19, Nas11, PK15a, PP13c, PMBP20, RJSJ16, RHD12, RS11, RAB12, RKBB14, RRP16, Sal18, SN15c, SKG17, Sub20, Sul17, TV10, USK17, VS10, Ven14b, WDAJ15, YRA23, ZUDS16, ZEfNF13, ZHUR16, KAS12]. Sedimentation [AJM^{+}15]. Seebeck [Joh14]. Seeds [CM15a, HST19]. Seeker [BS16]. SEER [RR16]. Segment [AH13c, BK14]. Segmentation [AD24, AAH21, Anh15a, AKS^{+}13, AP15, AMK12, AMH11, BB12a, BL15, BS14b, BT14b, CA11, MRTB16, DB13a, FTA18, GK15b, IKY22, JZS13, KSRS12, MPR14, KS14a, KKBS12, KS15d, Kau15, KPG13, KPR14, KY15, KS16e, KKSS12, KS14c, KA12g, KK15c, MM16, MNB13a, MNB13b, NRP12, Neb13, PB15, PMH10, RB14a, RRB12, RPR12, SAW16, SABA12, SNB18, SDB15b, SGT^{+}11, SS12g, SRS12, SR17, SCK17, SMR16, TRR16, TV13, TVGT18, TM18, UR12, Vya14, WC15, ZvAS12, ZEF17, ZD15, AVH10]. Segmented [SB13b]. Segmenting [CD13, SVMLb13]. Segregated [BG13]. SEIR [HII15, LRL17]. SEIR-Epidemic [LRL17]. Seismic [JGZ13, LsAAA13]. Seizure [NMN15, SMN^{+}22]. Seizures [ZEB17]. Select [SLK^{+}13]. Selected [MK19, PRT^{+}16, RSB12, RFAH14]. Selecting [GK15a, OAR16, MMK11]. Selection [AM15, APU16, Abu14, AVRN12, ASC15, AlS24d, ARK13, AAY15, AO13b, AHAPT15, AAS13, BAK14, BSS12, BA23, BMRA12, DRS13, DS13, ETG17, eFKS13, GEA16, GSD14, Gho15, GPS14, GRL20, GG11, HM18, IR16, JFJN20, KK10a, KS14b, KJ12b, KA12f, KJKK12, KGK13, KB16c, KS13c, LCA12, LW16, LAM14, MM14a, MK13b, MAE16a, MAI12, MR12b, MG17b, MTM12, ML13, NmBV13, PDK14, RR11, RAA14, RMG13, ST16, SSA11, Sin16, TBN^{+}15, TAM19, WKAL24, YJ11, ZPAQP11, NS24]. Selective [ARHERZ10, CJK16, KD22, MDM17, NYS15, NSMK17, SU12, TR15]. Selectivity [El 20]. Self [MRTB16, Dum20, FY12, IEME12, KS14a, Kum14b, MK12, MBZA16, Mos11, NJAK15, NKD14, Ngu24a, PKP13, SS10b, SK12a, SSK^{+}10, SS12b, SAR18, SDS15, UA14, UR16, VS12, SAS24, MMA13]. Self-Adaptive [Ngu24a, SAR18, UA14]. Self-Adjustable [MMA13]. self-broadcasting [SAS24]. Self-excited [SS12b]. Self-Organized [SS10b]. Self-Organizing [Dum20, FY12, MK12]. Self-Training [IEME12]. Self-Tuning [MBZA16, SK12a]. Selfish [PP13a]. Semantic [AEk13, AS12c, AGH14, BEE13, CPG15, EnFn14, Far14a, Far14b, Far17, FH13, FNH13, yGL13, IHD12, IM12, JG12, MM11b, NF15, NK14b, SU14, AaEeK11, ASI17, MK13b, Ojh11, SZM18, Zar24]. Semantic-based [FH13]. Semi [BS10, COWK14, CCG12, GS12b, Hus15, Jas13, KKR^{+}14, PKP13, PPK13, VK13, RPMH24]. Semi-Automatic [COWK14]. Semi-Optimal [Jas13]. semi-parallel [RPMH24]. Semi-Projective [PPK13]. Semi-structured [GS12b]. Semi-total [BS10]. Semiconductor [AM14]. Sense [Eg13, SG15]. Sensed [MJ14]. Sensing [AMJ16, BS17, BLL13, BBB16, BCC12, DCM15, EF15, GKM^{+}18, JVG15, KZ16a, KSSM24, KS14c, MSS14, PPB15, RJC17, SIAI14, SG14, SSSS12]. Sensitive [Afr17, BLL13, BHK13, BOK15, HHEY19, JPJ18, PA24]. Sensitivity [Fen18, WSW^{+}13]. Sensor [AII13, AAY21, AOO20, AWA12, ABB20, AU23, AmM12, AT15, AV13, BSM17, BS14a, BBHL12, BCC12, BJPP15, BC22, BKS14, CC17, DJG12, DS10, DSA11, DAB12, DKS14, DGP14, ETLE13, EE16b, EG15, FKFU12, FR12, GEA16, GSD14, GGI14, HZS14, sNIALY13, JI10a, JYP10, JTKC14, JSK17, KS16a, KMMM12a, KMMM12b, KCS14, KC11, KPV12, KA13b, KR16, LGM12a, LGM12b, LSS20, LA15, MM14a, MK13a, MFA13, MMP21, MP10, MK15a, MJS15, MMK21, MNN10, NM16, OC20, OHI15, PS15a, PSS12, PPU11, PS15d, RS13a, RMVP14, RKB19, RP14, RV12b, RB15, RR12, DBS15, SE12b, SE16, Sal17, SOO15, SMJ12, SPP11, SBHM13, SA10, SK10b, SU12, SS12e, SM13d, SSS16, SL11, SV14b, TBR15, TK14, TMC17, TB14, TOO16, UR16, UO23, XRL14, YYS^{+}23a, ZBI14, ZCDR13]. sensor [SKS16, YSH12]. Sensor-based [BC22]. Sensorless [AH16]. Sensors [AIC19, BT14a, EhMA13, FTLK14, KFEN21, OAA13, RR12, SVB13, TAMA13]. Sentence [MS16, BSM^{+}12, GK10]. Sentences [FAt13]. Sentiment [BBB16, DFGG15, HASH17, JK14a, Rok20, WRJ23]. SEP [BBHL12]. Separability [MS12]. Separating [SB21]. Separation [ASN12, AAR13, CM15a, HO15c, MK14, SET15]. Separator [DK18, HF16]. SEPIC [MH12b, RURS13]. SEPICS [ER12]. Septic [VR15b]. Sequence [AAP13, AN16a, AS18b, BMF11, BYT14, FLLR20, KSS12, KEBF13, Lah16, MJM17, NW13, RD13, SAK18, SVB17, TR14a]. Sequence-Dependent [BMF11]. Sequences [BBD13, MRTB16, Hos12, NK15, OKS14, RSA16, Sin13, TKC^{+}13]. Sequencing [Bon15, JHP12, TB12b]. Sequential [BRS16, BO21, CTG12, EmErB13, NP13]. Serang [RMH20]. Serial [MBKC15, Zon17]. Sericulture [MSA18]. Series [AMAY20, EBAS11, ER12, mH14b, ISM^{+}17, KK13a, KR10, KRM12, MC17b, Mal13c, PKB13, Poo22, PMM18, Raf14, RSP11, SK15c, AD13]. series-series [AD13]. series-shunt [AD13]. Serious [CHG12, DDBD22, EK14a]. Seroprevalence [SP11b]. Served [MGA20, SSN18]. Server [Al 16, AS13d, BLQ21, CHW16, HKL16, KS18, MG12, MM14b, SDB15a, JS14]. Servers [AMG^{+}12, Car18]. Service [AOB12, Abd14a, ATA14, ASC15, AICN14, AzSKM12, AMG^{+}12, Ara15, ASG10, AS13d, Cai17, CSKB11, CDC18, DDL18, DS13, DA18, EzHEgT13, EBB17, GRL20, GM13, KUM13a, KDT16, KC15a, MK13b, MN24, MMTS13, MSP12, MNN10, NG10, Nor13, OAA14, Ojh11, PUW13, PG11, RS16, Sm13b, She12, SSP15b, SAM14b, TTA10, UJ14, WT13, YEE14]. Service-based [RS16]. Service-oriented [MN24]. Services [AM11, AHH13, AWK19, ASN12, BS11b, CS18, Elm22, ES12b, FLM11, KR11a, KC15b, MK15a, NNR16, RKGK15, RKGK16, RGK17, Roc17, SP16, SH15, SS14e, She12, TA10, WDAJ15, PWA12, VMA23, CDC18]. Servomotor [NK14a]. Servqual [PUW13]. Sessions [Cha15]. Set [AS18a, AGL17, DB17, DAG16, DLN12, Eba13, Eis13, GHH15, GK15a, HAA12, JA12, JSSA20, KKBS12, KS15d, KKSS12, LAZ17, MMPH15, MB17, NM20, ST11, SGKS12, SL15, VG12, ABF14, KSRS12, NMS10]. Set-Valued [GHH15]. Sets [ANM17, CRN10, Eis14, ESAA18, MB12, MP15, NYU14, Pil13, RPV10, S22, SNB18, XCY11]. Setting [ARAA11, AAI13]. Settings [am12a, AA14b]. Setup [KC15a, MJM17]. Setups [BMF11]. Seven [SS12d, VSN13]. Seven- [SS12d]. Seventh [VR15b]. Several [Ker13, RFAH14]. SFF [MAARIH16]. SFL [YEE14]. SFLA [Ngu16, Ngu18, Ngu21, Ngu24b]. SFLA-Based [Ngu24b]. SFM [Sun19]. SG [EK14a]. SG-LOM [EK14a]. SGA [HB14b]. SHA256 [BC22]. Shadow [MRTB16, KA13a]. Shadowed [DM14]. Shadowing [NAH15, SSMP13]. Shafer [AB14a, RMD10]. Shallow [RÇ16]. Shamir [IPR15]. Shannon [MM18]. Shape [AZM13, GP15, GPC16, HSS11, KNGT10, RMD10, SB13b, TMK^{+}10, WG14]. Shape-based [AZM13]. Shaped [AI13a, JT15, SRM13]. Shaper [JMM21]. Shaping [JMM21]. Shapley [RKGK16]. Share [SAA24]. Shareable [FNH13]. Shared [CLBA^{+}12, DN12, HSMb12, MM15c, RB16]. Shared-Memory [CLBA^{+}12, HSMb12]. Shares [SSA14]. Sharing [AC18, AMG13, AAC13, BNA14, ER12, JM16, PA14, RKGK16, RJC17, SK10c, CS11]. Sharp [SE14]. Sharpening [PSR13]. Sharpness [LG13a]. Shear [HSEA15]. Shearlet [SC15b]. Shedding [AY18]. Sheet [PRK17]. Shell [LSM16]. Shielding [Say22]. Shift [KC14a, MY13, TJ12]. Shift-Phase [KC14a]. Shifter [PMK12]. Ship [PP15]. Shock [YS11]. Shocks [MRH15]. Shop [AF15, AMTM11, BMF11, GN14, GF17, GSA12b, wHsYwT12, LK16a, MJM17, RGCGCMFJ13, TB12c]. Shopping [AS18a, ENC12, JKM^{+}17]. Short [Ali16, Cuf14, GF14, JPDB14, PKJ15, QKRB11, SMN^{+}22, SAA24, She11]. Short-Term [JPDB14, SMN^{+}22, SAA24]. Shortcut [NS15a]. Shortened [KK15b]. Shorter [AT17]. Shortest [Kum14b, RK17]. Shots [GP12]. Show [MKT17]. Shrink [CT13]. Shrinkage [KJK12, SD16]. Shuffled [Ngu24a]. Shuffling [rPGM14, Ngu24a]. Shunt [AY18, SGB16, SK15c, AD13]. Side [STG12]. SIDIKA [AIP18]. Sierpinski [SSK13b]. SIFT [SB11a]. Sigma [LGM12a]. Sigma-Delta [LGM12a]. Sign [AZM13, MS16, CS18, EEE14, EA18, NS13a, SP18c, TSA12]. Sign-On [CS18]. Signal [AzSKM12, AAoS14, AHM14b, AMH11, BNS17, CCG19, DC22, GC18, HS15a, IEG14, ITK21, KHAE19, KST16, KB15, KR11b, LsAAA13, LT13, ML12, NRR12, RKB19, SD16, SG13b, SN14b, SA15c, SM13c, TR14a, WWSS12, GBM10]. Signals [BSS15, BK13, EAB13, HW18, JAB22, LSK17, MVC10, OSA14, PT13, PMA13, TT14, TN14a, NDC10]. Signature [AD11, AMK13, BB12a, BHC12, DI10, KTS10, KGK10, RRVP12, RMAM19, S14, SS14b, SB13b, US14]. Signatures [Ban11, VNM15]. Signcryption [Abd17, AR17a, KBAQ15]. Significance [AVH10, ATM18, CKT10, PMA13]. Significant [EAF^{+}16, GEB12, MA13b, OG16, Sab16, SD17]. Signs [IHD12, TBKA13]. Silica [AHZ15b]. Silico [NYS15]. Silkworm [MSA18]. Silo [SGD^{+}20]. SIMD [MNB13c]. Simevents [YJ14]. Similar [RKR17, FYM11]. Similarity [AAP13, AlS23b, AYS19, JFJN20, PDR11, RB11, RKR11, S24, SAW16, TY13b, REA12]. Similarity-Based [S24]. SIMO [SPXT17]. Simple [ABYB22, BA23, DDL16, Jas13, KJ15, SB21, SFVA14, Sud10].

Simplex [CCG12]. Simplified [AZ17, Cha17, IhZ12]. Simulate [AMK13, FK15]. Simulated [BOS15, FD14, GMAR14, GEG23, GB18, TB12c, TA10]. Simulating [KCS14, ZSR14]. Simulation [AKK11, AS15b, AAoS14, AVS13, AKS^{+}13, BSM17, BAIs^{+}14, CPSP13, DF15, DK18, ETKE20, Elg13, FM13, FDG17, FR15, GA15, IhZ12, JC14, KC15c, MBM14, MS10b, MBKH16, MB13c, Ngu22, OAR16, PSD13, PPR^{+}12, Raf15a, Raf15b, RSS^{+}11, SDW21, SHS23, SKJ15, SC15c, SPXT17, SZUSNK19, SET15, Suj14, TAM14b, TB15, VR14, YR12, ZKN10]. Simulations [AE12b, AAC13, Eja12]. Simulator [CN10, JSD14, KDB17]. Simulink [AB13a, CBhM12, KSP11]. Simultaneous [MNP11, DL23, MJM17]. Simultaneously [CM13]. Sine [TK15, YT15]. Single [Ahm20, Alg19, ASK17, CS18, DJG12, FPS16, Jag15b, KJA15, KPM16, LSB12, MAI15, Nah13, MK14, NSMK17, PN10, PJ12, RKB19, RC14, SDB15a, SVRB15, VGJ12]. Single-Bit [FPS16]. Single-cell [Ahm20]. Single-Server [SDB15a]. Singleton [HAPG17]. Singular [AAEK18, AFW14, KVN10, KK12a, KAP11, LCP21, PKP13]. Singularity [Rag15]. Sinhala [HPM20]. Sink [KS15c, NBL15, NM16, UR16, VM14a]. Sink-based [UR16]. Sinkhole [SC16]. Sinusoid [VGJ12]. Sinusoidal [KM16a]. SIP [MNYMGG12]. SIR [ZRE16]. SIRS [HLZ^{+}24]. SIRS-ARIMA [HLZ^{+}24]. SISO [AMN12, AZS16]. SISO-OFDM [AMN12]. SISO-OFDM/WOFDM [AMN12]. SISOTOOL [AM17c]. Site [BC12, DBM14, ESH^{+}14, IR16, RSB12, AA18a]. Sites [AHA14, AA18a, Ras17]. Siting [LRRR11]. Situation [aKNP^{+}15]. Situations [PWA12]. Six [RJA23, SS12b]. Six-phase [SS12b]. Size [AA13a, AHM17, BM13a, DA10, Eba13, KVKR16, NR12, OMM14, SBHM13, SAk13b]. Sized [PS15d, PMP14]. Sizes [KKGG12]. Sizing [AY18, BMF11, GN14, KK14, LRRR11, MBL15a, NHH10, YAT15a]. Skeletal [MB11a]. Skew [JKC11]. Skills [AARA19, AK18a, CHG12, Mar12]. Skin [EAEE13, MMK11, NDC10, NPA^{+}24, SO13]. Skip [KS17]. Skull [AVH10]. Sky [ABL17]. SLA [AsJ12]. SLAM [ESR13]. Slave [CM13, Ngu24b]. Sleep [GB15, GKG16]. Sleep-wake [GKG16]. Slew [KRMB10]. Slice [AAY17]. Slicing [AmAr14]. Sliding [AA12a, ANE^{+}13, ABS^{+}12, GAA15, GG11, HZ14, SFA14, SKM16]. SLIM [SFVA14]. Slip [HIU11]. Sloping [Raf15a]. Slot [PB16, SSG14, TRDB24]. Slots [LK16b]. Slotted [MT13b]. Smali [DBKN20]. Small [AAA14a, AmAdS12, BA12a, DDL16, GS14, GK15a, KSP11, Kam17, KNA12, LT13, OK14, PGPM23, PMP14, RHM15, RHT19, SKZ22, SVB23]. Small- [PMP14]. Small-Disjunct [AmAdS12]. Small-Scale [SKZ22, SVB23]. Smart [AMT^{+}17, AIC19, APS18, BBB16, BCC12, BA16, CMKS15, GEAE18, HS24, HA17, Ibr16, JS14, JKM^{+}17, JTI19, KDM17, MBM14, MDCS22, MSA18, OD13, RRudS12, RR18, RRAFAM19, RKBB14, RMH20, SK15a, SSA^{+}23, SL14, SC19b, SK15d, TAMA13, TVS19, You18]. Smartcard [MMG16]. Smartphone [DRBR24, GK15c, KS15a, Kad16, WNE21].

Smartphone-Based [DRBR24]. Smartphones [AA17b, EBO13, EB13]. SMC [ANE^{+}13]. SME [OK14, PMP14]. SMEH [SB15]. Smells [GSV20]. SMES [RJC15, BA12a]. SMF [AJ10]. Smoke [NC14]. Smoking [KKI16]. Smooth [Anh15b, Ori14, PSR13, KAS12]. Smoothed [HA16]. Smoothing [BI13, JLCV16, LZW15, SVRB15]. SMOTe [RAR24]. SMS [SB18a]. SMS-based [SB18a]. Smut [JA22]. Snippet [TY13b]. Snort [MAE16a, MSR16, GD13]. Snort-Negative [MAE16a]. SOA [AsJ12, MR15, ND17, TO13]. SOA-MZI [ND17]. Sobel [NK12]. Social [AAAA17, ATM18, BADR16, BM16a, BH17b, DMS13, ISM^{+}17, IM16, KVR14, MK15b, NHH10, NM19, Rok20, SAA24, SS14d, SW17, Tit16, USR17, WK16, YBM16, SAS24, Far17]. Socio [KTA17]. Socio-Cultural [KTA17]. SoCs [AB12a, RGS15]. Sodium [GMRK12]. Soft [BV14, JSSA20, Jah18, KDBT12, KETESH17, KT12a, MLRU12, MAM12, PG15b, SY12, SSS15b, TS16, You13]. Soft-DF-based [PG15b]. Soft-Max [SSS15b]. Soft-Set [JSSA20]. Software [ADH17, AHZAH16, AAYM13, AZY18a, AAA20, AZA23, AS15c, ANM17, AA18b, AYYB15, AI18, AKA^{+}18, ADS16, BSS12, BT16, Bou12a, BK11, BA12c, CK12, CS15a, hC13, DLBM12, DR15a, DSDD12, DHM15, EsTeEh^{+}13, FCS11, FR15, GM20, GSA12a, GL16, GAZ13, GVW24, GC12, HB14a, HK19, HZBG13, IAQ11, Ira12, KRR11, KA16a, KHSE12, Kas17, KS16c, KJKK12, KS13c, LSA^{+}17, MR12a, MHE15, MJP15, MT13a, MR12b, MB13b, ML13, NP12, OMM14, OFE15, OV13, PS14a, PS16b, RRR13, REUHTG18, RGK17, RP24, RH14, RKS16, RKS17, RS16, RPS12, SFA14, SATS13, SOT18, SS12d, SC12, SM12b, TO13, VV18, WH21, YS11, BAA24, GT11]. Software-based [AAA20]. Soil [JYP10]. Solar [AN16b, ANAK17, DD20, GA15, IM15, KC14b, LAA^{+}17, MB13c, RKE23, RBDD16, SC13, Tal11, ZAAS18]. Solid [Job17a]. Soliton [KZ16b]. Solomon [Ish14, KG11]. Soltuions [KZ16b]. Solution [AMT^{+}17, AHH20, ASII11, CDC17, EEEE16, Gha16, GAZ13, GCS14, HF20, HII15, JV14, KSZK15, Kub17, MMG16, MG14c, NAH15, Ojh11, RSM15, TNKS15, VR15b, WDS13, MDCS22]. Solutions [FHM16, JI10a, KA14, MREd14]. Solve [AI18, AMTM11, BP15, HBA13, KKI16, SDS15, SS13f, Eg13, TA12]. Solver [EDD19, RA12b]. Solvers [LNM24]. Solving [MNP11, AMS13, AFW14, AMAHH19, DMM14a, DD13, DVS13, EBAS11, ESZED15, EG14, GMAR14, GV15, wHsYwT12, JSY14, LSB17, MS14a, MB15b, OON17, OESKF16, RGCGCMFJ13, SK13a, ŞAG17, VGo12, FD14]. SOM [AkC13, Dum20, VN10]. Some [AD15, EFA16, ES15c, KBG11, Tal11, ADFAJ15, EEB16, PR10a, RSC16]. Sonar [NRL17]. Soret [SH17]. Sorting [ASCR11, CHW16, PS15b, SWH16]. Sound [ETKE20, HR22, MO17]. Soundex [She11]. Source [AA18c, AB16b, AKB16, BT16, DLR19, EhMA13, GM15a, GS14, HR22, HO15c, KS16c, KNS17, MK14, RURS13, SS16a, VM14a, CS12, GAA17]. Sources [CR15]. South [RSK15b, IND13, SP11b]. Sow [BML22]. SP [CKD12]. Space [AP15, BM15c, Cai18, CS11, CKD12, DP12, GK12b, HJE14, IEG14, KR14, Kum14b, LOAO23, MN16, NA16, PDK12, PW23, SS10e, TR15, VK11]. Space-sharing [CS11]. Space-Time [TR15, VK11]. Spaced [AA16a]. Spaces [ACK12, KETESH17, KTDG12, MMK11, MP15, TR13]. Spam [AM13a, KNA12]. Spare [PRK17, VR14]. Spark [GMM21]. Sparse [AAR13, Bha17, FML20, MM16, OSA14, PK19, SLU14]. Sparsity [GCK17]. SPARTEN [SFH17]. Spatial [AAEK18, BS12b, BIE^{+}12, EAF^{+}16, IHH14, JVG15, JSS^{+}17b, NTL18, Ota13, PP10, PSG12, KTA17]. Spatially [RT13]. Spatio [FAR16, LABV21, MH13b]. Spatio-Temporal [FAR16, MH13b, LABV21]. Spatiotemporal [AE12a, Hik13]. Spawn [EMF14]. Speaker [FAt13, GJ13a, IR10, JHB17, KKGG12, KK10b, YB15]. Specialized [RC12]. Species [LEE13, PC15a, RACI24]. Specific [CS14, DL23, ML13, Roc17, SS10b, TK^{+}17, ZAER12]. Specific-domain [ZAER12]. Specification [BB10, BBC12b, DS12c]. Specifications [HE15]. Specify [AHZAH16]. Speckle [MRR21, PBVD22, Sam13, SN10b]. SPECT [RMD10]. Spectral [AAEK18, BNS17, BIE^{+}12, KA10, PN10, Tha14, VDM22, ZCZ24]. Spectral-Spatial [AAEK18]. Spectrograms [KKGG12]. Spectroscopy [NJ16]. Spectrum [DB10, DCM15, EF15, GKM^{+}18, KB12a, LK15, LCP21, MVC10, Pil13, RJC17, SIAI14, SG14, SSSS12, SB12a]. Speech [AH14, AJI15, AAS13, CCJ14, Deb20, DB10, GCK17, GMS12, HS15a, HW18, Isl18, JHB17, KGK10, KAZA13, MP12a, NRR12, PN14, RSHI12, RK16, SV14a, SD16, SU18, SC14a, SAI16, US12, WK16, KBA10]. Speed [AA12a, AZM13, AAEEA15, CPSP13, DV14, EAM14, Ena12, GAA15, GGI14, GPR14, HR16, HAA16, HB13, HSEA15, JTA19, KPC16, KTS10, KKT15, KT10, KKS14, Mis14, NK14a, PCW10, PIB11, Ras16, SO14b, SSP15a, VGJ12, Ali12]. Speeding [ElZA13]. Speeding-up [ElZA13]. Spell [MAt12]. Spell-Checker [MAt12]. Spending [RD17b]. Sphere [Job17a, RPN14]. Spikes [KSV10]. Spiking [ASC14, FK15]. Spills [SABA12]. Spine [KB12b]. Spinous [PCC12]. Spiral [Zaf11]. SPL [EsTeEh^{+}13]. Spline [BI13, SAK13a, VR12, TOMGBSS13]. Splines [HSS11, VR15b]. Split [KKS14, MVC10, RKR17]. Splitter [RHU12]. Splitting [UNM18b]. Spoken [AS12a, Bha10, FAt13, SSJ12]. Sports [SA15a]. Spots [JA22]. Spotting [SP15c]. Spread [BI12, LABV21, PRK17]. Spreading [IAA13]. Spreadsheet [SGKS12]. SPS [HB14b]. SPSS [Pra14]. SPV [SGS15]. Spyware [SZEM16]. SQL [AK16, BLQ21, Bas12b, DSB12, NKG12, Pet22]. SQLIA [CJK17]. SQP [AMN15b]. Square [Ae14a, BWS16, SLK^{+}13, SS23b]. Squares [AlS24e, AlS24h, AlS24i, DK11, SErEs^{+}12, Tah17, Zyo12]. Squeezer [REA12]. Squirrel [RUB11]. SRAM [HAS17, Jah18, RSK15a]. SRAM-based [Jah18]. SRGM [Sar16a]. Sri [MJP15]. SRM [BBL13]. SS [HKA15]. SSAW [HF16]. SSAWs [SET15]. ST [Mah16]. ST-coded [Mah16]. Stability [AmH13, EEKG17, Eln14, IK11, KRM12, MBL15a, MEE16, Nad14, OAS19, PVA13, S22, S23, SMVK11, SK15c, TY16]. Stabilization [AVT13, CZW15b, FTBA16]. Stabilize [FYM11]. Stabilizer [AA12b, AA12c, RT20]. Stable [AA12f, BBHL12, MDM17, WKAL24]. Stack [MM20b, MB15c, MBR16]. Stack- [MB15c, MBR16]. Stacked [RGS15, SSD^{+}13]. Staff [OAO18, Qte14]. Stage [AEEO24, ASH13, ABBG14, GSA12b, Kei14, LGM12b, LRRR11, PBK14, RRR11, RKS16, RKS17, VSS16]. Stages [Hab13a]. Stamps [HS18]. Standard [AsJ12, AAOA23, HMB13, JM14, Jas13, KG15b, KB16b, MB11b, PBG12, Sar13, SSC15a, VRS12, BM20b, SQ18]. standardization [MJM16]. Standardizing [urPP13]. Standards [AKH15, BNS15, PB19, SKG17]. Standby [BKM15, Es12a, MAA13, MHE15, MRH15, Poo22]. Standing [SET15]. STAP [EMA11]. Star [AAA14a, DM14, SEA16, OHT13]. Started [HD14]. Starter [EG14]. STATCOM [AJ19, RJC15, Ali19]. State [AKK11, AYAI17a, BK10a, HO15c, VNM15, JV14, KS18, KNS17, KRR15, Liu24, MHE13, NP13, NUK16, PS15e, RUB11, RRB16, SAK18, TBA12, TN14b, Poo22, BI12]. State-control [JV14]. State-of-Art [HO15c]. State-of-the-Art [SAK18, NUK16]. Stateless [Meg10]. States [A18, EKA13, HKA15]. Static [BJ12, DBB12, JKTM15, KZ16a, MM17, NS13a, NC14, SAKS24]. Station [AOO20, DSA11, HB12a, KDTD11]. Stationary [GA15, Gha13, MVC10]. Stations [OO14]. Statistic [SSY11]. Statistic-Fuzzy [SSY11]. Statistical [ADT17, AHZ14, AA16b, CPD13, CS11, DRJ14, Eba13, GK21, HFSM12, KMS14, Moh14, NM20, NMN15, PM15, SN15a, SM23, SB12b]. Statistics [HKE^{+}14, MAaM14, ST14]. Status [SH23]. STBC [PT15, TR15, Tha14]. STBC-OFDM [Tha14]. Steady [KRR15]. Steam [SZOK24]. Steel [ABD22]. Steerable [KS16b]. Steganalysis [AH13a]. Steganalyzer [AM11]. Steganographic [PG13a]. Steganography [ASH16a, AAH21, AM19b, AS12d, AS16, ADVB10, AHAT16, AA17b, Bas12b, EAF^{+}16, EAO15, GEB12, HMO17, Kha11, Lah16, MS10a, MMR14a, NA20, NmEM13, OG16, RS14b, SS10c, SS13c, SK16, Sha18, TAS^{+}16, TS14, TRSK15]. Stego [AGRR10]. Stego- [AGRR10]. Stem [Abb12a]. Stemming [DB12a, Yac15]. Stenosed [HS14]. Step [AAA19, CCG12, FML20, IAM16, MAI15, SD16, SA13b, SB18b, Zaf11]. Steps [Naj12]. Stepwise [PDK14]. Stereo [AOK14, Oua12, Rez12, SM19, TH14]. Stereoscopic [PKDS17]. Stethoscope [PP16]. Stie [HS24]. Stiffened [EH16]. Still [EAMAES16]. Stimulated [TIT14]. Stitching [PR16]. Stochastic [BKM15, FYM11, KM12, KK22, KR11b, LABV21, MRH15, PVA13, SK12a, SB15, YB12]. Stock [BWH15, DH17, FEbO13, GT13, GK21, Jab14, KGAM10, Lah11, Lic20, NMS10, RK13, Raf14, SAA24, SMK14]. Stock-Dependent [SMK14]. Stone [AAN^{+}19]. Storage [AC18, AYYB15, ARL16, BH14, Bat18, CMKS15, EB21, JVF11, JMM21, MR15, SGD^{+}20, TGR12, TYSS15, UDA13]. Store [BB14b]. Stories [HS15c]. Storing [BBL15, Pet23]. Story [Asa13, ADKsO14]. STQP [FAR16]. Straightness [NH13]. strata [NRL17]. Strategic [AA18c, ES15b]. Strategies [BSG15, BA11, Elg13, Has24, JI10a, KAR10, KBEA17, MBKC15, MNB13a, SZOK24, VSN13]. Strategy [AK16, ARZ11, BM14a, BFA13, ER12, FSER20, GGI14, HS15c, HMAA13, HAb14, JM18, KM15b, KRSB12, Kum13b, NEA17, PT15, SM11, SGD^{+}20, RSK15b]. Stratified [HR13, ILG15]. Stream [AMHAh13, AYYB15, AS14d, BRD12, EEGHS15, EA11, JP15, KSZK15, SPBP17, ABA^{+}10, KWS^{+}11]. Streamflow [HHHK21]. Streaming [Alt13, HASH17, HKL16, SPP13]. Streams [MS14b, PRR15]. Strength [RKB19, RAC12, SA15c, SM13c, Sou15]. Strengths [MR13b]. Stress [AE13, KAL15, TD21]. Stress/Emotion [TD21]. Stretching [BH17a, PSR13]. String [GSA12b, NW13, Ode14, PRT^{+}16]. Strip [RB11, SKJ15]. Stroke [PNIS16, SGJ12]. Structural [AH11, BH17b, BB13b, DR15b, IM12, ISH12b, LBB11, Say22, STG13, TERCM15]. Structure [DMS13, DHM15, EAS16, HIE14, QT15, RM12, RP10, RAHH16, RL14, Sar13, SRM13, VK13, TMC17]. Structure-based [RL14]. Structured [HAS20, JPJ18, KW14, SA12b, GS12b]. Structureh [GM14a]. Structures [KE16, LCP21, NH13, PA14, SDD14]. Structuring [AOK18]. STTRAM [KKT15]. Student [AARA19, AA14c, AAO11, LG13b, DUP13, DTO23, ED19, KA14, SD21]. Student-Centric [LG13b]. Students [AS14b, Add12, AD14a, AAEA16, Ash19, BT16, CAS14, GSP18, GF14, Mar12, urPP13, SS14d]. Studies [CHG12, EDR15, EK14b, FMS13, GSP18, NYS15]. Studio [Kar11]. Study [AAA14a, AHZ14, Ahm20, ATAAAF19, AA13d, Al 13, AR17b, AAA21, ATB12, AI13b, Ali19, AAA^{+}17b, Ans14, AF15, AC13, AE14b, ARR20, BEE13, BME19, BGS11, BML22, Bou13, CKS13, CV13, CAS14, CS18, CKA18, DKP21, DC22, DDBD22, Eba13, ES12b, GP13, GBPTEQ^{+}13, GAA17, GK12b, GAZ13, Ghe16, yGL13, GG12, HMGV12, HE15, Hik13, HS24, HB12b, IPB14, JSD14, Jag15b, JSY14, Job17a, KDT16, KS10, KS13c, MSEEs13, MKSB22, MEAT11, MB11b, MMTS13, MAP13, MJ14, MEE16, MTM12, MB12, MHV^{+}14, MRG19, NJAK15, NAH11, OAG17, OC13, OJ11, PHK^{+}15, PBG12, PD14, PG11, PP13d, PK19, PKJ14, RVKB12, RFAH14, RPV10, RS11, RMD10, Red19, RKGK15, RKBB14, SJA13, SWP19, SJK21, SS14d, SKS10, SAK13a, SNA13, SSZ16, TTA10, TGR12, TOD13, TR14a]. Study [UA14, VS12, WBS16, YAE11, YBM16, YAT15b, ZD15, ASNS10, ABA^{+}10, HAA12, OK14, SN10b]. Studying [TSS12]. Sub [ES15a, IM12, NRL17, ND15, PPP11, RM16, SDG18, SM15a, TERCM15]. Sub-band [PPP11]. Sub-gigahertz [SDG18]. Sub-graph [IM12, ND15]. Sub-Image [SM15a]. Sub-Ontologies [TERCM15]. Sub-strata [NRL17]. Subbands [SCL14]. Subchannel [VT15]. Subgraphs [KPM16]. Subject [VNM15, KN12c, KM12]. Subject-Role [KN12c]. Subjected [EN13, Job17a]. Subjective [HEEG15]. Subjects [AE13, RA10b]. Submucous [PCC12]. Subrings [AS13b]. Subscriber [KDTD11]. Subset [Has15, Sin16]. Subsidiary [KAFS15]. Subspace [AS15b, KK17b, MRM13]. Substantial [KA12b]. Substitute [GYP^{+}15]. Substituted [PS12a]. Substitution [MAk12, VS10]. Substrate [KNGT10, PB16]. Substring [AAP13]. Subtitle [PH23]. Subtitled [Far14a]. Subtraction [KA13a, MKK13]. Subtractive [AHA14, Ber11]. Subversion [BO17]. Success [DR15a, KMG20, Qte14, Sul17]. Succinylation [AA18a]. Sudan [ZAA13]. Sudoku [EDD19]. Sufficient [KIH15]. Suffix [ABF14, JG12, PRT^{+}16]. Sugar [EUEI23]. Sugeno [GSCT12]. Suggesting [JBP16]. Suggestion [KM13b]. SUI [SNJ16]. Suitability [SVMLb13]. Suitable [FR12, KMB14]. Suite [MA22, BM14a, KNS17]. Suited [SZUSNK19]. Sulawesi [PPR17]. Sum [DSJ11b, KS13a, MMA13]. Summarization [ASA14, BEE13, GEG23, MD13, OO21, SS17a, SO14c]. Summarizer [BSM^{+}12]. Summary [KSS18]. Sunlight [Moh13b]. Super [DT22, GGX12, MESNH17, SR17, TAH13, TY16]. Super-Resolution [GGX12, MESNH17]. Supercapacitors [EB21]. Superior [DLN12]. Superiority [SSS12]. Supervise [PTPS19]. Supervised [AkC13, AA11b]. Supervision [ED19]. Supply [AB12a, FS11, MABE16, UAMN17]. Support [AA14b, AGFA16, AA17a, ASNS10, AA18a, AIG12, AO13b, BBC12a, BK13, CMKS15, DR12, DB17, Eis13, EFA16, EEA19, FKT12, FAA13, Gho11, GOCA15, HEEG15, IEME12, IR16, IAA13, JV12, JBP16, KH20, KDB17, KJTM15, MH12a, MJ15, MMP21, MCA14, MNN10, NWS15, NI15, OOO19, OAG17, PNIS16, PSG12, PG13c, RÇ16, SB18a, YA18, ZAA13, Zyo12]. Supported [ADFAJ15]. Supporting [AA18c, KA13b, Sha14]. Supports [MM18]. Suppression [HCL15, MKB13, NL15]. Surface [BH17a, HAAA15, Haz14, ISH12b, KM16a, KMS12, MDM17, SET15, TRM23]. Surfaces [DPS21, HSS11]. Surfer [HG11]. Surgery [HO15b]. Surrogate [HBA13]. Surround [AN14]. Surveillance [AZA23, DR12, aKNP^{+}15, MOM17, PM21, SG13b]. Survey [AGN14, ASASHF17, AJ16, Ash19, BAS12a, BA12a, BNS15, CP15b, CAS14, DSD15, EHA18, EK14b, EMS19, HAS20, HEBR11, ISM^{+}17, JTS15, KM16b, KK11, KW14, KBK12, KC11, MSM17, MV12a, MK15a, MNN10, NG10, MK14, NM19, PP13a, PK15a, Rah13, RRB16, SB11b, SMGS17, SK10a, SG13b, SOO15, SHR11, SAK18, SSP13b, SK14b, TR14b, TS11, VGP15, WAM12, HJB^{+}14, MKCV13, SA12a]. SUSAN [SS11]. Susceptibility [Obe23]. Suspended [JT11]. Suspension [AMA15, AAA17a]. Suspicious [AD24]. Sustainability [Amb24, AZMP17, AZY18b]. Sustainable [AZMP17, AZY18b, WBL^{+}23]. Suzuki [ACK12]. Suzuki-Type [ACK12]. SVC [AmH13, ATM12, AJC14, KWS^{+}11, Mos11, SMVK11, SD15]. SVD [EAMAES16, NP11]. SVM [BG16, BSS15, BIE^{+}12, CJK16, DS19, EE13a, KP15a, KA10, Lah11, NS13a, NP12, NP15, OOO19, PY19, PA12]. SVR [KJTM15]. Swahili [Tes15]. Swarm [AS17, AHAS^{+}17, AMSS13, ASS10, AA12c, AM17a, AHBG19, AMAHH19, AMAY20, AI18, AZ17, Ber11, BM13c, DCH16, DB12b, ED18, FR14, GEB12, GQD^{+}22, Has16, wHsYwT12, JS15, KA16a, KK17a, LGC11, MG14a, MG14b, Man10, ME15, OO18, PCR12, RHM14, SAY16, SM18, SMHS10, SDB15b, TY16, VR15a, Zyo12, VJR11]. Swarm-based [AMAHH19]. Swift [AAA21]. Swing [PSD13]. Switch [CM15b, PV17]. Switched [AHK12, DDC12, MLRU12, NSMK17]. Switched-Capacitor [NSMK17]. Switches [MK19]. Switching [AICN14, BCC12, Es12a, EKA13, KKT15, LG11, MM17, TAM19]. SWOT [VB10, Zar24]. Sybil [SS16d]. Symbolic [GVW24, HH13, Kha20, RA10b]. Symmetric [AN14, BC22, EMS19, RS13c, S24, SKGP12, UC15]. Symmetrical [VSN13]. Symmetry [rPGM14, KPM16, Kha13]. Symptomatic [MJP18]. Symptoms [CRBJ12]. SYN [AM17a]. Synaptic [FYM11]. Synchronization [AJ14, Dar12, MKSB22, SRS14a]. Synchronizing [TO13]. Synchronous [AA12a, Als18, BSI12, HR16, HK11, JPAK11, OA12, SHSA12, VSA15]. Syndication [ABYB22]. Syndrome [BML22, DOC^{+}19, LSA^{+}17]. Synergistic [RMD10]. Syntactic [LR12]. Syntesis [SFAS19]. Synthesis [AAS13, FR14, HW18, HMAA13, KAZA13, OS16, PZ10, PN10, Rah14, RK16, SPK16, YAA14, ZYW^{+}18]. Synthesized [SU18]. Synthetic [Has14]. SysRisk [RT11]. System [AmH13, AS17, AJ15, AHAS^{+}17, ATH12, AEE15, AVT13, Abd14a, AJT17, ANE^{+}13, Abu14, AB14b, AMN15b, AA14b, AGFA16, AYE13, AZA23, Afr14, AA14c, Agn15, AA12b, AA12c, AB16a, ATA19, AKN13, AU23, AAA11, AAO11, AD13, AB16b, AMA15, AQ15, AIC19, Al 16, AAA17a, AIRR19, AAA16, AER18, AERE21, ASM13, Alt11, AMA14c, AOSM12, AA16b, ARL16, ASS15b, AB14d, AO13b, AIP18, Ara15, AM10, ABS^{+}12, AAMA20, AAaN15, AA11b, ASI17, ur13, AT14, ARDS18, AAA19, ASG10, LEO15, BS18, BAF15, BAO16, BECM11, BLA12, Bef18, BBB16, BBL15, BA12b, BKM15, BD18, BK10a, BDN13, BA16, BNS15, BWH15, BJPP15, BYKO17, BB13a, BKB11, BIAE24, BOK15, CW23, Cai18, Car18, CAL^{+}14, CHW16, CSM21, CT14, CP15b, CSP14, CSN15, CBhM12, CB16, CS18]. System [CDC17, Dar13, DSJ11a, DSS18, De 21, DDL18, DLR19, DM10, DT22, DPJB12, DR12, DS12a, DBW19b, DOC^{+}19, DNR15, CB12, ESE13, EEH14, EmdHNmAW13, Es12a, EAE19, ETKE20, EhMA13, EAEE13, ENE14, El 14, EFA16, EAM14, EB13, EE16a, ED19, EOO22, EN13, Far12, Far11, Far24, FM13, FZM12, GS12a, GD13, GP13, GS12b, GC24, GB11, GP15, GK10, GDD13, GM13, GKK13, GOCA15, HF16, HSMb12, HJE14, HPC13, HE15, HEEG15, HII15, HB16, HS15a, HS10, HK15, HEN13, HMMN13, Ibr14, Ibr17, Ibr18, Ibr16, IS21, IGS17, IA18, IR16, IPR15, IRT11, IAM16, Iss13, IAA13, IM16, JIB14, JHP12, JA12, JD13, JBK13, JJN14, JJKD14, JAsA13, JH18, JP15, JJA14, JBP16, JPDB14, JIA13, JK14b, Joh14, JTI19, KA11, KSBM17, KM15a, KM15b, KH20]. System [KUM13a, KT10, KGK10, KA14, hKhHL13, KMB14, Kha11, KAqC14, KAZA13, KBEA17, KM16c, KDM17, KAA20, KN15, KC12, KG11, KJKK12, KM12, KVR14, KS15e, KD15, KNR15, KK16b, KNV10, KMG20, LJ16, LEE13, LK15, LRRR11, LSM16, LHKR16, LSB12, LC17a, LB11, LAA^{+}17, MWS20, MBM14, MBA20, MO17, Mah13, MN15, MAA13, MHE15, MRH15, MH12a, MG12, MC17b, MSM17, MKSB22, MA16, MOM17, MKAR22, Mes16, MAARIH16, MFE13, MH14a, MBZA16, MNR14, MBL15a, MZAE16, MMK21, MA17, MEE16, MPJ18, MTKP15, MCA14, MI13, MJP18, Mun14, MHV^{+}14, MEH14, NKH12, NRR^{+}10, NSA14, NMS10, NJAK15, NS13b, NRV12, NMN15, MSA18, NAD19, NMK15, NSMK17, NWS15, NI15, NU18, NB11, OJO11, OAA16, ONK24, OOE18, OFI17, Ojh11, OAA24, OEA18, Okp17, OO21, OAO18, OS14, OC13]. System [PAK16, PT15, PPR17, PMK12, Par14a, Par14b, PC15a, PSG12, Poo22, PA19, PDK12, PS12b, PA23b, QCERM15, Qte14, QKRB11, RMP11, RUB11, RÇ16, RSP11, RT11, RPN14, RAE15, RGR15, RV12b, RKR11, RW20, RRudS12, RR18, RRAFAM19, RJC15, RM11, RBDD16, RG21, RBB11, RO11, RKS16, RMS14b, RR22, RGJ11, RGNA13, RA23, SCD14, SS17a, SR15a, SMAM15, Sal18, Sal19, SSN18, SR13a, SK13a, SMS16, SK16, Sha14, SZEM16, STD14, SWH16, SMK17, SSG11, SS12e, SM13c, SGD13, SS15c, SSA16, SP18c, SNA13, SO14c, SB18a, SPSR12, SMC17, SS10f, SZS15, SN16, SB11c, TAMA13, TBN^{+}15, TTA10, TUS20, TY13a, TD18, TBA12, TS15, TY15a, TE16, TMP11, TJ12, TVY17, TO13, TD21, TY13b, TV10, TB15, TY16, Tua17, UAA16, UI16, USC12, UO23, VK11, VM14b]. System [VD10, VST17, WBL^{+}23, WNE21, Yac15, YB15, YAT15a, YMA16, YBA12, YA18, YB12, Zae13, ZEB17, ZEfNF13, ZHUR16, ZYW^{+}18, ZAER12, ZCZ24, ZAA13, Zon17, GSCT12, JVJ12, LLSL12, Mug24, PC11a]. Systematic [AOK18, BLA12, KR16, RS16, SKG17]. SystemC [HK11]. Systems [AAI13, AA14b, Ade13a, Ade13b, Agg13, ASS10, AKH15, AS15c, AFW14, AZS16, AMAB15, ATB12, AA12f, Alz12, AERS11, AB12b, ASS15b, ASB12, AAaAD12, BM13a, Bad13b, BAM13, BSI12, BH14, BS11b, BC22, BKB11, BA12c, CP15b, CKT10, CWC15, CZW15a, CZW15b, DL13, DDC^{+}24, DBW19a, Eis13, ESEMD21, ENEBES15, EBR12, EJC16, ESE16, EE13b, ETI14, FAZ13, Fen18, FK12, FTBA16, GGM23, HZK12, HS12, Has24, HHS14, Hus16, Hus12, IEES16, Jov15, KHSE12, KAR12, aKNP^{+}15, KTS17, KN17b, KKDC12, KNT11, KK14, Kum14b, LMZ12, MBAR12, MC17a, MAK13, MOOOOMT18, MB13b, MBE10, MHE13, MY13, NKR14, NRL17, Nas11, NBKM13, NYU14, Ngu24b, NF15, NMK14, NTL18, OS16, PS21, PN12, PHK^{+}15, PK12a, Par14a, RKE23, RS16, S22, S23, SSM12, SK12a, SRF12, SAK18, SRS14a, Sin14]. Systems [SPW15, SV18b, SE11, SPXT17, Sou15, Sta22, SJ19, SS13f, TB12a, TGR12, TG18, Tha14, TNC10, VV18, WDD16, WAG11, Wu21, Wu23a, Wu23b, YMR^{+}12, You18, BK11, Dar12, TB12b]. Systems/Cloud [SSM12]. Systems/IoT [BC22]. Systolic [Kum19].

T [SJV13]. T-F-T [SJV13]. Table [KJJ15, RÇ16]. Tables [BK10a, GHH15, SEA16]. Tablet [HS15c]. Tabu [JTA19, TB12c]. Tactic [HBS15]. Tactics [AXKT19]. Tag [MB13a, NMSA12]. Tagging [PY19, THT11]. Tags [SS13b]. Taif [Nad17]. Tailored [Far24]. Takagi [GSCT12]. Tallying [De 21]. TAM [ASS17]. Tamil [GP13, KKBS12, SS12a, SS12g]. Taming [PRK17]. Tamper [MIAEFS15]. Tandem [KRR15, RRR11, SRSV14]. Tangent [JFJN20]. Tangible [DOC^{+}19]. Tanh [RS14a]. Tank [AVS13, EOO22, SB18c, GSCT12]. TANNN [ASH16b]. Tanzania [Ana15, GPB22, NS14, Sha14, VM14b]. Tanzanian [KJTM15]. Tao [OAE16]. Tardiness [MJM17]. Target [CTG12, CK12, DMS13, FM20, Kum10, TN14b]. Targeting [AHZAH16]. Task [AAZN21, ALH^{+}24, CJ10, DBB12, ENEBES15, EHI17, Hem12, KSMM15, LS13, MT13a, MEH14, Ngu24b, RBF12, SZZR11, Sin14]. Tasks [AA14c, GQD^{+}22, KKR18, SE15]. Tax [GPB22]. Taxation [LR11]. Taxonomy [BBZG^{+}17, NS23, urPP13, QJ13]. Taylor [PMM18]. TB [KV16a]. TB-LEACH [KV16a]. TBDSA [CS15c]. TBDSA-Token [CS15c]. TCAD [YR12]. TCAM [KKT15]. TCDGP [DSD15]. TCP [HEs12, AQ14b, BJS13, CN10, DDC12, HEs12, JTS15, MKN12, NKR14, OMH11, PIB11, QA14, SN10a, SG12a, VjS12]. TCP/AQM [SN10a]. TCP/FTP [SG12a]. TCSC [KR10, LT13, MG14a, NHH10]. TE [AICN14]. Tea [CT13]. Teacher [ur13]. Teachers [BT16]. Teaching [AbAAw12, AAR14, DD21, GM20, HZLL15, KDB17, NAH11, PS14a, WDS13]. Team [BMRA12]. Teams [AIG12]. TEC [KDBT12]. Technical [AKN13, AER18, AERE21, CHG12, FEbO13, GAA17, Lah11, MOOOOMT18, MJP15, Maz21, OAG17]. Technique [ASH16a, AAH21, AYAB18, AOO20, Afr14, Agn15, APK11, AB14c, AD11, AA20, AS12d, AVS13, AS18d, ASK17, BHC12, BK13, BI13, CP15a, CS14, CM12, DS19, EUT17, EAMAES16, EAF^{+}16, EAO15, EOO22, GCP13, GS16a, GRP13, GR14, HKU19, Hos12, Ibr14, IHB^{+}15, JSS17a, JSK17, JTI19, JM16, MPR14, KM15a, KDBT12, KKS14, KA15, KA10, KPG13, aKNP^{+}15, KAAA13, KN17b, KM15d, KD22, Kub17, KC14c, Kum19, KR16, KMG20, LC17b, LKS14, MAE16a, MAI12, MB11a, Man10, MRJN14, MSA17, MBZA16, MAM19, MBR16, NTY19, NP11, NA16, NSM13, Okp17, OESKF16, PV17, PG13b, PS15d, RB14a, RAJ^{+}12, RBN^{+}16, RPN14, VGS11, RSK15b, RB16, SsTM12, SV12a, DBS15, SV18a, SB21, SSA11, Sar14, SK18, SMHS10, SM10, SCL14, SAS13, SM15a, SC14b, SSS16, SKS16, SVJ15, SL15, SVJ14]. Technique [TAM19, THAP19, UC15, YB15, YK16, YJ14, YMA13, You18, GBM10, RPMH24]. Techniques [AAN^{+}19, AMBS16, AARA19, AS14b, ARZ15, AD18, ASB15, AkAA14, ASH15, ASS15a, AER19, AAOA23, AC13, AS12e, AMJ16, BKK15, BH16, BS17, Bej16, BS21b, CAL^{+}14, CKT10, DR14, DS15, DC22, Eba13, GB15, GK13, GK15b, HS11, HNW17, HB12b, ITK21, JG11, Jag15b, JH18, Jay15, KV15, KM15c, KTBK12, KMS14, KAFS15, KKK13, KM15e, KB16c, LR11, MR11, MSD12, MJ12, MEKS15, ML13, NL15, NJ16, NAK^{+}14, NM19, ONJ13, PPP16a, Pet23, PSM15, PAH17, RRA17, RMAM19, RS11, RS14b, RRB16, RPS12, SP18a, SN22, SP13, SY12, SYAS11, SU18, SAW16, SAAM17, SG14, SC14a, SN14b, SAI16, SA15d, SG15, SNJ16, SAR16b, SK14b, SS12g, SCK17, SKGP12, TSE17, TCS14, TD16a, VGP15, VR11, Wis13, You13, ZEE15, Mug24, VJR11]. Technologies [HMMN13, MMG16, PS15a, PV17, SAR16b, SK15d]. Technology [AA18c, ASS17, am12a, AnAA14, ASS15b, Ari23, BLL13, CS15b, HZLL15, KKS14, Kum14a, LAA^{+}17, NRRM12, OAA24, PK12a, Pra14, PA23b, RKBB14, SP18b, SKG17, SLU14, SGA13, SMK17, SAH12, SB18a, SS23a, TOD13, TV10, DRJ14]. Teed [YWT12]. Telecom [TTA10]. Telecommunication [RHU12]. Telecommunications [LLSL12]. Telecoms [KHK13]. Telegram [APS18]. Telemedical [TAMA13]. Telemedicine [SL14, US14, WNE21]. Teleoperation [Muk13, TSA12]. Telescope [ESH^{+}14]. Television [LOAO23, Rab14, RT20]. Telugu [Gum19]. Temperature [ASD19, Afr17, AM17b, BLL13, DJ12, HD14, IM15, Job17a, KAA13, RSK15a]. Temperatures [AS18c]. Template [CS15a, DSJ11b, KC14a]. Temporal [AM19b, AMA14b, Far14a, FAR16, MN16, MH13b, PP17b, SK10a, LABV21]. Temporary [KP10a]. Tendency [GK12b]. Tenders [GDS21]. Tensile [RAC12]. Tentative [KNS20]. Tera [SHS23]. Term [AlS23a, Elc11, JPDB14, PS13, QKRB11, SP15a, SMN^{+}22, SAA24]. Terminal [AR16]. Terminological [TERCM15]. Terms [Ans14, OMH11, KK13c, Kum14a]. Ternary [AAGM15, KTBK12]. Terrained [MNB^{+}22]. Terrestrial [RKB19, VjS12]. Territories [MOM17]. Terrorists [Cur20]. Tertiary [ED19]. Test [AYAB18, BM14a, DGPP15, JVJ12, KP13, KN17b, KNS17, KJ12b, MMHP17, Raf15b, RHS18, RAC12, RGS15, SFA14, SCD10, Sm13b]. Test-bed [DGPP15]. Testing [AZ16, Ali12, BM15a, EADEA17, EsTeEh^{+}13, GENERM16, GVW24, HAH15, Jay15, KHSE12, Kas17, KA21, MMHA11, Mal13b, MKT17, MK18b, ML13, PP12b, RHS18, RPS12, SSO14, SMM11, SAM14b, JHP12, JVJ12]. Tests [AER19, AKS16]. Text [Ae14a, AMA14a, AlS23b, AAA16, BEE13, BG16, Bas12b, BSM^{+}12, DD20, DB12a, Elc11, GMM21, GJ13b, GEG23, Gum19, HJT14, JHB17, KM15c, Kha11, KWM16, MC17a, MJ15, MAE16b, OO21, PS21, PS13, PG13b, PY19, RS11, SBKS16, SPL15, SK13b, SGT^{+}11, SP15e, SK14b, SS12g, TB13, VS10, Ver17, EEE14, EA18]. Text-independent [JHB17]. Text-to-Image [AMA14a]. Texts [AK19]. Textual [Far14a, IHD12]. Textural [BMM13, PCC12]. Texture [ADZ11, AHBH11, DJG12, GPWS18, HAAA15, KTBK12, KS16e, KB12b, MGB11, PP13d, PG13c, RKR17, RC15, RKR11, SRS12, TS12, TBKA13]. Texture-based [KS16e]. Textured [ISH12b]. TF [AlS23a, Kam17, KSS18]. TF-IDF [AlS23a, Kam17, KSS18]. TFIDF [Ae14a]. TFR [BBL13]. Thailand [TTA10]. THD [MH12b]. their [AR17b, AYS19, GEG23, GS11, JFJN20, KKK13, LSK17, RMS14a]. Theme [KMT^{+}24]. Theorem [LAZ17, STM13]. Theorems [ES15c]. theoretic [MB12]. Theoretical [AKA^{+}18, BBZG^{+}17, HHEY19, MG14c]. Theory [AB14a, AYM16, Apo19, DM10, EG15, JM14, KM13a, MABE16, Pil13, RSA16, RKGK16, RGJ11, ST11, SL15, TRSK15, TS16, XCY11]. Therapy [AAA16, TD21]. there [GM15b]. Thereof [GCS14]. Thermal [AKSEB15, BH17a, DSP^{+}21, EH14, HD14, Haz14, HGK15, Ibr17, ILG15, KJA15, PN10, RGS15, SMSC14, SH17, SMHS10, SK15b, TH15, WBS16]. Thermally [HR13, ILG15, SJK21]. Thermo [EN13]. Thermo-Hydraulic [EN13]. Thermoelectric [Joh14]. Thermography [KMG21]. Thin [BI13, BAIs^{+}14, RRB12]. Things [ANAH20, MMN18, MSA18, PMBP20, PC15b, PA23b, SZM18, Zar24].

Thinking [SGKS12]. Thinning [CAL^{+}14, KEV14]. Thiourea [KJA15, PN10]. Third [NI14]. Third-Party [NI14]. Thomas [KUM13a]. Thorough [SC14a]. THR [IEL18]. Thread [DL23, NRR^{+}10]. Thread-Specific [DL23]. Threading [DL23]. Threads [Akg13]. Threats [ATB12, GCS14, PMBP20, SG16b]. Three [Abd14b, AYM16, AE12c, BFA13, BWS16, Ena12, FS11, GBPTEQ^{+}13, GSA12b, KMG21, KRR15, LEE13, Mis14, PS16b, Ras16, Ras17, SKM16, SM17a, SRSV14, SBM^{+}15, SYR13, SPSR12, VSS16, Wu23a, GSCT12]. Three-component [SPSR12]. Three-Echelon [FS11]. Three-Level [AE12c]. Three-Node [SRSV14]. Three-Phase [KMG21, PS16b, SM17a]. three-tank [GSCT12]. Threshold [DRS13, FYM11, HBS12, KJA15, LC17a, PDR11, RAA14, RB15, SSD^{+}13, SABA12, STD14, SS17c]. Thresholding [CT13, BG16, BL15, DPT14, İHŞ17, Jas13, KJK12, KK15c, MRR21, SD16, SS13e, TM18]. Throughput [AIK11, CKA18, DDC12, KA12b, MAM19, PPU11]. Thruster [FAK14, SZ15]. Thunderstorms [JIF12]. Thyristor [KR10, KRM12, PMK12, SSAZ13]. Ticketing [KSBM17]. Tie [PMK12]. Tier [HIS15, NSS15, DM14, HKU19]. Tifinagh [OMF^{+}14]. Tiles [ISH12b]. Tiller [PPR^{+}12]. Tilt [AN16b]. Time [AOO20, ASNS10, AMAY20, AR16, ADS16, AS13a, AK15, AAC13, AAR14, AAaAD12, AMTM11, AG13, AAY17, BLA12, BS14b, BR13, CM13, CKT10, DMM14b, DGPP15, EBAS11, Es12a, EJC16, EHHM17, FEE13, GK13, GK15c, HYA12, HKR17, ISM^{+}17, Isl18, IM16, JD15, JND15, KS15a, KK13a, KSS18, KBN18, Ker13, Key12, KDC12, KD22, KC15a, KG15c, LJ16, LRL17, MHE15, MG12, Mal13c, MS11, MSM15, MGA20, MK18b, MBZA16, MJM17, MNR14, MAM16, MA13b, NVV16, NKR14, NS13b, Neg13, PG13a, PKB13, PC15b, PMA13, PDK12, QT15, Raf14, RP14, SK13a, SCH21, SPBP17, SA12c, SS10e, SJV13, SSA16, SP18c, TR15, TCS19, TBA12, TG18, YRMB15, YAE11, Alz12, BAA24, EhMA13, JSS^{+}17b, LB11, MN24, NS24, SPP13, SGD^{+}20, THEE11, VK11]. Time-Efficient [AAR14]. Time-Series [KK13a]. Time-Sharing [AAC13]. Time-Triggered [TCS19]. Timekeeping [AnAA14]. Timeline [LSS20]. Times [KM12, RGCGCMFJ13]. Timetable [ESZED15, OOE18, TNKS15]. Timetabling [DRFF15, SB12c]. Timing [AKK11]. Tiny [Obe23]. TiO2 [AHZ15b]. Tissue [AVH10, SCK17]. Tissues [AS18c, KS14a, PSM15]. Titanium [SPP16]. TLBO [LGZX17, RB14b]. TLM [ZKN10]. TLM-2.0 [ZKN10]. TMHF [ARZ11]. TmSS [WNE21]. TOA [Cai18]. Tobago [Nor13]. Toddlers [SH23]. TODV [MS11]. TOFD [RVKB12]. Tofu [PMP14]. Together [IhZ12]. Token [ASY17, AS24, CS15c, sNIALY13, LR12]. Token-Based [AS24]. Tokenization [CJK17]. Tokens [AAS24]. Tolerance [ASS15a, AT14, FR12, KRSB12, SK18]. Tolerant [BK10a, DBB12, HZK12, JM15, NEA17, OJR11, Rah14, MP17, RB16, Sar13, STE14, SRC12, SC14b]. Toll [KAA20, VSS16]. Tomasulo [KDB17]. Tombariri [KDT16]. Tomography [NMK14, NMK15, RVGP20, Sam13, RPMH24]. Tool [AHAPT15, Ba13, CGKP13, DD21, ENE13, Fad18, FAA13, GS14, JARM13, JVF11, Kau15, MKM17, MD14a, NF15, NS23, OCZ13, RSS^{+}11, RDD^{+}12, VSA15, ZAER12, HS15c]. Toolkit [AbAAw12]. Tools [DFGG15, ES15b, JKS^{+}12, KC20, KA21, NMAB14, RRB16, SMM11, SSP13b, SBC12]. Tooth [PBVD22, BD10b]. Top [BRS16]. Top- [BRS16]. Topic [AN16a, CS16b, OFE15]. Topics [AEk13, IM16]. Topological [MP15, ND15, TR13]. Topologies [ÇGDT18, HS15b]. Topology [AAA14a, Alt13, BT14a, MNHB12, Mus13, MNN10, PPM12, TMS15, DB13b]. TOPSIS [eFKS13, IR16]. TORA [LS13, RSRT20]. Toroidal [Sal15]. Torque [BFA13, BBC12a, BBL13, Ena12, SV12a]. Torque-Speed [Ena12]. Torus [Ela11]. Torvik [ŞAG17]. TOSSIM [DGP14]. Total [AD15, Alg19, ANAH20, KDBT12, LZW15, SSK11, TY15a, BS10]. Touch [HS15c, PB13]. Tourism [TERCM15]. Tourist [Cai17, PPR17]. Tournament [GG11]. TPC [TM13]. TPC-IRBM [TM13]. TPL [HBS12]. TPS [AMN15a]. Traceback [KAR10]. Tracing [AKSEB15]. Track [GKK13, SB13b]. Tracker [MAK13, SP15e, SVB23]. Tracking [ATB12, AIG12, AH13c, BD10a, BNS15, COWK14, EAEE15, EHHM17, EBO13, EB13, FML20, GBPTEQ^{+}13, HMC12b, Key12, Kum10, MKK13, MBE16, MCaC14, OC20, SA15a, SAY16, Sha13, SC13, SVB23, TX20, TL16, TN14b, YYS^{+}23b, YMR^{+}12, Abd18, Alz12]. Traction [KTN12]. Trade [TS14]. Trade-off [TS14]. Tradeoff [HITA20, SH15]. Traditional [GK21, KM16b, SAP16]. Traffic [AICN14, ASII11, AzSKM12, AMBM11, AAoS14, AEKA11, AYM16, AH12c, ASA14, AK15, AIK11, BAO16, DB17, HZS14, JMM21, KP13, KCS14, KA17, LcCm13, MKN12, NU18, OOJ17, PZN21, RN24, SG13b, SG12a, SKG14, SP18c, THEE11, WAmMS14, WBL^{+}23, CS12]. Traffic-Signal [AAoS14]. Trafficking [PKJ14]. Train [Ibr16]. Trained [DCH16, Ela12, MMR14c]. Training [AO13a, AB13c, CHG12, DH19, EEAE17, IEME12, KSS14a, SSP13a, Sud10, TMP11, ZZW24]. Trait [Mar12]. Traits [NU12]. Trajectory [AMN15a, EHHM17, MBE16, RBM19]. Transaction [ASS17, DP11, RD17b, SYP10, UI16]. Transactional [AG14]. Transactions [RD17b, VMA23]. Transceiver [CPSP13]. Transconductance [LGM12b]. Transfer [AHZ15b, Anh15b, BM13c, BWS16, FS11, FDG17, HIU11, HAT16, Haz14, HY13, HR13, HJ13, KSA20, MCaC14, SM17b, SJK21, SK15b, TH15, TRM23, KM16a]. Transform [AAEK18, AYAB18, AMHAh13, AB14c, Aa13c, AAaNJ13, AHM14b, CT13, AE12c, AA21, CCG19, CT10, DR12, DDG16, Gaa16, GB18, HAAT16, HH13, JA11, JNR19, JKC11, KS16b, KK12a, KP15b, MKA17, MH11, NA20, NK19, NP11, NAK^{+}14, PKP13, PGS13, RT13, RAE15, RG21, RDE^{+}15, SS13c, SKT14, SGD13, SS12f, Sta22, SN16, SR17, TIA15, TBKA13, TK15, YB15, ZvAS12, LZW15, SGD13]. Transformation [AAO13, BM13b, KC14a, KP15a, KLS13, KK10c, MN16, NVV16, OLT14, PWG15, VGS11, SAM14a]. Transformed [SJV13, SAk13b, Tes15, YT15]. Transformer [KJKM21, SC15c, SBC12, SK12c, YA18]. Transformers [II16, KMG21].

Transforming [HKS12]. Transforms [AHS15, KTBK12, PP13b, SSJ12, YT15]. Transgender [BV21]. Transient [EN13, GA15, HD14, Nim15, RSSV14, RRR11]. Transistors [CN12, SSD^{+}13]. Transition [BK10a, SE14]. Transitions [HW18]. Translation [EEE14, EA18, EE16a, GK10, HPM20, MH13a, PFS12, SM23]. Transliteration [DDD12, HFSM12]. Transmission [ASD16, AJC14, AAOA23, BR11a, BOS15, DPS21, ES13, EAa14, FSG17, FIS16, GGI14, GB14, HB12a, Mah16, MSS13, NmEM13, PDK12, SBM^{+}15, SVB13, YWT12, YYS^{+}23b]. Transmit [ARZ11, JBK13, SZS15, TR15]. Transmitter [HZBG13]. Transparent [Ela11]. Transport [BG20, MBL15b, SK15d]. Transportation [Abd14a, ASR15, Nad17, Ngu24b, OESKF16, Sal18, TUS20]. Transpose [IMM12]. Transradial [RRA17]. Transverse [HS14]. Trapezoidal [PKJ14]. Travel [MP12b]. Traveler [ANA11]. Traveling [CW23, OJ11]. Travellers [MBL15b]. Travelling [AS14c, OON17, SGS15]. Treatment [ATA19, DPB14, El 13, Jag15b, KW16, LSM16]. Tree [Alb15, Bad13b, Bef18, EE13a, HS24, JSD14, Mee12, NS15a, PS15c, PA19, RB14a, RACI24, SN15a, SS16a, ST11, SVJ14, TMS15, TK15, NMS10, PN12, RBT10, Tri10]. tree- [NMS10]. Trees [Abb12b, ABF14, XL21, YK12]. Trellis [SS10e, Tha11]. Trend [Lah11, MMTS13, NMS10, SVRB15]. Trending [IM16]. Trends [DD20, Liu24, PS15a, SMGS17, SO14c]. Tri [GK15c]. Tri-Axial [GK15c]. Triangle [AB14c]. Triangular [BWS16]. Tribes [HSP10]. Tricorn [DLN12]. Tricycle [PR15]. Triggered [TCS19]. Trinidad [Nor13]. Trinity [KM15d]. Triple [GR14, SFH17]. TRIPLE-DES [SFH17]. Trivial [Par11]. TRMM [AAqA14]. TRS [SRF12]. TRSGM [DP11]. Trucking [TUS20]. True [MRP15]. True-Motion [MRP15]. Truncated [RRR17]. Truncation [KTDG12]. Trust [DS10, Far13, GJ15b, KM16b, KSlAM12, KEA19, Liu24, RS13d, SN15b, SN15c, SC14b, TOO16, Ven14b]. Trust-based [GJ15b]. Trustable [AAYM13]. Trusted [EzHEgT13]. Trustworhty [OAA14]. Trustworthiness [KEA19]. TSK [RBB11]. TSP [JSY14]. Tsunami [SGS18]. TSV [JM15]. TSV-based [JM15]. TTCM [AhAT13, Tha14]. TTCM-assisted [Tha14]. TTPS [Nor13]. TTS [MBE10]. Tube [CKS13, DJ12, IM15]. Tuberculosis [ATA19]. Tubes [EAMD14]. Tumor [AVH10, ESAA18, GJ15a, KKR^{+}14, RV16, Sha16, ZSR14]. Tumors [ASH15, MPR14, PS10]. Tumour [NJ16]. Tuned [AmH13, SAY16, SM12b, AMA15]. Tuning [AS17, Al 16, AS13a, BM13c, Eja12, IAH14, LM20, LGZX17, MBZA16, Ngu16, RKY12, SK12a, AA12c]. Tunnel [NSM14, Raf15b]. Tunnelling [RHD12]. Tunnels [Raf15a]. Tuples [BM16a]. Turbine [AB14b, RHM15, SOT18]. Turbines [BC12]. Turbo [PDK12, VK11]. Turbogenerator [TArM14]. Turbulences [CSP14]. Turbulent [SA15d]. Turning [HZ15]. Tutorial [ENE14]. Tutoring [CP15b, KA14]. TV [JM14, SCV13]. TVAC [Par14a]. TVAC-PSO [Par14a]. Twenty [NSM13]. Twenty-Five [NSM13]. Twin [JAB22]. Twitrends [IM16]. Twitter [IM16, SPBP17, SB18b]. Two [AJC14, L13, BSZ13, BB13a, Eba13, Es12a, EAMD14, FS11, HDMK14, HIS15, wHsYwT12, Ish12a, Job17b, KSSM24, LGM12b, LRRR11, MAA13, MHE15, MG17a, MAE16b, MRM13, MIAEFS15, MBR16, MG14c, Nim15, NMK14, PBK14, RAE15, RRR11, SK15a, SD16, SRF12, SRSV14, SSK11, SAK13a, SB18b, TYSS15, TM13, TArM14, VR11, YRV17, Kum14b]. Two-Area [SK15a]. Two-dimensional [MRM13]. Two-Equation [Nim15]. Two-layer [EAMD14]. Two-Phase [wHsYwT12]. Two-plants [HDMK14]. Two-Ring [BSZ13]. Two-Stage [LGM12b]. Two-Step [SD16]. Two-Tier [HIS15]. Two-Unit [MHE15]. Two-Warehouse [HDMK14]. TX [SVB23]. tXSchema [SP15a]. Type [ACK12, AKS16, BNS17, BIAE24, EM23, FAA13, Kum13b, MAaM14, SJV13, KB14a, Wu21, AHH20, GC18, HJBJ11, HSK14, LK15, RBB11, SK18, SS12e]. Type-1 [HSK14, RBB11]. Type-2 [HJBJ11, HSK14, LK15, RBB11, SK18, SS12e]. Type-I [AHH20]. Types [Ara15, Es12a, HB14b, Job17b, KC15a, KK22, MAA13, MAI15, TIT14]. Typesetting [hKhHL13].

U [PB16, TRDB24]. U-Slot [PB16, TRDB24]. UA [DGH16, DDL16]. Uang [TUS20]. UAV [KK20]. UAVs [ABL17]. Ubiquitous [KM13a, MB13a]. UCD [CGKP13]. Uda [SKS13, SS13a]. UDDI [PG11]. UDP [SG12a]. UDP/CBR [SG12a]. UHF [RT20]. Ultimate [IFT14]. Ultra [MFE13, NRP12, VRS10, VMT13]. Ultrasonic [BI13, Cai18]. Ultrasound [BK17, HS11, Has14, MPR14, MRR21, MSD12, RVKB12, RV12a, SN10b]. UML [KNS17, MMHA11, OLT14, VS12]. Un-Correlated [EE13b]. Unbounded [Nad14]. Uncertain [CWC15, CZW15a, CZW15b, SDZ15]. Uncertainties [Nad14, HZ14]. Uncertainty [Ahm20, AMA14b, DA12b, Fen18]. Unconstrained [CCG12, JND15, PK19]. Undecimated [RT13]. Under-Graduates [urPP13]. Undergraduate [AAEA16]. Undergraduates [MJP15]. undersampling [RPMH24]. Understanding [AI13b, HHS15, RKR10]. Understudy [MH13a]. Underwater [BKK15, KMMM12b, NRL17]. Undistortion [IUN10]. Unguided [JTS15]. Uni [HYA12]. Uni-Processor [HYA12]. Unicode [ASCR11]. Unified [AKB16, KSlAM12, SDZ15]. UnifiedSTPF [KSlAM12]. Uniform [BH17a, JVF11, Job17a, KAP15, KSZK15, JTKC14, SE12a]. Uniformity [BCO13, SMSC14]. Uniformity-Aware [SMSC14]. Uniformly [AA16a]. Uninterrupted [WBK22]. Union [MB12, TS16]. Union-closed [MB12]. Unique [AAaN15]. Uniqueness [PVA13, SKA13]. Unit [AB12a, AAS13, Bou13, GSN14, MAA13, MHE15, SAS21, TSS12]. Units [FKT12, GSB12, SRTB11]. Unity [JSD14]. Univariate [AKS16]. Universal [AH13a, SN14a]. Universities [ATAAAF19, Alt11, KM13b, SKS10]. University [GAA17, Nad17, NWS15, OOE18, SS14d, SB12c, UDA13, Yak18, YBM16, Ana15, OAG17, SU20]. Unknown [AYAB18, Alz12, AAY17, HZ14]. Unmanaged [Kar11]. Unmanned [ABL17, KN17a, MOM17, SA15a]. Unpredictable [GM16]. Unrealized [BD13]. Unrelated [GN14].

Unreliable [KS18, KC15a]. Unsteady [BH17a, CDC17, DGJ12, HIU11, ILG15, SDJ13, TH15]. Unstructured [AaEeK11, MA22, NP19, RS12]. Unsupervised [AkC13, BL15, Cur20, GP12, KM15d, MMR14b, PRR15, VM12a]. Up* [RA10a]. Up*/Down* [RA10a]. UP6 [Ngu18, Ngu23]. update [KT10]. Updated [IOO20]. Updating [FYM11]. UPFC [AD13, PC11a]. Upgrade [PCS18]. UPH [OFE15]. Uplink [KDTD11, Uzo13]. upon [MB13b]. Upper [MB15c]. UPQC [AKB16]. Urban [BIE^{+}12, BG20, aKNP^{+}15, OE14]. Urdu [AJI15, BSM^{+}12, MH13a]. URLs [HK10]. Usability [KEA19, NAH11]. Usage [Aly13, KBN18, MB15a, RKR10, SsTM12, SR15a, SSP13b, SV12b, YRV14]. USB [BA12b]. Use [ASS17, AGG16, BA12a, DA12a, EE16b, MOOOOMT18, MM12b, MBMA11, Qte14, SGD^{+}20, VV18, CHG12, KM13b]. Used [BS17, EB21, Mis14, PS15a, BD13, GG11, HK18, LGM12b, PSS12, SJA13, SAk13b]. User [AAA11, AWK19, ATM18, BCBS14, BBD13, BA12c, Cuf14, EF15, ETG17, GE12, GRC13, HG13, JHP12, LHKR16, LHM15, MTD14, NNR16, Pra14, Ran13, RKR10, RW20, SJ12, SSC15a, TGL13, USR17, Vad20, ZRP16, SS17c]. User-centered [AAA11]. Users [Far24, GK12b, HHEY19, HHS14, TE16]. Uses [MKAR22, SP15a]. Using [AlS24d, AlS24g, ASH15, AI18, BNS17, BBC12a, BD10b, CW23, CN10, DM10, Eg13, EsGK14, EAE14, GC24, GK10, HB12a, HSP10, HH13, HBA13, KAR10, KGK10, KR10, KNT11, KNV10, Las10, LNM24, MS10a, NBR10, NK14a, ONK24, OAA24, OA12, PS10, RN24, RGR15, RSA11, Roc17, S24, SMHS10, SSK11, SMK17, SEA16, SAM14a, SA12e, SS10f, SV10, TV10, VN10, VR11, VD10, Zar24, AAP13, AN14, AB14a, AN16a, A17, ASD19, AAN^{+}19, AM17c, AA18c, ACB17, Abb12b, Abb12a, AmH13, AS17, ADZ11, AAN12, AMSS13, AM11, ADH17, AARA19, AAY21, ASS17, ARAA11, AA13a, Abd14b, AH12b, AK18a, ATM12, ALO17, Ae14a, AZM13, AB13a, AB14b, AS14b, HAPG17, AMN15b, AMNA18, AYE13, AZA23, AD18, AA17a, AHA14, Agg17, ASN11, AS12b]. using [ABA^{+}10, AA12b, AM13a, AAEEA15, ASC15, AM17a, ARMS16, ATA19, AM19b, AICN14, AU23, Akg13, ABF14, AH17a, ASB15, AMA15, AY18, AkAA14, AZS16, Al 14, AIC19, AA18a, AlS23a, AlS23b, AlS24c, AlS24a, AlS24b, AlS24e, AlS24f, AlS24h, AlS24i, AS14c, AR17a, AS12c, Alb15, AAA17a, AJ16, AR17b, AA12d, AW20, AA20, ATB12, AAA14b, AAqA14, ASH16b, ANAH20, AHH14, AAoS14, AS16, ASM13, AASAA15, AKG20, AA11a, AAG13, AOSM12, AEKA11, AP16, AA12g, AB14d, AS13a, AV10, AMB11, AYAI17a, AAGM15, AHAT16, AHM14b, AKS^{+}13, AG14, ASK17, AAMA20, CT13, AKB16, AAaAD12, Asi15, ASI17, ur13, ARDS18, AP15, AS12e, AZMP17, AMH11, AE12c, AA21, AMJ16, AAS13, LEO15, MA22, BSAb12, BCC13, BR11a, BJ12, Bad13b, BNMM21, BLL13, BM15b, BCBS14, BGS11]. using [BG16, BI12, Bar16, BS12b, BSS15, Bas12b, BRD15, Baz18, BM16a, BKM12, Bej16, BIE^{+}12, BBD13, Ber11, BK17, BG13, Bha17, Bha10, BHC12, BM13c, BK10a, BYT14, BDN13, BSP11, BV21, BWH15, BP15, BS21b, BBL13, BOS15, BA15b, BIAE24, BMRA12, BSM^{+}12, CHW16, CP15a, CA11, CP12, CTG12, CJK^{+}12, MRTB16, CCJ14, COWK14, CJK16, Cha12, CT14, CM15a, CPSP13, CV13, CMKS15, Cha15, CS11, CT10, CB16, CS15b, DK13, DK11, DD15, DI10, DH17, DPS21, DSD15, DP15b, DSS18, De 21, DLR19, Deb20, DRS13, DH12, DL13, DPJB12, DB17, DS19, DB10, DVS13, DRS15, DUP13, DR15b, DGPP15, DDL16, DD21, DS13, Dix11, DM17, DTS16, Dum20, DTO23, DPB14, DGP14, DV14, ES13, EAJ^{+}15, EBAS11, Eis13, Eis14, ED18, EHSA23, sMTEkEdAr12]. using [EmErB13, EBR12, EAEE13, EAEE14, ElR16, EEA19, EEGHS15, EMA11, Elm22, EM23, EE13a, EB13, ESE16, EOO22, EG15, FJM16, FML20, Far14b, FLM11, FEbO13, FAK14, FSNR18, FSA12, Fer14, FK15, FM20, FPS16, FR14, FH13, FPA13, GD13, GM20, GPWS18, GJ15a, GC18, GSD14, GSA12a, GJ15b, GVRP10, GCP13, GDS21, GJP13, GS16a, GCK17, GRP13, GMM21, GEB12, Gha16, Gha13, GMAR14, GEG23, GB18, GPS14, GRS22, GRL20, GGK14, GDS16, GS11, GRC13, GGM23, GPR14, GK15c, GK21, Hab13a, HAA16, HO15a, HO15b, HB13, HKE^{+}14, HF20, HTM14, HS14, Hem22, HPM20, HR21, HPR17, HS24, Hos12, HKF12, Hus16, HSS11, HMAA13, Hus15, HMMN13, VNM15, Ibr14, Ibr17, Ibr18, IEG14, Ibr16, IEL18, IEME12, IHH11, IUN10, II16, IR16]. using [ITK21, IPR15, Ish12a, ISH12b, IHB^{+}15, IM15, IAA13, JM14, Jab14, JV14, JSSA20, JT15, JPJ18, JNR19, JD15, JND15, JH18, JP15, Jay15, JLCV16, JYP10, JA22, JBT23, JARM13, JVF11, JK14b, JKC11, Job17b, JP16, JGZ13, JI10b, JS15, JKTM15, JS18b, JS12d, KSRS12, KA16a, KKR18, KPC16, KSBM17, KV14, Kal11, KSKS23, KSP11, KK10a, KH20, KTS10, KUM13a, KRMB10, KRSB12, KBA10, KJ14, KK20, KV15, KKS14, KS14b, KS15d, Kau15, KM15c, KK16a, KV16b, KS16b, KS16a, KS16c, KTDG12, KSS12, KTBK12, KKGG12, KPG13, KMS14, KSS14a, KMMM12a, KMMM12b, KC14b, KEBF13, KSZK15, KH12, KN12b, KJK12, KSV10, KVK19, KZ16b, Kha20, KK17a, KDC12, KCS14, KKDC12, KJTM15, KKR^{+}14, KA21, KW14, KN12c, KKM18, KP15b, KJJ15, KK10b, KC12, KGAM10, KAAM11, KR11a, KPB12, KAS12]. using [KT12b, KRM12, KA13a, KPV13, KKK13, KA13b, KAS13, KS13b, KPR14, KAR15, KVKR16, KB16c, KS16e, Kum19, KNS20, KSSM24, KB12b, KKSS12, KS14c, KK17b, KR16, KR11b, KA12g, KK15c, KC15c, LGM12a, LsAAA13, Lah16, Lah11, LR11, LRRR11, LKO16, LAZ17, LA15, LSP17, LC17a, LCP21, LHM15, LP16, LG11, LKS14, LAA^{+}17, LGZX17, MKN12, MM13, MLRU12, MK13a, MDGS13, MMN18, MREd14, MZEE13, MAK13, MKK13, MH13a, MB15a, MC17b, MG14a, MB11a, MK18a, MSM15, MDD21, MP12a, MM18, MM14b, MM16, MBKH16, Mee12, MKAR22, MNB13b, Mes23, MAE16b, MSD12, MSP12, MSS14, MT13b, MM15c, MK15a, MEE15, ML12, MBMB12, Moh13a, MAk12, MIS12, MRM13, MD14b, MMR14c, MBZA16, MK15b, MAM16, MBE16, MV12b, MBL15a, MJS15, MMK21, MGL22, MR13b, Mos11, MNB13c]. using [Mug24, MBR16, MGB11, MHV^{+}14, MSR16, MRG19, NS15a, NA20, NKM11, NKH12, NS13a, NJ16, Nah13, NK15, NJAK15, NRP12, NM24, NC14, NAK^{+}14, NDC10, NSM13, NA18, NSS17, NmEM13, NP12, NLC12, Neg13, NP15, Ngu18, NMN15, Nim15, NRR12, NPA^{+}24, NF15, NS13c, NMK14, NHMR13, NM16, Nor13, NO18, NU18, OAA16, Obe23, OO18, OOO19, Ojh11, OP18, OEA18, OO21, OAO18, OJR11, OON17, OV13, OKS18, PP13b, PKP13, PG15a, PP12a, PN14, PT15, PP17a, PKB13, PC11a, Par11, PPM12, Par14a, PSA18, Pat14, PR16, PSR13, PP13c, PPP16b, PP17b, PS13, PVL11, PFS12, PBVD22, PH23, PC11b, Pet22, PRK17, PB15, PSG12, PB14, Poo22, PK16, PG13b, PY19, PPP22, PKJ14, PSM15, PP12b, Pri13, PWG15, Ps14b, PMD13, PG13c, PMP14]. using [RRVP12, RT13, RB14a, RR16, RKR17, RUB11, RBF12, Raf14, RÇ16, Rag15, RMS12, RVGP20, RHU12, RHS18, RKB19, RMAM19, RV12a, RC14, RP14, VGS11, RJS^{+}13, RK16, REA12, RKR10, RV13, RR18, RJC15, RMD10, RS14b, RSK15b, RKGK16, RGK17, RPMH24, RG21, RA10b, Rez12, RKS16, RKS17, RR22, Rok20, RL14, SE12a, MS12, SV12a, SPP13, SE14, DBS15, SOME16, SP16, SP18a, SV18a, SN22, SA15a, SZM18, SKM16, SDS20, STM13, SND13, SS16a, SO13, SCE14, SF15a, SS10b, SMAM15, SSA11, Sal19, SD16, SAA24, SCD10, SU18, SAS24, SAW16, SRTB11, Sar16a, SFH17, STS19, SP11a, SP15c, SAAM17, SS12c, SSS15a, SM18, SH15, SS21, SDG18, SS15a, Sel22, SR23, SGM^{+}11, SK16, SM10, SBR18, SGB16, SSS15b, SP15d, Sha16, SBHM13, SS17b, Sha18, SCL14, SSK13a, SR13b]. using [SS14e, SBS13, SABA12, SM15b, SQ18, SAKS24, SSR14, SJV13, SP11b, SF15b, SCV13, SK10c, SS12e, SK12b, SSS13c, SSG14, SRS14b, Sin14, SVRB15, SS15c, SD15, SNJ16, SKS16, SS16c, Sin16, SP18c, SJ14, SAH12, SV14b, SSK13b, SM19, SET15, ST14, SC16, SA12d, SR15b, SS12g, SU14, SRS12, Sri12, Sta22, SAM14b, Sud10, SGJ12, SB12b, SZS15, SVJ14, SLK^{+}13, SS17c, SMR16, SB12c, SKZ22, SVB23, TIA15, TSE17, TAM14a, TK14, TBN^{+}15, TL16, TKC^{+}13, TOMGBSS13, TSA12, TERCM15, TMS15, TV13, TMK^{+}10, TK15, TMP11, TD16a, TVY17, TVGT18, Tit16, TY15b, Tri10, TB15, TArM14, THAP19, UC15, USK17, UJ14, UDA13, UO23, Vad20, VR14, VK13, VK11, VMM14, VR15a, VM14a, VD15, Ver17, VKJ^{+}15, VV12, VST17, VS12, VDM22, WKG^{+}18, WDN^{+}20]. using [Win14, WWSS12, XL21, YWT12, YB15, YK16, YMA16, YJ14, YSH12, You18, ZYL^{+}21, ZYM13, ZRP16, Zyo12]. USSD [GPB22]. USSD-based [GPB22]. Utility [GPC16, KPS11, PN11]. Utilization [DD14, Jag15b, MDP12, RS11, WKG^{+}18, ZJR^{+}19]. Utilizing [BLQ21, JM14, MZ18]. Uttarakhand [VB10]. UWB [JD13, KVRS14, Mal10, RK15, TSCD15, VMT13]. UX [HG13, PW23].

V [HAA16, LG11]. V&V [SA12c]. V/ [HAA16]. Vacation [ASG10, KS18]. Vacations [AS13d]. Vacuum [AHH20]. Vague [PP12a]. Validated [LM20]. Validating [AHA14]. Validation [GPB22, JS18b, NHS11, NYS15, SJ12, VG12, YJ14]. Value [AAEK18, MRTB16, DMM14b, EBAS17, JT11, KVN10, KK12a, KAP11, KNI14, PKP13, RP15, RKGK16, SAK13a, TSE17, VR12, VR15b, YMA16].

Valued [GHH15, KB14a, LAZ17]. Values [KBK10, MRP16, SS17c, VG12]. Valve [MB15b]. VANET [AM15, Abd13, KK13c, PIB11, RD17a]. VANETS [DSD15, LS17, DS15, RMR14, SWP19]. Variability [KKK13, KR11b, RA10b, TD16b]. Variable [BH17a, HB13, HD14, Haz14, HGK15, ILG15, KSS13, LGM12a, MSP12, SH17, SSC11, SS17c, TH15, TRM23, VD15]. Variable-Bandwidth [VD15]. Variables [PDK14, SK12c]. Variance [AH13a, Hus16, KR11b]. Variant [Abd17, GJS17, SS10c, TMP11]. Variants [CN10, KK15a, OMH11, WAG11]. Variation [AS18c, HS15b, JSY14, LZW15, MSEEs13]. Variational [LAZ17]. Variations [ACK14]. Variety [Rus15]. Various [AR17b, AAGM15, ASB12, ARR20, BR11b, KMS14, KJK12, MSP12, PBG12, PP13d, PAH17, SAS13, SN14b, SCK17, TIT14, GJP13, SSC11]. Varying [AJSG11, BLA12, KKGG12, MS17, MAM16, PIB11, PDR11, STD14, ASNS10]. Vb.Net [Ade11]. VBA [DD21]. VBLAST [SB11c]. VCG [MK17]. VDRC [SPW15]. Vector [AH16, AA17a, AA18a, ASD16, BBC12a, BK13, DR12, DB17, EEA19, FKT12, Gho11, HAA16, IEME12, KSS12, KPG13, KSS14a, KJTM15, KK10b, KR11b, MJ15, MS11, NK19, NW13, OOO19, PNIS16, PG13c, SND13, SP11a, SK12b, Sud10, WKC11, Zyo12]. Vectors [Ms13, MN15]. Vedic [RJS^{+}13]. Vegetables [TM18]. Vehicle [AAA17a, CM13, DDB11, EAM14, FS11, GPR14, KSP11, KTS17, Mis14, MPJ18, NKM12, OOJ17, RN24, SA15a, SK13a]. Vehicle-Pedestrian [OOJ17]. Vehicles [AzSKM12, ABL17]. Vehicular [HTPB11, RMR14, RD17a, SWP19]. Vein [DM10]. Velocity [EB13, HZ15, Raf15a]. Ventilated [MBR16]. Ventilation [AM17b, Raf15b]. Ventricles [HSD13]. Ventricular [WWSS12]. Vents [MB15c]. Verb [EE16a]. Verifiable [De 21]. Verification [AD11, AVH10, AAR11, BHC12, GC24, LC17b, MR11, NAK^{+}14, RRVP12, RMAM19, RD17b, TGL13, UAA16, VV12, YA14]. Verify [DRBR24]. Verifying [BH14]. Version [ABD22, EEL11, MB12, SHR14]. Versioning [AIG12]. Versions [KK15b]. Versus [AJ18, MRG19, FKT12, Mal13a, MAE19, THEE11, UJ13]. Vertebrae [SR15c]. Vertical [CP14, FEE13, GDJ12, HD14, Haz14, ILG15, IFT14, KA12c, MM20b, MA10, MZ18, MB15c, SDJ13, TRM23]. Vertically [HGK15, SA15b, SM17b]. Very [MMP21]. Vessels [KK15c, RRB12]. VESTAL [SR15c]. VHDL [DNR15, KM15b, NK12, SA12e]. VHF [NAhA13]. via [AZ16, AsS13, AAA21, AAMG13, BA12b, DRFF15, EAS16, EBO13, Far14a, HAb14, KE16, Kum14a, LEE13, NSA14, NCL11, OSA14, PCS18, RACI24, SS12c, ZPAQP11]. Vibration [EH16, KM16a, OS16, SKM16, SYR17]. Video [Abd14a, AM19b, AsS13, AE12a, Alt13, AJC14, AHAT16, AK15, AIK11, MRTB16, DR12, DA18, ES13, EME13, EAEE15, Elm22, Far14a, FIS16, GP12, GRC13, HMC12b, JD15, JKS^{+}12, KPB12, KB15, LBNP12, LAM14, MKA17, MA12, MS14b, MBMB12, MNB13c, NA20, NC14, PB19, PM21, RMP11, RT20, SPP13, SR13a, SG13b, SSV14, Sub24, TK15, YT15, YYS^{+}23b, ZAER12].

Video-Centric [Sub24]. Video/Voice [AIK11]. Videos [PP13b]. Vienna [MA15]. View [Ibr18, IHH14, PA23a, TUS20, USC12, XRL14, She12]. View-based [IHH14]. ViewModel [PA23a]. Views [Dix11, VPGF10]. Vigenere [AS14d]. Vigilance [Kum12]. Villages [KDT16]. ViNET [AAAA21]. Viola [EAEE13, EAEE14]. Violence [Deb20]. Virgin [JAB22, MM20a, MM20b]. Virtex [ZTM11, DV14]. Virtual [AAO22, Ana15, AG13, CJK^{+}12, DE13, DD21, Dix11, HHS15, JHP12, JAB22, KAS13, MB15a, MK13d, PS16b, RK21, SATS13, SV10, USK17, YSH12, Tri10, YMA13]. Virtual-Lab [DD21]. Virtualization [KBBM15, MMHP17, SAM14b]. Virus [BML22]. Visco [CP14]. Visco-Elastic [CP14]. Viscoelastic [BB13a]. Viscosity [BH17a, HD14, Haz14, HGK15, ILG15, SH17, TH15]. Viscous [HD14, Haz14, HGK15, SDJ13]. Visible [Pil13]. Vision [ATH12, AYE13, AMS14, BA12b, DSP^{+}21, DMB21, Elm22, GK13, JD15, MBA20, SSO14, SG13b, SM19, TH14]. Vision-based [ATH12]. Visits [yGL13]. Visual [BBB16, BNA14, ESR13, FML20, FSNR18, IR10, JND15, JM16, Oua12, SA15a, SAY16, SSA14, TL16, Kar11]. Visualization [AJB13, AB16b, BMD16, HSS11, KK11, MTJ13, MM16, NHMR13, SSZ16, USK17]. Visualize [Wu23a]. Visualizing [ANM17]. Visually [SS23a]. Visushrink [CFB11]. VITERBI [PY19, DV14]. VLR [ARHERZ10]. VLSI [AC13, CKT10, Jay15, KKT15]. VM [SMS16]. Vocal [MK13c, Sak15, VKJ^{+}15]. Vocational [EAE14]. VoD [RMP11]. Voice [AIK11, CS15b, DA18, ESE16, KP13, MSM15, NNR16, OCZ13, RDD^{+}12, SU18, SC19b]. VOIP [ASA14, ARK13, BME19, MSP12, RBM19]. Voltage [ANE^{+}13, AKB16, ER12, GSB12, IK11, KC12, KR10, KRM12, MBL15a, MEE16, RRAFAM19, SMT19, SMVK11, SSD^{+}13, SK15c, TAH13, TArM14, TY16, KN12b]. Voltages [OAA13]. Volterra [BFD15, LEE13]. Volume [ANM17, DSM^{+}20, HS18, KV16b, SKG14]. Volunteers [MOOOOMT18]. Voting [De 21, KS15e, LC17a, LC17b, Okp17, SY15, Yak18, AAMA20]. Voting-based [LC17a, LC17b]. Vowels [NRR12]. VP9 [PB19]. VPRS [US15]. VQ [KSS14a, YB15]. vs [AD14a, AkC13, AB13c, AMK12, RGK17, SM14b]. VSC [AKB16]. VSM [GM20]. VTOL [SA15a]. Vulnerabilities [GSV20, Sul17]. Vulnerability [AZ16, AMY23, AKA^{+}18, IAA13].

W [KM12]. Wait [LK16a, MJM17]. Waiting [Neg13]. wake [GKG16]. Walfisch [AMM21]. Walk [BNMM21]. Walking [DRBR24]. Wall [HR13]. Wallet [BM14b]. Walsh [KSS12, TK15, YT15]. WAN [AICN14]. WAP [APK11]. Warehouse [BAF15, HDMK14, MC16, SD12, SEA16, Win14]. Warehouses [DMM14b, TS11]. Warm [Es12a, MAA13]. Warning [ARAA11]. Warnings [NC14]. Warping [MSM15]. Warshall [SM14b]. Washing [AERE21]. Wastage [CM13, MC16]. Waste [LSM16]. Water [AIRR19, AAZN21, EN13, KSZK15, KNV10, LSB17, MMK21, NA18, PTPS19, PP15, RÇ16, SDG18, SB18c, WKG^{+}18, YAT15a, TMC17]. Water-Tank [SB18c]. Watering [PA23b]. Watermark [ESKM20].

Watermarking [AQ15, AS13c, BM16a, CM12, DBB24, EAMAES16, ESKM20, EzHEgT13, EME13, Ela11, Gaa16, GPC16, HKF12, HB12b, JG11, KM15c, KK12a, KH12, KPB12, KVS16, LBNP12, MKA17, MA12, MIAEFS15, NTY19, NK19, NP11, NRV12, PP13b, PKP13, PVL11, RB14a, VGS11, RDE^{+}15, SA13a, SS15a, SGB14, SP10, TC16, YBA12]. Watershed [SNB18, SR17, SVMLb13, ZTM11]. Wave [AAZN21, GK15c, HSEA15, LSB17, Moh13b]. Waveforms [SR13b]. Waveguide [PB16]. Wavelength [AM14]. Wavelet [AAEK18, AR10, AMHAh13, AQ15, AJM^{+}15, AAaNJ13, AHM14b, AAO13, CT13, AAaN15, AE12c, AA21, Cha12, CCG19, DL19, DR14, DP15b, DR12, DSM14, DDG16, Gha13, HS12, JG11, JJN14, JT15, JA11, JNR19, KM17, KP15a, KA16b, KK12a, KTBK12, KH12, KJK12, KS16d, KKM18, KP15b, Lah16, LSK17, LA17, MKA17, MRR21, MD14b, MGB11, NK19, OJR11, PVL11, PWG15, PGS13, SNA15, SS13c, SCL14, SKT14, SK12b, SS13e, SErEs^{+}12, SSJ12, TIA15, TT14, TN14a, VDM22, WWSS12, YB15]. Wavelet-ANFIS [AJM^{+}15]. Wavelet-based [AAaN15, DL19]. Wavelet-Packets [Gha13]. Wavelets [Las10, RFAH14, DVS13]. Waves [BB12b, EAMD14, SET15]. Wavy [TRM23]. Way [AAT16, BBA12, DAB12, TFT11, BJ12]. Waypoint [MKN12]. WBANs [AKG20]. WBCT [KPB12]. WCDMA [KA12a, Uzo13]. WDM [ASK17, CSP14, CSN15, RK19]. WDM-PI [CSN15]. WDM-Polarization [CSP14]. WDM-PON [ASK17]. WE-MQS [NNR16]. Weakly [DAG16]. Weaknesses [MR13b]. Weapon [Agg13]. Wearable [Cai17, SSA^{+}23]. Weather [AIC19, BL15, EKES20, SB18a, Sta22, YK16]. Web [IHH14, LHM15, RL14, She12, SA13b, SRM12, Zar24, PWA12, MKT17, AaEeK11, ABYB22, AR15, AHA14, ABF14, AAO11, AS12c, ASN12, AMG^{+}12, AAS24, AS24, ASS15b, AIP18, ASI17, BH16, BN14, Bou12a, CHW16, CS14, Cha15, CS18, DBM14, DS13, DM17, CB12, EBB17, Elg13, ESB13, ES12b, ETD13, Far14b, Far17, FLM11, GS16b, GENERM16, GOCA15, HG11, HB16, HMMN13, JS18a, JVF11, KH20, KNA12, Kas17, KAqC14, KM15d, KN15, KP10b, KN11, KR11a, KA12g, MK13b, MB13a, MM14b, Mee12, MK18b, MM12b, MJP18, NG10, NKM12, NF15, Ojh11, OCZ13, PG11, PK15b, PSG12, PP12b, RKR10, RV13, RKGK16, RGK17, RO11, RS16, RBO19, SS12a, SJA13, SP15b, SJ12, SC15a, SA12b, SHR14, SGS18, SSP13b, SRM12, SB18a, SL15, SV12b, TGL13, TA10, UJ14]. Web [WT13, WDAJ15, Yac15, YMA13]. Web-Based [Far14b, ABYB22, AAO11, Bou12a, CHW16, CB12, Elg13, GOCA15, HB16, Kas17, MJP18, NF15, PSG12, SGS18, SB18a, TGL13, Yac15]. Web-Cloud-based [WDAJ15]. Web-Services [She12]. Website [Elg13, Far13, KEA19, MBL15b]. Websites [Bej16, Dak23, Has15, PB19]. Weed [AA12b, AA12c]. Weeding [VR11]. Week [SKG14]. Weibull [MC17b, RA18, TYSS15]. Weighing [PS13]. Weight [BSM^{+}12, VR15b]. Weighted [DD13, KRSB12, LHM15, LZW15, Obe23, OHI15, PMD13, DBS15, SS13e, SK14c, TOO16]. Weighting [BA23, GPWS18, PT13, Elc11].

Weil [SSS13b]. Weiner [ITK21]. Welfare [NHH10]. Well [TMC17]. Wells [BS11a, Naj12]. went [RK13]. Western [SKS10]. Westwood [HEs12]. WestwoodNew [HEs12]. Weyuker [MR13a]. Whale [JM18]. Wheel [Sin13]. Wheelchair [GRS22]. Wheelchairs [TVY17]. Wheeled [HM13, KNI14]. Wheels [SKM16]. Whip [PRK17]. Whip-Lash [PRK17]. Whispered [JHB17]. White [LOAO23]. WI [DNR15, ARP17, AZ16, BJS13, DV16, DSS13, EBO13]. WI-FI [DNR15, ARP17, AZ16, DV16, EBO13]. Wi-Max [DSS13]. Wi-TCP [BJS13]. Wide [AK18b, AICN14, BSM17, KK10c, MDGS13, MM17, MFE13, PS16a, SS12a, VMT13, ATA14]. Wide-Area [KK10c]. Wideband [FSG17, TRDB24]. Widgets [KN12c]. Width [MAI12, SE14, Suj14]. Wiener [BFD15, Sam13]. WiFi [FSG17, IRT11, THEE11, SR23]. Wiki [NK14b, NK14b]. WiMAX [AMN12, MFE13, VG16, CS16a, FM13, GDD13, Kum14a, MS17, PM13, RBM19, SKG17, SSG11, SKJ15, SB12a, TOD13, IRT11, DD14, SZUSNK19]. WiMax/WiFi [IRT11]. Wind [AB14b, BC12, DD20, HB13, JJN14, KMB14, RSB12, RHM15, Ras16, Ras17, SS12b, SS15c, SK15c, Tua17]. Windings [SMC17]. Window [CM13, GG11, IAH14, MAI12, MKT17, MSS13, SK13a, SN14b, RMS14a]. Windows [JT15, RTD10]. Windup [Mis14]. Wing [HA16]. Wired [AQ14b]. Wireframe [VPGF10]. Wireless [AII13, AAY21, AOAG13, AOO20, ASNS10, ABB20, AmM12, AsS13, AzSKM12, ARHERZ10, AKG20, AT15, CK13, AV13, BT14a, BS14a, BBHL12, BJPP15, BKS14, BJS13, CPSP13, CA16, DS10, DV16, DSA11, DAB12, DA18, DKS14, DGP14, EsGK14, EhMA13, ETLE13, EE16b, EG15, FKFU12, FEE13, FR12, FTLK14, GEA16, GSD14, GMRK12, GK12a, GMAR14, GPS14, GGI14, GDS16, HZS14, HS12, HR21, Hir12, sNIALY13, JI10a, JJKD14, JKS^{+}12, JYP10, JTKC14, JSK17, KA12c, KFEN21, KS16a, KMMM12a, KMMM12b, KAAA13, KNGT10, KCS14, KC11, KPV12, LGM12a, LGM12b, LCA12, LSS20, LA15, MM14a, MK13a, MFA13, Mal10, MG17a, MMP21, MP10, MK15a, MKCV13, MJS15, MMK21, MAM19, MCB12, MNHB12, MNN10, NS15a, NOO16, NM16, OC20, OHI15, PS15a, PT15, PVR20, PPU11, PK12b]. Wireless [PSS11, PS15d, PAH17, RS13a, RMVP14, RR11, RP14, RV12b, RB15, RR12, RKG10, DBS15, SE12b, SE16, Sal17, SOO15, SMJ12, SR23, SPP11, SA10, SK10b, SU12, SS12e, SSS12, SM13c, SM13d, SSS16, SKS16, SL11, SV14b, SS16d, SJV^{+}15, SA14, TBKG13, TBR15, TK14, TMC17, TB14, TCS14, TOO16, UR16, UO23, XRL14, YRV17, YK13, YYS^{+}23b, YYS^{+}23a, ZBI14, ZCDR13]. Wirelessly [IOO20]. Wirelessly-Updated [IOO20]. Wisdom [SA13b]. Wise [AB14c, HZS14]. within [BEE13, BBL15, RGNA13, ST12]. without [AAaN15, AMK12, AH13c, OAA13, SSA14, ZZ11]. WLAN [KG15b, MDM17, PS16a, SKJ15]. WLAN/WiMAX [SKJ15]. WOFDM [AMN12, AMN12]. Wolf [ALH^{+}24, YMAS17]. Wood [MV12b]. Word [AlS23b, BSM^{+}12, Eg13, SGT^{+}11, SG15, HLZ^{+}24]. Word2vec [EEA19].

Words [ASCR11, SSJ12]. Work [MD14a, WDN^{+}20, YBM16]. Workers [ONK24]. Workflow [AS11, GSA11, KAR15, RGNA13, TAP15]. Workflows [KKR18, SV18b]. Working [KS18]. Workload [GM16, HYA12, MC12, MM14b, RKGK16]. Workloads [BM13b]. Workspace [PKDS17]. Workstations [RD13]. World [El 15, KC14c, SJ14, VD15, Zar24, SS12a]. Worm [ARAA11, AA13a]. Worm Secure [KK12b]. Wormhole [AMY23, JSK17, KK12b, NS16, Par11, PC15b, RSM15, VR11]. WOT [TVS19]. Wound [İHŞ17]. WPAN [CS10]. WPD [Sar14]. WPSO [mH14b]. WRAN [MJM16]. Wrapping [RAC12]. Writer [RMAM19]. Writer-independent [RMAM19]. Writing [CKS13, GK12b, SS12f]. Written [PC11b]. wrong [RK13]. WSN [AKK11, JI10a, KVKR16, MNR14, NBL15, RK17, SK18, TSRM10, TY13a, Tha13, VM14a]. WSNs [AkAA14, ATB12, AV15, CC17, KP16, KBAQ15, LC17a, LC17b, SHHL13, SC16].

x [LA17, BLQ21, SSK11]. X-PRESS [BLQ21]. X-ray [SSK11]. XGFX [Bor19]. Xiaoping [SS14b]. Xilinx [DV14]. XMI [KV15]. XML [Asi15, SG13a, Win14]. XML-DOS [SG13a]. XMPP [Bor19]. XQuery [KPV12, AB11]. XSG [GC18]. Xylenol [KJA15].

Y-Splitter [RHU12]. YAA [AP16]. Yagi [SKS13, SS13a]. Yagi-Uda [SKS13]. Yanlin [SS14b]. Yaoundé [MDCS22]. Yaw [NVR12]. Yawn [ARDS18]. YCbCr [MA12]. Year [Add12]. Yield [RÇ16, SCH21]. Yields [AN16b]. Yogyakarta [HS24]. Yola [UI16]. YOLO [ABAO24, ONK24]. YOLOv4 [SKZ22]. Yoruba [AAA19]. Young [PNIS16]. YouTube [Far14a].

Z [JM15, AM17c]. Z-Domain [AM17c]. Z4 [JS12d]. Z4-linearity [JS12d]. ZAG [VP10]. Zernike [MEE15]. Zero [AAaAD12, SSV11, SSS13b, Liu24]. ZIG [VP10]. ZIG-ZAG [VP10]. ZigBee [ÇGDT18, DKS14, HS15d, CPSP13, NS15a]. Zinc [PN10]. ZIP [KC12]. Zn [PS12a]. Zone [CTG10, JC14, KBBM15, SI13, WDD16]. Zone-Aware [CTG10]. Zones [YRV17]. ZRP [MS10b]. ZTA [Liu24]. ZUC [Sul17].

References

A:2017:RFP

[A17] Vijayalakshmi A. Recognizing faces with partial occlusion using inpainting. International Journal of Computer Applications, 168(??):20–24, June 2017. CODEN ???? ISSN

               0975-8887. URL https://www.ijcaonline.org/archives/
               volume168/number13/27944-2017914570/.

A:2018:MCR

[A18] Moustapha B A. The Markov chain resulting from the states of the Bitcoin. International Journal of Computer Applications, 181(??):1–7, July 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number4/29701-2018917520/.

Alnaji:2011:NCA

[AA11a] Khaled W. Alnaji and Wesam M. Ashour. A novel clustering algorithm using K-means (CUK). International Journal of Computer Applications, 25(??):25–30, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume25/number1/2995-4025/.

Assogba:2011:FCE

[AA11b] M. Kokou Assogba and Amine Nait Ali. Fingerprint characteristic extraction by ridge orientation: an approach for a supervised contactless biometric system. International Journal of Computer Applications, 16(??):14–19, February 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume16/number6/2018-
2722/.

Abdelsalam:2012:DFS

[AA12a] Mohamed M. Abdelsalam and Marwa F. Areed. Decoupled fuzzy sliding mode control for a synchronous motor speed control. International Journal of Computer Applications, 47(??):29–35, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume47/number11/7233-0190/.

Ahmed:2012:OPS

[AA12b] Ashik Ahmed and B. M. Ruhul Amin. Optimization of power system stabilizer for multi-machine power system using invasive weed optimization algorithm. International Journal of Computer Applications, 39(??):28– 34, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume39/
number7/4834-7092/.

Ahmed:2012:PCI

[AA12c] Ashik Ahmed and B. M. Ruhul Amin. Performance comparison of invasive weed optimization and particle swarm optimization algorithm for the tuning of power system stabilizer in multi-machine power system. International Journal of Computer Applications, 41(??):29–36, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume41/number16/5626-7943/.

Alhanjouri:2012:NMQ

[AA12d] Mohammed Alhanjouri and Ayman M. Al Derawi. A new method of query over encrypted data in database using hash map. International Journal of Computer Applications, 41 (??):46–51, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume41/
number4/5533-7580/.

Ali:2012:PRI

[AA12e] Behloul Ali and Aksa Abla. PC FFIW: a robust image matching algorithm. International Journal of Computer Applications, 49(??):20–24, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume49/number21/7895-1230/.

Alwidian:2012:SCB

[AA12f] Sanaa A. Alwidian and Alaa N. Alslaity. A stable CoveringSet-based leader election algorithm (SCLEA) for mobile ad hoc distributed systems. International Journal of Computer Applications, 54(??):25–32, September 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume54/number2/8539-
2083/.

Anand:2012:EAD

[AA12g] Surbhi Anand and Rinkle Rani Aggarwal. An efficient algorithm for data cleaning of log file using file extensions. International Journal of Computer Applications, 48 (??):13–18, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number8/7367-0097/.

Abdulrasool:2013:RCS

[AA13a] A. S. Abdulrasool and S. M. Abbas. Reservoir computing: Size and connectivity optimization using the “Worm algorithm”. International Journal of Computer Applications, 69(??):18–22, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number4/11830-7532/.

Aghaei:2013:OAC

[AA13b] Ashkan Aghaei and Sassan Azadi. Optimizing azadi controller with COA. International Journal of Computer Applications, 61(??):22–26, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume61/number8/9949-4594/.

Al-azzawi:2013:RPF

[Aa13c] Nemir A. Al-azzawi. Rigid point feature registration for MRI brain images based on fuzzy transform. International Journal of Computer Applications, 71(??):1–6, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume71/number4/12343-8626/.

AlHammadi:2013:DME

[AA13d] Dina Abdulaziz Al Hammadi and Mehmet Sabih Aksoy. Data mining in education- an experimental study. International Journal of Computer Applications, 62(??):31– 34, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number15/10158-5035/.

Abbas:2014:QAB

[AA14a] Nizar Hadi Abbas and Haitham Saadoon Aftan. Quantum artificial bee colony algorithm for numerical function optimization. International Journal of Computer Applications, 93(??):28–33, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number9/16244-5800/.

Addah:2014:BCD

[AA14b] Jerome Addah and Stephen Apanga. Building clinical decision support systems for clinical practice in poor settings: Critical system features. International Journal of Computer Applications, 91(??):1–4, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume91/number1/15842-4718/.

Agbonifo:2014:FLS

[AA14c] Oluwatoyin Catherine Agbonifo and Olumide Sunday Adewale. A framework of e-learning system for students’ engagement with cognitive tasks to enhance learning performance. International Journal of Computer Applications, 98(??):1–7, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number2/17152-7203/.

Alyaseri:2014:PBH

[AA14d] Sana Alyaseri and Alaa Aljanaby. Population based heuristic approaches for grid job scheduling. International Journal of Computer Applications, 91(??):45–50, April 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume91/number5/15881-4853/.

Abimbola:2016:DAB

[AA16a] Raji A. Abimbola and Amusa K. Akinwale. Design and analysis of broadside arrays of uniformly spaced linear elements. International Journal of Computer Applications, 156(??):19–24, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume156/number6/26713-2016912477/.

Alzubaydi:2016:PNN

[AA16b] Dhia Alzubaydi and Shyma Akram Alrubaie. Probabilistic neural network with GLCM and statistical measurements for increasing accuracy of iris recognition system. International Journal of Computer Applications, 136(??):44–51, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume136/number12/24209-2016908628/.

Afroge:2017:HMR

[AA17a] Shyla Afroge and Boshir Ahmed. A hybrid model for recognizing handwritten Bangla characters using support vector machine. International Journal of Computer Applications, 174(??):41–46, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number1/28375-2017915312/.

Apau:2017:DIS

[AA17b] Richard Apau and Clement Adomako. Design of image steganography based on RSA algorithm and LSB insertion for Android smartphones. International Journal of Computer Applications, 164(??):13–22, April 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume164/number1/27446-2017913557/.

AlMehediHasan:2018:PSL

[AA18a] Md. Al Mehedi Hasan and Shamim Ahmad. predSucc-Site: Lysine succinylation sites prediction in proteins by using support vector machine and resolving data imbalance issue. International Journal of Computer Applications, 182(??):8–13, September 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number15/29937-2018917787/.

Alabajee:2018:EAL

[AA18b] Marrwa Abd-AlKareem Alabajee and Taghreed Riyadh Alreffaee. Exploring ant lion optimization algorithm to enhance the choice of an appropriate software reliability growth model. International Journal of Computer Applications, 182(??):1–8, July 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number4/29747-2018917499/.

ALhumayan:2018:SSP

[AA18c] Alhanoof Mohammed ALhumayan and Abdullah Saad AL-Malaise. Supporting strategic planning decision-making in higher education by developing IDSS using open source dashboard technology. International Journal of Computer Applications, 181(??):12–14, November 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number29/30123-2018918118/.

Alhashmi:2020:GIA

[AA20] Mohammad Alhashmi and Abdulaziz Alroomi. Graph isomorphism algorithm using pieces patching puzzle technique (ppp — technique). International Journal of Computer Applications, 176(??):1–7, April 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number12/31250-2020920044/.

Azeroual:2021:GIF

[AA21] Assma Azeroual and Karim Afdel. GPU implementation of Faber Schauder discrete wavelet transform using CUDA. International Journal of Computer Applications, 183(??):1–8, December 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number42/32208-2021921815/.

Akhigbe:2011:AMI

[AAA11] Bernard Ijesunor Akhigbe, Babajide Samuel Afolabi, and Emmanuel Rotimi Adagunodo. Assessment of measures for information retrieval system evaluation: a user-centered approach. International Journal of Computer Applications, 25(??):6–12, July 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume25/
number7/3046-4138/.

Absar:2014:PSS

[AAA14a] Nurul Absar, Mohammed Jahangir Alam, and Tasnuva Ahmed. Performance study of star topology in small Internetworks. International Journal of Computer Applications, 107(??):45–53, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number2/18727-9961/.

Ali:2014:DOP

[AAA14b] Ramzy S. Ali, Ammar A. Aldair, and Ali K. Almousawi. Design an optimal PID controller using artificial bee colony and genetic algorithm for autonomous mobile robot. International Journal of Computer Applications, 100(??):8–16, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number16/17607-8016/.

Ali:2016:NTA

[AAA16] A. M. Ali, A. E. Amin, and A. F. Abdo. Natural therapy advisory system based on fusion of image processing and text processing. International Journal of Computer Applications, 154(??):1–6, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume154/number10/26524-2016912231/.

Aldair:2017:DPC

[AAA17a] Ammar A. Aldair, Turki Y. Abdalla, and Eman Badee Alsaedee. Design of PID controller using ABC for full vehicle nonlinear active suspension system with passenger seat. International Journal of Computer Applications, 167(??):18–27, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number11/27815-2017914370/.

Alzahrani:2017:CSD

[AAA^{+}17b] Sara Alzahrani, Hanan Ahmed, Hebah Alnahdi, Roaa Alamoudi, and Hemalatha M. Comparative study for different antenna designs for the 4G networks. International Journal of Computer Applications, 167(??):1–4, June 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume167/number8/27788-2017914320/.

Ayogu:2019:FST

[AAA19] Ikechukwu I. Ayogu, Adebayo O. Adetunmbi, and Bosede A. Ayogu. A first step towards the development of Yoruba named entity recognition system. International Journal of Computer Applications, 182(??):1–4, February 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume182/number41/30362-2019918465/.

Adenowo:2020:SBD

[AAA20] Adetokunbo A. Adenowo, Adesoji A. Awobajo, and Sheriff Alimi. Software-based diagnostic approach for detection of malaria parasite in blood. International Journal of Computer Applications, 175(??):17–23, August 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume175/number12/31504-2020920600/.

Alhazmi:2021:MMO

[AAA21] Elaf Alhazmi, Abdulwahab Aljubairy, and Ahoud Alhazmi. Memory management via ownership concept Rust and Swift: Experimental study. International Journal of Computer Applications, 183(??):1–10, August 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number22/32054-2021921572/.

Alenezi:2017:IIS

[AAAA17] Khalid Abdulkareem Alenezi, Imad Fakhri Taha Al Shaikhli, Sufyan Salim Mahmood AlDabbagh, and Khaled Alhatem. The influence of Internet and social media on purchasing decisions in Egypt and a comparison between Egypt and Kuwait. International Journal of Computer Applications, 178(??):26–32, November 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number5/28672-2017915826/.

Alam:2021:VIM

[AAAA21] Md Nakibul Alam, Md Abdullah Al Faruk, Rifat Tasnim Anannya, and D. M. Shafin Ahmed. ViNET infrastructure model for Internet connectivity in the scenario of dhaka city. International Journal of Computer Applications, 183(??):40–53, October 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number32/32141-2021921715/.

Asif:2012:PED

[AAaAD12] R. Asif, R. A. Abd-alhameed, O. O. Anoh, and Y. A. S. Dama. Performance evaluation of DWT-OFDM and FFT-OFDM for multicarrier communications systems using time domain zero forcing equalization. International Journal of Computer Applications, 51(??):34–38, August 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume51/number4/8032-1314/.

Asif:2015:UWB

[AAaN15] Rameez Asif, Raed Abd-alhameed, and J. M. Noras. A unique wavelet-based multicarrier system with and without MIMO over multipath channels with AWGN. International Journal of Computer Applications, 117(??):31–40, May 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume117/number9/20585-3015/

.

Anoh:2013:WPT

[AAaNJ13] Kelvin O. O. Anoh, R. A. Abd-alhameed, J. M. Noras, and S. M. R. Jones. Wavelet packet transform modulation for multiple input multiple output applications. International Journal of Computer Applications, 63(??):46– 51, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number7/10482-5221/.

Akoto-Adjepong:2020:ENC

[AAAOG20] Vivian Akoto-Adjepong, Michael Asante, and Steve Okyere-Gyamfi. An enhanced non-cryptographic hash function. International Journal of Computer Applications, 176(15): 10–17, ???? 2020. ISSN 0975-8887. URL https://bit.

ly/4gf7me4;   https://www.ijcaonline.org/archives/
volume176/number15/31276-2020920014.

Aidoo:2022:MPA

[AAB22] Anthony Aidoo, Frank Arthur, and Gloria A. Botchway. A morphological pyramids approach to grayscale image enhancement. International Journal of Computer Applications, 184(??):1–10, October 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number31/32509-2022922363/.

Arora:2013:SSE

[AAC13] Swinky Arora, Ankit Arora, and Gursharanjit Singh Cheema. Scheduling simulations: an experimental approach to time-sharing multiprocessor scheduling schemes. International Journal of Computer Applications, 63(??): 29–35, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number11/10512-5476/.

Ahirwar:2016:BBA

[AAC16] Mahendra Kumar Ahirwar, Manish Kumar Ahirwar, and Uday Chourasia. Black box anomaly detection in multicloud environment. International Journal of Computer Applications, 144(??):31–37, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number2/25155-2016910125/.

Anand:2013:ICC

[AAD13] Abhineet Anand, Abhijeet, and D P Das. Improving convergence of congestion control algorithm. International Journal of Computer Applications, 81(??):1– 4, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume81/
number2/13981-1977/.

AlMurtadha:2016:FIA

[AAEA16] Yahya M. AlMurtadha, Khaled M. Alhawiti, Abdelrahman Osman Elfaki, and Osman A. Abdalla. Factors influencing academic achievement of undergraduate computing students. International Journal of Computer Applications, 146(??):23–28, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number3/25379-2016910658/.

Ahmed:2015:BDM

[AAEEA15] Ahmed M. Ahmed, Amr Ali-Eldin, Mohamed S. Elksasy, and Faiz F. Areed. Brushless DC motor speed control using both PI controller and fuzzy PI controller. International Journal of Computer Applications, 109(??):29–35, January 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume109/number10/19227-
0933/.

AbdEl-atey:2011:CSW

[AaEeK11] Amira AbdEl-atey, Sherif El-etriby, and Arabi Kishk. Constructing Semantic Web form from unstructured Web page. International Journal of Computer Applications, 30(??):34– 41, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume30/
number10/3675-5137/.

AbdelFattah:2018:SSH

[AAEK18] M. AbdelFattah, L. F. AbdelAal, and R. El-Khoribi. Spectral-spatial hyperspectral image classification based on randomized singular value decomposition and 3-dimensional discrete wavelet transform. International Journal of Computer Applications, 180(??):1–10, April 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number30/29231-2018916753/.

Alsharif:2013:ECA

[AAG13] Ouais Alsharif, Deema Alshamaa, and Nada Ghneim. Emotion classification in Arabic poetry using machine learning. International Journal of Computer Applications, 65 (??):10–15, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number16/11006-6300/.

Anwar:2015:DTC

[AAGM15] Mohammed Khaleel Anwar, Mohammed Al-Gawagzeh, and Syed Mujeebuddin. Design of ternary codes for various lengths using MSAA. International Journal of Computer Applications, 116(??):20–26, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number6/20340-2130/.

Abdelbar:2021:TIP

[AAH21] Mahmoud S. Abdelbar, Abdelmgeid A. Ali, and M. Hasabala. A technique for increasing payload capacity of RGB images steganography based on mod factor and segmentation. International Journal of Computer Applications, 183(??):29–35, July 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number13/31988-2021921439/.

Addah:2013:AMC

[AAI13] Jerome Addah, Stephen Apanga, and Mustapha Issahaku. Adopting medical computer systems: the perspectives of physicians in a poor setting. International Journal of Computer Applications, 75(??):43–47, August 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume75/number18/13352-0929/.

Arthur:2020:SCB

[AAMA20] John Kingsley Arthur, Kofi Sarpong Adu-Manu, and Charles Adjetey. A secured cloud-based e-voting system using information dispersal algorithm. International Journal of Computer Applications, 175(??):27–32, September 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume175/number20/31570-2020920734/.

Amirrudin:2013:MPM

[AAMG13] Nurul‘ain Amirrudin, Sharifah H. S. Ariffin, N. N. N. Abd Malik, and N. Effiyana Ghazali. Mobility prediction via Markov model in LTE femtocell. International Journal of Computer Applications, 65(??):40–44, March 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume65/number18/11027-6404/.

Abdalla:2012:RCC

[AAN12] Abdulkareem Y. Abdalla, Turki Y. Abdalla, and Khloud A. Nasar. Routing with congestion control in computer network using neural networks. International Journal of Computer Applications, 57(??):34–41, November 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume57/number2/9089-3031/.

A:2019:MPR

[AAN^{+}19] Oladeji F A, Idowu P A, Egejuru N, Faluyi S G, and Balogun J A. Model for predicting the risk of kidney stone using data mining techniques. International Journal of Computer Applications, 182(??):36–56, January 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume182/number38/30317-2019918404/.

Akintoye:2011:DWB

[AAO11] K. A. Akintoye, O. T. Arogundade, and Oluwakemi Oke. Development of a Web-based student–lecturer relationship information system (e-assessment). International Journal of Computer Applications, 25(??):43–47, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume25/number8/3049-4145/.

Arai:2013:IOP

[AAO13] Kohei Arai, Indra Nugraha Abdullah, and Hiroshi Okumura. Identification of ornamental plant functioned as medicinal plant based on redundant discrete wavelet transformation. International Journal of Computer Applications, 61 (??):34–38, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number1/9895-4468/.

Attar:2014:RRP

[aAO14] Noha El. attar, Wael Awad, and Fatma Omara. RPOAWLB: Resource provisioning optimization approach based on RPOA with load balance. International Journal of Computer Applications, 105(??):34–41, November 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume105/number7/18392-9650/.

Agangiba:2022:CMA

[AAO22] William Akotam Agangiba, Millicent Agangiba, and George Essah Yaw Okai. Computational modeling and assessment of acceptance level of virtual learning platforms. International Journal of Computer Applications, 184(??):49–54, May 2022. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume184/number12/32380-
2022922108/.

Assa-Agyei:2023:OPA

[AAOA23] Kwame Assa-Agyei, Funminiyi Olajide, and Temitope Alade. Optimizing the performance of the advanced encryption standard techniques for secured data transmission. International Journal of Computer Applications, 185(??):31–36, July 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume185/number21/32818-
2023922941/.

Aljaafreh:2014:ATS

[AAoS14] Ahmad Aljaafreh, Naeem Al-oudat, and Ma’en Saleh. Adaptive traffic-signal control using discrete event simulation model. International Journal of Computer Applications, 101(??):7–12, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number12/17737-8910/.

A:2013:FEH

[AAP13] Robinson Silvester A, J. Cruz Antony, and M. Pratheepa. Fast and efficient hashing for sequence similarity search using substring extraction in DNA sequence databases. International Journal of Computer Applications, 78(??):13– 17, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume78/
number9/13516-1295/.

Ali:2014:TMU

[AAqA14] Hussain Zaydan Ali, Salih M. Al-qaraawi, and Ghusoon Eidan Arab. TRMM mapping using different models in Iraq. International Journal of Computer Applications, 96 (??):33–37, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number14/16865-6753/.

Asnawi:2011:FVL

[AAR11] Rustam Asnawi, Wan Fatimah Wan Ahmad, and Dayang Rohaya Awang Rambli. Formalization and verification of a live multimedia presentation model. International Journal of Computer Applications, 20(??):32–41, April 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume20/number2/2404-3198/.

Anbumalar:2013:SNN

[AAR13] S. Anbumalar, R. Anandanatarajan, and P. Rameshbabu. Sparse non-negative matrix factorization and its application in overlapped chromatograms separation. International Journal of Computer Applications, 63(??):1– 10, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number21/10587-5199/.

Asaduzzaman:2014:TPP

[AAR14] A. Asaduzzaman, R. Asmatulu, and M. Rahman. Teaching parallel programming for time-efficient computer applications. International Journal of Computer Applications, 90 (??):18–25, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume90/
number7/15585-4264/.

Abdelrazek:2019:DCC

[AARA19] Elsaeed Abdelrazek, Doaa M. Abu-Rashed, and Asmaa E. Ahmed. Developing the curriculum of computer mathematics using mobile educational techniques to improve students’ skills dealing with logic gates. International Journal of Computer Applications, 178(??):6–17, September 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number47/30865-2019919348/.

Azmy:2013:AUS

[AAS13] Waleed M. Azmy, Sherif Abdou, and Mahmoud Shoman. Arabic unit selection emotional speech synthesis using blending data approach. International Journal of Computer Applications, 81(??):22–28, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume81/number8/14033-1887/.

Abdulrahim:2014:NIR

[AAS14] Abdulrazaq Abdulrahim, Saleh E. Abdullahi, and Junaidu B. Sahalu. A new improved round robin (NIRR) CPU scheduling algorithm. International Journal of Computer Applications, 90(??):27–33, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume90/number4/15563-4277/.

Amer:2024:BJW

[AAS24] Mohamed Amer, Mohamed Amer, and Tarek S. Sobh. Biometric JSON Web tokens (BJWT): Enhancing Web API security with biometric key exchange and OTP-JWT authentication. International Journal of Computer Applications, 186(??):15–21, September 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number39/biometric-json-web-tokens-bjwt-
enhancing-web-api-security-with-biometric-key-exchange-
and-otp-jwt-authentication/.

Almomani:2015:PEP

[AASAA15] Omar Almomani, Mahmoud Al-Shugran, Jafar A. Alzubi, and Omar A. Alzubi. Performance evaluation of positionbased routing protocols using different mobility models in MANET. International Journal of Computer Applications, 119(??):43–48, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number3/21050-3692/.

Ahammad:2016:TAB

[AAT16] Tanvir Ahammad, Md. Sajib Al Mamun, and Mehnaz Tabassum. Towards the application of big data: a new way to make data driven healthcare decision. International Journal of Computer Applications, 134(??):15–21, January 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume134/number14/23982-
2016908094/.

Anand:2015:NES

[AAY15] Smit Anand, Nishat Afreen, and Shama Yazdani. A novel and efficient selection method in genetic algorithm. International Journal of Computer Applications, 129(??):7–12, November 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume129/number15/23147-
2015907067/.

Ayeni:2017:IDT

[AAY17] Bakare K. Ayeni, Obiniyi A. Afolayan, and Nurat Yusuf. Improved dynamic time slice round robin scheduling algorithm with unknown burst time. International Journal of Computer Applications, 167(??):21–28, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number7/27784-2017914305/.

Abdo:2021:FLA

[AAY21] Mohammed Ameen A. Abdo, Ala Eldin Abdallah Awouda, and Yousif Elfatih Yousif. A fuzzy logic adaptive image compression level using cross-layering in wireless multimedia sensor networks. International Journal of Computer Applications, 174(??):18–24, April 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number28/31853-2021921200/.

Abdulrazaq:2013:TSA

[AAYM13] Abdulrazaq Nathim Abdulrazaq, Mohd Sharifuddin Ahmad, Mohd Zaliman M. Yusoff, and Moamin A. Mahmoud. A trustable software agent framework for the examination paper preparation and moderation process (EPMP). International Journal of Computer Applications, 83(??):23– 30, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number15/14525-2917/.

Amer:2021:NTS

[AAZN21] Dina A. Amer, Gamal Attiya, Ibrahim Ziedan, and Aida A. Nasr. A new task scheduling algorithm based on water wave optimization for cloud computing. International Journal of Computer Applications, 183(??):65–75, May 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume183/number3/31912-2021921320/.

Andamuthu:2011:DXM

[AB11] T. P. Andamuthu and P. Balasubramanie. Developing Xquery model for distributed query processing in mobile networks. International Journal of Computer Applications, 25(??):36–41, July 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume25/
number4/3018-4082/.

Aminzadeh:2012:CLD

[AB12a] Hamed Aminzadeh and Ali Bakhshi. Circuit-level design of a power supply unit with extra low-noise output for portable integrated SoCs. International Journal of Computer Applications, 50(??):1–6, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number14/7836-0746/.

Ammari:2012:PAP

[AB12b] Mohamed L. Ammari and Anis Bdiri. Peak-to-average power ratio analysis for golden coded MIMO systems. International Journal of Computer Applications, 52(??):32– 38, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number6/8208-1616/.

Acakpovi:2013:NPM

[AB13a] Amevi Acakpovi and Essel Ben Hagan. Novel photovoltaic module modeling using Matlab/Simulink. International Journal of Computer Applications, 83(??):27– 32, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number16/14535-2978/.

AlAwadi:2013:MPI

[AB13b] Aymen Hasan Rashid Al Awadi and Bahari Belaton. Multiphase IRC botnet and botnet behavior detection model. International Journal of Computer Applications, 66(??): 41–51, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number15/11164-6289/.

Asoodeh:2013:CPE

[AB13c] Mojtaba Asoodeh and Parisa Bagheripour. Core porosity estimation through different training approaches for neural network: Back-propagation learning vs. genetic algorithm. International Journal of Computer Applications, 63 (??):11–15, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number5/10461-5172/.

A:2014:IRU

[AB14a] Phuke V A and H. N. Bharathi. Information retrieval using Dempster–Shafer theory. International Journal of Computer Applications, 102(??):33–37, September 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume102/number13/17879-8876/.

Acakpovi:2014:WTS

[AB14b] Amevi Acakpovi and Essel Ben Hagan. A wind turbine system model using a doubly-fed induction generator (DFIG). International Journal of Computer Applications, 90(??):6–11, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume90/
number15/15794-4477/.

Ahmed:2014:TWM

[AB14c] Rustam Ali Ahmed and Bhogeswar Borah. Triangle wise mapping technique to transform one face image into another face image. International Journal of Computer Applications, 87(??):1–8, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume87/number6/15209-3714/.

Ananthi:2014:HID

[AB14d] P. Ananthi and P. Balasubramanie. Hybrid intrusion detection system using FCRM mechanism. International Journal of Computer Applications, 105(??):25– 29, November 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume105/number9/18406-9676/.

Ahuja:2016:DDH

[AB16a] Lincon Ahuja and Prashant Bhardwaj. Design and development of health monitored quick appointment system. International Journal of Computer Applications, 152(??):14–19, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume152/number7/26331-2016911893/.

Al-Barakati:2016:NSE

[AB16b] Abdullah Al-Barakati. News source evaluation and visualization system. International Journal of Computer Applications, 154(??):1–12, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume154/number8/26508-2016912178/.

Adda:2017:DHA

[AB17] Chahrazed Adda and Abou Elhassen Benyamina. Design of the H264 application and implementation on heterogeneous architectures. International Journal of Computer Applications, 180(??):23–31, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number7/28812-2017916056/.

Asiedu:2020:ESA

[AB20] William Asiedu and K. O. Boateng. Efficient and secure AAA for mobile router in network mobility environment (A3MR-NEMO). International Journal of Computer Applications, 175(??):12–19, September 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number17/31543-2020920676/.

Ahmad:2010:CSB

[ABA^{+}10] Shish Ahmad, Mohd. Rizwan Beg, Qamar Abbas, Jameel Ahmad, and Mohd Atif. Comparative study between stream cipher and block cipher using RC4 and Hill Cipher. International Journal of Computer Applications, 1(??):9–12, February 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume1/number25/465-
770/.

Al-barhamtoshy:2012:TTA

[AbAAw12] Hassanin Al-barhamtoshy, Sherif Abdou, and Fakhraddin A. Al-wajih. A toolkit for teaching Arabic handwriting. International Journal of Computer Applications, 49(??): 17–23, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number23/7943-1273/.

Apene:2024:DMM

[ABAO24] Oghenevovwero Zion Apene, Nachamada Vachaku Blamah, Gilbert Imuetinyan Osaze Aimufua, and Morufu Olalere. Development of a mathematical model for crime detection based on YOLO network architecture. International Journal of Computer Applications, 186(??):17–24, May 2024. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.
org/archives/volume186/number20/development-of-a-
mathematical-model-for-crime-detection-based-on-yolo-
network-architecture/.

Abbas:2012:APC

[Abb12a] Mohamed Abd Elhamid Abbas. Accurate prediction of cancers stem cells incident using enhanced Adaboost algorithm. International Journal of Computer Applications, 53 (??):19–25, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number1/8385-1990/.

Abbas:2012:ACH

[Abb12b] Mohamed Abd Elhamid Abbas. Anticipating of cardiovascular heart diseases using computer based poly trees model. International Journal of Computer Applications, 59(??):19– 26, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number13/9608-4239/.

Aikins-Bekoe:2020:ECD

[ABB20] Stephen Aikins-Bekoe and James Ben Hayfron-Acquah. Elliptic curve Diffie–Hellman (ECDH) analogy for secured wireless sensor networks. International Journal of Computer Applications, 176(??):1–8, April 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number10/31234-2020920015/.

Atalla:2014:LBM

[ABBG14] Shadi Atalla, Andrea Bianco, Robert Birke, and Luca Giraudo. A load balancer for a multi-stage router architecture. International Journal of Computer Applications, 89 (??):1–7, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number3/15479-4193/.

Abdulhussain:2013:EMC

[Abd13] Sadiq H. Abdulhussain. Enhanced management of certificate caching and revocation lists in VANET. International Journal of Computer Applications, 83(??):38– 47, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number12/14503-2393/.

Abduljalil:2014:VCS

[Abd14a] Fekri M. Abduljalil. Video capture service in the intelligent transportation system based on cloud computing. International Journal of Computer Applications, 97 (??):35–41, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number5/17007-7159/.

Abdulwahhab:2014:MPE

[Abd14b] Omar Waleed Abdulwahhab. Mobile position estimation based on three angles of arrival using an interpolative neural network. International Journal of Computer Applications, 100(??):1–5, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number7/17534-8109/.

Abdelfatah:2017:NPS

[Abd17] Roayat Ismail Abdelfatah. A novel proxy signcryption scheme and its elliptic curve variant. International Journal of Computer Applications, 165(??):36–43, May 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume165/number2/27549-2017913804/.

Abdalla:2018:AFF

[Abd18] Turki Y. Abdalla. Adaptive fuzzy FOPID control scheme for path tracking of mobile robot. International Journal of Computer Applications, 181(??):1–5, October 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number22/30014-2018917887/.

AssisAmorim:2022:MVC

[ABD22] Antonio Eduardo Assis Amorim, Líria Baptista De Rezende, and Suzana De Almeida Prado Pohl Sanzovo. A modified version of Canny filter applied to identify oxidation points in steel plate. International Journal of Computer Applications, 184(??):45–52, May 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number11/32372-2022922091/.

Akhter:2014:MWA

[ABF14] Manira Akhter, Ashin Ara Bithi, and Abu Ahmed Ferdaus. Mining Web access patterns using root-set of suffix trees. International Journal of Computer Applications, 94(??):23–29, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number9/16372-5818/.

Atole:2017:ESR

[ABL17] Ronnel R. Atole, Leo Constantine S. Bello, and John Rey S. Lirag. Eyes in the sky: a review of civilian unmanned aerial vehicles (UAVs). International Journal of Computer Applications, 173(??):36–41, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume173/number6/28343-2017915349/.

Areed:2012:DSM

[ABS^{+}12] Fayez G. Areed, Mohamed A. Badr, Sabry F. Saraya, Mohammed S. Elksasy, and Mohamed M. Abdelsalam. Decoupled sliding mode control for a multivariable nonlinear system. International Journal of Computer Applications, 55(??):25–32, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume55/number6/8760-2674/.

Abuhaija:2014:SSA

[Abu14] Belal Abuhaija. System selection algorithm for heterogeneous radio access networks. International Journal of Computer Applications, 88(??):35–40, February 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume88/number13/15416-4065/.

Abuzanouneh:2017:HMO

[Abu17] Khalil Ibrahim Mohammad Abuzanouneh. Hybrid multiobjectives genetic algorithms and immigrants scheme for dynamic routing problems in mobile networks. International Journal of Computer Applications, 164(??):49–57, April 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume164/number5/27483-2017913641/

.

Abdul-Barik:2022:DEW

[ABYB22] Alhassan Abdul-Barik, Farudeen Yussif, and Malik Sumaila Bipembi. Deployment of an enhanced Web-based really simple syndication (RSS) aggregator. International Journal of Computer Applications, 184(??):1–5, July 2022. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number23/32454-2022922210/.

Aswale:2013:CSD

[AC13] P. S. Aswale and S. S. Chopade. Comparative study of different low power design techniques for reduction of leakage power in CMOS VLSI circuits. International Journal of Computer Applications, 70(??):16–24, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume70/number11/12005-6866/.

Alzughaibi:2015:EEB

[AC15] Arwa Darwish Alzughaibi and Zenon Chaczko. Evaluation of entropy-based edge detector methods for human object. International Journal of Computer Applications, 129(??):1–5, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume129/number13/23130-2015907077/.

A:2018:KAC

[AC18] Kulkarni Mayuri A and V. R. Chirchi. Key-aggregate cryptosystem based on elliptic curve cryptography for data sharing in cloud storage with result and analysis. International Journal of Computer Applications, 179(??):22–25, June 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume179/number48/
29501-2018917265/.

Aaronson:2017:RFD

[ACB17] Robert Yao Aaronson, Wu Chen, and Ben-Bright Benuwa. Robust face detection using convolutional neural network. International Journal of Computer Applications, 170(??):14–20, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume170/number6/28073-2017914855/.

Aggarwal:2012:STF

[ACK12] Madhu Aggarwal, Renu Chugh, and Raj Kamal. Suzukitype fixed point results in G-metric spaces and applications. International Journal of Computer Applications, 47 (??):14–17, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number12/7239-0073/.

Anitha:2014:LPA

[ACK14] D. Anitha, K. Manjunatha Chari, and P. Satish Kumar. Low power ALU design considering PVT variations. International Journal of Computer Applications, 104(??):19–23, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number17/18300-9442/.

Al-Dulaimi:2011:HSV

[AD11] Khamael Abbas Al-Dulaimi. Handwritten signature verification technique based on extract features. International Journal of Computer Applications, 30(??):42–46, September 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume30/number2/3612-
5024/.

Akter:2013:CPI

[AD13] Samima Akter and Priyanath Das. Comparison of the performance of IPFC (series-series) and UPFC (series-shunt) FACTS controller in power system. International Journal of Computer Applications, 67(??):42–48, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume67/number2/11371-7020/.

Ade:2014:IBV

[AD14a] Roshani Ade and P. R. Deshmukh. Instance-based vs batchbased incremental learning approach for students classification. International Journal of Computer Applications, 106(??):37–41, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number3/18504-9580/.

Amin:2014:PDR

[AD14b] Abu Bony Amin and Choton Kanti Das. Prominence of different routing protocols in MANET for heterogeneous communication environment. International Journal of Computer Applications, 95(??):9–15, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number1/16557-1753/.

Abdo:2015:TIS

[AD15] Hosam Abdo and Darko Dimitrov. The total irregularity of some composite graphs. International Journal of Computer Applications, 122(??):1–9, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number21/21846-5170/.

Afriyie:2018:MBE

[AD18] Yaw Afriyie and M. I. Daabo. Multiple bits error detection and correction in RRNS architecture using the MRC and HD techniques. International Journal of Computer Applications, 180(??):18–23, May 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number39/29387-2018917030/.

A:2024:ABC

[AD24] Divya A and Janaki Sathya D. Artificial bee colony (ABC) optimization algorithm based automatic segmentation and detection of suspicious lesions in lung CT images. International Journal of Computer Applications, 186(??):7–13, September 2024. CODEN ????

ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume186/number41/artificial-bee-colony-
abc-optimization-algorithm-based-automatic-segmentation-
and-detection-of-suspicious-lesions-in-lung-ct-images/

.

Addah:2012:CLL

[Add12] Jerome Addah. Computer literacy and e-learning: Attitudes among first year students in a Ghanaian medical school. International Journal of Computer Applications, 51 (??):21–25, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number22/8343-1960/.

Adenugba:2011:VNC

[Ade11] D. A. Adenugba. A Vb.Net class library and a client application for radio refractivity. International Journal of Computer Applications, 33(??):6–13, November 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume33/number3/3998-5665/.

Adeogun:2013:CBC

[Ade13a] Ramoni O. Adeogun. Cluster based channel model and performance analysis for MIMO satellite formation flying communication systems. International Journal of Computer Applications, 71(??):1–5, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number11/12399-9084/.

Adeogun:2013:RMB

[Ade13b] Ramoni O. Adeogun. A robust MUSIC based scheme for interference location in satellite systems with multibeam antennas. International Journal of Computer Applications, 82 (??):1–6, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume82/
number12/14165-2322/.

Al-deen:2015:DAD

[ADFAJ15] Anhar Khairy Al-Deen, Kubais Saeed Fahady, and Reem Ali Al-Jarah. Design advance database supported with some of data mining ensemble for early deduction of osteoporosis. International Journal of Computer Applications, 121(??):18– 29, July 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume121/number16/
21625-4928/.

Abdelghany:2017:THA

[ADH17] Abdelghany Salah Abdelghany, Nagy Ramadan Darwish, and Hesham Ahmed Hefny. Towards a hybrid approach for software project management using ontology alignment. International Journal of Computer Applications, 168(??):12–19, June 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume168/number6/27878-2017914438/.

Asabere:2014:AEB

[ADKsO14] Nana Yaw Asabere, Vida Doku, Simonov Kusi-sarpong, and David Oppong. Adopting electronic business in Ghana: Story of the hospitality industry. International Journal of Computer Applications, 85(??):1–9, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume85/number4/14826-3067/.

Anand:2016:IFE

[ADS16] Adarsh Anand, Deepika, and Ompal Singh. Incorporating features enhancement archetype in software reliability growth modeling and optimal release time determination. International Journal of Computer Applications, 139(??):1–6, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number4/24475-2016909137/.

Abdullah:2017:NRM

[ADT17] Mohammed Najm Abdullah, Afrah Salman Dawood, and Ali Kamal Taqi. Network resource management optimization for SDN based on statistical approach. International Journal of Computer Applications, 177(??):

5–13, November 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume177/number6/28628-2017915766/.

Amirtharajan:2010:PBP

[ADVB10] R. Amirtharajan, Adharsh D, Vignesh V, and R. John Bosco Balaguru. PVD blend with pixel indicator — OPAP composite for high fidelity steganography. International Journal of Computer Applications, 7(??):31–37, October 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume7/number9/1275-1801/.

Abdalla:2011:MDD

[ADZ11] Al Mutaz M. Abdalla, Safaai Dress, and Nazar Zaki. Masses detection in digital mammogram by gray level reduction using texture coding method. International Journal of Computer Applications, 29(??):19–23, September 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume29/number4/3555-4887/.

Ali:2012:ESM

[AE12a] Mohammad Athar Ali and Eran A. Edirisinghe. Efficient spatiotemporal matching for video copy detection in H.264/ AVC video. International Journal of Computer Applications, 41(??):1–7, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume41/
number15/5614-7890/.

Allali:2012:DNS

[AE12b] Karam Allali and Kamal El Karouni. Direct numerical simulations of reaction fronts propagation under quasi-periodic gravitational modulation. International Journal of Computer Applications, 50(??):32–36, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number6/7778-0863/.

Azarbad:2012:ECU

[AE12c] Milad Azarbad and Ataollah Ebrahimzadeh. ECG compression using the three-level quantization and wavelet transform. International Journal of Computer Applications, 59 (??):28–38, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number1/9515-3916/.

Abbass:2013:GLB

[AE13] Mohamed A. Abbass and Emad El Samahy. A generic luczak-based cardiovascular model for healthy subjects under physical stress. International Journal of Computer Applications, 66(??):29–35, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume66/number16/11170-6347/.

Abu-errub:2014:ATC

[Ae14a] Aymen Abu-errub. Arabic text classification algorithm using TFIDF and chi square measurements. International Journal of Computer Applications, 93(??):40–45, May 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume93/number6/16223-5674/.

Augustine:2014:SEA

[AE14b] Anupriya Augustine and Jubin Sebastian E. A study of efficient anonymous routing protocols in MANET. International Journal of Computer Applications, 91(??): 1–7, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume91/
number8/15898-4876/.

Arumugam:2017:DMB

[AE17] P. Arumugam and R. Ezhilarasi. Data mining based neural network model for rainfall forecasting. International Journal of Computer Applications, 170(??):30–33, July 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume170/number4/28061-2017914831/

.

Ajemni:2017:DHB

[AEA17] H. Ajemni, R. El Harabi, and M. N. Abdelkrim. Directed hypergraph-based models for the fault monitoring of chemical reaction kinetics. International Journal of Computer Applications, 166(??):1–8, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume166/number12/27718-2017914040/.

Abdeltwab:2015:PIC

[AEE15] Marwa Abdelfatah Abdeltwab, K. A. Elbarbary, and Adel Elhenawy. Performance improvement of coded OFDM communication system in AWGN channel. International Journal of Computer Applications, 110(??):17–23, January 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume110/number11/19361-
0969/.

Abunajm:2024:DLA

[AEEO24] Saleh Abunajm, Nelly Elsayed, Zag Elsayed, and Murat Ozer. Deep learning approach for early stage lung cancer detection. International Journal of Computer Applications, 186(??):11–17, February 2024. CODEN ????

ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume186/number8/deep-learning-approach-
for-early-stage-lung-cancer-detection/.

Abdelbary:2013:STM

[AEk13] Haasan Abdelbary and Abeer El-korany. Semantic topics modeling approach for community detection. International Journal of Computer Applications, 81(??):50– 58, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume81/
number6/14020-2177/.

Alzaher:2011:RTM

[AEKA11] Mohamed S. Abd Alzaher, Hussein A. Elsayed, Somaia I. Kayed, and Wagdy R. Anis. Road traffic modeling using data communication networks. International Journal of Computer Applications, 26(??):24–31, July 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume26/number4/3091-4241/.

Aboshosha:2015:SLF

[AEKE15] Ashraf Aboshosha, Kamal A. El Dahshan, Eman A. Karam, and Ebeid A. Ebeid. Score level fusion for fingerprint, iris and face biometrics. International Journal of Computer Applications, 111(??):47–55, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number4/19530-1171/.

Alkotby:2018:ESD

[AER18] Mohamed R. Alkotby, Elsaeed Elsaeed Mohamed Abd Elrazek, and M. Z. Rashad. An expert system to diagnose and fix common car breakdowns for industrial technical education in Egypt. International Journal of Computer Applications, 182(??):30–37, August 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number7/29835-2018917648/.

Arafa:2019:PAB

[AER19] Rasha Kandel M. Arafa, Elsaeed Elsaeed Mohamed Abd Elrazek, and M. Z. Rashad. A proposed application based on mobile learning techniques to manage electronic tests. International Journal of Computer Applications, 181(??):51–57, April 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume181/number49/
30493-2019918699/.

Alkotby:2021:ISD

[AERE21] Mohamed R. Alkotby, Elsaeed E. Mohamed Abd Elrazek, M. Z. Rashad, and M. M. R. Elbasyouny. Intelligent system for diagnosing and fixing automatic washing machine malfunctions for industrial technical education in Egypt. International Journal of Computer Applications, 174(??):19–28, March 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume174/number27/
31845-2021921198/.

Amin:2011:MIL

[AERS11] Mohamed Amin, Nabawia El-Ramly, and Doaa Shebl. Modeling intelligent e-learning systems based on adaptive fuzzy higher order Petri nets. International Journal of Computer Applications, 25(??):7–14, July 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume25/number10/3151-4355/.

Ariyasingha:2015:PSM

[AF15] I. D. I. D. Ariyasingha and T. G. I. Fernando. A performance study for the multi-objective ant colony optimization algorithms on the job shop scheduling problem. International Journal of Computer Applications, 132(??):1–8, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number14/23659-2015907638/.

Afroze:2014:IIA

[Afr14] Tamanna Afroze. Improved image analysis technique for embedded system design. International Journal of Computer Applications, 86(??):1–4, January 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume86/number3/14962-3136/.

Afroze:2017:TSM

[Afr17] Tamanna Afroze. Temperature sensitive microprocessor design to reduce heat generation and improve performance. International Journal of Computer Applications, 164(??):1–8, April 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume164/number3/27460-2017913596/.

Aisha:2014:NLM

[AFW14] H. A. Aisha, W. L. Fatima, and M. Y. Waziri. Newton’slike method for solving systems of nonlinear equations with singular Jacobian. International Journal of Computer Applications, 98(??):1–3, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume98/number13/17240-7574/.

Agarkar:2010:FBC

[AG10] A. M. Agarkar and A. A. Ghatol. FFANN based cost effective major infant disease management. International Journal of Computer Applications, 7(??):29–33, October 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume7/number11/1289-1755/.

Awari:2013:DLC

[AG13] Gaurav S. Awari and Minal S. Ghute. Design of low cost virtual patient for real time ECG analyzers. International Journal of Computer Applications, 68(??):12– 16, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume68/
number24/11726-7315/.

Arora:2014:IPC

[AG14] Indu Arora and Anu Gupta. Improving performance of cloud based transactional applications using in-memory data grid. International Journal of Computer Applications, 107(??):14–19, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number13/18811-0387/.

Adeyemo:2016:OSS

[AGFA16] O. A. Adeyemo, O. N. Gidi-Fanimokun, and J. O. Alabi. Online support system for diabetes management. International Journal of Computer Applications, 152(??):6–11, October 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume152/number10/
26353-2016911454/.

Aggarwal:2013:HAH

[Agg13] Remica Aggarwal. Hybrid analytic hierarchy process fuzzy linear programming approach for weapon systems screening. International Journal of Computer Applications, 65 (??):7–13, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number7/10934-5883/.

AlvesMorais:2016:UAN

[AGG16] Romulo Alves Morais, Denise Gomes Alves, and Warley Gramacho Da Silva. Use of artificial neural network to optimize osmotic dehydration process of cashew from Cerrado (Anacardium occidentale). International Journal of Computer Applications, 156(??):1–7, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume156/number7/26718-2016912467/.

Aggarwal:2017:CAH

[Agg17] Saurabh Aggarwal. Comparative analysis of hedge funds in financial markets using machine learning models. International Journal of Computer Applications, 163(??):25–29, April 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume163/number3/27377-2017913489/

.

Attia:2014:SIR

[AGH14] Zeinab E. Attia, Ahmed M. Gadallah, and Hesham A. Hefny. Semantic information retrieval model: Fuzzy ontology approach. International Journal of Computer Applications, 91(??):9–14, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume91/number13/15940-5156/.

Alves:2017:CMS

[AGL17] Cézanne Alves, Warley Gramacho Da Silva, and Rafael Lima De Carvalho. Connectivity maintenance of a set of agents through MST-based algorithm. International Journal of Computer Applications, 166(??):18–23, May 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume166/number11/27713-2017914149/.

Abbasifard:2014:SNN

[AGN14] Mohammad Reza Abbasifard, Bijan Ghahremani, and Hassan Naderi. A survey on nearest neighbor search methods. International Journal of Computer Applications, 95 (??):39–52, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number25/16754-7073/.

Agnihotri:2015:IMM

[Agn15] Akshay Agnihotri. Interference mitigation in MC-CDMA system with CFO compensation based modified MMSE-FDE technique. International Journal of Computer Applications, 119(??):7–10, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number6/21070-3743/.

Amirtharajan:2010:OPB

[AGRR10] Rengarajan Amirtharajan, Aishwarya G, Madhumita Rameshbabu, and John Bosco Balaguru Rayappan. Optimum pixel & bit location for colour image stego- a distortion resistant approach. International Journal of Computer Applications, 10 (??):17–24, November 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume10/
number7/1494-2011/.

Asaker:2015:DPM

[AGZ15] A. A. Asaker, R. S. Ghoname, and A. A. Zekry. Design of a planar MIMO antenna for LTE-advanced. International Journal of Computer Applications, 115(??):27–33, April 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume115/number12/
20207-2465/.

Alhanjouri:2011:DSH

[AH11] Mohammed Alhanjouri and Hana Hejazi. DIC structural HMM based IWAK-means to enclosed face data. International Journal of Computer Applications, 18(??):43– 50, March 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume18/
number4/2269-2923/.

Abdulsattar:2012:NMA

[AH12a] Mahmood A. Abdulsattar and Zahir A. Hussein. New multiple antennas method based energy detector for cognitive radio over fading channels. International Journal of Computer Applications, 52(??):20–25, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume52/number5/8198-1588/.

Abidin:2012:NNB

[AH12b] Zaenal Abidin and Agus Harjoko. A neural network based facial expression recognition using fisherface. International Journal of Computer Applications, 59(??):30– 34, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number3/9531-3956/.

Anisheh:2012:DAN

[AH12c] Seyed Mahmoud Anisheh and Hamid Hassanpour. Designing an approach for network traffic anomaly detection. International Journal of Computer Applications, 37 (??):49–55, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume37/
number3/4592-6039/.

Alimoradi:2013:EVD

[AH13a] Dariush Alimoradi and Maryam Hasanzadeh. The effect of variance difference of dyadic quantized histograms on universal steganalysis. International Journal of Computer Applications, 62(??):19–24, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume62/number8/10100-4746/.

Amirshekari:2013:PFM

[AH13b] Samira Amirshekari and Ali Harounabadi. Proposing a formal model for performance improvement in grid environment. International Journal of Computer Applications, 71 (??):33–39, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number12/12413-9179/.

Azzou:2013:TLS

[AH13c] S. Ait Kaci Azzou and M. C. Baba Hnini. Tracking line segment without knowledge of camera motion. International Journal of Computer Applications, 73(??): 1–8, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number6/12742-9650/.

AlSaad:2014:SEB

[AH14] Saad Najim Al Saad and Eman Hato. A speech encryption based on chaotic maps. International Journal of Computer Applications, 93(??):19–28, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume93/number4/16203-5488/.

Abdulmelik:2016:SVC

[AH16] Mustafa B. Abdulmelik and Turki K. Hassan. Sensorless vector controlled multilevel inverter fed BLDC motor. International Journal of Computer Applications, 141(??):16–20, May 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume141/number11/
24828-2016909849/.

Akomolafe:2017:PMA

[AH17a] Oladeji P. Akomolafe and Attah H. Honesty. Protecting mobile agent using enhanced reference monitor based security framework. International Journal of Computer Applications, 161(??):17–22, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number8/27168-2017913245/.

Al-Hamami:2017:PFP

[AH17b] Mohammad Alaa Hussain Al-Hamami. A proposed framework for photos copyright protection in Facebook. International Journal of Computer Applications, 162(??):6–10, March 2017. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume162/number1/27205-
2017913179/.

Afshoon:2014:AVQ

[AHA14] Ramin Afshoon, Ali Harounabadi, and Javad Mir Abedini. Assessment and validating the quality of educational Web sites using subtractive clustering. International Journal of Computer Applications, 98(??):42–47, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume98/number4/17175-7264/.

Appiah:2015:TSF

[AHAPT15] Felix Appiah, J. B. Hayfron-Acquah, Joseph K. Panford, and Frimpong Twum. A tool selection framework for cross platform mobile app development. International Journal of Computer Applications, 123(??):14–19, August 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume123/number2/21931-2015905246/.

Abdel-Halim:2017:ICS

[AHAS^{+}17] Hala A. Abdel-Halim, Othman E A, A. A. Sakr, A. A. Zaki, and A. A. Abouelsoud. An intelligent control system design for an evaporator based on particle swarm optimization. International Journal of Computer Applications, 166(??):17–29, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume166/number9/27698-2017914113/.

Al-hilali:2013:CAC

[AhAT13] Riyadh A. Al-hilali, Abdulkareem S. Abdallah, and Raad H. Thaher. Chaotic adaptive control of non-binary TTCM decoding algorithm. International Journal of Computer Applications, 75(??):12–20, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume75/number5/13106-0414/.

Apau:2016:EDS

[AHAT16] Richard Apau, J. B. Hayfron-Acquah, and Frimpong Twum. Enhancing data security using video steganography, RSA and Huffman code algorithms with LSB insertion. International Journal of Computer Applications, 143(??):28–36, June 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume143/number4/25068-2016910156/.

Al-Hiary:2011:FAD

[AHBAR^{+}11] H. Al-Hiary, S. Bani-Ahmad, M. Reyalat, M. Braik, and Z. ALRahamneh. Fast and accurate detection and classification of plant diseases. International Journal of Computer Applications, 17(??):31–38, March 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume17/number1/2183-2754/.

Al-Hegami:2019:PSB

[AHBG19] Ahmed Sultan Al-Hegami and Abeer Saleh Hamdi Bin-Ghodel. A particle swarm based approach for classification of cancer based on CT scan. International Journal of Computer Applications, 178(??):26–31, May 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number12/30582-2019918862/.

Ahmed:2011:CLB

[AHBH11] Faisal Ahmed, Emam Hossain, A. S. M. Hossain Bari, and Md. Sakhawat Hossen. Compound local binary pattern (CLBP) for rotation invariant texture classification. International Journal of Computer Applications, 33(??):5– 10, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume33/
number6/4022-5724/.

Abduljalil:2013:NAH

[AHH13] Sahar Mohammed Abduljalil, Osman Hegazy, and Ehab E. Hassanein. A novel approach for handling security in cloud computing services. International Journal of Computer Applications, 69(??):9–14, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume69/number5/11837-7562/.

Alikhani:2014:PME

[AHH14] Payam Alikhani, Seyyed Mohammadreza Hesami, and Abdolnabi Hashemi. Prediction of microbial enhanced oil recovery using an artificial intelligence method based on experimental data. International Journal of Computer Applications, 103(??):24–28, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number15/18152-9417/.

Al-Haysah:2020:EVS

[AHH20] Ahmed M. Al-Haysah and A. H. Hasmani. The exact vacuum solution for Kasner metric from Bianchi Type-I cosmological model. International Journal of Computer Applications, 177(??):1–5, March 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number45/31206-2020919969/.

Arakat:2012:ANF

[AHK12] Wafaa A. Arakat, Amira Y. Haikal, and Ayman H. Kassem. Adaptive neuro-fuzzy controller for multi-layered switched reluctance motor. International Journal of Computer Applications, 44(??):20–25, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume44/number1/6228-8304/.

Ahmad:2014:PPD

[AHK14] Syed Md. Tarique Ahmad, Shameemul Haque, and Prince Shoeb Khan. Privacy preserving in data mining by normalization. International Journal of Computer Applications, 96 (??):14–18, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number6/16797-6509/.

Ahmed:2014:BPG

[Ahm14a] Waseem Ahmed. Bridging the performance gap between manual and automatic compilers with intent-based compilation. International Journal of Computer Applications, 86 (??):1–7, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number9/15010-3303/.

Ara:2014:BDR

[AHM14b] Iffat Ara, Md. Najmul Hossain, and S. M. Yahea Mahbub. Baseline drift removal and de-noising of the ECG signal using wavelet transform. International Journal of Computer Applications, 95(??):15–17, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number16/16678-6783/.

Amer:2017:DAF

[AHM17] Ibrahim M. Amer, Salma Hamdy, and Mostafa G. M. Mostafa. Deep Arabic font family and font size recognition. International Journal of Computer Applications, 176(??):1–6, October 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume176/number4/28537-2017915589/.

Ahmed:2020:EPU

[Ahm20] Sumon Ahmed. Effects of pseudotime uncertainty in downstream analysis for single-cell data: a case study. International Journal of Computer Applications, 175(??): 1–6, September 2020. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume175/number18/31549-2020920696/.

Awad:2015:CIE

[AHS15] Abdullah M. Awad, Rehab F. Hassan, and Ali M. Sagheer. Chaos image encryption based on DCT transforms and Henon map. International Journal of Computer Applications, 127(??):1–7, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number11/22770-2015906532/.

Ahmad:2014:SAC

[AHZ14] Mohd. Kalamuddin Ahmad, Mohd. Husain, and A. A. Zilli. A statistical analysis and comparatively study of embedding hypercube. International Journal of Computer Applications, 103(??):1–7, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number16/18155-9254/.

Ahmad:2015:FEH

[AHZ15a] Mohd. Kalamuddin Ahmad, Mohd. Husain, and A. A. Zilli. A framework for embedded hypercube interconnection networks: Based on neural network approach. International Journal of Computer Applications, 120(??):25–34, June 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume120/number2/21201-3873/

.

Al-Hafidh:2015:HTI

[AHZ15b] Manal H. Al-Hafidh and Khawla A. Al Zubaidy. Heat transfer investigation of TiO2 nanofluid in silica aerogel/glass fiber composite material enclosure. International Journal of Computer Applications, 112(??):16–23, February 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume112/number8/19685-1425/.

Abdul-Hadi:2016:SBM

[AHZAH16] Alaa M. Abdul-Hadi, Abdullah M. Zyarah, and Haider M. Abdul-Hadi. Software based method to specify the extreme learning machine network architecture targeting hardware platforms. International Journal of Computer Applications, 138(??):49–53, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume138/number7/24394-2016908967/.

Ahmed:2013:SMP

[AI13a] Razin Ahmed and Md. Fokhrul Islam. E-shaped microstrip patch antenna for Ku band. International Journal of Computer Applications, 80(??):15–19, October 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume80/number6/13864-1719/.

Ali:2013:MIL

[AI13b] Siti Rahaimah Bt Hj Ali and Dato’ Noraini Idris. A model to identify the level of numeracy understanding of primary school pupils: a case study. International Journal of Computer Applications, 67(??):41–49, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume67/number5/11395-6694/.

Almshhadany:2018:UMO

[AI18] Sarah E. Almshhadany and Laheeb M. Ibrahim. Using multiobjective artificial fish swarm algorithm to solve the software project scheduling problem. International Journal of Computer Applications, 181(??):6–13, September 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number16/29904-2018917753/.

AlMarouf:2019:IBS

[AIC19] Ahmed Al Marouf, Shariful Islam, and Narayan R. Chakraborty. IoT-based smart meeting room weather detection system using Arduino and relative sensors. International Journal of Computer Applications, 178(??):44–51, June 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number17/
30630-2019918990/.

Ajao:2014:IQS

[AICN14] Adewale A. Ajao, Dike U. Ike, Ndujiuba Charles, and John S. Ndueso. Improvement of quality of service (QoS) over a wide area network (WAN) using multiprotocol label switching traffic engineering (MPLS-TE). International Journal of Computer Applications, 88(??):14–18, February 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume88/number10/15387-3736/.

Alyahya:2012:MVA

[AIG12] Sultan Alyahya, Wendy K. Ivins, and W. A. Gray. Managing versioning activities to support tracking progress of distributed agile teams. International Journal of Computer Applications, 39(??):12–19, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number16/4903-7405/.

Abbas:2013:OEC

[AII13] Nizar Hadi Abbas, Tarik Zeyad Ismaeel, and Rassim Nooraldin Ibrahim. Optimization of energy consumption in wireless sensor networks based on nature-inspired algorithms. International Journal of Computer Applications, 77(??):32– 39, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number14/13553-1345/.

Aziz:2011:TPE

[AIK11] Md. Tariq Aziz, Mohammad Saiful Islam, and Md. Nazmul Islam Khan. Throughput performance evaluation of video/voice traffic in IPv4/IPv6 networks. International Journal of Computer Applications, 35(??):5–12, December 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume35/number2/4371-
6033/.

Amini:2011:FPI

[AIO11] Bahram Amini, Roliana Ibrahim, and Mohd Shahizan Othman. A framework for personalized information integration in higher education institutes. International Journal of Computer Applications, 23(??):6–12, June 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume23/number4/2879-3749/.

Anwar:2018:ISA

[AIP18] Nuril Anwar, Khoiriyah Isni, and Lovandri Dwanda Putra. Intelligent system application of health consultation based on Web and Android apps “SIDIKA”. International Journal of Computer Applications, 181(??):6–11, November 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume181/number27/30106-
2018918046/.

Alfurati:2019:DCF

[AIRR19] Israa Subri Alfurati, Bayadir A. Issa, Osama T. Rashid, and Abdulmuttalib T. Rashid. Design and construction a falling water digital display system. International Journal of Computer Applications, 177(??):16–24, November 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number16/30982-2019919549/.

Agarkar:2010:PDP

[AJ10] A. M. Agarkar and Prajakta Joharapurkar. PMD & DGD performance analysis in SMF due to fiber irregularities. International Journal of Computer Applications, 12(??):6– 11, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume12/
number6/1685-2246/.

Abdullahi:2013:EFM

[AJ13] Ibrahim Abdullahi and S. B. Junaidu. Empirical framework to mitigate problems in longer job first scheduling algorithm LJF+CBT. International Journal of Computer Applications, 75(??):9–14, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume75/number14/13178-0763/.

Abdulsattar:2014:DBS

[AJ14] Mahmood A. K. Abdulsattar and Raed Sattar Jebur. DSP-based synchronization algorithm implementation for OFDM. International Journal of Computer Applications, 107(??):22–29, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number9/18780-0104/.

Abbasi:2015:PBC

[AJ15] Zeeshan A. Abbasi and Zainul Abdin Jaffery. Pilot based channel estimation and SCGNN based equalization for MIMO-OFDM system. International Journal of Computer Applications, 131(??):11–15, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume131/number3/23428-2015907205/.

Alfarjat:2016:SSI

[AJ16] Ahmed Hweishel A. Alfarjat and Hanumanthappa J. A survey over the security issues of Bluetooth using elliptic curve cryptography. International Journal of Computer Applications, 150(??):14–17, September 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume150/number8/26112-2016911591/.

Abouhawwash:2018:KPM

[AJ18] Mohamed Abouhawwash and M. A. Jameel. KKT proximity measure versus augmented achievement scalarization function. International Journal of Computer Applications, 182(??):1–7, October 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number24/30078-2018917986/.

Alsammak:2019:EES

[AJ19] Ahmed Nasser Alsammak and Sarmad Arkan Janderma. Enhancement effects of the STATCOM on the distance relay protection. International Journal of Computer Applications, 182(??):10–14, February 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number40/30356-2019918461/.

Abderrahim:2013:IMV

[AJB13] Zeineb Abderrahim, Hanen Jeder, and Mohamed Salim Bouhlel. Interactive multiresolution visualization of 3D mesh. International Journal of Computer Applications, 67 (??):33–39, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number14/11465-7074/.

Anzola:2014:EQT

[AJC14] John Petearson Anzola, Andrés Camilo Jiménez, and Juan Camilo Cuellar. Evaluation of QoS in the transmission of video H.264/SVC for ad hoc networks of two jumps. International Journal of Computer Applications, 108(??):1–8, December 2014. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume108/number19/
19016-0539/.

Ali:2015:ASR

[AJI15] Hazrat Ali, An Jianwei, and Khalid Iqbal. Automatic speech recognition of Urdu digits with optimal classification approach. International Journal of Computer Applications, 118(??):1–5, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number9/20770-3275/.

Alizdeh:2015:WAM

[AJM^{+}15] Mohamad Javad Alizdeh, Pejman Mohammadnia Joneyd, Meysam Motahhari, Farid Ejlali, and Hamed Kiani. A wavelet-ANFIS model to estimate sedimentation in dam reservoir. International Journal of Computer Applications, 114(??):19–25, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume114/number9/20007-1958/.

Al-Jammas:2018:IIF

[AJN18] Mohamed Al-Jammas and Rusul Nabeel. IBBB inter frame prediction of H.264/AVC. International Journal of Computer Applications, 181(??):1–4, September 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number14/29887-2018917575/.

Agrawal:2011:MBP

[AJSG11] Sudhir Agrawal, Sanjeev Jain, Sanjeev Sharma, and Roopam Gupta. Mobility based performance analysis of AODV and DYMO under varying degree of node misbehavior. International Journal of Computer Applications, 30(??):36– 41, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume30/
number7/3652-5104/.

Abdulsada:2017:IOD

[AJT17] Mohammed Abdulla Abdulsada, Mohanad Azeez Joodi, and Firas M. Tuaimah. Investigation of one day ahead load forecasting for Iraqi power system. International Journal of Computer Applications, 163(??):24–29, April 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume163/number1/27360-2017913450/.

Aslam:2011:CAC

[AK11] Farhad Aslam and Gagandeep Kaur. Comparative analysis of conventional, p, PI, PID and fuzzy logic controllers for the efficient control of concentration in CSTR. International Journal of Computer Applications, 17(??): 12–16, March 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume17/
number6/2226-2836/.

Al-khafaji:2013:ICB

[Ak13] Ghadah Al-khafaji. Image compression based on quadtree and polynomial. International Journal of Computer Applications, 76(??):31–37, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume76/number3/13229-0658/.

Aly:2014:ACS

[AK14] Walid Mohamed Aly and Hany Atef Kelleny. Adaptation of cuckoo search for documents clustering. International Journal of Computer Applications, 86(??):4– 10, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number1/14947-3041/.

Archana:2015:MHA

[AK15] P. Archana and Subhash S Kulkarni. Meta-heuristic approach for resource optimization in mobile real time video traffic. International Journal of Computer Applications, 110(??):30–36, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number14/19387-0928/.

Alazab:2016:NSM

[AK16] Ammar Alazab and Ansam Khresiat. New strategy for mitigating of SQL injection attack. International Journal of Computer Applications, 154(??):1–10, November 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume154/number11/26532-
2016911974/.

Aboalela:2018:CAK

[AK18a] Rania Aboalela and Javed Khan. Computation of assessing the knowledge in one domain by using cognitive skills levels. International Journal of Computer Applications, 180(??):18–25, January 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number12/28914-2018916241/.

Ahire:2018:CCW

[AK18b] D. D. Ahire and G. K. Kharate. Corner cut wide band microstrip patch antenna for biomedical applications. International Journal of Computer Applications, 180(??):12–16, January 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number15/28936-2018916283/.

Alshemali:2019:TMA

[AK19] Basemah Alshemali and Jugal Kalita. Toward mitigating adversarial texts. International Journal of Computer Applications, 178(??):1–7, September 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number50/30888-2019919384/.

Amankwah:2018:TFS

[AKA^{+}18] Richard Amankwah, Patrick Kwaku Kudjo, Beatrice Korkor Agyemang, Kofi Mensah, Bright Brew, and Samuel Yeboah Antwi. A theoretical framework for software vulnerability detection based on cascaded refinement network. International Journal of Computer Applications, 182(??):12–15, November 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number25/30131-
2018918078/.

Al-khayat:2014:WPC

[AkAA14] Omar N. Al-khayat, Siddeeq Y. Ameen, and Mohammed N. Abdallah. WSNs power consumption reduction using clustering and multiple access techniques. International Journal of Computer Applications, 87(??):33–39, February 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume87/number9/15239-3780/

.

Ashraf:2016:PQI

[AKB16] Saima Ashraf, Anita Khosla, and Abdul Hamid Bhat. Power quality improvement using voltage source converter (VSC) based unified power quality conditioner (UPQC). International Journal of Computer Applications, 148(??): 21–27, August 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume148/number1/25722-2016910981/.

Al-ketbi:2013:SAV

[AkC13] Omar Al-ketbi and Marc Conrad. Supervised ANN vs. unsupervised SOM to classify EEG data for BCI: Why can GMDH do betterh. International Journal of Computer Applications, 74(??):37–44, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume74/number4/12876-9901/.

Akgun:2013:PEP

[Akg13] Devrim Akgun. Performance evaluations for parallel image filter on multi-core computer using Java threads. International Journal of Computer Applications, 74(??): 13–19, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number11/12928-9836/.

Almuhammad:2020:DRP

[AKG20] Kafaa Almuhammad, Souheil Khawatmi, and Marwan Ghabally. Developing a routing protocol in wireless body area networks (WBANs) using multi-criteria decision making (MCDM) methods. International Journal of Computer Applications, 177(??):26–34, January 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number33/31115-2020919814/.

Aghajanzadeh:2015:CME

[AKH15] Nasser Aghajanzadeh and Alireza Keshavarz-Haddad. A concise model to evaluate security of SCADA systems based on security standards. International Journal of Computer Applications, 111(??):1–9, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number14/19603-1450/.

Agarkar:2011:LPC

[AKK11] Seema Ajay Agarkar, K. D. Kulat, and R. V. Kshirsagar. Low power considerations and timing simulation analysis of recent 8-bit embedded controllers including AT89C5130A/ 31A-M-frozen-in-idle-state MCUs for WSN applications. International Journal of Computer Applications, 15(??):25– 36, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume15/
number5/1944-2599/.

Ajay:2013:NCB

[AKN13] Amarendra Kumar Ajay, Swapna S Kolhatkar, and Atul B Naik. Need of cloud based system for higher & technical education in India. International Journal of Computer Applications, 63(??):38–41, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume63/number12/10522-5502/.

Kester:2015:HCE

[aKNP^{+}15] Quist aphetsi Kester, Laurent Nana, Anca Christine Pascu, Sophie Gire, Jojo M. Eghan, and Nii Narku Quaynor. A hybrid cryptographic encryption technique for securing surveillance digital images in mobile urban situation awareness systems. International Journal of Computer Applications, 119(??):16–22, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number7/21079-3756/.

Arefin:2013:SCI

[AKS^{+}13] Md. Gauhar Arefin, Mithun Kumar P K, Abu Sayem, Md Delowar Hossain, and Md Motiur Rahman. Segmentation of CT image using MAP-model and simulation annealing. International Journal of Computer Applications, 61 (??):45–48, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number2/9904-4492/.

Ayinde:2016:EIT

[AKS16] Kayode Ayinde, John Olatunde Kuranga, and Gbenga Sunday Solomon. Empirical investigation of type 1 error rate of univariate tests of normality. International Journal of Computer Applications, 148(??):24–31, August 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume148/number8/25778-2016911253/.

Al-khalidi:2015:TRI

[AKSEB15] Farah Al-Khalidi, Reza Saatchi, Heather Elphick, and Derek Burke. Tracing the region of interest in thermal human face for respiration monitoring. International Journal of Computer Applications, 119(??):42–46, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume119/number4/21059-3716/.

AlMasud:2013:SAS

[Al 13] Shah Murtaza Rashid Al Masud. Study and analysis of scientific scopes and issues towards developing an efficient LECIM. International Journal of Computer Applications, 74(??):46–56, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number6/12893-0061/.

AlBdour:2014:EBI

[Al 14] Nashat G. Al Bdour. Encoding binary images using cellular automata for data compression. International Journal of Computer Applications, 92(??):25–30, April 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume92/number4/15998-4965/.

AlSadi:2016:TON

[Al 16] Ghania Al Sadi. Tuning and optimizing network file system server performance. International Journal of Computer Applications, 134(??):25–29, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number10/23952-2016908130/.

Arora:2012:HIB

[ALA12] Shweta Arora, Narendra D. Londhe, and Anuja Kumar Acharya. Human identification based on iris recognition for distant images. International Journal of Computer Applications, 45(??):32–39, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number16/6866-9470/.

Alam:2016:DNM

[Ala16] Tauhidul Alam. Decentralized and nondeterministic multirobot area patrolling in adversarial environments. International Journal of Computer Applications, 156(??): 1–8, December 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume156/number2/26678-2016912367/.

Alberg:2015:ITA

[Alb15] Dima Alberg. An interval tree approach to predict forest fires using meteorological data. International Journal of Computer Applications, 132(??):17–22, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume132/number4/23582-2015907398/.

Algamluoli:2019:THD

[Alg19] Ammar Falah Algamluoli. Total harmonic distortion of single phase mid-point cycloconverter comparison between MOSFET, IGBT and GTO power devices. International Journal of Computer Applications, 181(??):10–15, January 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume181/number35/30257-
2019918300/.

Alharbi:2012:MOH

[Alh12] Fahd Alharbi. Multi objectives heuristic algorithm for grid computing. International Journal of Computer Applications, 46(??):39–45, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number18/7045-9691/.

Aminu:2024:EGW

[ALH^{+}24] Jafar Aminu, Rohaya Latip, Zurina Mohd Hanafi, Shafinah Kamarudin, Bashar Umar Kangiwa, and Ayuba Liman. An enhanced grey wolf optimization algorithm for efficient task scheduling in mobile edge computing. International Journal of Computer Applications, 186(??):39–44, December 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number56/
an-enhanced-grey-wolf-optimization-algorithm-for-
efficient-task-scheduling-in-mobile-edge-computing/

.

Ali:2012:HSP

[Ali12] Ikbal Ali. High-speed peer-to-peer communication based protection scheme implementation and testing in laboratory. International Journal of Computer Applications, 38 (??):16–24, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume38/
number4/4676-6793/.

Ali:2016:PKE

[Ali16] Fairouz Sher Ali. Public key encryption with ’fixed and short length’ keyword search. International Journal of Computer Applications, 154(??):1–8, November 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume154/number4/26476-2016912086/.

Ali:2019:SBD

[Ali19] Abdul-Jabbar Fathel Ali. Study the behavior of the distribution network with the application of Statcom device. International Journal of Computer Applications, 181(??):36–44, March 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume181/number44/
30415-2019918553/.

Abou-Loukh:2017:CPA

[ALO17] Sadiq J. Abou-Loukh and Arwa Ra’ad Obaid. Comparison of preprocessing algorithms using an affordable EEG headset. International Journal of Computer Applications, 160(??):25–31, February 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume160/number1/27039-2017912949/.

Alsammak:2018:FLB

[Als18] Ahmed Nasser B. Alsammak. Fuzzy logic based power factor control of synchronous machine. International Journal of Computer Applications, 182(??):33–42, September 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number14/29934-
2018917812/.

AlShammari:2023:IKE

[AlS23a] Ahmad Farhan AlShammari. Implementation of keyword extraction using term frequency-inverse document frequency (TF-IDF) in Python. International Journal of Computer Applications, 185(??):9–14, September 2023. CODEN ????

ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number35/32916-2023923137/.

AlShammari:2023:ITS

[AlS23b] Ahmad Farhan AlShammari. Implementation of text similarity using word frequency and cosine similarity in Python. International Journal of Computer Applications, 185(??):54–59, October 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number36/32927-2023923160/.

AlShammari:2024:ICUa

[AlS24a] Ahmad Farhan AlShammari. Implementation of classification using K-nearest neighbors (KNN) in Python. International Journal of Computer Applications, 186(??):19–24, July 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number33/
implementation-of-classification-using-k-nearest-
neighbors-knn-in-python/.

AlShammari:2024:ICUb

[AlS24b] Ahmad Farhan AlShammari. Implementation of clustering using K-means in Python. International Journal of Computer Applications, 186(??):12–17, September 2024. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.
org/archives/volume186/number40/implementation-of-
clustering-using-k-means-in-python/.

AlShammari:2024:ICF

[AlS24c] Ahmad Farhan AlShammari. Implementation of curve fitting using polynomial regression in Python. International Journal of Computer Applications, 186(??):27–32, January 2024. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume186/number6/33076-2024923400/

.

AlShammari:2024:IFS

[AlS24d] Ahmad Farhan AlShammari. Implementation of feature selection using correlation matrix in Python. International Journal of Computer Applications, 186(??):29–34, December 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number58/
implementation-of-feature-selection-using-correlation-
matrix-in-python/.

AlShammari:2024:ILRa

[AlS24e] Ahmad Farhan AlShammari. Implementation of linear regression using least squares and gradient descent in Python. International Journal of Computer Applications, 186(??):52–57, February 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number9/implementation-of-linear-regression-
using-least-squares-and-gradient-descent-in-python/

.

AlShammari:2024:ILRb

[AlS24f] Ahmad Farhan AlShammari. Implementation of logistic regression using gradient descent in Python. International Journal of Computer Applications, 186(??):41–46, March 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number13/
implementation-of-logistic-regression-using-gradient-
descent-in-python/.

AlShammari:2024:IME

[AlS24g] Ahmad Farhan AlShammari. Implementation of model evaluation using confusion matrix in Python. International Journal of Computer Applications, 186(??):42–48, November 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number50/
implementation-of-model-evaluation-using-confusion-
matrix-in-python/.

AlShammari:2024:IMR

[AlS24h] Ahmad Farhan AlShammari. Implementation of multiple regression using least squares and gradient descent in Python. International Journal of Computer Applications, 186(??):4–9, May 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number21/
implementation-of-multiple-regression-using-least-
squares-and-gradient-descent-in-python/.

AlShammari:2024:IPR

[AlS24i] Ahmad Farhan AlShammari. Implementation of polynomial regression using least squares and gradient descent in Python. International Journal of Computer Applications, 186(??):20– 26, June 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number27/
implementation-of-polynomial-regression-using-least-
squares-and-gradient-descent-in-python/.

Altameem:2011:CML

[Alt11] T. Altameem. Contextual mobile learning system for Saudi Arabian universities. International Journal of Computer Applications, 21(??):21–26, May 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume21/number4/2499-3377/.

Altuhafi:2013:RPP

[Alt13] Ammar Waysi Altuhafi. A review on peer-to-peer live video streaming topology. International Journal of Computer Applications, 68(??):6–14, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number5/11573-6881/.

Aly:2013:ECS

[Aly13] Walid M. Aly. Evaluation of cuckoo search usage for model parameters estimation. International Journal of Computer Applications, 78(??):1–6, September 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume78/number11/13530-1072/.

Alzohairy:2012:NIM

[Alz12] T. A. A. Alzohairy. Neural internal model control for tracking unknown nonaffine nonlinear discrete-time systems under external disturbances. International Journal of Computer Applications, 40(??):19–26, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume40/number6/4960-7217/.

Aramudhan:2010:NSC

[AM10] M. Aramudhan and K. Mohan. New secure communication protocols for mobile e-health system. International Journal of Computer Applications, 8(??):10–15, October 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume8/number4/1201-1713/.

Abdelfattah:2011:MAH

[AM11] Eman Abdelfattah and Ausif Mahmood. A multi-algorithm, high reliability, extensible steganalyzer using services oriented architecture. International Journal of Computer Applications, 27(??):18–26, August 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume27/number5/3297-4503/.

Abdullah-al-mamun:2012:TES

[am12a] Md. Abdullah al mamun. Technology in educational settings in the Government Polytechnic Institutes of Bangladesh: a critical analysis. International Journal of Computer Applications, 54(??):32–40, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume54/number13/8628-2502/.

Amalarethinam:2012:EDG

[AM12b] D. I. George Amalarethinam and P. Muthulakshmi. Enhanced DAGitizer for grid computing through the discovery of least cost path. International Journal of Computer Applications, 55(??):1–7, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume55/number4/8740-2612/.

Ahmed:2013:ISD

[AM13a] Alaa H. Ahmed and Mohammad Mikki. Improved spam detection using DBSCAN and advanced digest algorithm. International Journal of Computer Applications, 69(??): 11–16, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number25/12126-8300/.

Al-marakeby:2013:MCP

[Am13b] A. Al-marakeby. Multi core processors for camera based OMR. International Journal of Computer Applications, 68 (??):1–5, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume68/
number13/11636-7116/.

Aulakh:2014:ICP

[AM14] Areet Aulakh and Kulwinder Singh Malhi. To investigate the characteristics parameters of semiconductor optical amplifier based on wavelength converters for all optical networks. International Journal of Computer Applications, 95 (??):37–41, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number19/16706-6856/.

Abada:2015:IRS

[AM15] Driss Abada and Abdellah Massaq. Improving relay selection scheme for connecting VANET to Internet over IEEE 802.11p. International Journal of Computer Applications, 132(??):1–8, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number2/23563-2015907297/.

Ahmed:2017:SFA

[AM17a] Zonayed Ahmed and Maliha Mahbub. SYN flood attack prevention using particle swarm optimization in cloud computing environment. International Journal of Computer Applications, 177(??):27–33, November 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number1/28593-2017915661/.

AlTuarihi:2017:ENI

[AM17b] Muna Hameed Al Tuarihi and Ala’a Abbas Mahdi. Experimental and numerical investigation on air flow and temperature distribution under different geometries of displacement ventilation devices in an office room. International Journal of Computer Applications, 169(??):15–26, July 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume169/number4/27973-2017914652/.

ALgamluoli:2017:CBC

[AM17c] Ammar ALgamluoli and Muatafa Falah Mahmood. Controller for boost converter in S-domain and Z-domain by using SISOTOOL MATLAB. International Journal of Computer Applications, 174(??):3–7, September 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume174/number9/28433-2017915244/.

Adigun:2019:MPB

[AM19a] Taiwo Adigun and Angela Makolo. Multilayer perceptron based model of large-scale gene regulatory network. International Journal of Computer Applications, 178(??):6–15, August 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number42/
30815-2019919148/.

Aiman:2019:VSU

[AM19b] Fariha Aiman and G. R. Manjula. Video steganography using convolutional neural network and temporal residual method. International Journal of Computer Applications, 178(??):24–29, September 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number46/30860-2019919375/.

Ahmad:2014:DTI

[AMA14a] Abusukhon Ahmad, Talib Mohammad, and Hani Mahmoud Almimi. Distributed text-to-image encryption algorithm. International Journal of Computer Applications, 106(??):1–5, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number1/18481-9518/.

AlMamun:2014:ETU

[AMA14b] Md. Al Mamun, Md. Nazrul Islam Mondal, and Boshir Ahmed. Evaluating temporal uncertainty of multi-temporal images for geographical deviance. International Journal of Computer Applications, 103(??):14–18, October 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume103/number14/18141-
9339/.

Alwadan:2014:PCS

[AMA14c] Tariq Alwada’n, Adel Hamdan Mohammad, and Hamza Aldabbas. Privacy in cloud system. International Journal of Computer Applications, 93(??):1–4, May 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume93/number2/16184-2255/.

Al-Mutar:2015:QCA

[AMA15] Wissam H. Al-Mutar and Turki Y. Abdalla. Quarter car active suspension system control using fuzzy controller tuned by PSO. International Journal of Computer Applications, 127(??):38–43, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number2/22704-2015906334/.

Algabalawy:2015:OMC

[AMAB15] M. Algabalawy, S. F. Mekhamer, A. Y. Abdelaziz, and M. A. L. Bader. Optimal multi-criteria design of hybrid power generation systems: a new contribution. International Journal of Computer Applications, 129(??): 13–24, November 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume129/number2/23044-2015906819/.

Al-maqaleh:2012:EAA

[AmAdS12] Basheer M. Al-maqaleh, Mohammed A. Al-dohbai, and Hamid Shahbazkia. An evolutionary algorithm for automated discovery of small-disjunct rules. International Journal of Computer Applications, 41(??):33–37, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume41/number8/5563-7643/.

Al-Manaseer:2019:SBA

[AMAHH19] Esraa Salem Al-Manaseer, Hind Mousa Al-Hamadeen, and Abdelaziz I. Hammouri. A swarm-based algorithm for solving economic load dispatch problem. International Journal of Computer Applications, 177(??):11–18, December 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume177/number26/31060-
2019919719/.

Al-mahmood:2014:LIC

[AmAr14] Haider Al-mahmood and Zainab Al-rubaye. Lossless image compression based on predictive coding and bit plane slicing. International Journal of Computer Applications, 93(??):1–6, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number1/16176-2068/.

Al-Mekhlafy:2020:TSF

[AMAY20] Muneer Abdullah Saeed Hazaa Al-Mekhlafy, Muneef Abdulkarim Farea Ahmed, and Mohamed Ghaleb Yosuef. Time series forecasting: a hybrid neuro-fuzzy-particle swarm optimization model. International Journal of Computer Applications, 176(??):1–8, July 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number40/31464-2020920457/.

Anjomshoa:2011:EDO

[AMB11] Mehdi Anjomshoa, Ali Mahani, and Mostafa Esmaeil Beig. Evolutionary design and optimization of digital circuits using imperialist competitive algorithm. International Journal of Computer Applications, 32(??):14–19, October 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume32/number1/3868-
5404/.

Ambasht:2024:LAA

[Amb24] Anshumali Ambasht. Leveraging AI and analytics in climate science: Enhancing predictions and sustainability practices. International Journal of Computer Applications, 186(??):10–16, June 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number24/leveraging-ai-and-analytics-in-
climate-science-enhancing-predictions-and-sustainability-
practices/.

AlMamunShohag:2011:OAI

[AMBM11] Shoyeb Al Mamun Shohag and Mozammel-Bin-Motalab. An observation and analysis of IPTV and multicasting traffic. International Journal of Computer Applications, 30(??): 1–6, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume30/
number1/3610-5023/.

Abdellah:2016:PPA

[AMBS16] Mohamed Refaat Abdellah, H. Aboelseoud M, Khalid Shafee Badran, and M. Badr Senousy. Privacy preserving association rule hiding techniques: Current research challenges. International Journal of Computer Applications, 136(??):11–17, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume136/number6/24156-2016908446/.

Alomari:2012:BBD

[AMG^{+}12] Esraa Alomari, Selvakumar Manickam, B. B. Gupta, Shankar Karuppayah, and Rafeef Alfaris. Botnet-based distributed denial of service (DDoS) attacks on Web servers: Classification and art. International Journal of Computer Applications, 49(??):24–32, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume49/number7/7640-0724/.

Agarwal:2013:FLS

[AMG13] Ratish Agarwal, Mahesh Motwani, and Roopam Gupta. A framework for load sharing in clustered ad hoc networks. International Journal of Computer Applications, 78(??):31– 37, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume78/
number8/13512-1284/.

Azami:2011:ISS

[AMH11] Hamed Azami, Karim Mohammadi, and Hamid Hassanpour. An improved signal segmentation method using genetic algorithm. International Journal of Computer Applications, 29(??):5–9, September 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume29/number8/3586-4967/.

Abbas:2013:PSS

[AmH13] Nizar Hadi Abbas and Ruba Al mula Humadi. Power system stability enhancement using SVC with modified PSO tuned PID controller. International Journal of Computer Applications, 71(??):15–22, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number3/12337-8610/.

Ahmad:2013:RDH

[AMHAh13] Mostafa A. Ahmad, Aly M. Meligy, Amal H. Hashim, and Ayoub K. Al-hamadi. Reversible data hiding by integer wavelet transform with lossless EZW bit-stream. International Journal of Computer Applications, 66(??): 8–15, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number2/11054-5952/.

Aziz:2016:LCC

[AMJ16] Najam Aziz, Nasru Minallah, and Ahmad Junaid. Land cover classification and crop estimation using remote sensing techniques. International Journal of Computer Applications, 155(??):1–6, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume155/number2/26574-2016912259/.

Aydi:2012:ACE

[AMK12] Walid Aydi, Nouri Masmoudi, and Lotfi Kamoun. Active contour without edges vs GVF active contour for accurate pupil segmentation. International Journal of Computer Applications, 54(??):25–36, September 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume54/number4/8555-2118/.

Alam:2013:EBS

[AMK13] Kazi Md. Rokibul Alam, Saifuddin Mahmud, and Mohammad Nazmul Alam Khan. An exploitation of blind signature schemes to simulate privacy-related applications. International Journal of Computer Applications, 66(??): 23–27, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number3/11065-5980/.

Ali:2020:HRR

[AMK20] Khaji Faizan Ali, Abhijeet Marikal, and Kakelli Anil Kumar. A hybrid round robin scheduling mechanism for process management. International Journal of Computer Applications, 177(??):14–19, February 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number36/31139-2020919851/.

Al-mahmud:2012:IBA

[AmM12] Abdullah Al-mahmud and Matei Ciobanu Morogan. Identitybased authentication and access control in wireless sensor networks. International Journal of Computer Applications, 41(??):18–24, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume41/
number13/5602-7858/.

Ayorinde:2021:RBW

[AMM21] Ayotunde Ayorinde, Hisham Muhammed, and A. Ike Mowete. On the response of basic Walfisch–Ikegami and Walfisch–Bertoni models to QMM calibration. International Journal of Computer Applications, 174(??):53–64, February 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume174/number18/31781-
2021921100/.

Adeel:2012:PBA

[AMN12] Ahsan Adeel, Shahid Mehmood, and Mian Hammad Nazir. PAPR and bandwidth analysis of SISO-OFDM/WOFDM and MIMO OFDM/WOFDM (WiMAX) for multi-path fading channels. International Journal of Computer Applications, 44(??):38–44, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume44/number22/6414-8891/.

Adami:2015:MDOb

[AMN15a] Amirhossein Adami, Mahdi Mortazavi, and Mehran Nosratollahi. Multidisciplinary design optimization of a deorbit maneuver considering propulsion, TPS, and trajectory. International Journal of Computer Applications, 116(??):1–6, April 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume116/number23/
20496-2109/.

Adami:2015:MDOa

[AMN15b] Amirhossein Adami, Mahdi Mortazavi, and Mehran Nosratollahi. Multidisciplinary design optimization of hydrogen peroxide monopropellant propulsion system using GA and SQP. International Journal of Computer Applications, 113(??):14–21, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume113/number9/19853-1774/.

Adewale:2018:RRD

[AMNA18] Adeyinka A. Adewale, Victor O. Matthews, Charles N. Ndujiuba, and Adeyemi M. Adenrele. Reduction of routing delay in an enterprise network using dynamic multipoint private network. International Journal of Computer Applications, 179(??):1–6, January 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number9/28825-2018915427/.

Afkhami:2013:BHS

[AMS13] Sepideh Afkhami, Omid R. Ma’rouzi, and Ali Soleimani. A binary harmony search algorithm for solving the maximum clique problem. International Journal of Computer Applications, 69(??):38–43, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume69/number12/11897-7956/.

Aziz:2014:CAP

[AMS14] Nurulnadwan Aziz, Ariffin Abdul Mutalib, and Siti Mahfuzah Sarif. Critical analysis in proposing a conceptual design model of assistive courseware for low vision (AC4LV) learners. International Journal of Computer Applications, 92(??):18–25, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume92/
number10/16044-5173/.

Alzohairy:2016:IMB

[AMS16] Tamer A. A. Alzohairy, Emad El-Dein H. A. Masameer, and Mahmoud S. Sayed. Image mosaicing based on neural networks. International Journal of Computer Applications, 136(??):25–31, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume136/number1/24118-2016908338/.

Abdelfatah:2013:FGC

[AMSS13] Alaa Abdelfatah, Sahar A. Mokhtar, Alaa Sheta, and Basma Solaiman. Forecast global carbon dioxide emission using swarm intelligence. International Journal of Computer Applications, 77(??):1–5, September 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume77/number12/13443-1343/.

A:2017:ECS

[AMT^{+}17] Sharaaf N A, Hijaz A J M, Kiroshan T, Suresh R, and S. G. S. Fernando. Easy clean — a smart solution for garbage finding and collecting. International Journal of Computer Applications, 169(??):1–7, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume169/number3/27962-2017914617/.

Attar:2011:NIC

[AMTM11] S. F. Attar, M. Mohammadi, and R. Tavakkoli-Moghaddam. A novel imperialist competitive algorithm to solve flexible flow shop scheduling problem in order to minimize maximum completion time. International Journal of Computer Applications, 28(??):27–32, August 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume28/number10/3426-4781/.

Alifo:2023:WAV

[AMY23] Ferdinand Alifo, Doe Martin, and Mustapha Awinsongya Yakubu. Wormhole attack vulnerability assessment of MANETs: Effects on routing protocols and network performance. International Journal of Computer Applications, 185(??):24–29, December 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number48/33015-2023923312/.

A:2014:FBI

[AN14] Meenakshi Sundaram A and C. Nandini. Feature based image authentication using symmetric surround saliency mapping in image forensics. International Journal of Computer Applications, 104(??):43–51, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number13/18266-9289/.

A:2016:TMP

[AN16a] Deepak N A and Sinha U N. Topic model for person identification using gait sequence analysis. International Journal of Computer Applications, 133(??):1–6, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number7/23795-2016907892/.

Al-Najjar:2016:EHC

[AN16b] Hussein M. Taqi Al-Najjar. Electric, heating and cooling yields of solar collectors for different atmospheric conditions and tilt angles. International Journal of Computer Applications, 141(??):1–10, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume141/number10/24817-2016909805/.

Asadi:2011:NHA

[ANA11] Alireza Arab Asadi, Ali Naserasadi, and Zeinab Arab Asadi. A new hybrid algorithm for traveler salesman problem based on genetic algorithms and artificial neural networks. International Journal of Computer Applications, 24 (??):6–9, June 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume24/
number5/2945-3926/.

Anatory:2015:ESV

[Ana15] Justinian Anatory. The establishment of the School of Virtual Education at the University of Dodoma in Tanzania: Personal experience and reflections. International Journal of Computer Applications, 115(??):1–10, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume115/number16/20232-2164/.

Ami-narh:2014:ABF

[AnAA14] James Ami-narh, Lawrence K Aziale, and Albert Akanferi. The adoption of biometric fingerprint timekeeping technology in the Ghanaian business community — effectiveness and impact. International Journal of Computer Applications, 85(??):32–39, January 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume85/number9/14871-3246/.

Ali:2020:FTQ

[ANAH20] Ali A. Majeed Ali, Osama Abdulhak M. Nasher, and Ahmed Sultan Al-Hegami. A framework for total quality management of diesel generator fuel consumption using machine learning and Internet of Things (IoT). International Journal of Computer Applications, 176(??):43–52, May 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume176/number22/
31334-2020920234/.

Al-Najjar:2017:EPH

[ANAK17] Hussein M. Taqi Al-Najjar and Akram A. Abbood Al-Khazzar. Experimental prediction of hourly diffuse solar radiation with clearness index in Baghdad (Iraq). International Journal of Computer Applications, 158(??): 20–28, January 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume158/number7/26921-2017912849/.

Abouda:2013:PSM

[ANE^{+}13] Salim Abouda, Frederic Nollet, Najib Essounbouli, Abdessattar Chaari, and Yassine Koubaa. PID and sliding mode control-SMC for a motor-pump voltage in a photovoltaic-PV system based on MPPT controller. International Journal of Computer Applications, 73(??):44–49, July 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume73/number3/12725-9952/.

Anh:2014:IDA

[Anh14a] Dao Nam Anh. Image denoising by addaptive non-local bilatetal filter. International Journal of Computer Applications, 99(??):4–10, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume99/number12/17423-8275/.

Anh:2014:IBF

[Anh14b] Dao Nam Anh. Iterative bilateral filter and non-local mean. International Journal of Computer Applications, 106(??):33–38, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number11/18566-9818/.

Anh:2015:SIC

[Anh15a] Dao Nam Anh. Segmentation by incremental clustering. International Journal of Computer Applications, 111(??):23–30, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number12/19591-1360/.

Anh:2015:SCB

[Anh15b] Dao Nam Anh. Smooth context based color transfer. International Journal of Computer Applications, 116(??):29–37, April 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume116/number15/
20413-2825/.

Ahmed:2017:RSE

[ANM17] Basaeir Y. Ahmed, Safa A. Najim, and Widad A. Mansour. Role of software engineering in visualizing large volume of hyperspectral and medical data sets. International Journal of Computer Applications, 176(??):1–5, October 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number9/28582-2017915580/.

Ansari:2014:PAE

[Ans14] Nadia Mustaqim Ansari. Performance analysis of ETTH and FTTH in terms of cost and scalability: a comparative study carried out on MATLAB. International Journal of Computer Applications, 98(??):1–5, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume98/number11/17224-7546/.

Adeoti:2013:ETA

[AO13a] Olatunde A. Adeoti and Peter A. Osanaiye. Effect of training algorithms on the performance of ANN for pattern recognition of bivariate process. International Journal of Computer Applications, 69(??):8–12, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume69/number20/12085-8031/.

Anwar:2013:DSS

[AO13b] Muhammad Naveed Anwar and Michael Philip Oakes. Decision support system for the selection of an ITE or a BTE hearing aid. International Journal of Computer Applications, 76(??):37–42, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume76/number14/13318-0936/.

Asiligwa:2016:RAG

[AO16] Bonface I. Asiligwa and Elijah I. Omwenga. A roadmap for the adoption of government e-payments in Kenya. International Journal of Computer Applications, 144(??):8–18, June 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume144/number1/25141-2016910062/

.

Adeyelu:2013:MCJ

[AOAG13] Adekunle Adeyelu, Emmanuel Olajubu, Adesola Aderounmu, and Tivlumun Ge. A model for coordinating jobs on mobile wireless computational grids. International Journal of Computer Applications, 84(??):17– 24, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number13/14636-1637/.

A:2012:CSA

[AOB12] Oladeji F A, Oyetunji M O, and Okunoye O B. cWRR: a scheduling algorithm for maximizing the performance of quality of service network router. International Journal of Computer Applications, 41(??):30–34, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume41/number2/5515-7516/.

Altantawy:2014:NGF

[AOK14] Doaa A. Altantawy, Marwa Obbaya, and Sherif S. Kishk. New global formulation for a bilateral based stereo matching algorithm. International Journal of Computer Applications, 98(??):21–28, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume98/number8/17205-7419/.

Amutsama:2018:SRB

[AOK18] Wilson Ambale Amutsama, Isaiah Mulang Onando, and Andrew Mwaura Kahonge. A systematic RDF-based approach for structuring government open data to enhance accessibility. International Journal of Computer Applications, 182(??):11–17, November 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number28/30155-2018918152/.

Abbadi:2012:DIN

[AOM12] B. Abbadi, A. Oulefki, and M. Mostefai. Development and implementation of a new dynamic face detection operator. International Journal of Computer Applications, 49 (??):18–23, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number20/7887-1218/.

Adeyelu:2020:ETD

[AOO20] Adekunle A. Adeyelu, Onaji J. Onah, and Iwuese J. Orban. An enhanced-time difference of arrival technique for estimating mobile station position in wireless sensor networks. International Journal of Computer Applications, 175(??):5–11, October 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number22/31581-2020920181/.

Aly:2012:MBS

[AOSM12] Ola M. Aly, Houda M. Onsi, Gouda I. Salama, and Tarek A. Mahmoud. Multimodal biometric system using iris, palmprint and finger-knuckle. International Journal of Computer Applications, 57(??):1–6, November 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume57/number16/9195-3717/.

Awate:2015:SOD

[AP15] Rajesh B. Awate and Archana B. Patankar. Segmentation of optic disc and optic cup to calculate CDR using Kekre’s LUV color space for detecting glaucoma. International Journal of Computer Applications, 127(??):7–11, October 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume127/number17/
22819-2015906712/.

Amponsah:2016:EDM

[AP16] Anokye Acheampong Amponsah and Kwaku Agyepong Pabbi. Enhancing direct marketing using data mining: a case of YAA Asantewaa Rural Bank Ltd. in Ghana. International Journal of Computer Applications, 153(??):6–12, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume153/number7/26413-2016912092/.

Agrawal:2011:IWT

[APK11] Shorya Agrawal, Nirved K. Pandey, and Amit Kanskar. Implementation of WAP through an innovative and efficient technique. International Journal of Computer Applications, 36(??):22–27, December 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume36/number4/4482-6305/.

Apollos:2019:MTM

[Apo19] Ezeogu Apollos. Memristor theory and mathematical modelling. International Journal of Computer Applications, 178(??):1–8, June 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number27/30703-2019919089/.

Ardika:2018:TBI

[APS18] Ngurah Made Ardika, Nyoman Piarsa, and Arya Sasmita. Telegram bot integration with face recognition as smart home features. International Journal of Computer Applications, 182(??):42–47, September 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number13/29925-2018917778/.

Absar:2016:MCA

[APU16] Nurul Absar, Abhijit Pathak, and Md. Abbas Uddin. Multicriteria analysis for the best location selection in Chittagong city Area, Bangladesh. International Journal of Computer Applications, 143(??):10–18, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume143/number8/25096-2016910030/.

Al-quzwini:2014:DIF

[Aq14a] Mahmoud M. Al-quzwini. Design and implementation of a fiber to the home FTTH access network based on GPON. International Journal of Computer Applications, 92 (??):30–42, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume92/
number6/16015-5050/.

Alisa:2014:FBT

[AQ14b] Zainab T. Alisa and Sara Raad Qasim. A fuzzy based TCP congestion control for wired networks. International Journal of Computer Applications, 89(??):36– 42, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number4/15494-4401/.

Al-qdah:2015:HSS

[AQ15] Majdi Al-Qdah. Hybrid security system based on wavelet domain watermarking and chaotic map cryptography. International Journal of Computer Applications, 110(??):1–6, January 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume110/number13/
19373-9190/.

Adhikary:2010:ICB

[AR10] Munesh Chandra Adhikary and Ashanta Ranjan Routray. Image compression based on wavelet and quantization with optimal Huffman code. International Journal of Computer Applications, 5(??):6–9, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume5/number2/893-1267/.

Adautin:2015:FRA

[AR15] Esther D. Adautin and Nagoor Meeran A R. Forensic reconstruction and analysis of residual artifacts from portable Web browser. International Journal of Computer Applications, 128(??):19–24, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number18/22974-2015906741/.

Alkama:2016:FTT

[AR16] Meryem Alkama and Mostafa Rachik. Free terminal time in optimal control approach of Chikungunya model. International Journal of Computer Applications, 140(??):9–16, April 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume140/number13/24671-
2016909536/.

Alagarsamy:2017:EMM

[AR17a] Sumithra Alagarsamy and S. P. Rajagopalan. Exponentiated multiple message communication using certificateless signcryption for mobile network security. International Journal of Computer Applications, 178(??): 13–24, November 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume178/number7/28692-2017915705/.

Algharib:2017:CSV

[AR17b] Hanan M. S. Algharib and Shafqat Ur Rehman. A case study on various preprocessing methods and their impact on face recognition using random forest. International Journal of Computer Applications, 175(??): 5–14, October 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume175/number9/28562-2017915642/.

Arablou:2015:DAP

[Ara15] Fathollah Dadgar Arablou. Dynamic analysis of priority with linear decreasing function of a priority in the service system with four types of application. International Journal of Computer Applications, 124(??):6–12, August 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume124/number15/22178-2015905555/.

Abdulla:2011:SWA

[ARAA11] Shubair A. Abdulla, Sureswara Ramadass, Altyeb Altaher, and Amer Al Nassiri. Setting a worm attack warning by using machine learning to classify NetFlow data. International Journal of Computer Applications, 36(??):49– 56, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume36/
number2/4467-6258/.

Al-rousan:2013:NCB

[ArAn13] Nadia Al-rousan and Hazem Al-najjar. The network coding based on marked packets. International Journal of Computer Applications, 77(??):18–23, September 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume77/number6/13398-1046/.

Awasekar:2018:DFD

[ARDS18] Pranali Awasekar, Menaka Ravi, Shivani Doke, and Zaheed Shaikh. Driver fatigue detection and alert system using nonintrusive eye and yawn detection. International Journal of Computer Applications, 180(??):1–5, May 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number44/29438-2018917140/.

Ali:2010:SAC

[ARHERZ10] Md. Mohsin Ali, Mst. Shakila Khan Rumi, Md. Hasnat-E-Rabbi, and M. M. Zakariya. A selective aperiodic checkpointing approach for VLR failure restoration in wireless mobile networks. International Journal of Computer Applications, 11(??):10–16, December 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume11/number9/1612-2167/.

Arifial:2023:IAR

[Ari23] Ryu Arifial. Implementation of augmented reality technology as a media for introducing cultural diversity in Indonesia. International Journal of Computer Applications, 185(??):1–5, December 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number50/33027-2023923332/.

Ali:2013:SVC

[ARK13] Malik Ahsan Ali, Imran Rashid, and Adnan Ahmed Khan. Selection of VoIP CODECs for different networks based on QoS analysis. International Journal of Computer Applications, 84(??):38–44, December 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume84/number5/14575-2702/.

Amirullah:2016:FAC

[ARL16] Amirullah Amirullah, Imam Riadi, and Ahmad Luthfi. Forensics analysis from cloud storage client application on proprietary operating system. International Journal of Computer Applications, 143(??):1–7, June 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume143/number1/25038-2016907696/.

Ahuja:2016:DLS

[ARMS16] Raaj Ahuja, Mukesh Ramrakhyani, Bunty Manchundiya, and Sonal Shroff. Dual layer secured password manager using Blowfish and LSB. International Journal of Computer Applications, 143(??):5–10, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume143/number3/25055-2016910048/.

Abadi:2017:WFS

[ARP17] Anis Susila Abadi, Imam Riadi, and Yudi Prayudi. Wi-Fi security level analysis for minimizing cybercrime. International Journal of Computer Applications, 164(??):35–39, April 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume164/number7/27499-2017913667/

.

Azam:2020:ESV

[ARR20] Md. Shafiul Azam, Md. Abu Raihan, and Humayan Kabir Rana. An experimental study of various machine learning approaches in heart disease prediction. International Journal of Computer Applications, 175(??):16–21, September 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume175/number21/
31576-2020920741/.

Akram:2013:ILL

[ARS13] Sk. Mahammad Akram, V. Leela Rani, and K. Sailaja. Implementation of low leakage and high performance 8-bit ALU for low power digital circuits. International Journal of Computer Applications, 82(??):24–28, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume82/number18/14265-2411/.

Ayub:2011:TTM

[ARZ11] Qaisar Ayub, Sulma Rashid, and M. Soperi Mohd Zahid. TMHF: Transmit max hop first forwarding strategy to optimize the performance of epidemic routing protocol. International Journal of Computer Applications, 18 (??):40–45, March 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume18/
number5/2277-2949/.

Adnan:2015:EDM

[ARZ15] Saifullah Adnan, Naveed Ur Rehman, and Mohammad Irshad Zahoor. Effect of different modulation techniques comparison of linear MIMO receivers. International Journal of Computer Applications, 121(??):36–40, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume121/number20/21658-5061/.

Amalarethinam:2011:EDO

[AS11] D. I. George Amalarethinam and F. Kurus Malai Selvi. An efficient dual objective grid workflow scheduling algorithm. International Journal of Computer Applications, 33 (??):7–12, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume33/
number1/3983-5625/.

Adam:2012:SEA

[AS12a] T. B. Adam and Md Salam. Spoken English alphabet recognition with mel frequency cepstral coefficients and back propagation neural networks. International Journal of Computer Applications, 42(??):21–27, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number12/5745-7946/.

Ahadpour:2012:CBI

[AS12b] Sodeif Ahadpour and Yaser Sadra. A chaos-based image encryption scheme using chaotic coupled map lattices. International Journal of Computer Applications, 49 (??):15–18, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number2/7599-0311/.

Alam:2012:WSR

[AS12c] Mansaf Alam and Kishwar Sadaf. Web search result clustering using heuristic search and latent semantic indexing. International Journal of Computer Applications, 44 (??):28–33, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number15/6342-8633/.

Ali:2012:NTE

[AS12d] Abdelmgeid Amin Ali and Al-Hussien Seddik Saad. New technique for encoding the secret message to enhance the performance of MSLDIP image steganography method (MPK encoding). International Journal of Computer Applications, 59(??):47–52, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume59/number15/9628-4277/.

Ayub:2012:ADI

[AS12e] Shahanaz Ayub and J. P. Saini. Abnormality detection in Indian ECG using correlation techniques. International Journal of Computer Applications, 58(??):33– 38, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number14/9352-3683/.

Andrews:2013:ETN

[AS13a] J. Andrews and T. Sasikala. Evaluation of tuning and normalized tuning time using an effective and automated framework pertained to benchmark applications. International Journal of Computer Applications, 69(??):48–53, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume69/number9/11874-7679/.

Anitha:2013:HMA

[AS13b] B. Anitha and D. Sivakumar. On (h, m)-anti-fuzzy subrings. International Journal of Computer Applications, 75 (??):45–47, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number8/13135-0499/.

Arya:2013:GRM

[AS13c] Meenakshi S. Arya and Rajesh Siddavatam. Geometric robust multimodal biometric watermarking scheme for copyright protection of digital images. International Journal of Computer Applications, 72(??):40–52, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume72/number9/12525-9056/.

Ayyappan:2013:AMG

[AS13d] G. Ayyappan and K. Sathiya. Analysis of M^{[}x] /G/1 queue with service interruption and extended server vacations with Bernoulli schedule. International Journal of Computer Applications, 68(??):1–7, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number21/11700-7325/.

Abbas:2014:DOA

[AS14a] Mohamed Nadhim Abbas and Hadi K. Shamkhi. Diabolo optical antenna for enhancing and confining electric field resonance. International Journal of Computer Applications, 97(??):29–33, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number23/17324-7696/.

Acharya:2014:EPS

[AS14b] Anal Acharya and Devadatta Sinha. Early prediction of students performance using machine learning techniques. International Journal of Computer Applications, 107(??):37–43, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number1/18717-9939/.

AlZoubi:2014:TSP

[AS14c] Reem Al Zoubi and Asiya Abdus Salam. Travelling salesman problem using dynamic approach. International Journal of Computer Applications, 94(??):20–23, May 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume94/number16/16443-6070/.

Ali:2014:ESV

[AS14d] Fairouz Mushtaq Sher Ali and Falah Hassan Sarhan. Enhancing security of Vigenere cipher by stream cipher. International Journal of Computer Applications, 100(??):1–4, August 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume100/number1/17486-7998/.

Agrawal:2015:OMS

[AS15a] Akansha Agrawal and Shreya Sharma. Optimizing k-means for scalability. International Journal of Computer Applications, 120(??):20–24, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number17/21320-4337/.

Agrawal:2015:MSS

[AS15b] Pallavi Agrawal and Madhu Shandilya. MATLAB simulation of subspace based high resolution direction of arrival estimation algorithm. International Journal of Computer Applications, 130(??):22–27, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number15/23286-2015907175/.

Ahmad:2015:RSS

[AS15c] Shahnawaz Ahmad and Mohd. Sadiq. Recommender systems for software requirements negotiation and prioritization. International Journal of Computer Applications, 117(??):1–7, May 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume117/number13/
20611-3261/.

Azer:2015:PMC

[AS15d] Marianne A. Azer and Noha Gamal El-Din Saad. Prevention of multiple coordinated jellyfish attacks in mobile ad hoc networks. International Journal of Computer Applications, 120(??):12–20, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number20/21342-4354/.

Almazaydeh:2016:ISU

[AS16] Wa’El Ibrahim A. Almazaydeh and H. S. Sheshadri. Image steganography using LSB, LSB+Huffman code, and LSB+Arithmetic code. International Journal of Computer Applications, 155(??):1–7, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume155/number11/26646-2016911478/.

Abbas:2017:TPC

[AS17] Nizar Hadi Abbas and Ahmed Ramz Sami. Tuning of PID controllers for quadcopter system using hybrid memory based gravitational search algorithm — particle swarm optimization. International Journal of Computer Applications, 172(??):9–18, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume172/number4/28238-2017915125/.

Al-Shalabi:2018:OSA

[AS18a] Luai A. Al-Shalabi. Online shopping adoption factors in Kuwait market based on data mining rough set approach. International Journal of Computer Applications, 180(??):10– 17, April 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number32/
29249-2018916832/.

Alazzam:2018:PDS

[AS18b] Hadeel Alazzam and Ahmad Sharieh. Parallel DNA sequence approximate matching with multi-length sequence aware approach. International Journal of Computer Applications, 180(??):1–6, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number26/29117-2018916594/.

Ali:2018:TVD

[AS18c] Y. E. Mohammed Ali and A. G. Saber. Temperatures variation in different human tissues according to blood flow coefficient. International Journal of Computer Applications, 180(??):10–14, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number28/29151-2018916668/.

Argiddi:2018:NAJ

[AS18d] R. V. Argiddi and Disha Rajan Shah. New approach for joint multilabel classification with community-aware label graph learning technique. International Journal of Computer Applications, 180(??):1–7, April 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number36/29295-2018916892/.

Almurayziq:2019:PFQ

[AS19] Awatif Almurayziq and Shaimaa Salama. Proposed framework for quality assessment of e-government portals in Saudi Arabia. International Journal of Computer Applications, 177(??):13–24, November 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number20/31014-2019919633/.

Amer:2024:NFS

[AS24] Mohamed Amer and Tarek S. Sobh. New framework for securing Web APIs token-based authentication/authorization with auto expire auto refresh (AEAR) features. International Journal of Computer Applications, 186(??):8–14, June 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number25/
new-framework-for-securing-web-apis-token-based-authentication-
authorization-with-auto-expire-auto-refresh-aear-
features/.

Asabere:2013:BCM

[Asa13] Nana Yaw Asabere. Benefits and challenges of mobile learning implementation: Story of developing nations. International Journal of Computer Applications, 73(??): 23–27, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number1/12706-9504/.

Anwar:2014:DAS

[ASA14] Uzma Anwar, Ghulam Shabbir, and Malik Ahsan Ali. Data analysis and summarization to detect illegal VOIP traffic with call detail records. International Journal of Computer Applications, 89(??):1–7, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume89/number8/15519-2724/.

Al-Shqeerat:2017:CCS

[ASASHF17] Khalil H. A. Al-Shqeerat, Faiz M. A. Al-Shrouf, Mohammad R. Hassan, and Hassen Fajraoui. Cloud computing security challenges in higher educational institutions

— a survey. International Journal of Computer Applications, 161(??):22–29, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number6/27154-2017913217/.

Arora:2012:CBV

[ASB12] Kanika Arora, Abhishek Singhal, and Abhay Bansal. Correlation between various quality characteristics for aspectoriented systems. International Journal of Computer Applications, 48(??):11–21, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume48/number9/7375-0150/.

Akyol:2015:AIH

[ASB15] Kemal Akyol, Baha Sen, and Safak Bayir. Automatic identification of hard exudates in retinal fundus images using techniques of information retrieval. International Journal of Computer Applications, 120(??):11–16, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume120/number13/21286-4236/.

Alhawarat:2014:ICS

[ASC14] M. Alhawarat, T. Olde Scheper, and N. T. Crook. Investigation of a chaotic spiking neuron model. International Journal of Computer Applications, 99(??):1–8, August 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume99/number17/17462-8258/.

Ahmed:2015:MCS

[ASC15] Arif Ahmed, Abadhan S Sabyasachi, and Ananya Choudhury. Mobile cloud service selection using back propagation neural network. International Journal of Computer Applications, 129(??):1–5, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume129/number14/23138-2015907092/.

Amin:2011:EUB

[ASCR11] Md. Ruhul Amin, Asif Mohammed Samir, Madhusodan Chakraborty, and Md. Mahfuzur Rahaman. An efficient Unicode based sorting algorithm for Bengali words. International Journal of Computer Applications, 24(??): 11–16, June 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume24/
number7/2967-3966/.

Ansari:2016:ATP

[ASD16] M. Izharul Hasan Ansari, S. P. Singh, and M. N. Doja. Adaptive-transmission-power ad hoc on-demand distance vector routing protocol for mobile ad hoc network. International Journal of Computer Applications, 138(??):1–10, March 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume138/number2/24348-2016908721/.

A:2019:ATP

[ASD19] Akolo J A, Felix A S, and Okoh D. Air temperature prediction using artificial neural network for Anyigba, North-Central Nigeria. International Journal of Computer Applications, 177(??):29–34, October 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number10/30935-2019919506/.

Ayyappan:2010:MMR

[ASG10] G. Ayyappan, Gopal Sekar, and A. Muthu Ganapathi. M/M/1 retrial queueing system with vacation interruptions under Erlang-K service. International Journal of Computer Applications, 2(??):52–57, May 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume2/number2/626-882/.

Anandhakumar:2011:ABC

[ASG11] R. Anandhakumar, S. Subramanian, and S. Ganesan. Artificial bee colony algorithm to generator maintenance scheduling in competitive market. International Journal of Computer Applications, 31(??):44–53, October 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume31/number9/3856-5374/.

Arif:2013:DMA

[ASH13] Fahmi Arif, Nanna Suryana, and Burairah Hussin. A data mining approach for developing quality prediction model in multi-stage manufacturing. International Journal of Computer Applications, 69(??):35–40, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume69/number22/12106-8375/.

Ali:2015:UDM

[ASH15] Eman M. Ali, Ahmed F. Seddik, and Mohamed H. Haggag. Using data mining techniques for children brain tumors classification based on magnetic resonance imaging. International Journal of Computer Applications, 131(??):36–42, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume131/number2/23424-2015907268/.

A:2016:CAS

[ASH16a] Abdelmged A A, Al-Hussien Seddik Saad, and Nada Hussien. A combined approach of steganography and cryptography technique based on parity checker and Huffman encoding. International Journal of Computer Applications, 148(??):26–32, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume148/number2/25731-2016911031/.

Ali:2016:ADC

[ASH16b] Eman M. Ali, Ahmed F. Seddik, and Mohamed H. Haggag. Automatic detection and classification of Alzheimer’s disease from MRI using TANNN. International Journal of Computer Applications, 148(??):30–34, August 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume148/number9/25787-2016911320/.

Ashraf:2019:ARI

[Ash19] Syeda Nazia Ashraf. Awareness regarding importance of SDLC among IT professionals and students: a survey. International Journal of Computer Applications, 178(??):38–46, June 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number18/
30638-2019919026/.

Asija:2015:ESP

[Asi15] Ruchika Asija. Enhancing security and privacy of healthcare data using XML schema. International Journal of Computer Applications, 116(??):1–6, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number12/20385-2684/.

Atina:2017:IRS

[ASI17] Vihi Atina, Eko Sediyono, and R. Rizal Isnanto. Information retrieval system for Indonesian manuscript using Semantic Web. International Journal of Computer Applications, 170(??):29–34, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume170/number8/28092-2017914930/.

Al-Shargabi:2011:ICT

[ASII11] Mohammed A. Al-Shargabi, Abdul Samad Ismail, and Sevia M. Idrus. Internet compressed traffic: a solution for the explosion of the Internet. International Journal of Computer Applications, 21(??):12–15, May 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume21/number6/2517-3419/.

Al-sagaf:2012:SMD

[AsJ12] Alawi Abdullah Ahmed Al-sagaf and Dayang Norhayati Abang Jawawi. Standard monitor design for SLA parameters in SOA. International Journal of Computer Applications, 45(??):43–48, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number9/6812-9165/.

Askari:2014:EPA

[Ask14] Mahmoud Askari. Evaluate the prediction accuracy and confidence intervals of Intel Nehalem base on regression model. International Journal of Computer Applications, 92(??):6–9, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume92/
number3/15987-4927/.

Anitha:2015:NMM

[ASK15] S. Anitha, T. Subhashini, and M. Kamaraju. A novel multimodal medical image fusion approach based on phase congruency and directive contrast in NSCT domain. International Journal of Computer Applications, 129(??):30–35, November 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume129/number10/23111-
2015907014/.

Arshad:2017:RBN

[ASK17] Muhammad Arshad, Nasir Saleem, and Shahzada Alamgir Khan. A Rayleigh backscattering noise resilient and cost effective single fiber WDM-PON scheme using DQPSK and intensity re-modulation technique. International Journal of Computer Applications, 170(??):24–30, July 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume170/number7/28085-2017914919/.

Almohammad:2013:HIS

[ASM13] Manhal S. Almohammad, Gouda I. Salama, and Tarek A. Mahmoud. Human identification system based on face using active horizontal levels (AHLs) feature. International Journal of Computer Applications, 61(??):27– 32, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number20/10197-4972/.

Aghababa:2011:EIC

[ASN11] Mohammd Pourmahmood Aghababa, Amin Mohammadpour Shotorbani, and Easa Narimani. Estimation of image corruption inverse function and image restoration using a PSO-based algorithm. International Journal of Computer Applications, 13(??):30–35, January 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume13/number1/1744-2374/.

Alipour:2012:ACM

[ASN12] Hadiseh Seyyed Alipour, Mehdi Sabbari, and Eslam Nazemi. An access control model for Web services with dynamic separation of duty rules. International Journal of Computer Applications, 42(??):6–13, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number6/5695-7469/.

Agrawal:2010:ESC

[ASNS10] Neeraj Agrawal, Sanjeev Sharma, Arun Nahar, and Ramjeevan Singh. Empirical study of CSMA, MACA and EMACA protocols to support QoS under varying conditions of pause time in ad-hoc wireless network by resources of DSR routing protocol. International Journal of Computer Applications, 8(??):29–35, October 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume8/number3/1191-1666/.

Abreu:2017:NGH

[ASP17] Ivanildo Abreu, Rildenir Silva, and Luan Pereira. A neural genetic hybrid model for eigenstructure allocation in the LQR project in DFIG. International Journal of Computer Applications, 162(??):9–15, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume162/number12/27294-2017913419/.

Al-Sobky:2015:GPR

[ASR15] Al-Sayed Ahmed Al-Sobky and Ibrahim M. I. Ramadan. A generalized piecewise regression for transportation models. International Journal of Computer Applications, 129 (??):16–22, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume129/number17/23165-2015907139/.

Al-saedi:2012:CMR

[AsRA^{+}12] Karim Hashim Al-saedi, Sureswaran Ramadass, Ammar Almomani, Selvakumar Manickam, and Wafaa A. H. Ali Alsalihy. Collection mechanism and reduction of IDS alert. International Journal of Computer Applications, 58(??):40– 48, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number4/9274-3530/.

Aghababa:2010:APS

[ASS10] M. Pourmahmood Aghababa, A. M. Shotorbani, and R. M. Shotorbani. An adaptive particle swarm optimization applied to optimum controller design for AVR power systems. International Journal of Computer Applications, 11 (??):22–29, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number10/1618-2176/.

Al-suhail:2013:NNB

[AsS13] Ghaida A. Al-suhail and Sarah A. Subber. Neural networkbased video quality via adaptive FEC in wireless environment. International Journal of Computer Applications, 65 (??):11–15, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number6/10927-5870/.

Amin:2015:RFT

[ASS15a] Zeeshan Amin, Harshpreet Singh, and Nisha Sethi. Review on fault tolerance techniques in cloud computing. International Journal of Computer Applications, 116(??):11–17, April 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume116/number18/
20435-2768/.

Anand:2015:SWM

[ASS15b] Vijay Karan Singh Anand, Jagjit Singh, and Harmeet Singh Saimbhi. Scrutinizing Web mining technology and multi agent systems into intelligent information retrieval system (IIRS). International Journal of Computer Applications, 126(??):17–20, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume126/number13/22611-2015906193/.

Abdulgader:2017:UIT

[ASS17] Sami Ghait Abdulgader, Suryono, and Jatmiko Endro Suseno. The use of information technology toward online transaction in micro enterprise using technology acceptance model (TAM) framework. International Journal of Computer Applications, 177(??):5–10, November 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number4/28612-2017915520/.

Abbasimehr:2011:NFC

[AST11] Hossein Abbasimehr, Mostafa Setak, and M. J. Tarokh. A neuro-fuzzy classifier for customer churn prediction. International Journal of Computer Applications, 19(??): 35–41, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume19/
number8/2379-3138/.

Adedokun:2017:TDM

[ASY17] E. A. Adedokun, Shaibu Idris Salawu, and Basira Yahaya. Towards the development of a modified token based congestion control with adaptive forwarding for opportunistic networks. International Journal of Computer Applications, 174(??):5–10, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number7/28417-2017915424/.

Aung:2014:ENF

[AT14] Ohnmar Aung and Thandar Thein. Enhancing NameNode fault tolerance in Hadoop distributed file system. International Journal of Computer Applications, 87(??):41– 47, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number12/15264-4020/.

Amal:2015:HCM

[AT15] Mohammed Amal and Ben Bella S. Tawfik. High connectivity with minimum BER in wireless sensor networks. International Journal of Computer Applications, 121(??):10–17, July 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume121/number17/
21631-4947/.

Alslaity:2017:ECB

[AT17] Alaa Alslaity and Thomas Tran. An enhanced cryptosystem based on shorter keys and new security component. International Journal of Computer Applications, 166(??):44–50, May 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume166/number8/27693-2017914150/.

Agyekum:2014:QSE

[ATA14] Kwame A. P. Agyekum, Eric T. Tchao, and Emmanuel Affum. Quality of service evaluation of a deployed 3G data network for a campus wide e-learning platform. International Journal of Computer Applications, 91(??): 14–19, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume91/
number16/15964-5263/.

Ahulu:2019:MCT

[ATA19] Boaz Ahulu, Frimpong Twum, and J. B. Hayfron Acquah. Mapping cases of tuberculosis distribution and treatment outcomes using geographical information system — the case of Birim Central Municipality, E/R— Ghana. International Journal of Computer Applications, 182(??):45–54, January 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number35/30294-2019918349/.

Al-Tamimi:2022:SPA

[ATA22] Abdul-Gabbar T. Al-Tamimi and Bushra Abduh Mohammed Aljafary. Security and performance analysis of chaos-based image encryption schemes. International Journal of Computer Applications, 184(??):1–20, October 2022. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number32/32517-2022922384/.

Al-Taweel:2019:LAO

[ATAAAF19] Sadik Ali Al-Taweel, Abdurahman H. Al-Abdali, and Belal Al-Fuhaidi. E-learning adoption in open learning programs: a case study of public and private universities in jeddah city. International Journal of Computer Applications, 178(??):14–22, May 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number10/30566-2019918823/.

Alhmiedat:2012:STD

[ATB12] Tareq Alhmiedat, Anas Abu Taleb, and Mohammad Bsoul. A study on threats detection and tracking systems for military applications using WSNs. International Journal of Computer Applications, 40(??):12–18, February 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume40/number15/5055-7347/.

Abdelhedi:2012:DAV

[ATH12] Slim Abdelhedi, Khaled Taouil, and Bassem Hadjkacem. Design of automatic vision-based inspection system for monitoring in an olive oil bottling line. International Journal of Computer Applications, 51(??):39–46, August 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume51/number21/8330-1968/.

Abood:2012:MSC

[ATM12] Afaneen A. Abood, Firas M. Tuaimah, and Aseel H. Maktoof. Modeling of SVC controller based on adaptive PID controller using neural networks. International Journal of Computer Applications, 59(??):9–16, December 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume59/number6/9551-4007/.

Apadile:2018:ASI

[ATM18] Botlhokang Apadile, Edwin Thuma, and Gontlafetse Mosweunyane. Assessing the significance of incorporating user profiles in social book search. International Journal of Computer Applications, 179(??):1–8, February 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number23/29005-2018916459/.

Aji:2023:HRM

[AU23] Beni Mustiko Aji and Wahyu Sri Utami. Heart rate monitoring system using max30102 sensor and Gaussian naive Bayes algorithm. International Journal of Computer Applications, 185(??):7–12, December 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number47/33004-2023923249/.

Anita:2010:BHA

[AV10] E. A. Mary Anita and V. Vasudevan. Black hole attack prevention in multicast routing protocols for mobile ad hoc networks using certificate chaining. International Journal of Computer Applications, 1(??):21–28, February 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume1/number12/267-426/.

Ansari:2013:PAC

[AV13] Mohammad Hasan Ansari and Vahid Tabataba Vakili. Performance analysis and classification of clone attack detection procedures in mobile wireless sensor networks. International Journal of Computer Applications, 71(??): 5–12, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number21/12608-9417/.

Ashok:2015:LLO

[AV15] Malvika Ashok and Rohit Vaid. LOKS: Low-overhead forward and backward key secrecy in WSNs. International Journal of Computer Applications, 127(??):36–41, October 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume127/number1/22697-2015906317/

.

Arulmary:2018:GFE

[AV18] M. Arulmary and S. P. Victor. Geometrical feature extraction for glaucoma detection. International Journal of Computer Applications, 180(??):1–5, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number27/29141-2018916640/.

Anitha:2010:FBE

[AVH10] J. Anitha, C. Kezi Selva Vijila, and D. Jude Hemanth. Fuzzy based experimental verification of significance of skull tissue removal in brain tumor image segmentation. International Journal of Computer Applications, 1(??):56–61, February 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume1/number29/572-245/.

Acha:2012:ASM

[AVRN12] Rathnakar Acha, V. Vityanathan, Pethur Raj, and S. Nagarajan. Antenna selection in MIMO cognitive radio.

International Journal of Computer Applications, 41(??): 36–39, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume41/
number19/5653-8136/.

Aravind:2013:SBM

[AVS13] P. Aravind, M. Valluvan, and M. Saranya. Simulation based modeling and implementation of adaptive control technique for non linear process tank. International Journal of Computer Applications, 68(??):1–6, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume68/number16/11660-7242/.

Abdo:2013:FSL

[AVT13] Maher Abdo, Ahmad Reza Vali, and Ali Reza Toloei. Fuzzy stabilization loop of one axis gimbal system. International Journal of Computer Applications, 77(??):6– 13, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number3/13372-0975/.

Alhasan:2020:PDA

[AW20] Abdul-Wadud Alhasan and Sonjie Wei. Predicting DDoS anomaly patterns in SDN controller using hidden Markov model. International Journal of Computer Applications, 175(??):33–41, December 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number39/31710-2020920961/.

Adnan:2012:AMF

[AWA12] Nazrul H. Adnan, Khairunizam WAN, and Shahriman AB. Accurate measurement of the force sensor for intermediate and proximal phalanges of index finger. International Journal of Computer Applications, 45(??):59–65, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume45/number15/6860-9492/.

Ali:2019:DEU

[AWK19] Syed Asim Ali, Urooj Waheed, and Hira Anwar Khan. Developing elements of user experience for cellular phones and services: Evidence from Pakistan. International Journal of Computer Applications, 178(??):1–5, June 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number26/30695-2019918994/.

Ahmad:2019:DDT

[AXKT19] Waqas Ahmad, Hongwei Xie, Ammad Khan, and Mubashir Tariq. Data duplication tactics with Hadoop. International Journal of Computer Applications, 177(??):6–12, November 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume177/number15/30973-
2019919548/.

Al-Yousif:2018:ESL

[AY18] Ahmed M. H. Al-Yousif. Estimating sizing and location of shunt compensation using load shedding. International Journal of Computer Applications, 182(??):7–10, December 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number32/30232-
2018918230/.

Abu-Youssef:2018:NTU

[AYAB18] S. E. Abu-Youssef, Nahd S. A. Ali, and M. E. Bakr. A new test for unknown age class of life distribution based on Laplace transform technique. International Journal of Computer Applications, 180(??):21–27, April 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number33/29259-2018916833/.

Anwaar:2017:PPIb

[AYAI17a] Haris Anwaar, Yin YiXin, Muhammad Ammar Ashraf, and Salman Ijaz. Point to point ILC with initial state learning using neural networks. International Journal of Computer Applications, 178(??):35–38, November 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number3/28657-2017915794/.

Anwaar:2017:PPIa

[AYAI17b] Haris Anwaar, Yin YiXin, Muhammad Ammar Ashraf, and Salman Ijaz. Point to point ILC with receding horizon optimization approach. International Journal of Computer Applications, 165(??):19–24, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number11/27618-2017914074/.

Adly:2013:RMJ

[AYE13] Mahmoud Sedky Adly, Aliaa A. A. Youssif, and Ahmed Sharaf Eldin. Recording and measuring of jaw movements using a computer vision system. International Journal of Computer Applications, 81(??):38–43, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume81/number18/14226-2466/.

Anass:2016:IDT

[AYM16] Rghioui Anass, Hernafi Yassine, and Bouhorma Mohammed. IoT for ITS: a dynamic traffic lights control based on the Kerner three phase traffic theory. International Journal of Computer Applications, 145(??):40–48, July 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume145/number1/25245-2016910557/.

Almrayat:2019:EIG

[AYS19] Sondus Almrayat, Rana Yousef, and Ahmad Sharieh. Evaluating the impact of GUI similarity between Android applications to measure their functional similarity. International Journal of Computer Applications, 178(??):31–38, June 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number21/
30661-2019919075/.

Alhudhaif:2015:CPS

[AYYB15] Adi Alhudhaif, Maryam Yammahi, Tong Yan, and Simon Berkovich. A cyber-physical stream algorithm for intelligent software defined storage. International Journal of Computer Applications, 109(??):21–25, January 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume109/number5/19185-0672/.

Al-Zadjali:2016:PTV

[AZ16] Buthaina Mohammed Al-Zadjali. Penetration testing of vulnerability in Android Linux kernel layer via an open network (Wi-Fi). International Journal of Computer Applications, 134(??):40–43, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number6/23922-2016907998/.

Attiya:2017:SPS

[AZ17] Ibrahim Attiya and Xiaotong Zhang. A simplified particle swarm optimization for job scheduling in cloud computing. International Journal of Computer Applications, 163(??):34–41, April 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume163/number9/27426-2017913744/.

Adom:2023:CCU

[AZA23] Worlanyo Adom, Ping Zhang, and Frederick A. Adrah. Combating cybercrime using a prototype PC surveillance and monitoring software system. International Journal of Computer Applications, 185(??):34–39, May 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number9/32732-2023922756/.

Abukhait:2013:SSR

[AZM13] Jafar Abukhait, Imad Zyout, and Ayman M. Mansour. Speed sign recognition using shape-based features. International Journal of Computer Applications, 84(??):31– 37, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number15/14655-2964/.

Azadnia:2017:AII

[AZMP17] Mohammad Azadnia, Shamsossadat Zahedi, Abdorreza Majjedin, and Mohammad Reza Pourabedy. Analysis of the impact of ICT on sustainable development using sustainability indicators. International Journal of Computer Applications, 169(??):13–24, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume169/number6/27989-2017914766/.

Al-zohairy:2016:MFC

[AZS16] Tamer A. Al-Zohairy and Khaled S. Salem. Model following control of SISO nonlinear systems using PID neural networks. International Journal of Computer Applications, 140(??):12–17, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume140/number10/24629-2016909465/.

Al-zoubi:2012:WMP

[AzSKM12] Hussein R. Al-zoubi, Sahar Z. Shatnawi, Alaa I. Kalaf, and Balqees A. Mohammad. A wireless mobile-phone approach to traffic signal preemption for faster service of emergency vehicles. International Journal of Computer Applications, 46(??):35–41, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume46/number3/6891-9225/.

Adelizadeh:2018:IFP

[AZY18a] Mostafa Adelizadeh, Zahra Salehi Zanjani, and Yossef Yasi. Identification the first priority over effective factors in occurrence of building construction accidents with employing the analytical hierarchy process method, and expert choice software. International Journal of Computer Applications, 180(??):13–18, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number25/29111-2018916581/.

Azadnia:2018:CME

[AZY18b] Mohammad Azadnia, Shamsossadat Zahedi, and Alireza Yari. A conceptual model for evaluating impact of the local search engine projects (LOSEPs) on sustainable development based on sustainability indicators. International Journal of Computer Applications, 179(??):1–11, May 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number42/29360-2018916987/.

Bani-Ahmad:2011:CER

[BA11] Sulieman Bani-Ahmad. Comparative evaluation of requestpartitioning-based processor allocation strategies in 2D meshbased multicomputers. International Journal of Computer Applications, 26(??):40–48, July 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume26/number7/3113-4275/.

Baruah:2012:SUA

[BA12a] Nomi Baruah and Ashima. A survey of the use of agile methodologies in different Indian small and medium scale enterprises (SMEs). International Journal of Computer Applications, 47(??):38–44, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume47/number20/7307-0521/.

Bernal:2012:VSU

[BA12b] Alejandra Cruz Bernal and Gilberto Moreno Aguilar. Vision system via USB for object recognition and manipulation with Scorbot-ER 4U. International Journal of Computer Applications, 56(??):10–15, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume56/number18/8998-3174/.

Butt:2012:OSD

[BA12c] Saad Masood Butt and Wan Fatimah Wan Ahmad. Overview of systems design and development with regards to the involvement of user, HCI and software engineers. International Journal of Computer Applications, 58(??):1– 4, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number7/9291-3501/.

Ba-alwi:2013:KAT

[Ba13] Fadl Mutaher Ba-alwi. Knowledge acquisition tool for learning membership function and fuzzy classification rules from numerical data. International Journal of Computer Applications, 64(??):24–30, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume64/number13/10694-5603/.

B:2015:SSC

[BA15a] Rekha B and D V Ashoka. SCIDR: a scalable cluster based inter-domain routing protocol for heterogeneous MANET. International Journal of Computer Applications, 122(??):33–40, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number4/21690-4795/.

Brahmane:2015:EAD

[BA15b] Anilkumar V. Brahmane and Amruta Amune. An efficient approach for dynamic distributed network intrusion detection using online adaboost-based parameterized methods. International Journal of Computer Applications, 117(??):7–13, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number18/20652-3186/.

Bhuyar:2016:DIS

[BA16] Renuka Bhuyar and Saniya Ansari. Design and implementation of smart office automation system. International Journal of Computer Applications, 151(??): 37–42, October 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume151/number3/26217-2016911716/.

Banat:2023:ISA

[BA23] Maoren G. F. K. Banat and Joko Aryanto. Implementation of simple additive weighting method to analyze the selection of Rover/Ranger Mate Scout. International Journal of Computer Applications, 185(??):8–12, November 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number46/32998-2023923271/.

Botchway:2024:HAF

[BAA24] Ivy Botchway, Felix Larbi Aryeh, and Boniface Kayode Alese. Hybrid ANN and Fireworks algorithm for realtime software reliability prediction. International Journal of Computer Applications, 186(??):27–34, December 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number59/
hybrid-ann-and-fireworks-algorithm-for-real-time-
software-reliability-prediction/.

Badr:2013:OGL

[Bad13a] E. M. Badr. Odd graceful labeling of the revised friendship graphs. International Journal of Computer Applications, 65(??):1–6, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number11/10965-6017/.

Badr:2013:ALA

[Bad13b] Sherif M. Badr. Adaptive layered approach using C5.0 decision tree for intrusion detection systems (ALIDS). International Journal of Computer Applications, 66(??):18– 22, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number22/11247-5956/.

Barua:2016:FSN

[BADR16] Sraboni Barua, Tamim Al Mahmud, Samrat Kumar Dey, and Md. Motiur Rahman. Future of social network with collaboration of cloud computing and brain computer interface. International Journal of Computer Applications, 145(??):34–37, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume145/number14/25349-2016910863/.

Bahgat:2017:PEC

[BADS17] Waleed M. Bahgat, Mansoor M. Al-Dabaa, and Hassan H. Soliman. Performance evaluation of cooperative relaying for MAC protocols. International Journal of Computer Applications, 158(??):35–49, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume158/number1/26875-2017912727/.

Bakkali:2015:MOS

[BAF15] Hajira Bakkali, Abdellah Azmani, and Abdelhadi Fennan. A multi-ontology system for modeling the information flow in a warehouse. International Journal of Computer Applications, 120(??):26–35, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number23/21402-4417/.

Boucetta:2014:ASS

[BAIs^{+}14] E. Boucetta, A. Amghar, H. Idrissi-saba, A. M. Gue, D. Esteve, and M. Djafari-rouhani. Atomic scale simulation of dislocation loops formation in thin foil under high energy electron irradiation. International Journal of Computer Applications, 96(??):13–19, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume96/number10/16828-6588/.

Bader:2014:SEM

[BAK14] M. A. L. Bader, Ahmed M Atallah, and Mahmoud M. Kashef. A selection of electrical machine as an electromechanical battery for earth orbit satellite. International Journal of Computer Applications, 90(??):1–6, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume90/number2/15543-4094/.

Brasil:2020:EMB

[BAK20] Antonio Ricardo Alexandre Brasil, Jefferson Oliveira Andrade, and Karin Satie Komati. Eye movements biometrics: a bibliometric analysis from 2004 to 2019. International Journal of Computer Applications, 176(??):1–9, May 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume176/number24/31344-
2020920243/.

Banda:2013:IID

[BAM13] Juan M. Banda, Rafal A. Angryk, and Petrus C. Martens. ImageFARMER: Introducing a data mining framework for the creation of large-scale content-based image retrieval systems. International Journal of Computer Applications, 79 (??):8–13, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number13/13799-1777/.

Banday:2011:EPD

[Ban11] M. Tariq Banday. Easing PAIN with digital signatures. International Journal of Computer Applications, 29(??):46– 56, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume29/
number2/3533-4822/.

Bansah:2016:IID

[BAO16] Isaac Bansah, Tonny Montana Adegboyega, and Stephen Brako Oti. Implementation of intrusion detection system and traffic analysis — a case of a Linux platform. International Journal of Computer Applications, 141(??):21–29, May 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume141/number7/24798-2016909709/

.

Baridam:2014:ACB

[Bar14] Barilee Baridam. Application of the codon-based scoring method in motif detection. International Journal of Computer Applications, 92(??):1–4, April 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume92/number9/16034-4811/.

Barnouti:2016:FRU

[Bar16] Nawaf Hazim Barnouti. Face recognition using PCA-BPNN with DCT implemented on Face94 and Grimace databases. International Journal of Computer Applications, 142(??):8–13, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number6/24898-2016909813/.

Baghel:2012:SMA

[BAS12a] Malti Baghel, Shikha Agrawal, and Sanjay Silakari. Survey of metaheuristic algorithms for combinatorial optimization. International Journal of Computer Applications, 58(??): 21–31, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number19/9391-3813/.

Bassil:2012:GBT

[Bas12b] Youssef Bassil. A generation-based text steganography method using SQL queries. International Journal of Computer Applications, 57(??):27–31, November 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume57/number12/9169-3571/.

Batamuliza:2018:EPP

[Bat18] Jennifer Batamuliza. Efficient and privacy preserving protocol against insider attack for data storage in cloud computing. International Journal of Computer Applications, 179(??):27–36, April 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number34/29220-2018916731/.

Baz:2018:DIE

[Baz18] Mohammed Baz. Digital image encryption using logistic chaotic key-based RC6. International Journal of Computer Applications, 182(??):17–23, July 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number2/29734-2018917453/.

Benammar:2010:HMA

[BB10] M. Benammar and F. Belala. How to make AADL specification more precise. International Journal of Computer Applications, 8(??):16–23, October 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume8/number10/1243-1756/.

BenAbdelghani:2012:NPM

[BB12a] I. Abroug Ben Abdelghani and N. Essoukri Ben Amara. A neuronal planar modeling for handwriting signature based on automatic segmentation. International Journal of Computer Applications, 49(??):29–34, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume49/number8/7648-0739/.

Bhasin:2012:CGR

[BB12b] Anuradha Bhasin and Bhupesh Bhatia. Comparison of growth rate of electromagnetic waves in pre-bunched Cerenkov free electron laser and free electron laser. International Journal of Computer Applications, 52(??):1– 3, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number17/8291-1482/.

Bose:2013:IVF

[BB13a] Dhiman Bose and Uma Basu. Incompressible viscoelastic flow of a generalised Oldroyed-B fluid through porous medium between two infinite parallel plates in a rotating system. International Journal of Computer Applications, 79 (??):13–20, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number1/13704-1452/.

Boulmakoul:2013:SAF

[BB13b] Azedine Boulmakoul and Zineb Besri. Structural analysis foundations for enterprise organizational reengineering. International Journal of Computer Applications, 74 (??):38–44, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number13/12949-0239/.

Badgujar:2014:CRA

[BB14a] Suhas Yuvraj Badgujar and Anand Bone. Cloud resource allocation as preemptive scheduling approach. International Journal of Computer Applications, 88(??):14– 18, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume88/
number18/15452-3989/.

Bharthan:2014:REM

[BB14b] Ankit Bharthan and Devesh Bharathan. RelationalJ-SON, an enriched method to store and query JSON records. International Journal of Computer Applications, 98(??):3–6, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number7/17193-7382/.

Berisha:2012:CNI

[BBA12] Artan Berisha, Behar Baxhaku, and Artan Alidema. A class of non invertible matrices in GF(2) for practical one way hash algorithm. International Journal of Computer Applications, 54(??):19–20, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume54/number18/8667-2574/.

BenAhmed:2016:SCS

[BBB16] Kaoutar Ben Ahmed, Mohammed Bouhorma, and Mohamed Ben Ahmed. Smart citizen sensing: a proposed computational system with visual sentiment analysis and big data architecture. International Journal of Computer Applications, 152(??):20–27, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume152/number6/26324-2016911880/.

BenSalem:2012:SVM

[BBC12a] Samira Ben Salem, Khmais Bacha, and Abdelkader Chaari. Support vector machine-based decision for induction motor fault diagnosis using air-gap torque frequency. International Journal of Computer Applications, 38(??):27– 33, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume38/
number5/4686-6812/.

Benlahrache:2012:MCB

[BBC12b] Nadira Benlahrache, Faiza Belala, and Taha A. Cherfia. Model checking BRS based AADL specification. International Journal of Computer Applications, 52(??):43– 52, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number21/8338-1955/.

Benyarou:2013:MUD

[BBD13] Mourad Benyarou, Fethi Tarik Bendimerad, and Fatima Debbat. Multi-user detection by MMSE adaptive algorithm for multi-beam-MIMO-MC-CDMA using sequences of references. International Journal of Computer Applications, 69(??):40–46, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number19/12081-8267/.

Benkirane:2012:DBS

[BBHL12] Said Benkirane, Abderrahim Benihssane, M. Lahcen Hasnaoui, and Mohamed Laghdir. Distance-based stable election protocol (DB-SEP) for heterogeneous wireless sensor network. International Journal of Computer Applications, 58(??):9–15, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number16/9364-3804/.

Bouchareb:2013:TAF

[BBL13] Ilhem Bouchareb, Amar Bentounsi, and Abdesselam Lebaroud. Torque analysis of faulty SRM using FEA-TFR approach. International Journal of Computer Applications, 61(??): 28–35, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number4/9916-4517/.

Berber:2015:FFO

[BBL15] El-Mehdi Berber, Beldjilali Bouziane, and Myriam Lamolle. A framework for fuzzy ontology storing onto relational database within an a priori data integration system. International Journal of Computer Applications, 122(??):10–18, July 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume122/number15/
21775-5045/.

Babaeizadeh:2014:KDA

[BBM14] Mahnoush Babaeizadeh, Majid Bakhtiari, and Mohd Aizaini Maarof. Keystroke dynamic authentication in mobile cloud computing. International Journal of Computer Applications, 90(??):29–36, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume90/number1/15541-4274/.

Bashir:2013:PEC

[BBMA13] Osama Ahmed Bashir, Md Asri Bin Ngadi, Yahia Abdalla Mohamed, and Mohamed Awad. Performance evaluation of controlling high bandwidth flows by RED-PD. International Journal of Computer Applications, 67(??): 19–23, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number13/11456-7061/.

Ben-Bright:2017:TTM

[BBZG^{+}17] Benuwa Ben-Bright, Yongzhao Zhan, Benjamin Ghansah, Richard Amankwah, Dickson Keddy Wornyo, and Ernest Ansah. Taxonomy and a theoretical model for feedforward neural networks. International Journal of Computer Applications, 163(??):39–49, April 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume163/number4/27386-2017913513/.

Bencherif:2012:AWG

[BC12] M. Bencherif and Ali Chermitti. Assessment of wind generation potentiality and site matching of wind turbines in Algeria. International Journal of Computer Applications, 57 (??):40–47, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number19/9226-3785/.

Blankson:2022:SSS

[BC22] Henry Blankson and Rajan Chattamvelli. A symmetric scheme for securing data in cyber-physical systems/IoT sensor-based systems based on AES and SHA256. International Journal of Computer Applications, 184(??):12–17, August 2022. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume184/number24/
32459-2022922278/.

Banerjee:2014:EUA

[BCBS14] Indradip Banerjee, Dipankar Chatterjee, Souvik Bhattacharyya, and Gautam Sanyal. Establishing user authentication using face geometry. International Journal of Computer Applications, 92(??):1–7, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume92/number16/16090-5259/.

Bhattacharjee:2012:ESE

[BCC12] Dipanjan Bhattacharjee, Gangotri Chakraborty, and Sushabhan Choudhury. Environment sensing and energy efficient smart sensor switching. International Journal of Computer Applications, 50(??):12–18, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number23/7962-1268/.

Babatunde:2013:FMU

[BCC13] Iwasokun Gabriel Babatunde, Akinyokun Oluwole Charles, and Angaye Cleopas. Fingerprint matching using neighbourhood distinctiveness. International Journal of Computer Applications, 66(??):1–8, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume66/number21/11237-5800/.

Babatunde:2013:ULA

[BCO13] Iwasokun Gabriel Babatunde, Akinyokun Oluwole Charles, and Olabode Olatunbosun. Uniformity level approach to fingerprint ridge frequency estimation. International Journal of Computer Applications, 61(??):26–32, January 2013. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume61/number22/10229-4809/

.

Banerjee:2010:EMT

[BD10a] Anuradha Banerjee and Paramartha Dutta. An efficient method for tracking nodes (EMTN) in mobile ad hoc networks. International Journal of Computer Applications, 11 (??):38–41, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number2/1552-2069/.

Bishnoi:2010:CCU

[BD10b] P. K. Bishnoi and P. K. Das. Check in check out using Blue Tooth ad hoc network (CICO-BT). International Journal of Computer Applications, 7(??):21–27, September 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume7/number3/1145-1498/

.

Baghel:2013:MRA

[BD13] Ranjan Baghel and Maitreyee Dutta. Multi-relational algebra and its application to unrealized datasets used in C4.5. International Journal of Computer Applications, 73 (??):25–28, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number7/12755-9710/.

Bhatti:2018:BDD

[BD18] Dhaval Bhatti and Anuradha Deshpande. Bad data detection and data filtering in power system. International Journal of Computer Applications, 182(??):36–39, December 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number30/30222-
2018918263/.

Bittencourt:2017:LFC

[BDC17] Joao Carlos Bittencourt, Wagner Luiz De Oliveira, and Ricardo Chaves. Low-footprint CLEFIA FPGA implementations with full-key expansion. International Journal of Computer Applications, 174(??):1–8, September 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume174/number5/28400-2017915392/.

Bhuiyan:2013:OPH

[BDN13] Mahmud Abdul Matin Bhuiyan, Anik Deb, and Arefin Nasir. Optimum planning of hybrid energy system using HOMER for rural electrification. International Journal of Computer Applications, 66(??):45–52, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume66/number13/11148-6239/.

Battou:2011:ALS

[BECM11] Amal Battou, Ali El Mezouary, Chihab Cherkaoui, and Driss Mammass. An adaptive learning system architecture based on a granular learning object framework. International Journal of Computer Applications, 32(??):18– 27, October 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume32/
number5/3900-5465/.

Badry:2013:TSW

[BEE13] Rasha Mohammed Badry, Ahmed Sharaf Eldin, and Doaa Saad Elzanfally. Text summarization within the latent semantic analysis framework: Comparative study. International Journal of Computer Applications, 81(??):40– 45, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume81/
number11/14060-2366/.

Befekadu:2018:EPN

[Bef18] Abeselom Befekadu. Enhancing the performance of network intrusion detection system by combining naïve Bayes, decision tree and K-nearest neighbors algorithms. International Journal of Computer Applications, 180(??):48–53, June 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number49/
29573-2018917351/.

Bejju:2016:SAC

[Bej16] Anurag Bejju. Sales analysis of e-commerce websites using data mining techniques. International Journal of Computer Applications, 133(??):36–40, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number5/23786-2016907812/.

Berneti:2011:DFS

[Ber11] Shahram Mollaiy Berneti. Design of fuzzy subtractive clustering model using particle swarm optimization for the permeability prediction of the reservoir. International Journal of Computer Applications, 29(??):33– 37, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume29/
number11/3687-5117/.

Beyer:2014:MSL

[Bey14] Cassandra Beyer. Mobile security: a literature review. International Journal of Computer Applications, 97(??): 9–11, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number8/17025-7315/.

Belgacem:2013:DTC

[BFA13] A. Belgacem, M. A. Fnaiech, and M. N. Abdelkrim. Direct torque control strategy for dual three phase induction motors. International Journal of Computer Applications, 65(??):33–37, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume65/number25/11275-6524/.

Bacchini:2014:EAG

[BFD14] Pedro H. F. Bacchini, Jose Olimpio Ferreira, and Talles Marcelo G. De A.Barbosa. Emopad: an affective gamepad.

International Journal of Computer Applications, 102(??): 5–12, September 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume102/number15/17889-8885/.

BenAhmed:2015:IFO

[BFD15] Zouhour Ben Ahmed, Gérard Favier, and Nabil Derbel. Identification of fifth-order Wiener and Hammerstein channels based on the estimation of an associated Volterra kernel. International Journal of Computer Applications, 123(??):1–5, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number7/21968-2015905405/.

Bharathi:2013:SHC

[BG13] V. C. Bharathi and M. Kalaiselvi Geetha. Segregated handwritten character recognition using GLCM features. International Journal of Computer Applications, 84(??):1– 7, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number2/14545-2644/.

Bansal:2016:CTT

[BG16] Manisha Bansal and Naresh Kumar Garg. Crossbreed thresholding text extraction procedure for images using DWT domain and SVM classifier. International Journal of Computer Applications, 134(??):23–28, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume134/number5/23911-2016907957/.

Bhoite:2017:HPC

[BG17] Kranti Bhoite and Sachin Gengaje. Handoff prioritize call admission control policy for integrated macrocell-femtocell network. International Journal of Computer Applications, 164(??):21–26, April 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume164/number2/27457-2017913593/.

BritoBatista:2020:MOA

[BG20] Mariana Brito Batista and Warley Gramacho Da Silva. Metaheuristics for optimization analysis on urban public transport. International Journal of Computer Applications, 176(??):1–6, April 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number17/31290-2020920124/.

Bansal:2011:ASP

[BGS11] Rajan Bansal, Himani Goyal, and Parminder Singh. Analytical study the performance evaluation of mobile ad hoc networks using AODV protocol. International Journal of Computer Applications, 14(??):34–37, January 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume14/number4/1829-2442/.

Barsoum:2014:VOR

[BH14] Ayad F. Barsoum and M. Anwar Hasan. Verifying outsourced replicated data in cloud computing storage systems. International Journal of Computer Applications, 99 (??):1–13, August 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume99/
number7/17382-7921/.

Bader:2016:RWD

[BH16] Waseem I. Bader and Abdelaziz I. Hammouri. Responsive Web design techniques. International Journal of Computer Applications, 150(??):18–27, September 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume150/number2/26066-2016911463/.

Baruah:2017:EVV

[BH17a] Indira Baruah and G. C. Hazarika. Effects of variable viscosity and thermal conductivity of unsteady micropolar fluid under mixed convection in presence of uniform magnetic field on stretching surface. International Journal of Computer Applications, 166(??):17–25, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume166/number3/27650-2017913949/.

Bhattacharya:2017:SSN

[BH17b] Munmun Bhattacharya and Bratati Hui. Securing social network graph against structural attack based on ant colony optimization. International Journal of Computer Applications, 167(??):21–26, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number9/27801-2017914372/.

Bhattacharjee:2010:SKI

[Bha10] Utpal Bhattacharjee. Search key identification in a spoken query using isolated keyword recognition. International Journal of Computer Applications, 5(??):14–21, August 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume5/number8/933-1310/.

Bharti:2017:FCU

[Bha17] Priya Bharti. Fingerprint compression using sparse representation. International Journal of Computer Applications, 179(??):32–36, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number3/28719-2017915908/.

Bhattacharya:2012:IAT

[BHC12] Tanmay Bhattacharya, Sirshendu Hore, and S. R. Bhadra Chaudhuri. An image authentication technique by handwritten signature verification using DWT and ANN. International Journal of Computer Applications, 47(??): 35–40, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number21/7476-0575/.

Buaba:2013:OLF

[BHK13] Ruben Buaba, Abdollah Homaifar, and Eric Kihn. Optimal load factor for approximate nearest neighbor search under exact Euclidean locality sensitive hashing. International Journal of Computer Applications, 69(??): 22–31, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number21/12096-8258/.

Baridam:2012:PPS

[BI12] Barileé Baridam and Chukwuma Irozuru. The prediction of prevalence and spread of HIV/AIDS using artificial neural network — the case of Rivers State in the Niger Delta, Nigeria. International Journal of Computer Applications, 44 (??):42–45, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number2/6239-8584/.

BinHakim:2013:UEG

[BI13] Safayat Bin Hakim and Kazi Khairul Islam. Ultrasonic elastogram generation by 2D thin plate smoothing spline based mathematical interpolation technique. International Journal of Computer Applications, 75(??):1–7, August 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume75/number2/13080-0321/.

Bhole:2019:ERR

[BI19] Leena Bhole and Maya Ingle. Estimating range and relationship of EEG frequency bands for emotion recognition. International Journal of Computer Applications, 178(??):16–21, May 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number13/
30590-2019918896/.

Bridget:2024:GDP

[BIAE24] Kilakime Warikaramu-Ere Bridget, Kingsley Ezebunwo Igbudu, Olukunle Oladipupo Amoo, and Obot Israel Emaekop. Gastroenteritis disease in pediatrics detection system using fuzzy logic type 2. International Journal of Computer Applications, 186(??):22–29, March 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number11/
gastroenteritis-disease-in-pediatrics-detection-system-
using-fuzzy-logic-type-2/.

Bekkari:2012:SSC

[BIE^{+}12] Aissam Bekkari, Soufian Idbraim, Azeddine Elhassouny, Driss Mammass Danielle Ducrot, Mostapha El Yassa, and Danielle Ducrot. Spectral and spatial classification of high resolution urban satellites images using Haralick features and SVM with SAM and EMD distance metrics. International Journal of Computer Applications, 46(??): 28–37, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number11/6956-9585/.

Babu:2012:OWS

[BJ12] T. K. Ramesh Babu and Nanda Dulal Jana. An optimized way for static channel allocation in mobile networks using genetic algorithms. International Journal of Computer Applications, 45(??):48–52, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number19/7029-9829/.

Bindal:2015:PAS

[BJPP15] Amit Kumar Bindal, Anuj Jain, Devendra Prasad, and R. B. Patel. Power ambulance system for energy harvested wireless sensor network. International Journal of Computer Applications, 117(??):9–13, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number16/20637-3313/.

Bonde:2013:WTT

[BJS13] Padma Bonde, Bhavana Jharia, and A. K. Shrivastav. Wi-TCP: a TCP in wireless environment. International Journal of Computer Applications, 73(??):30–34, July 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume73/number8/12763-9731/.

Brar:2014:MEA

[BJS14] Gurinderpal Singh Brar, V. K. Jain, and Amanpreet Singh. Modeling and economic analysis of energy generation from biomass energy. International Journal of Computer Applications, 108(??):1–4, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number20/19024-8579/.

Bhirud:2010:NAI

[BK10a] S. G. Bhirud and Vijay Katkar. Novel architecture for intrusion-tolerant distributed intrusion detection system using packet filter firewall and state transition tables. International Journal of Computer Applications, 8(??):29–32, October 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume8/number11/1248-
1631/.

Bhowmik:2010:NAC

[BK10b] Biswajit Bhowmik and Abhishek Kumar. A new approach of complier design in context of lexical analyzer and parser generation for NextGen languages. International Journal of Computer Applications, 6(??):21–25, September 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume6/number11/1116-

Bushehrian:2011:ESA

[BK11] Omid Bushehrian and Robert Khaldar. An extendable software architecture for personalized e-learning systems. International Journal of Computer Applications, 33(??):1– 6, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume33/
number10/4053-5819/.

Bhuvaneswari:2013:SVM

[BK13] P. Bhuvaneswari and J. Satheesh Kumar. Support vector machine technique for EEG signals. International Journal of Computer Applications, 63(??):1–5, February 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume63/number13/10523-5503/.

Bagal:2014:PMF

[BK14] S. B. Bagal and U. V. Kulkarni. Pruned modified fuzzy hyperline segment neural network and its application to pattern classification. International Journal of Computer Applications, 93(??):43–50, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume93/number12/16271-6008/.

Bhan:2017:RPH

[BK17] Sarita Zutshi Bhan and Pragati Kapoor. Reduction of power in high frequency oscillators using active elements for focused ultrasound application. International Journal of Computer Applications, 173(??):12–17, September 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume173/number5/28330-2017915108/.

Boucherit:2011:RLB

[BKB11] Ammar Boucherit, Abdallah Khebaba, and Faiza Belala. Rewriting logic based approach for the formalization of critical systems based on multi-agent system. International Journal of Computer Applications, 13(??):6–13, January 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume13/number2/1755-
2392/.

Bade:2013:REA

[BKK13] Shilpa Bade, Meeta Kumar, and Pooja Kamat. A reactive energy-alert algorithm for MANET and its impact on node energy consumption. International Journal of Computer Applications, 71(??):1–6, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number18/12583-9106/.

Baajwa:2015:ERG

[BKK15] Davinder Singh Baajwa, S. A. Khan, and Jaideep Kaur. Evaluating the research gaps of underwater image enhancement techniques. International Journal of Computer Applications, 117(??):19–23, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number20/20672-3369/.

Beg:2012:BAB

[BKM12] Parveen Beg, Iqbal A. Khan, and Sudhanshu Maheshwari. Biphase amplifier based precision rectifiers using current conveyors. International Journal of Computer Applications, 42 (??):14–18, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number3/5672-7704/.

Bhardwaj:2015:PAS

[BKM15] R. K. Bhardwaj, Komaldeep Kaur, and S. C. Malik. Performance analysis of a stochastic system with standby failure and maintenance. International Journal of Computer Applications, 125(??):6–11, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number1/22394-2015904139/.

Boaz:2014:WSN

[BKS14] Lawrence Boaz, Shubi Kaijage, and Ramadhani Sinde. Wireless sensor node for gas pipeline leak detection and location. International Journal of Computer Applications, 100(??):29–33, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number18/17627-8394/.

Buragohain:2015:ABD

[BKSB15] Chaitanya Buragohain, Manash Jyoti Kalita, Santosh Singh, and Dhruba K. Bhattacharyya. Anomaly based DDoS attack detection. International Journal of Computer Applications, 123(??):35–40, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number17/22055-2015905786/.

BenIdder:2015:UML

[BL15] Hassan Id Ben Idder and Nabil Laachfoubi. Unsupervised multi-level thresholding method for weather satellite cloud segmentation. International Journal of Computer Applications, 118(??):1–5, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number16/20826-3544/.

Bedoui:2012:NMS

[BLA12] Saida Bedoui, Majda Ltaief, and Kamel Abderrahim. New method for the systematic determination of the model’s base of time varying delay system. International Journal of Computer Applications, 46(??):13–19, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume46/number1/6871-8967/.

Bleed:2018:QD

[Ble18] Salma Omar Bleed. L-quadratic distribution. International Journal of Computer Applications, 179(??):6–11, January 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume179/number13/28858-
2018916151/.

Bakkali:2013:GPT

[BLL13] Asma Bakkali, Youssef Lagmich, and Abdelouahid Lyhyaoui. Green passive temperature sensing technology using sensitive ceramic capacitor. International Journal of Computer Applications, 71(??):37–42, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number15/12436-9166/.

Bantog:2021:CPR

[BLQ21] Jimmy Z. Bantog, Luisito Lolong Lacatan, and Mary Ann F. Quioc. Cross-platform relational data extraction utilizing SQL server (X-PRESS). International Journal of Computer Applications, 183(??):34–41, October 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume183/number31/32132-2021921703/.

B:2013:CSR

[BM13a] Govindarajalu B and K. M. Mehata. Code size reduction in embedded systems with redesigned ISA for RISC processors. International Journal of Computer Applications, 64 (??):38–45, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number12/10690-5594/.

Banerjee:2013:ACE

[BM13b] Joydipto Banerjee and Biswajit Mohapatra. An approach to cost-effective transformation of workloads towards cloud delivery models. International Journal of Computer Applications, 82(??):47–53, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume82/number14/14232-2354/.

Bhattacharyya:2013:MRB

[BM13c] Anindya Bhattacharyya and N. Murali. Model reference based tuning of PID controller using Bode’s ideal transfer function and constrained particle swarm optimization. International Journal of Computer Applications, 71(??): 21–27, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number17/12450-9207/.

Badhera:2014:FBT

[BM14a] Usha Badhera and Annu Maheshwari. Fault based test suite prioritization based on minimal MUMCUT strategy. International Journal of Computer Applications, 89(??): 33–38, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number16/15717-4613/.

Bhatti:2014:CAI

[BM14b] Arif M. Bhatti and Mehedi Masud. Context aware intelligent wallet for healthcare. International Journal of Computer Applications, 96(??):31–36, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume96/number15/16871-6769/.

Balasubramanian:2015:MDT

[BM15a] Pad Balasubramanian and Prasanth Malla. De-mystifying data testing and applying automation. International Journal of Computer Applications, 126(??):1–5, September 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume126/number11/22593-
2015906212/.

Balpande:2015:PGI

[BM15b] Vijaya Balpande and Anjali Mahajan. Parallelization of graph isomorphism using OpenMP. International Journal of Computer Applications, 117(??):33–41, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume117/number8/20576-2982/.

Bobade:2015:SOM

[BM15c] Sunil Devidas Bobade and Vijay R. Mankar. Space optimized multiplier architecture for embedded cryptoprocessor. International Journal of Computer Applications, 113(??):26–32, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume113/number14/19897-1982/.

Bedi:2016:ITU

[BM16a] Rajneeshkaur Bedi and Anjali Mahajan. Identification of Ktuples using K-anonymity algorithm to the watermarking of social network database. International Journal of Computer Applications, 142(??):1–7, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number14/24961-2016909968/.

Bera:2016:EPM

[BM16b] Somnath Bera and Kalpana Mahalingam. Extending Parikh q-matrices. International Journal of Computer Applications, 134(??):1–7, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number4/23899-2016907909/.

Bhowmik:2020:ENC

[BM20a] Awnon Bhowmik and Unnikrishnan Menon. Enhancing the NTRU cryptosystem. International Journal of Computer Applications, 176(??):46–53, June 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number29/31388-2020920320/.

Bhowmik:2020:MME

[BM20b] Awnon Bhowmik and Unnikrishnan Menon. MES — Modern Encryption Standard. International Journal of Computer Applications, 176(??):21–27, July 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number36/31435-2020920479/.

Bharadwaj:2016:IDV

[BMD16] Naina Bharadwaj, Prachi Mishra, and Prajyoti Dsilva. Interactive data visualization for census data. International Journal of Computer Applications, 149(??):20– 26, September 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume149/number4/25985-2016911387/.

Bahnasse:2019:SEV

[BME19] Ayoub Bahnasse, Zakariae Malainine, and Hamza El Azzaoui. Study and evaluation of VoIP scalability performances. International Journal of Computer Applications, 182(??):10–14, April 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number47/30507-2019918702/.

Babaei:2011:LSS

[BMF11] M. Babaei, M. Mohammadi, and S. M. T. FatemiGhomi. Lot sizing and scheduling in flow shop with sequence-dependent setups and backlogging. International Journal of Computer Applications, 30(??):52–59, September 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume30/number8/3658-4957/.

Blair:2022:CSA

[BML22] Benjamin W. Blair, Sasidhar Malladi, and James F. Lowe. A case study of the application of artificial intelligence to aid in the clinical detection of porcine reproductive and respiratory syndrome virus in sow farms. International Journal of Computer Applications, 184(??):13–20, April 2022. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number9/32356-2022922061/.

BenAyed:2012:RCE

[BMM12] Norh‘ene Gargouri Ben Ayed, Alima Damak Masmoudi, and Dorra Sellami Masmoudi. Robust contrast enhancement for digital mammography. International Journal of Computer Applications, 56(??):15–19, October 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume56/number9/8918-2976/.

BenAyed:2013:RTA

[BMM13] Norhene Gargouri Ben Ayed, Alima Damak Masmoudi, and Dorra Sellami Masmoudi. Robust textural approach for false positive reduction in mammographic mass. International Journal of Computer Applications, 68(??): 1–5, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume68/
number6/11580-6902/.

Burney:2012:GAT

[BMRA12] S. M. Aqil Burney, Nadeem Mahmood, Kashif Rizwan, and Usman Amjad. A generic approach for team selection in multi-player games using genetic algorithm. International Journal of Computer Applications, 40(??):11– 17, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume40/
number17/5071-7440/.

Begum:2014:QAR

[BMU14] Mahbuba Begum, Md. Golam Moazzam, and Mohammad Shorif Uddin. Quantitative analysis on robustness of FLD and PCA-based face recognition algorithms. International Journal of Computer Applications, 99(??):10– 14, August 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume99/
number19/17480-8222/.

Barboucha:2014:CPP

[BN14] H. Barboucha and M. Nasri. Calculating page pertinence for Web referencing: a new approach. International Journal of Computer Applications, 98(??):1–5, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume98/number19/17288-7254/.

Buckley:2014:VSS

[BNA14] N. Buckley, A. K. Nagar, and S. Arumugam. Visual secret sharing between remote participants. International Journal of Computer Applications, 103(??): 8–17, October 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume103/number2/18044-8931/.

Bakiri:2021:NCM

[BNMM21] Hussein Bakiri, Hamisi Ndyetabura, Libe Massawe, and Hellen Maziku. A novel cleansing method for random-walk data using extended multivariate nonlinear regression: a data preprocessor for load forecasting mechanism. International Journal of Computer Applications, 183(??):49–57, July 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume183/number16/
32014-2021921503/.

Bibi:2015:ASA

[BNS15] Sana Bibi, Umber Nazir, and Mehreen Sishar. Analytical survey for assuring quality standards in GPS based tracking system. International Journal of Computer Applications, 111(??):30–36, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number8/19561-1316/.

Basumatary:2017:STD

[BNS17] Nayan Basumatary, Bhabesh Nath, and Nityananda Sarma. Signal type detection in CRN: a machine learning framework using spectral correlation feature. International Journal of Computer Applications, 177(??):1–7, November 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume177/number7/28682-2017915792/

.

Bukola:2016:AID

[BO16] Olubadeji Bukola and Adetunmbi A O. Auto-immunity dendritic cell algorithm. International Journal of Computer Applications, 137(??):10–17, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume137/number2/24246-2016908689/.

Bukola:2017:PAP

[BO17] Olubadeji Bukola and Adetunmbi A O. Purging/antirobustness as a panacea for attack subversion and autoimmunity on the network. International Journal of Computer Applications, 173(??):1–8, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume173/number8/28352-2017914214/.

Boutaounte:2021:SPC

[BO21] Mehdi Boutaounte and Youssef Ouadid. Sequential primitive classifier combination for amazigh alphabets OCR. International Journal of Computer Applications, 174(??):6–13, March 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume174/number24/
31820-2021921140/.

Bissoto:2017:ACD

[BOB17] Moises Bruno L. Bissoto, Ary Henrique M. Oliveira, and Glenda M. Botelho. Analysis of communities detection algorithms in complex networks. International Journal of Computer Applications, 173(??):1–7, September 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume173/number7/28344-2017915356/.

Bukola:2015:DCE

[BOK15] Olubadeji Bukola, Adetunmbi A O, and Alese B K. DCA as context (environment) sensitive system. International Journal of Computer Applications, 131(??): 39–46, December 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume131/number1/23417-2015907236/.

Bonny:2015:POD

[Bon15] Talal Bonny. Performance optimization of the database sequencing applications. International Journal of Computer Applications, 112(??):1–8, February 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume112/number5/19659-1315/.

Borikar:2019:XEG

[Bor19] Aditya Borikar. XGFX: eXtensible game framework through XMPP. International Journal of Computer Applications, 178(??):45–51, September 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number51/30903-2019919434/.

Boudkhil:2015:AFP

[BOS15] Abdelhakim Boudkhil, Asmaa Ouzzani, and Belabbes Soudini. Analysis of fundamental photodetection noises and evaluation of PIN and APD photodiodes performances using an optical high debit transmission chain simulated by optisystem. International Journal of Computer Applications, 115(??):21–29, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number18/20252-2620/.

Bouket:2012:NWB

[Bou12a] Ali Chenari Bouket. NEMIDSOFT: a Web-based software as an aid to the identification of nematodes. International Journal of Computer Applications, 57(??):16– 21, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number4/9102-3239/.

Boutasseta:2012:PPB

[Bou12b] N. Boutasseta. PSO-PI based control of photovoltaic arrays. International Journal of Computer Applications, 48 (??):36–40, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number17/7444-0557/.

Boum:2013:OBQ

[Bou13] A. T. Boum. Observer based and quadratic dynamic matrix control of a fluid catalytic cracking unit: a comparison study. International Journal of Computer Applications, 80(??):1–8, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume80/number3/13838-1668/.

Bhargava:2011:CMD

[BP11] Sanjay Bhargava and G. N. Purohit. Construction of a minimal deterministic finite automaton from a regular expression. International Journal of Computer Applications, 15 (??):16–27, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume15/
number4/1938-2589/.

Biswas:2015:FGP

[BP15] Papun Biswas and Bijay Baran Pal. Fuzzy goal programming approach to solve linear multilevel programming problems using genetic algorithm. International Journal of Computer Applications, 115(??):10–19, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number3/20130-2215/.

Basak:2016:IBF

[BP16] Saikat Basak and Ranjan Parekh. An improved bagof-features approach for object recognition from natural images. International Journal of Computer Applications, 151(??):5–11, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume151/number10/26267-2016911916/.

Baboo:2011:PCP

[BR11a] K. Chandrasekar Baboo and N. V. Ramana. Performance comparison of DE, PSO and GA approaches in transmission power loss minimization using FACTS devices. International Journal of Computer Applications, 33(??):58– 62, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume33/
number5/4020-5717/.

Babu:2011:EBA

[BR11b] A. Sudhir Babu and K. V. Sambasiva Rao. Evaluation of BER for AWGN, Rayleigh and Rician fading channels under various modulation schemes. International Journal of Computer Applications, 26(??):23–28, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume26/number9/3132-4317/.

Bouanaka:2013:PFF

[BR13] F. Bouanaka and S. Rebiai. Pressure and free flight time effects on glow discharge characteristics. International Journal of Computer Applications, 72(??):38–44, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume72/number1/12461-8817/.

Bharat:2015:SAD

[BR15] Sharma Bharat and Mandre B R. A secured and authorized data deduplication in hybrid cloud with public auditing. International Journal of Computer Applications, 120(??):19–24, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number16/21311-4286/.

Belmeguenai:2012:SCA

[BRD12] Aissa Belmeguenai, Mohammed Redjimi, and Nadir Derouiche. A stream cipher algorithm based on nonlinear combination generator with enhanced security for image encryption. International Journal of Computer Applications, 56 (??):17–23, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number1/8855-2807/.

Bavkar:2015:GAH

[BRD15] S. S. Bavkar, J. S. Rangole, and V. U. Deshmukh. Geometric approach for human emotion recognition using facial expression. International Journal of Computer Applications, 118(??):17–22, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number14/20814-3174/.

Bethel:2016:KDA

[BRR16] G. N. Beena Bethel, T. V. Rajinikanth, and S. Viswanadha Raju. A knowledge driven approach for efficient analysis of heart disease dataset. International Journal of Computer Applications, 147(??):39–46, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume147/number9/25685-2016911187/.

Bhaskar:2016:PCG

[BRS16] K. Vijay Bhaskar, K. Thammi Reddy, and S. Sumalatha. Pushing constraints to generate top-K closed sequential graph patterns. International Journal of Computer Applications, 137(??):34–42, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume137/number7/24290-2016908818/.

Bibi:2010:IDS

[BS10] K. Ameenal Bibi and R. Selvakumar. The inverse domination in semi-total block graphs. International Journal of Computer Applications, 8(??):5–7, October 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume8/number8/1231-1806/.

Berneti:2011:ICA

[BS11a] Shahram Mollaiy Berneti and Mehdi Shahbazian. An imperialist competitive algorithm artificial neural network method to predict oil flow rate of the wells. International Journal of Computer Applications, 26(??):47–50, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume26/number10/3137-4326/.

Bhandari:2011:TOB

[BS11b] Amit Bhandari and Manpreet Singh. Towards ontologybased services interfaces in services systems based environment. International Journal of Computer Applications, 15 (??):19–24, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume15/
number6/1953-2611/.

B:2012:ADD

[BS12a] Sujithkumar S B and Vipula Singh. Automatic detection of diabetic retinopathy in non-dilated RGB retinal fundus images. International Journal of Computer Applications, 47 (??):26–32, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number19/7297-0511/.

Barua:2012:EDB

[BS12b] Hrishav Bakul Barua and Sauravjyoti Sarmah. An extended density based clustering algorithm for large spatial 3D data using polyhedron approach. International Journal of Computer Applications, 58(??):4–15, November 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume58/number2/9252-3418/.

Bushehrian:2012:EEI

[BS12c] Omid Bushehrian and Mojtaba Shahkolahi. An extendible emulator for investigating configuration algorithms in mobile ad-hoc networks. International Journal of Computer Applications, 59(??):19–23, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume59/number16/9632-4348/.

BenAmmar:2014:NAB

[BS14a] Mohsen Ben Ammar and Rachid Souissi. A new approach based on wireless sensor network and fuzzy logic for forest fire detection. International Journal of Computer Applications, 89(??):48–55, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume89/number2/15477-4169/.

Bhatnagar:2014:EMI

[BS14b] Monika Bhatnagar and Prashant Kumar Singh. An efficient method of image segmentation for harvest time identification. International Journal of Computer Applications, 87 (??):31–34, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number7/15222-3734/.

Berwal:2016:DMA

[BS16] Parul Berwal and Kamna Solanki. Data mining application for health seeker and provider. International Journal of Computer Applications, 149(??):15–23, September 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume149/number8/26017-2016911525/.

Bajpai:2017:AIE

[BS17] Kriti Bajpai and Rishi Soni. Analysis of image enhancement techniques used in remote sensing satellite imagery. International Journal of Computer Applications, 169(??):1–11, July 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume169/number10/28018-
2017914884/.

Babiker:2018:RCS

[BS18] Abdelaziz Babiker and Abd-El-Kader Sahraou. Requirements changing in system engineering and impact on safety. International Journal of Computer Applications, 180(??):5–9, April 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number37/
29304-2018916915/.

Basha:2021:DBM

[BS21a] Noor Basha and Ashok Kumar P S. Distance-based Kmeans clustering algorithm for anomaly detection in categorical datasets. International Journal of Computer Applications, 183(??):9–14, June 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number11/31970-2021921415/.

Bond:2021:MDC

[BS21b] Koby Bond and Alaa Sheta. Medical data classification using machine learning techniques. International Journal of Computer Applications, 183(??):1–8, June 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume183/number6/31928-2021921339/.

Baareh:2012:FBO

[BSAb12] Abdel Karim Baareh, Alaa F. Sheta, and Mohammad S. Al-batah. Feature based 3D object recognition using artificial neural networks. International Journal of Computer Applications, 44(??):1–7, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume44/number5/6256-8402/.

Bais:2015:RBC

[BSG15] Varsha Bais, Amit Sinhal, and Bhupesh Gour. Rate base congestion control in multipath routing strategies under MANET. International Journal of Computer Applications, 112(??):29–34, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number13/19729-1417/.

Banu:2012:IBA

[BSI12] Nazreen Banu, Samia Souissi, and Taisuke Izumi. An improved Byzantine agreement algorithm for synchronous systems with mobile faults. International Journal of Computer Applications, 43(??):1–7, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number22/6400-8878/.

Burney:2012:UTS

[BSM^{+}12] Aqil Burney, Badar Sami, Nadeem Mahmood, Zain Abbas, and Kashif Rizwan. Urdu text summarizer using sentence weight algorithm for word processors. International Journal of Computer Applications, 46(??):38–45, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume46/number19/7053-9731/.

Balavalad:2017:DSM

[BSM17] Kirankumar B. Balavalad, B. G. Sheeparamatti, and Veekshit B. Math. Design and simulation of MEMS capacitive pressure sensor array for wide range pressure measurement. International Journal of Computer Applications, 163(??):39–46, April 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume163/number6/27403-2017913554/.

Bhushan:2011:PAF

[BSP11] Bharat Bhushan, Madhusudan Singh, and Prem Prakash. Performance analysis of field oriented induction motor using fuzzy PI and fuzzy logic based model reference adaptive control. International Journal of Computer Applications, 17(??):5–12, March 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume17/number4/2211-2811/.

Bakshi:2012:NIA

[BSS12] Tuli Bakshi, Bijan Sarkar, and Subir Kumar Sanyal. A novel integrated AHP-QFD model for software project selection under fuzziness. International Journal of Computer Applications, 54(??):36–43, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume54/number7/8581-2327/.

Basim:2015:AES

[BSS15] N Mohammed Abu Basim, P Sathyabalan, and P Suresh. Analysis of EEG signals and facial expressions to detect drowsiness and fatigue using Gabor filters and SVM linear classifier. International Journal of Computer Applications, 115(??):9–14, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number11/20194-2433/.

Bakura:2013:NGM

[BSZ13] Mala Umar Mustapha Bakura, Bala Alhaji Salihu, and Suleiman Zubair. A novel geometrical mixed-bounce tworing deterministic model for MIMO mobile-to-mobile nonisotropic communication channel. International Journal of Computer Applications, 70(??):1–6, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume70/number28/12252-7820/.

Bahazaq:2014:EEC

[BT14a] Fahad Mohammed Bahazaq and Vijey Thayananthan. Energy efficiency with computational topology and routing in wireless sensors networks. International Journal of Computer Applications, 89(??):37–42, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume89/number12/15685-4548/.

Burney:2014:MCA

[BT14b] S. M. Aqil Burney and Humera Tariq. K-means cluster analysis for image segmentation. International Journal of Computer Applications, 96(??):1–8, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume96/number4/16779-6360/.

Bandyopadhyay:2016:IEO

[BT16] Soma Bandyopadhyay and S. S. Thakur. ICT in education: Open source software and its impact on teachers and students. International Journal of Computer Applications, 151(??):19–24, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume151/number6/26238-2016911915/.

Bhavsar:2014:DHS

[BV14] Namrata P. Bhavsar and Brijesh Vala. Design of hard and soft decision decoding algorithms of LDPC. International Journal of Computer Applications, 90(??):10– 15, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume90/
number16/15803-4653/.

Bhuvaneswari:2021:TFR

[BV21] R. Bhuvaneswari and S. Ganesh Vaidyanathan. Transgender face recognition using ROI based convolutional neural network. International Journal of Computer Applications, 183(??):1–4, June 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number8/31944-2021921360/.

Billah:2015:PCS

[BWH15] Mustain Billah, Sajjad Waheed, and Abu Hanifa. Predicting closing stock price using artificial neural network and adaptive neuro fuzzy inference system (ANFIS): the case of the dhaka stock exchange. International Journal of Computer Applications, 129(??):1–5, November 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume129/number11/23114-2015906952/.

Bahri:2012:FIM

[BWS^{+}12] Nejmeddine Bahri, Imen Werda, Amine Samet, Mohamed Ali Ben Ayed, and Nouri Masmoudi. Fast intra mode decision algorithm for H264/AVC HD baseline profile encoder. International Journal of Computer Applications, 37 (??):8–13, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume37/
number6/4610-6601/.

Boulahia:2016:NIM

[BWS16] Zoubair Boulahia, Abderrahim Wakif, and Rachid Sehaqui. Numerical investigation of mixed convection heat transfer of nanofluid in a lid driven square cavity with three triangular heating blocks. International Journal of Computer Applications, 143(??):37–45, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume143/number6/25083-2016910227/.

Borandag:2017:EPR

[BYKO17] Emin Borandag, Fatih Yucalar, Deniz Kilinc, and Akin Ozcift. The emergency position reporting system and a sample application. International Journal of Computer Applications, 167(??):34–36, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number2/27746-2017914188/.

Bhola:2014:MLB

[BYT14] Amit Bhola, Sanjeev Kumar Yadav, and Arvind Kumar Tiwari. Machine learning based approach for protein function prediction using sequence derived properties. International Journal of Computer Applications, 105(??):17–21, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number12/18429-9789/.

Chaabane:2011:SMB

[CA11] Lamiche Chaabane and Moussaoui Abdelouahab. Segmentation of MR brain images using a data fusion approach. International Journal of Computer Applications, 36(??):27– 32, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume36/
number12/4553-6450/.

Cuzanauskas:2016:IWM

[CA16] Tomas Cuzanauskas and Aurimas Anskaitis. Innovative wireless medium control algorithm for wireless devices in ISM bands. International Journal of Computer Applications, 146(??):1–5, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number6/25399-2016910780/.

Cai:2017:IWE

[Cai17] Dawei Cai. An innovative wearable electronic tourist guide device and automatic guide service. International Journal of Computer Applications, 180(??):1–7, December 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume180/number4/28785-2017916010/

.

Cai:2018:IUP

[Cai18] Dawei Cai. An indoor ultrasonic positioning system with fast TOA approach in distance space. International Journal of Computer Applications, 182(??):1–6, September 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number17/29951-
2018917868/.

Carneiro:2014:TBT

[CAL^{+}14] Romulo Ferrer L. Carneiro, Jessyca Almeida Bessa, Jermana Lopes De Moraes, Edson Cavalcanti Neto, and Auzuir Ripardo De Alexandria. Techniques of binarization, thinning and feature extraction applied to a fingerprint system. International Journal of Computer Applications, 103(??):1–8, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number10/18107-9291/.

Cardoso:2018:SPE

[Car18]  Fabiano Stingelin Cardoso.  System and process of

electric energy cogeration for data centers environment servers. International Journal of Computer Applications, 181(??):30–38, July 2018. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume181/number1/29682-2018917397/.

Chitkara:2014:SHH

[CAS14] Neha Chitkara, Priya Aggarwal, and Priya Singhal. Study habits of higher performing engineering students: a survey. International Journal of Computer Applications, 97 (??):33–37, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number2/16982-7029/.

Dwi:2012:PLP

[CB12] Andharini Dwi C and Ari Basuki. Personalized learning path of a Web-based learning system. International Journal of Computer Applications, 53(??):17–22, September 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume53/number7/8434-
2206/.

Choubey:2016:RAI

[CB16] Poonam Choubey and Rupali Bhartiya. Result assessment to intrusion detection system using factors analysis in MANET. International Journal of Computer Applications, 152(??):20–25, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume152/number5/26315-2016911758/.

Chermitti:2012:DLC

[CBhM12] Ali Chermitti, Omar Boukli-hacene, and Samir Mouhadjer. Design of a library of components for autonomous photovoltaic system under Matlab/Simulink. International Journal of Computer Applications, 53(??):13–19, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number14/8488-2434/.

Chung:2017:FLB

[CC17] Won-Jin Chung and Tae-Ho Cho. Fuzzy logic based efficient route determination method for improving the energy efficiency of sensor networks in FAP-based WSNs. International Journal of Computer Applications, 172(??): 35–39, August 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume172/number5/28250-2017915164/.

Uzochukwu:2020:HPM

[CCE20] Onwuachu Uzochukwu C, Ugwu C, and Williams Edem. High performance model for handling machine breakdown in identical parallel machines. International Journal of Computer Applications, 177(??):11–19, March 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number48/31232-2020919965/.

Choudhury:2012:SAF

[CCG12] Raja Roy Choudhury, Arundhati Roy Choudhury, and Mrinal Kanti Ghose. Semi analytical formulation of dispersion characteristics for a step and parabolic-index optical fiber: Introduction of Nelder–Mead simplex method for nonlinear unconstrained minimization. International Journal of Computer Applications, 49(??):38–44, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume49/number22/7906-1292/.

Chen:2019:FSG

[CCG19] Jun Chen, Dingkai Chen, and Shuangcheng Ge. Filtering of seawall GPR signal by means of multi-wavelet transform. International Journal of Computer Applications, 178(??):13–18, August 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number38/30784-2019919255/.

Chandrasekar:2014:ERS

[CCJ14] Purnima Chandrasekar, Santosh Chapaneri, and Deepak Jayaswal. Emotion recognition from speech using discriminative features. International Journal of Computer Applications, 101(??):31–36, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number16/17775-8913/.

Chowdhury:2013:ESC

[CD13] M. M. Aftab Chowdhury and Kaushik Deb. Extracting and segmenting container name from container images. International Journal of Computer Applications, 74 (??):18–22, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number19/13001-0039/.

Chafik:2016:MPI

[CD16] Sanaa Chafik and Cherki Daoui. A modified policy iteration algorithm for discounted reward Markov decision processes. International Journal of Computer Applications, 133(??):28–33, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume133/number10/23823-2016908033/.

Chutia:2017:NSU

[CDC17] Muhim Chutia, Tabendra Nath Das, and Pranob Jyoti Chetia. Numerical solution of unsteady hydromagnetic Couette flow in a rotating system bounded by porous plates with Hall effects. International Journal of Computer Applications, 171(??):1–6, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume171/number2/28150-2017914967/.

Cico:2018:IDE

[CDC18] Orges Cico, Zamir Dika, and Betim Cico. Integrated development environment as a service (IDEaaS) — models and architecture part of the Google Cloud Core Services. International Journal of Computer Applications, 182(??):44–50, July 2018. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume182/number6/29770-2018917534/.

Chaabouni:2011:VPI

[CFB11] Imen Chaabouni, Wiem Fourati, and M. Salim Bouhlel. Visushrink pretreatment for image compression. International Journal of Computer Applications, 23(??):10–16, June 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume23/number6/2893-
3785/.

Cetin:2018:PAZ

[ÇGDT18] Gürcan Çetin, Hüseyin Gürüler, Mehmet Osman Devrim, and Sefa Tuncay. Performance analysis of Zig-Bee protocol in the large-scale network topologies. International Journal of Computer Applications, 181(??): 11–15, August 2018. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume181/number8/29791-2018917611/.

Chowdhury:2013:FDI

[CGKP13] A Chowdhury, H Gebretsadik, S Kumar, and R M Punekar. Feasibility of development of an immunization management tool for mobile phones by UCD approach. International Journal of Computer Applications, 61(??):26– 31, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number18/10030-5057/.

Charniya:2012:ACL

[Cha12] Nadir Charniya. Automated car license plate localization using wavelet analysis. International Journal of Computer Applications, 42(??):41–45, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number2/5669-7702/.

Chaturvedi:2014:AQI

[Cha14] Jyoti Chaturvedi. Adaptive quantum inspired genetic algorithm for combinatorial optimization problems. International Journal of Computer Applications, 107(??): 34–42, December 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume107/number4/18743-9996/.

Chawla:2015:OCW

[Cha15] Suruchi Chawla. Optimization of clusters of Web query sessions using genetic algorithm for effective personalized Web search. International Journal of Computer Applications, 122(??):9–17, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number9/21726-4885/.

Chawda:2017:SEC

[Cha17] Pradeep Chawda. A simplified equivalent circuit model of MEMS electrostatic actuator. International Journal of Computer Applications, 160(??):17–23, February 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume160/number9/27101-2017913097/.

Cieutat:2012:ALB

[CHG12] Jean-Marc Cieutat, Olivier Hugues, and Nehla Ghouaiel. Active learning based on the use of augmented reality outline of possible applications: Serious games, scientific experiments, confronting studies with creation, training for carrying out technical skills. International Journal of Computer Applications, 46(??):31–36, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume46/number20/7059-9800/.

Castillo:2016:MAC

[CHW16] Edward A. Castillo, Sheng-Jen Hsieh, and Yunyi Wang. Miniaturized automated color sorting system using programmable logic control and Web-based server. International Journal of Computer Applications, 143(??):12–20, June 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume143/number2/25048-2016910015/

.

Chauhan:2010:QGH

[CJ10] Sameer Singh Chauhan and R. C. Joshi. QoS guided heuristic algorithms for grid task scheduling. International Journal of Computer Applications, 2(??):24–31, June 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume2/number9/694-975/.

Carlos:2019:MAD

[CJC19] Kele Mbbe Ripaul Carlos, Nkenlifack Marcellin Julius, and Kamla Vivien Corneille. Multi-agent decision-making model for the construction of schools. International Journal of Computer Applications, 178(??):4–13, August 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number40/30799-2019919281/.

Chandak:2012:DLB

[CJK^{+}12] Akshay Chandak, Krishnakant Jaju, Akshay Kanfade, Amit Joshi, and Pushkar Lohiya. Dynamic load balancing of virtual machines using QEMU-KVM. International Journal of Computer Applications, 46(??):10–14, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume46/number6/6910-9263/.

Charaya:2016:CBI

[CJK16] Sahil Charaya, Sonika Jindal, and Bhavneet Kaur. Content based image retrieval using selective region matching with region of interest and SVM. International Journal of Computer Applications, 137(??):28–33, March 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume137/number3/24257-2016908667/.

Choudhary:2017:MQT

[CJK17] Sudhakar Choudhary, Arvind Kumar Jain, and Anil Kumar. Model query, tokenization and character matching: a combined approach to prevent SQLIA. International Journal of Computer Applications, 162(??):13–18, March 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume162/number9/27270-2017913357/.

Chandrasekaran:2012:ALR

[CK12] R. Chandrasekaran and R. Venkatesh Kumar. Application of logistic regression to predict over target baseline of software projects. International Journal of Computer Applications, 44(??):1–6, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume44/number10/6296-8493/.

Ananth:2013:CLB

[CK13] Anbu Ananth C and Selvakumar K. A cross layer based channel assignment algorithm in multi radio multi channel wireless mesh networks. International Journal of Computer Applications, 61(??):7–13, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume61/number21/10202-4900/.

Ciirah:2018:CCE

[CKA18] Stephen M. Ciirah, Andrew M. Kahonge, and Elisha O. Abade. Cloud computing: an empirical study of the correlation between scalability and throughput of cloud-based applications. International Journal of Computer Applications, 179(??):10–17, April 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number33/29209-2018916741/.

Chugh:2012:CAR

[CKD12] Renu Chugh, Vivek Kumar, and Ombir Dahiya. Comparative analysis of rate of convergence of Agarwal et al., Noor and SP iterative schemes for complex space. International Journal of Computer Applications, 41(??):6–15, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume41/number11/5583-7819/.

Chahal:2013:SWM

[CKS13] Nidhi Chahal, Vinod Karar, and Manpreet Singh. Study of writing methodologies for graphical display on flat cathode ray tube. International Journal of Computer Applications, 74(??):29–32, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume74/number7/12900-9826/.

Chede:2010:SVT

[CKT10] Santosh Chede, Kishore Kulat, and Rajesh Thakare. A significance of VLSI techniques for low power real time systems. International Journal of Computer Applications, 1(??):86–91, February 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume1/number22/437-
666/.

Chtioui:2012:DHC

[CLBA^{+}12] H. Chtioui, S. Niar Lamih, R. Ben-Atitallah, M. Zahran, Jl. Dekeyser, and M. Abid. A dynamic hybrid cache coherency protocol for shared-memory MPSoC architectures. International Journal of Computer Applications, 47 (??):45–50, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number3/7172-9801/.

Chitla:2012:AIT

[CM12] Arathi Chitla and M. Chandra Mohan. Authentication of images through lossless watermarking (LWM) technique with the aid of elliptic curve cryptography (ECC). International Journal of Computer Applications, 57(??):17– 25, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number6/9118-3275/.

Chand:2013:RTV

[CM13] Padmabati Chand and J. R. Mohanty. Real time vehicle routing problem with time window simultaneously delivery products and pick up wastage materials with proposed masterslave genetic algorithm. International Journal of Computer Applications, 72(??):39–46, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume72/number19/12653-9359/.

Chaugule:2015:DSO

[CM15a] Archana Chaugule and Suresh N. Mali. Digital separation of occluded seeds using image analysis. International Journal of Computer Applications, 123(??): 30–37, August 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume123/number4/21950-2015905284/.

Chaurasia:2015:RDR

[CM15b] Anil Kumar Chaurasia and Rajesh Mehra. Robust design of RF MEMS switch design with reduced buckeling effect. International Journal of Computer Applications, 119(??):1–6, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number24/21381-4376/.

Chavan:2015:SCS

[CMKS15] Akshay Chavan, Shantanu Mudagi, G. V. Raghu Teja Kumar, and Prithvi Sharma. Smart cloud storage using home automation support. International Journal of Computer Applications, 112(??):33–40, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number10/19706-1552/.

Chiplunkar:2010:SBP

[CN10] Niranjan N. Chiplunkar and Yuvaraju B N. Scenario based performance analysis of variants of TCP using NS2 — simulator. International Journal of Computer Applications, 4(??):20–24, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume4/number9/855-1197/.

Chandrasekara:2012:GNN

[CN12] Bagavathi Chandrasekara and K A Narayanankutty. Gallium nitride nanotube and its application as transistors. International Journal of Computer Applications, 47(??): 39–44, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number14/7259-0347/.

Christodoulou:2019:ENP

[CNP19] Dimitrios Christodoulou, Nikitas Nikitakos, and Dimitrios Papachristos. The effects of the naval profession on the seafarers’ mental health. International Journal of Computer Applications, 178(??):7–14, June 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number24/30680-2019919035/.

Chao:2014:SAN

[COWK14] Chen Chao, Wataru Ohyama, Tetsushi Wakabayashi, and Fumitaka Kimura. Semi-automatic noninvasive tracking of local myocardial motion using dynamic programming based elastic model. International Journal of Computer Applications, 101(??):32–38, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number5/17686-8546/.

Chaki:2012:PLR

[CP12] Jyotismita Chaki and Ranjan Parekh. Plant leaf recognition using Gabor filter. International Journal of Computer Applications, 56(??):26–29, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume56/number10/8927-3000/.

Choudhury:2014:VEE

[CP14] Rita Choudhury and Saswati Purkayastha. Visco-elastic effects on convection flow in a vertical rotating channel partially filled with a porous medium. International Journal of Computer Applications, 104(??):9–17, October 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume104/number4/18189-9098/

.

Ch:2015:BDC

[CP15a] Laxmiprasanna Ch and Ramesh Palakeerthi. BLDC drive control using artificial intelligence technique. International Journal of Computer Applications, 118(??):5–9, May 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume118/number4/20731-3100/

.

Chaturvedi:2015:SES

[CP15b] D. K. Chaturvedi and Atul Prakash Prajapati. Survey of expert systems and the cognitive approaches towards an effective tutoring system. International Journal of Computer Applications, 132(??):1–7, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number5/23587-2015907372/.

Chettri:2013:SDC

[CPD13] Sarat Kumar Chettri, Bonani Paul, and Ajoy Krishna Dutta. Statistical disclosure control for data privacy preservation. International Journal of Computer Applications, 80 (??):38–43, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume80/
number10/13899-1880/.

CostaDeAraujo:2015:SMR

[CPG15] Jucélio Costa De Araújo, José Maria Parente De Oliveira, and Leonardo Garcia Marques. Semantic method of recovering of medical articles. International Journal of Computer Applications, 128(??):26–32, October 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume128/number16/22959-2015906750/.

Chauhan:2013:WPA

[CPSP13] Jeetender Singh Chauhan, Gyan Prabhakar, Sunil Semwal, and Atul Kumar Pandey. Wireless personal area network based simulation and design to control the speed of permanent magnet DC motor using Zigbee transceiver protocol. International Journal of Computer Applications, 69 (??):6–13, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number23/12109-8303/.

Chaturvedi:2015:QIG

[CQM15] D. K. Chaturvedi, Tanveer Qamar, and O. P. Malik. Quantum inspired GA based neural control of inverted pendulum. International Journal of Computer Applications, 122(??):46–52, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number23/21869-5210/.

Chelvi:2012:PPA

[CR12] T. Kalai Chelvi and P. Rangarajan. Principal pattern analysis: a combined approach for dimensionality reduction with pattern categorization. International Journal of Computer Applications, 41(??):35–41, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume41/number6/5548-7616/.

Chandrasekhar:2015:OCP

[CR15] M. Chandrasekhar and G. S. N. Raju. Optimized cosecant patterns from arrays of discrete sources. International Journal of Computer Applications, 113(??): 11–16, March 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume113/number8/19846-1709/.

Chugh:2012:CLM

[CRA12] Renu Chugh, Mamta Rani, and Ashish. On the convergence of logistic map in NOOR orbit. International Journal of Computer Applications, 43(??):1–4, April 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume43/number18/6200-8739/.

Chattopadhyay:2012:MIS

[CRBJ12] Subhagata Chattopadhyay, Fethi A. Rabhi, Sangeeta Bhanjachaudhuri, and Sarat K. Jena. Mining important symptoms of adult depression. International Journal of Computer Applications, 57(??):5–9, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume57/number5/9108-3258/.

Chauhan:2010:NJS

[CRN10] Yashwant S. Chauhan, Rajeshri Rana, and Ashish Negi. New Julia sets of Ishikawa iterates. International Journal of Computer Applications, 7(??):34–42, October 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume7/number13/1321-1675/.

Choudhary:2010:PEL

[CS10] Neeraj Choudhary and Ajay K. Sharma. Performance evaluation of LR-WPAN for different path-loss models. International Journal of Computer Applications, 7(??):22–28, October 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume7/number10/1284-
1698/.

Chhabra:2011:SME

[CS11] Amit Chhabra and Gurvinder Singh. Statistical modeling and evaluation of parallel space-sharing job scheduling algorithms for PC-cluster using design of experiments (DOE). International Journal of Computer Applications, 25(??):17–24, July 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume25/
number11/3158-4370/.

Chaudhary:2012:ARM

[CS12] Manish Singh Chaudhary and Varsha Singh. An adaption of reactive measure for enhancement and performance characterization of routing protocol under different traffic source. International Journal of Computer Applications, 44 (??):30–33, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number23/6430-8928/.

Carter:2013:DIP

[CS13] S. B. Ron Carter and A. Selvaraj. Design and implementation of PLC based elevator. International Journal of Computer Applications, 68(??):4–10, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number7/11589-6927/.

Chandra:2014:ITW

[CS14] Vivek Chandra and Nidhi Saxena. An improved technique for Web page classification in respect of domain specific search. International Journal of Computer Applications, 102(??):7–10, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume102/number4/17801-8615/.

Chomal:2015:STE

[CS15a] Vikas S. Chomal and Jatinderkumar R. Saini. Software template for evaluating and scoring software project documentations. International Journal of Computer Applications, 116(??):15–27, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number1/20300-2333/.

Choudhury:2015:VCB

[CS15b] Nupur Choudhury and Chingtham Tejbanta Singh. Voice controlled BOEbot using PointBug algorithm for human robot interaction using Android technology. International Journal of Computer Applications, 119(??):7–13, June 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume119/number13/21125-
3705/.

Chuchra:2015:MIT

[CS15c] Rimmy Chuchra and R. K. Seth. Modeling implementation of TBDSA-token based data security algorithm in cloud computing. International Journal of Computer Applications, 114(??):12–17, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume114/number7/19989-1619/.

Chandrakala:2016:HHC

[CS16a] K. Chandrakala and D. Saraswady. HCFA: Hybrid coupled flat architecture for QoS enhancement of multimedia applications in mobile WiMAX broadband access network. International Journal of Computer Applications, 148(??):31–37, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume148/number13/25820-2016911303/.

Chao:2016:AMG

[CS16b] Brandon Chao and Ankit Sirmorya. Automated movie genre classification with LDA-based topic modeling. International Journal of Computer Applications, 145(??):1–5, July 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume145/number13/25335-
2016910822/.

Choukse:2018:SAA

[CS18] Dharmendra Choukse and Umesh Kumar Singh. To study the assign authorization to access the object through single sign-on Web services in library management system (LMS). International Journal of Computer Applications, 179(??):34–39, February 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number20/28988-2018916364/.

Chanda:2011:CEC

[CSKB11] Jayeeta Chanda, Sabnam Sengupta, Ananya Kanjilal, and Swapan Bhattacharya. CA-ESB: Context aware enterprise service bus. International Journal of Computer Applications, 30(??):1–8, September 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume30/number3/3626-5062/.

Cavalcante:2021:QEQ

[CSM21] Haniel G. Cavalcante, Jéferson N. Soares, and José E. B. Maia. Question expansion in a question-answering system in a closed-domain system. International Journal of Computer Applications, 183(??):1–5, September 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume183/number23/32064-2021921621/.

Chaudhary:2015:GLR

[CSN15] Sushank Chaudhary, Abhishek Sharma, and Neetu. 6 × 20 Gbps long reach WDM-PI based high altitude platform intersatellite communication system. International Journal of Computer Applications, 122(??):41–45, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume122/number22/21861-5192/.

Chaudhary:2014:RTW

[CSP14] Sushank Chaudhary, Saurabh Sharma, and Parveen. Role of turbulences in WDM-polarization interleaving scheme based inter-satellite communication system. International Journal of Computer Applications, 104(??):1–7, October 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume104/number10/18235-
9224/.

Chitaliya:2010:EMF

[CT10] N. G. Chitaliya and A. I. Trivedi. An efficient method for face feature extraction and recognition based on contourlet transform and principal component analysis using neural network. International Journal of Computer Applications, 6(??):28–34, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume6/number4/1066-1260/.

Arunpriya:2013:FST

[CT13] Arunpriya C and Antony Selvadoss Thanamani. Fuzzy shrink thresholding based tea leaf image enhancement using wavelet transform. International Journal of Computer Applications, 63(??):1–5, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume63/number18/10563-5441/.

Chaturvedi:2014:ERU

[CT14] Akanksha Chaturvedi and Alpika Tripathi. Emotion recognition using fuzzy rule- based system. International Journal of Computer Applications, 93(??):25–28, May 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume93/number11/16260-5920/.

Chaturvedi:2010:AZA

[CTG10] Nitin Chaturvedi, Jithin P. Thoma, and S Gurunarayanan. Adaptive zone-aware multi-bank on chip last level L2 cache partitioning for chip multiprocessors. International Journal of Computer Applications, 7(??):19–23, September 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume7/number1/1131-1482/

.

Chand:2012:TOS

[CTG12] Chetna Chand, Amit Thakkar, and Amit Ganatra. Target oriented sequential pattern mining using recency and monetary constraints. International Journal of Computer Applications, 45(??):12–18, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number10/6815-9169/.

Cufoglu:2014:UPS

[Cuf14] Ayse Cufoglu. User profiling — a short review. International Journal of Computer Applications, 108(??): 1–9, December 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume108/number3/18888-0179/.

Curia:2020:UHA

[Cur20] Francesco Curia. Unsupervised hybrid algorithm to detect anomalies for predicting terrorists attacks. International Journal of Computer Applications, 176(??):1–8, July 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume176/number35/31424-
2020920432/.

Chavan:2013:MFD

[CV13] Anupama A Chavan and Vijay Kumar Verma. Mining functional dependency in relational databases using FUN and dep-miner: a comparative study. International Journal of Computer Applications, 78(??):34– 36, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume78/
number15/13602-1377/.

Cahyani:2023:TSP

[CW23] Catharina Adinda Mega Cahyani and Trianggoro Wiradinata. Traveling salesman problem multi-destination route recommendation system using genetic algorithm and Google Maps API. International Journal of Computer Applications, 185(??):35–43, June 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number18/32798-2023922907/.

Chen:2015:ANN

[CWC15] Zhenfeng Chen, Zhongsheng Wang, and Jian Cen. Adaptive neural network control for a class of MIMO uncertain pure-feedback nonlinear systems. International Journal of Computer Applications, 124(??):1–5, August 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume124/number16/22192-2015905776/.

Chua:2015:FRD

[CWT15] Shing Chyi Chua, Eng Kiong Wong, and Alan Wee Chiat Tan. Fingerprint ridge distance estimation: a mathematical modeling. International Journal of Computer Applications, 126(??):24–29, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume126/number15/22629-2015906308/.

Chen:2015:DAC

[CZW15a] Zhenfeng Chen, Xuhong Zhang, and Zhongsheng Wang. Direct adaptive control for a class of uncertain nonlinear systems. International Journal of Computer Applications, 119(??):11–15, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number16/21150-4168/.

Chen:2015:OFS

[CZW15b] Zhenfeng Chen, Xuhong Zhang, and Zhongsheng Wang. Output feedback stabilization of a class of MIMO uncertain nonaffine nonlinear systems. International Journal of Computer Applications, 124(??):1–5, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number17/22196-2015905777/.

Deshmukh:2010:PEA

[DA10] Rajesh Deshmukh and Asha Ambhaikar. Performance evaluation of AODV and DSR with reference to network size. International Journal of Computer Applications, 11(??):27– 32, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number8/1601-2150/.

Dinu:2012:NMO

[DA12a] Simona Dinu and Gabriela Andrei. A new multi-objective genetic algorithm for use in investment management. International Journal of Computer Applications, 58(??):1– 8, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number13/9339-3656/.

Dutta:2012:HMD

[DA12b] Palash Dutta and Tazid Ali. A hybrid method to deal with aleatory and epistemic uncertainty in risk assessment. International Journal of Computer Applications, 42 (??):37–43, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number11/5740-7930/.

Duru:2018:IQS

[DA18] Chinedu Duru and Mamilus Ahaneku. Improved quality of service for video, data and voice integrated wireless cellular networks through the queue dynamic channel reservation scheme. International Journal of Computer Applications, 180(??):17–26, January 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number9/28890-2018916123/.

Dama:2013:NAO

[DAaGZ13] Y. A. S. Dama, R. A. Abd-alhameed, T. S. Ghazaany, and S. Zhu. A new approach for OSTBC and QOSTBC. International Journal of Computer Applications, 67(??): 45–48, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number6/11403-6719/.

Delavar:2012:KEW

[DAB12] Arash Ghorbannia Delavar, Abootorab Alirezaie, and Amir Abbas Baradaran. KGAWSN: an effective way to reduce energy consumption in wireless sensor networks by K-means and genetic algorithms. International Journal of Computer Applications, 48(??):25–28, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.
org/archives/volume48/number12/7401-0328/.

Dahab:2016:AIR

[DAB^{+}16] Mohamed Y. Dahab, Ablah AlAmri, Bayan Bagasi, Ebtesam AlMalki, and Ola AlBeshri. Automatic identifying rhythm of Arabic poem. International Journal of Computer Applications, 153(??):21–36, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume153/number1/26368-2016911956/.

Dhanalakshmi:2016:IFM

[DAG16] P. Dhanalakshmi, S. Arunpriya, and S. Gayathiriprabha. On intuitionistic fuzzy multi weakly generalized closed set. International Journal of Computer Applications, 146(??):21–25, July 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume146/number13/
25458-2016910905/.

Dake:2023:ORF

[Dak23] Delali Kwasi Dake. Online recruitment fraud detection: a machine learning-based model for Ghanaian job websites. International Journal of Computer Applications, 184(??):20– 28, March 2023. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume184/number51/
32652-2023922639/.

Daryasafar:2012:SCE

[Dar12] Navid Daryasafar. Synchronization and channel estimation in MIMO-OFDM systems. International Journal of Computer Applications, 45(??):12–18, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number13/6839-9364/.

Darvishi:2013:NDL

[Dar13] Mostafa Darvishi. A novel developed linear and nonlinear system identification for an industrial dryer. International Journal of Computer Applications, 74(??): 27–35, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number10/12921-9908/.

Dev:2010:RFN

[DB10] Amita Dev and Poonam Bansal. Robust features for noisy speech recognition using MFCC computation from magnitude spectrum of higher order autocorrelation coefficients. International Journal of Computer Applications, 10(??):36– 38, November 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume10/
number8/1499-2016/.

Dilekh:2012:INH

[DB12a] Tahar Dilekh and Ali Behloul. Implementation of a new hybrid method for stemming of Arabic text. International Journal of Computer Applications, 46(??):14–19, May 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume46/number8/6927-
9344/.

Djemame:2012:CCA

[DB12b] Safia Djemame and Mohamed Batouche. Combining cellular automata and particle swarm optimization for edge detection. International Journal of Computer Applications, 57 (??):16–22, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number14/9182-3602/.

Devi:2013:PAF

[DB13a] O. Mema Devi and Shahin Ara Begum. Performance analysis of fuzzy competitive learning algorithms for MR image segmentation. International Journal of Computer Applications, 71(??):6–13, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number9/12384-8736/.

Drif:2013:EED

[DB13b] Ahlem Drif and Abdellah Boukerram. Energy efficient DSR algorithm based on topology for mobile ad-hoc network. International Journal of Computer Applications, 83 (??):1–4, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number11/14489-2849/.

Deshpande:2017:PIT

[DB17] Minal Deshpande and Preeti Bajaj. Performance improvement of traffic flow prediction model using combination of support vector machine and rough set. International Journal of Computer Applications, 163(??):31–35, April 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume163/number2/27370-2017913473/

.

Delavar:2012:BSAa

[DBB12] Arash Ghorbannia Delavar, Ali Reza Khalili Boroujeni, and Javad Bayrampoor. A balanced scheduling algorithm with fault tolerant and task migration based on primary static mapping (PSM) in grid. International Journal of Computer Applications, 52(??):10–21, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume52/number8/8221-1649/.

Dzebu:2024:RPP

[DBB24] Crystal Dzebu, Bernard Adjei Buckman, and Evans Boye. A robust privacy-preserving architecture for image copyright authentication based on the blind watermarking approach and elliptic curve cryptography. International Journal of Computer Applications, 186(??):53–60, December 2024. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume186/number55/a-
robust-privacy-preserving-architecture-for-image-
copyright-authentication-based-on-the-blind-watermarking-
approach-and-elliptic-curve-cryptography/.

Dahdouh:2020:PEO

[DBKN20] Marwa Dahdouh, Amer Bouchi, Souheil Khawatmi, and Mouhamad Ayman Naal. Programmatic effect of optimized smali code on saving energy of Android applications. International Journal of Computer Applications, 177(??):33–41, March 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume177/number42/
31186-2020919928/.

Deghaidy:2014:WRF

[DBM14] M. Maged M. Deghaidy, Khaled Mahmoud Badran, and Gouda Ismail Mohamed. Web recommendation framework based on association rules coverage to be applied for site modification. International Journal of Computer Applications, 91(??):28–33, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume91/number2/15854-4754/.

S:2015:EOU

[DBS15] Ruthvic S D, Ravi B, and Udaya Kumar Shenoy. Energy optimization using neighborhood based weighted rendezvous technique for wireless sensor networks. International Journal of Computer Applications, 120(??):1–6, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume120/number8/21244-4025/.

Doringin:2019:PMC

[DBW19a] Fanny Doringin, Marson James Budiman, and Stephy Walukow. 3 phase monitoring and control control systems based on Labview. International Journal of Computer Applications, 178(??):21–24, June 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number16/30619-2019918940/.

Doringin:2019:BAC

[DBW19b] Fanny Doringin, Marson James Budiman, and Stephy Walukow. Balance automatic control system load 3 phase. International Journal of Computer Applications, 177(??):33–37, October 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number9/30927-2019919472/.

Dias:2022:ECS

[DC22] Cleber Gustavo Dias and Rodrigo Cardozo De Jesus. An experimental comparative study between machine learning and signal processing techniques to detect broken rotors bars in induction motors. International Journal of Computer Applications, 184(??):1–8, October 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number30/32501-2022922351/.

Dang:2016:DPP

[DCH16] Tuan Linh Dang, Thang Cao, and Yukinobu Hoshino. Data pre-processing for a neural network trained by an improved particle swarm optimization algorithm. International Journal of Computer Applications, 154(??): 1–8, November 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume154/number1/26452-2016912022/.

Dhurvey:2015:NMM

[DCM15] Hemendra Dhurvey, Hirdesh Chack, and O. P. Meena. A novel multi-minislot cooperative spectrum sensing scheme in cognitive radio networks. International Journal of Computer Applications, 125(??):21–26, September 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume125/number10/22469-2015906036/.

Deb:2013:ASF

[DD13] Moumita Deb and P. K. De. Algorithm for solving fuzzy multiobjective linear fractional programming problem by additive weighted method. International Journal of Computer Applications, 80(??):1–6, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume80/number2/13830-1428/.

Danbatta:2014:DPS

[DD14] Kafilu Barau Danbatta and Umar Garba Danbatta. Dynamic pricing scheme for effective bandwidth utilization in mobile Wimax. International Journal of Computer Applications, 104(??):7–12, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number15/18276-8982/.

Dandavate:2015:DLD

[DD15] P. P. Dandavate and S. S. Dhotre. Data leakage detection using image and audio files. International Journal of Computer Applications, 115(??):1–4, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number8/20168-2306/.

Dias:2020:TMA

[DD20] Cleber Gustavo Dias and Alaydes Mikaelle De Morais. A text mining approach in patents and papers: Analyzing wind and solar energy trends. International Journal of Computer Applications, 175(??):1–12, December 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number37/31690-2020920947/.

Dias:2021:VLT

[DD21] Cleber Gustavo Dias and Luiz Carlos Da Silva. A virtual-lab tool for teaching the fundamentals of a DC motor-generator operation using excel-VBA. International Journal of Computer Applications, 183(??):1–8, September 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume183/number24/32072-2021921627/.

Dashtban:2011:NAV

[DDB11] Muhammad H. Dashtban, Zahra Dashtban, and Hassan Bevrani. A novel approach for vehicle license plate localization and recognition. International Journal of Computer Applications, 26(??):22–30, July 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume26/number11/3167-4382/.

Domingues:2022:CIS

[DDBD22] Andreia M. Domingues, Sabrinna Delgado, Marcia A. S. Bissaco, and Sidnei A. De Araújo. Computational intelligence in serious games: a case study to identify patterns in a game for children with learning disabilities. International Journal of Computer Applications, 184(??):40–44, April 2022. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number7/32344-2022922044/.

Datta:2012:CTA

[DDC12] Subhasis Datta, Avijan Dutta, and Subhrabrata Choudhury. Cost-throughput analysis of a modified TCP for optical burst switched networks. International Journal of Computer Applications, 60(??):1–7, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume60/number16/9773-4312/.

Dias:2024:MAD

[DDC^{+}24] Roberto Alexandre Dias, Mário De Noronha Neto, Gregory Chagas Da Costa Gomes, Fernando Benites, Stefan Gürtler, Siqueira De Moraes Neto, and Níkola Zaia De Figueiredo. Methodology for anomaly detection and alert generation in photovoltaic systems. International Journal of Computer Applications, 186(??):9–15, August 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number36/
methodology-for-anomaly-detection-and-alert-generation-
in-photovoltaic-systems/.

Dhore:2012:IHE

[DDD12] M. L. Dhore, S. K. Dixit, and Ruchi M. Dhore. Issues in Hindi to English and Marathi to English machine transliteration of named entities. International Journal of Computer Applications, 51(??):37–44, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume51/number14/8112-1728/.

Dey:2012:PDA

[DDDD12] Arijit Dey, Kunal Das, Debashis De, and Mallika De. Probabilistic defect analysis model for quantum dot cellular automata design at analytical phase. International Journal of Computer Applications, 55(??):33–41, October 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume55/number7/8768-
2693/.

Dixit:2016:DCM

[DDG16] Anuja Dixit, Rahul Dixit, and R. K. Gupta. Detection of copy-move forgery exploiting LBP features with discrete wavelet transform. International Journal of Computer Applications, 153(??):1–10, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume153/number3/26380-2016911979/.

Dias:2016:OBS

[DDL16] Roberto Alexandre Dias, Gregory Das Chagas Gomes, and Marcelo Lobo Heldwin. Optimizer in a box: Simple economic optimization of power dispatch in small microgrids using OPC-UA. International Journal of Computer Applications, 156(??):27–33, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume156/number9/26739-2016912569/.

DeBorba:2018:MPD

[DDL18] Bruno De Borba, Roberto Alexandre Dias, and Fabrizio Leal Freitas. MQTT protocol discovery service for an IoTbased gensets monitoring system. International Journal of Computer Applications, 180(??):9–14, March 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number29/29175-2018916695/.

DeVega:2023:DDD

[DDM^{+}23] Bonaparte A. De Vega, Jairus S. Delos Reyes, Mateo Madrine F. Magalong, Michael M. Mislang, and Jenniea A. Olalia. Driver drowsiness detection and notification through facial pattern analysis. International Journal of Computer Applications, 185(??):6–10, May 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number12/32751-2023922783/.

Dias:2014:EAL

[DDN14] Roberto Alexandre Dias, Tiago Emanoel De Souza, and Valdir Noll. Experimental analysis of the Linux RTpatched for data acquisition applied to power sector. International Journal of Computer Applications, 101(??): 43–49, September 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume101/number6/17695-8668/.

Darold:2020:MLE

[DDS20] Denilton Luiz Darold, Carlos Roberto Da Rolt, and Andrea Sabbioni. Machine learning to estimate the floating population in florianopolis. International Journal of Computer Applications, 175(??):1–6, October 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number27/31620-2020920812/.

DeBarros:2020:BPS

[De 20] Charles F. De Barros. Building provably secure block ciphers from cryptographic hash functions. International Journal of Computer Applications, 176(??):1–7, April 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number16/31282-2020920123/.

DeBarros:2021:VEV

[De 21] Charles F. De Barros. A verifiable electronic voting system with homomorphic tallying using elliptic-curve cryptography. International Journal of Computer Applications, 183(??):1–8, December 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number44/32224-2021921848/.

Daniel:2013:SFI

[DE13] Uchenna Peter Daniel and Gregory Epiphaniou. Safeguarding forensic integrity of virtual environment evidence. International Journal of Computer Applications, 82(??):43– 52, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume82/
number6/14123-2240/.

Debnath:2020:CDM

[Deb20] Tanmay Debnath. A control device to monitor domestic violence using speech analysis. International Journal of Computer Applications, 176(??):12–16, May 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number25/31354-2020920214/.

DeMirandaPereira:2015:SMH

[DF15] Ivana Cardial De Miranda Pereira and Nelson Francisco Favilla Ebecken. A simulation model and a hybrid genetic algorithm for energy-aware MANET routing and planning. International Journal of Computer Applications, 124(??):42–50, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number11/22152-2015905682/.

DAndrea:2015:ATA

[DFGG15] Alessia D’Andrea, Fernando Ferri, Patrizia Grifoni, and Tiziana Guzzo. Approaches, tools and applications for sentiment analysis implementation. International Journal of Computer Applications, 125(??):26–33, September 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume125/number3/22414-2015905866/

.

Dias:2016:EOU

[DGH16] Roberto Alexandre Dias, Gregory Das Chagas Gomes, and Marcelo Lobo Heldwin. Evaluation of OPC-UA framework applied to microgrid energy dispatch management. International Journal of Computer Applications, 154(??):20–27, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume154/number7/26504-2016912213/.

Das:2012:REU

[DGJ12] S. Das, S. K. Guchhait, and R. N. Jana. Radiation effects on unsteady MHD free convective Couette flow of heat generation/absorbing fluid. International Journal of Computer Applications, 39(??):42–51, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number3/4804-7036/.

Dutta:2014:IGR

[DGP14] Raju Dutta, Shishir Gupta, and Debraj Paul. Improvement on gossip routing protocol using TOSSIM in wireless sensor networks. International Journal of Computer Applications, 97(??):41–47, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume97/number21/17135-7809/.

Dhakal:2015:PCA

[DGPP15] Dependra Dhakal, Kiran Gautam, Bishal Pradhan, and Ashis Pradhan. Performance comparision of AODV and DSR routing protocols using real time test-bed. International Journal of Computer Applications, 112(??): 1–7, February 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume112/number2/19635-1209/.

Delavar:2012:BSAb

[DH12] Arash Ghorbannia Delavar and Ayden Halakouie. BGPS: a scheduling algorithm for batch pipeline resources using GA. International Journal of Computer Applications, 57(??): 25–31, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number15/9192-3617/.

Darwish:2017:PSM

[DH17] Ahmed Haj Darwish and Aliaa Hilal. Prediction of stock market using C-means clustering and particle filter. International Journal of Computer Applications, 179(??):12–19, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number2/28707-2017915876/.

Dang:2019:IPA

[DH19] Tuan Linh Dang and Yukinobu Hoshino. Improved PSO algorithm for training of neural network in co-design architecture. International Journal of Computer Applications, 182(??):1–7, March 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number44/30442-2019918583/.

Desai:2021:BPS

[DH21] Pankti Desai and Thaier Hamid. Best practices for securing financial data and PII in public cloud. International Journal of Computer Applications, 183(??):1–6, December 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number40/32192-
2021921737/.

Dhavale:2012:DDB

[Dha12] Sunita V. Dhavale. DWT and DCT based robust iris feature extraction and recognition algorithm for biometric personal identification. International Journal of Computer Applications, 40(??):33–37, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume40/number7/4978-7235/.

Dobosz:2015:HOM

[DHM15] Rafal Dobosz, Richard Hurley, and Sabine Mcconnell. A hybrid OpenMP-MPI parallelization of structure software. International Journal of Computer Applications, 118(??):1–9, May 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume118/number11/
20786-3434/.

Dhar:2013:FRR

[DHT13] Mrinal Kanti Dhar, Quazi M. Hasibul Haque, and Md. Tanjimuddin. Face recognition by radial basis function network (RBFN). International Journal of Computer Applications, 78(??):21–26, September 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume78/number3/13470-1143/.

Daramola:2010:OSR

[DI10] S. Adebayo Daramola and T. Samuel Ibiyemi. Offline signature recognition using hidden Markov model (HMM). International Journal of Computer Applications, 10(??):17– 22, November 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume10/
number2/1454-1967/.

Dixon:2011:DVV

[Dix11] Mark B. Dixon. Defining virtual views of electronic resources using declarative queries. International Journal of Computer Applications, 29(??):27–34, September 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume29/number5/3560-4899/.

Dewangan:2012:BET

[DJ12] Bhoopendra Dewangan and S. A. K. Jilani. Behaviour and effect of temperature over pressure tube of nuclear reactor. International Journal of Computer Applications, 53(??): 11–17, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number11/8464-2381/.

Deshmukh:2013:BDB

[DJA13] Amit A. Deshmukh, Rakesh Jondhale, and Ishitva Ajmera. Broadband and dual band modified rectangular microstrip antennas. International Journal of Computer Applications, 82(??):33–38, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume82/number13/14179-2422/.

Dale:2012:SSH

[DJG12] M. P. Dale, M. A. Joshi, and H. J. Galiyawala. A single sensor hand geometry and palm texture fusion for person identification. International Journal of Computer Applications, 42(??):11–16, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number7/5704-7726/.

Das:2013:FFC

[DJJ13] Sanatan Das, Mrinal Jana, and Rabindra Nath Jana. Free and forced convective flow in a horizontal channel embedded in a porous medium. International Journal of Computer Applications, 65(??):36–41, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume65/number3/10908-5836/.

Damarla:2011:ICD

[DK11] Seshu Kumar Damarla and Madhusree Kundu. Identification and control of distillation process using partial least squares based artificial neural network. International Journal of Computer Applications, 29(??):29– 35, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume29/
number7/3576-4936/.

Damak:2013:PEC

[DK13] Nawel Damak and Samira Kamoun. Parameter estimation and control of induction machine using a new recursive algorithm. International Journal of Computer Applications, 67(??):1–5, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number12/11444-7050/.

Dehghani:2018:BSB

[DK18] Alireza Dehghani and Alireza Kashaniniya. A bioparticle separator based on dielectrophoretic integrated with acoustic force: Design and simulation. International Journal of Computer Applications, 184(??):6–11, January 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number3/32312-2022921987/.

Dhane:2012:GNL

[DKB12] Pranali Dhane, Krishnan Kutty, and Sachin Bangadkar. A generic non-linear method for fisheye correction. International Journal of Computer Applications, 51(??):58– 65, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number10/8082-1483/.

Dubey:2010:NCB

[DKP10] Ashutosh K. Dubey, Ganesh Raj Kushwaha, and Jay Prakash. A novel CSUA based data mining approach for mobile computing environments. International Journal of Computer Applications, 7(??):23–26, September 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume7/number4/1151-1506/.

DeSilva:2021:ASC

[DKP21] Dilshan I. DeSilva, Saluka R. Kodituwakku, and Amalka J. Pinidiyaarachchi. An analytical study of cognitive codelevel object-oriented complexity measures. International Journal of Computer Applications, 183(??):8–14, December 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number45/32233-
2021921851/.

Dutta:2014:AMI

[DKS14] Dushyanta Dutta, Arindam Karmakar, and Dilip Kr. Saikia. An analytical model for IEEE 802.15.4/ZigBee wireless sensor networks with duty cycle mechanism for performance prediction and configuration of MAC parameters to achieve QoS and energy efficiency. International Journal of Computer Applications, 102(??):1–9, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume102/number5/17808-8629/.

Delimi:2013:RIP

[DL13] Rima Delimi and M. H. Latreche. Resolution of inverse problems for electromagnetic levitated and guided systems using the finite element method and the genetic algorithms. International Journal of Computer Applications, 76 (??):30–34, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume76/
number7/13261-0740/.

Dahwah:2019:NWB

[DL19] M. S. Dahwah and A. N. Leukhin. A new wavelet-based algorithm for R-peak detection in ECG and it’s a comparison with the currently existing algorithms. International Journal of Computer Applications, 181(??):22–25, March 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume181/number46/30430-
2019918604/.

Dai:2021:DLF

[DL21] Jingyuan Dai and Shahram Latifi. A deep learning framework for prediction of the mechanism of action. International Journal of Computer Applications, 183(??):1–7, June 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number12/31976-
2021921383/.

Do:2023:ITS

[DL23] An Do and Wei-Ming Lin. Intelligent thread-specific rename register allocation for simultaneous multi-threading processors based on cache behavior. International Journal of Computer Applications, 185(??):1–9, August 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number29/32873-2023923037/.

Damak:2012:SHA

[DLBM12] Taheni Damak, Hassen Loukil, Ahmed Ben Attitala, and Nouri Masmoudi. Software and hardware architecture of H.264/AVC decoder. International Journal of Computer Applications, 59(??):20–27, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume59/number19/9800-4366/.

Dimri:2012:RSJ

[DLN12] Priti Dimri, Shashank Lingwal, and Ashish Negi. A relative superior Julia set and relative superior tricorn and multicorns of fractals. International Journal of Computer Applications, 43(??):29–36, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number6/6109-8327/.

DeGuzman:2019:RHS

[DLR19] Rachell V. De Guzman, Clarissa Joy M. Lubrido, and Claire N. Reyes. Remote home security system using open-source electronic platform. International Journal of Computer Applications, 178(??):5–8, June 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number22/30664-2019918941/.

Deepamalar:2010:EPV

[DM10] M. Deepamalar and M. Madheswaran. An enhanced palm vein recognition system using multi-level fusion of multimodal features and adaptive resonance theory. International Journal of Computer Applications, 1(??):95–101, February 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume1/number20/414-612/.

Dutta:2014:EAT

[DM14] Sourjya Dutta and Iti Saha Misra. Error analysis of 2tier M-ary star QAM modulation in shadowed fading channels. International Journal of Computer Applications, 88 (??):9–16, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume88/
number1/15314-3629/.

Doshi:2017:WPM

[DM17] Poonam P. Doshi and Emmanuel M. Web pattern mining using ECLAT. International Journal of Computer Applications, 179(??):9–14, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number8/28755-2017916009/.

Devi:2018:MDF

[DM18] M. S. Vimala Devi and B. Maheswari. Minimal dominating functions of corona product graph of a path with a cycle. International Journal of Computer Applications, 179(??):9–15, February 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number19/28975-2018916323/.

De:2022:SEN

[DM22] Soumitra De and Jaydev Mishra. Satisfying essential normalization properties of neutrosophic decomposed relations. International Journal of Computer Applications, 184(??):18–22, September 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number29/32500-2022922365/.

DeSouza:2021:MVA

[DMB21] Jerri B. De Souza, Miréia Florêncio Maio, and Peterson A. Belan. A machine vision for automatic hydrometers calibration. International Journal of Computer Applications, 183(??):14–18, November 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number35/32155-2021921729/.

Daham:2014:MHS

[DMM14a] Bnar Faisal A. Daham, Mohammed N. Mohammed, and Kanar Sh. Mohammed. Modified harmony search algorithm for solving the four-color mapping problem. International Journal of Computer Applications, 91(??):34– 38, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume91/
number6/15888-5085/.

Dey:2014:MWI

[DMM14b] Jayanta Kumar Dey, Shyamal Kumar Mandal, and Manoranjan Maiti. Multi-warehouses inventory problem of deteriorating items with fuzzy lead-time and partial lost sales under inflation and time value of money. International Journal of Computer Applications, 92(??):11–19, April 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume92/number13/16067-4937/.

Das:2013:TAB

[DMS13] Umasankar Das, Girija Prasad Mohapatra, and Kunal Shrivastava. Target advertisement based on cohesive structure in a social network. International Journal of Computer Applications, 77(??):1–6, September 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume77/number2/13363-0957/.

Dash:2012:CMD

[DN12] Tirtharaj Dash and Tanistha Nayak. Chain multiplication of dense matrices: Proposing a shared memory based parallel algorithm. International Journal of Computer Applications, 58(??):11–16, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume58/number1/9245-3408/.

De:2015:MEC

[DNP15] Nilanjan De, Sk. Md. Abu Nayeem, and Anita Pal. Modified eccentric connectivity index and polynomial of corona product of graphs. International Journal of Computer Applications, 132(??):1–5, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number9/23619-2015907536/.

Duggani:2015:VIR

[DNR15] Lingaraj Duggani, Udaykumar L. Naik, and Vijay Rayar. VHDL implementation of robust WI-FI positioning system. International Journal of Computer Applications, 121(??):32–37, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number6/21545-4560/.

DosSantosDourado:2019:TIS

[DOC^{+}19] Gabriel Dos Santos Dourado, Josef Augusto Oberdan Souza Silva, Ariane Ruben Calaça Di Menezes, Kárita Cecília Rodrigues Silva, Juliana Santos De Souza Hannum, and Talles Marcelo Gonçalves De Andrade Barbosa. A tangible interface system for people with Down syndrome. International Journal of Computer Applications, 178(??): 16–21, September 2019. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume178/number45/30850-2019919360/.

Dharwa:2011:DMH

[DP11] Jyotindra N. Dharwa and Ashok R. Patel. A data mining with hybrid approach based transaction risk score generation model (TRSGM) for fraud detection of online financial transaction. International Journal of Computer Applications, 16(??):18–25, February 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume16/number1/1977-2651/.

Das:2012:IIR

[DP12] Abhijit Das and Ranjan Parekh. Improved iris recognition in 2D eigen space. International Journal of Computer Applications, 52(??):1–6, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume52/number19/8314-1247/.

Das:2015:IBC

[DP15a] Ajit Kumar Das and Sudarsan Padhy. An improved binary classification framework for investment class rating. International Journal of Computer Applications, 119(??):33–40, June 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume119/number21/
21363-4384/.

De:2015:IEI

[DP15b] Shubhankar De and Ranjan Parekh. Illumination and expression invariant automatic human face recognition using wavelet, eigen and Fisher analysis. International Journal of Computer Applications, 117(??):28–35, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume117/number21/20681-3528/.

Datta:2016:IFF

[DP16] Suchana Datta and Chandan Pan. An intelligent forensic framework towards cloud: Its ontological aspects. International Journal of Computer Applications, 138(??):1–8, March 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume138/number9/24404-2016908934/

.

Devi:2017:DDR

[DP17] S. Nirmala Devi and A. Pethalakshmi. DRR: Dynamic resource reservation in grid computing. International Journal of Computer Applications, 159(??):27–32, February 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume159/number6/27008-2017912967/

.

Duodu:2014:DAM

[DPB14] Quashie Duodu, Joseph Kobina Panford, and James Ben Hayfron-acquah. Designing algorithm for malaria diagnosis using fuzzy logic for treatment (AMDFLT) in Ghana. International Journal of Computer Applications, 91(??): 22–28, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume91/
number17/16102-5353/.

Deshpande:2012:DDE

[DPJB12] Asavari Deshpande, Anup Patil, Saurabh Joshi, and Suraj Bothara. DBCrypto: a database encryption system using query level approach. International Journal of Computer Applications, 45(??):27–32, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number8/6801-9137/.

Daud:2021:EDC

[DPS21] Julianus Gesuri Daud, Muchdar Dg. Patabo, and Josephine Sundah. Early detection coronavirus transmission on object surfaces using kite bones method. International Journal of Computer Applications, 174(??):35–41, January 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume174/number14/31748-
2021921037/.

Dehghan:2014:NFI

[DPT14] Neda Dehghan, Pedram Payvandy, and Mohamad Ali Tavanaei. Nano fiber images thresholding based on imperial competitive algorithm. International Journal of Computer Applications, 99(??):37–41, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume99/number6/17380-7916/.

Deshpande:2012:IVS

[DR12] Meghana M. Deshpande and Jaideep G. Rana. Intelligent video surveillance system based on wavelet transform and support vector machine. International Journal of Computer Applications, 48(??):42–45, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume48/number14/7419-0453/.

Dalal:2014:HEW

[DR14] Sandeep Dalal and Poonam Rani. Histogram equilization on wavelet based compression techniques. International Journal of Computer Applications, 98(??):1–5, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume98/number14/17248-7593/.

Darwish:2015:MDS

[DR15a] Nagy Ramadan Darwish and Nancy M. Rizk. Multidimensional success factors of agile software development projects. International Journal of Computer Applications, 118(??):23–30, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number15/20823-3453/.

Dey:2015:MOS

[DR15b] Samir Dey and Tapan Kumar Roy. Multi-objective structural design optimization using fuzzy optimization programming based on T -Norm. International Journal of Computer Applications, 117(??):20–26, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number12/20607-3229/.

Dutta:2024:SBG

[DRBR24] Sandip Dutta, Soumen Roy, Purba Banerjee, and Utpal Roy. Smartphone-based gait authentication for health insurance incentives: a data-driven approach to verify walking compliance. International Journal of Computer Applications, 186(??):1–13, November 2024. CODEN ????

ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume186/number46/smartphone-based-gait-
authentication-for-health-insurance-incentives-a-
data-driven-approach-to-verify-walking-compliance/

.

DeOliveira:2015:CTG

[DRFF15] Ricardo T. A. De Oliveira, Filippo César G. Régis, Paulo Renato A. Firmino, and Tiago A. E. Ferreira. Course timetabling via genetic algorithms: a real case. International Journal of Computer Applications, 131(??):1–5, December 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume131/number10/23482-
2015907400/.

Devi:2014:SMD

[DRJ14] M. Uma Devi, C. Raghvendra Rao, and M. Jayaram. Statistical measures for differentiation of photocopy from print technology forensic perspective. International Journal of Computer Applications, 105(??):1–7, November 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume105/number15/18450-9792/.

Deepa:2013:ATS

[DRS13] K. Deepa, R. Rangarajan, and M. Senthamil Selvi. Automatic threshold selection using PSO for GA based duplicate record detection. International Journal of Computer Applications, 62(??):22–27, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume62/number4/10068-4674/.

Devi:2015:GNB

[DRS15] G. S. K. Gayatri Devi, G. S. N. Raju, and P. V. Sridevi. Generation of narrow beams using differential evolution algorithm from circular arrays. International Journal of Computer Applications, 112(??):20–27, February 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume112/number3/19646-1235/.

Dananjayan:2010:PAT

[DS10] P. Dananjayan and P. Samundiswary. Performance analysis of trust based AODV for wireless sensor networks. International Journal of Computer Applications, 4(??):6–13, August 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume4/number12/879-1250/

.

Dev:2012:MBA

[DS12a] Harsh Dev and Amit Seth. MDA based approach towards design of database for banking system. International Journal of Computer Applications, 49(??):32–37, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume49/number16/7713-1129/.

Dhara:2012:ACB

[DS12b] Mousumi Dhara and K. K. Shukla. Advanced cost based graph clustering algorithm for random geometric graphs. International Journal of Computer Applications, 60(??):20– 34, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number4/9681-4111/.

Doreswamy:2012:NDS

[DS12c] Doreswamy and Hemanth K S. A novel design specification distance (DSD) based K-mean clustering performance evaluation on engineering materials’ database. International Journal of Computer Applications, 55(??):26– 33, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume55/
number15/8832-3043/.

Dixit:2013:WSS

[DS13] Saurabh Kumar Dixit and Divya Sachan. A Web service selection model using cognitive parameters. International Journal of Computer Applications, 73(??): 16–20, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number21/13017-0054/.

Dhawan:2015:EDD

[DS15] Bhawna Dhawan and Tanu Preet Singh. Efficient data dissemination techniques in VANETs: a review. International Journal of Computer Applications, 116(??):22–32, April 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume116/number7/20349-2535/

.

Deshwal:2019:BCD

[DS19] Vishal Deshwal and Mukta Sharma. Breast cancer detection using SVM classifier with grid search technique. International Journal of Computer Applications, 178(??):18–23, July 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number31/
30735-2019919157/.

Dawood:2011:EEW

[DSA11] M. Sheik Dawood, S. Sadasivam, and G. Athisha. Energy efficient wireless sensor networks based on QoS enhanced base station controlled dynamic clustering protocol. International Journal of Computer Applications, 13 (??):44–49, January 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume13/
number4/1765-2420/.

Das:2012:RBD

[DSB12] Debasish Das, Utpal Sharma, and D. K. Bhattacharyya. Rule based detection of SQL injection attack. International Journal of Computer Applications, 43(??):15– 24, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume43/
number19/6210-8812/.

Dawande:2015:SAE

[DSD15] Jyotsna Rao Dawande, Sanjay Silakari, and Anjna Deen. A survey of all existing clustering protocols in VANETS but main emphasis of survey laid on currently using protocol i.e TCDGP. International Journal of Computer Applications, 118(??):22–31, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number6/20751-3146/.

Datta:2012:SBA

[DSDD12] N. S. Datta, R. Sarker, H. S. Dutta, and M. De. Software based automated early detection of diabetic retinopathy on non dilated retinal image through mathematical morphological process. International Journal of Computer Applications, 60(??):20–24, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume60/number18/9793-4395/.

Safwat:2014:EAD

[dSEdZ14] Nour El din Safwat, M. A. El-dakroury, and Abdelhalim Zekry. The evolution of aircraft data networks. International Journal of Computer Applications, 94(??): 27–32, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number11/16389-5968/.

Das:2011:HEM

[DSJ11a] S. Das, B. C. Sarkar, and R. N. Jana. Hall effects on MHD Couette flow in rotating system. International Journal of Computer Applications, 35(??):22– 30, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume35/
number13/4561-6464/.

Dawoud:2011:FTM

[DSJ11b] Nadir Nourain Dawoud, Brahim Belhaouari Samir, and Josefina Janier. Fast template matching method based optimized sum of absolute difference algorithm for face localization. International Journal of Computer Applications, 18 (??):30–34, March 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume18/
number8/2302-2912/.

Dhandra:2014:IIL

[DSM14] B. V. Dhandra, Shridevi Soma, and Gururaj Mukarambi. Identification of institutional logo based on wavelet features. International Journal of Computer Applications, 107(??):9–13, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number15/18826-0255/.

DeLira:2020:MNR

[DSM^{+}20] Maria Camila S. De Lira, Ademir B. Santos Neto, Maria C. Moraes Batista, Roberta Macedo M. Gouveia, and Tiago Alessandro E. Ferreira. Multidimensional and non-relational data models: a comparison with a big volume of data. International Journal of Computer Applications, 175(??):1–7, December 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume175/number36/
31682-2020920927/.

DaSilva:2021:CTI

[DSP^{+}21] Rodes Angelo B. Da Silva, Borko Stosic, Héliton Pandorfi, Gledson Luiz P. De Almeida, Marcos Vinícius Da Silva, and Pedro Henrique Dias Batista. Classification of thermal images of bovine mastitis by computer vision. International Journal of Computer Applications, 174(??):41–45, April 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume174/number31/
31880-2021921256/.

Dubey:2013:LBR

[DSS13] Madhukar Dubey, Piyush Singh, and Gaurav Shrivastava. Location base routing with energy awareness scheme under Wi-Max in MANET for minimization of routing overhead. International Journal of Computer Applications, 64 (??):29–34, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number3/10615-5333/.

DeAraujoArlindo:2018:IAU

[DSS18] Jardeson De Araujo Arlindo, Renata Imaculada Soares Pereira, and Sandro César Silveira Jucá. IoT arcade using Linux embedded system. International Journal of Computer Applications, 181(??):35–38, August 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number12/29828-2018917710/.

Dahi:2014:QGA

[DT14] Soumaya Dahi and Sami Tabbane. Quadratic greedy algorithm for multichannel access in cognitive radio network. International Journal of Computer Applications, 91 (??):28–33, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume91/
number15/15958-5331/.

Deffo:2022:FFR

[DT22] Lionel Landry Sop Deffo and Elie Fute Tagne. FaRes: a face recognition system based on motion detection and image super resolution. International Journal of Computer Applications, 184(??):28–33, March 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number4/32321-2022921997/.

Dung:2023:PSA

[DTO23] Solomon Tok Dung, Godwin Thomas, and Yinka Oyerinde. Predicting students’ academic performance using artificial neural networks. International Journal of Computer Applications, 185(??):1–7, June 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number19/32801-2023922889/.

Duhan:2016:PPF

[DTS16] Neelam Duhan, Parul Tomar, and Amit Siwach. PFIMII: Parallel frequent itemset mining using interval intersection. International Journal of Computer Applications, 156(??):10–15, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume156/number13/26768-2016912586/.

Dumebi:2020:CRU

[Dum20] Okpor Margaret Dumebi. Character recognition using neural network self-organizing map (NN-SOM). International Journal of Computer Applications, 175(??):20–24, October 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume175/number23/31591-
2020920761/.

Dewi:2013:IPS

[DUP13] Athanasia O. P. Dewi, Wiranto H. Utomo, and Sri Yulianto J P. Identification of potential student academic ability using comparison algorithm K-means and farthest first. International Journal of Computer Applications, 63 (??):18–26, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number17/10558-5631/.

Dureha:2014:AAG

[Dur14] Anukriti Dureha. An accurate algorithm for generating a music playlist based on facial expressions. International Journal of Computer Applications, 100(??): 33–39, August 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume100/number9/17557-8163/.

Dutta:2014:VFI

[DV14] Ritam Dutta and Charles Van Der Mast. Virtex 4 FPGA implementation of Viterbi decoded 64-bit RISC for high speed application using Xilinx. International Journal of Computer Applications, 88(??):30–35, February 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume88/number14/15422-4011/.

Daund:2016:WBA

[DV16] Aditi Daund and Pankaj Kumar Vyas. Wireless broadband access with the application of IEEE 802.11b based Wi-Fi model. International Journal of Computer Applications, 154(??):39–44, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume154/number5/26491-2016912143/.

Devi:2013:SDE

[DVS13] Meenu Devi, S. R. Verma, and M. P. Singh. Solving differential equations of second order using quadratic Legendre multi-wavelets (QLMW) with operational matrix of integration. International Journal of Computer Applications, 75 (??):43–49, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number15/13190-0912/.

DaSOliveira:2023:APR

[DXT23a] Rafael Da S.Oliveira, José A. M. Xexéo, and Renato H. Torres. An analysis of the performance and randomness of lattice-based public key cryptography algorithms. International Journal of Computer Applications, 185(??):11–17, November 2023. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume185/number42/
32967-2023923227/.

DosSRocha:2023:AIA

[DXT23b] Bruno Dos S.Rocha, José A. M. Xexéo, and Renato H. Torres. Artificial intelligence applied to the identification of block ciphers under CBC mode. International Journal of Computer Applications, 185(??):1–8, September 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number34/32907-2023923114/.

Esmaeili:2011:SDM

[EA11] Mohammadjafar Esmaeili and Arwa Almadan. Stream data mining and anomaly detection. International Journal of Computer Applications, 34(??):39–42, November 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume34/number9/4130-
5984/.

ElAlfi:2018:IAS

[EA18] A. E. E. El Alfi and S. M. EL Atawy. Intelligent Arabic Sign Language to Arabic text translation for easy deaf communication. International Journal of Computer Applications, 180(??):19–26, May 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number41/29403-2018917081/.

Elkailani:2014:TPO

[EAa14] Omima Elkailani and Mohmoud Al-aujali. Transmission performance of optical add/drop multiplexers in metro optical access communication networks. International Journal of Computer Applications, 96(??):5–8, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume96/number8/16812-6553/.

Ekhi:2013:CES

[EAB13] Ziya Ekhi, Akif Akgul, and Mehmet Recep Bozkurt. The classificaton of EEG signals recorded in drunk and nondrunk people. International Journal of Computer Applications, 68(??):40–44, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number10/11619-7018/.

El-Arishy:2017:MIN

[EADEA17] S. M. El-Arishy, L. S. Diab, and E. S. El-Atfy. Moments inequalities for NBRUrp distributions with hypotheses testing applications. International Journal of Computer Applications, 159(??):34–40, February 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume159/number2/26976-2017912879/.

ElHaji:2014:UKM

[EAE14] Essaid El Haji, Abdellah Azmani, and Mohamed El Harzli. Using KNN method for educational and vocational guidance. International Journal of Computer Applications, 100(??):24–30, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number12/17578-8335/.

ElAlami:2019:BSB

[EAE19] M. E. ElAlami, Hosnia M. M. Ahmed, and Mahy E. Elemam. Building a system based on intelligent agencies for assisting in classification and circulation of books and periodicals automatically. International Journal of Computer Applications, 178(??):22–30, May 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number9/30559-2019918803/.

Elalami:2012:PIF

[EAEa12] M. E. Elalami, A. E. Amin, and A. E. El-alfi. A personal identification framework based on facial image and fingerprint fusion biometric. International Journal of Computer Applications, 51(??):41–48, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume51/number7/8058-1411/.

ElMaghraby:2013:HFD

[EAEE13] Amr El Maghraby, Mahmoud Abdalla, Othman Enany, and Mohamed Y. El Nahas. Hybrid face detection system using combination of Viola–Jones method and skin detection. International Journal of Computer Applications, 71 (??):15–22, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number6/12361-8681/.

ElMaghraby:2014:DAF

[EAEE14] Amr El Maghraby, Mahmoud Abdalla, Othman Enany, and Mohamed Y. El Nahas. Detect and analyze face parts information using Viola–Jones and geometric approaches. International Journal of Computer Applications, 101(??):23–28, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number3/17667-8494/.

ElMaghraby:2015:DTM

[EAEE15] Amr El Maghraby, Mahmoud Abdalla, Othman Enany, and Mohamed Y. El Nahas. Detecting and tracking of multiple people in video based on hybrid detection and human anatomy body proportion. International Journal of Computer Applications, 109(??):10–14, January 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume109/number17/19416-0656/.

Elshazly:2016:FIR

[EAF^{+}16] E. A. Elshazly, Safey A. S. Abdelwahab, R. M. Fikry, S. M. Elaraby, O. Zahran, and M. El-Kordy. FPGA implementation of robust image steganography technique based on least significant bit (LSB) in spatial domain. International Journal of Computer Applications, 145(??):43–52, July 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume145/number12/25334-
2016910796/.

E:2015:MBA

[EAJ^{+}15] Manjunathswamy B E, Appaji M. Abhishek, Thriveni J, Venugopal K R, and L. M. Patnaik. Multimodal biometric authentication using ECG and fingerprint. International Journal of Computer Applications, 111(??):33–39, February 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume111/number13/19601-
1452/.

Elfadil:2014:SCV

[EAM14] Nazar Elfadil, Sultan Aljahdali, and Mohammed Moawad. System for communicating a vehicle position and speed during accident. International Journal of Computer Applications, 94(??):13–18, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume94/number19/16498-5585/.

El-Assy:2016:IBD

[EAMAES16] A. M. El-Assy, M. A. Mohamed, and M. E. A. Abou-El-Seoud. Improved BEMD-DWT-DCT-SVD robust watermarking technique for still images. International Journal of Computer Applications, 150(??):13–20, September 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume150/number9/26120-2016911620/

.

Elhanaoui:2014:GWP

[EAMD14] Abdelkader Elhanaoui, Elhoucein Aassif, Gerard Maze, and Dominique Decultot. Guided waves propagation in twolayer tubes: Effects of the inner part in contact with air in the cavity. International Journal of Computer Applications, 102(??):34–43, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume102/number2/17790-8581/.

Emam:2015:MIS

[EAO15] Marwa M. Emam, Abdelmgeid A. Aly, and Fatma A. Omara. A modified image steganography method based on LSB technique. International Journal of Computer Applications, 125(??):12–17, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number5/22427-2015905908/.

E:2016:MDP

[EAS16] Yoro R E, Adekunle Y A, and Ebiesuwa Seun. Modeling diffusion problems via graph-based structure. International Journal of Computer Applications, 139(??):29–35, April 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume139/number11/
24536-2016909434/.

Eltoum:2013:VBT

[EB13] Ibrahim Abdallah Hag Eltoum and Mohammed Bouhorma. Velocity based tracking and localization system using smartphones with GPS and GPRS/3G. International Journal of Computer Applications, 76(??):47–51, August 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume76/number11/13295-0875/.

Elyaqouti:2021:CMS

[EB21] Mustapha Elyaqouti and Lahoussine Bouhouch. Characterization and modeling of supercapacitors used for hybrid electric energy storage. International Journal of Computer Applications, 183(??):9–15, September 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number26/32090-2021921639/.

Ebad:2013:CSQ

[Eba13] Shouki A. Ebad. A comparative study of query-set size and fixed-data perturbation as two techniques to secure statistical databases. International Journal of Computer Applications, 68(??):7–10, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number4/11565-6856/.

Eghnam:2011:SCT

[EBAS11] Karam M. Eghnam, Sulieman Bani-Ahmad, and Alaa F. Sheta. Solving capelin time series ecosystem problem using hybrid artificial neural networks- genetic algorithms model. International Journal of Computer Applications, 19 (??):8–12, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume19/
number2/2336-3046/.

El-bassiouny:2015:ELD

[EBAS15] A. H. El-Bassiouny, N. F. Abdo, and H. S. Shahen. Exponential Lomax distribution. International Journal of Computer Applications, 121(??):24–29, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number13/21602-4713/.

El-Bassiouny:2017:NGE

[EBAS17] A. H. El-Bassiouny, Mohamed Abouhawwash, and H. S. Shahen. New generalized extreme value distribution and its bivariate extension. International Journal of Computer Applications, 173(??):1–10, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume173/number3/28312-2017915276/.

ElBitar:2017:CBM

[EBB17] Ibrahim El Bitar, Zouhair Bazzal, and Fatima-Zahra Belouadha. Case-based matching algorithm for dynamic Web service discovery. International Journal of Computer Applications, 165(??):10–16, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number3/27551-2017913805/.

Eltoum:2013:TMO

[EBO13] Ibrahim Abdallah Hag Eltoum, Mohammed Bouhorma, and Tarik Ait Otman. Tracking a moving object according to the angle and the direction via a smartphones over GPS plus Wi-Fi coverage with GPRS/3G. International Journal of Computer Applications, 81(??):37–41, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume81/number17/14219-2480/.

ElBitar:2012:LAA

[EBR12] Ibrahim El Bitar, Fatima-Zahra Belouadha, and Ounsa Roudiës. A logic and adaptive approach for efficient diagnosis systems using CBR. International Journal of Computer Applications, 39(??):1–5, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number15/4893-7393/.

El-dosuky:2018:OPF

[ED18] M. A. El-Dosuky. Optimal power flow using cockroach swarm optimization. International Journal of Computer Applications, 179(??):1–6, April 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number36/29271-2018916297/.

Emmenuel:2019:CBS

[ED19] Ekwonwune Emmenuel and Michael O. Daniel. A cloudbased student’ project supervision and allocation system for Nigeria tertiary institutions. International Journal of Computer Applications, 178(??):27–35, June 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number19/30643-2019918963/.

Eva:2019:HBP

[EDD19] Onokpasa Eva, Bisandu Desmond, and Bakwa Dunka. A hybrid backtracking and pencil and paper sudoku solver. International Journal of Computer Applications, 181(??):39–43, April 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume181/number47/
30472-2019918642/.

El-Damcese:2015:SPE

[EDR15] M. A. El-Damcese and Dina A. Ramadan. Studies on properties and estimation problems for modified extension of exponential distribution. International Journal of Computer Applications, 125(??):21–28, September 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume125/number4/22421-2015905891/.

Elsayad:2013:DBC

[EE13a] Alaa M. Elsayad and H. A. Elsalamony. Diagnosis of breast cancer using decision tree models and SVM. International Journal of Computer Applications, 83(??):19– 29, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number5/14445-2604/.

Emmanuel:2013:CAM

[EE13b] Affum Emmanuel and Addo Emmanuel. On comparison analysis of multiple interference MIMO systems with uncorrelated Rician fading channels. International Journal of Computer Applications, 76(??):1–5, August 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume76/number17/13336-0291/.

Eludiora:2016:DVG

[EE16a] Safiriyu Eludiora and Gabriel Elufidodo. Development of a verb group machine translation system. International Journal of Computer Applications, 153(??):39– 52, November 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume153/number4/26393-2016912005/.

Emmanuel:2016:UCI

[EE16b] Eronu Majiyebo Emmanuel and Ibrahim Tashiwa Emmanuel. Use of contextual information to optimize wireless sensor node’s energy consumption rate during reprogramming procedures. International Journal of Computer Applications, 143(??):28–39, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume143/number11/25123-2016910473/.

Elalfi:2019:AES

[EEA19] A. E. E. Elalfi, A. F. Elgamal, and N. A. Amasha. Automated essay scoring using Word2vec and support vector machine. International Journal of Computer Applications, 177(??):20–29, December 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number25/31053-2019919707/.

Elglaad:2017:PMI

[EEAE17] Mohamed Elglaad, A. A. Ewees, and E. E. Abd-Elrazek. A prediction method to improve training management. International Journal of Computer Applications, 166(??):39–43, May 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume166/number4/27661-2017914033/

.

Elabbasy:2016:ABS

[EEB16] E. M. Elabbasy and A. A. El-Biaty. Asymptotic behavior of some rational difference equations. International Journal of Computer Applications, 136(??):18–24, February 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume136/number8/24174-2016908208/

.

ElAlfi:2014:IAT

[EEE14] A. E. E. El Alfi, M. M. R. El Basuony, and S. M. El Atawy. Intelligent Arabic text to Arabic sign language translation for easy deaf communication. International Journal of Computer Applications, 92(??):22–29, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume92/number8/16030-5126/.

Elsharkawi:2016:CES

[EEEE16] Abdelfattah Elsharkawi, Kamal A. ElDahshan, Eman K. Elsayed, and Mahmoud Eltaher. Contrast enhancement satellite images: a hybrid solution for cloud removal. International Journal of Computer Applications, 138(??):42–49, March 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume138/number8/24403-2016909053/.

Elganzoury:2015:ESC

[EEGHS15] Hisham S. Elganzoury, Talaat A. El-Garf, A. A. Hafez, and Ahmed Safwat. Enhanced stream cipher algorithm using consecutive nonlinear functions. International Journal of Computer Applications, 123(??):33–37, August 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume123/number10/21997-2015905546/.

Eisa:2014:ISC

[EEH14] Mohamed Eisa, E. I. Esedimy, and Alaa Halawani. Improving scalability of cloud system based on routing algorithms. International Journal of Computer Applications, 101(??):40–46, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number11/17734-8847/.

ElDahshan:2017:ASR

[EEKG17] Kamal ElDahshan, Afaf Abd El-Kader, and Nermeen Ghazy. Achieving stability in the round robin algorithm. International Journal of Computer Applications, 172(??):15–20, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume172/number6/28255-2017915161/.

ElMelhaoui:2011:ANR

[EEL11] Ouafae El Melhaoui, Mohamed El Hitmy, and Fairouz Lekhal. Arabic numerals recognition based on an improved version of the loci characteristic. International Journal of Computer Applications, 24(??):36–41, June 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume24/number1/2912-3830/.

Ezzat:2012:FCD

[EEM12] Eman M. Ezzat, Doaa S. Elzanfaly, and Mostafa A. Mostafa. Fly over the clouds or drive through the crowd: a cloud adoption framework in action. International Journal of Computer Applications, 56(??):1–8, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume56/number4/8876-2856/.

Eisa:2014:ECC

[EER14] Mohamed Eisa, E. I. Esedimy, and M. Z. Rashad. Enhancing cloud computing scheduling based on queuing models. International Journal of Computer Applications, 85 (??):17–23, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number2/14813-3032/.

Elaskary:2015:OSS

[EF15] Ramez Moh. Elaskary and Yasmine Fahmy. Optimizing spectrum sensing for cognitive radio networks to limit interference on primary user. International Journal of Computer Applications, 123(??):1–6, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number12/22008-2015903658/.

Elalfi:2016:DID

[EFA16] A. E. E. Elalfi, M. A-H. Fouda, and A. A. Atta. Developing an intelligent decision support system for the diagnosis of some children’s diseases. International Journal of Computer Applications, 151(??):32–38, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume151/number2/26208-2016911688/.

Fattah:2013:TBM

[eFKS13] I. M. Abd el Fattah, W. I. Khedr, and K. M. Sallam. A TOPSIS based method for gene selection for cancer classification. International Journal of Computer Applications, 67(??):39–44, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number17/11490-7195/.

El-gedawy:2013:UFS

[Eg13] Madeeh Nayer El-gedawy. Using fuzzifiers to solve word sense ambiguation in Arabic language. International Journal of Computer Applications, 79(??):1–8, October 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume79/number2/13710-1465/.

Esmaelzadeh:2014:ASS

[EG14] Reza Esmaelzadeh and Hossein Ghadiri. Appropriate starter for solving the Kepler’s equation. International Journal of Computer Applications, 89(??):31–38, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume89/number7/15517-4394/.

Esmaeeli:2015:EEP

[EG15] Mahnaz Esmaeeli and Seyed Ali Hosseini Ghahroudi. An energy-efficiency protocol in wireless sensor networks using theory of games and fuzzy logic. International Journal of Computer Applications, 126(??):8–13, September 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume126/number1/22514-2015905969/

.

ElDahshan:2015:DBP

[EGES15] Kamal A. ElDahshan, Ahmed A. A. Gad-ElRab, and Ahmed Sobhi. A distance-based predicted region policy for cache replacement in mobile environments. International Journal of Computer Applications, 126(??): 1–10, September 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume126/number7/22561-2015906086/.

Ezzat:2014:IAN

[EH14] Akram W. Ezzat and Ikhlas M. Hasan. Investigation of alumina nano fluid thermal conductivity. International Journal of Computer Applications, 102(??):15–23, September 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume102/number11/17858-
8721/.

El-Helloty:2016:FVA

[EH16] Arafa El-Helloty. Free vibration analysis of stiffened laminated composite plates. International Journal of Computer Applications, 156(??):12–23, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume156/number1/26672-2016912340/.

El-Henawy:2018:MHA

[EHA18] Ibrahim El-Henawy and Nagham Ahmed Abdelmegeed. Meta-heuristics algorithms: a survey. International Journal of Computer Applications, 179(??):45–54, February 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume179/number22/29004-
2018916427/.

Elsayed:2017:RTT

[EHHM17] Mohammed Elsayed, Abdallah Hammad, Ashraf Hafez, and Hala Mansour. Real time trajectory tracking controller based on Lyapunov function for mobile robot. International Journal of Computer Applications, 168(??):1–6, June 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume168/number11/27916-2017914540/.

Erfan:2017:OTS

[EHI17] Md. Erfan, Farhana Huq, and Md. Shariful Islam. An optimal task scheduling mechanism for mobile cloud computing. International Journal of Computer Applications, 175(??):1–8, October 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume175/number6/28489-2017915592/.

ElGhor:2013:RTO

[EhMA13] Hussein El Ghor and El hadi M. Aggoune. Real-time operating system for wireless sensors powered by renewable energy source. International Journal of Computer Applications, 81(??):1–7, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume81/number12/14061-1961/.

El-hadary:2023:IMA

[EHSA23] Sally Zaki El-Hadary, Sara A. Shehab, and Hatem Said Ahmed. Improving MSAProbs algorithm performance and parallel computing using GPU. International Journal of Computer Applications, 185(??):14–18, April 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number1/32668-2023922532/.

Elhassouny:2011:CDF

[EIB^{+}11] Azeddine Elhassouny, Soufiane Idbraim, Aissam Bekkari, Driss Mammass, and Danielle Ducrot. Change detection by fusion/contextual classification based on a hybrid DSmT model and ICM with constraints. International Journal of Computer Applications, 35(??):28–40, December 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume35/number8/4422-
6154/.

Eisa:2013:IGD

[Eis13] Mohamed Eisa. Improving group decision support systems using rough set. International Journal of Computer Applications, 69(??):9–13, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume69/number2/11812-7473/.

Eisa:2014:NAE

[Eis14] Mohamed Eisa. A new approach for enhancing image retrieval using neutrosophic sets. International Journal of Computer Applications, 95(??):12–20, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume95/number8/16613-6453/.

Ejaz:2012:AGA

[Eja12] Muhammad Ejaz. Application of genetic algorithm for tuning of reduced ordered robust PID controller and computer simulations. International Journal of Computer Applications, 42(??):20–24, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number17/5785-7742/.

Elloumi:2016:IAC

[EJC16] Khouloud Elloumi, Mohamed Jemel, and Mohamed Chtourou. Indirect adaptive control for discrete-time nonlinear systems based on T -S fuzzy model. International Journal of Computer Applications, 143(??):43–49, June 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume143/number9/25109-2016910357/.

ElBorji:2014:ILA

[EK14a] Yassine El Borji and Mohamed Khaldi. An IEEE LOM application profile to describe serious games “SG-LOM”. International Journal of Computer Applications, 86(??): 1–8, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number13/15042-3404/.

ElBouanani:2014:AAS

[EK14b] Sara El Manar El Bouanani and Ismail Kassou. Authorship analysis studies: a survey. International Journal of Computer Applications, 86(??):22–29, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume86/number12/15038-3384/.

ElGhadouali:2013:ICS

[EKA13] Asmae El Ghadouali, Oulaid Kamach, and Benaissa Amami. Identification of compatible states in switching mode. International Journal of Computer Applications, 72(??): 42–45, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number20/12661-9447/.

EL-Kader:2020:DNF

[EKES20] Hatem Mohamed Abd El-Kader and Amr Ibrahim Awd El-Shora. Developing a neuro-fuzzy model for weather prediction. International Journal of Computer Applications, 177(??):15–26, January 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number34/31121-2020919818/.

Ejegwa:2016:AIF

[EKI16] P. A. Ejegwa, L. N. Kwarkar, and K. N. Ihuoma. Application of intuitionistic fuzzy multisets in appointment process. International Journal of Computer Applications, 135(??):1–4, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number1/24010-2016908167/.

ElHouby:2013:AAC

[El 13] Enas M. F. El Houby. Analysis of associative classification for prediction of HCV response to treatment. International Journal of Computer Applications, 63(??):38– 44, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number15/10545-5542/.

ElNahal:2014:MMB

[El 14] Waleed El Nahal. Mobile multimodal biometric system for security. International Journal of Computer Applications, 89(??):12–16, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume89/number10/15666-4041/.

ElHouby:2015:WGO

[El 15] Enas M. F. El Houby. World geographical ontology model. International Journal of Computer Applications, 120(??):25– 33, June 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume120/number15/
21305-4081/.

ElAallaoui:2020:MIF

[El 20] Mohamed El Aallaoui. Medical image fusion algorithm based on adaptive selectivity reconstruction and pulse coupledneural network. International Journal of Computer Applications, 175(??):1–9, September 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number19/31557-2020920721/.

Elazhary:2011:FBT

[Ela11] Hanan Elazhary. A fast, blind, transparent, and robust image watermarking algorithm with extended torus automorphism permutation. International Journal of Computer Applications, 32(??):34–41, October 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume32/number4/3895-5458/.

Elalami:2012:DAR

[Ela12] M. E. Elalami. Destructive algorithm for rule extraction based on a trained neural network. International Journal of Computer Applications, 42(??):8–14, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume42/number21/5833-8119/.

Elci:2011:TCP

[Elc11] Atilla Elci. Text classification by PNN-based term reweighting. International Journal of Computer Applications, 29(??):7–13, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume29/
number12/3701-5188/.

Elouaham:2013:PEE

[ELDM13] S. Elouaham, R. Latif, A. Dliou, and F. M. R. Maoulainine. Periodogram and ensemble empirical mode decomposition analysis of electromyography processing. International Journal of Computer Applications, 65(??):33– 40, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number2/10898-5822/.

Elgamal:2013:EWB

[Elg13] A. F. Elgamal. Educational website based on integration between Web-based simulation and retrieval strategies. International Journal of Computer Applications, 73 (??):41–46, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number15/12821-0302/.

Elmaghraby:2022:ERI

[Elm22] Amr Elmaghraby. Extract rich information from images and video using custom vision cognitive services. International Journal of Computer Applications, 184(??):15–28, June 2022. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume184/number16/
32402-2022922152/.

Elnashar:2014:DNN

[Eln14] Gamal A. Elnashar. Dynamical nonlinear neural networks with perturbations modeling and global robust stability analysis. International Journal of Computer Applications, 85 (??):14–21, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number15/14917-3479/.

ElRakaiby:2016:GIE

[ElR16] Mohamed ElRakaiby. Grayscale image encryption using cellular automata. International Journal of Computer Applications, 142(??):1–3, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number1/24857-2016909655/.

El-laithy:2013:SPL

[ElZA13] Dina M. El-laithy, Abdelhalim Zekry, and Mohamed Abouelatta. Speeding-up phase-locked loops based on adaptive loop bandwidth. International Journal of Computer Applications, 61(??):1–6, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume61/number3/9905-4494/.

Elsamadony:2023:EHR

[EM23] Ola A. Elsamadony and Ahmed Mattar. Estimation the hazard rate function using Gumbel type 2 kernel. International Journal of Computer Applications, 185 (??):12–15, May 2023. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume185/number6/32704-2023922709/.

Elkamchouchi:2011:EME

[EMA11] Hassan Elkamchouchi, Darwish Mohamed, and Wael Ali. The effect of missing elements on the performance of D3LS STAP approach using real antenna elements. International Journal of Computer Applications, 28(??):13– 18, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume28/
number11/3429-4624/.

El-mongy:2013:DPS

[EmdHNmAW13] Ahmed Mohamed Abd El-mongy, Alaa El deen Hamouda, Nihal Nounou, and Abdel moneim A. Wahdan. Design of prediction system for key performance indicators in balanced scorecard. International Journal of Computer Applications, 72(??):10–14, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume72/number8/12512-6016/.

ElGamal:2013:FVW

[EME13] A. F. El Gamal, N. A. Mosa, and W. K. El Said. A fragile video watermarking algorithm for content authentication based on block mean and modulation factor. International Journal of Computer Applications, 80(??):21– 28, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume80/
number4/13850-1684/.

ElKafrawy:2015:ECM

[EME15] Passent El Kafrawy, Amr Mausad, and Heba Esmail. Experimental comparison of methods for multi-label classification in different application domains. International Journal of Computer Applications, 114(??):1–9, March 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume114/number19/20083-1666/.

El-matarawy:2013:PDC

[EmErB13] Ali El-matarawy, Mohammad El-ramly, and Reem Bahgat. Parallel and distributed code clone detection using sequential pattern mining. International Journal of Computer Applications, 62(??):25–31, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume62/number10/10118-4792/.

Embaby:2014:DSA

[EMF14] Sherif Abdelazim Embaby, Ahmed Shawky Moussa, and Ibrahim Farag. A dynamic scheduling algorithm for spawn processes in MPI-2 to improve and maintain load balancing. International Journal of Computer Applications, 106(??):1–6, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number12/18569-6055/.

Elgeldawi:2019:CAS

[EMS19] Enas Elgeldawi, Maha Mahrous, and Awny Sayed. A comparative analysis of symmetric algorithms in cloud computing: a survey. International Journal of Computer Applications, 182(??):7–16, April 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number48/30516-2019918726/.

Ezzat:2013:THB

[EN13] Akram W. Ahmed Ezzat and Emad W. Nayyef. Thermohydraulic behavior of water cooling channel subjected to constant heat flux during pressure reduction transient in its cooling system. International Journal of Computer Applications, 80(??):23–33, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume80/number1/13826-1564/.

Enany:2012:ETP

[Ena12] Mohamed A. Enany. Effects of three phase AC/DC converter drive on the torque-speed characteristic of DC motor. International Journal of Computer Applications, 58(??): 22–26, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number6/9287-3481/.

Eddahibi:2012:TID

[ENC12] M. Eddahibi, A. Nejeoui, and C. Cherkaoui. Towards an intelligent and deeply automated shopping bot. International Journal of Computer Applications, 45(??): 21–27, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume45/
number24/7113-9806/.

ElMhouti:2013:DTQ

[ENE13] Abderrahim El Mhouti, Azeddine Nasseh, and Mohamed Erradi. Development of a tool for quality assessment of digital learning resources. International Journal of Computer Applications, 64(??):27–31, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume64/number14/10703-5613/.

ElMhouti:2014:DIT

[ENE14] Abderrahim El Mhouti, Azeddine Nasseh, and Mohamed Erradi. Design of an intelligent tutorial system based on ontologies. International Journal of Computer Applications, 89 (??):30–34, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number1/15468-4085/.

El-nattat:2015:NTS

[ENEBES15] Amal El-Nattat, Nirmeen A. El-Bahnasawy, and Ayman El-Sayed. A new task scheduling algorithm for maximizing the distributed systems efficiency. International Journal of Computer Applications, 110(??):9–16, January 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume110/number9/19343-0279/.

Essawe-ndedi:2014:NSC

[EnFn14] Priso Essawe-ndedi and Marcel Fouda-ndjodo. On the notion of semantic correctness: a context-based approach. International Journal of Computer Applications, 94(??): 1–6, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number5/16336-5628/.

Enyindah:2022:SOT

[EOO22] P. Enyindah, C. Onukwugha, and E. N. Osegi. Securing oil tank level control system against adversary attacks using fuzzy logic technique. International Journal of Computer Applications, 184(??):30–41, April 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number6/32335-2022922017/.

Ezhilarasi:2012:FIV

[ER12] A. Ezhilarasi and M. Ramaswamy. FPGA implementation of voltage sharing strategy for series connected SEPICS. International Journal of Computer Applications, 39(??):11– 18, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume39/
number14/4887-7375/.

EsSaady:2011:ADA

[EREM11] Youssef Es Saady, Ali Rachidi, Mostafa El Yassa, and Driss Mammass. AMHCD: a database for amazigh handwritten character recognition research. International Journal of Computer Applications, 27(??):44–48, August 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume27/number4/3286-4475/.

El-sherbeny:2012:OSW

[Es12a] Mohamed Salah El-sherbeny. Optimal system for warm standby components in the presence of standby switching failures, two types of failures and general repair time. International Journal of Computer Applications, 44(??): 17–26, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number20/6379-8846/.

Esfandi:2012:SAC

[ES12b] Abolfazl Esfandi and Mehdi Sabbari. Study of access control issue in Web services. International Journal of Computer Applications, 49(??):11–16, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume49/number1/7589-7647/.

E:2013:IVT

[ES13] Shalini E and T. V. P. Sundararajan. Improved video transmission over mobile ad-hoc networks using AOMDV and MDC with path diversity. International Journal of Computer Applications, 62(??):35–42, January 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume62/number17/10176-5087/.

Eken:2015:BDF

[ES15a] Suleyman Eken and Ahmet Sayar. Big data frameworks for efficient range queries to extract interested rectangular sub regions. International Journal of Computer Applications, 119(??):36–39, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number22/21372-4423/.

ElDeenAhmed:2015:MSP

[ES15b] Mohamed Amal El Deen Ahmed and Marwa Mohamed Solayman. Maximizing strategic performance results: Adopting balanced scorecards and BI tools. International Journal of Computer Applications, 117(??):1–11, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume117/number10/20587-2368/.

Eswaran:2015:STF

[ES15c] K. Eswaran and Vishwajeet Singh. Some theorems for feed forward neural networks. International Journal of Computer Applications, 130(??):1–17, November 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume130/number7/23218-2015907021/.

Eltayeb:2018:EED

[ESAA18] Engy N. Eltayeb, Nancy M. Salem, and Walid Al-Atabany. Evaluating the efficiency of different feature sets on brain tumor classification in MR images. International Journal of Computer Applications, 180(??):1–7, May 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number38/29376-2018917008/.

Elyusufi:2013:CHC

[ESB13] Yasyn Elyusufi, Hamid Seghiouer, and Abdelhamid Benkaddour. Customization of human-computer interface guided by ontological approach in Web 2.0. International Journal of Computer Applications, 81(??):1–6, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume81/number5/14005-2070/.

Ebaid:2013:PMR

[ESE13] Nada Ebaid, Mona Shokair, and Samy El Dalil. Performance of multihop relaying system in cognitive radio networks. International Journal of Computer Applications, 81 (??):48–52, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume81/
number3/13996-2025/.

Elzaher:2016:SMV

[ESE16] Mahmoud F. Abd Elzaher, Mohamed Shalaby, and Salwa H. El Ramly. Securing modern voice communication systems using multilevel chaotic approach. International Journal of Computer Applications, 135(??):17–21, February 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume135/number9/24077-2016908497/

.

El-Sokkary:2021:PHA

[ESEMD21] Nesma El-Sokkary, Walaa Hussein El-Masry, and Nagy Ramadan Darwish. A proposed hybrid approach for developing healthcare information systems. International Journal of Computer Applications, 183(??):17–23, September 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number28/32107-
2021921664/.

El-salhy:2014:URP

[EsGK14] Walaa El-salhy, Fayed F. M. Ghaleb, and Ahmed M. Khedr. Using a random perturbation-based scheme for establishing authentic associations in wireless mesh network. International Journal of Computer Applications, 94 (??):18–22, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number20/16551-6198/.

Elhalkouj:2014:EEL

[ESH^{+}14] T. Elhalkouj, M. Sabil, A. Habib, Z. Benkhaldoun, My. Y. Elazhari, M. Lazrek, and O. Azagrouze. European Extremely Large Telescope: Isopistonic angle measurements at Aklim site. International Journal of Computer Applications, 99(??):1–4, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume99/number5/17366-7889/.

El-Sheikh:2020:HWS

[ESKM20] R. A. El-Sheikh, F. Khalifa, and M. A. Mohammed. A hybrid watermarking scheme for increasing watermark capacity and security. International Journal of Computer Applications, 177(??):35–47, February 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number35/31134-2020919861/.

Elgayar:2013:NGR

[ESR13] Sara Elgayar, Mohammed A.-M. Salem, and Mohamed I. Roushdy. Nested graph representation for visual SLAM based on local and global feature processing. International Journal of Computer Applications, 78(??):1–8, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume78/
number16/13604-1350/.

El-sofany:2013:ESQ

[EsTeEh^{+}13] Hosam F. El-sofany, Islam A. Taj-eddin, Hassan El-hoimal, Tourki Al-tourki, and Amjad Al-sadoon. Enhancing software quality by an SPL testing based software testing. International Journal of Computer Applications, 69(??): 5–13, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number6/11844-7574/.

El-Sherbiny:2015:GAS

[ESZED15] Mahmoud M. El-Sherbiny, Ramadan A. Zeineldin, and Abdallah M. El-Dhshan. Genetic algorithm for solving course timetable problems. International Journal of Computer Applications, 124(??):1–7, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number10/22144-2015905408/.

Ezatdoost:2013:ACG

[ETD13] Yasin Ezatdoost, Ali Tourani, and Amir Seyed Danesh. Available challenges and guidelines in the field of deep Web and intensive crawling. International Journal of Computer Applications, 77(??):1–5, September 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume77/number1/13355-0948/.

Emmanuel:2017:LBI

[ETG17] Ajagbonna Babatunde Emmanuel, Abdoulie Momodou Sunkary Tekanyi, and Mohammed Aliyu Gadam. Load balancing improvement in normal and clustered user distribution LTE advanced HetNets through a hybrid channel-gain access-aware cell selection scheme. International Journal of Computer Applications, 165(??):30–36, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number4/27564-2017913854/.

Erande:2014:CGH

[ETI14] Pratik Erande, Noshir Tarapore, and Vrushali Inamdar. Clustered graph hierarchical layout algorithm for systems biology models. International Journal of Computer Applications, 96(??):28–35, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume96/number24/16945-7012/.

ElAlami:2020:DSS

[ETKE20] M. E. ElAlami, S. M. K. Tobar, S. M. Kh, and Eman A. Esmaeil. Design of simulation system based on digital processing of sound to identify the genres of Arabic music. International Journal of Computer Applications, 175(??):36–39, October 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number26/31618-2020920814/.

Elazhari:2013:ELE

[ETLE13] Maryam Elazhari, Ahmed Toumanari, Rachid Latif, and Nadya El Moussaid. Enhanced low energy communication protocol for implantable wireless body sensor networks. International Journal of Computer Applications, 75 (??):26–29, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number11/13157-0854/.

Ekong:2023:MLB

[EUEI23] Anietie Ekong, Edward Udo, Otuekong Ekong, and Savior Inyang. Machine learning based model for the prediction of fasting blood sugar level towards cardiovascular disease control for the enhancement of public health. International Journal of Computer Applications, 184(??):5–12, March 2023. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume184/number52/
32657-2023922622/.

Egbugha:2017:DII

[EUT17] Malachi C. Egbugha, I. J. Umoh, and A. M. S. Tekanyi. Development of an improved intrusion detection based secured robust header compression technique. International Journal of Computer Applications, 161(??):22–29, March 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume161/number7/27161-2017913234/

.

El-zawawi:2013:DWS

[EzHEgT13] Nour El-zawawi, Mohamed Hamdy, Raina El-gohary, and Mohamed F. Tolba. A database watermarking service with a trusted authority architecture for cloud environment. International Journal of Computer Applications, 69 (??):1–9, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number13/11899-7964/.

Fahmy:2013:ABC

[FA13] Ashraf Anwar Fahmy and Fahad Al Raddady. ANN-based classifier for PAF prediction. International Journal of Computer Applications, 83(??):7–13, December 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume83/number17/14667-2811/.

Fayose:2019:DFB

[FA19] Taiwo Stephen Fayose and Kayode Ayinde. Different forms biasing parameter for generalized ridge regression estimator. International Journal of Computer Applications, 181(??):21–29, January 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number37/30274-2019918339/.

Fellaji:2013:CDS

[FAA13] Soumaya Fellaji, Abdellah Azmani, and Abdelhadi Akharif. Conception of decision support tool for the prevention and detection of type 2 diabetes. International Journal of Computer Applications, 76(??):24–27, August 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume76/number10/13284-0704/.

Fadhil:2018:BTG

[Fad18] Anfal A. Fadhil. Build tool to generate relational database from class diagram. International Journal of Computer Applications, 181(??):14–17, December 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number31/30185-2018918095/.

Fazlyab:2014:DAC

[FAK14] Alireza Fazlyab, Abbas Ajorkar, and Mansour Kabganian. Design of an adaptive controller of a satellite using thruster actuator. International Journal of Computer Applications, 102(??):6–12, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume102/number10/17849-8798/.

Faraj:2011:DED

[Far11] Khalid Faraj. Design error detection and correction system based on Reed–Muller matrix for memory protection. International Journal of Computer Applications, 34(??):42– 48, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume34/
number8/4123-5929/.

Farag:2012:HRS

[Far12] M. S. Farag. HandwrittenText recognition system for automatic reading of historical Arabic manuscripts. International Journal of Computer Applications, 60(??):31– 37, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number13/9754-4383/.

Farag:2013:BWA

[Far13] Mohamed S. Farag. Building a website architecture based on face authentication for developing trust between consumers in KSA. International Journal of Computer Applications, 72(??):45–53, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume72/number2/12469-8858/.

Farhadi:2014:CSA

[Far14a] Babak Farhadi. Creating a semantic academic lecture video search engine via enrichment textual and temporal features of subtitled YouTube EDU media fragments. International Journal of Computer Applications, 96(??): 13–18, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number13/16853-6719/.

Farhadi:2014:NFN

[Far14b] Babak Farhadi. NER-FL: a novel named entity recognizer of Farsi language using the Web-based natural language processors and semantic annotations. International Journal of Computer Applications, 98(??):34–37, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume98/number1/17150-7193/.

Fatima:2016:SST

[FAR16] Nikhat Fatima, Ayesha Ameen, and Syed Raziuddin. STQP: Spatio-temporal indexing and query processing. International Journal of Computer Applications, 150(??): 5–9, September 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume150/number10/26127-2016911514/.

Farhadi:2017:NCB

[Far17] Babak Farhadi. A novel connectivity-based reinforcement learning algorithm for ranking linked data on the Social Semantic Web. International Journal of Computer Applications, 171(??):6–12, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume171/number1/28143-2017914598/.

Faruque:2024:CAS

[Far24] Mohammad Gazi Golam Faruque. Creating an ADU system: Design and development of a mobile application tailored for users battling addiction. International Journal of Computer Applications, 186(??):20–24, May 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number19/creating-an-adu-system-design-
and-development-of-a-mobile-application-tailored-
for-users-battling-addiction/.

Faek:2013:SRN

[FAt13] Fatima K. Faek and Abdulbasit K. Al-talabani. Speaker recognition from noisy spoken sentences. International Journal of Computer Applications, 70(??):11–14, May 2013. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume70/number20/12182-8213/

.

Fathy:2013:MHB

[FAZ13] Ramy A. Fathy, Ahmed A. Abdelhafez, and Abdelhalim Zekry. Meta-heuristic based adaptation engine for cognitive radio systems. International Journal of Computer Applications, 64(??):53–60, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume64/number18/10738-5686/.

Freitas:2011:SNR

[FCS11] Fabrìcio G. Freitas, Daniel P. Coutinho, and Jerffeson T. Souza. Software next release planning approach through exact optimization. International Journal of Computer Applications, 22(??):1–8, May 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume22/number8/2607-3636/.

Flora:2014:AAA

[FCW14] Harleen K. Flora, Swati V. Chande, and Xiaofeng Wang. Adopting an agile approach for the development of mobile applications. International Journal of Computer Applications, 94(??):43–50, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume94/number17/16454-6199/.

Ferdous:2014:SAA

[FD14] S. M. Ferdous and Anindya Das. A simulated annealing approach for solving Minimum Manhattan Network Problem. International Journal of Computer Applications, 98 (??):1–6, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number22/17312-7654/.

Fernandes:2017:CPS

[FDG17] Sara P. Fernandes, Glêndara Aparecida De Souza Martins, and Warley Gramacho Da Silva. Computational program for simulation of mass and heat transfer of Cerrado fruits. International Journal of Computer Applications, 165(??):1–5, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number5/27566-2017913856/.

Fayek:2013:MOO

[FEbO13] Magda B. Fayek, Hatem M. El-boghdadi, and Sherin M. Omran. Multi-objective optimization of technical stock market indicators using GAs. International Journal of Computer Applications, 68(??):41–48, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number20/11698-7428/.

Fettouh:2013:IVH

[FEE13] Abdessadeq Fettouh, Najib El Kamoun, and Abdelaziz El Fazziki. Improving vertical handover performance of real time applications over heterogeneous wireless networks. International Journal of Computer Applications, 62(??):11– 18, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number21/10219-4936/.

Feng:2018:SAS

[Fen18] Geng Feng. Sensitivity analysis for systems under epistemic uncertainty with probability bounds analysis. International Journal of Computer Applications, 179(??):1–6, April 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume179/number31/29191-
2018915892/.

Fernando:2014:MFN

[Fer14] Subha Fernando. Modeling function of nectar foraging of honeybees using operant conditioning. International Journal of Computer Applications, 104(??):45–51, October 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume104/number7/18218-9248/

.

Fouad:2013:ADC

[FH13] Khaled M. Fouad and Moataz O. Hassan. Agent for documents clustering using semantic-based model and fuzzy. International Journal of Computer Applications, 62(??): 10–16, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number3/10059-4651/.

Fourati:2016:BSE

[FHM16] Sonia Fourati, Soumaya Hamouda, and B. T. Maharaj. Bargaining solutions for energy efficient and fair power allocation in cognitive D2D communications. International Journal of Computer Applications, 155(??):24– 31, December 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume155/number6/26609-2016912328/.

FadlElssied:2011:RFM

[FIA11] Nadir Omer FadlElssied, Othman Ibrahim, and Adil Ali A. Alaziz. A review of fuzzy mechanismsfor e-government security. International Journal of Computer Applications, 34 (??):16–22, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume34/
number7/4111-5914/.

Fowdur:2016:NPC

[FIS16] T. P. Fowdur, D. Indoonundon, and K. M. S. Soyjaudah. A novel prioritised concealment and flexible macroblock ordering scheme for video transmission. International Journal of Computer Applications, 150(??): 35–42, September 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume150/number6/26100-2016911554/.

Fakhye:2016:DRG

[FJM16] Rodnny Jesus Mendoza Fakhye, Geraldo Carvalho Brito Junior, and Roberto Dalledone Machado. Dynamic response of gravity dam model using isogeometric analysis. International Journal of Computer Applications, 151(??):7–14, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume151/number4/26219-2016911731/.

Fernando:2012:MCA

[FK12] T G I Fernando and Tatiana Kalganova. Multi-colony ant systems for multi-hose routing. International Journal of Computer Applications, 59(??):1–14, December 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume59/number2/9517-
3931/.

Fernando:2015:MHU

[FK15] Subha Fernando and Nishantha Kumarasinghe. Modeling a honeybee using spiking neural network to simulate nectar reporting behavior. International Journal of Computer Applications, 130(??):32–39, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number8/23233-2015907078/.

FarmanUllah:2012:ECC

[FKFU12] FarmanUllah, Aamir Khan, Hasan Farooq, and Imdad Ullah. Effect of channel coding with node cooperation for wireless passive sensor networks. International Journal of Computer Applications, 39(??):23–29, February 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume39/number10/4857-7127/.

Fergani:2012:CFF

[FkKB12] B. Fergani, Mohamed khireddine Kholladi, and M. Bahri. Comparison of FCM and FISODATA. International Journal of Computer Applications, 56(??):35–39, October 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume56/number8/8913-

Fahmi:2012:SVM

[FKT12] Ben Rejab Fahmi, Nouira Kaouther, and Abdelwahed Trabelsi. Support vector machines versus multi-layer perceptrons for reducing false alarms in intensive care units. International Journal of Computer Applications, 49(??): 41–47, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number11/7675-0969/.

Flayyih:2018:CAM

[Fla18] Wameedh Nazar Flayyih. Crosstalk aware multi-bit error detection with limited error correction coding for reliable on-chip communication. International Journal of Computer Applications, 179(??):1–8, May 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number40/29344-2018916934/.

Flach:2020:ABS

[FLLR20] Quezia N. Flach, Arthur F. Lorenzon, Marcelo C. Luizelli, and Fabio D. Rossi. Analysis of biological sequence search performance in NoSQL database. International Journal of Computer Applications, 176(??):1–6, July 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number42/31480-2020920416/.

Faycal:2011:DOP

[FLM11] Zaidi Faycal, Khamer Lazhar, and Zaidi Mohamed. A development-oriented process for building Web services ontology using OWL-S language: Application in medical Web services. International Journal of Computer Applications, 34 (??):8–14, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume34/
number5/4097-5940/.

Fatah:2013:EPL

[FM13] Safwan E. Abdul Fatah and Sami A. Mawjoud. Effect of physical layer limitations on the MAC layer capacity of mobile WiMAX system, analysis and simulation. International Journal of Computer Applications, 70 (??):31–37, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number8/11984-7859/.

Figueiredo:2020:CTO

[FM20] Levi P. Figueiredo and Jose E. B. Maia. Cooperative target observation moving over a planar graph using a modified hill climbing search. International Journal of Computer Applications, 176(??):1–5, May 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number19/31304-2020920134/.

Farahani:2020:VOT

[FML20] Mohamad Hosein Davoodabadi Farahani, Mohsen Khan Mohamadi, and Mojtaba Lotfizad. Visual object tracking using sparse representation and interest points in a double step approach. International Journal of Computer Applications, 175(??):1–9, August 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number10/31486-2020920563/.

Ferdous:2013:DSR

[FMS13] Saba Ferdous, Muhammad Usman Mirza, and Usman Saeed. Docking studies reveal phytochemicals as the long searched anticancer drugs for breast cancer. International Journal of Computer Applications, 67(??):1–5, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume67/number25/11740-7073/.

Fouad:2013:BLR

[FNH13] Khaled M. Fouad, Mostafa A. Nofal, and Jehan A. Hasan. Building learning repository based on semantic and shareable learning objects. International Journal of Computer Applications, 70(??):34–42, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume70/number17/12162-8205/.

Freitas:2013:OBI

[FPA13] Mario H. G. Freitas, Flavio L. C. Padua, and Guilherme T. Assis. Object-based image retrieval using local feature extraction and relevance feedback. International Journal of Computer Applications, 78(??):8–14, September 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume78/number7/13499-1239/.

Fitriani:2016:SBP

[FPS16] Wirda Fitriani, Andysah Putera, and Utama Siahaan. Singlebit parity detection and correction using Hamming code 7-bit model. International Journal of Computer Applications, 154(??):12–16, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume154/number2/26462-2016912031/.

Firoozbahrami:2012:SND

[FR12] Mahdie Firoozbahrami and Amir Masoud Rahmani. Suitable node deployment based on geometric patterns considering fault tolerance in wireless sensor networks. International Journal of Computer Applications, 60(??):49– 56, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number7/9708-4170/.

Florence:2014:SLA

[FR14] P. V. Florence and G. S. N. Raju. Synthesis of linear antenna arrays using accelerated particle swarm optimization algorithm. International Journal of Computer Applications, 103(??):43–49, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number3/18056-8970/.

Fuad:2015:HSC

[FR15] Kazi Ahmed Asif Fuad and Shahriyar Masud Rizvi. Hardware software co-simulation of Canny edge detection algorithm. International Journal of Computer Applications, 122(??):7–12, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number19/21806-5124/.

Farid:2011:AID

[FRR11a] Dewan Md. Farid, Mohammad Zahidur Rahman, and Chowdhury Mofizur Rahman. Adaptive intrusion detection based on boosting and naīve Bayesian classifier. International Journal of Computer Applications, 24(??): 12–19, June 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume24/
number3/2932-3883/.

Farid:2011:EAC

[FRR11b] Dewan Md. Farid, Mohammad Zahidur Rahman, and Chowdhury Mofizur Rahman. An ensemble approach to classifier construction based on bootstrap aggregation. International Journal of Computer Applications, 25(??): 30–34, July 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume25/
number5/3027-4098/.

Fard:2011:CBT

[FS11] Farzaneh Amiri Fard and Mostafa Setak. Comparison between two algorithms for multi-depot vehicle routing problem with inventory transfer between depots in a three-echelon supply chain. International Journal of Computer Applications, 28(??):39–45, August 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume28/number6/3388-4703/.

Ferhat:2012:IRU

[FSA12] Fillali Ferhat, Maza Sofiane, and Graini Abid. Image restoration using a network of reduced and regularized neural networks. International Journal of Computer Applications, 54 (??):1–6, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume54/
number8/8583-2331/.

Fouad:2020:COF

[FSER20] Mary M. Fouad, Ahmed I. Saleh, and Mohamed F. El-Rahamawy. A co-operative fog-based load balance (CF-BLB) strategy. International Journal of Computer Applications, 177(??):15–21, March 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number41/31176-2020919764/.

S:2013:IMB

[fSG13] Ahmed Abou El fetouh S and Mona Gamal. An intelligent model in bioinformatics based on rough-neural computing. International Journal of Computer Applications, 64 (??):43–48, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number2/10610-5324/.

Forouzannezhad:2017:CWD

[FSG17] Parisa Forouzannezhad, Sahereh SahandAbadi, and Morteza Azizi Ghoomi. Compact and wideband disc monopole antenna based on epsilon negative transmission line for WiFi applications. International Journal of Computer Applications, 175(??):1–5, October 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number3/28465-2017915471/.

Ferdous:2018:EHA

[FSNR18] Md Javedul Ferdous, A. F. M. Saifuddin Saif, Dip Nandi, and Mashiour Rahman. An efficient hybrid architecture for visual behavior recognition using convolutional neural network. International Journal of Computer Applications, 181(??):32–37, November 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number25/30094-2018918048/.

Francis:2018:BIS

[FTA18] Oppong-Twum Francis, Frimpong Twum, and J. B. Hayfron Acquah. Binary image segmentation through the Carl Friedrich Gauss equation. International Journal of Computer Applications, 179(??):1–9, May 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number41/29352-2018916969/.

Frej:2016:DOO

[FTBA16] Ghazi Bel Haj Frej, Assem Thabet, Mohamed Boutayeb, and Mohamed Aoun. Decentralized observers for optimal stabilization of large class of nonlinear interconnected systems. International Journal of Computer Applications, 137(??):1–7, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume137/number14/24445-2016909036/.

Fute:2014:MAP

[FTLK14] Elie Tagne Fute, Emmanuel Tonye, Fabrice Lauri, and Abderrafiaa Koukam. Multi-agent patrolling: Multi-objective approach of the event detection by a mobile wireless sensors network. International Journal of Computer Applications, 88(??):1–8, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume88/number12/15401-3845/.

Fekihal:2012:SOM

[FY12] Mabruk Ali Fekihal and Jabar H. Yousif. Self-organizing map approach for identifying mental disorders. International Journal of Computer Applications, 45(??):25–30, May 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume45/number7/6793-
9120/.

Fernando:2011:NTU

[FYM11] Subha Fernando, Koichi Yamada, and Ashu Marasinghe. New threshold updating mechanism to stabilize activity of Hebbian neuron in a dynamic stochastic ’multiple synaptic’ network, similar to homeostatic synaptic plasticity process. International Journal of Computer Applications, 36(??): 29–37, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume36/
number3/4473-6277/.

Feki:2012:DCA

[FZM12] Elyes Feki, M. Aymen Zermani, and Abdelkader Mami. Decoupling control approach for neonate incubator system. International Journal of Computer Applications, 47 (??):49–57, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number2/7164-9851/.

Gallegos:2015:AST

[GA15] Carlos D. Rodríguez Gallegos and Manuel S. Alvarez Alvarado. Analysis of the stationary and transient behavior of a photovoltaic solar array: Modeling and simulation. International Journal of Computer Applications, 127(??):26–33, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number4/22719-2015906374/.

Gameel:2015:SMC

[GAA15] Ghada K. Gameel, Fayez F. Areed, and Mahmoud E. Abdalla. Sliding mode control for speed of an induction motor with backlash. International Journal of Computer Applications, 125(??):54–58, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number6/22440-2015905943/.

Gaata:2016:EIW

[Gaa16] Methaq T. Gaata. An efficient image watermarking approach based on Fourier transform. International Journal of Computer Applications, 136(??):8–11, February 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume136/number9/24180-2016908559/.

Gbedawo:2017:IDC

[GAA17] Victor Gbedawo, Komi Agbesi, and Theophilus Adukpo. Intrusion detection on campus network, the open source approach: Accra technical university case study. International Journal of Computer Applications, 164(??):20–27, April 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume164/number6/27488-2017913664/

.

Gharehchopogh:2013:AOS

[GAZ13] Farhad Soleimanian Gharehchopogh, Esmail Amini, and Behnam Zebardast. Aspect-oriented software development based solution for intervention concerns problems: Case study. International Journal of Computer Applications, 63 (??):16–25, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number4/10453-5157/.

Gashayija:2011:DDI

[GB11] Jean Marie Gashayija and Elmarie Biermann. Design and development of an image classification and recognition system for CubeSat constellation. International Journal of Computer Applications, 36(??):26–30, December 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume36/number11/4534-6444/.

Gondha:2014:EMP

[GB14] Ankur V. Gondha and Arjav A. Bavarva. An extension of MAC protocol for mobility detection and efficient data transmission. International Journal of Computer Applications, 94(??):28–32, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume94/number6/16348-5692/.

Garg:2015:ICT

[GB15] Vijay Kumar Garg and R. K. Bansal. Intelligent computing techniques for the detection of sleep disorders: a review. International Journal of Computer Applications, 110(??):27–46, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number1/19283-0701/.

Girmay:2018:DCU

[GB18] Gebremichael Girmay and D. Lalitha Bhaskari. Data compression using simulated circular indexing transform (SCIT). International Journal of Computer Applications, 179(??):1–9, May 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume179/number43/
29368-2018917126/.

Gawich:2012:MOB

[GBHI12] M. Gawich, A. Badr, A. Hegazy, and H. Ismail. A methodology for ontology building. International Journal of Computer Applications, 56(??):39–45, October 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume56/number2/8867-2834/.

Gupta:2010:MDT

[GBM10] Meenakshi Gupta, Manisha Bharadia, and Richa Maheshwari. Multirate DSP and its technique to reduce the cost of the analog signal conditioning filters. International Journal of Computer Applications, 4(??):27–34, August 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume4/number10/862-1211/.

Garcia-b:2013:CST

[GBPTEQ^{+}13] Abel Garcia-B., Cristian G. Perez-T., Eduardo S. Espinoza-Q., Francisco R. Trejo-M., Luis E. Ramos-V., Hugo Romero-T., and Jesus M. Munoz-P. Comparative study of three different path tracking controls in mobile robots. International Journal of Computer Applications, 74(??): 47–51, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number5/12885-9823/.

Gulia:2012:NPI

[GC12] Preeti Gulia and Rajender S. Chillar. New proposed inheritance metrics to measure the software complexity. International Journal of Computer Applications, 58(??):1– 4, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number21/9403-3486/.

Gaikwad:2018:DID

[GC18] Kaustubh M. Gaikwad and Mahesh Chavan. Design and implementation of digital cheybshev Type I filter using XSG for noise reduction in ECG signal. International Journal of Computer Applications, 179(??):20–24, February 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number18/28968-2018916314/.

Gangwar:2024:BBA

[GC24] Shivam Gangwar and Anushka Chaurasia. Blockchain-based authentication and verification system for academic certificate using QR code and decentralized applications. International Journal of Computer Applications, 186(??):1–10, July 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number26/
blockchain-based-authentication-and-verification-
system-for-academic-certificate-using-qr-code-and-
decentralized-applications/.

Gayathri:2017:AHM

[GCK17] A. Gayathri, G. Chenchamma, and K. V. V. Kumar. Autoregressive hidden Markov model based speech enhancement using sparsity. International Journal of Computer Applications, 157(??):29–39, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume157/number2/26807-2017912640/.

Gargi:2013:PID

[GCP13] Rajesh Gargi, Yogesh Chaba, and R. B. Patel. Performance improvement of DYMO routing protocol using gateway authentication technique. International Journal of Computer Applications, 66(??):1–7, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume66/number20/11197-5802/.

Ghosh:2014:ATR

[GCS14] Amrit Ghosh, Prasun Chakrabarti, and Pierluigi Siano. Approach towards realizing the security threats for mobile IPv6 and solution thereof. International Journal of Computer Applications, 90(??):26–30, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume90/number10/15758-4417/.

G:2013:DHI

[GD13] Prathibha P G and Dileesh E D. Design of a hybrid intrusion detection system using Snort and Hadoop. International Journal of Computer Applications, 73(??): 5–10, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number10/12775-9226/.

Gupta:2013:CCE

[GDD13] Shilpi Gupta, Mehul V. Desai, and Upena D. Dalal. Comparison of channel estimation in mobile WiMAX (IEEE 802.16e) DL-PUSC system. International Journal of Computer Applications, 84(??):18–21, December 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume84/number7/14589-2824/.

Guchhait:2012:CEH

[GDJ12] S. K. Guchhait, S. Das, and R. N. Jana. Combined effect of hall current and rotation on MHDmixed convection oscillating flow in a rotating vertical channel. International Journal of Computer Applications, 49(??):1–11, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume49/number13/7684-0990/.

Goswami:2016:PCA

[GDS16] Abhinandan Goswami, Srikanta Dash, and Yang Saring. Packet combining and APC using decode and forward relay over wireless networks. International Journal of Computer Applications, 140(??):45–49, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume140/number3/24578-2016909268/.

GarridoMartinez-Llop:2021:PPR

[GDS21] Pablo Garrido Martínez-Llop, Juan De Dios Sanz Bobi, and Carlos Simancas Cuevas. Price prediction in railway tenders using machine learning. International Journal of Computer Applications, 183(??):9–13, May 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number2/31898-2021921291/.

Ghiduk:2012:DIU

[GE12] Ahmed S. Ghiduk and Mohammed Elashiry. Design and implementation of the user interfaces and the applications for mobile devices. International Journal of Computer Applications, 46(??):12–21, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume46/number12/6960-9423/.

Gad-Elrab:2017:AHM

[GE17] Ahmed A. A. Gad-Elrab. Adaptive hybrid methods for cooperative caching consistency in mobile ad-hoc networks. International Journal of Computer Applications, 169(??):34–45, July 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume169/number2/27960-2017914613/.

Gad-ElRab:2016:ESS

[GEA16] Ahmed A. A. Gad-ElRab and Doaa M. Alhilaly. Efficient selection scheme for data processing in wireless sensor networks. International Journal of Computer Applications, 139(??):8–16, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number8/24509-2016909230/.

Gad-ElRab:2015:AMB

[GEAE15] Ahmed A. A. Gad-ElRab, T. A. A. Alzohairy, and Farouk A. Emara. Application method-based efficient offloading scheme in mobile cloud computing. International Journal of Computer Applications, 132(??):1–8, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume132/number3/23571-2015907304/.

Gad-ElRab:2018:AAR

[GEAE18] Ahmed A. A. Gad-ElRab, T. A. A. Alzohairy, and Amr T. A. Elsayed. An adaptive activity recognition approach in smart environments. International Journal of Computer Applications, 181(??):15–23, August 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number13/29878-2018917738/.

Gerami:2012:LSB

[GEB12] Parisa Gerami, Subariah Ebrahim, and Morteza Bashardoost. Least significant bit image steganography using particle swarm optimization and optical pixel adjustment. International Journal of Computer Applications, 55(??):20–25, October 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume55/number2/8727-
2602/.

Girgis:2023:AET

[GEG23] Moheb R. Girgis, Marina Esam, and Mamdouh M. Gomaa. Automated extractive text summarization using genetic and simulated annealing algorithms and their hybridization. International Journal of Computer Applications, 185(??):34–43, September 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number33/32904-2023923105/.

Girgis:2016:ANW

[GENERM16] Moheb R. Girgis, Alaa I. El-Nashar, Tarek A. Abd El-Rahman, and Marwa A. Mohammed. An ASP.NET Web applications data flow testing approach. International Journal of Computer Applications, 153(??):6– 18, November 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume153/number8/26421-2016912117/.

Gomaa:2014:ASA

[GF14] Wael Hassan Gomaa and Aly Aly Fahmy. Arabic short answer scoring with effective feedback for students. International Journal of Computer Applications, 86(??):35– 41, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number2/14961-3177/.

Ghasemi:2017:PFG

[GF17] Maryam Ghasemi and Ali Farzan. Pareto-front generation by classical and meta-heuristic methods in flexible job shop scheduling with multiple objectives. International Journal of Computer Applications, 165(??):15–21, May 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume165/number1/27537-2017913476/.

Ghazouani:2014:ISR

[GFMS14] Mohamed Ghazouani, Sophia Faris, Hicham Medromi, and Adil Sayouti. Information security risk assessment — a practical approach with a mathematical formulation of risk. International Journal of Computer Applications, 103(??):36–42, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number8/18097-9155/.

Goyal:2011:SWU

[GG11] Sachin Goyal and Roopam Gupta. Sliding window used for robustness and fidelity optimization employing neighborhood concept and genetic algorithm with tournament selection. International Journal of Computer Applications, 20 (??):36–42, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume20/
number8/2452-3310/.

Gupta:2012:SEP

[GG12] Kapil Gupta and P. K. Ghosh. Study of error performance of rotated PSK modulation in Nakagami-q (Hoyt) fading channel. International Journal of Computer Applications, 41(??):37–43, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume41/number17/5635-8018/.

Gola:2014:IRA

[GGI14] Kamal Kr. Gola, Bhumika Gupta, and Zubair Iqbal. Implementation and result analysis of secure strategy for high speed transmission and efficient collection of data in wireless sensor network. International Journal of Computer Applications, 108(??):19–27, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number12/18963-0297/.

Goswami:2014:AOR

[GGK14] Anil Kumar Goswami, Shalini Gakhar, and Harneet Kaur. Automatic object recognition from satellite images using artificial neural network. International Journal of Computer Applications, 95(??):33–39, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number10/16633-6502/.

Guitouni:2023:SAM

[GGM23] Zied Guitouni, Mohammed Ali Ghaieb, and Mohsen Machhout. Security analysis of medical image encryption using AES modes for IoMT systems. International Journal of Computer Applications, 185(??):15–21, April 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number2/32677-2023922668/.

Guangling:2012:ISR

[GGX12] Sun Guangling, Li Guoqing, and Jiang Xiaoqing. Image super-resolution reconstruction based on multi-groups of coupled dictionary and alternative learning. International Journal of Computer Applications, 41(??):22– 31, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume41/
number10/5577-7683/.

Ghodrati:2014:PNM

[GH14] Masoomeh Ghodrati and Ali Harounabadi. Provide a new mapping for deadlock detection and resolution modeling of distributed database to colored Petri net. International Journal of Computer Applications, 95(??): 1–7, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number5/16587-6289/.

Ghamry:2013:FIH

[Gha13] Nivin Ghamry. FPGA implementation of hearing aids using stationary wavelet-packets for denoising. International Journal of Computer Applications, 69(??):30– 36, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number15/11921-8125/.

Ghafil:2016:IAS

[Gha16] Hazim Nasir Ghafil. Inverse acceleration solution for robot manipulators using harmony search algorithm. International Journal of Computer Applications, 144(??):1–7, June 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume144/number6/25180-2016910297/

.

G:2012:DOB

[GHBN12] Rangaraju H G, Vinayak Hegde, Raja K B, and Muralidhara K N. Design and optimization of n-bit reversible binary comparator. International Journal of Computer Applications, 55(??):22–30, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume55/number18/8992-3072/.

Gbegbe:2017:EBR

[GHD17] Raymond Gbegbe, Soumaya Hamouda, and Guy Landry Diety. Energy-based RRH assignment scheme for heterogeneous cloud radio access networks. International Journal of Computer Applications, 180(??):33–39, December 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number2/28774-2017915948/.

Gherboudj:2016:CSC

[Ghe16] Amira Gherboudj. Comparative study of cuckoo inspired metaheuristics applying to knapsack problems. International Journal of Computer Applications, 155(??):25–31, December 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume155/number12/26658-
2016912508/.

Giang:2015:DBA

[GHH15] Nguyen Long Giang, Pham Minh Ngoc Ha, and Nguyen Manh Hung. Distance based attribute reduction in set-valued decision tables. International Journal of Computer Applications, 117(??):8–13, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number24/20701-3378/.

Ghodselahi:2011:HSV

[Gho11] Ahmad Ghodselahi. A hybrid support vector machine ensemble model for credit scoring. International Journal of Computer Applications, 17(??):1–5, March 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume17/number5/2220-2829/.

Ghose:2015:AEF

[Gho15] Supratip Ghose. Analysis and effect of feature selection with generic classifier over activity recognition dataset for daily life activity recognition. International Journal of Computer Applications, 131(??):33–39, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume131/number16/23536-2015907622/.

G:2013:CFM

[GJ13a] Nagaraja B G and H. S. Jayanna. Combination of features for multilingual speaker identification with the constraint of limited data. International Journal of Computer Applications, 70(??):1–6, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume70/number6/11963-7823/.

Ghai:2013:TED

[GJ13b] Deepika Ghai and Neelu Jain. Text extraction from document images- a review. International Journal of Computer Applications, 84(??):40–48, December 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume84/number3/14559-2661/.

Gupta:2014:NIE

[GJ14] Ritu Gupta and Anchal Jain. A new image encryption algorithm based on DNA approach. International Journal of Computer Applications, 85(??):27–31, January 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume85/number18/14942-3440/

.

Gaikwad:2015:BTC

[GJ15a] Sonali B. Gaikwad and Madhuri S. Joshi. Brain tumor classification using principal component analysis and probabilistic neural network. International Journal of Computer Applications, 120(??):5–9, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number3/21205-3885/.

Gandhi:2015:PFM

[GJ15b] Jenish R. Gandhi and Rutvij H. Jhaveri. Packet forwarding misbehaviour isolation using fuzzy trust-based secure routing in MANET. International Journal of Computer Applications, 122(??):30–35, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number3/21684-4786/.

Gaikwad:2016:FPD

[GJG16] Suhas Machhindra Gaikwad, Rahul Joshi, and Shishir Machhindra Gaikwad. Find out pen drive location with the help of mobile GPS. International Journal of Computer Applications, 135(??):14–18, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number2/24020-2016908322/.

Goswami:2016:NNA

[GJM16] Anil Kumar Goswami, Heena Joshi, and S. P. Mishra. Neural network approach for automatic landuse classification of satellite images: One-against-rest and multi-class classifiers. International Journal of Computer Applications, 134(??):35–42, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number11/23961-2016908148/.

Gaurav:2013:AEE

[GJP13] Kunal Gaurav, Pooja Jha, and K. S. Patnaik. Analyzing effort estimation in multistage based FL-COCOMO II framework using various fuzzy membership functions. International Journal of Computer Applications, 66(??): 35–39, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number6/11092-6051/.

Garg:2017:CAV

[GJS17] Diksha Garg, Ankita Jiwan, and Shailendra Singh. Computational approaches for variant identification. International Journal of Computer Applications, 165(??):18–24, May 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume165/number8/27594-2017913970/

.

Gangwani:2023:IDI

[GJUL23] Pranav Gangwani, Santosh Joshi, Himanshu Upadhyay, and Leonel Lagos. IoT device identity management and blockchain for security and data integrity. International Journal of Computer Applications, 184(??):49–55, January 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume184/number42/32593-
2023922529/.

Goyal:2010:DHP

[GK10] Vishal Goyal and Pardeep Kumar. Development of Hindi– Punjabi parallel corpus using existing Hindi–Punjabi machine translation system and using sentence alignments. International Journal of Computer Applications, 5(??):15–19, August 2010. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume5/number9/941-
1319/.

Gharehchopogh:2012:AED

[GK12a] Farhad Soleimanian Gharehchopogh and Zeinab Abbasi Khalifehlou. Analysis and evaluation of dynamic load balancing in IEEE 802.11b wireless local area networks. International Journal of Computer Applications, 47(??): 9–12, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number22/7486-0193/.

Gharehchopogh:2012:DMA

[GK12b] Farhad Soleimanian Gharehchopogh and Seyyed Reza Khaze. Data mining application for cyber space users tendency in blog writing: a case study. International Journal of Computer Applications, 47(??):40–46, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume47/number18/7291-0509/.

Ghotkar:2013:VBR

[GK13] Archana S. Ghotkar and Gajanan K. Kharate. Vision based real time hand gesture recognition techniques for human computer interaction. International Journal of Computer Applications, 70(??):1–8, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume70/number16/12148-8103/.

Grosek:2015:SSS

[GK15a] Jacob Grosek and J. Nathan Kutz. Selecting a small set of optimal gestures from an extensive lexicon. International Journal of Computer Applications, 119(??):1–8, June 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume119/number5/21060-3722/

.

Gupta:2015:PEP

[GK15b] Rajeev Gupta and Ashok Kumar. Performance evaluation of proposed segmentation framework with existing techniques for noisy iris images. International Journal of Computer Applications, 114(??):1–6, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume114/number1/19939-1723/.

Gupta:2015:HAR

[GK15c] Sarthak Gupta and Ajeet Kumar. Human activity recognition through smartphone’s tri-axial accelerometer using time domain wave analysis and machine learning. International Journal of Computer Applications, 127(??):22–26, October 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume127/number18/22831-
2015906733/.

Gyamerah:2021:PSM

[GK21] Seth Gyamerah and Dennis Redeemer Korda. Prediction of stock market returns using LSTM model and traditional statistical model. International Journal of Computer Applications, 183(??):57–61, November 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number37/32175-2021921773/.

Ginish:2016:ODC

[GKG16] Ginish, Munish Kumar, and Amit Grover. Optimal duty cycling with sleep-wake schedule between paired nodes and flexible routing across pairs. International Journal of Computer Applications, 144(??):20–24, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number8/25200-2016910419/.

Gupta:2013:RTF

[GKK13] Anand Kr. Gupta, Sushant Katiyar, and Nitin Kumar. Railway track finding system with RFID application. International Journal of Computer Applications, 83(??):24– 30, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number7/14461-2738/.

Gafoor:2018:ASS

[GKM^{+}18] Rishbiya Abdul Gafoor, Riya Kuriakose, Sibila M, Lakshmi C K, and Reshmi S. Adaptive spectrum sensing in cognitive radio networks. International Journal of Computer Applications, 179(??):10–16, May 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number45/29434-2018917117/.

Garg:2016:SMA

[GL16] Nidhi Garg and Sachin Lakra. A software metric for the activeness of a local area network. International Journal of Computer Applications, 156(??):7–14, December 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume156/number10/26743-
2016912526/.

Gupta:2013:FES

[GM13] Munish Gupta and Jagmohan Mago. Fuzzy expert system to evaluate the quality of service provider in the implementation of e-government. International Journal of Computer Applications, 67(??):16–22, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume67/number18/11495-7205/.

G:2014:HRJ

[GM14a] Ravi (Ravinder) Prakash G and Kiran M. How replicated join expressions equal map phase or reduce phase in a MapReduce structureh. International Journal of Computer Applications, 107(??):43–50, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number12/18807-0431/.

G:2014:HMM

[GM14b] Ravinder Prakash G and Kiran M. How minimal are MapReduce arrangements for binning expressionsh. International Journal of Computer Applications, 99(??):7– 14, August 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume99/
number11/17415-8195/.

G:2015:CRE

[GM15a] Ravi (Ravinder) Prakash G and Kiran M. Characterization of randomized external source output map reduce expressions. International Journal of Computer Applications, 123(??):9–16, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number14/22025-2015905403/.

G:2015:DTE

[GM15b] Ravi (Ravinder) Prakash G and Kiran M. Does there exist pruning decomposition for MapReduce expressions arrangements? International Journal of Computer Applications, 125(??):41–48, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number12/22488-2015906155/.

G:2015:CFP

[GM15c] Ravi (Ravinder) Prakash G and Kiran M. On chain folding problems of chain mapper and chain reducer meta expressions. International Journal of Computer Applications, 116(??):35–42, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number16/20424-2740/.

G:2015:JMM

[GM15d] Ravi (Ravinder) Prakash G and Kiran M. On job merging MapReduce meta expressions for multiple decomposition mapping and reducing. International Journal of Computer Applications, 118(??):14–21, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number13/20804-3501/.

G:2016:WMU

[GM16] Ravi (Ravinder) Prakash G and Kiran M. What is a minimum of unpredictable workload pattern over all elastic scaling in cloud computing? International Journal of Computer Applications, 140(??):44–53, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume140/number4/24586-2016909294/.

Gabisa:2020:TID

[GM20] Getu Gabisa and Mengesha Mamo. Teaching the implementation of digital control using Proteus VSM software. International Journal of Computer Applications, 177(??):16–21, January 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number29/31083-2020919766/.

Girgis:2014:SWM

[GMAR14] Moheb R. Girgis, Tarek M. Mahmoud, Bahgat A. Abdullatif, and Ahmed M. Rabie. Solving the wireless mesh network design problem using genetic algorithm and simulated annealing optimization methods. International Journal of Computer Applications, 96(??):1–10, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume96/number11/16835-6680/.

Gerakidis:2021:EBT

[GMM21] Sergios Gerakidis, Sofia Megarchioti, and Basilis Mamalis. Efficient big text data clustering algorithms using Hadoop and spark. International Journal of Computer Applications, 174(??):13–21, January 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number15/31753-2021921030/.

Gautam:2012:CAS

[GMRK12] Deepraj Gautam, Holytiana Macha, Rajdeep, and Mradul Kumar. A cognitive approach for safe sodium phosphate enema administration by implementing wireless body area network. International Journal of Computer Applications, 57 (??):6–11, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number13/9172-3573/.

Gomathy:2012:GCC

[GMS12] M. Gomathy, K. Meena, and K. R. Subramaniam. Gender clustering and classification algorithms in speech processing: a comprehensive performance analysis. International Journal of Computer Applications, 51(??):9– 17, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number20/8156-1533/.

Gangraj:2014:MAB

[GN14] Ebrahim Asadi Gangraj and Nasim Nahavandi. A metaheuristic approach for batch sizing and scheduling problem in flexible flow shop with unrelated parallel machines. International Journal of Computer Applications, 97(??): 31–36, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number6/17013-7292/.

Ghaffarzadeh:2016:KDC

[GNN16] Parvin Ghaffarzadeh, Mohammad H. Nadimi, and Akbar Nabiollahi. KMGEM: Data clustering by combination of K-means and grenade explosion algorithm. International Journal of Computer Applications, 147(??):21–29, August 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume147/number14/25826-
2016911333/.

Guvenc:2015:DWB

[GOCA15] Ercument Guvenc, Osman Ozkaraca, Gurcan Cetin, and Gurbuz Akcay. Development of a Web-based decision support system for pediatric patients. International Journal of Computer Applications, 114(??):15–20, March 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume114/number6/19982-1924/.

Geetha:2012:UAR

[GP12] M. Kalaiselvi Geetha and S. Palanivel. Unsupervised approach for retrieving shots from video. International Journal of Computer Applications, 60(??):1–8, December 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume60/number6/9693-
4144/.

G:2013:IDL

[GP13] Bharadwaja Kumar G and Melvin Jose Johnson Premkumar. Issues in developing LVCSR system for Dravidian languages: an exhaustive case study for Tamil. International Journal of Computer Applications, 70(??):1–7, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume70/number19/12172-8180/.

Ghosal:2014:LFG

[GP14] Balaram Ghosal and Asim Kumar Panda. Li-Fi: a green energy initiative. International Journal of Computer Applications, 95(??):1–4, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number11/16635-6507/.

Ghosh:2015:ACL

[GP15] Souvik Ghosh and Ranjan Parekh. Automated color logo recognition system based on shape and color features. International Journal of Computer Applications, 118(??):13–20, May 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume118/number12/
20796-3457/.

Gration:2022:IVU

[GPB22] Justuce Muhoza Gration, Kweyamba Maximillian Pastory, and Mawazo Baruti. Implementation and validation of USSD-based mobile phone application prototype for enforcementof property-related tax compliance: the case of Tanzania. International Journal of Computer Applications, 184(??):15–20, November 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number37/32555-2022922464/.

Gupta:2016:WSD

[GPC16] Anshika V. Gupta, B. M. Patil, and V. M. Chandode. Watermarking shape datasets with utility and distance preservation. International Journal of Computer Applications, 133(??):4–9, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume133/number16/23868-2016908092/.

Gupta:2014:ESV

[GPR14] Pratishtha Gupta, G N Purohit, and Manisha Rathore. Estimating speed of vehicle using centroid method in MAT-LAB. International Journal of Computer Applications, 102(??):1–8, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume102/number14/17880-8451/.

Goel:2014:NSN

[GPS14] Nitin Goel, Neetesh Purohit, and B. R. Singh. A new scheme for network selection in heterogeneous wireless network using fuzzy logic. International Journal of Computer Applications, 88(??):1–5, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume88/number3/15329-3650/.

Gaffar:2018:TFE

[GPWS18] Achmad Fanany Onnilita Gaffar, Arief Bramanto Wicaksono Putra, Agusma Wajiansyah, and Supriadi. Texture feature extraction based on local weighting pattern (LWP) using fuzzy logic approach. International Journal of Computer Applications, 179(??):1–8, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number28/29133-2018916614/.

Gyamfi:2022:GIP

[GQD^{+}22] Enoch Opanin Gyamfi, Zhiguang Qin, Juliana Mantebea Danso, Daniel Adu-Gyamfi, Nelson Opoku-Mensah, Noble Arden Elorm Kuadey, and Daniel Konin. Geometricinspired particle swarm optimization (PSO) for classification tasks. International Journal of Computer Applications, 183(??):32–40, February 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number53/32292-2022921954/.

Gugnani:2014:TIE

[GR14] Shashank Gugnani and Rajendra Kumar Roul. Triple indexing: an efficient technique for fast phrase query evaluation. International Journal of Computer Applications, 87 (??):9–13, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number13/15266-3970/.

Graciano:2013:MOE

[GRC13] Marcio L. Graciano, Alexandre R. S. Romariz, and Jose Camargo Da Costa. Methodology for objective evaluation of video broadcasting quality using a video camera at the user’s home. International Journal of Computer Applications, 63(??):37–42, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume63/number16/10554-5750/.

Gomes:2020:CSB

[GRL20] Raphael Gomes, Geovany Rodrigues, and Gilberto Lobo. A cloud service broker for cost effective infrastructure selection using multiple deployment options. International Journal of Computer Applications, 175(??):1–8, November 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume175/number32/31654-
2020920874/.

George:2013:EMD

[GRP13] Sudhish N George, Arun Raj R, and Deepthi P P. An efficient multimedia DRM technique using secure arithmetic coding. International Journal of Computer Applications, 61 (??):1–8, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number17/10017-4593/.

Golrou:2022:WCE

[GRS22] Aminollah Golrou, Nasrin Rafiei, and Mahdieh Sabouri. Wheelchair controlling by eye movements using EOG based human machine interface and artificial neural network. International Journal of Computer Applications, 184(??):12–18, December 2022. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume184/number38/
32562-2022922465/.

Goyal:2011:PAO

[GS11] Shefali Goyal and Manjit Singh. Performance analysis of optical networks and their improvement using power saving method. International Journal of Computer Applications, 25(??):25–28, July 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume25/
number6/3036-4116/.

G:2012:PON

[GS12a] Saranya K G and G. Sudha Sadhasivam. A personalized online news recommendation system. International Journal of Computer Applications, 57(??):6–14, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number18/9212-3758/.

Ganguly:2012:ECM

[GS12b] Rita Ganguly and Anirban Sarkar. Evaluations of conceptual models for semi-structured database system. International Journal of Computer Applications, 50(??): 5–12, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume50/
number18/7869-1145/.

Gill:2014:OSE

[GS14] Rupali Gill and Jaiteg Singh. An open source ETL tool — medium and small scale enterprise ETL(MaSSEETL). International Journal of Computer Applications, 108(??):

15–22, December 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume108/number4/18899-0190/.

Gayathri:2016:ATU

[GS16a] B. M. Gayathri and C. P. Sumathi. An automated technique using Gaussian naïve Bayes classifier to classify breast cancer. International Journal of Computer Applications, 148(??):16–21, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume148/number6/25761-2016911146/.

Ghafarian:2016:FEP

[GS16b] Ahmad Ghafarian and Seyed Amin Hosseini Seno. Forensics evaluation of privacy of portable Web browsers. International Journal of Computer Applications, 147(??): 5–11, August 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume147/number8/25671-2016911009/.

Ghanem:2011:MBA

[GSA11] Ahmed A. Ghanem, Ahmed I. Saleh, and Hesham A. Ali. MAHEFT-based adaptive grid workflow scheduling approach. International Journal of Computer Applications, 35 (??):22–31, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume35/
number7/4414-6133/.

Gambo:2012:SPQ

[GSA12a] Ishaya Gambo, Abimbola Soriyan, and Philip Achimugu. Software performance quality evaluation of MINPHIS architecture using ATAM. International Journal of Computer Applications, 46(??):30–36, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume46/number23/7107-9782/.

Gupta:2012:MAT

[GSA12b] Deepak Gupta, Sameer Sharma, and Shefali Aggarwal. A mathematical approach for three stage flow shop production schedule with jobs in a string of disjoint job block. International Journal of Computer Applications, 44(??): 41–48, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number19/6375-8836/.

Ghosh:2012:LFB

[GSB12] Nibedita Ghosh, Sharmistha Sharma, and Subhadeep Bhattacharjee. A load flow based approach for optimum allocation of distributed generation units in the distribution network for voltage improvement and loss minimization. International Journal of Computer Applications, 50 (??):15–22, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume50/
number15/7847-1075/.

Ghorbel:2012:RFD

[GSCT12] H. Ghorbel, M. Souissi, M. Chaabane, and F. Tadeo. Robust fault detection for Takagi–Sugeno discrete models: Application for a three-tank system. International Journal of Computer Applications, 44(??):1–7, April 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume44/number18/6360-6599/.

Gajjar:2014:CHS

[GSD14] Sachin Gajjar, Mohanchur Sarkar, and Kankar Dasgupta. Cluster head selection protocol using fuzzy logic for wireless sensor networks. International Journal of Computer Applications, 97(??):38–43, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume97/number7/17022-7310/.

Grosek:2013:EGR

[GSK13] Jacob Grosek, Peizhe Shi, and J. Nathan Kutz. Enhanced gesture recognition performance through improved pre-processing. International Journal of Computer Applications, 62(??):1–8, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume62/number9/10105-4757/.

G:2012:DOR

[GSN12] Rangaraju H G, Aakash Babu Suresh, and Muralidhara K N. Design and optimization of reversible multiplier circuit. International Journal of Computer Applications, 52 (??):44–50, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number10/8242-1523/.

G:2014:DER

[GSN14] Rangaraju H G, Arpitha H S, and Muralidhara K N. Design of efficient reversible multiply accumulate (MAC) unit. International Journal of Computer Applications, 85 (??):1–12, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number16/14922-3338/.

Gangele:2018:DMA

[GSP18] Sharad Gangele, Kirti Soni, and Sunil Patil. Data mining approach towards students behavior assessment methods for higher studies. International Journal of Computer Applications, 181(??):11–14, November 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number30/30171-2018918099/.

G:2021:PPI

[GSR21] Padmashree M G, Arunalatha J S, and Venugopal K R. PIECC: Point inversion algorithm for elliptic curve cryptology to secure IoT data communication. International Journal of Computer Applications, 183(??):1–9, September 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number27/32096-
2021921655/.

Gupta:2020:EER

[GSV20] Aakanshi Gupta, Bharti Suri, and Vijin Vincent. An empirical examination of the relationship between code smells and vulnerabilities. International Journal of Computer Applications, 176(??):1–9, June 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number32/31405-2020920362/.

Ganesh:2011:IIS

[GT11] N. Ganesh and S. Thangasamy. Issues identified in the software process due to barriers found during eliciting requirements on agile software projects: Insights from India. International Journal of Computer Applications, 16 (??):7–12, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume16/
number5/2011-2713/.

Govindasamy:2013:PFL

[GT13] V. Govindasamy and P. Thambidurai. Probabilistic fuzzy logic based stock price prediction. International Journal of Computer Applications, 71(??):28–32, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume71/number5/12356-8669/.

Gumma:2019:ETT

[Gum19] Siva Rama Sastry Gumma. Extracting text from Telugu color documents by removing dither patterns. International Journal of Computer Applications, 181(??):1–7, April 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume181/number48/30476-
2019918653/.

Guo:2018:EMR

[Guo18] Xiaohui Guo. An efficient method for removal of high-density salt-and-pepper noise. International Journal of Computer Applications, 180(??):26–29, April 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number31/29244-2018916813/.

Gonidakis:2015:BAA

[GV15] Dimitrios Gonidakis and Aristidis Vlachos. Bat algorithm approaches for solving the combined economic and emission dispatch problem. International Journal of Computer Applications, 124(??):1–7, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number1/22065-2015905288/.

Ganesan:2010:ANN

[GVRP10] N. Ganesan, K. Venkatesh, M. A. Rama, and A. Malathi Palani. Application of neural networks in diagnosing cancer disease using demographic data. International Journal of Computer Applications, 1(??):76–85, February 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume1/number26/476-783/.

Glockenmeier:2024:SEB

[GVW24] Rachel Glockenmeier, Varghese Vaidyan, and Yong Wang. Symbolic execution-based code coverage framework for augmented software testing. International Journal of Computer Applications, 186(??):1–7, September 2024. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume186/number37/symbolic-execution-based-
code-coverage-framework-for-augmented-software-testing/

.

Gunathilake:2015:CMR

[GYP^{+}15] P. M. P. C. Gunathilake, Y. P. R. D. Yapa, A. J. Pinidiyaarachchi, J. V. Wijayakulasooriya, and R. Ranaweera. Computerized 3D modeling for rapid artificial bone substitute manufacturing. International Journal of Computer Applications, 119(??):1–5, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number17/21156-4177/.

Hussain:2016:ACC

[HA16] Ihsan Y. Hussain and Mustafa S. Abood. Aerodynamic characteristics of CLARK-y smoothed inverted wing with ground effects. International Journal of Computer Applications, 136(??):42–50, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume136/number7/24169-2016908502/.

Htiouech:2017:SAM

[HA17] Skander Htiouech and Ameur Alzaidi. Smart agents for the multidimensional multi-choice knapsack problem. International Journal of Computer Applications, 174(??): 5–9, September 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume174/number6/28409-2017915404/.

Husain:2012:SRI

[HAA12] Sadia Husain, Yasir Ahmad, and M. Afshar Alam. A study on the role of intuitionistic fuzzy set in decision making problems. International Journal of Computer Applications, 48 (??):35–41, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number19/7458-0501/.

Habbi:2016:SCI

[HAA16] Hanan Mikhael D. Habbi, Hussein Jalil Ajeel, and Inaam Ibrahim Ali. Speed control of induction motor using PI and v/F scalar vector controllers. International Journal of Computer Applications, 151(??):

36–43, October 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume151/number7/26248-2016911831/.

Hameed:2015:CMS

[HAAA15] Abdul Adeem Zaily Hameed, Faik Hammad Anter, and Muzhir Shaban Al-Ani. Composite material surface analysis based image texture analysis. International Journal of Computer Applications, 129(??):22–26, November 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume129/number1/23037-2015906810/.

Hassoon:2016:FIB

[HAAT16] Ali Fattah Hassoon, Maher K. Al-Azzawi, and Tariq Tashan. Face identification based on contourlet transform and multilayer perceptron classifier. International Journal of Computer Applications, 133(??):12–18, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number2/23757-2016907384/.

Habashy:2013:IDR

[Hab13a] Shahira M. Habashy. Identification of diabetic retinopathy stages using fuzzy C-means classifier. International Journal of Computer Applications, 77(??):7–13, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number9/13420-1096/.

Haque:2013:ANB

[HaB13b] Md. Mozammel Haque and Md. Al amin Bhuiyan. ART network based face recognition with Gabor filters. International Journal of Computer Applications, 77(??):8– 12, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number13/13541-0861/.

Hussein:2014:MLE

[HAb14] Karim Q. Hussein and Maha A. Al-bayati. Mobile learning: the effective role of multimedia “Analytical approach via an object oriented strategy”. International Journal of Computer Applications, 90(??):1–6, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume90/number8/15591-4106/.

Hridoy:2015:RTB

[HAH15] Syed Akib Anwar Hridoy, Faysal Ahmed, and Md. Shazzad Hosain. Regression testing based on Hamming distance and code coverage. International Journal of Computer Applications, 120(??):1–5, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number14/21292-4308/.

Halder:2013:DCE

[Hal13] Amiya Halder. Dynamic contrast enhancement algorithm. International Journal of Computer Applications, 74(??):1–4, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number12/12934-9879/.

Hameed:2014:FFE

[Ham14] Sundos A. Hameed. Facial features extraction by relative geometrical position. International Journal of Computer Applications, 98(??):37–40, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume98/number15/17263-7619/.

Acquah:2017:EQI

[HAPG17] J. B. Hayfron-Acquah, J. K. Panford, and Y. Poakwah Gyimah. Enhancing quality of images using fuzzy logic and singleton parameters. International Journal of Computer Applications, 165(??):7–13, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number13/27727-2017913991/.

Hassan:2014:SAU

[Has14] Mawia Ahmed Hassan. Synthetic aperture ultrasound image reconstruction. International Journal of Computer Applications, 95(??):17–22, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number3/16574-6259/.

Hassan:2015:DME

[Has15] Doaa Hassan. On determining the most effective subset of features for detecting phishing websites. International Journal of Computer Applications, 122(??):1–7, July 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume122/number20/21813-
5191/.

Hashim:2016:SIA

[Has16] Firas Ali Hashim. Swarm intelligent application in networks routing problem. International Journal of Computer Applications, 133(??):25–28, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume133/number1/23752-2016907730/.

Haruna:2017:IFL

[HAS17] O. E. Haruna, K. A. Abubilal, and A. T. Salawudeen. An improved frame level redundancy scrubbing algorithm for SRAM based FPGA. International Journal of Computer Applications, 170(??):31–38, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume170/number5/28068-2017914844/.

Hasan:2020:SCC

[HAS20] Md. Tareq Hasan, Md. Kawshik Ahmed, and Amathul Hadi Shakara. A survey on cloud computing: Structured organization of data center. International Journal of Computer Applications, 176(??):32–36, June 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number33/31420-2020920410/.

Hasan:2024:ACS

[Has24] Akm Hasan. Advancing cybersecurity strategies: a holistic exploration of operations, data systems, and machine intelligence. International Journal of Computer Applications, 186(??):12–16, March 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number12/advancing-cybersecurity-strategies-
a-holistic-exploration-of-operations-data-systems-
and-machine-intelligence/.

Hossain:2017:IHP

[HASH17] F. M. Tanvir Hossain, Maruf Ahmed, Anik Saha, and Khandaker Tabin Hasan. Identifying human personalized sentiment with streaming data. International Journal of Computer Applications, 160(??):26–31, February 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume160/number7/27087-2017913088/.

Hasan:2016:NAM

[HAT16] Mushtaq Ismael Hasan, Abdul Jabbar F. Ali, and Rhan S. Tufah. Numerical analysis of the magnetohydrodynamic flow and heat transfer in microchannel. International Journal of Computer Applications, 149(??):1–7, September 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume149/number2/25966-2016911349/

.

Hazarika:2014:VVT

[Haz14] G. C. Hazarika. Variable viscosity and thermal conductivity on convective heat transfer in a dusty fluid over a vertical permeable surface with radiation and viscous dissipation. International Journal of Computer Applications, 103(??):39–45, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number11/18121-9439/.

Habibvand:2012:UGA

[HB12a] Golnaz Habibvand and Reza Mossayebi Behbahani. Using genetic algorithm for fuel consumption optimization of a natural gas transmission compressor station. International Journal of Computer Applications, 43(??):1–6, April 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume43/number1/6064-
8180/.

Hussein:2012:CSD

[HB12b] Ensaf Hussein and Mohamed A. Belal. A comparative study of digital watermarking techniques in frequency domain. International Journal of Computer Applications, 52 (??):43–50, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number20/8321-1961/.

Hamdi:2013:ARP

[HB13] N. Hamdi and A. Bouzid. Active and reactive power control of a DFIG for variable speed wind energy conversion using a new controller. International Journal of Computer Applications, 67(??):1–6, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume67/number16/11476-6876/.

Hanchate:2014:SPC

[HB14a] Dinesh B. Hanchate and Rajankumar S. Bichkar. Software project contacts by GRGA scheduling and EVM. International Journal of Computer Applications, 97(??): 1–26, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number13/17064-7502/.

Hanchate:2014:SCC

[HB14b] Dinesh Bhagwan Hanchate and Rajankumar S. Bichkar. SPS by combination of crossover types and changeable mutation SGA. International Journal of Computer Applications, 94(??):1–11, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number10/16376-5882/.

Hassan:2016:WBS

[HB16] Md Shabbir Hassan and Arif M. Bhatti. Web-based system for assessment and continuous improvement. International Journal of Computer Applications, 141(??):13–18, May 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume141/number14/24851-
2016909879/.

Htiouech:2013:USI

[HBA13] Skander Htiouech, Sadok Bouamama, and Rabeh Attia. Using surrogate information to solve multidimensional multichoice knapsack problem. International Journal of Computer Applications, 73(??):1–7, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume73/number13/12798-9883/.

Houda:2012:AMH

[HBN12] Chouiref Houda, Boussaid Boumedyen, and Abdelkrim M. Naceur. Application of moving horizon parameter estimator in fault diagnosis of broken bars in induction motor. International Journal of Computer Applications, 50 (??):19–23, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume50/
number17/7863-1125/.

Hamatta:2012:TNE

[HBS12] Hatem S. A. Hamatta, Mohammad U. Bokhari, and Shams Tabrez Siddigui. TPL: NOVEL efficient threshold power consumption scheme on MANET. International Journal of Computer Applications, 58(??):38–42, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number15/9362-3708/.

Hossain:2015:NTM

[HBS15] Mohammad Alamgir Hossain, Alwi Mohammad A Bamhdi, and Goutam Sanyal. A new tactic to maintain privacy and safety of imagery information. International Journal of Computer Applications, 110(??):6–12, January 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume110/number5/19310-0764/.

Chou:2013:MES

[hC13] Chen huei Chou. Metrics in evaluating software defects. International Journal of Computer Applications, 63(??):23– 29, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number3/10447-5147/.

Hanji:2015:NFF

[HCL15] Geeta Hanji, Basaveshwari C, and M. V. Latte. Novel fuzzy filters for noise suppression from digital grey and color images. International Journal of Computer Applications, 125(??):29–37, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number15/22511-2015906236/.

Hazarika:2014:EVV

[HD14] G. C. Hazarika and Sachin Doley. Effects of variable viscosity and thermal conductivity with chemical reaction on a transient MHD flow past an impulsively started vertical plate with ramped temperature and concentration with viscous dissipation. International Journal of Computer Applications, 107(??):9–19, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number21/19138-9943/.

Hazari:2014:IMT

[HDMK14] Samar Hazari, Jayanta Kumar Dey, K. Maity, and Samarjit Kar. An interactive method for two-plants production inventory control with two-warehouse facility under imprecise environment. International Journal of Computer Applications, 96(??):1–13, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume96/number21/16915-5497/.

Hassan:2015:CSO

[HE15] M. S. Hassan and Adel A. Elbaset. A comparative study for optimum design of grid connected PV system based on actual system specifications. International Journal of Computer Applications, 116(??):19–34, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number3/20316-2375/.

Hazman:2011:SOL

[HEBR11] Maryam Hazman, Samhaa R. El-Beltagy, and Ahmed Rafea. A survey of ontology learning approaches. International Journal of Computer Applications, 22(??):36–43, May 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume22/number9/2610-
3642/.

Hassan:2015:DSS

[HEEG15] Mohamed Hassan, Abeer Badr Eldin, and Alaa El-Ghazali. A decision support system for subjective forecasting of new product sales. International Journal of Computer Applications, 126(??):25–30, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume126/number2/22526-2015905997/.

Hemamalini:2012:RGT

[Hem12] M. Hemamalini. Review on grid task scheduling in distributed heterogeneous environment. International Journal of Computer Applications, 40(??):24–30, February 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume40/number2/4929-

Hemalatha:2022:DPD

[Hem22] M. Hemalatha. Diabetic patient’s data classification and prediction using machine learning ensemble algorithm. International Journal of Computer Applications, 184(??):14–18, June 2022. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume184/number19/
32426-2022922209/.

Hussein:2013:RAH

[HEN13] Sherif E. Hussein and Mahmoud Abo El-Nasr. Resources allocation in higher education based on system dynamics and genetic algorithms. International Journal of Computer Applications, 77(??):40–48, September 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume77/number10/13434-1136/.

Hagag:2012:ETW

[HEs12] Shimaa Hagag and Ayman El-sayed. Enhanced TCP Westwood congestion avoidance mechanism (TCP Westwood-New). International Journal of Computer Applications, 45 (??):21–29, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume45/
number5/6776-9071/.

Hamdache:2014:OCI

[HES14] Amine Hamdache, Ilias Elmouki, and Smahane Saadi. Optimal control with an isoperimetric constraint applied to cancer immunotherapy. International Journal of Computer Applications, 94(??):31–37, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume94/number15/16421-6073/.

Hewahi:2011:RBK

[Hew11] Nabil M. Hewahi. Rule-based knowledge representation inspired from finite automata. International Journal of Computer Applications, 21(??):14–17, May 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume21/number3/2493-3367/.

Haddadi:2016:NAM

[HF16] Bahareh Haddadi and Morteza Fathipour. Numerical analysis of 3D model of the SSAW separator system. International Journal of Computer Applications, 141(??):7–12, May 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume141/number12/24834-
2016909914/.

Hartono:2020:DMA

[HF20] Hartono and Devi Fitrianah. A data mining approach for product promotion and inventory solution using FPgrowth algorithm. International Journal of Computer Applications, 177(??):37–44, February 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number38/31157-2020919889/.

Htun:2012:ITM

[HFSM12] Ohnmar Htun, Andrew Finch, Eiichiro Sumita, and Yoshiki Mikami. Improving transliteration mining by integrating expert knowledge with statistical approaches. International Journal of Computer Applications, 58(??):12– 22, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number17/9373-3821/.

Hareshkumar:2011:RWS

[HG11] Navadiya Hareshkumar and Deepak Garg. Random Web surfer PageRank algorithm. International Journal of Computer Applications, 35(??):36–41, December 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume35/number11/4448-6214/.

Hasan:2013:DUE

[HG13] Zahid Hasan and Rathindra Chandra Gope. Dynamics of user experience (UX). International Journal of Computer Applications, 81(??):18–24, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume81/number16/14207-2443/.

Hazarika:2015:EVV

[HGK15] G. C. Hazarika, Kalpana Goswami, and Jadav Konch. Effects of variable viscosity and thermal conductivity on MHD flow past a vertically moving porous plate with viscous and Joule dissipation. International Journal of Computer Applications, 123(??):40–48, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number9/21990-2015905631/.

Hiremath:2013:FRU

[HH13] P. S. Hiremath and Manjunath Hiremath. 3D face recognition using Radon transform and symbolic LDA. International Journal of Computer Applications, 67(??): 1–4, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number4/11380-6666/.

Haruna:2019:TMD

[HHEY19] Charles R. Haruna, MengShu Hou, and Barbie Eghan-Yartel. Theoretical model for detecting sensitive data items of users in data publication. International Journal of Computer Applications, 178(??):1–8, August 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number37/30774-2019919239/.

Hasan:2021:ASR

[HHHK21] Syed H. Hasan, Syeda Huyam Hasan, Syed Hamid Hasan, and Salman Khalid. Assessing search and rescue optimization based DNN model for streamflow data prediction. International Journal of Computer Applications, 183(??):11–16, August 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number19/32031-2021921493/.

Heidarinia:2014:IAM

[HHS14] Neda Heidarinia, Ali Harounabadi, and Mehdi Sadeghzadeh. An intelligent anti-money laundering method for detecting risky users in the banking systems. International Journal of Computer Applications, 97(??):35–39, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume97/number22/17141-7780/.

Hwang:2015:IVE

[HHS15] Wu-Yuin Hwang, Jung-Lung Hsu, and Siao-Han Syu. The impact of virtual experience and informative control on the enhancement of product understanding. International Journal of Computer Applications, 111(??):16–21, February 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume111/number10/19574-

Haridas:2014:MLF

[HIE14] Nisha Haridas, Aravind Illa, and Elizabeth Elias. Multiplierless farrow structure based linear phase low pass interpolation filter. International Journal of Computer Applications, 95(??):1–6, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number14/16659-6645/.

Hassan:2015:NSS

[HII15] A. A. M. Hassan, S. A. Hoda Ibrahim, and Mohamed G. M. Ibrahim. Numerical solution of a system SEIR nonlinear ODEs by Runge–Kutta fourth order method. International Journal of Computer Applications, 124(??):1–6, August 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume124/number3/22081-2015903852/

.

Hikal:2013:CES

[Hik13] Noha A. Hikal. A comparative evaluation study of automated gait recognition based on spatiotemporal feature and different neural network classifiers. International Journal of Computer Applications, 68(??):36–42, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number9/11610-6991/.

Hire:2012:SWD

[Hir12] Dnyanda Namdeo Hire. Secured wireless data communication. International Journal of Computer Applications, 54 (??):27–30, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume54/
number1/8532-2064/.

Hoque:2015:TTC

[HIS15] Fazlul Hoque, Md. Jahidul Islam, and Swakkhar Shatabda. A two-tier classification model for financial fraud detection. International Journal of Computer Applications, 118(??):1–8, May 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume118/number19/
20850-3586/.

Hossain:2020:TBM

[HITA20] Mohammad Asif Hossain, Md. Imdadul Islam, Fahima Tabassum, and M. R. Amin. The tradeoff between mean delay and energy saving factor under DRX scheme. International Journal of Computer Applications, 177(??):40–47, March 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume177/number43/
31195-2020919946/.

Hamza:2011:UHT

[HIU11] M. M. Hamza, B. Y. Isah, and H. Usman. Unsteady heat transfer to MHD oscillatory flow through a porous medium under slip condition. International Journal of Computer Applications, 33(??):12–17, November 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume33/number4/4007-5679/.

Hussain:2013:NCHb

[HJ13] Ihsan Y. Hussain and Nasr A. Jabbar. Natural convection heat transfer in cylindrical enclosure of porous media with periodic boundary conditions. International Journal of Computer Applications, 65(??):43–50, March 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume65/number9/10956-5918/.

Hanamsagar:2014:FAM

[HJB^{+}14] Ameya Hanamsagar, Ninad Jane, Bhagyashree Borate, Aditi Wasvand, and S. A. Darade. Firewall anomaly management: a survey. International Journal of Computer Applications, 105(??):1–5, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number18/18474-9363/.

H:2011:ABI

[HJBJ11] Saima H, J. Jaafar, S. Belhaouari, and T. A. Jillani. ARIMA based interval Type-2 fuzzy model for forecasting. International Journal of Computer Applications, 28 (??):17–21, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume28/
number3/3369-4652/.

Hadriche:2014:DMP

[HJE14] Abir Hadriche, Nawel Jmail, and Ridha Elleuch. Different methods of partitioning the phase space of a dynamic system. International Journal of Computer Applications, 93(??):1–5, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number15/16288-5931/.

Hadriche:2013:RCN

[HJG^{+}13] Abir Hadriche, Nawel Jmail, Hamadi Ghariani, Abdennaceur Kachouri, and Laurent Pezard. Reduction of continuous neuronal model to discrete binary automata. International Journal of Computer Applications, 69(??): 5–10, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number24/12116-8140/.

Hanini:2014:TMA

[HJT14] Muna Hanini, Nidal A. M. Jabari, and Radwan Tahboub. Text modeling in adaptive educational chat room. International Journal of Computer Applications, 103(??): 33–37, October 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume103/number5/18073-9022/.

Hati:2010:AIU

[HK10] Debashis Hati and Amritesh Kumar. An approach for identifying URLs based on division score and link score in focused crawler. International Journal of Computer Applications, 2(??):48–53, May 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume2/number3/643-899/.

Hocine:2011:ESM

[HK11] Riadh Hocine and Hamoudi Kalla. Encoding SystemC models in formal synchronous formalism. International Journal of Computer Applications, 34(??):26–32, November 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume34/number3/4080-
5876/.

Huq:2015:RBS

[HK15] A. M. Anisul Huq and Mosharraf Hossain Khan. A reputation based system to overcome malicious behavior in peerto-peer networks. International Journal of Computer Applications, 128(??):1–6, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number6/22874-2015906545/.

H:2018:CIC

[HK18] Kondekar Vipul H and Bodhe S K. A comprehensive investigation of color models used in image processing. International Journal of Computer Applications, 180(??):19–24, February 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number22/29065-2018916507/.

Hasan:2019:SDM

[HK19] Syed Hamid Hasan and Usman Ali Khan. Software development methods — properties and advances. International Journal of Computer Applications, 178(??):26–32, September 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume178/number53/30913-
2019919436/.

Halhalli:2015:BSB

[HKA15] Suresh R. Halhalli, Subhash Kulkarni, and K. S. R. Anjaneyulu. BW and SS based handover analysis of four states of mobile node in a five node network model. International Journal of Computer Applications, 118(??):30–35, May 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume118/number8/20767-3215/

.

Hammad:2014:DCM

[HKE^{+}14] M. E. Hammad, H. Kasban, Sayed M. Elaraby, Moawad I. Dessouky, O. Zahran, and Fathi E. Abd El-Samie. Distillation column malfunctions identification using higher order statistics. International Journal of Computer Applications, 108(??):7–13, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number1/18873-0129/.

Husain:2012:DIW

[HKF12] Farooq Husain, Ekram Khan, and Omar Farooq. Digital image watermarking using combined DWT and DFRFT. International Journal of Computer Applications, 60(??):17– 25, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number11/9736-4287/.

Ke:2013:DIP

[hKhHL13] Chih horng Ke, Chao hsien Hsieh, and Chiang Lee. Design and implementation of a portable name card typesetting system. International Journal of Computer Applications, 79(??):35–42, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume79/number3/13723-1519/.

Hsieh:2016:ILF

[HKL16] Chao-Hsien Hsieh, Chih-Horng Ke, and Chiang Lee. IPS: a lightweight framework for cross-platform multimedia streaming server. International Journal of Computer Applications, 139(??):1–8, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number7/24499-2016909176/.

Hashmi:2017:FOB

[HKR17] Md. Farukh Hashmi, M. Pramod Kumar, and K. S. Rao. A framework for optimization of the boot time on embedded Linux environment with Raspberry Pi platform. International Journal of Computer Applications, 162(??):17–26, March 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume162/number3/27222-2017913304/.

Hachichi:2012:TDD

[HKS12] H. Hachichi, I. Kitouni, and D. E. Saidouni. Transforming DATA* with dotty format to aggregate region automaton. International Journal of Computer Applications, 37 (??):35–42, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume37/
number10/4647-6721/.

Habib:2019:ECT

[HKU19] A. Habib, Waseem Ullah Khan, and Tahir Ullah. An efficient cross-tier interference mitigation technique in LTE femtocell environment. International Journal of Computer Applications, 178(??):37–43, September 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number49/30887-2019919424/.

Hossain:2015:HDI

[HKY15] Ferdous Hossain, Mithun Kumar P K, and Mohammad Abu Yousuf. Hardware design and implementation of adaptive Canny edge detection algorithm. International Journal of Computer Applications, 124(??): 31–38, August 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume124/number9/22133-2015905446/.

Hu:2024:RWG

[HLZ^{+}24] Junjun Hu, Xiaoyan Li, Yongkuo Zhang, Xiajie Ai, and Lei Chen. The research on word game based on SIRS-ARIMA model and machine learning algorithm. International Journal of Computer Applications, 186(??):26–36, April 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number15/
the-research-on-word-game-based-on-sirs-arima-model-
andmachine-learning-algorithm/.

Hacene:2013:ANO

[HM13] Nacer Hacene and Boubekeur Mendil. Autonomous navigation and obstacle avoidance for a wheeled mobile robots: a hybrid approach. International Journal of Computer Applications, 81(??):34–37, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume81/number7/14027-2285/.

Hein:2018:PBF

[HM18] Chit La Pyae Myo Hein and Khin Mar Myo. Permissionbased feature selection for Android malware detection and analysis. International Journal of Computer Applications, 181(??):29–39, September 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number19/29974-2018917902/.

Hussein:2013:NSS

[HMAA13] M. H. Hussein, H. Moselhy, S. Aly, and M. E. Awad. A new strategy to synthesis an optimum controllable HEN by using fuzzy analogical gates. International Journal of Computer Applications, 83(??):41–55, December 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume83/number3/14432-2578/.

Hassouna:2013:SMB

[HMB13] Mohammed Hassouna, Nashwa Mohamed, and Eihab Bashier. A secure mobile banking scheme based on certificateless cryptography in the standard security model. International Journal of Computer Applications, 74(??): 1–6, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number9/12910-9890/.

Hameurlaine:2012:ANH

[HMC12a] Messaoud Hameurlaine, Abdelouahab Moussaoui, and Hadda Cherroun. AntMeans: a new hybrid algorithm based on ant colonies for complex data mining. International Journal of Computer Applications, 60(??):6– 12, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number17/9782-4314/.

Housni:2012:MOT

[HMC12b] Khalid Housni, Driss Mammass, and Youssef Chahir. Moving objects tracking in video by graph cuts and parameter motion model. International Journal of Computer Applications, 40(??):20–27, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume40/number10/5001-7283/.

Hanmandlu:2012:CSP

[HMGV12] Madasu Hanmandlu, Neha Mittal, Ankit Gureja, and Ritu Vijay. A comprehensive study of palmprint based authentication. International Journal of Computer Applications, 37 (??):17–24, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume37/
number2/4580-6499/.

Hussny:2013:LCI

[HMMN13] Hasnat Ahmad Hussny, Ahmed Mateen, Tasleem Mustafa, and Muhammad Murtaza Nayyer. E-learner’s collective intelligent system framework: Web mining for personalization in E-Learning 2.0 ecosystem using Web 2.0 technologies. International Journal of Computer Applications, 66 (??):1–9, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number4/11069-5987/.

Haque:2018:APF

[HMN18] Md Jayedul Haque, Khairul Mottakin, and Nazmun Nahar. Automatic pattern forecasting from banking financial data. International Journal of Computer Applications, 182(??):13–19, August 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number8/29839-2018917632/.

Halidou:2017:RSB

[HMO17] Aminou Halidou, Youssoufa Mohamadou, and Georges Delort Olle O. Robust steganography based on matching pixel locations. International Journal of Computer Applications, 168(??):3–10, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume168/number12/27925-2017914252/.

Hamid:2012:PAC

[HMY12] Thaier Hamid, Carsten Maple, and Yong Yue. A parallel algorithm to calculate the costrank of a network. International Journal of Computer Applications, 44(??): 17–22, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number3/6243-8161/.

Hidayat:2016:NFD

[HN16] Nurul Hidayat and Azhari S N. New formulation of dynamic collaborative learning to effectuate of Indonesia Integrated and Open Online Learning (PDITT) programme. International Journal of Computer Applications, 137(??):22–26, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume137/number1/24240-2016908650/.

Hakim:2017:EQO

[HNW17] Luqman Hakim, Suryanto Nugroho, and Sigit Hadi Waluyo. Experiment of query optimization techniques to get the efficient query. International Journal of Computer Applications, 170(??):6–10, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume170/number10/28104-2017914888/.

Hadi:2015:PAP

[HO15a] Zainab Saad Hadi and Buthaina Mosa Omran. Peak-to-average power reduction using repeated frequency domain filtering and clipping in OFDM. International Journal of Computer Applications, 122(??):6–10, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume122/number11/21742-4958/.

Haithem:2015:AFR

[HO15b] Hermessi Haithem and Mourali Olfa. Aided 3D facial restoration for plastic surgery planning using elastic radial curves. International Journal of Computer Applications, 127(??):7–11, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number7/22739-2015903074/.

Houda:2015:BAS

[HO15c] Abouzid Houda and Chakkor Otman. Blind audio source separation: State-of-art. International Journal of Computer Applications, 130(??):1–6, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number4/23194-2015906491/.

Hossein:2012:CEA

[Hos12] Md. Syed Mahamud Hossein. A compression and encryption algorithms on DNA sequences using R^{2}CP and modified Huffman technique. International Journal of Computer Applications, 57(??):1–10, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume57/number1/9075-6994/.

Harekal:2013:MMP

[HPC13] Divakar Harekal, Vijaykumar B P, and R. Chandrasekhar. mHealth mobile phone based patient compliance system. International Journal of Computer Applications, 79(??): 24–29, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number8/13762-1602/.

Henry:2010:CCB

[HPM10] Rabinder Henry, Bhushan Patil, and Praveen Malav. Compact CPLD board designing and implemented for digital clock. International Journal of Computer Applications, 3(??):7–10, July 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume3/number11/783-1108/.

Hettiarachchi:2020:TBI

[HPM20] Nethmi Hettiarachchi, Indrachapa Pathiraja, and Dilum Madawala. Translation of Brahmi inscriptions to Sinhala using natural language processing. International Journal of Computer Applications, 177(??):19–25, January 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume177/number30/31090-
2020919776/.

Hikmatyar:2017:NFF

[HPR17] Missi Hikmatyar, Yudi Prayudi, and Imam Riadi. Network forensics framework development using interactive planning approach. International Journal of Computer Applications, 161(??):41–48, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number10/27188-2017913352/.

Hussain:2013:NCHa

[HR13] Ihsan Y. Hussain and Basim K. Raheem. Natural convection heat transfer from a plane wall to thermally stratified porous media. International Journal of Computer Applications, 65(??):33–41, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume65/number1/10891-5792/.

Hojati:2014:EED

[HR14a] Marziyeh Hojati and Mehdi Rezaeian. Empirical evaluation of dissimilarity measures for content-based image retrieval. International Journal of Computer Applications, 105(??):1–7, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number11/18418-9573/.

Hunaish:2014:PBI

[HR14b] Ahmed Sadeq Hunaish and Jawad Radhi. PLC based implementation of fuzzy controller for boost converter. International Journal of Computer Applications, 97(??): 36–44, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number11/17055-7359/.

Habbi:2016:PFS

[HR16] Hanan Mikhael D. Habbi and Sara T. Rashed. PI and fuzzy speed controllers for PM synchronous motor drive. International Journal of Computer Applications, 149(??):1–9, September 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume149/number1/25958-2016911222/.

Hidayat:2021:OWS

[HR21] Taufik Nur Hidayat and Imam Riadi. Optimation wireless security IEEE 802.1X using the extensible authentication protocol-protected extensible authentication protocol (EAP-PEAP). International Journal of Computer Applications, 174(??):25–30, January 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number11/31722-2021920988/.

Hidayat:2022:DBC

[HR22] Ahmad Syarif Hidayat and Wahyu Kusuma Raharja. Design and build a camera directional movement control device based on sound source position detection. International Journal of Computer Applications, 183(??):40–47, February 2022. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number51/32276-
2022921931/.

Hrisheekesha:2010:EAB

[HS10] P. N. Hrisheekesha and Jaydev Sharma. Evolutionary algorithm based optimal control in distribution system with dispersed generation. International Journal of Computer Applications, 1(??):31–37, February 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume1/number14/305-471/.

Hafizah:2011:CEU

[HS11] Wan Mahani Hafizah and Eko Supriyanto. Comparative evaluation of ultrasound kidney image enhancement techniques. International Journal of Computer Applications, 21 (??):15–19, May 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume21/
number7/2524-3432/.

Hasan:2012:PAF

[HS12] Md. Mahmudul Hasan and S. S. Singh. PAPR analysis of FFT and wavelet based OFDM systems for wireless communications. International Journal of Computer Applications, 60(??):38–42, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume60/number15/9771-4425/.

Hazarika:2014:BFT

[HS14] G. C. Hazarika and Aparajita Sarmah. Blood flow through a stenosed artery with the effect of transverse magnetic field using a non-Newtonian model. International Journal of Computer Applications, 101(??):5–8, September 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume101/number1/17649-8435/.

Hato:2015:LRC

[HS15a] Eman Hato and Dalya Shihab. Lorenz and Rossler chaotic system for speech signal encryption. International Journal of Computer Applications, 128(??):25–33, October 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume128/number11/22919-
2015906670/.

Henna:2015:LCV

[HS15b] Shagufta Henna and Muhammad Bilal Saleemi. Least channel variation multi-channel MAC (LCV-MMAC) under mobile random topologies in mobile adhoc networks. International Journal of Computer Applications, 131(??): 1–6, December 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume131/number4/23434-2015907207/.

Hussain:2015:CHS

[HS15c] Hashiroh Hussain and Norshuhada Shiratuddin. Components of heuristic strategy: Evaluation of stories with DST tool on touch screen tablet. International Journal of Computer Applications, 130(??):6–10, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number11/23251-2015907106/.

Hussein:2015:CLE

[HS15d] Abla Hussein and Ghassan Samara. Coordinator location effects in AODV routing protocol in ZigBee mesh network. International Journal of Computer Applications, 127(??):1–7, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number8/22746-2015906373/.

Hetharie:2018:ICR

[HS18] Renee Yosua Hetharie and Eko Sediyono. Implementation of customer relationship management (CRM) to increase customer loyalty and product sales volume at Alexander Stamps. International Journal of Computer Applications, 181(??):7–11, November 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number26/30098-2018918066/.

Hirwono:2024:CIS

[HS24] Bowo Hirwono and Suhirman. Classification of Indonesian smart card scholarship recipients with principal component analysis using the naive Bayes and decision tree methods case study: Stie Pariwisata API Yogyakarta. International Journal of Computer Applications, 186(??): 13–21, January 2024. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume186/number5/33068-2024923375/.

Hussain:2013:DHL

[HSD13] S. Javeed Hussain, T. Satya Savitri, and P. V. Sree Devi. Detection of hydrocephalus lateral ventricles quantitatively in brain MRI images of infants. International Journal of Computer Applications, 83(??):12–15, December 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume83/number9/14475-2698/.

Hassan:2015:NSW

[HSEA15] Mohammed A. Hassan, Nancy M. Salem, and Mohamed I. El-Adawy. A new shear wave speed estimation method for shear wave elasticity imaging. International Journal of Computer Applications, 125(??):48–53, September 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume125/number8/22456-2015906130/.

Helal:2017:HAA

[HSH17] Alaa Eldeen M. Helal, Ahmed Farag Seddik, and Ayat Allah F. Hussein. A hybrid approach for artifacts removal from EEG recordings. International Journal of Computer Applications, 168(??):10–19, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume168/number4/27861-2017914301/.

Hndoosh:2014:FMM

[HSK14] Rana Waleed Hndoosh, M. S. Saroa, and Sanjeev Kumar. Fuzzy mathematical models of Type-1 and Type-2 for computing the parameters and its applications. International Journal of Computer Applications, 104(??):17–28, October 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume104/number14/18270-
9333/.

Hadian:2012:FGP

[HSMb12] Ali Hadian, Saeed Shahrivari, and Behrouz Minaei-bidgoli. Fine-grained parallel ant colony system for shared-memory architectures. International Journal of Computer Applications, 53(??):8–13, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume53/number8/8439-2223/.

Hegadi:2010:LEH

[HSP10] Rajendra Hegadi, Ramesh Shahabadkar, and Shamshekhar Patil. Lightweight encryption of hierarchical access control mechanism key using tribes of Farey fractions and its analysis. International Journal of Computer Applications, 6(??):23–28, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume6/number3/1060-1385/.

Hussain:2011:SPS

[HSS11] Malik Zawwar Hussain, Muhammad Sarfraz, and Ayesha Shakeel. Shape preserving surfaces for the visualization of positive and convex data using rational bi-quadratic splines. International Journal of Computer Applications, 27 (??):12–20, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume27/
number10/3338-4594/.

Honeyily:2019:CBE

[HST19] Honeyily, Disha Sugha, and Vinay Thakur. Conditional based edge detection algorithm enhancing the quality of clustered grain seeds. International Journal of Computer Applications, 178(??):12–15, July 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number33/30750-2019919203/.

Hassainia:2014:MCI

[HTM14] S. Hassainia, R. Toufouti, and S. Meziane. Modeling and control of induction motor using causal informational graph. International Journal of Computer Applications, 104(??):34–38, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number1/18170-9051/.

Hiware:2011:DHM

[HTPB11] Rahul K. Hiware, Abhijit S. Titarmare, Dinesh Padole, and Preeti Bajaj. Design of homogeneous multi-core based vehicular embedded controller. International Journal of Computer Applications, 28(??):40–47, August 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume28/number5/3380-4687/.

Hussein:2012:DDQ

[Hus12] Mohammad Hussein. Decentralized dynamic query optimization based on mobiles agents for large scale data integration systems. International Journal of Computer Applications, 60(??):7–14, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume60/number8/9710-4172/.

Hussein:2015:EID

[Hus15] Sarmad A. Abdal Hussein. Experimental investigation of double pipe heat exchanger by using semi circular disc baffles. International Journal of Computer Applications, 115(??):13–17, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number4/20138-2237/.

Husain:2016:MVC

[Hus16] Anas Jebreen Atyeeh Husain. On minimum variance CPUscheduling algorithm for interactive systems using goal programming. International Journal of Computer Applications, 135(??):51–59, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number11/24097-2016908550/.

Herath:2018:MPT

[HW18] H. M. L. N. K. Herath and J. V. Wijayakulasooriya. Modeling of phoneme transitions for natural synthesis of speech signals. International Journal of Computer Applications, 181(??):22–29, October 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number23/30026-2018918008/.

Harahap:2013:MRB

[HWT^{+}13] Erwin Harahap, Janaka Wijekoon, Rajitha Tennekoon, Fumito Yamaguchi, Shinichi Ishida, and Hiroaki Nishi. Modeling of router-based request redirection for content distribution network. International Journal of Computer Applications, 76(??):37–46, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume76/number13/13310-0857/.

Hussain:2013:IFC

[HY13] Ihsan Y. Hussain and Aya A. Yaseen. An investigation into forced convection heat transfer through porous media. International Journal of Computer Applications, 66 (??):15–22, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number1/11048-5940/.

Hamdy:2012:RTW

[HYA12] Abeer Hamdy, Ahmed E. Youssef, and Reda Ammar. Real-time workload allocation on a uni-processor. International Journal of Computer Applications, 53(??):17– 24, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number12/8473-2401/.

Hamidzadeh:2014:ACC

[HZ14] S. M. Hamidzadeh and Arman Zarringhalam. Attitude control of chaotic satellite with unknown input and uncertainties based on sliding control. International Journal of Computer Applications, 97(??):32–36, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume97/number3/16990-7092/.

Hassan:2015:OTV

[HZ15] A. B. M. Khalid Hassan and Ebad Zahir. Optimizing the turning velocity in a line follower robot. International Journal of Computer Applications, 117(??):14–17, May 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume117/number3/20534-2878/

.

Hassan:2013:SDR

[HZBG13] Sara M. Hassan, A. Zekry, M. A. Bayomy, and G. Gomah. Software defined radio implementation of LTE transmitter physical layer. International Journal of Computer Applications, 74(??):41–46, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume74/number8/12909-0065/.

Hammi:2012:FTM

[HZK12] Yosra Hammi, Nadia Zanzouri, and Mekki Ksouri. Faulttolerant model predictive control of hybrid systems. International Journal of Computer Applications, 43(??):29– 35, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume43/
number10/6142-8387/.

Hu:2015:CMT

[HZLL15] Yingjie Hu, Jingzhi Zhang, Yuchen Liang, and Zhuohao Li. Configuration monitoring technology based PLC practical teaching device. International Journal of Computer Applications, 126(??):13–18, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume126/number8/22571-2015906133/.

Harizi:2014:TAW

[HZS14] Mehdi Harizi, Mehdi Mohtasham Zadeh, and Aref Saiahi. Traffic analysis, a wise approach to detect compromised nodes in wireless sensor networks. International Journal of Computer Applications, 97(??):6–9, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume97/number14/17073-7511/.

Ismanto:2015:MAD

[IA15] Heru Ismanto and Azhari. Multi attribute decision making method for determination of regional development plan priority in Papua Province based on the indicator of Gross Regional Domestic Product (GRDP). International Journal of Computer Applications, 117(??):21–26, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume117/number7/20567-2957/.

Inusah:2018:ESA

[IA18] Fuseini Inusah and Anokye Acheampong Amponsah. An expert system to assist businesses in financial decision making to enhance efficiency. International Journal of Computer Applications, 181(??):32–39, August 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number7/29787-2018917592/.

Issa:2013:VAF

[IAA13] Sanae Khali Issa, Abdellah Azmani, and Benaissa Amami. Vulnerability analysis of fire spreading in a building using fuzzy logic and its integration in a decision support system. International Journal of Computer Applications, 76 (??):48–53, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume76/
number6/13255-0732/.

Ibrahim:2014:FBT

[IAH14] Tharwat Ibrahim, Gamal Attiya, and Ahmed Hamad. Fuzzy based tuning congestion window for improving end-to-end congestion control protocols. International Journal of Computer Applications, 87(??):1–8, February 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume87/number1/15169-1385/.

Ismail:2016:HSD

[IAM16] Esraa M. Ismail, Isam M. Abdulbaq, and Ammar I. Majeed. High step up DC-DC converter fed from photovoltaic system. International Journal of Computer Applications, 139(??):6–13, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number3/24468-2016909427/.

Iqbal:2011:CAP

[IAQ11] M. Aqeel Iqbal, Farooque Azam, and Ahmed Ali Qureshi. Critical aspects of project management causing major impacts on software development. International Journal of Computer Applications, 27(??):39–47, August 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume27/number9/3326-4576/.

Ibraheem:2014:DLF

[Ibr14] Ibraheem K. Ibraheem. Damping low frequency oscillations in power system using quadratic Gaussian technique based control system design. International Journal of Computer Applications, 92(??):18–23, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume92/number11/16052-5204/.

Ibrahim:2016:STC

[Ibr16] Dogan Ibrahim. Smart train collision detection system using a microcontroller. International Journal of Computer Applications, 152(??):16–22, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume152/number2/26291-2016911771/.

Ibraheem:2017:SIT

[Ibr17] Ibraheem Kasim Ibraheem. System identification of thermal process using Elman neural networks with no prior knowledge of system dynamics. International Journal of Computer Applications, 161(??):38–46, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number11/27194-2017913316/.

Ibraheem:2018:AIF

[Ibr18] Ibraheem Kasim Ibraheem. Adaptive IIR and FIR filtering using evolutionary LMS algorithm in view of system identification. International Journal of Computer Applications, 182(??):31–39, August 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number11/29867-2018917740/.

IbnElAhrache:2016:PDS

[IBS16] Sara Ibn El Ahrache, Hassan Badir, and Abderrahmane Sbihi. Persistent data security in public clouds. International Journal of Computer Applications, 155(??):

31–36, December 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume155/number9/26635-2016912430/.

Ibrahim:2012:ECA

[IE12] Lamiaa Fattouh Ibrahim and Manal El Harby. Enhancing clustering algorithm to plan efficient mobile network. International Journal of Computer Applications, 59(??):18– 24, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number18/9648-4401/.

Ibrahim:2016:IRM

[IEES16] Shaimaa ElSayed Ibrahim, K. A. ElBarbary, and Ragab M. El-Sagheer. ICI reduction in multicarrier systems. International Journal of Computer Applications, 133(??):18–26, January 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume133/number13/
23847-2016908116/.

Ibrahim:2014:SSP

[IEG14] Mohamed Ismail Ibrahim, Khairy A. Elbarbary, and Abdelhamid Gaafar. Space signal processing for helicopter satellite communication using 2 × 1 Alamouti scheme. International Journal of Computer Applications, 87(??):1– 5, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number8/15225-3740/.

Ibtissam:2018:EBS

[IEL18] Zaaj Ibtissam, Chaouki Brahim El Khalil, and Masmoudi Lhoussaine. Extraction of building in satellite image THR using features detection. International Journal of Computer Applications, 181(??):23–27, August 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number10/29809-2018917649/.

Iggane:2012:STU

[IEME12] M’bark Iggane, Abdelatif Ennaji, Driss Mammass, and Mostafa El Yassa. Self-training using a K-nearest neighbor as a base classifier reinforced by support vector machines. International Journal of Computer Applications, 56 (??):43–46, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number6/8899-2925/.

Ismail:2014:UBC

[IFT14] R. E. S. Ismail, A. S. Fahmy, and N. M. Tawfik. Ultimate behavior of composite castellated beams under vertical loads. International Journal of Computer Applications, 108(??):40–46, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number5/18911-0214/.

Innocent:2017:CAR

[IGS17] Fale Mantim Innocent, Abdulsalam Ya’U Gital, and Iliya Nengak Sitlong. Course advisory and results expert system (CARES): an implementation of FMI course autoscheduling algorithm. International Journal of Computer Applications, 179(??):6–13, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number5/28730-2017915936/.

Islam:2015:ADT

[IHB^{+}15] Mohammad Rakibul Islam, Rifad Hossain, Md. Ziaul Haque Bhuiyan, Tahmeed Ahmed Margoob, Md. Taslim Reza, and Kazi Khairul Islam. Arrhythmia detection technique using basic ECG parameters. International Journal of Computer Applications, 119(??):11–15, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number10/21102-3819/.

Ibraheem:2012:TSR

[IHD12] Basem Ibraheem, Abeer Hamdy, and Nevien Darwish. Textual signs reading for indoor semantic map construction. International Journal of Computer Applications, 53(??):36– 43, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number10/8459-2268/.

Imen:2011:BCD

[IHH11] Haj Brahim Imen, Soufien Hajji, and Soufien Hajji. Backstepping controller design using a high gain observer for induction motor. International Journal of Computer Applications, 23(??):1–6, June 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume23/number3/2873-3730/.

Ibrahim:2014:CVB

[IHH14] Mennatallah H. Ibrahim, Hesham A. Hefny, and Nermin Hamza. Contextual view-based access control model for spatial data on web. International Journal of Computer Applications, 87(??):38–42, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume87/number15/15288-4046/.

Ilkin:2017:CGH

[İHŞ17] Sümeyya İlkin, Fatma Selin Hangişi, and Suhap Şahin. Comparison of global histogram-based thresholding methods that applied on wound images. International Journal of Computer Applications, 165(??):23–28, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number9/27603-2017914002/.

Ismael:2012:DSI

[IhZ12] Mohammad Abdulmoniem Ali Ismael and Abd El haleem Zekry. Design, simulation and implementation of the Ethernet simplified MAC together with the P2MP emulation layer for EPON ONU devices. International Journal of Computer Applications, 55(??):57–64, October 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume55/number10/8795-2792/.

Ingle:2016:ARS

[II16] Vikal R. Ingle and V. T. Ingole. Absolute and relative scaling of power transformers using grey incidence and relational analysis. International Journal of Computer Applications, 150(??):36–42, September 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume150/number1/26060-2016911453/.

Injeti:2011:OPD

[IK11] S. Kumar Injeti and Navuri P Kumar. Optimal planning of distributed generation for improved voltage stability and loss reduction. International Journal of Computer Applications, 15(??):40–46, February 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume15/number1/1910-2545/.

Irianto:2022:ICP

[IKY22] Suhendro Y. Irianto, Sri Karnila, and Dona Yuliawati. Improvement CBIR performance of region-based segmentation on DCT images. International Journal of Computer Applications, 184(??):24–29, July 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number20/32434-2022922220/.

Iranian:2015:UMN

[ILG15] D. Iranian, P. Loganathan, and P. Ganesan. Unsteady MHD natural convective flow over vertical plate in thermally stratified media with variable viscosity and thermal conductivity. International Journal of Computer Applications, 121(??):18–24, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number3/21520-4498/.

Idarrou:2012:SCM

[IM12] Ali Idarrou and Driss Mammass. Structural clustering multimedia documents: an approach based on semantic sub-graph isomorphism. International Journal of Computer Applications, 51(??):14–21, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume51/number1/8005-1343/.

Islam:2015:PPS

[IM15] M. P. Islam and T. Morimoto. Performance prediction of solar collector adsorber tube temperature using a nonlinear autoregressive model with eXogenous input. International Journal of Computer Applications, 114(??):24–32, March 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume114/number12/20031-
2129/.

Ivan:2016:TRT

[IM16] Cosmina Ivan and Andrei Moldovan. Twitrends: a real time trending topics detection system for Twitter social network. International Journal of Computer Applications, 152(??):16–25, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume152/number4/26307-2016911785/.

Islam:2011:EAR

[IMIA11] Md. Shohidul Islam, Md. Niaz Morshed, SK. Shariful Islam, and Md. Mejbahul Azam. An experimental analysis of random early discard (RED) queue for congestion control. International Journal of Computer Applications, 15 (??):18–21, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume15/
number2/1920-2563/.

Ismail:2012:CMT

[IMM12] I. A. Ismail, G. S. Mokaddis, and Mariam K. Metry. Computing a matrix transpose of multithreading for queuing parallel in Matlab programming. International Journal of Computer Applications, 49(??):41–47, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume49/number5/7627-0694/.

Ismaeel:2015:NCB

[IN15] Tarik Zeyad Ismaeel and Ahmed Saad Names. A new cryptosystem based on fingerprint features. International Journal of Computer Applications, 132(??):31–36, December 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume132/number12/23649-
2015907662/.

Isong:2013:NAP

[IND13] Bassey Isong, Ndwamato Netshipise, and Gilbert Dzawo. A new approach for protecting elections integrity in South Africa. International Journal of Computer Applications, 62 (??):10–18, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number20/10210-5042/.

Idiaghe:2020:DWU

[IOO20] E. O. Idiaghe, R. I. Osaze, and E. Osazee. Design of a wirelessly-updated digital clock with Android interface. International Journal of Computer Applications, 175(??):6–10, August 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume175/number15/
31527-2020920641/.

IbnAbdulrahman:2014:FLA

[IPB14] Umar Farouk Ibn Abdulrahman, Joseph Kobina Panford, and James Ben Hayfron-acquah. Fuzzy logic approach to credit scoring for micro finance in Ghana: a case study of KWIQPLUS money lending. International Journal of Computer Applications, 94(??):11–18, May 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume94/number8/16362-5772/.

Irfan:2015:IEU

[IPR15] Pahrul Irfan, Yudi Prayudi, and Imam Riadi. Image encryption using combination of chaotic system and Rivest Shamir Adleman (RSA). International Journal of Computer Applications, 123(??):11–16, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number6/21962-2015905344/.

Islam:2010:HFD

[IR10] Md. Rabiul Islam and Md. Fayzur Rahman. Hybrid feature and decision fusion based audio-visual speaker identification in challenging environment. International Journal of Computer Applications, 9(??):9–15, November 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume9/number5/1384-1864/.

Irawan:2016:DSS

[IR16] Agus Irawan and Rochmawati. Decision support system for alternative site selection of Kkn using Topsis. International Journal of Computer Applications, 150(??): 1–6, September 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume150/number7/26102-2016911538/.

Irani:2012:DPN

[Ira12] Gholamali Nejad Hajali Irani. Decentralized principles: New modular software development principles, a robust object oriented approach. International Journal of Computer Applications, 44(??):26–33, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume44/number13/6324-8671/.

Islam:2011:PAW

[IRT11] Shaffatul Islam, Mamunur Rashid, and Mohammed Tarique. Performance analysis of WiMax/WiFi system under different codecs. International Journal of Computer Applications, 18(??):13–19, March 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume18/number6/2290-2973/.

Indiarto:2021:AEP

[IS21] Mahdif Indiarto and Hartono Siswono. Analysis of equalizer performance against bit error rate in filter bank multicarrier system with AWGN and multipath channel. International Journal of Computer Applications, 183(??):21–25, July 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume183/number15/
32003-2021921476/.

Ishaq:2012:SMU

[Ish12a] Zeba Ishaq. SECURE MANET using two head cluster in hierarchal cooperative IDS. International Journal of Computer Applications, 57(??):10–13, November 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume57/number3/9093-3126/.

Islam:2012:EAS

[ISH12b] Md. Maidul Islam, Md. Rowshan Sahriar, and Md. Belal Hossain. An enhanced automatic surface and structural flaw inspection and categorization using image processing both for flat and textured ceramic tiles. International Journal of Computer Applications, 48(??):1–10, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume48/number3/7325-9159/.

Ishengoma:2014:ADH

[Ish14] Fredrick R. Ishengoma. The art of data hiding with Reed–Solomon error correcting codes. International Journal of Computer Applications, 106(??):28–31, November 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume106/number14/18590-

Islam:2018:NSR

[Isl18] M. Babul Islam. Noisy speech recognition by mel-LPC based AR-HMM with power and time derivative parameters. International Journal of Computer Applications, 180(??):1–5, May 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number42/
29408-2018917149/.

Islam:2017:CST

[ISM^{+}17] Md Rafiqul Islam, Naznin Sultana, Mohammad Ali Moni, Prohollad Chandra Sarkar, and Bushra Rahman. A comprehensive survey of time series anomaly detection in online social network data. International Journal of Computer Applications, 180(??):13–22, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number3/28780-2017915989/.

Issa:2013:OEC

[Iss13] Rateb H. Issa. Optimal efficiency controller of AC drive system. International Journal of Computer Applications, 62 (??):40–46, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number12/10135-4935/.

Isupov:2016:AMC

[Isu16] Konstantin Isupov. An algorithm for magnitude comparison in RNS based on mixed-radix conversion II. International Journal of Computer Applications, 141(??):1–4, May 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume141/number5/24777-2016909626/

.

Ireeta:2021:MDS

[ITK21] Winston Tumps Ireeta, Achire Tito, and Yusto Kaahwa. Multirate digital signal processing for image reconstruction using error optimization techniques: the combined adaptive Weiner filter and the MDBUTMF. International Journal of Computer Applications, 183(??):1–15, August 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume183/number20/32038-2021921548/.

C:2018:NAC

[IU18] Ejiofor C I and Okon E. Uko. A neuroph approach for classifying ecoli data. International Journal of Computer Applications, 179(??):19–26, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number24/29080-2018916309/.

Inban:2010:DUD

[IUN10] Sakonthawat Inban, Nopadol Uchaipichat, and Nitikarn Nimsuk. The development of undistortion digital image application using backpropogation neural network model. International Journal of Computer Applications, 12(??):6– 9, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume12/
number8/1701-2230/.

James:2011:EBA

[JA11] Esther Annlin Kala James and S. Annadurai. An efficient Bayesian approach to face recognition based on wavelet transform. International Journal of Computer Applications, 15 (??):22–26, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume15/
number8/1969-2636/.

Jadhav:2012:CBI

[JA12] Seema Haribhau Jadhav and Shah Aqueel Ahmed. Content based image retrieval system with hybrid feature set and recently retrieved image library. International Journal of Computer Applications, 59(??):46–55, December 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume59/number5/9548-4001/.

Jabbar:2016:DPP

[JA16a] A. I. A. Jabbar and Fawaz Y. Abdullah. Delay performance of packet scheduling algorithms in LTE based machine-to-machine communications. International Journal of Computer Applications, 146(??):38–43, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number10/25438-2016910951/.

Jabbar:2016:ERP

[JA16b] A. I. A. Jabbar and Fawaz Y. Abdullah. Effect of RACH procedure on the performance of LTE-based M2M communication. International Journal of Computer Applications, 147(??):12–17, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume147/number5/25648-2016911071/.

Jayasooriya:2022:DBL

[JA22] G. R. I. L. Jayasooriya and Samantha Mathara Arachchi. Diagnosis of bacterial leaf blight, brown spots, and leaf smut rice plant diseases using light GBM. International Journal of Computer Applications, 183(??):53–58, January 2022. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number48/32259-
2022921895/.

Jabin:2014:SMP

[Jab14] Suraiya Jabin. Stock market prediction using feed-forward artificial neural network. International Journal of Computer Applications, 99(??):4–8, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume99/number9/17399-7959/.

Jimenez:2022:VTV

[JAB22] A. Jiménez, M. P. Aguilera, and G. Beltrán. Virtual twin of the virgin olive oil elaboration process from near infrared signals processing by artificial neural networks. International Journal of Computer Applications, 184(??):31–35, June 2022. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume184/number15/
32397-2022922145/.

Jabbar:2015:AFL

[JAG15a] A. I. A. Jabbar and Ahmed I. Al-Ghannam. Applying fuzzy logic principles to improve the performance of the random early detection algorithm. International Journal of Computer Applications, 122(??):27–31, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number17/21795-5172/.

Jagtap:2015:CSU

[Jag15b] Rajesh Jagtap. A comparative study of utilization of single and hybrid data mining techniques in heart disease diagnosis and treatment plan. International Journal of Computer Applications, 123(??):1–6, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number8/21976-2015905556/.

Jahanirad:2018:CEA

[Jah18] Hadi Jahanirad. Co-evolutionary approach to reduce soft error rate of implemented circuits on SRAM-based FPGA. International Journal of Computer Applications, 180(??):42– 49, May 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number43/
29422-2018917136/.

Jesmin:2013:BCR

[JARM13] Tasnuba Jesmin, Kawsar Ahmed, Md. Zamilur Rahman, and Md. Badrul Alam Miah. Brain cancer risk prediction tool using data mining. International Journal of Computer Applications, 61(??):22–27, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume61/number12/9980-4810/.

Jassim:2013:SOE

[Jas13] Firas A. Jassim. Semi-optimal edge detector based on simple standard deviation with adjusted thresholding. International Journal of Computer Applications, 68(??):43– 48, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume68/
number2/11555-6834/.

Jasim:2013:FLB

[JAsA13] Mahmood K Jasim, Anwar M Al-saleh, and Alaa Aljanaby. A fuzzy logic based handwritten numeral recognition system. International Journal of Computer Applications, 83(??):36–43, December 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume83/number10/14487-2796/.

Jayaweera:2012:AFI

[Jay12] Y. D. Jayaweera. Automatic file indexing framework: an effective approach to resolve dangling file pointers. International Journal of Computer Applications, 49(??): 6–11, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number15/7701-1052/.

Jayagowri:2015:TLP

[Jay15] R. Jayagowri. Techniques for low power and area optimized VLSI testing using novel scan flip-flop. International Journal of Computer Applications, 113(??): 22–28, March 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume113/number5/19824-1664/.

Jain:2013:DBE

[JBK13] Sandeep Kumar Jain, Manoranjan Rai Bharti, and Amardip Kumar. Distance based an efficient transmit power control scheme in cognitive radio system with multiple antennas. International Journal of Computer Applications, 72 (??):32–37, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number21/12668-9382/.

Jayaram:2016:DSS

[JBP16] M. A. Jayaram, Amrutha P B, and Chaithanya D P. Decision support system for detecting, diagnosing and suggesting remedial measures for building cracks. International Journal of Computer Applications, 137(??):9–15, March 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume137/number6/24278-2016908777/.

Jena:2023:PFA

[JBT23] Dinesh Kumar Jena, Abhyarthana Bisoyi, and Aruna Tripathy. Predictive framework for advanced customer churn prediction using machine learning. International Journal of Computer Applications, 185(??):43–51, June 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number13/32760-2023922818/.

Jain:2014:SBP

[JC14] Neha Jain and Yogesh Chaba. Simulation based performance analysis of zone routing protocol in Manet. International Journal of Computer Applications, 88(??):47– 52, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume88/
number4/15344-3684/.

Jadhav:2013:BEC

[JD13] J. B. Jadhav and P. J. Deore. Bandwidth enhancement of compact monopole antennas for application in UWB communication system. International Journal of Computer Applications, 76(??):48–52, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume76/number9/13279-0898/.

Jariwala:2015:PAE

[JD15] Krupa Jariwala and Upena Dalal. Performance analysis of eye localization methods for real time vision interface using low grade video camera. International Journal of Computer Applications, 114(??):33–40, March 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume114/number2/19952-1777/.

Jafar:2020:SMT

[JFJN20] Muhammad Naveed Jafar, Asma Farooq, Komal Javed, and Nazia Nawaz. Similarity measures of tangent, cotangent and cosines in neutrosophic environment and their application in selection of academic programs. International Journal of Computer Applications, 177(??):17–24, March 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number46/31217-2020919980/.

Jabade:2011:LRW

[JG11] Vaishali S. Jabade and Sachin R. Gengaje. Literature review of wavelet based digital image watermarking techniques. International Journal of Computer Applications, 31 (??):28–35, October 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume31/
number7/3838-5334/.

Janruang:2012:SSN

[JG12] Jongkol Janruang and Sumanta Guha. Semantic suffix net clustering for search results. International Journal of Computer Applications, 59(??):1–8, December 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume59/number7/9557-4017/.

Joonaki:2013:IAP

[JGZ13] Edris Joonaki, Shima Ghanaatian, and Ghassem Zargar. An intelligence approach for porosity and permeability prediction of oil reservoirs using seismic data. International Journal of Computer Applications, 80(??):19– 26, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume80/
number8/13881-1778/.

Jasim:2018:HAR

[JH18] Wala’a N. Jasim and Esra J. Harfash. Human activity recognition system to benefit healthcare field by using HOG and Harris techniques with K-NN model. International Journal of Computer Applications, 180(??):7–14, May 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number40/29393-2018917045/.

Jawarkar:2017:TIS

[JHB17] Naresh P. Jawarkar, Raghunath S. Holambe, and Tapan Kumar Basu. Text-independent speaker identification in emotional and whispered speech environments. International Journal of Computer Applications, 175(??):18–27, October 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume175/number5/28484-2017915465/

.

Jabari:2012:IAP

[JHP12] Nidal A. M. Jabari, Mohammad Hariadi, and Mauridhi Hery Purnomo. Intelligent adaptive presentation and e-testing system based on user modeling and course sequencing in virtual classroom. International Journal of Computer Applications, 50(??):23–31, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number9/7800-0921/.

Jaichandran:2010:FAA

[JI10a] R Jaichandran and A. Anthony Irudhayaraj. A futuristic analysis on the application problems of wireless sensor networks (WSN) with effective strategies and optimal solutions. International Journal of Computer Applications, 1(??):10–15, February 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume1/number8/189-326/.

Joshi:2010:PIL

[JI10b] Swapnil D. Joshi and V. S. Inamdar. Performance improvement in large graph algorithms on GPU using CUDA: an overview. International Journal of Computer Applications, 10(??):10–14, November 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume10/number10/1520-1992/.

Jeyachitra:2013:QED

[JIA13] J. Jeyachitra, K. Iyakutti, and K. Alagarsamy. Quality engineering of distributed system with hassle free load curve optimization. International Journal of Computer Applications, 77(??):38–45, September 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume77/number11/13441-1219/.

J:2014:ESD

[JIB14] Ayangbekun Oluwafemi J, Olatunde Ayoola I, and Felix O. Bankole. An expert system for diagnosis of blood disorder. International Journal of Computer Applications, 100(??):36–40, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number3/17509-8061/.

J:2012:ANN

[JIF12] Litta A J, Sumam Mary Idicula, and C. Naveen Francis. Artificial neural network model for the prediction of thunderstorms over Kolkata. International Journal of Computer Applications, 50(??):50–55, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number11/7819-1135/.

Jauro:2014:FAI

[JJA14] Fatsuma Jauro, S. B. Junaidu, and S. E. Abdullahi. Falcon-AO++: an improved ontology alignment system. International Journal of Computer Applications, 94(??): 1–7, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number2/16312-5541/.

Jalundhwala:2014:WFA

[JJKD14] Aarushi Jalundhwala, Pratik Jhaveri, Sandeep Khudanpur, and Amit Deshmukh. Wireless fingerprint attendance marking system. International Journal of Computer Applications, 108(??):1–5, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number8/18928-0294/.

Jain:2014:WBP

[JJN14] Bhavna Jain, Shailendra Jain, and R. K. Nema. Wavelet based power quality monitoring in grid connected wind energy conversion system. International Journal of Computer Applications, 98(??):7–15, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume98/number18/17281-7140/.

Jawadwala:2014:FRS

[JK14a] Maryam K. Jawadwala and Seema Kolkur. Feature ranking in sentiment analysis. International Journal of Computer Applications, 94(??):42–49, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume94/number13/16407-6150/.

Jeyepalan:2014:ABP

[JK14b] D. P. Jeyepalan and E. Kirubakaran. Agent based parallelized intrusion detection system using ant colony optimization. International Journal of Computer Applications, 105(??):1–6, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number10/18410-9687/.

Jipeng:2011:SCC

[JKC11] Tian Jipeng, G. Hemantha Kumar, and H. K. Chethan. Skew correction for Chinese character using Hough transform. International Journal of Computer Applications, 22 (??):33–36, May 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume22/
number2/2554-3499/.

Jayakody:2017:SHD

[JKM^{+}17] J. A. D. C. A. Jayakody, P. D. Kudagamage, H. A. W. P. Madushanka, S. S. Pathirana, and H. B. D. N. Dushanthi. Smart home delivery e-shopping cart to facilitate disable individuals. International Journal of Computer Applications, 180(??):25–32, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number6/28806-2017916047/.

Javed:2012:AVC

[JKS^{+}12] Wajahat Javed, Aamir Khan, Khurram Shahzad, Muhammad Asif, and Faisal Munir. Analysis of video coding and error resilience tools of H.264/AVC in wireless environment. International Journal of Computer Applications, 50(??):1–7, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume50/
number1/7732-0680/.

Joshi:2015:SHG

[JKTM15] Tejashri J. Joshi, Shiva Kumar, N. Z. Tarapore, and Vivek Mohile. Static hand gesture recognition using an Android device. International Journal of Computer Applications, 120(??):48–53, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number21/21356-4348/.

Jayakody:2016:IPN

[JLCV16] J. A. D. C. Anuradha Jayakody, Shashika Lokuliyana, Dinusha Chathurangi, and Demini Vithana. Indoor positioning: Novel approach for Bluetooth networks using RSSI smoothing. International Journal of Computer Applications, 137(??):26–32, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume137/number13/24337-2016909028/.

Jabbar:2014:UTF

[JM14] A. I. A. Jabbar and Emad A. Mohammed. Utilizing TV frequency bands using coalitional game theory according to the standard of IEEE 802.22. International Journal of Computer Applications, 85(??):10–16, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume85/number6/14844-3089/.

Jouybari:2015:FZO

[JM15] Hoda Naghibi Jouybari and Karim Mohammadi. FT-Z-OE: a fault tolerant and low overhead routing algorithm on TSVbased 3D network on chip links. International Journal of Computer Applications, 115(??):23–29, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume115/number2/20123-2191/.

Judith:2016:AHS

[JM16] I. Diana Judith and G. J. Joyce Mary. An advance halftone secure secret sharing scheme with error diffusion technique in visual cryptography. International Journal of Computer Applications, 142(??):40–46, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number7/24912-2016909952/.

Jain:2018:OBS

[JM18] Kavita Jain and Tanuj Manglani. Optimal bidding strategy for profit maximization of generation companies based on whale optimization algorithm in day ahead market. International Journal of Computer Applications, 179(??):29–41, April 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume179/number38/
29327-2018916867/.

Joseph:2021:HOS

[JMM21] Kithinji Joseph, Makau S. Mutua, and Gitonga D. Mwathi. HPDDRR: Optimized scheduler shaper for bandwidth management and traffic shaping in Internet protocol storage area networks. International Journal of Computer Applications, 183(??):20–32, November 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number36/32163-2021921746/.

Jariwala:2015:RTR

[JND15] Krupa Jariwala, Aakash Nandi, and Upena Dalal. A real time robust eye center localization using geometric eye model and edge gradients in unconstrained visual environment. International Journal of Computer Applications, 128(??):22–27, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number1/22838-2015906415/.

Jan:2019:MIC

[JNR19] Nazir Jan, Nasruminallah, and Rehmatullah. Multidimensional image compression through discrete wavelet transform in Matlab using Daubechies wavelet. International Journal of Computer Applications, 178(??):13–16, May 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number8/30549-2019918778/.

Jobair:2017:ANS

[Job17a] Hussein K. Jobair. Analytical and numerical study of the temperature distribution for a solid sphere subjected to a uniform heat generation. International Journal of Computer Applications, 168(??):30–37, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume168/number2/27850-2017914304/.

Jobair:2017:IPC

[Job17b] Hussein K. Jobair. Improving of photovoltaic cell performance by cooling using two different types of fins. International Journal of Computer Applications, 157(??): 6–15, January 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume157/number5/26825-2017912691/.

John:2014:HES

[Joh14] T. Stephen John. High efficient Seebeck thermoelectric device for power system design and efficiency calculation: a review of potential household appliances. International Journal of Computer Applications, 97(??):37–42, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume97/number18/17109-7763/.

Jovanovic:2015:FMS

[Jov15] Slobodan Jovanović. Flexible manufacturing systems and quantitative anlysis of flexible manufacturing systems. International Journal of Computer Applications, 132(??): 6–14, December 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume132/number1/23556-2015907243/.

Jasuja:2015:CCS

[JP15] Bobby Jasuja and Abhishek Pandya. Crypto-compression system: an integrated approach using stream cipher cryptography and entropy encoding. International Journal of Computer Applications, 116(??):34–41, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume116/number21/20463-2831/.

Jog:2016:GPA

[JP16] Pratik P. Jog and Archana B. Patankar. Graphical password authentication using a fake cursor approach. International Journal of Computer Applications, 148(??):35–40, August 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume148/number11/25804-
2016911298/.

Jaganathan:2011:SGM

[JPAK11] S. Jaganathan, S. Palaniswami, R. Adithya, and M. Naresh Kumaar. Synchronous generator modelling and analysis for a microgrid in autonomous and grid connected mode. International Journal of Computer Applications, 13(??): 3–7, January 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume13/
number5/1779-2454/.

Jena:2014:STH

[JPDB14] C. Jena, C. K. Panigrahi, M. Das, and M. Basu. Shortterm hydrothermal scheduling by artificial immune system algorithm. International Journal of Computer Applications, 107(??):5–14, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number18/18849-0365/.

Jain:2018:RAS

[JPJ18] Ruby Bhuvan Jain, Manimala Puri, and Umesh Jain. A robust approach to secure structured sensitive data using non-deterministic random replacement algorithm. International Journal of Computer Applications, 179(??):17–21, June 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume179/number50/
29515-2018917306/.

Jayasree:2012:DAE

[JPP12] P. V. Y. Jayasree, G. Raghu Poojita, and J. Chaitanya Priya. Design of active electromagnetic interference filter to eliminate common-mode noise in conducted interference. International Journal of Computer Applications, 43 (??):10–15, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume43/
number11/6146-8518/.

J:2012:ECG

[JS12a] Jacinth Salome J and R M Suresh. Efficient clustering for gene expression data. International Journal of Computer Applications, 47(??):30–35, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume47/number5/7186-9925/.

Jaiganesh:2012:KEL

[JS12b] V. Jaiganesh and P. Sumathi. Kernelized extreme learning machine with Levenberg–Marquardt learning approach towards intrusion detection. International Journal of Computer Applications, 54(??):38–44, September 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume54/number14/8638-2577/.

Jamaludin:2012:CPB

[JS12c] Nor Azliana Akmal Jamaludin and Shamsul Sahibuddin. Challenges of a project-based learning approach towards requirement engineering. International Journal of Computer Applications, 50(??):1–5, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number3/7748-0802/.

Jouhari:2012:IEC

[JS12d] Houda Jouhari and El Mamoun Souidi. Improving embedding capacity by using the Z4-linearity of Preparata codes. International Journal of Computer Applications, 53 (??):1–6, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number18/8518-2553/.

Jain:2014:ESM

[JS14] Toshi Jain and Sandeep Pratap Singh. An efficient and secure multi-server smart card based authentication scheme. International Journal of Computer Applications, 93(??):1–7, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number13/16272-6025/.

Joshi:2015:NAC

[JS15] A. D. Joshi and J. V. Shinde. A novel approach for color image feature extraction using swarm intelligence. International Journal of Computer Applications, 111(??): 29–35, February 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume111/number7/19553-1276/.

Jahan:2018:RPP

[JS18a] Tasneem Jahan and Amit Saxena. A robust privacy preserving approach of outsourced data by modified frequent Web access pattern. International Journal of Computer Applications, 182(??):7–11, July 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number5/29756-2018917438/.

Joshi:2018:VIC

[JS18b] Nikhilesh Joshi and Tanuja K. Sarode. Validation of image compression algorithms using neural network. International Journal of Computer Applications, 179(??):1–8, January 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume179/number12/
28849-2018916127/.

Jimson:2012:NMB

[JSC12] L. Jimson, P. Senthil, and D. Chandrakanth. A new methodology for blind image deconvolution. International Journal of Computer Applications, 43(??):1–5, April 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume43/number20/6216-8686/.

Jadon:2014:MSC

[JSD14] Shruti Jadon, Anubhav Singhal, and Suma Dawn. Military simulator — a case study of behaviour tree and unity based architecture. International Journal of Computer Applications, 88(??):26–29, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume88/number5/15350-3691/.

Johal:2010:HFB

[JSK10] Navdeep Kaur Johal, Samandeep Singh, and Harish Kundra. A hybrid FPAB/BBO algorithm for satellite image classification. International Journal of Computer Applications, 6(??):31–36, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume6/number5/1074-1403/.

Johnson:2017:WAD

[JSK17] Marcus Okunlola Johnson, Arish Siddiqui, and Amin Karami. A wormhole attack detection and prevention technique in wireless sensor networks. International Journal of Computer Applications, 174(??):1–8, September 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume174/number4/28392-2017915376/

.

Jana:2017:DIP

[JSS17a] Saikat Jana, Jayashree Singha, and Souvik Singha. Design and implementation of path establishment and maintenance technique for MANETs. International Journal of Computer Applications, 158(??):5–10, January 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume158/number3/26886-2017912762/.

Jayakody:2017:RTI

[JSS^{+}17b] J. A. D. C. Anuradha Jayakody, R. H. Suriarachchi, O. W. Samarasinghe, B. A. Chamain, and A. H. M. C. M. Abeyrathna. Real-time indoor map generation to identify spatial environment changes. International Journal of Computer Applications, 180(??):38–43, December 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number5/28800-2017916045/.

Jafar:2020:ASS

[JSSA20] Muhammad Naveed Jafar, Muhammad Saqlain, Muhammad Saeed, and Qamar Abbas. Application of soft-set relations and soft matrices in medical diagnosis using Sanchez’s approach. International Journal of Computer Applications, 177(??):7–11, January 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number32/31105-2020919692/.

Jasser:2014:ACO

[JSY14] Muhammed Basheer Jasser, Mohamad Sarmini, and Rauf Yaseen. Ant colony optimization (ACO) and a variation of bee colony optimization (BCO) in solving TSP problem: a comparative study. International Journal of Computer Applications, 96(??):1–8, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume96/number9/16819-6587/.

Jayabal:2011:DSB

[JT11] Raymond J. Jayabal and Lau Chiew Tong. Deriving the suspended backoff counter value distribution in CSMA/ CA. International Journal of Computer Applications, 28 (??):14–20, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume28/
number9/3417-4768/.

Jain:2015:AEP

[JT15] Paras Jain and Vipin Tyagi. Adaptive edge-preserving image denoising using arbitrarily shaped local windows in wavelet domain. International Journal of Computer Applications, 114(??):33–45, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume114/number16/20065-2141/.

Jarjees:2019:ATS

[JTA19] Amer Mohammad Jarjees, Omar Turath Tawfeeq, and Ali Abbawi Mohammed Alabbawi. Adaptive tabu search optimization applied on fuzzy logic DC motor speed control. International Journal of Computer Applications, 182(??):30– 36, March 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume182/number45/
30456-2019918601/.

Joseph:2019:DNA

[JTI19] Mom Joseph, Olanipekun Tobi, and Gabriel A. Igwue. Development of a new adaptive beam-forming technique for smart antenna system. International Journal of Computer Applications, 178(??):43–49, May 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number11/30578-2019918845/.

Jha:2014:RRN

[JTKC14] Vivekanand Jha, Neeharika Taneja, Harshpreet Kaur, and Elisha Chauhan. Reliable relay node placement in large scale dense and non-uniform wireless sensor network. International Journal of Computer Applications, 86(??):1– 8, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number19/15101-2720/.

Jain:2015:STC

[JTS15] Monal Jain, Deepak Singh Tomar, and Shiv Kumar Singh. A survey on TCP congestion control schemes in guided media and unguided media communication. International Journal of Computer Applications, 118(??):20–24, May 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume118/number3/20726-3081/

.

J:2012:EMP

[JV12] Hannah Monisha J and Rhymend Uthariaraj V. Enhanced MAC parameters to support hybrid dynamic prioritization in MANETs. International Journal of Computer Applications, 45(??):35–41, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number18/7021-9754/.

Jaddu:2014:CFS

[JV14] Hussein Jaddu and Milan Vlach. Closed form solution of nonlinear-quadratic optimal control problem by state-control parameterization using Chebyshev polynomials. International Journal of Computer Applications, 91 (??):1–7, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume91/
number10/15914-5281/.

Jathar:2017:DIH

[JV17] Nidhi Jathar and Abhilasha Vyas. Design and implementation of hierarchical multi-class emotion classification model. International Journal of Computer Applications, 171(??):23–28, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume171/number8/28202-2017915149/.

Jeyalatha:2011:DIT

[JVF11] S. Jeyalatha, B. Vijayakumar, and Munawwar Firoz. Design and implementation of a tool for Web data extraction and storage using Java and uniform interface. International Journal of Computer Applications, 22(??):1–6, May 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume22/number4/2575-
3551/.

Jayababu:2015:PDC

[JVG15] Y. Jayababu, G. P. S. Varma, and A. Govardhan. Probabilistic distance clustering: a spatial clustering approach for remote sensing imagery. International Journal of Computer Applications, 126(??):9–14, September 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume126/number14/22619-2015906289/.

Jain:2012:MTM

[JVJ12] Neelesh Jain, Ashish Verma, and Shalini Jain. Multilevel test method for testing microcontroller based ECG system. International Journal of Computer Applications, 46 (??):31–35, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number9/6940-9395/.

Jayashree:2010:CEF

[JYP10] L. S. Jayashree, V. K. Yamini, and R. Manju Priya. A communication efficient framework for soil moisture monitoring using wireless sensor networks. International Journal of Computer Applications, 1(??):16–23, February 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume1/number16/348-528/.

Jaffery:2013:PAI

[JZS13] Zainul Abdin Jaffery, Zaheeruddin, and Laxman Singh. Performance analysis of image segmentation methods for the detection of masses in mammograms. International Journal of Computer Applications, 82(??):44– 50, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume82/
number2/14092-2100/.

Kavitha:2010:NFD

[KA10] K. Kavitha and S. Arivazhagan. A novel feature derivation technique for SVM based hyper spectral image classification. International Journal of Computer Applications, 1(??):25–31, February 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume1/number15/327-496/.

Kaba:2011:DSA

[KA11] Serkan Kaba and Murat ACAR. Designing a system to analyze portfolio risks and to determine optimum margin requirements. International Journal of Computer Applications, 32(??):34–40, October 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume32/number7/3918-5516/.

K:2012:LFA

[KA12a] Affum Emmanuel A K and Ed Danso Ansong. Load factor assessment and evaluation on the performance MIMO WCDMA Rayleigh interference channel. International Journal of Computer Applications, 52(??):12–17, August 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume52/number7/8213-
1631/.

K:2012:DAS

[KA12b] Ramesh K and Zameer Ahmed. Dynamically avoiding the substantial throughput penalty of FRR. International Journal of Computer Applications, 51(??):25–28, August 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume51/number9/8070-
1464/.

Kadhim:2012:NAV

[KA12c] Dheyaa Jasim Kadhim and Sanaa Shaker Abed. A new approach of vertical handoff in the heterogeneous wireless networks. International Journal of Computer Applications, 57 (??):1–5, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number9/9139-3356/.

Kalaichelvi:2012:ANN

[KA12d] V. Kalaichelvi and Ahammed Shamir Ali. Application of neural networks in character recognition. International Journal of Computer Applications, 52(??):1–6, August 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume52/number12/8251-1770/.

Khaleel:2012:ERS

[KA12e] Turkan Ahmed Khaleel and Manar Younis Ahmed. The enhancement of routing security in mobile ad-hoc networks. International Journal of Computer Applications, 48 (??):41–48, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number16/7435-0480/.

Kumar:2012:MEF

[KA12f] V. Arul Kumar and L. Arockiam. MFSPFA: an enhanced filter based feature selection algorithm. International Journal of Computer Applications, 51(??):27– 31, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number12/8096-1682/.

Kuppusamy:2012:MPK

[KA12g] K. S. Kuppusamy and G. Aghila. A model for personalized keyword extraction from Web pages using segmentation. International Journal of Computer Applications, 42 (??):21–26, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number4/5682-7720/.

Kumar:2013:EHA

[KA13a] Kshitij Kumar and Suneeta Agarwal. An efficient hierarchical approach for background subtraction and shadow removal using adaptive GMM and color discrimination. International Journal of Computer Applications, 75(??): 1–7, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number12/13160-0752/.

Kumar:2013:IKM

[KA13b] Vijay Kumar and Priyanka Ahlawat. Improving key management scheme supporting node mobility in heterogeneous sensor networks using collision-keys. International Journal of Computer Applications, 70(??):40–44, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume70/number23/12211-8361/.

Kazi:2014:AIS

[KA14] Hameedullah Kazi and Asia Kainat Awan. Accepting inferred student solutions by tutoring system in an ill-defined domain. International Journal of Computer Applications, 94(??):8–11, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number18/16457-5544/.

Kaur:2015:NHT

[KA15] Sarabjeet Kaur and Ada. A new hybrid technique for iris recognition. International Journal of Computer Applications, 122(??):11–18, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number13/21759-4993/.

K:2016:NMO

[KA16a] Mallikharjuna Rao K and K. Anuradha. A new method to optimize the reliability of software reliability growth models using modified genetic swarm optimization. International Journal of Computer Applications, 145(??):1–8, July 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume145/number5/25271-2016910610/

.

Kaur:2016:CCN

[KA16b] Mandeep Kaur and Aarti. Color channel normalization and fuzzy filter based image fusion in discrete wavelet domain. International Journal of Computer Applications, 151(??):33–37, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume151/number5/26233-2016911820/.

Kolhar:2017:FSB

[KA17] Manjur Kolhar and Abdalla Alameen. Fairly scheduled bandwidth for the data centre traffic. International Journal of Computer Applications, 166(??):38–43, May 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume166/number5/27668-2017914102/.

Kissi:2021:PTI

[KA21] Michael Kyei Kissi and Michael Asante. Penetration testing of IEEE 802.1X port-based authentication protocols using Kali Linux hacking tools. International Journal of Computer Applications, 174(??):19–26, March 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume174/number26/31838-2021921186/.

Kaiser:2013:ITC

[KAA13] M. Shamim Kaiser, Shamim Al Mamun, and Kazi M. Ahmed. Interference temperature constraint-based radio resource allocation for the network coded cognitive cooperative network (NCCCN). International Journal of Computer Applications, 67(??):1–6, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume67/number7/11404-6728/.

Kommey:2020:GBF

[KAA20] B. Kommey, C. Acquaye, and T. A. Abeka. GhanaToll: Barrier-free RFID toll collection and management system for Ghana. International Journal of Computer Applications, 176(??):22–29, June 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number27/31368-2020920278/.

Khan:2013:PES

[KAAA13] Shumaila Khan, Syed M. Anwar, Sobia Arshad, and Asjad Amin. Power efficient and secure IPV6 robust header compression technique for low power wireless personal area networks. International Journal of Computer Applications, 62 (??):10–14, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number18/10179-4982/.

Kumar:2011:IBA

[KAAM11] Mohit Kumar, K. K. Agrawal, Deepak Arora, and Reena Mishra. Implementation and behavioural analysis of graph clustering using restricted neighborhood search algorithm. International Journal of Computer Applications, 22(??): 15–20, May 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume22/
number6/2589-3581/.

Kadhim:2016:CHO

[Kad16]  Hussam Jawad Kadhim.  Controlling home and of-

fice appliances with the Bluetooth of smartphone. International Journal of Computer Applications, 152(??): 25–28, October 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume152/number3/26301-2016911821/.

Khedr:2015:MBS

[KAFS15] Ayman E. Khedr, Manal A. Abdel-Fattah, and Marwa M. Solayman. Merging balanced scorecards and business intelligence techniques: an applied case on the IT subsidiary company in the Egyptian civil aviation. International Journal of Computer Applications, 121(??):26–51, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume121/number11/21586-4658/.

Kalhoodashti:2011:CCC

[Kal11] Hossein Ebadi Kalhoodashti. Concentration control of CSTR using NNAPC. International Journal of Computer Applications, 26(??):34–38, July 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume26/number6/3106-4265/.

Kayode:2015:EPR

[KAL15] A. A. Kayode, J. A. Adewale, and N. T. A. Lawal. An exploration of prevalence of repetitive stress injuries among computer operators in Nigeria. International Journal of Computer Applications, 109(??):1–5, January 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume109/number1/19149-0568/.

Kamath:2017:RIO

[Kam17] Vidya Kamath. Results and inference obtained from a small implementation of the DF-ICF — the modified TF-IDF. International Journal of Computer Applications, 166(??):20–23, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume166/number1/27633-2017913877/.

Khan:2011:IEN

[KAP11] Nafis Uddin Khan, K. V. Arya, and Manisha Pattanaik. Image enhancement and de-noising by diffusion based singular value decompositiong. International Journal of Computer Applications, 32(??):14–22, October 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume32/number8/3923-5538/.

K:2015:CUD

[KAP15] Nageswara Rao K, Germina K A, and Shaini P. A characterization of k-uniform DCSL graphs. International Journal of Computer Applications, 125(??):1–5, September 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume125/number7/22441-2015905956/

.

Khaliq:2014:FEW

[KAqC14] Noor Abdul Khaliq, Salih Mahdy Al-qaraawi, and Muayad Sadik Croock. Fetal electrocardiography and Web based outof-hospital monitoring system. International Journal of Computer Applications, 93(??):36–41, May 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.
org/archives/volume93/number3/16198-5468/.

Karthik:2010:IDG

[KAR10] S. Karthik, V. P. Arunachalam, and T. Ravichandran. An investigation of 2n direction geographical traceback using direction ratio sampling algorithm (DRSA) & IP traceback strategies. International Journal of Computer Applications, 1(??):59–64, February 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume1/number3/82-178/.

Karaci:2011:PCM

[Kar11] Abdulkadir Karaci. Performance comparison of managed C# and Delphi Prism in Visual Studio and unmanaged Delphi 2009 and C++ Builder 2009 languages. International Journal of Computer Applications, 26(??):9–15, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume26/number1/3070-4199/.

Kaur:2012:AEB

[KAR12] Mandeep Kaur, P. Ahmed, and M. Qasim Rafiq. Analyzing EEG based neurological phenomenon in BCI systems. International Journal of Computer Applications, 57 (??):40–49, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number17/9209-3755/.

Kumar:2015:AWS

[KAR15] P. Kumar, V. Anandarangan, and A. Reshma. An approach to workflow scheduling using priority in cloud computing environment. International Journal of Computer Applications, 109(??):32–38, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume109/number11/19236-1008/.

Kumar:2012:NMA

[KAS12] K. Ravi Kumar, S. Anand, and M. Sydulu. A novel multi agent based PSO approaches for security constrained optimal power flows using smooth and non-smooth cost functions. International Journal of Computer Applications, 41 (??):14–21, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume41/
number3/5521-7552/.

Kumar:2013:OAM

[KAS13] Narander Kumar, Shalini Agarwal, and Vipin Saxena. Overload avoidance model using optimal placement of virtual machines in cloud data cetres. International Journal of Computer Applications, 73(??):18–25, July 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume73/number11/12786-9963/.

Kasurinen:2017:EWB

[Kas17] Jussi Kasurinen. Experiences from a Web-based course in software testing and quality assurance. International Journal of Computer Applications, 166(??):11–16, May 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume166/number2/27640-2017913918/

.

Kaur:2015:HET

[Kau15] Ramandeep Kaur. Histogram equalization tool: Brightness preservation and contrast enhancement using segmentation with opening-by-reconstruction. International Journal of Computer Applications, 111(??):11–23, February 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume111/number2/19509-1126/

.

Khan:2013:DPS

[KAZA13] Muhammad Akbar Ali Khan, Sahibzada Abdur Rehman Abid, Fatima Tuz Zuhra, and Nasir Ahmad. The development of Pashto speech synthesis system. International Journal of Computer Applications, 71(??):49– 53, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number24/12695-9478/.

Kamoun:2012:NGM

[KB12a] Karim Kamoun and Sadok Ben Yahia. A novel global measure approach based on ontology spectrum to evaluate ontology enrichment. International Journal of Computer Applications, 39(??):23–30, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number17/4913-7466/.

Kumaran:2012:SMI

[KB12b] N. Kumaran and R. Bhavani. Spine MRI image retrieval using texture features. International Journal of Computer Applications, 46(??):1–7, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume46/number24/7120-9178/.

Kaushik:2014:EIV

[KB14a] Rajeev Kaushik and Rakesh K Bajaj. On exponential interval valued intuitionistic fuzzy entropy of order a and type b and its applications in decision making. International Journal of Computer Applications, 95(??):1–6, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume95/number12/16643-6615/.

Khormuji:2014:PNE

[KB14b] Morteza Kolali Khormuji and Mehrnoosh Bazrafkan. Persian named entity recognition based with local filters. International Journal of Computer Applications, 100(??):1–6, August 2014. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume100/number4/17510-
8062/.

Kumar:2015:LFR

[KB15] B. Senthil Kumar and S. Santhosh Baboo. Linking and familiarity rating method classifies the music, video assessment responses of EEG-signal. International Journal of Computer Applications, 116(??):6–11, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number10/20370-2577/.

Kaur:2016:CRM

[KB16a] Jashandeep Kaur and Rekha Bhatia. Customer relationship management classification by hybridizing genetic algorithm and fuzzy K-nearest neighbor. International Journal of Computer Applications, 147(??):13–17, August 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume147/number13/25713-2016911286/.

Kumar:2016:LMA

[KB16b] K. J. Jegadish Kumar and R. Balasubramanian. Lightweight mixcolumn architecture for advanced encryption standard. International Journal of Computer Applications, 136(??):31–36, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume136/number11/24200-2016908583/.

Kumar:2016:NID

[KB16c] Koushal Kumar and Jaspreet Singh Batth. Network intrusion detection with feature selection techniques using machine-learning algorithms. International Journal of Computer Applications, 150(??):1–13, September 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume150/number12/26144-2016910764/.

Karaci:2010:RHS

[KBA10] Abdulkadir Karaci, Ibrahim Buyukyazici, and Muharrem Aktumen. Recognition of human speech using q-Bernstein polynomials. International Journal of Computer Applications, 2(??):22–28, June 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume2/number5/666-934/.

Klugah-brown:2015:SSC

[KBAQ15] Benjamin Klugah-Brown, John Bosco Aristotle Kanpogninge Ansuura, and Xia Qi. A signcryption scheme from certificateless to identity-based environment for WSNs into IoT. International Journal of Computer Applications, 120(??):16–23, June 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume120/number9/21255-4099/.

Kasapovic:2015:PVD

[KBBM15] Suad Kasapovic, Esmir Biberkic, and Lejla Banjanovic-Mehmedovic. Performances of virtualization in design of fully collapsed demilitarized zone. International Journal of Computer Applications, 131(??):1–6, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume131/number13/23506-2015907475/.

Khelifa:2017:SIO

[KBEA17] Mohamed O. M. Khelifa, Mostafa Belkasmi, Yahya O. Mohamed Elhadj, and Yousfi Abdellah. Strategies for implementing an optimal ASR system for Quranic recitation recognition. International Journal of Computer Applications, 172(??):35–41, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume172/number9/28282-2017915209/.

Khairy:2018:PEN

[KBEA18] Dalia Khairy, Gamal Behery, A. A. Ewees, and Elsaeed AbdElrazek. Performance evaluation of neural network and deep neural network for human activity recognition. International Journal of Computer Applications, 180(??):44–50, February 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume180/number21/29060-
2018916503/.

Kumar:2011:SPG

[KBG11] Tanuj Kumar, Rakesh Kumar Bajaj, and Nitin Gupta. On some parametric generalized measures of fuzzy information, directed divergence and information improvement. International Journal of Computer Applications, 30(??):5– 10, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume30/
number9/3666-5185/.

Kamal:2010:AMV

[KBK10] Reet Kamal, Sukhwinder Bir, and Amanjot Kaur. Approximation of missing values in DNA microarray gene expression data. International Journal of Computer Applications, 4(??):20–25, July 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume4/number3/810-1150/.

Kulshrestha:2012:FPR

[KBK12] Megha Kulshrestha, V. K. Banga, and Sanjeev Kumar. Finger print recognition: Survey of minutiae and Gabor filtering approach. International Journal of Computer Applications, 50(??):17–21, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number4/7759-0823/.

Kidoudou:2022:NRM

[KBM22] Peter Arnaud Kidoudou, Regis F. Babindamana, and Benjamin Mampassi. A new rank metric codes based identification scheme. International Journal of Computer Applications, 184(??):14–17, October 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number34/32534-2022922409/.

Katuuk:2018:AFL

[KBN18] Ronny Katuuk, Marson James Budiman, and Donald B. Noya. Application of fuzzy logic algorithm for hybrid car usage time. International Journal of Computer Applications, 181(??):1–4, October 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number21/30006-2018917918/.

Kuadey:2021:BHA

[KBT^{+}21] Noble Arden Kuadey, Lily Bensah, Baidenger Agyekum Twumasi, Carlos Ankora, Gerald Tietaa Maale, and Anthony Mawuena Kuadey. Black hole attack in mobile ad hoc networks: Challenges and directions. International Journal of Computer Applications, 183(??):1–7, October 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number29/32110-
2021921672/.

Karthikeyani:2012:CDM

[KBTB12] V. Karthikeyani, I. Parvin Begum, K. Tajudin, and I. Shahina Begam. Comparative of data mining classification algorithm (CDMCA) in diabetes disease prediction. International Journal of Computer Applications, 60 (??):26–31, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number12/9745-4307/.

Kumar:2011:SSA

[KC11] Sumit Kumar and Siddhartha Chauhan. A survey on scheduling algorithms for wireless sensor networks. International Journal of Computer Applications, 20(??):7–13, April 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume20/number5/2432-
3271/.

Kumaar:2012:PSS

[KC12] T. A. Ramesh Kumaar and I. A. Chidambaram. Power system security enhancement using FACTS devices in a power system network with voltage dependent loads and ZIP loads. International Journal of Computer Applications, 45 (??):26–39, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume45/
number4/6769-9054/.

Kanagalakshmi:2014:NSP

[KC14a] K. Kanagalakshmi and E. Chandra. Novel shift-phase transformation based cancelable and irrevocable biometric template generation for fingerprints. International Journal of Computer Applications, 89(??):41–53, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume89/number20/15751-4708/.

Ketkar:2014:EPE

[KC14b] Maitreyi Ketkar and A. M. Chopde. Efficient parameter extraction of solar cell using modified ABC. International Journal of Computer Applications, 102(??): 1–6, September 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume102/number1/17776-8535/.

Kumar:2014:MCL

[KC14c] Narander Kumar and Priyanka Chaudhary. Minimize cyber losses in cyber world through the optimization technique. International Journal of Computer Applications, 96 (??):18–22, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number20/16910-6993/.

Kirupa:2015:UBA

[KC15a] K. Kirupa and K. Udaya Chandrika. Unreliable batch arrival retrial queue with negative customers, multi-types of heterogeneous service, setup time and reserved time. International Journal of Computer Applications, 127(??):38–42, October 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume127/number10/22769-
2015906529/.

Kumar:2015:MFB

[KC15b] A. Ramesh Kumar and C. Chandrasekar. Machine flow based energy-power approximation on elastic cloud services. International Journal of Computer Applications, 128(??):1–8, October 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume128/number8/22890-2015906616/.

Kushwaha:2015:PAA

[KC15c] Shalini Kushwaha and Deepak Chaudhary. Performance analysis of AODV, CBRP, AOMDV, DSR MANET routing protocol using NS2 simulation. International Journal of Computer Applications, 116(??):12–18, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume116/number5/20331-1893/.

Kanakam:2020:PCT

[KC20] Prathyusha Kanakam and ASN Chakravarthy. Packet crafting tools for cyber crime security attacks. International Journal of Computer Applications, 176(??):28–30, June 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume176/number31/31402-
2020920360/.

Kapusuz:2013:ANN

[KCD13] K. Y. Kapusuz, S. Can, and B. Dagdeviren. Artificial neural network based bandwidth estimation of a CPW-fed patch antenna. International Journal of Computer Applications, 69(??):37–40, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume69/number11/11890-7938/.

Khlaif:2014:STL

[KCS14] Abdulmomen Kadhim Khlaif, Muayad Sadik Croock, and Shaimaa Hameed Shaker. Simulating traffic lights control using wireless sensor networks. International Journal of Computer Applications, 104(??):1–8, October 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume104/number12/18251-9445/.

Kumar:2015:INS

[KD15] Manish Kumar and Debasis Dwibedy. Implementation of a novel system for cross platform communication of diversified applications over network. International Journal of Computer Applications, 128(??):38–44, October 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume128/number9/22904-2015906631/.

Kiran:2022:SSI

[KD22] Kiran and Parameshachari B D. Secure selective image encryption technique for real time applications. International Journal of Computer Applications, 184(??): 15–20, March 2022. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume184/number5/32327-2022922009/.

Kehagias:2017:ABS

[KDB17] Dimitris Kehagias and V. Douskas-Bertlviser. Androidbased simulator to support tomasulo algorithm teaching and learning. International Journal of Computer Applications, 170(??):24–29, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume170/number2/28043-2017914703/.

Kalita:2012:SCT

[KDBT12] S. Kalita, M. Devi, A. K. Barbara, and P. H. Talukdar. Soft computing technique for recognition of earthquake precursor from low latitude total electron content (TEC) profiles. International Journal of Computer Applications, 44 (??):11–14, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number17/6354-8775/.

Khawandi:2012:RTR

[KDC12] Shadi Khawandi, Bassam Daya, and Pierre Chauvet. Real time recognition of elderly daily activity using fuzzy logic through fusion of motion and location data. International Journal of Computer Applications, 54(??):55–60, September 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume54/number3/8549-
2109/.

Kissoon:2017:SIM

[KDM17] Dishay Kissoon, Hinouccha Deerpaul, and Avinash Mungur. A smart irrigation and monitoring system. International Journal of Computer Applications, 163(??):39–45, April 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume163/number8/27418-2017913688/

.

Kolhe:2010:HPL

[KDS10] K. R. Kolhe, P. R. Devale, and P. Shrivastava. High performance lossless multimedia data compression through improved dictionary. International Journal of Computer Applications, 10(??):29–35, November 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume10/number1/1444-1953/.

Kapoh:2016:DDP

[KDT16] Harson Kapoh, Fanny Doringin, and Meiny M. Taju. Design and development of the public service information: a case study of villages in District East Tombariri, Minahasa-Indonesia. International Journal of Computer Applications, 156(??):6–11, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume156/number12/26759-2016912493/.

Kulkarni:2011:CAU

[KDTD11] Sandhya Kulkarni, Shwetha D, Devaraju J T, and D. Das. Channel aware uplink scheduler for a mobile subscriber station of IEEE 802.16e. International Journal of Computer Applications, 35(??):15–22, December 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume35/number6/4404-6115/.

Khalifa:2014:MDA

[KE14] Amal Khalifa and Dina Elsayad. Microarrays data analysis for cancer disease on a cluster of computers. International Journal of Computer Applications, 95(??): 13–20, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number20/16709-6864/.

Kaplan:2016:PMS

[KE16] N. H. Kaplan and I. Erer. Pansharpening of multispectral satellite images via lattice structures. International Journal of Computer Applications, 140(??):9–14, April 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume140/number7/24605-2016909366/.

Kumar:2019:SPU

[KEA19] Rakesh Kumar, Afshan Ejaz, and Syed Asim Ali. Satisfaction, perceived usability and trust impact on website trustworthiness. International Journal of Computer Applications, 178(??):50–57, June 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number23/30678-2019919106/.

Khaled:2013:APD

[KEBF13] H. Khaled, R. El Gohary, N. L. Badr, and H. M. Faheem. Accelerating pairwise DNA sequence alignment using the CUDA compatible GPU. International Journal of Computer Applications, 84(??):25–31, December 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume84/number1/14542-2619/.

Keivanian:2014:MAP

[Kei14] Farshid Keivanian. Minimization of average power consumption in 3 stage CMOS ring oscillator based on MSFLA, Fuzzy-MSFLA, GA, and fuzzy-GA. International Journal of Computer Applications, 104(??):30–37, October 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume104/number16/18289-9441/.

Kermia:2013:SAD

[Ker13] Omar Kermia. Schedulability analysis of distributed realtime applications under dependence and several latency constraints. International Journal of Computer Applications, 62 (??):1–7, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number14/10145-4978/.

Kandil:2017:PSC

[KETESH17] Ali Kandil, Osama El-Tantawy, Sobhy El-Sheikh, and Shawqi Ahmed Hazza. Pairwise soft connected in soft bitopological spaces. International Journal of Computer Applications, 169(??):12–27, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume169/number11/28028-2017914614/.

Kinjarapu:2014:REF

[KEV14] Ananda Kumar Kinjarapu, Jagadeeswara Rao Enni, , and Kamakshi Prasad Valurouthu. Robust and efficient fully parallel 2D thinning algorithm. International Journal of Computer Applications, 85(??):1–6, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume85/number5/14834-3092/.

Keyrouz:2012:FMP

[Key12] Fakheredine Keyrouz. A fast-multiplying PSO algorithm for real-time multiple object tracking. International Journal of Computer Applications, 60(??):1–6, December 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume60/number3/9669-4098/.

Kamdjou:2021:LRH

[KFEN21] Hugues M. Kamdjou, Elie T. Fute, Adnen El Amraoui, and Armand Nzeukou. A low-resources hardware-based audio data compression scheme for wireless sensors networks. International Journal of Computer Applications, 174(??):13– 18, April 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume174/number32/
31883-2021921258/.

Kumar:2011:BER

[KG11] Sanjeev Kumar and Ragini Gupta. Bit error rate analysis of Reed–Solomon code for efficient communication system. International Journal of Computer Applications, 30(??): 11–15, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume30/
number12/3710-5174/.

Kankariya:2015:EAR

[KG15a] Chetana Kankariya and Sudesh Gupta. An efficient algorithm for removal of salt and pepper noise from images. International Journal of Computer Applications, 118(??):24–28, May 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume118/number18/20846-
3502/.

Kaur:2015:CEI

[KG15b] Amandeep Kaur and Amit Grover. Capacity estimation of an infrastructure network carrying HTTP and e-mail data based on IEEE 802.11b WLAN standard. International Journal of Computer Applications, 123(??):36–40, August 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume123/number16/22047-
2015905767/.

Kushwaha:2015:RTR

[KG15c] Monika Kushwaha and Saurabh Gupta. Response time reduction and performance analysis of load balancing algorithms at peak hours in cloud computing. International Journal of Computer Applications, 128(??):26–31, October 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume128/number17/22967-
2015906762/.

Kumar:2010:FMA

[KGAM10] P. N. Kumar, Rahul Seshadri G, Hariharan A, and V. P. Mohandas. Financial market analysis of Bombay Stock Exchange using an agent based model. International Journal of Computer Applications, 8(??):37–42, October 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume8/number13/1304-1642/.

Kaur:2010:MBS

[KGK10] Manvjeet Kaur, Akshay Girdhar, and Mandeep Kaur. Multimodal biometric system using speech and signature modalities. International Journal of Computer Applications, 5(??):13–16, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume5/number12/962-1339/.

Kumar:2013:ACB

[KGK13] Rakesh Kumar, Girdhar Gopal, and Rajesh Kumar. Alpha cut based novel selection for genetic algorithm. International Journal of Computer Applications, 70(??): 13–17, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number21/12190-8268/.

Khalifa:2012:RNB

[KH12] Amal Khalifa and Safwat Hamad. A robust non-blind algorithm for watermarking color images using multi-resolution wavelet decomposition. International Journal of Computer Applications, 37(??):33–39, January 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume37/number8/4631-6666/.

Kolyaie:2013:MAD

[KH13] Elmira Kolyaie and Ali Harounabadi. Mapping activity diagram to fuzzy GSPN. International Journal of Computer Applications, 84(??):24–29, December 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume84/number4/14566-2684/.

Khokhar:2015:PAF

[KH15] Nasir Mahmud Khokhar and Wail Harasani. Performance analysis of fast block matching motion estimation algorithms. International Journal of Computer Applications, 122(??):15–20, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number6/21704-4817/.

Kamalapurkar:2020:WSS

[KH20] Savitha Kamalapurkar and Samyama Gunjal G H. Web support system for prediction of heart disease using k-nearest neighbor algorithm. International Journal of Computer Applications, 175(??):1–6, August 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number14/31518-2020920638/.

Khairullah:2011:NTS

[Kha11] Md. Khairullah. A novel text steganography system in cricket match scorecard. International Journal of Computer Applications, 21(??):43–47, May 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume21/number9/2537-3462/.

Khairullah:2013:ASC

[Kha13] Md. Khairullah. An analysis of scan converting a line with multi symmetry. International Journal of Computer Applications, 61(??):30–33, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume61/number15/10006-4871/.

Khatib:2020:ADS

[Kha20] Mohamed M. Khatib. Automated discovery of symbolic approximation formulae using genetic programming. International Journal of Computer Applications, 176(??):29–34, April 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume176/number13/
31262-2020920053/.

Kamel:2019:LCR

[KHAE19] Walaa Esame Kamel, Ashraf Hafez, and Hala Mohamed Abd-Elkader. A low cost remotely analysis of ECG signal based on LabVIEW. International Journal of Computer Applications, 177(??):42–46, December 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number21/31024-2019919680/.

Kirui:2013:HCI

[KHK13] Clement Kirui, Li Hong, and Edgar Kirui. Handling class imbalance in mobile telecoms customer churn prediction. International Journal of Computer Applications, 72 (??):7–13, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number23/12680-9446/.

Kashyap:2012:MBT

[KHSE12] Abhinav Kashyap, Thomas Holzer, Shahryar Sarkani, and Tim Eveleigh. Model based testing for software systems: an application of Markov modulated Markov process. International Journal of Computer Applications, 46 (??):13–20, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number14/6976-9491/.

Khan:2015:MAP

[KIA15] Risala Tasin Khan, Md. Imdadul Islam, and M. R. Amin. Medium access probability of cognitive radio network at 1900 MHz and 2100 MHz. International Journal of Computer Applications, 124(??):42–45, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number5/22103-2015905490/.

Khaliluzzaman:2015:SCA

[KIH15] Md. Khaliluzzaman, Md. Monirul Islam, and Md. Monjur Hasan. Sufficient condition and algorithm for Hamiltonian in 3-connected 3-regular planar bipartite graph. International Journal of Computer Applications, 117(??):6–10, May 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume117/number11/20596-
3096/.

Kaur:2012:ACB

[KJ12a] Livjeet Kaur and Mohinder Pal Joshi. Analysis of chemotaxis in bacterial foraging optimization algorithm. International Journal of Computer Applications, 46(??):18–23, May 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume46/number4/6895-

Kumar:2012:EAS

[KJ12b] Rakesh Kumar and Jyotishree. Effect of annealing selection operators in genetic algorithms on benchmark test functions. International Journal of Computer Applications, 40 (??):38–46, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume40/
number3/5027-7174/.

Karami:2014:CDP

[KJ14] Amin Karami and Ronnie Johansson. Choosing DB-SCAN parameters automatically using differential evolution. International Journal of Computer Applications, 91 (??):1–11, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume91/
number7/15890-5059/.

Kumar:2015:SER

[KJ15] Avenash Kumar and Tahseen A. Jilani. A simple and efficient roadmap to process fingerprint images in frequency domain. International Journal of Computer Applications, 112(??):19–25, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number4/19655-1292/.

Kannan:2015:IXO

[KJA15] R. Kannan, D. Jayaraman, and S. Aravindhan. Influence of xylenol orange dye on optical, thermal, dielectric, laser damage threshold and mechanical properties of L-alanine thiourea (LATU) single crystals. International Journal of Computer Applications, 113(??):30–39, March 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume113/number6/19832-1684/.

Kudesia:2015:EAB

[KJJ15] Richa Kudesia, Ankur Jain, and Bharti Jain. Effect analysis of black hole attack of AODV protocol in MANET using table driven approach. International Journal of Computer Applications, 119(??):28–35, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number9/21097-3800/.

Khan:2012:IDB

[KJK12] Shamaila Khan, Anurag Jain, and Ashish Khare. Image denoising based on adaptive wavelet thresholding by using various shrinkage methods under different noise condition. International Journal of Computer Applications, 59 (??):13–17, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number20/9818-4397/.

Kumar:2012:OCS

[KJKK12] Deepak Kumar, P. C. Jha, P. K. Kapur, and U. Dinesh Kumar. Optimal component selection problem for cots based software system under consensus recovery block scheme: a goal programming approach. International Journal of Computer Applications, 47(??):9–14, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume47/number4/7174-9814/.

Kazi:2021:TBN

[KJKM21] Taaha Kazi, Sameer Joshi, Steeve Kaitharath, and Imran Ali Mirza. Transformer based neural joke generator. International Journal of Computer Applications, 183(??):1–4, October 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume183/number34/
32150-2021921724/.

Kichonge:2015:PTE

[KJTM15] Baraka Kichonge, Geoffrey R. John, Thomas Tesha, and Iddi S. N. Mkilaha. Prediction of Tanzanian energy demand using support vector machine for regression (SVR). International Journal of Computer Applications, 109(??): 34–39, January 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume109/number3/19172-0643/.

Kalyani:2010:ARU

[KK10a] P. Kalyani and M. Karnan. Attribute reduction using forward selection and relative reduct algorithm. International Journal of Computer Applications, 11(??):8– 12, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number3/1564-1499/.

Kulkarni:2010:PCS

[KK10b] Vaishali Kulkarni and H. B. Kekre. Performance comparison of speaker recognition using vector quantization by LBG and KFCG. International Journal of Computer Applications, 3(??):32–37, July 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume3/number10/773-1086/.

Kumawat:2010:OTA

[KK10c] Santosh Kumawat and Ajay Khunteta. Operational transformation algorithms for wide-area collaborative applications. International Journal of Computer Applications, 6(??):25–30, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume6/number7/1088-1420/.

Khan:2011:DIV

[KK11] Muzammil Khan and Sarwar Shah Khan. Data and information visualization methods, and interactive mechanisms: a survey. International Journal of Computer Applications, 34 (??):1–14, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume34/
number1/4061-5722/.

Kejgir:2012:LWT

[KK12a] Sushma G. Kejgir and Manesh Kokare. Lifting wavelet transform with singular value decomposition for robust digital image watermarking. International Journal of Computer Applications, 39(??):10–18, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number18/5078-7193/.

Kumar:2012:WPW

[KK12b] Vijay Kumar and Ashwani Kush. Worm Secure protocol for wormhole protection in AODV routing protocol. International Journal of Computer Applications, 44(??): 15–21, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number4/6251-8394/.

Kampf:2013:HTL

[KK13a] Mirko Kämpf and Jan W. Kantelhardt. Hadoop.TS: Large-scale time-series processing. International Journal of Computer Applications, 74(??):1–8, July 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume74/number17/12974-0233/.

Kumar:2013:RHN

[KK13b] Chandan Jyoti Kumar and Sanjib Kumar Kalita. Recognition of handwritten numerals of Manipuri script. International Journal of Computer Applications, 84(??):1– 5, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number17/14674-2835/.

Kumar:2013:ESV

[KK13c] Opinder Kumar and Mohinder Kumar. Enhancing security in VANET in terms of confidentiality and authentication. International Journal of Computer Applications, 67 (??):26–29, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number24/11738-7371/.

Kumar:2014:INO

[KK14] T Vinod Kumar and I Satish Kumar. Implementation of novel optimization algorithm for optimal placement and sizing of capacitor banks in radial distribution systems for power loss minimization and net savings maximization. International Journal of Computer Applications, 108(??):19–25, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number11/18955-0233/.

Khurana:2015:VDL

[KK15a] Mehak Khurana and Meena Kumari. Variants of differential and linear cryptanalysis. International Journal of Computer Applications, 131(??):20–28, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume131/number18/23549-2015907534/.

Kulshrestha:2015:SHB

[KK15b] Preeti Kulshrestha and Arun Kumar. Secret handshakes based on shortened versions of DSS. International Journal of Computer Applications, 132(??):13–18, December 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume132/number15/23669-
2015907565/.

Kuri:2015:ASR

[KK15c] Saumitra Kumar Kuri and Jayant V. Kulkarni. Automated segmentation of retinal blood vessels using optimized Gabor filter with local entropy thresholding. International Journal of Computer Applications, 114(??):37–42, March 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume114/number11/20026-
2112/.

Kaur:2016:AIF

[KK16a] Ramandeep Kaur and Sukhpreet Kaur. An approach for image fusion using PCA and genetic algorithm. International Journal of Computer Applications, 145(??):54–59, July 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume145/number6/25286-2016910816/

.

Kumar:2016:OPA

[KK16b] Divya Kumar and Arun Kumar. Outage probability analysis of multihop-multicast cooperative system. International Journal of Computer Applications, 148(??):1–5, August 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume148/number5/25750-2016911104/

.

Khawaldeh:2017:DOB

[KK17a] Mohammad K. Khawaldeh and Saed Khawaldeh. Design of optimal beamforming using particles swarm optimization. International Journal of Computer Applications, 174(??):25–32, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number2/28381-2017915327/.

Kumari:2017:RDA

[KK17b] Ch. Kusuma Kumari and G. Radha Kumari. Robust direction of arrival estimation using subspace methods. International Journal of Computer Applications, 159(??): 24–29, February 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume159/number1/26968-2017912686/.

Kaleem:2018:EAF

[KK18] Abdullah Mohammed Kaleem and Rajendra D. Kokate. An efficient approach for fetal ECG extraction. International Journal of Computer Applications, 182(??):1–5, December 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number33/30239-
2018918258/.

Karve:2020:MUP

[KK20] Dhruv Karve and Farhan Kapadia. Multi-UAV path planning using modified Dijkstra’s algorithm. International Journal of Computer Applications, 175(??):26–33, October 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume175/number28/31629-
2020920816/.

Kumar:2022:SAM

[KK22] Manoj Kumar and Shiv Kumar. Stochastic analysis of Markov chain model with different types of repair facility. International Journal of Computer Applications, 184(??):34–40, August 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number25/32470-2022922306/.

Karthik:2012:LSM

[KKBS12] S. Karthik, Hemanth V K, V. Balaji, and K. P. Soman. Level set methodology for Tamil document image binarization and segmentation. International Journal of Computer Applications, 39(??):7–12, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number9/4846-7117/.

Khoukhi:2012:FDC

[KKDC12] A. Khoukhi, H. Khalid, R. Doraiswami, and L. Cheded. Fault detection and classification using Kalman filter and hybrid neuro-fuzzy systems. International Journal of Computer Applications, 45(??):7–14, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number22/7079-9312/.

Kekre:2012:SIU

[KKGG12] H. B. Kekre, Vaishali Kulkarni, Prashant Gaikar, and Nishant Gupta. Speaker identification using spectrograms of varying frame sizes. International Journal of Computer Applications, 50(??):27–33, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number20/7921-1228/.

Khalid:2016:PIA

[KKI16] M. Khalid, Fareeha Sami Khan, and Azhar Iqbal. Perturbationiteration algorithm to solve fractional giving up smoking mathematical model. International Journal of Computer Applications, 142(??):1–6, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number9/24921-2016909891/.

Koceski:2012:DEC

[KKK12] Saso Koceski, Natasa Koceska, and Ivica Kocev. Design and evaluation of a cell phone pointing interface for interaction with large projector based displays. International Journal of Computer Applications, 51(??):27–32, August 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume51/number3/8024-1157/.

Kumar:2013:HVA

[KKK13] Harish Kumar, Kamaldeep Kaur, and Gurpreet Kaur. Heart variability analysis by using non-linear techniques and their comparison. International Journal of Computer Applications, 65(??):33–36, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume65/number20/11043-6385/.

Kora:2018:HAD

[KKM18] Padmavathi Kora, Ch. Usha Kumari, and K. Meenakshi. Heart arhythmia detection using wavelet coherence and firefly algorithm. International Journal of Computer Applications, 179(??):1–8, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number27/29125-2018916539/.

Kishore:2014:TIC

[KKR^{+}14] P. V. V. Kishore, Sampath Kumar, Y. Sahithi Reddy, J. K. Chaithanya, and M. V. D. Prasad. Tumor identification in CT medical images using semi automatic active contour models. International Journal of Computer Applications, 96(??):8–14, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number17/16884-6892/.

K:2018:SPN

[KKR18] Vinay K, S. M. Dilip Kumar, and Venugopal K R. Scheduling preemptive and non-preemptive tasks of scientific workflows using hybrid instances in cloud environment. International Journal of Computer Applications, 181(??):1–6, September 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume181/number15/29895-
2018917817/.

Kaur:2014:DHS

[KKS14] Gurleen Kaur, Arvind Kumar, and Jatinder Singh. Design of high speed full adder using improved differential split logic technique for 130nm technology and its implementation in making ALU. International Journal of Computer Applications, 96(??):40–47, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume96/number18/16898-6952/.

Kumaravel:2012:HFI

[KKSS12] M. Kumaravel, S. Karthik, P. Sivraj, and K. P. Soman. Human face image segmentation using level set methodology. International Journal of Computer Applications, 44 (??):16–22, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number12/6315-8658/.

Karpagam:2015:LPV

[KKT15] G. Karpagam, E. Konguvel, and M. Thangamani. A low power VLSI implementation of STTRAM based TCAM for high speed switching circuits. International Journal of Computer Applications, 115(??):38–42, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume115/number7/20167-2299/.

Kuhn:2013:ATH

[KLS13] Matthias Kuhn, Joachim Lippold, and Horst Salzwedel. Automatic transformation of hospital processes into an executable model with EPML. International Journal of Computer Applications, 80(??):20–30, October 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume80/number9/13890-1818/.

Kumar:2012:SMC

[KM12] Ashish Kumar and S. C. Malik. Stochastic modeling of a computer system with priority to PM over s/W replacement subject to maximum operation and repair times. International Journal of Computer Applications, 43(??): 27–34, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume43/
number3/6084-8228/.

Kamble:2013:DTB

[KM13a] Priyanka N. Kamble and Parikshit N. Mahalle. Decision theory based auto-delegation (DTA-d) scheme for ubiquitous computing. International Journal of Computer Applications, 79(??):1–6, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume79/number7/13750-1571/.

Kumar:2013:CCU

[KM13b] Sarvesh Kumar and Omkara Murthy. Cloud computing for universities: a prototype suggestion and use of cloud computing in academic institutions. International Journal of Computer Applications, 70(??):1–6, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume70/number14/12027-7633/.

Kahoul:2015:APM

[KM15a] Nabil Kahoul and Mounira Mekki. Adaptive P&O MPPT technique for photovoltaic buck-boost converter system. International Journal of Computer Applications, 112(??): 23–27, February 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume112/number12/19719-0773/.

Kalloor:2015:VRH

[KM15b] John Kalloor and S. Meyyappan. VHDL realization of hybrid control strategy for a nonlinear system. International Journal of Computer Applications, 113(??):17–21, March 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume113/number12/19878-

Kaur:2015:PEN

[KM15c] Manmeet Kaur and Kamna Mahajan. Performance evaluation of natural language text watermarking using encryption techniques. International Journal of Computer Applications, 129(??):22–28, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume129/number3/23055-2015906789/.

Khodade:2015:UTW

[KM15d] Sayali Khodade and Nilav Mukherjee. Unsupervised technique for Web data extraction: Trinity. International Journal of Computer Applications, 115(??):43–48, April 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume115/number19/20263-
2668/.

Kumar:2015:CIT

[KM15e] Kalagotla Satish Kumar and T. Sita Mahalakshmi. Computational intelligence techniques for classification of cancer data. International Journal of Computer Applications, 132(??):14–21, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number10/23629-2015907607/.

Kadhim:2016:IVF

[KM16a] Zena K. Kadhim and Hadi O. Mery. Influence of vibration on free convection heat transfer from sinusoidal surface. International Journal of Computer Applications, 136(??):1–6, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume136/number4/24138-2016908252/.

Kavecky:2016:STA

[KM16b] Slavomír KaveckÝ and Penka Martincová. A survey on trust aware security and scheduling in traditional and ad hoc grids. International Journal of Computer Applications, 133(??):1–13, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume133/number12/23835-2016908058/.

Khormehr:2016:NFE

[KM16c] Arash Khormehr and Vafa Maihami. A novel fuzzy expert system design for predicting heart diseases. International Journal of Computer Applications, 138(??): 33–38, March 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume138/number4/24369-2016908789/.

Kirimi:2016:ADM

[KM16d] John M. Kirimi and Christopher A. Moturi. Application of data mining classification in employee performance prediction. International Journal of Computer Applications, 146(??):28–35, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number7/25412-2016910883/.

Kagoiya:2017:HAN

[KM17] Kenneth Kagoiya and Elijah Mwangi. A hybrid and adaptive non-local means wavelet based MRI denoising method with bilateral filter enhancement. International Journal of Computer Applications, 166(??):1–7, May 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume166/number10/27702-2017914121/.

Kerrouche:2014:SPC

[KMB14] K. D. E. Kerrouche, A. Mezouar, and L. Boumedien. The suitable power control of wind energy conversion system based doubly fed induction generator. International Journal of Computer Applications, 87(??):35– 44, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number3/15191-3570/.

Kwao:2020:ISM

[KMG20] Lazarus Kwao, Richard Millham, and Enoch Opanin Gyamfi. An integrated success model for adopting biometric authentication technique for District Health Information Management System 2, Ghana. International Journal of Computer Applications, 177(??):1–16, February 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number40/31168-2020919686/.

Kimande:2021:AIT

[KMG21] Jethro Kimande, Elijah Mwangi, and Michael Gatari. Application of infrared thermography in fault detection and preventive maintenance in three-phase distribution transformers. International Journal of Computer Applications, 174(??):7–12, January 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number12/31728-2021920995/.

Keshtgary:2012:AME

[KMMM12a] Manijeh Keshtgary, Mohammad Reza Mansouri, Reza Mohammadi, and Mohammad Mahmoudi. Analytical model of energy consumption in cluster based wireless sensor networks with data aggregation using m/m/1 queuing model. International Journal of Computer Applications, 53(??):1– 6, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number13/8478-2410/.

Keshtgary:2012:ECE

[KMMM12b] Manijeh Keshtgary, Reza Mohammadi, Mohammad Mahmoudi, and Mohammad Reza Mansouri. Energy consumption estimation in cluster based underwater wireless sensor networks using m/m/1 queuing model. International Journal of Computer Applications, 43(??):6–10, April 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume43/number24/6435-8362/.

Khatri:2012:OPP

[KMS12] Neha Khatri, Vinod Mishra, and Rama Gopal V Sarepaka. Optimization of process parameters to achieve nano level surface quality on polycarbonate. International Journal of Computer Applications, 48(??):39–44, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume48/number13/7412-0507/.

Kekre:2014:CVF

[KMS14] H. B. Kekre, Aditi Mehta, and Paulami Shah. Comparison of various feature extraction techniques in CBIR using statistical parameters. International Journal of Computer Applications, 93(??):1–6, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume93/number18/16432-5600/.

Kavitha:2015:HSH

[KMS15] R. Kavitha, M. V. Ramana Murthy, and Mahaboob Sharief Shaik. Hybrid scheduling for heterogeneous asymmetric mobile communication. International Journal of Computer Applications, 119(??):36–39, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number19/21179-4204/.

Kadam:2019:BHP

[KMS19] Sarika Kadam, Akshata Meshram, and Sheetal Suryawanshi. Blockchain for healthcare: Privacy preserving medical record. International Journal of Computer Applications, 178(??):5–9, July 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number36/30773-2019919092/.

Kimbal:2024:DEG

[KMT^{+}24] Anthon Arie Kimbal, Olga Engelien Melo, Robby Tangkudung, Harson Kapoh, and Roby Lumbu. Development of 3D educational game with Pancasila theme based on Android. International Journal of Computer Applications, 186(??):34–43, January 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number1/33040-2024923342/.

Kumar:2011:DIS

[KN11] M. Sunil Kumar and P. Neelima. Design and implementation of scalable, fully distributed Web crawler for a Web search engine. International Journal of Computer Applications, 15(??):8–13, February 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume15/number7/1963-2629/.

Kamal:2012:NFA

[KN12a] Md. Sarwar Kamal and Sonia Farhana Nimmy. New fuzzy algorithm to inspect adenoids. International Journal of Computer Applications, 43(??):6–13, April 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume43/number21/6385-8463/.

Khan:2012:CMP

[KN12b] Iqbal A. Khan and Ahmed M. Nahhas. Current mode programmable analog modules using low voltage digitally controlled CMOS CCII. International Journal of Computer Applications, 48(??):38–44, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume48/number4/7340-0200/.

Kokil:2012:DCP

[KN12c] Asha Kokil and Leckraj Nagowah. Developing cross platform secured mobile widgets using subject-role based access control mechanism. International Journal of Computer Applications, 57(??):34–41, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume57/number22/9423-3839/.

Kudlikar:2015:RPW

[KN15] Anuradha K. Kudlikar and Meghana B. Nagori. Refinement in personalize Web search system with privacy protection. International Journal of Computer Applications, 117(??):1–6, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number6/20555-2935/.

Khakshour:2017:RNF

[KN17a] Meisam Khakshour and Nima Nourani. Robust nonlinear fuzzy formation control of unmanned quadrotors. International Journal of Computer Applications, 172(??): 18–24, August 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume172/number7/28263-2017915179/.

Khatun:2017:ATS

[KN17b] Afrina Khatun and Naushin Nower. Automatic test scenario generation technique for medical cyber physical systems. International Journal of Computer Applications, 176(??):1–8, October 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number5/28545-2017915585/.

Karimpour:2012:WSD

[KNA12] Jaber Karimpour, Ali A. Noroozi, and Somayeh Alizadeh. Web spam detection by learning from small labeled samples. International Journal of Computer Applications, 50 (??):1–5, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume50/
number21/7924-0993/.

Khare:2010:NMS

[KNGT10] Anubhuti Khare, Rajesh Nema, Puran Gour, and Rajeev Kumar Thakur. New multiband e-shape microstrip patch antenna on RT DUROID 5880 substrate and RO4003 substrate for pervasive wireless communication. International Journal of Computer Applications, 9(??):6–14, November 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume9/number8/1408-1901/

.

Khatoon:2014:EVA

[KNI14] Shahida Khatoon, Kaukab Naz, and Ibraheem. Eigen value analysis of optimal controller design of wheeled autonomous mobile robot. International Journal of Computer Applications, 103(??):13–18, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number17/18293-9211/.

Kumar:2015:ODR

[KNR15] Parvesh Kumar, Shiv Narayan, and Jetesh Raheja. Optimal design of robust fractional order PID for the flight control system. International Journal of Computer Applications, 128(??):31–35, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number14/22944-2015906759/.

Khatun:2017:UDS

[KNS17] Afrina Khatun, Naushin Nower, and Kazi Sakib. UML diagrams and source based automatic test suite regeneration for improving state model coverage. International Journal of Computer Applications, 176(??):1–8, October 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number7/28564-2017915633/.

Kumar:2020:CTE

[KNS20] B. K. Praveen Kumar, Gundamaraju Nithya, and K. Santhi Sree. Covid-19: a tentative estimation of fatality rates using random forest algorithm. International Journal of Computer Applications, 176(??):17–22, May 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number23/31338-2020920197/.

Kadir:2011:NNA

[KNSS11] A. Kadir, L. E. Nugroho, A. Susanto, and P. I. Santosa. Neural network application on foliage plant identification. International Journal of Computer Applications, 29(??):15– 22, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume29/
number9/3592-4981/.

Kumar:2011:MOR

[KNT11] D. Kranthi Kumar, S. K. Nagar, and J. P. Tiwari. Model order reduction of interval systems using mihailov criterion and cauer second form. International Journal of Computer Applications, 32(??):17–21, October 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume32/number6/3907-5483/.

Kumaravel:2010:DCP

[KNV10] S. Kumaravel, P. Neelamegam, and R. Vasumathi. Distributed chloride prediction system using neural network and PIC18F452 microcontrollers in water analysis. International Journal of Computer Applications, 8(??):15–20, October 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume8/number14/1316-
1800/.

Kongmunvattana:2015:LBP

[Kon15] Angkul Kongmunvattana. Load balancing for parallel motif discoveries. International Journal of Computer Applications, 124(??):29–34, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number13/22166-2015905745/.

Kortas:2023:IAC

[Kor23] Nawel Kortas. IoT and the autonomy of communicating objects lead toward an adapted future. International Journal of Computer Applications, 185(??):18–24, April 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number4/32693-2023922692/.

Kothari:2010:DTP

[KP10a] Amit D. Kothari and Ashok R. Patel. On demand temporary parallel route recovery for frequent link failure in adhoc networks. International Journal of Computer Applications, 11(??):1–6, December 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume11/number11/1630-2191/.

Kumar:2010:EMR

[KP10b] Arvind Kumar and Km. Pooja. An effective method for ranking of changed Web pages in incremental crawler. International Journal of Computer Applications, 8(??):38–41, October 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume8/number7/1219-
1760/.

Kharche:2013:PTC

[KP13] Shubhangi Kharche and Pramod B. Patil. Performance test comparison of Internet layer protocols on CentOS and 2007 for data, DNS and voice traffic. International Journal of Computer Applications, 62(??):8–13, January 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume62/number2/10050-4637/.

Kaur:2015:WCT

[KP15a] Maninder Kaur and Pooja. Wavelet and curvelet transformation based image fusion with ANFIS and SVM. International Journal of Computer Applications, 121(??):13–19, July 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume121/number14/
21607-4639/.

Kowsalya:2015:EAP

[KP15b] S. Kowsalya and D. Shanmuga Priyaa. An enhanced approach for preprocessing of mammogram images using inverse Daubechies wavelet transform and non-linear diffusion. International Journal of Computer Applications, 130(??):6–12, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number5/23203-2015906955/.

Katiyar:2016:MAI

[KP16] Shivangi Katiyar and Devendra Prasad. Mobile agent initiated energy efficient data aggregation in WSNs: MAEDA. International Journal of Computer Applications, 140(??):10– 15, April 2016. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume140/number1/24557-
2016909170/.

Kumar:2014:MME

[KPAB14] Surender Kumar, M. Prateek, N. J. Ahuja, and Bharat Bhushan. MEECDA: Multihop energy efficient clustering and data aggregation protocol for HWSN. International Journal of Computer Applications, 88(??):28– 35, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume88/
number9/15383-4047/.

Kumar:2012:AHV

[KPB12] K. Kishore Kumar, Movva Pavani, and V. Seshu Babu. Adaptive hybrid video watermarking using WBCT and NMF. International Journal of Computer Applications, 56 (??):18–25, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number16/8975-3162/.

Kadam:2016:ISE

[KPC16] Sapna R. Kadam, B. M. Patil, and V. M. Chandode. Improve speed efficiency and maintain data integrity of dynamic big data by using map reduce. International Journal of Computer Applications, 137(??):5–12, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume137/number8/24293-2016908687/.

Kekre:2013:SBU

[KPG13] H. B. Kekre, Archana B. Patankar, and Hema Ramesh Galiyal. Segmentation of blast using vector quantization technique. International Journal of Computer Applications, 72(??):20–23, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number15/12571-9186/.

Kavitha:2016:FFS

[KPM16] D. Kavitha, V. Kamakshi Prasad, and J. V. R. Murthy. Finding frequent subgraphs in a single graph based on symmetry. International Journal of Computer Applications, 146(??):5–8, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number11/25440-2016910895/.

Kumar:2014:RIS

[KPR14] Pushpendra Kumar, Rekha Pandit, and Vineet Richhariya. Retinal image segmentation by using gradient descent method. International Journal of Computer Applications, 86(??):1–7, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number10/15018-3306/.

Kannimuthu:2011:IIF

[KPS11] S. Kannimuthu, K. Premalatha, and S. Shankar. iFUM — improved fast utility mining. International Journal of Computer Applications, 27(??):32–36, August 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume27/number11/3343-4602/.

Kumar:2012:XBQ

[KPV12] Manish Kumar, Monalisa Panigrahi, and Shekhar Verma. XQuery based query processing architecture in wireless sensor networks. International Journal of Computer Applications, 43(??):5–10, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number23/6417-8856/.

Kumar:2013:FRD

[KPV13] Padmaja Vijay Kumar, M. N. Giri Prasad, and Padmaja K. V. Face recognition of database of compressed images using local binary patterns. International Journal of Computer Applications, 74(??):10–17, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume74/number16/12968-9591/.

Kumar:2010:BFA

[KR10] M. Senthil Kumar and P. Renuga. Bacterial foraging algorithm based enhancement of voltage profile and minimization of losses using thyristor controlled series capacitor (TCSC). International Journal of Computer Applications, 7(??):21–27, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume7/number2/1137-1488/.

Kumar:2011:MLS

[KR11a] S. Britto R. Kumar and S. Albert Rabara. Multi-level security for integrated financial mobile Web services using federated ESB. International Journal of Computer Applications, 32(??):39–45, October 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume32/number10/3948-5292/.

Kuntamalla:2011:EAN

[KR11b] Srinivas Kuntamalla and L. Ram Gopal Reddy. The effect of aging on nonlinearity and stochastic nature of heart rate variability signal computed using delay vector variance method. International Journal of Computer Applications, 14(??):40–44, January 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume14/number5/1837-2470/.

Kottawar:2014:CFE

[KR14] Vinayak Gajanan Kottawar and A. M. Rajurkar. Color feature extraction in content based image retrieval based on quaternion space. International Journal of Computer Applications, 107(??):17–21, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number20/18867-0286/.

Karthikeyan:2015:GDF

[KR15] V. Karthikeyan and M. Rajasekar. Generalized directable fuzzy automata. International Journal of Computer Applications, 131(??):1–5, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume131/number12/23498-2015907441/.

Kundu:2016:SAD

[KR16] Sujata Kundu and Chayan Ranjit. A systematic approach of data fusion technique in RFID sensor network using neurofuzzy technique. International Journal of Computer Applications, 139(??):7–14, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number9/24516-2016909045/.

Kapoor:2017:CDA

[KR17] Prashant Kapoor and Paresh Rawat. Comparison of detection accuracy and effect of JPEG2000 compression on iris recognition. International Journal of Computer Applications, 162(??):37–42, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume162/number4/27235-2017913281/.

Kiran:2013:CAR

[KRH13] K. V. D. Kiran, L. S. S. Reddy, and N. Lakshmi Haritha. A compartive analysis on risk assessment information security models. International Journal of Computer Applications, 82(??):41–47, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume82/number9/14148-2314/.

Kumar:2012:OLR

[KRM12] M. Senthil Kumar, P. Renuga, and K. Maharaja. Optimal location and rating of thyristor controlled series capacitor for enhancement of voltage stability using fast voltage stability index (FVSI) approach. International Journal of Computer Applications, 46(??):19–23, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume46/number10/6944-9232/.

Kar:2010:MSE

[KRMB10] R. Kar, K. Ramakrishna, Ashis K. Mal, and A. K. Bhattacharjee. A model for slew evaluation for on-chip RC interconnects using gamma distribution function. International Journal of Computer Applications, 1(??):68–73, February 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume1/number10/219-368/.

Kashyap:2012:ELB

[KRP12] Indu Kashyap, R. K. Rathy, and Diwaker Pandey. An efficient location based reactive multi-path routing protocol for Manet. International Journal of Computer Applications, 40(??):24–29, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume40/number9/4994-7258/.

K:2011:PAF

[KRR11] Seetharam K, Venkata Subba Reddy, and Laxmi B. Rananavare. PCA analysis of few parameters role in software development. International Journal of Computer Applications, 14(??):15–20, February 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume14/number6/1889-2506/.

Kumaran:2013:NFA

[KRR13] J. Kumaran, G. Ravi, and R. Rajkumar. Neuralfuzzy approach for power load forecasting analysis. International Journal of Computer Applications, 69(??): 31–35, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number16/12048-8116/.

Kumar:2015:SSA

[KRR15] G. Rajendra Kumar, Kuda Nageswara Rao, and K. Srinivas Rao. Steady state analysis of a three node tandem communication network model with DBA having homogeneous compound Poisson direct bulk arrivals at node 1 and node 2. International Journal of Computer Applications, 129(??):8–16, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume129/number4/23059-2015906799/.

Kar:2012:NAG

[KRSB12] Chinmoy Kar, Vineet Kumar Rakesh, Tapas Samanta, and Sreeparna Banerjee. A new approach to grid scheduling using random weighted genetic algorithm with fault tolerance strategy. International Journal of Computer Applications, 48(??):42–47, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume48/number23/7524-0632/.

Krishnamoorthy:2010:ICS

[KS10] Shivsubramani Krishnamoorthy and K. P. Soman. Implementation and comparative study of image fusion algorithms. International Journal of Computer Applications, 9(??):25–35, November 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume9/number2/1357-1832/.

Kanikar:2012:EAE

[KS12a] Prashasti Kanikar and Ketan Shah. An efficient approach for extraction of actionable association rules. International Journal of Computer Applications, 54(??):5– 10, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume54/
number11/8608-2458/.

Kmuar:2012:EDD

[KS12b] Santosh Kmuar and S. C. Sharma. Evaluation of direct and derived mobility metrics of mobility models and its impact on reactive routing protocols. International Journal of Computer Applications, 48(??):40–49, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume48/number21/7484-0590/.

Kulkarni:2013:FCB

[KS13a] U. V. Kulkarni and S. V. Shinde. A fuzzy classifier based on product and sum aggregation reasoning rule. International Journal of Computer Applications, 62(??):9– 14, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number5/10075-4687/.

Kumar:2013:OPP

[KS13b] Purshottam Kumar and Ranjit Singh. Optimizing process parameters of rotary furnace using bio-fuels: an interactive ANN approach. International Journal of Computer Applications, 62(??):25–31, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume62/number1/10046-4631/.

Kumari:2013:SBS

[KS13c] A. Charan Kumari and K. Srinivas. Search-based software requirements selection: a case study. International Journal of Computer Applications, 64(??):28–34,

February 2013.  CODEN ????  ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number21/10760-5715/.

Kalaiselvi:2014:KBS

[KS14a] T. Kalaiselvi and K. Somasundaram. Knowledge based self initializing FCM algorithms for fast segmentation of brain tissues in magnetic resonance images. International Journal of Computer Applications, 90(??):19–26, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume90/number14/15788-4511/.

Kaur:2014:ILU

[KS14b] Navjot Kaur and Gagandeep Singh. Improved leach using highest remaining energy first based cluster head selection. International Journal of Computer Applications, 103(??):1–6, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number7/18083-8935/.

Kumari:2014:EPR

[KS14c] Rozy Kumari and Narinder Sharma. Enhanced the performance of remote sensing color image segmentation by using L0 gradient minimization and DBPTGMF. International Journal of Computer Applications, 105(??):43–47, November 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume105/number16/18465-
9834/.

K:2015:RTH

[KS15a] Neelasagar K and K. Suresh. Real time 3D-handwritten character and gesture recognition for smartphone. International Journal of Computer Applications, 123(??):1–8, August 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume123/number13/
22023-2015905662/.

Kasana:2015:CEQ

[KS15b] Reena Kasana and Saurabh Sharma. Cost efficient query optimization in mobile environment. International Journal of Computer Applications, 120(??):30–38, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume120/number22/21394-4443/.

Kaur:2015:MSA

[KS15c] Navpreet Kaur and Jatinder Pal Sharma. Mobile sink and ant colony optimization based energy efficient routing algorithm. International Journal of Computer Applications, 121(??):23–31, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number1/21505-4377/.

Kaur:2015:FSI

[KS15d] Sarabpreet Kaur and J. S. Sahambi. A framework for segmentation of inhomogeneous live cell images using fractional derivatives and level set method. International Journal of Computer Applications, 127(??):1–8, October 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume127/number3/22706-2015906357/.

Kumar:2015:BSB

[KS15e] Sanjay Kumar and Manpreet Singh. Biometric security based intelligent e-voting system. International Journal of Computer Applications, 117(??):33–41, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume117/number4/20546-2918/.

Kaur:2016:OMA

[KS16a] Harveen Kaur and Mandeep Singh Sra. Optimization of mobile agent using genetic algorithm in wireless sensor network. International Journal of Computer Applications, 134(??):31–46, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number2/23890-2016907831/.

Kaur:2016:MFI

[KS16b] Kiranjeet Kaur and Gurinder Singh. Multi-focus image fusion in transform domain using steerable pyramids. International Journal of Computer Applications, 138(??):14–20, March 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume138/number5/24374-2016908826/.

Kaur:2016:PBO

[KS16c] Uttamjit Kaur and Gagandeep Singh. Predicting the behaviour of open source software using object oriented metrics. International Journal of Computer Applications, 150(??):10–22, September 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume150/number5/26088-2016911501/.

Khedr:2016:INR

[KS16d] Wael M. Khedr and Mohamed S. Sawah. An improved NAS-RIF algorithm based on wavelet packet. International Journal of Computer Applications, 135(??): 10–13, February 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume135/number8/24068-2016908416/.

Kumar:2016:RIS

[KS16e] Upendra Kumar and Omjeet Singh. Retinal image segmentation by using texture-based Gabor filter optimized by gradient descent followed by evolutionary algorithm.

International Journal of Computer Applications, 148(??): 37–47, August 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume148/number4/25749-2016911096/.

Kaur:2017:NIC

[KS17] Upinder Kaur and Pushpa Rani Suri. A new improved circular skip list with priority search. International Journal of Computer Applications, 161(??):7–14, March 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume161/number2/27118-2017913112/.

Kalyanaraman:2018:MWV

[KS18] R. Kalyanaraman and A. Sundaramoorthy. A Markovian working vacation queue with server state dependent arrival rate and with unreliable server. International Journal of Computer Applications, 179(??):1–7, May 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number39/29336-2018916926/.

Kakde:2020:OCC

[KSA20] Aditya Kakde, Durgansh Sharma, and Nitin Arora. Optimal classification of COVID-19: a transfer learning approach. International Journal of Computer Applications, 176(??):25– 31, May 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume176/number20/
31316-2020920165/.

Kadge:2017:MTS

[KSBM17] Sanam Kadge, Saniya Sayyed, Bhavesh Buwa, and Burhanuddin Madraswala. M-ticketing system using QR codes for Mumbai local. International Journal of Computer Applications, 162(??):17–24, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume162/number11/27287-2017913403/.

Kumar:2011:KDD

[KSC11] Pardeep Kumar, Vivek Kumar Sehgal, and Durg Singh Chauhan. Knowledge discovery in databases (KDD) with images: a novel approach toward image mining and processing. International Journal of Computer Applications, 27 (??):10–13, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume27/
number6/3307-4531/.

Kumar:2014:IOB

[KSK14] Sandeep Kumar, Vivek Kumar Sharma, and Rajani Kumari. Improved onlooker bee phase in artificial bee colony algorithm. International Journal of Computer Applications, 90 (??):31–39, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume90/
number6/15579-4304/.

Kalla:2023:PDI

[KSKS23] Dinesh Kalla, Fnu Samaah, Sivaraju Kuraku, and Nathan Smith. Phishing detection implementation using databricks and artificial intelligence. International Journal of Computer Applications, 185(??):1–11, May 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number11/32742-2023922764/.

Kuhn:2018:AMR

[KSL^{+}18] Matthias Kühn, Thomas Severin, Joachim Lippold, Horst Salzwedel, and Volker Nissen. Adaptive mutation rate at genetic algorithms with multi-chromosome representation in multi-department hospital process optimization. International Journal of Computer Applications, 182(??):41–52, October 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume182/number21/
30060-2018917980/.

Khattak:2012:EUS

[KSlAM12] Zubair Ahmad Khattak, Suziah Sulaiman, and Jamalul lail Ab. Manan. Evaluation of unified security, trust and privacy framework (UnifiedSTPF) for federated identity and access management (FIAM) mode. International Journal of Computer Applications, 54(??):12–19, September 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume54/number6/8569-2290/.

Kumar:2015:GAA

[KSMM15] Dilip Kumar, Bibhudatta Sahoo, Bhaskar Mondal, and Tarni Mandal. A genetic algorithmic approach for energy efficient task consolidation in cloud computing. International Journal of Computer Applications, 118(??):1–6, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume118/number2/20714-3066/.

Kaloko:2011:DDS

[KSP11] Bambang Sri Kaloko, Soebagio, and Mauridhi Hery Purnomo. Design and development of small electric vehicle using MAT-LAB/Simulink. International Journal of Computer Applications, 24(??):19–23, June 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume24/number6/2960-3940/.

K:2012:HCS

[KSRS12] Manjusha K, Sachin Kumar S, Jolly Rajendran, and K. P. Soman. Hindi character segmentation in document images using level set methods and non-linear diffusion. International Journal of Computer Applications, 44(??): 42–49, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number16/6351-8745/.

Kekre:2012:NCA

[KSS12] H. B. Kekre, Tanuja K. Sarode, and Jagruti K. Save. New clustering algorithm for vector quantization using Walsh sequence. International Journal of Computer Applications, 39 (??):4–9, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume39/
number1/4782-6985/.

Kuhn:2013:VMR

[KSS13] Matthias Kuhn, Thomas Severin, and Horst Salzwedel. Variable mutation rate at genetic algorithms: Introduction of chromosome fitness in connection with multi-chromosome representation. International Journal of Computer Applications, 72(??):31–38, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume72/number17/12636-9343/.

Kekre:2014:GET

[KSS14a] H. B. Kekre, Tanuja K. Sarode, and Jagruti K. Save. Generation of an effective training feature vector using VQ for classification of image database. International Journal of Computer Applications, 108(??):15–23, December 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume108/number18/19011-0511/.

Kishore:2014:NMS

[KSS^{+}14b] M. V. Kishore, G. Pandit Samuel, N. Aditya Sundar, M. Enayath Ali, and Y. Lalitha Varma. A novel methodology for secure communications and prevention of forgery attacks. International Journal of Computer Applications, 96(??):5–12, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number22/16924-6762/.

Kannan:2018:RTE

[KSS18] Jeyakumar Kannan, Ar Md Shanavas, and Sridhar Swaminathan. Real time event detection adopting incremental TF-IDF based LSH and event summary generation. International Journal of Computer Applications, 180(??):22–30, January 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume180/number13/28922-
2018916252/.

Kumar:2024:ESR

[KSSM24] Sachin Kumar, Shiwalika Sambyal, Sourabh Shastri, and Vibhakar Mansotra. Enhancing satellite remote sensing image classification using two layer convolutional neural network. International Journal of Computer Applications, 186(??):18– 23, July 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number30/
enhancing-satellite-remote-sensing-image-classification-
using-two-layer-convolutional-neural-network/.

Khan:2016:MCA

[KST16] Muzaffar Khan, Jaikaran Singh, and Mukesh Tiwari. A multiclassifier approach of EMG signal classification for diagnosis of neuromuscular disorders. International Journal of Computer Applications, 133(??):13–18, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number4/23773-2016907710/.

Khanwani:2010:AED

[KSV10] Payal Khanwani, Susmita Sridhar, and K. Vijaylakshmi. Automated event detection of epileptic spikes using neural networks. International Journal of Computer Applications, 2(??):14–17, June 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume2/number4/660-928/.

Khalid:2015:SWQ

[KSZK15] M. Khalid, Mariam Sultana, Faheem Zaidi, and Javed Khan. A solution for a water quality model in a uniform stream channel using new iterative method. International Journal of Computer Applications, 115(??):1–4, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume115/number6/20152-2310/.

Karthikeyan:2010:DDU

[KT10] B. Karthikeyan and M. Tamileniyan. Dynamic data update for intelligent speed adaptation (ISA) system. International Journal of Computer Applications, 11(??):8– 13, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number1/1549-2062/.

Khaire:2012:IED

[KT12a] Pushpajit A. Khaire and Nileshsingh V. Thakur. Image edge detection based on soft computing approach. International Journal of Computer Applications, 51(??):12– 14, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number8/8061-1426/.

Kumar:2012:NSU

[KT12b] Rajesh Kumar and I. M. Talwar. Network security using firewall and cryptographic authentication. International Journal of Computer Applications, 57(??):13– 19, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number23/9438-3834/.

Karim:2014:AAI

[KT14] Tajbia Karim and Tasnuva Tasneem. Analytical adjustment of image contrast. International Journal of Computer Applications, 98(??):44–49, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume98/number20/17303-7794/.

Kumar:2015:CCB

[KT15] Ashok S Kumar and Sudha T. Clustering and cooperative beamforming in green cognitive radio networks with relays. International Journal of Computer Applications, 121(??):1–6, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number21/21821-5066/.

Kaladzavi:2017:GSD

[KTA17] Guidedi Kaladzavi, Yaya Traore, and Ado Adamou Abba Ari. Geo-spatial domain ontology: the case of the socio-cultural infrastructures. International Journal of Computer Applications, 178(??):42–48, November 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number1/28642-2017915720/.

Kekre:2012:PCG

[KTBK12] H. B. Kekre, Sudeep D. Thepade, Varun K. Banura, and Sanchit Khandelwal. Performance comparison of gradient mask texture based image retrieval techniques using global and local hybrid wavelet transforms with ternary image maps. International Journal of Computer Applications, 46 (??):27–35, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number2/6881-9184/.

Kekre:2012:ICU

[KTDG12] H. B. Kekre, Sudeep Thepade, Rik Kamal Kumar Das, and Saurav Ghosh. Image classification using block truncation coding with assorted color spaces. International Journal of Computer Applications, 44(??):9–14, April 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume44/number6/6265-8418/.

Kacem:2012:CAD

[KTN12] Moez Hadj Kacem, Souhir Tounssi, and Rafik Neji. Control of an actuator DC energy-saving dedicated to the electric traction. International Journal of Computer Applications, 54(??):20–25, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume54/number10/8602-2368/.

Kacem:2013:OCF

[KTN13] Moez Hadj Kacem, Souhir Tounsi, and Rfik Neji. Optimization of cutting frequency electromagnetic interuptor. International Journal of Computer Applications, 67 (??):23–27, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number21/11520-7400/.

Kanaujiya:2010:ISS

[KTS10] Sunita Kanaujiya, S. P. Tripathi, and N. C. Sharma. Improving speed of the signature scanner using BMH algorithm. International Journal of Computer Applications, 11 (??):27–31, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number4/1569-2096/.

Khan:2017:BRD

[KTS17] Fahad Khan, Yu Tang, and Gong Sicheng. A bidirectional resonant DC-DC converter for electrical vehicle charging/discharging systems. International Journal of Computer Applications, 177(??):20–24, November 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number5/28623-2017915756/.

Kubba:2017:EMR

[Kub17] Hassan A. Kubba. An efficient and more reliable second order power flow solution method with interpolation technique. International Journal of Computer Applications, 161(??):17–27, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number1/27113-2017912940/.

Kumaresh:2010:EAD

[Kum10] N. Kumaresh. An empirical approach — distributed mobility management for target tracking in MANETs. International Journal of Computer Applications, 2(??):35–39, June 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume2/number6/671-943/.

Kumar:2012:MBF

[Kum12] K. S. Chidanand Kumar. Morphology based facial feature extraction and facial expression recognition for driver vigilance. International Journal of Computer Applications, 51 (??):17–24, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number2/8014-1142/.

Kaparang:2013:LSD

[KUM13a] Daniel Riano Kaparang, Wiranto H. Utomo, and Danny Manongga. Logistics system design based on service oriented architecture using Thomas Erl model. International Journal of Computer Applications, 68(??):10–19, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume68/number1/11542-6803/.

Kumar:2013:SDB

[Kum13b] Manish Kumar. Strategy for design and building multimedia data type. International Journal of Computer Applications, 73(??):50–52, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume73/number4/12733-9589/.

Kumar:2014:PAL

[Kum14a] Hitesh Kumar. Performance analysis of LTE and IEEE 802.16 WiMAX in terms of attenuation and BER via MAT-LAB; scenario based ROF amalgamation as backhaul technology. International Journal of Computer Applications, 97 (??):32–36, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number20/17127-7787/.

Kumar:2014:SPB

[Kum14b] Manoj Kumar. Shortest paths between two points in space for self managed digital systems. International Journal of Computer Applications, 100(??):1–3, August 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume100/number13/17582-8284/.

Kumar:2019:DIS

[Kum19] Manoj Kumar. Design of IIR systolic array architecture by using linear mapping technique. International Journal of Computer Applications, 182(??):14–19, February 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number39/30349-
2019918463/.

Kakollu:2014:NPC

[KV14] Ramesh Kakollu and E. Vargil Vijay. A novel pattern classification using granular reflex fuzzy min-max neural network. International Journal of Computer Applications, 107(??):29–33, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number19/18862-0562/.

Kaundal:2015:NAU

[KV15] Gurpreet Kaundal and Sheveta Vashisht. A novel approach using XMI log generation cryptography and mining techniques for to enhance security of data. International Journal of Computer Applications, 115(??):25–32, April 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume115/number10/20190-
2418/.

Kaur:2016:PAH

[KV16a] Gaganpreet Kaur and Gagandeep Kaur Virk. Performance analysis of hybrid TB-LEACH enegy based multihop protocol. International Journal of Computer Applications, 143(??):15–21, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume143/number5/25072-2016910163/.

Kaur:2016:CVP

[KV16b] Mandeep Kaur and Prince Verma. Comment volume prediction using regression. International Journal of Computer Applications, 151(??):1–9, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume151/number1/26194-2016910131/.

Khatavkar:2019:MPA

[KVK19] Vaibhav Khatavkar, Makarand Velankar, and Parag Kulkarni. Multi-perspective analysis of news articles using machine learning algorithms. International Journal of Computer Applications, 181(??):22–26, February 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number42/30344-2019918503/.

Kumar:2016:ADS

[KVKR16] Ajay Kumar, Rohit Vaid, Sandhya Katiyar, and Shumaila Rizwan. Adaptive data size compressed algorithm to reduce energy and provide route security using AOMDV protocol in WSN. International Journal of Computer Applications, 147(??):32–38, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume147/number2/25627-2016911003/.

Kaur:2010:RFE

[KVN10] Mandeep Kaur, Rajeev Vashisht, and Nirvair Neeru. Recognition of facial expressions with principal component analysis and singular value decomposition. International Journal of Computer Applications, 9(??):36–40, November 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume9/number12/1434-
1933/.

Kumar:2014:CAS

[KVR14] Rajeev Kumar, B. K. Verma, and Shyam Sunder Rastogi. Context-aware social popularity based recommender system. International Journal of Computer Applications, 92 (??):37–42, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume92/
number2/15985-4907/.

Keerthana:2014:AAF

[KVRS14] P. Keerthana, V. Vaithianathan, J. Raja, and R. Srinivasan. Analysis of active feedback and its influence on UWB low noise amplifier. International Journal of Computer Applications, 89(??):29–35, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume89/number18/15733-4645/.

Kumar:2016:ECC

[KVS16] Manoj Kumar, O. P. Verma, and Archana Saxena. Elliptic curve cryptography (ECC) based relational database watermarking. International Journal of Computer Applications, 154(??):20–28, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume154/number6/26496-2016912159/.

Kokare:2014:SNL

[KW14] Rohini Kokare and Kirti H. Wanjale. A survey of natural language query builder interface for structured databases using dependency parsing. International Journal of Computer Applications, 107(??):9–14, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number5/18745-9998/.

Kiragu:2016:CBR

[KW16] Mark K. Kiragu and Peter W. Waiganjo. Case based reasoning for treatment and management of diabetes. International Journal of Computer Applications, 145(??):20–29, July 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume145/number4/25266-2016910552/

.

Kwale:2016:TCC

[KWM16] F. M. Kwale, P. W. Wagacha, and A. Mwaura. A text clustering comparison methodology. International Journal of Computer Applications, 139(??):12–19, April 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume139/number13/24550-2016909515/.

Kessentini:2011:HSP

[KWS^{+}11] Amina Kessentini, Imen Werda, Amine Samet, Mohamed Ali Ben Ayed, and Nouri Masmoudi. H.264/SVC performance and encoder bit-stream analysis. International Journal of Computer Applications, 36(??):30–37, December 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume36/number8/4513-
6381/.

Kumar:2015:MOS

[KY15] Satrughan Kumar and Jigyendra Sen Yadav. Moving object segmentation in dynamic environment by reducing impulsive noise from background model. International Journal of Computer Applications, 118(??):43–48, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume118/number22/20881-3632/.

Kalui:2016:PPA

[KZ16a] Dorothy Kalui and Dezheng Zhang. Profile privacy aware framework for static objects participatory sensing. International Journal of Computer Applications, 153(??): 1–9, November 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume153/number9/26428-2016910805/.

Khater:2016:SSN

[KZ16b] Mostafa M. A. Khater and Emad H. M. Zahran. Soliton soltuions of nonlinear evolutions equation by using the extended exp(−ϕ(ξ)) expansion method. International Journal of Computer Applications, 145(??):1–5, July 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume145/number3/25255-2016910516/.

B:2013:PTM

[L13] Muralidhara B L. Parallel two master method to improve BLAST algorithm’s performance. International Journal of Computer Applications, 63(??):8–14, February 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume63/number1/10428-5097/.

Leabi:2015:EER

[LA15] Safa Khudair Leabi and Turki Younis Abdalla. Energy efficient routing protocol for maximizing lifetime in wireless sensor networks using fuzzy logic. International Journal of Computer Applications, 122(??):19–25, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume122/number1/21665-4740/.

Lal:2017:LWE

[LA17] Shyam Lal and Anavaruddin Ansari. Legendre wavelet expansion of a function f(x, y) and its approximation. International Journal of Computer Applications, 171(??):1–7, August 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume171/number6/28182-2017914819/.

Lotfy:2017:DOP

[LAA^{+}17] Ahmed Lotfy, Wagdy R. Anis, M. A. Atalla, Joseph V. M. Halim, and M. Abouelatta. Design an optimum PV system for the satellite technology using high efficiency solar cells. International Journal of Computer Applications, 168(??):6–11, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume168/number3/27853-2017914327/.

Lartey:2023:AML

[LAAI23] Benjamin Lartey, Kelvin Adrah, Frederick Adrah, and Joan Isichei. Application of machine learning for predicting the occurrence of nephropathy in diabetic patients. International Journal of Computer Applications, 185(??):38–43, July 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume185/number24/32843-
2023923002/.

Landry:2021:STS

[LABV21] Clara Landry, Catherine Abadie, François Bonnot, and Jean Vaillant. A spatio-temporal stochastic model for an emerging plant disease spread in a heterogeneous landscape. International Journal of Computer Applications, 174(??):1–7, January 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume174/number16/
31757-2021921051/.

Lahmiri:2011:CPS

[Lah11] Salim Lahmiri. A comparison of PNN and SVM for stock market trend prediction using economic and technical information. International Journal of Computer Applications, 29 (??):24–30, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume29/
number3/3545-4860/.

Lahiri:2016:ASU

[Lah16] Sounak Lahiri. Audio steganography using echo hiding in wavelet domain with pseudorandom sequence. International Journal of Computer Applications, 140(??):16–22, April 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume140/number2/24564-2016909223/

.

Lotfi:2014:AAH

[LAM14] Elaachak Lotfi, Belahbib Amine, and Bouhorma Mohammed. Application of analytic hierarchical process method for video game genre selection. International Journal of Computer Applications, 96(??):30–37, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume96/number16/16881-6888/.

Lashkari:2010:NNB

[Las10] AmirEhsan Lashkari. A neural network based method for brain abnormality detection in MR images using Gabor wavelets. International Journal of Computer Applications, 4(??):9–15, July 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume4/number7/841-1140/.

Lazaiz:2017:CFP

[LAZ17] Samih Lazaiz, Mohamed Aamri, and Omar Zakary. Caristi’s fixed point theorem and Ekeland’s variational principle for set valued mapping using the LZ-functions. International Journal of Computer Applications, 158(??): 1–6, January 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume158/number2/26877-2017912755/.

Litayem:2011:ILR

[LB11] Nabil Litayem and Slim Ben Saoud. Impact of the Linux realtime enhancements on the system performances for multicore Intel architectures. International Journal of Computer Applications, 17(??):17–23, March 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume17/number3/2202-2796/.

Lmati:2016:TAG

[LBA16] Imane Lmati, El Habib Benlahmar, and Naceur Achtaich. Towards adaptive generation of mathematical exercises. International Journal of Computer Applications, 147(??): 33–38, August 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume147/number6/25660-2016911138/.

Layeb:2011:HGA

[LBB11] Abdesslem Layeb, Imen Bensetira, and Kenza Bouaroudj. A hybrid genetic algorithm for RNA structural alignment. International Journal of Computer Applications, 19(??): 41–47, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume19/
number7/2370-3120/.

Laouamer:2012:MJV

[LBNP12] Lamri Laouamer, Abdelhamid Benhocine, Laurent Nana, and Anca Pascu. Motion JPEG video authentication based on quantization matrix watermarking: Application in robotics. International Journal of Computer Applications, 47(??):1–5, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number24/7501-0353/.

Lim:2017:ATR

[LC17a] Sang-Hyeok Lim and Tae-Ho Cho. Automatic threshold reset scheme using a double fuzzy system for improvement of detection rate in a probabilistic voting-based filtering scheme of WSNs. International Journal of Computer Applications, 176(??):38–43, October 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number1/28520-2017915513/.

Lim:2017:RVT

[LC17b] Sang-Hyeok Lim and Tae-Ho Cho. Report verification technique for improvement of the energy efficiency in a probabilistic voting-based filtering scheme of WSNs. International Journal of Computer Applications, 171(??): 21–25, August 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume171/number3/28161-2017914992/.

Lahby:2012:NON

[LCA12] Mohamed Lahby, Leghris Cherkaoui, and Abdellah Adib. New optimized network selection decision in heterogeneous wireless networks. International Journal of Computer Applications, 54(??):1–7, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume54/number16/8647-2550/.

Leiva-cochachin:2013:ELA

[LcCm13] Andres M. Leiva-cochachin and Fredy Chalco-mendoza. Embedding Linux with ability to analyze network traffic on a development board based on FPGA. International Journal of Computer Applications, 77(??):39– 44, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number17/13618-1430/.

LimaDeMenezes:2021:SGU

[LCP21] Moisés Lima De Menezes, Reinaldo Castro Souza, and José Francisco Pessanha. Scenarios generation using bootstrap in the multichannel singular spectrum analysis approach and PAR (p) structures: Application to affluent natural energy. International Journal of Computer Applications, 174(??):15–24, January 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number13/31737-2021921018/.

Liu:2022:ANN

[LDB^{+}22] Haibo Liu, Tessa De Korte, Sylvain Bernasconi, Christophe Bleunven, Damiano Lombardi, and Muriel Boulakia. Artificial neural network comparison on hERG channel blockade detection. International Journal of Computer Applications, 184(??):1–9, May 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number14/32389-2022922119/.

Lagmiri:2013:CLV

[LEE13] Souad Najoua Lagmiri, El Houssine El Mazoudi, and Nouredinne Elalami. Control of Lotka–Volterra three species system via a high gain observer design. International Journal of Computer Applications, 77(??):1– 6, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number15/13556-1378/.

B:2015:EAS

[LEO15] Muhammad-Bello B L, Alhassan M E, and Ganiyu S O. An enhanced ATM security system using second-level authentication. International Journal of Computer Applications, 111(??):8–15, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number5/19533-1181/.

Lohani:2011:VSC

[LG11] Rajesh B. Lohani and Jaya V. Gaitonde. I-V and switching characteristics of back illuminated OPFET using finite difference methods. International Journal of Computer Applications, 31(??):5–11, October 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume31/number4/3810-5261/.

V:2013:INI

[LG13a] Jaya V L and R. Gopikakumari. IEM: a new image enhancement metric for contrast and sharpness measurements. International Journal of Computer Applications, 79 (??):1–9, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number9/13766-1620/.

B:2013:CAN

[LG13b] Muralidhara B L and Chitty Babu G. Centralized admission: a novel student-centric e-governance process. International Journal of Computer Applications, 66(??):41– 46, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number23/11260-6556/.

Labed:2011:MHP

[LGC11] Said Labed, Amira Gherboudj, and Salim Chikhi. A modified hybrid particle swarm optimization algorithm for multidimensional knapsack problem. International Journal of Computer Applications, 34(??):11–16, November 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume34/number2/4070-5586/.

Laajimi:2012:LPV

[LGM12a] Radwene Laajimi, Nawfil Gueddah, and Mohamed Masmoudi. Low power variable gain amplifier with bandwidth of 80—300 MHz using for sigma-delta analogue to digital converter in wireless sensor receiver. International Journal of Computer Applications, 43(??):35–37, April 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume43/number8/6126-8344/.

Laajimi:2012:NDM

[LGM12b] Radwene Laajimi, Nawfil Gueddah, and Mohamed Masmoudi. A novel design method of two-stage CMOS operational transconductance amplifier used for wireless sensor receiver. International Journal of Computer Applications, 39(??):1–11, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number11/4861-7093/.

Luo:2017:TPC

[LGZX17] Zhiyong Luo, Qi Guo, Jie Zhao, and Shensheng Xu. Tuning of a PID controller using modified dynamic group based TLBO algorithm. International Journal of Computer Applications, 157(??):17–23, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume157/number1/26795-2016912593/.

Laskar:2016:ADP

[LHKR16] Md. Tahmid Rahman Laskar, Md. Tahmid Hossain, Abu Raihan Mostofa Kamal, and Nafiul Rashid. Automated disease prediction system (ADPS): a user input-based reliable architecture for disease prediction. International Journal of Computer Applications, 133(??):24–29, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number15/23864-2016908193/.

Liraki:2015:PUN

[LHM15] Zeynab Liraki, Ali Harounabadi, and Javad Mirabedini. Predicting the users’ navigation patterns in web, using weighted association rules and users’ navigation information. International Journal of Computer Applications, 110(??):16–21, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number12/19368-1047/.

Licht:2020:MCM

[Lic20] Fabio Lopes Licht. Monte Carlo method and Brownian movement applied to future stock market analysis. International Journal of Computer Applications, 175(??):1–6, December 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume175/number34/31666-
2020920898/.

Liu:2024:AMF

[Liu24] Yuanyuan Liu. Analysis of Multi-factor Authentication (MFA) schemes in Zero Trust Architecture (ZTA): Current state, challenges, and future trends. International Journal of Computer Applications, 186(??):30–36, December 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number57/
analysis-of-multi-factor-authentication-mfa-schemes-
in-zero-trust-architecture-zta-current-state-challenges-
and-future-trends/.

Lad:2016:DPR

[LJ16] Hiteshkumar Lad and Vibhutikumar Joshi. DRC protocol for the response time reduction in CAN based distributed embedded system. International Journal of Computer Applications, 142(??):16–20, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number13/24957-2016909975/.

Lakhekar:2015:ITF

[LK15] Mahesh V. Lakhekar and Shirish L. Kotgire. Interval Type-2 fuzzy logic system for dynamic spectrum access in cognitive radio. International Journal of Computer Applications, 131(??):34–40, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume131/number8/23472-2015907439/.

Louaqad:2016:MIL

[LK16a] Saad Louaqad and Oulaid Kamach. Mixed integer linear programs for blocking and no wait job shop scheduling problems in robotic cells. International Journal of Computer Applications, 153(??):1–7, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume153/number10/26436-2016912158/.

Luthra:2016:DRP

[LK16b] Parveen Luthra and Kiranbir Kaur. A design of rectangular patch antenna with fractal slots for multiband applications. International Journal of Computer Applications, 138(??):19–23, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume138/number6/24384-2016908851/.

Latha:2016:IRU

[LKO16] P. Latha, V. Vijaya Kumar, and A. Obulesu. Image retrieval using quantized local binary pattern. International Journal of Computer Applications, 155(??): 7–15, December 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume155/number5/26599-2016912308/.

Lokare:2014:CGH

[LKS14] Deepali Lokare, A. M. Kanthe, and Dina Simunic. Cooperative gray hole attack discovery and elimination using credit based technique in MANET. International Journal of Computer Applications, 88(??):13–22, February 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume88/number15/15427-3850/.

Lin:2012:ASN

[LLSL12] Chen-Chun Lin, Chun-Ling Lin, Joseph Z. Shyu, and Chin-Teng Lin. The ANFIS system for nonlinear combined forecasts in the telecommunications industry. International Journal of Computer Applications, 37(??):30– 35, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume37/
number12/4740-6958/.

LemosFonseca:2020:MTG

[LM20] Thiago Henrique Lemos Fonseca and Alexandre Cesar Muniz De Oliveira. Metaheuristic tuning generalisation by crossvalidated racing. International Journal of Computer Applications, 177(??):1–9, March 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number47/31222-2020919991/.

Listrovoy:2012:ISH

[LMZ12] S. V. Listrovoy, S. V. Minukhin, and S. V. Znakhur. Investigation of the scheduler for heterogeneous distributed computing systems based on minimal cover method. International Journal of Computer Applications, 51(??):35– 44, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number19/8154-1948/.

Lelon:2024:FEM

[LNM24] Kiptoo Lelon, Waweru Njeri, and Joseph Muguro. Finite element mesh refinement for EM-field solvers modeling using genetic algorithm. International Journal of Computer Applications, 186(??):14–23, December 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number52/finite-element-mesh-refinement-
for-em-field-solvers-modeling-using-genetic-algorithm/

.

Luka:2023:GLD

[LOAO23] Matthew K. Luka, Okpo U. Okereke, Ejike C. Anene, and Elijah E. Omizegba. A geo-location database framework for television white space administration in Nigeria. International Journal of Computer Applications, 185(??):14–20, November 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume185/number44/32983-
2023923253/.

Lodh:2016:CAI

[LP16] Avishikta Lodh and Ranjan Parekh. Computer aided identification of flags using color features. International Journal of Computer Applications, 149(??):1–7, September 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume149/number11/26038-
2016911587/.

Lalam:2012:EQC

[LPP12] Satya Kumar Lalam, Monica Pragada, and Sudheer Palempati. EGIS 1.0: QR code screening mechanism to prevent airline luggage loss. International Journal of Computer Applications, 52(??):47–52, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume52/number1/8170-1434/.

Lakshmi:2011:MLA

[LR11] R. Deepa Lakshmi and N. Radha. Machine learning approach for taxation analysis using classification techniques. International Journal of Computer Applications, 12(??): 1–6, January 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume12/
number10/1723-2322/.

Lalitha:2012:AAT

[LR12] D. Lalitha and K. Rangarajan. An application of array token Petri nets to clustering analysis for syntactic patterns. International Journal of Computer Applications, 42 (??):21–25, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number16/5777-8033/.

Lhous:2017:FOT

[LRL17] Mustapha Lhous, Mostafa Rachik, and Abdelilah Larrache. Free optimal time control problem for a SEIR-epidemic model with immigration of infective. International Journal of Computer Applications, 159(??):1–5, February 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume159/number3/26978-2017912886/.

Lalitha:2011:TSM

[LRRR11] M. Padma Lalitha, V. C. Veera Reddy, N. Sivarami Reddy, and V. Usha Redddy. A two stage methodology for siting and sizing of DG for minimum loss in radial distribution system using RCGA. International Journal of Computer Applications, 25(??):10–16, July 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume25/number2/3005-4048/.

Lahon:2013:TOR

[LS13] M. Lahon and Y. Jayanta Singh. Task oriented risk assessment (TORA). International Journal of Computer Applications, 66(??):28–31, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume66/number7/11098-6060/.

Lengliz:2017:EVR

[LS17] Ilhem Lengliz and Afef Slama. Enhancing VANETs’ routing operation with the route lifetime policy. International Journal of Computer Applications, 164(??):35–40, April 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume164/number9/27515-2017913758/

.

Lima:2017:SBA

[LSA^{+}17] D. Lima, M. Sousa, R. Araújo, J. Hannum, A. Da Rocha, and T. Barbosa. Software with biofeedback to assist people with down syndrome. International Journal of Computer Applications, 158(??):31–37, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume158/number5/26906-2017912847/.

Laasri:2013:SSC

[LsAAA13] El Hassan Ait Laasri, Es said Akhouayri, Driss Agliz, and Abderrahman Atmani. Seismic signal classification using multi-layer perceptron neural network. International Journal of Computer Applications, 79(??):35– 43, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number15/13821-1950/.

Lawan:2012:HSP

[LSB12] S. H. Lawan, D. S. Shu’aibu, and S. A. Babale. Hyperbolicsecant pulse propagation in a single mode optical fiber system. International Journal of Computer Applications, 58(??): 23–27, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number12/9335-3645/.

Lakshminarasimman:2017:WWO

[LSB17] L. Lakshminarasimman, M. Siva, and R. Balamurugan. Water wave optimization algorithm for solving multi-area economic dispatch problem. International Journal of Computer Applications, 167(??):19–27, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number5/27768-2017914247/.

Lal:2017:HWE

[LSK17] Shyam Lal, K. K. Shukla, and Sarika Keshri. Haar wavelet expansions of signals and their applications in image processing. International Journal of Computer Applications, 173(??):1–23, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume173/number9/28360-2017915317/.

Langue:2016:MCF

[LSM16] Maureen Langue, Sukandar Sawidin, and Jusuf L. Mappadang. Monitoring the charcoal furnace control system of coconut shell waste treatment becoming charcoal with AT-Mega 328P microcontroller. International Journal of Computer Applications, 156(??):22–28, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume156/number8/26730-2016912497/.

Lestari:2017:DMA

[LSP17] Ika Winda Lestari, I Made Sukarsa, and I Ketut Adi Purnawan. Developing manufacturing application using enterprise resource planning concept. International Journal of Computer Applications, 164(??):19–24, April 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume164/number8/27503-2017913709/.

Laloo:2020:TIH

[LSS20] Nassirah Laloo, Mohammad Sameer Sunhaloo, and Raja K. Subramanian. Timeline improvement of hierarchical clustering algorithms for wireless sensor networks. International Journal of Computer Applications, 176(??):1–7, June 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume176/number30/31389-
2020920329/.

Lima:2020:CMB

[LSSD20] Marilia N. C. A. Lima, Wedson L. Soares, Iago R. R. Silva, and Roberta A. De A.Fagundes. A combined model based on clustering and regression to predicting school dropout in higher education institution. International Journal of Computer Applications, 176(??):1–8, June 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number34/31421-2020920396/.

M:2012:DMC

[LT12] Chayadevi M L and Raju G T. Data mining, classification and clustering with morphological features of microbes. International Journal of Computer Applications, 52 (??):1–5, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number4/8187-1547/.

Lather:2013:HTC

[LT13] J. S. Lather and V K Tayal. Hh TCSC controller and fuzzy PSS design in mitigating small signal oscillations. International Journal of Computer Applications, 68(??): 6–11, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume68/
number14/11645-7136/.

Latifi:2016:MDB

[LW16] Shahram Latifi and Steven Wilson. Maximum distance band selection of hyperspectral images. International Journal of Computer Applications, 133(??):36–43, January 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume133/number17/24009-
2016908214/.

Lumentut:2023:IBC

[LWL^{+}23] Ventje Lumentut, Stephy Beatrix Walukow, Fitria Claudya Lahinta, Fanny Doringin, and Maksy Sendiang. IoT based customer KWH meter design. International Journal of Computer Applications, 184(??):6–10, April 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number53/32665-2023922625/.

Liu:2015:LTW

[LZW15] Qiegen Liu, Li Zhu, and Jianhua Wu. Log-transform weighted total variation for image smoothing. International Journal of Computer Applications, 124(??):35–40, August 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume124/number2/22080-2015905411/

.

Meenakshi:2010:NAF

[MA10] A. Meenakshi and K. Alagarsamy. A novelty approach for finding frequent itemsets in horizontal and vertical layout- HVCFPMINETREE. International Journal of Computer Applications, 10(??):20–27, November 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume10/number5/1478-1995/.

Masoumi:2012:HCD

[MA12] Majid Masoumi and Shervin Amiri. A high capacity digital watermarking scheme for copyright protection of video data based on YCbCr color channels invariant to geometric and non-geometric attacks. International Journal of Computer Applications, 51(??):13–20, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume51/number13/8101-1693/.

Mohammed:2013:FIB

[MA13a] Samir Jasam Mohammed and Hayder Fadhil Abdulsada. FPGA implementation of 3 bits BCH error correcting codes. International Journal of Computer Applications, 71 (??):35–42, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number7/12373-8698/.

Muhammad:2013:JEP

[MA13b] Ario Muhammad and Ratih Nur Esti Anggraini. Java Earthquake Program (JEP) to predict the recurrence time of significant earthquake in Indonesia. International Journal of Computer Applications, 64(??):6–12, February 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume64/number9/10660-5428/.

Manojmanimaran:2015:DID

[MA15] S. Manojmanimaran and C. Amutha. Design and implementation of dsPIC fed Vienna rectifier for induction heating appliances. International Journal of Computer Applications, 114(??):20–25, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume114/number15/20054-2012/.

Maraud:2016:OBC

[MA16] Hussain S. Maraud and Isam M. Abdulbaqi. Optimal battery charger fed by photovoltaic system based on decreased charging current method. International Journal of Computer Applications, 139(??):12–21, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number12/24541-2016909419/.

Mosa:2017:PED

[MA17] Buthaina Mosa and Aya Falah Algamluoli. Performance of energy detector for cognitive radio system over AWGN and Rayleigh channel. International Journal of Computer Applications, 167(??):30–34, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number3/27753-2017914220/.

B:2022:ESU

[MA22] Parashiva Murthy B M and Sumithra Devi K A. An efficient security for unstructured big data using a reconfigurable security suite. International Journal of Computer Applications, 184(??):42–46, June 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number18/32419-2022922198/.

Mahmoud:2013:RUW

[MAA13] M. A. W. Mahmoud, M. A. Attallah, and A. Alharby. On reliability of (n + 1)-unit warm standby system based on imperfect repair facility and two types of failures. International Journal of Computer Applications, 70 (??):1–6, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number1/11923-7704/.

Mahmoud:2014:CIE

[MAaM14] M. A. W. Mahmoud, Y. Abdel-aty, and N. M. Mohamed. Characterization of the inverse exponential-type distribution based on recurrence relations for dual generalized order statistics. International Journal of Computer Applications, 91(??):1–7, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume91/number14/15946-4911/.

Mobin:2016:IFD

[MAARIH16] Md Iftekharul Mobin, Md Abid-Ar-Rafi, Md Neamul Islam, and Md Rifat Hasan. An intelligent fire detection and mitigation system safe from fire (SFF). International Journal of Computer Applications, 133(??):1–7, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number6/23787-2016907858/.

Mouatassim:2016:MBH

[MABE16] Salma Mouatassim, Mustapha Ahlaqqach, Jamal Benhra, and My Ali El Oualidi. Model based on hybridized game theory to optimize logistics: Case of blood supply chain. International Journal of Computer Applications, 145(??):37–48, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume145/number15/25358-2016910910/.

Mahmoud:2016:HSN

[MAE16a] Tarek M. Mahmoud, Abdelmgeid A. Ali, and Hussein M. Elshafie. A hybrid snort-negative selection network intrusion detection technique. International Journal of Computer Applications, 146(??):24–31, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number5/25395-2016910703/.

Mikhail:2016:THD

[MAE16b] Mervat Mikhail, Yasmine Abouelseoud, and Galal Elkobrosy. Text hiding in a digital cover image using two dimensional indexing based on chaotic maps. International Journal of Computer Applications, 138(??):1–6, March 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume138/number12/24428-2016909037/.

Mohamed:2019:DVD

[MAE19] M. A. Mohamed, M. I. Fath Allah, and L. W. Elbon. DCT versus DWT chaotic based color image encryption. International Journal of Computer Applications, 177(??):24–41, December 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number24/31046-2019919706/.

Mahgoub:2013:INS

[Mah13] Hany Mahgoub. IARMMD: a novel system for incremental association rules mining from medical documents. International Journal of Computer Applications, 64(??):28– 35, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number1/10599-5299/.

Mohammad:2015:NLC

[MAH15] Amina Mohammad, Bassant Abdelhamid, and I. M. Hafez. A new linearly combined Nyquist pulses for PAPR reduction in IFDMA. International Journal of Computer Applications, 126(??):38–43, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume126/number6/22559-2015906073/.

Mahasneh:2016:DSC

[Mah16] Hassan Mahasneh. In designing of ST-coded in-home PLC transmission scheme. International Journal of Computer Applications, 141(??):1–5, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume141/number2/24753-2016909554/.

Mahmood:2019:IMM

[Mah19] Radhwan Basher Mahmood. Impact of mobility models on performance of ad hoc position based routing protocol LAR. International Journal of Computer Applications, 181(??):27–31, February 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number40/30328-2019918427/.

Maidamisa:2012:CAW

[MAI12] Ahmad A. Maidamisa, Rohanin Ahmad, and Abdul A. M. Ismail. A comparative analysis of window width selection technique in data envelopment analysis. International Journal of Computer Applications, 41(??):21–26, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume41/number5/5537-7589/.

Mohammad:2015:CAD

[MAI15] Adnan Mohammad, Abdullah Al Manzur, and Ashique Ibni Aziz. Comparisons among different types of single phase step down AC to DC converters. International Journal of Computer Applications, 121(??):13–17, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume121/number5/21535-4540/.

Mohamed:2012:DHL

[MAk12] Marghny H. Mohamed and Hussein I. Abul-kasim. Data hiding by LSB substitution using gene expression programming. International Journal of Computer Applications, 45 (??):13–20, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume45/
number14/6848-9408/.

Makhloufi:2013:MPP

[MAK13] Mohamed Tahar Makhloufi, Yassine Abdessemed, and Mohamed Salah Khireddine. Maximum power point tracker for photovoltaic systems using on-line learning neural networks. International Journal of Computer Applications, 72 (??):29–36, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number10/12530-9000/.

Malhotra:2010:ERP

[Mal10] Jyoteesh Malhotra. Error rate performance in high data rate UWB channels for wireless personal area networks. International Journal of Computer Applications, 10(??):31– 37, November 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume10/
number6/1484-2001/.

Malhotra:2013:MCB

[Mal13a] Deepti Malhotra. Minimum CDT based scheduling algorithm versus FCFS in grid environment. International Journal of Computer Applications, 68(??):18–24, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume68/number11/11622-6505/.

Mall:2013:EFR

[Mal13b] Sober Mall. Estimating a function of real demand for money in Pakistan: an application of bounds testing approach to cointegration. International Journal of Computer Applications, 79(??):32–50, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume79/number5/13740-1548/.

Mall:2013:FDI

[Mal13c] Sober Mall. Foreign direct investment inflows in Pakistan: a time series analysis with autoregressive distributive lag (ARDL) approach. International Journal of Computer Applications, 78(??):7–16, September 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume78/number5/13483-1186/.

Mostafa:2012:SCM

[MAM12] Salama A. Mostafa, Mohd Sharifuddin Ahmad, and M. A. Firdaus M. A soft computing modeling to casebased reasoning implementation. International Journal of Computer Applications, 47(??):14–21, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.
org/archives/volume47/number7/7199-9976/.

Mahmoud:2015:MAR

[MAM15] Tarek M. Mahmoud, Abdelmgeid A. Aly, and Omar Makram M. A modified AODV routing protocol to avoid black hole attack in MANETs. International Journal of Computer Applications, 109(??):27–33, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume109/number6/19195-0809/.

Mohammed:2016:MQC

[MAM16] Maysoon A. Mohammed, Mazlina AbdulMajid, and Balsam A. Mustafa. Multi-queue CPU process prioritization using a dynamic quantum time algorithm compared with varying time quantum and round-robin algorithms. International Journal of Computer Applications, 134(??):31–37, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number7/23929-2016907983/.

Mom:2019:DTA

[MAM19] Joseph M. Mom, George K. Azever, and Moses S. Maor. Data throughput assessment of 4T4R MIMO technique on LTE wireless cellular networks. International Journal of Computer Applications, 182(??):18–24, April 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number50/30538-2019918748/.

Munandar:2016:MAC

[MAMA16] Tb. Ai Munandar, Azhari Azhari, Aina Musdholifah, and Lincolin Arsyad. Modified agglomerative clustering by klassen disparities for identification hierarchical cluster of regional developing countries. International Journal of Computer Applications, 141(??):1–5, May 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume141/number9/24809-2016909730/.

Mohammed:2012:NAM

[MAMS12] Khudhair Abbas Mohammed, Mohd Sharifuddin Ahmad, Salama A. Mostafa, and M. A. Firdaus M. Sharifuddin. A nodal approach to modeling human-agents collaboration. International Journal of Computer Applications, 43 (??):33–40, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume43/
number12/6158-8555/.

Mangat:2010:SIB

[Man10] Veenu Mangat. Swarm intelligence based technique for rule mining in the medical domain. International Journal of Computer Applications, 4(??):19–24, July 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume4/number1/796-1131/.

Mohapatra:2013:FFC

[MAP13] Puspanjali Mohapatra, Munnangi Anirudh, and Tapas Kumar Patra. Forex forecasting: a comparative study of LL-WNN and NeuroFuzzy hybrid model. International Journal of Computer Applications, 66(??):46–53, March 2013. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume66/number18/11188-6451/

.

Martin:2012:EOO

[Mar12] V. Maria Antoniate Martin. Evaluation of object oriented programming skills of students with respect to trait emotional intelligence based on students performance. International Journal of Computer Applications, 48(??): 34–40, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number10/7387-0187/.

Muaidi:2012:TAS

[MAt12] Hasan Muaidi and Rasha Al-tarawneh. Towards Arabic spell-checker based on N-grams scores. International Journal of Computer Applications, 53(??):12–16, September 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume53/number3/8400-
2168/.

Mazen:2021:FOT

[Maz21] Tamer Sh. Mazen. A framework for optimization of technical indicators parameters for forex (foreign exchange) based on genetic algorithm. International Journal of Computer Applications, 183(??):6–10, September 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number25/32081-2021921623/.

Mangala:2011:NAR

[MB11a] T. Rajani Mangala and S. G. Bhirud. A new automatic road extraction technique using gradient operation and skeletal ray formation. International Journal of Computer Applications, 29(??):17–25, September 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume29/number1/3529-4813/.

Medouakh:2011:SSJ

[MB11b] S. Medouakh and Z-E. Baarir. Study of the standard JPEG2000 in image compression. International Journal of Computer Applications, 18(??):27–33, March 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume18/number1/2247-2877/.

Moussa:2012:CSG

[MB12] M. I. Moussa and E. M. Badr. A computational study for the graph-theoretic version of the union-closed sets conjecture. International Journal of Computer Applications, 50(??):1–5, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume50/
number12/7820-0945/.

Mandal:2013:WCA

[MB13a] Lopa Mandal and Pramatha Nath Basu. Web content authoring for ubiquitous and context aware learning — a tag based approach. International Journal of Computer Applications, 75(??):23–28, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume75/number7/13124-0480/.

Mittal:2013:PQF

[MB13b] Sonu Mittal and Pradeep Kumar Bhatia. Predicting quantitative functional dependency metric based upon the interface complexity metric in component based software systems: a new approach. International Journal of Computer Applications, 73(??):21–28, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume73/number2/12713-9527/.

Moustafa:2013:SOP

[MB13c] Bouzaki Mohammed Moustafa and Benyoucef Boumediene. Simulation and optimization of the performance in hit solar cell. International Journal of Computer Applications, 80 (??):43–47, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume80/
number13/13925-1930/.

Malik:2015:LMV

[MB15a] Vikas Malik and C. R. Barde. Live migration of virtual machines in cloud environment using prediction of CPU usage. International Journal of Computer Applications, 117(??):1–5, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number23/20691-3604/.

Mouassa:2015:ABC

[MB15b] Souhil Mouassa and Tarek Bouktir. Artificial bee colony algorithm for solving OPF problem considering the valve point effect. International Journal of Computer Applications, 112(??):45–53, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number1/19634-1208/.

Muhammad:2015:APA

[MB15c] A. L. Muhammad and A. B. Baffa. Airflow process across vertical vents induced by stack- driven effect with an opposing flow in one of the upper openings. International Journal of Computer Applications, 123(??):1–8, August 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume123/number1/21920-2015900728/.

Mukherjee:2016:CAA

[MB16] Jayabrata Mukherjee and Deepak Bhattacharya. Computation (abacus) aspects of the sahasralingam. International Journal of Computer Applications, 143(??):13–20, June 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume143/number13/25136-
2016910405/.

Mustapha:2017:RSA

[MB17] Khalid Shahabdeen Mustapha and Edem Kwedzo Bankas. RNS scaling algorithm for a new moduli set {2^{(2n+1)} + 1, 2^{(2n+1)}, 2^{(2n+1)} − 1}. International Journal of Computer Applications, 165(??):21–28, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number10/27609-2017913974/.

Macedo:2020:ECV

[MBA20] Robson A. G. Macedo, Peterson A. Belan, and Sidnei A. Araújo. An embedded computer vision system for beans quality inspection. International Journal of Computer Applications, 175(??):44–53, October 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number24/31603-2020920779/.

Mahmood:2012:DKM

[MBAR12] Nadeem Mahmood, Aqil Burney, Zain Abbas, and Kashif Rizwan. Data and knowledge management in designing healthcare information systems. International Journal of Computer Applications, 50(??):34–39, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume50/number2/7745-0798/.

MoreiraESilvaBernardes:2017:ODC

[MBCL17] Maurício Moreira E Silva Bernardes, Graziela Brunhari Kauling, Milena Chang Chain, and Lars Löfqvist. Original design in a copying-intensive industry. International Journal of Computer Applications, 159(??): 29–38, February 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume159/number7/27016-2017912998/.

Mnasri:2010:DDP

[MBE10] Zied Mnasri, Fatouma Boukadida, and Noureddine Ellouze. Design and development of a prosody generator for Arabic TTS systems. International Journal of Computer Applications, 12(??):24–31, December 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume12/number1/1641-2206/.

Mohammed:2016:TTC

[MBE16] Reham H. Mohammed, Fahmy Bendary, and Kamel Elserafi. Trajectory tracking control for robot manipulator using fractional order-fuzzy-PID controller. International Journal of Computer Applications, 134(??):22–29, January 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume134/number15/23992-
2016908155/.

Mahapatra:2015:MCS

[MBKC15] Abhijit Mahapatra, Kaustav Biswas, Amit Kumar, and Avik Chatterjee. Motion control strategies based on PD control for a four degree-of-freedom serial robotic manipulator to mimic human index finger articulations. International Journal of Computer Applications, 109(??):35–42, January 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume109/number13/19252-
0949/.

Mediouni:2016:MPE

[MBKH16] Nejib Mediouni, Samir Ben Abid, Oussama Kallel, and Salem Hasnaoui. Modeling and performance evaluation of 2D and 3D NoCs using discrete event simulation. International Journal of Computer Applications, 137(??):1–7, March 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume137/number12/24324-
2016908852/.

Mohandas:2015:OLS

[MBL15a] N. Mohandas, R. Balamurugan, and L. Lakshminarasimman. Optimal location and sizing of multiple DGs to enhance the voltage stability in the distribution system using a chaotic ABC algorithm. International Journal of Computer Applications, 128(??):10–26, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number2/22844-2015906432/.

Muhammad-bello:2015:DIW

[MBL15b] Muhammad-Bello and Bilkisu Larai. Design and implementation of a website for road transport operators and travellers in Nigeria. International Journal of Computer Applications, 110(??):37–44, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number2/19291-0720/.

Menaouer:2013:KME

[MBM13] Brahami Menaouer, Atmani Baghdad, and Nada Matta. Knowledge mapping evolution guided by data mining. International Journal of Computer Applications, 71(??): 32–39, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number1/12325-8554/.

Maatoug:2014:MSE

[MBM14] Abdelfettah Maatoug, Ghalem Belalem, and Kadda Mostefaoui. Modeling and simulation of energy management system for smart city with the formalism DEVS: Towards reducing the energy consumption. International Journal of Computer Applications, 90(??):38–43, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume90/number18/15824-4710/.

Mozammel-Bin-Motalab:2011:CCB

[MBMA11] Mozammel-Bin-Motalab and Shoyeb Al Mamun Shohag. Cloud computing and the business consequences of ERP use. International Journal of Computer Applications, 28 (??):31–37, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume28/
number8/3406-4751/.

Mofarreh-Bonab:2012:AVC

[MBMB12] Mostafa Mofarreh-Bonab and Mohamad Mofarreh-Bonab. Adaptive video compression using PCA method. International Journal of Computer Applications, 44(??):6– 11, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number21/6393-8866/.

Mohan:2016:DRD

[MBP16] K. Jagan Mohan, M. Balasubramanian, and S. Palanivel. Detection and recognition of diseases from paddy plant leaf images. International Journal of Computer Applications, 144(??):34–41, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number12/25234-2016910505/.

Muhammad:2016:ISD

[MBR16] A. L. Muhammad, A. B. Baffa, and M. Z. Ringim. Investigations of stack- driven airflow through rectangular cross- ventilated building with two openings using analytic technique. International Journal of Computer Applications, 141(??):5–11, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume141/number6/24786-2016909631/.

Mohammad:2016:RTS

[MBZA16] Mohammad Tarik Mohammad, Sara Basim, Allaa Zaki, and Maher Algreer. Real time self-tuning controller for position control of DC motor system using pole-placement technique. International Journal of Computer Applications, 155(??):16–21, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume155/number3/26585-2016912279/.

Mann:2012:HWC

[MC12] Deep Mann and Inderveer Chana. Heterogeneous workload consolidation for efficient management of data centers in cloud computing. International Journal of Computer Applications, 50(??):13–17, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number10/7806-0939/.

Manco:2013:EFI

[MC13] Angelo Manco and Vittorio Ugo Castrillo. Efficient FPGA implementation of the basic receiving functions for aeronautical reconfigurable data-link. International Journal of Computer Applications, 71(??):16–22, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume71/number20/12602-9390/.

Mateen:2016:RWD

[MC16] Ahmed Mateen and Lareab Chaudhary. Reduce the wastage of data during movement in data warehouse. International Journal of Computer Applications, 152(??): 20–24, October 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume152/number8/26340-2016911907/.

Majid:2017:NPB

[MC17a] Abdul Majid and Amit Chaturvadi. Non path-based mutual anonymity approach with secure cipher text to image encryption message-peer-to-peer systems. International Journal of Computer Applications, 161(??):14–18, March 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume161/number13/27330-2017913433/.

Malik:2017:RME

[MC17b] S. C. Malik and S. K. Chauhan. Reliability and MTSF evaluation of a parallel-series system using Weibull failure laws. International Journal of Computer Applications, 172(??):11–21, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume172/number2/28222-2017915069/.

Mousa:2014:FBD

[MCA14] Amthal K. Mousa, Muayad S. Croock, and Mohammed N. Abdullah. Fuzzy based decision support model for irrigation system management. International Journal of Computer Applications, 104(??):14–20, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number9/18230-9177/.

Mouhadjer:2014:ETO

[MCaC14] Samir Mouhadjer, Ali Chermitti, and Sid ali Chikhi. Energy transfer optimization in photovoltaic conversion by the maximum power point tracking. International Journal of Computer Applications, 106(??):16–20, November 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume106/number10/18556-7902/.

Mondal:2012:QDS

[MCB12] Bikromadittya Mondal, Sarit Chakraborty, and Sharmistha Bhattacharjee. Quasi-dynamic scheduling in wireless broadcast networks. International Journal of Computer Applications, 43(??):13–17, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number9/6130-8360/.

Malhotra:2013:EAN

[MD13] Shilpi Malhotra and Ashutosh Dixit. An effective approach for news article summarization. International Journal of Computer Applications, 76(??):5–10, August 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume76/number16/13329-0797/.

Minnie:2014:NFT

[MD14a] J. Befija Minnie and A. Victor Devados. A new fuzzy tool multi- impacts resolving fuzzy cognitive maps (MIR-FCM) to analyze the factors causing work life imbalance. International Journal of Computer Applications, 105(??):31–37, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number13/18440-9788/.

Mohamed:2014:AEF

[MD14b] Muhidin A. Mohamed and Mohamed A. Deriche. An approach for ECG feature extraction using Daubechies 4 (DB4) wavelet. International Journal of Computer Applications, 96(??):36–41, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number12/16850-6712/.

Michel:2022:DNI

[MDCS22] Deussom Djomadji Eric Michel, Magoudaya David, Feudjio Cyrille, and Michael Ekonde Sone. Design of an NB-IoT smart metering solution: Coverage and capacity planning: Case of yaoundé and douala. International Journal of Computer Applications, 184(??):20–30, March 2022. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number2/32305-2022921972/.

Masnik:2021:DMI

[MDD21] Paulo Vinicius Masnik, Roberto Alexandre Dias, and Mario De Noronha Neto. Detection of myocardial infarction in electrocardiograms using machine learning. International Journal of Computer Applications, 183(??):12–19, June 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number9/31954-2021921384/

.

Mahersi:2013:DWB

[MDGS13] Sami Mahersi, Mohamed Dhieb, Hamadi Ghariani, and Mounir Samet. Design of a wide band RF amplifier using scattering parameters. International Journal of Computer Applications, 66(??):1–4, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume66/number11/11125-6196/.

Moreira:2017:CSF

[MDM17] R. B. Moreira, A. F. Dos Santos, and R. H. C. Maniçoba. A compact and stable frequency selective surface for WLAN applications. International Journal of Computer Applications, 166(??):1–3, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume166/number7/27678-2017914056/.

Mousa:2012:OOP

[MDNA12] Allam Mousa, Yousef Dama, Mahmoud Najjar, and Bashar Alsayeh. Optimizing outdoor propagation model based on measurements for multiple RF cell. International Journal of Computer Applications, 60(??):5–10, December 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume60/number5/9686-
4121/.

Mehra:2012:RDG

[MDP12] Rajesh Mehra, Swapna Devi, and S. S. Pattnaik. Reconfigurable design of GSM digital down converter for enhanced resource utilization. International Journal of Computer Applications, 57(??):41–47, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume57/number11/9161-3318/.

Mazroui:2014:RMI

[ME14] A. Mazroui and A. Kerkour Elmiad. Recognition of multifont isolated Arabic characters by bézier curves. International Journal of Computer Applications, 100(??):1–7, August 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume100/number6/17526-8101/

.

Michel:2015:NPM

[ME15] Deussom Djomadji Eric Michel and Tonye Emmanuel. New propagation model optimization approach based on particles swarm optimization algorithm. International Journal of Computer Applications, 118(??):39–47, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume118/number10/20785-3430/.

Marghny:2011:EEC

[MEAT11] M. H. Marghny, Rasha M. Abd El-Aziz, and Ahmed I. Taloba. An effective evolutionary clustering algorithm: Hepatitis C case study. International Journal of Computer Applications, 34(??):1–6, November 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume34/number6/4092-5420/.

Meena:2012:EDF

[Mee12] Shyam Sundar Meena. Efficient discovery of frequent patterns using KFP-tree from Web logs. International Journal of Computer Applications, 49(??):15–18, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume49/number10/7662-0770/.

Mobarakeh:2015:FNN

[MEE15] Majid Iranpour Mobarakeh, Mehran Emadi, and Majid Emadi. FRBF neural network base for face recognition using Zernike moments and PCA. International Journal of Computer Applications, 125(??):10–14, September 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume125/number2/22403-2015905834/.

Mostafa:2016:PFS

[MEE16] Mostafa H. Mostafa, Mostafa A. Elshahed, and Magdy M. Elmarsfawy. Power flow study and voltage stability analysis for radial system with distributed generation. International Journal of Computer Applications, 137(??): 19–26, March 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume137/number9/24304-2016908919/.

Meghanathan:2010:EAG

[Meg10] Natarajan Meghanathan. An energy-aware greedy perimeter stateless routing protocol for mobile ad hoc networks. International Journal of Computer Applications, 9(??):30– 35, November 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume9/
number6/1388-1871/.

Mustafa:2014:IST

[MEH14] Lailah M. Mustafa, Mohamed K. Elmahy, and Mohamed H. Haggag. Improve scheduling task based task grouping in cloud computing system. International Journal of Computer Applications, 93(??):1–5, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume93/number8/16232-5561/.

Mousa:2015:ECB

[MEKS15] Farid Ali Mousa, Reda A. El-Khoribi, and Mahmoud E. Shoman. EEG classification based on machine learning techniques. International Journal of Computer Applications, 128(??):22–27, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number4/22861-2015906515/.

Meshref:2016:NIA

[Mes16] Hossam Meshref. Novel implementation of the artificial immune system memory. International Journal of Computer Applications, 155(??):36–41, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume155/number13/26669-2016912585/.

Meshref:2023:IIP

[Mes23] Hossam Meshref. Investigating the impact of prominent factors on the diagnosis of diabetes and its associated diseases using ensemble machine learning models. International Journal of Computer Applications, 184(??):19–30, February 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume184/number48/32630-
2023922598/.

Mohamed:2017:NIS

[MESNH17] Sayed A. Mohamed, A. S. El-Sherbeny, Ayman H. Nasr, and A. K. Helmy. A new image super-resolution restoration algorithm. International Journal of Computer Applications, 173(??):5–12, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume173/number10/28443-2017915355/.

Malekzadeh:2013:ESN

[MFA13] Mina Malekzadeh, Zohre Fereidooni, and M. H. Shahrokh Abadi. Effects of sensor nodes mobility on routing energy consumption level and performance of wireless sensor networks. International Journal of Computer Applications, 79(??):11–19, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume79/number10/13776-1645/.

Mohamed:2013:AWI

[MFE13] Sarah Fouad Mohamed, Yasmine Fahmy, and Abd El Halim Zekry. Avoiding WIMAX interference on ultra wide band MB-OFDM system by cognitive radio. International Journal of Computer Applications, 64(??):11– 15, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number6/10637-5379/.

Mohd:2012:BMM

[MFM12] Izzeldin Ibrahim Mohd, Chay Chin Fatt, and Muhammad N. Marsono. 32-bit N × N matrix multiplication: Performance evaluation for Altera FPGA, i5 Clarkdale and Atom Pineview-D Intel general purpose processors. International Journal of Computer Applications, 52(??):17– 23, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number11/8246-1757/.

Malik:2012:CBA

[MG12] S. C. Malik and Gitanjali. Cost-benefit analysis of a parallel system with arrival time of the server and maximum repair time. International Journal of Computer Applications, 46(??):39–44, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number5/6907-9303/.

Mandala:2014:OPT

[MG14a] Manasarani Mandala and C. P. Gupta. Optimal placement of TCSC for congestion management using modified particle swarm optimization. International Journal of Computer Applications, 88(??):34–40, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume88/number11/15400-4073/.

Mane:2014:HPS

[MG14b] Sandeep U. Mane and Pankaj G. Gaikwad. Hybrid particle swarm optimization (HPSO) for data clustering. International Journal of Computer Applications, 97 (??):1–5, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number19/17112-7514/.

Muley:2014:TSI

[MG14c] Tushar C. Muley and Sanjay S. Gharde. Theoretical solution for identifying and creating digital camouflage images by two scale decomposition. International Journal of Computer Applications, 106(??):8–13, November 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume106/number5/18514-9599/.

Mali:2017:MAT

[MG17a] G. U. Mali and D. K. Gautam. A model application of two phase commit protocol in wireless DTN. International Journal of Computer Applications, 162(??): 23–28, March 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume162/number5/27240-2017913297/.

Mostafa:2017:FSA

[MG17b] Manal Mostafa and Mohamed Gamal. Feature selection algorithm for enhancing modeling efficiency. International Journal of Computer Applications, 173(??): 1–7, September 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume173/number4/28320-2017915279/.

Mustafa:2019:NAH

[MG19] Hassan M. H. Mustafa and Satish Gajawada. Novel artificial human optimization field algorithms — the beginning. International Journal of Computer Applications, 182(??):19–34, February 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number42/30373-2019918488/.

MarquesFilgueirasTeixeira:2020:LPA

[MGA20] Silvana Marques Filgueiras Teixeira, Warley Gramacho Da Silva, and Glêndara Aparecida De Souza Martins. Linear programming applied the quantitative analysis of the nutrients of the menus served in full-time daycare centers. International Journal of Computer Applications, 175(??):22–26, December 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number35/31677-2020920906/.

Mukane:2011:SIT

[MGB11] Shailendrakumar M. Mukane, Sachin R. Gengaje, and Dattatraya S. Bormane. On scale invariance texture image retrieval using fuzzy logic and wavelet co-occurrence based features. International Journal of Computer Applications, 18 (??):10–17, March 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume18/
number3/2266-2917/.

Montellano:2022:RHF

[MGL22] C. David Betancourt Montellano, C. Onder Francisco Campos Garcia, and Roberto Oswaldo Cruz Leija. Recognition of handwritten flowcharts using convolutional neural networks. International Journal of Computer Applications, 184(??):37–41, March 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number1/32301-2022921969/.

Mohammed:2011:EHT

[MH11] Aree Ali Mohammed and Jamal Ali Hussein. Efficient hybrid transform scheme for medical image compression.

International Journal of Computer Applications, 27(??): 16–20, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume27/
number7/3313-4548/.

Mahmud:2012:ERB

[MH12a] Tanjim Mahmud and Mohammad Shahadat Hossain. An evidential reasoning-based decision support system to support house hunting. International Journal of Computer Applications, 57(??):51–58, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume57/number21/9242-3884/.

Muhammad:2012:AFE

[MH12b] MIEEE Muhammad and Md. Ashraful Hoque. On analysis of front end current of rectifier converter for low THD and high PF with SEPIC. International Journal of Computer Applications, 37(??):35–40, January 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume37/number4/4600-6558/.

Malik:2013:UEM

[MH13a] Asad Abdul Malik and Asad Habib. Urdu to English machine translation using bilingual evaluation understudy. International Journal of Computer Applications, 82(??):5– 12, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume82/
number7/14126-1040/.

Musdholifah:2013:MST

[MH13b] Aina Musdholifah and Siti Zaiton Mohd Hadhim. Mining spatio-temporal data of fatal accident. International Journal of Computer Applications, 63(??):40–46,

February 2013.  CODEN ????  ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number8/10490-5243/.

Mohamed:2014:FNN

[MH14a] Mohamed and A. H. A fuzzy neural network fault diagnostic system. International Journal of Computer Applications, 94(??):9–13, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume94/number1/16305-5531/.

Humadi:2014:OPS

[mH14b] Ruba Al mula Humadi. Optimal placement of series and compound compensators for power loss minimization based on WPSO algorithm. International Journal of Computer Applications, 97(??):1–7, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume97/number12/17056-7361/.

Mokhlis:2013:RMS

[MHE13] N. A. Mokhlis, N. A. Hassan, and E. M. El Sayed. On the reliability of multi-state m-consecutive-at least-k-out-of-n: F systems. International Journal of Computer Applications, 69(??):40–44, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume69/number7/11858-7629/.

Mahmoud:2015:CAT

[MHE15] M. A. W. Mahmoud, M. Y. Haggag, and A. E. B. Abd Elghany. Cost analysis of a two-unit cold standby system considering hardware, software failures and inspection with maximum repair time. International Journal of Computer Applications, 129(??):1–8, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume129/number5/23066-2015906910/.

Mishra:2015:EFM

[MHJ15] Arunodaya Raj Mishra, Dhara Singh Hooda, and Divya Jain. On exponential fuzzy measures of information and discrimination. International Journal of Computer Applications, 119(??):1–7, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number23/21373-3712/.

Murthy:2014:HII

[MHV^{+}14] Y. V. Srinivasa Murthy, Kalaga Harish, D. K. Vishal Varma, Karri Sriram, and B. V. S. S. Revanth. Hybrid intelligent intrusion detection system using Bayesian and genetic algorithm (BAGA): Comparitive study. International Journal of Computer Applications, 99(??):1–8, August 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume99/number2/17342-7808/.

Muangon:2013:CBR

[MI13] Weenawadee Muangon and Sarun Intakosum. Casebased reasoning for design patterns searching system. International Journal of Computer Applications, 70(??): 16–24, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number26/12231-8433/.

Mohamed:2015:TFE

[MIAEFS15] M. A. Mohamed, Abdelhameed Ibrahim, M. G. Abd-El-Fattah, and A. S. Samrah. Two fragile encryptionwatermarking combined schemes for image authentication with tamper detection and localization. International Journal of Computer Applications, 126(??):23– 33, September 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume126/number3/22533-2015906014/.

Mohamed:2012:OFA

[MIS12] Ashara Banu Mohamed, Norbik Bashah Idris, and Bharanidharan Shanmugum. An operational framework for alert correlation using a novel clustering approach. International Journal of Computer Applications, 54(??):23– 28, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume54/
number12/8618-2480/.

Mishra:2014:DIS

[Mis14] Prasun Mishra. Design and implementation of speed controller with anti-windup scheme for three phase induction motor used in electric vehicle. International Journal of Computer Applications, 89(??):19–25, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume89/number9/15531-4371/.

Mishra:2012:COM

[MJ12] Nidhi Mishra and C. K. Jha. Classification of opinion mining techniques. International Journal of Computer Applications, 56(??):1–6, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume56/number13/8948-3122/.

Montaseri:2014:FSA

[MJ14] S. Abbas Montaseri and Reza Javidan. A feasibility study on applying remotely sensed data for agricultural management in beyza land. International Journal of Computer Applications, 99(??):1–5, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume99/number8/17390-7932/.

Malviya:2015:NTC

[MJ15] Rajesh Malviya and Pranita Jain. A novel text categorization approach based on K-means and support vector machine. International Journal of Computer Applications, 130(??):1–7, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number14/23274-2015907164/.

Mohammed:2016:PAW

[MJM16] Emad A. Mohammed, A. I. A. Jabbar, and Khalid K. Mohammed. Performance analysis of WRAN at physical layer of IEEE 802.22 standardization. International Journal of Computer Applications, 133(??):4–8, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number14/23852-2016908086/.

Mohammad:2017:SCM

[MJM17] Rahmanidoust Mohammad, Zheng Jianguo, and Rabiee Meysam. Simultaneous considering of machine availability constraint, sequence dependent setup time and ready time in a no-wait hybrid flow shop scheduling problem to minimize mean tardiness. International Journal of Computer Applications, 169(??):30–37, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume169/number7/27997-2017914789/.

Manawadu:2015:ETC

[MJP15] C. D. Manawadu, Md Gapar Md Johar, and S. S. N. Perera. Essential technical competencies for software engineers: Perspectives from Sri Lankan undergraduates. International Journal of Computer Applications, 113(??):27–34, March 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume113/number17/19920-

Mujawar:2018:WBF

[MJP18] I. K. Mujawar, B. T. Jadhav, and Kapil Patil. Web-based fuzzy expert system for symptomatic risk assessment of diabetes mellitus. International Journal of Computer Applications, 182(??):5–12, July 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number3/29740-2018917482/.

Mohapatra:2015:ILW

[MJS15] Nibedita Priyadarshini Mohapatra, Jagruti Jena, and Sujata Kumari Sahu. Improving the life of the wireless sensor network using energy harvesting clustering. International Journal of Computer Applications, 114(??):37–43, March 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume114/number10/20018-
2028/.

Mahmood:2012:SON

[MK12] F. Mahmood and F. Kunwar. A self-organizing neural scheme for door detection in different environments. International Journal of Computer Applications, 60(??):13– 18, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number9/9719-3679/.

Maharana:2013:IAP

[MK13a] Rakesh Maharana and Pabitra Mohan Khilar. An improved authentication protocol for hierarchical wireless sensor networks using ECC. International Journal of Computer Applications, 67(??):23–30, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume67/number22/11528-7352/.

Maheswari:2013:ALS

[MK13b] S. Maheswari and G. R. Karpagam. Applying logical scoring preference method for Semantic Web service selection. International Journal of Computer Applications, 65 (??):38–46, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number19/11035-6259/.

Majidnezhad:2013:ABM

[MK13c] Vahid Majidnezhad and Igor Kheidorov. An ANNbased method for detecting vocal fold pathology. International Journal of Computer Applications, 62(??):1– 4, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number7/10089-4722/.

Mandal:2013:EVM

[MK13d] Sameer Kumar Mandal and Pabitra Mohan Khilar. Efficient virtual machine placement for on-demand access to infrastructure resourcesin cloud computing. International Journal of Computer Applications, 68(??):6–11, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume68/number12/11629-7101/.

Nandakumar:2014:ESN

[MK14] Mona Nandakumar M and Edet Bijoy K. An experimental survey on non-negative matrix factorization for single channel blind source separation. International Journal of Computer Applications, 100(??):1–6, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number5/17518-8071/.

Mittal:2015:QSP

[MK15a] Mohit Mittal and Krishan Kumar. Quality of services provisioning in wireless sensor networks using artificial neural network: a survey. International Journal of Computer Applications, 117(??):28–40, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number5/20553-2931/.

Mohammadi:2015:IMS

[MK15b] Afshin Mohammadi and Keyhan Khamforoosh. Influence maximization in social networks using learning automata. International Journal of Computer Applications, 129(??):4–10, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume129/number8/23091-2015906898/.

Malled:2017:AVA

[MK17] Gulnaz Malled and Logesh K. Advantages of VCG auction based relay algorithm for DTNs. International Journal of Computer Applications, 178(??):20–30, November 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume178/number4/28662-2017915810/

.

Manjunath:2018:RFD

[MK18a] T. G. Manjunath and Ashok Kusagur. Robust fault detection of multilevel inverter using optimized radial basis function based artificial neural network in renewable energy power generation application. International Journal of Computer Applications, 180(??):8–15, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number48/29559-2018917231/.

Mittal:2018:PTP

[MK18b] Kajol Mittal and Rizwan Khan. Performance testing and profiling of Web based application in real time. International Journal of Computer Applications, 180(??):30–34, June 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number46/
29548-2018917220/.

Muhammed:2019:SHE

[MK19] Ibrahim A. Muhammed and Mohamad N. Abdul Kadir. Selected harmonics elimination for modified multilevel inverter with bi-directional switches. International Journal of Computer Applications, 178(??):28–34, July 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number32/30744-2019919187/.

Malik:2017:WAV

[MKA17] Gunjan Malik, Tarun Kumar, and Gaurav Agarwal. Wavelet analysis in video watermarking through wavelet transform scheme. International Journal of Computer Applications, 161(??):29–36, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number4/27140-2017913177/.

Melo:2022:SSU

[MKAR22] Olga Engelien Melo, Harson Kapoh, Ventje Ferdy Aror, and Ali Akbar Ramchie. Security system uses CCTV camera as facial image recognition using face-API. International Journal of Computer Applications, 184(??):10–14, December 2022. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume184/number41/32585-
2022922496/.

Mohamed:2013:CSF

[MKB13] Mchiri Mohamed, Trabelsi Karim, and Safya Belghith. Chaos suppression in forced van der pol oscillator. International Journal of Computer Applications, 68(??):18– 23, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume68/
number23/11719-7388/.

Mohammadani:2013:SIW

[MKCV13] Khalid Mohammadani, Hameedullah Kazi, Ibraheem Channa, and Danish Vasan. A survey on integrated wireless network architectures. International Journal of Computer Applications, 79(??):4–9, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume79/number4/13727-1523/.

Malik:2013:ODT

[MKK13] Asad Abdul Malik, Amaad Khalil, and Hameed Ullah Khan. Object detection and tracking using background subtraction and connected component labeling. International Journal of Computer Applications, 75(??):1–5, August 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume75/number13/13168-0421/.

Maassarani:2017:DCT

[MKM17] Sara Maassarani, Mostafa R. A. Khalifa, and Nabil Mohareb. Developing a calculation tool for embodied energy in the conceptual design phase. International Journal of Computer Applications, 157(??):41–49, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume157/number6/26839-2017912730/.

Maakar:2012:PAP

[MKN12] Sunil Kr. Maakar, Sudesh Kumar, and Amit Nain. Performance analysis of proactive and reactive MANET routing protocols using TCP traffic for freeway and random waypoint mobility model. International Journal of Computer Applications, 43(??):13–18, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number7/6114-8294/.

B:2014:PAP

[MKOS14] Kuboye B M, Alese B K, Falaki S O, and Adewale O S. Performance analysis of a prioritized multilevel access allocation scheme for data calls in GSM networks. International Journal of Computer Applications, 97(??):1–6, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume97/number1/16968-6533/.

Manginsela:2022:DIS

[MKSB22] Antonius P. G. Manginsela, Marike A. S. Kondoj, Maksy Sendiang, and Muhammad K. Bakary. Development of integrated system and synchronization of data study program with higher education’s external quality assurance system. International Journal of Computer Applications, 183(??):28–33, February 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number52/32283-2022921940/.

Melo:2017:FBP

[MKT17] Olga Melo, Anthon Arie Kimbal, and Robby Tangkudung. Functional and basis path method for testing on the development of WEB show window of crysanthemun flowers. International Journal of Computer Applications, 175 (??):1–6, October 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume175/number8/28505-2017915599/.

Misra:2015:CED

[MKWK15] Neeraj Kumar Misra, Mukesh Kumar Kushwaha, Subodh Wairya, and Amit Kumar. Cost efficient design of reversible adder circuits for low power applications. International Journal of Computer Applications, 117(??):37–45, May 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume117/number19/
20665-3408/.

Moein:2012:IES

[ML12] Sara Moein and Rajasvaran Logeswaran. Intelligent ECG signal noise removal using PSONN. International Journal of Computer Applications, 45(??):9–17, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume45/number6/6783-9085/.

Mwambe:2013:SAS

[ML13] Othmar Othmar Mwambe and Oleksandr Lutsaievskyi. Selection and application of software testing techniques to specific conditions of software projects. International Journal of Computer Applications, 70(??):22–28, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume70/number18/12168-8143/.

Mahar:2012:PAI

[MLRU12] Mukhtiar Ahmed Mahar, Abdul Sattar Larik, Mohammad Rafiq, and Mohammad Aslam Uqaili. Performance analysis and intelligent power controller design of soft switched half bridge dc/dc converter using neural networks. International Journal of Computer Applications, 46 (??):21–27, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number15/6985-9550/.

Malempati:2011:GBA

[MM11a] Sreelatha Malempati and Shashi Mogalla. Grid based approach for data confidentiality. International Journal of Computer Applications, 25(??):1–5, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume25/number9/3063-4186/.

Mourad:2011:SMC

[MM11b] Brik Mourad and Touahria Mohamed. Semantic metadata for creating and describing educational resources. International Journal of Computer Applications, 19(??): 13–17, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume19/
number9/2390-3169/.

Mahmoudpour:2012:HAM

[MM12a] Saeed Mahmoudpour and Sattar Mirzakuchaki. Hardware architecture for a message hiding algorithm with novel randomizers. International Journal of Computer Applications, 37(??):46–53, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume37/
number7/4623-6632/.

Moturi:2012:UMD

[MM12b] Christopher A. Moturi and Silas K. Maiyo. Use of mapreduce for data mining and data optimization on a Web portal. International Journal of Computer Applications, 56 (??):39–43, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number7/8906-2945/.

Mahajan:2013:RPU

[MM13] Seema Mahajan and Himanshu Mazumdar. Rainfall prediction using neural net based frequency analysis approach. International Journal of Computer Applications, 84(??):7– 11, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number9/14602-2851/.

Maadi:2014:OCH

[MM14a] Marjan Maadi and Masoureh Maadi. Optimization of cluster heads selection by imperialist competitive algorithm in wireless sensor networks. International Journal of Computer Applications, 89(??):29–34, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume89/number19/15742-4697/.

Maurya:2014:WAR

[MM14b] Laxmi Shanker Maurya and Anil Kumar Malviya. Web application reliability assessment using error and workload data obtained from server error and access logs. International Journal of Computer Applications, 97(??): 6–9, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number15/17081-7530/.

Maneva:2015:FAM

[MM15a] Svetla Maneva and Krassimir Manev. Finite automata models in agro-ecosystem and plant protection. International Journal of Computer Applications, 119(??):1–6, June 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume119/number18/21164-

Marfo:2015:CDT

[MM15b] Joseph Marfo and Ezekiel Mensah Martey. Creating designs through mathematical functions. International Journal of Computer Applications, 123(??):24–30, August 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume123/number3/21941-2015905272/

.

Misso:2015:DOO

[MM15c] Angelina Misso and Alfred Mwambela. Development of online optical fiber communications experiments using the interactive ILabs shared architecture. International Journal of Computer Applications, 131(??):6–13, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume131/number11/23491-2015907450/.

Maurya:2016:SVA

[MM16] Swati R. Maurya and Ganesh M. Magar. Segmentation and visualization of airborne sparse forest LiDAR data using region growing approach. International Journal of Computer Applications, 156(??):16–22, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume156/number5/26705-2016912434/.

Mishra:2017:LPW

[MM17] Vivek Mishra and Vivek Kumar Modanwal. Low power wide fan-in control pulse operated domino multiplexor with static switching. International Journal of Computer Applications, 172(??):23–29, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume172/number8/28272-2017915194/.

Matto:2018:MFP

[MM18] George Matto and Joseph Mwangoka. Mining frequent patterns of crime using FP-growth with multiple minimum supports based on Shannon entropy. International Journal of Computer Applications, 180(??):45–52, March 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume180/number24/29108-2018916577/.

Marquez:2020:DMR

[MM20a] A. Jiménez Márquez and G. Beltrán Maza. Data mining for rules extraction to virgin olive oil optimization process.

International Journal of Computer Applications, 176(??):24– 30, April 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume176/number14/
31271-2020920065/.

Marquez:2020:MLA

[MM20b] A. Jiménez Márquez and G. Beltrán Maza. Machine learning application in process on extra virgin olive oil elaboration disk stack vertical centrifuge modeling. International Journal of Computer Applications, 176(??):30–35, July 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume176/number37/31446-
2020920511/.

Mitchell:2021:ICC

[MM21] D. A. Mitchell and H. B. Mitchell. Inter-color context classifier for high performance lossless Bayer image compression. International Journal of Computer Applications, 183(??):1–7, July 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number14/31991-2021921451/.

Millar:2023:AFK

[MM23] Katie Millar and Dinesh Mothi. An algorithm for faster keyword detection on a forensic image. International Journal of Computer Applications, 185(??):38–45, June 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number15/32773-2023922842/.

Madi:2013:NPA

[MMA13] Abdessalam Ait Madi, Anas Mansouri, and Ali Ahaitouf. A new PWL approximation for the “Self-Adjustable offset min-sum” decoding with a highly reduced-complexity. International Journal of Computer Applications, 61(??): 1–6, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number19/10033-5068/.

Mammass:2016:MAS

[MMG16] Mouad Mammass, Hafid Mammass, and Fattehallah Ghadi. A multiservice access solution based on S-OrBAC model, contactless smartcard and NFC technologies. International Journal of Computer Applications, 151(??):40–44, October 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume151/number8/26257-2016911856/

.

Mirfakhrai:2012:PIF

[MMH12] Yasaman Mirfakhrai, Mehran Mohsenzadeh, and Seyed Mohsen Hashemi. Proposing an interoperable framework among cloud providers. International Journal of Computer Applications, 55(??):40–46, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume55/number11/8803-3032/.

Madadpour:2011:TAO

[MMHA11] Somayeh Madadpour, Seyed-Hassan Mirian-Hosseinabadi, and Vahdat Abdelzad. Testing aspect-oriented programs with UML activity diagrams. International Journal of Computer Applications, 33(??):4–11, November 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume33/number8/4038-5783/.

Mohapatra:2017:PEV

[MMHP17] Subasish Mohapatra, Subhadarshini Mohanty, Arunima Hota, and Shradha Pattanayak. Performance evaluation of virtualization in a private cloud test bed. International Journal of Computer Applications, 162(??): 20–27, March 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume162/number7/27256-2017913386/.

Mousavi:2011:NSF

[MMK11] B. Somayeh Mousavi, Amin Shahraki Moghadam, and Mohammad Mojtaba Keikhayfarzaneh. A new skin and frontal face detector by selecting optimized color spaces. International Journal of Computer Applications, 24(??): 23–28, June 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume24/
number2/2922-3859/.

Murori:2016:NCB

[MMK16] Francis K. Murori, Elijah Mwangi, and Patrick M. Karimi. A new computer-based ferromagnetic metal detector for security applications. International Journal of Computer Applications, 144(??):37–44, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number11/25227-2016910481/.

Mokua:2021:RWQ

[MMK21] Nahshon Mokua, Ciira Wa Maina, and Henry Kiragu. A raw water quality monitoring system using wireless sensor networks. International Journal of Computer Applications, 174(??):35–42, February 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number21/31799-2021921113/.

Mahajan:2015:EGE

[MML15] Sonika Mahajan, Rajesh Mehra, and Parveen K. Lehana. Effect of glottal excitation interchange in Hindi and Dogri languages. International Journal of Computer Applications, 116(??):1–8, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number9/20361-2563/.

Mahlangu:2018:AUI

[MMN18] Gilbert Mahlangu, Samuel Musungwini, and Sindiso Nleya. An algorithm for using Internet of Things (IoTs) to improve load management in electric power grid. International Journal of Computer Applications, 179(??):10–17, April 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume179/number35/
29225-2018916786/.

Mamalis:2021:EML

[MMP21] Basilis Mamalis, Sotiris Mamalis, and Marios Perlitis. Efficient multi-level clustering for very large wireless sensor networks with gateways support and meta-heuristic integration. International Journal of Computer Applications, 183(??):30–38, June 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number7/31942-2021921362/.

Mishra:2015:RSA

[MMPH15] Sujogya Mishra, Shakti Prasad Mohanty, Sateesh Kumar Pradhan, and Radhanath Hota. Rough set approach in finding the cause of decline and down fall of jute industries and the remedy. International Journal of Computer Applications, 121(??):35–41, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number19/21650-4900/.

Mahmood:2014:PKS

[MMR14a] Nesir Rasool Mahmood, Ali Abdul Azeez Mohammad, and Zahraa Nesir Rasool. Public key steganography. International Journal of Computer Applications, 100(??): 29–32, August 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume100/number8/17547-8143/.

Mehta:2014:TUC

[MMR14b] R. G. Mehta, N. J. Mistry, and M. Raghuwanshi. Towards unsupervised and consistent high dimensional data clustering. International Journal of Computer Applications, 87(??): 40–44, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number2/15183-3532/.

Mohamed:2014:DLA

[MMR14c] Marghny H. Mohamed, Yasmeen T. Mahmoud, and Saad Z. Rida. Destructive learning analysis and constructive algorithm for rule extraction based on a trained neural network using gene expression programming. International Journal of Computer Applications, 89(??):18–26, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume89/number17/15723-4602/.

Mfenyana:2013:DFC

[MMTS13] S. I. Mfenyana, N. Moroosi, M. Thinyane, and S. M. Scott. Development of a Facebook crawler for opinion trend monitoring and analysis purposes: Case study of government service delivery in dwesa. International Journal of Computer Applications, 79(??):32–39, October 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume79/number17/13964-1934/.

Mahmood:2015:PED

[MN15] Maher K. Mahmood and Besma Nazar Nathem. Performance evaluation of DS-CDMA system based orthogonal chaotic vectors over Rayleigh fading channel. International Journal of Computer Applications, 114(??):8–14, March 2015.

CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume114/number3/19956-0060/

.

Mahadev:2016:IFT

[MN16] Preeti Mahadev and P. Nagabhushan. Incremental feature transformation for temporal space. International Journal of Computer Applications, 145(??):28–38, July 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume145/number8/25300-2016910737/.

Mahadev:2017:IDR

[MN17] Preeti Mahadev and P. Nagabhushan. Incremental dimensionality reduction in hyperspectral data. International Journal of Computer Applications, 163(??):21–34, April 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume163/number7/27407-2017913575/

.

Prakash:2022:MOD

[MN22] Madhura Prakash M and Krishnamurthy G N. Multiobject detection and localization of artefacts in endoscopy images. International Journal of Computer Applications, 184(??):28–33, October 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number33/32527-2022922422/.

Masinde:2024:SOA

[MN24] Robbinson C. Masinde and Wanjiku J. Nganga. Serviceoriented architecture for real-time notification of home death events and integration. International Journal of Computer Applications, 186(??):48–57, December 2024. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.
org/archives/volume186/number51/service-oriented-
architecture-for-real-time-notification-of-home-death-
events-and-integration/.

Merzougui:2013:EAE

[MNB13a] M. Merzougui, M. Nasri, and B. Bouali. Entropic approach and evolution strategies for optimizing the image segmentation by pixel classification: Application to quality control. International Journal of Computer Applications, 61 (??):22–28, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number13/9989-4834/.

Merzougui:2013:ISU

[MNB13b] M. Merzougui, M. Nasri, and B. Bouali. Image segmentation using Isodata clustering with parameters estimated by evolutionary approach: Application to quality control. International Journal of Computer Applications, 66 (??):25–30, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number19/11194-6345/.

Mudumbe:2013:MVC

[MNB13c] John Mduduzi Mudumbe, James Ntaganda, and Cho Yong Beom. Multiview video coding optimization using SIMD on portable devices. International Journal of Computer Applications, 71(??):12–17, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number16/12440-8946/.

Sindhusree:2022:OPP

[MNB^{+}22] Sindhusree M, Aishwarya N, Aditya T B, Pavithra G, and T C Manjunath. An optimized path-planning concept for a micro-mouse robot for autonomous applications in terrained environments. International Journal of Computer Applications, 184(??):26–34, August 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number26/32477-2022922323/.

Mountassir:2012:WMN

[MNHB12] Tarik Mountassir, Bouchaib Nassereddine, Abdelkrim Haqiq, and Samir Bennani. Wireless mesh networks topology auto planning. International Journal of Computer Applications, 52(??):27–33, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number2/8175-1494/.

Muthukarpagam:2010:DIT

[MNN10] S. Muthukarpagam, V. Niveditta, and S. Neduncheliyan. Design issues, topology issues, quality of service support for wireless sensor networks: Survey and research challenges. International Journal of Computer Applications, 1(??):1–4, February 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume1/number6/151-272/.

Abiodun:2011:EGA

[MNP11] Ikotun Abiodun M, Lawal Olawale N, and Adelokun Adebowale P. The effectiveness of genetic algorithm in solving simultaneous equations. International Journal of Computer Applications, 14(??):38–41, February 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume14/number8/1900-2534/.

Mohammed:2014:PMD

[MNR14] Samir Jasim Mohammed, Mahmoud Shaker Nasr, and Ali Jasim Ramadhan. Portable monitoring/diagnostic system based on the combining of real time WSN with GSM and Internet. International Journal of Computer Applications, 101(??):1–8, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number2/17656-8469/.

Mousavi-Nik:2012:PSS

[MNYMGG12] Samaneh Sadat Mousavi-Nik, M. H. Yaghmaee-Moghaddam, and M. B. Ghaznavi-Ghoushchi. Proposed secure SIP authentication scheme based on elliptic curve cryptography. International Journal of Computer Applications, 58(??):25– 30, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number8/9303-3524/.

Machanje:2017:ASM

[MO17] Daniel Isoso Machanje and Joseph Onderi Orero. Automatic security monitoring and alert system through sound analysis. International Journal of Computer Applications, 169(??):1–12, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume169/number8/28002-2017914686/.

Mohamad:2013:EEQ

[Moh13a] Imad J. Mohamad. End-to-end QoS improvement using IPv6 header reduction over MPLS. International Journal of Computer Applications, 80(??):27–32, October 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume80/number11/13907-1839/.

Mohammed:2013:PCI

[Moh13b] Zaid Mohammed. Properties of color images pictured on sunlight and different wave lengths. International Journal of Computer Applications, 83(??):27–41, December 2013. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume83/number4/14439-2592/

.

Mohammed:2014:SPK

[Moh14] B. E. Mohammed. Statistical properties of Kumaraswamygeneralized exponentiated exponential distribution. International Journal of Computer Applications, 94(??): 1–8, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number4/16328-5602/.

Marube:2017:AUA

[MOM17] Fred Ondieki Marube, Okelo Odongo, and Lawrence G. Muchemi. An autonomous unmanned aerial security surveillance system to enhance security in remote territories. International Journal of Computer Applications, 179(??):9–16, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number6/28739-2017915964/.

Mambo:2018:IHS

[MOOOOMT18] Susan Njoki Mambo, George W. Odhiambo-Otieno, George Ochieng’-Otieno, and Wanja Mwaura-Tenermbergen. Improving health systems: Influence of technical capacities of community health volunteers on use of community health information systems in Kenya. International Journal of Computer Applications, 181(??):30–37, July 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number3/29699-2018917458/.

Mosaad:2011:CSE

[Mos11] M. I. Mosaad. Control of self excited induction generator using ANN based SVC. International Journal of Computer Applications, 23(??):22–25, June 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume23/number5/2883-3755/.

Mosavi:2014:DMC

[Mos14] Amir Mosavi. Decision-making in complicated geometrical problems. International Journal of Computer Applications, 87(??):22–25, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume87/number19/15460-4057/.

Manikandan:2010:ERP

[MP10] K. Manikandan and T. Purusothaman. An efficient routing protocol design for distributed wireless sensor networks. International Journal of Computer Applications, 10(??):5– 10, November 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume10/
number4/1472-1988/.

Mathivanan:2012:ESE

[MP12a] M. Mathivanan and S. Chenthur Pandian. Efficient speech enhancement approach based on minima controlled recursive averaging through modified map criterion using hidden Markov model. International Journal of Computer Applications, 44(??):27–34, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume44/number14/6333-8708/.

Mishra:2012:AAT

[MP12b] Nitin Mishra and C Patvardhan. ATMA: Android travel mate application. International Journal of Computer Applications, 50(??):1–8, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number16/7852-1083/.

Murugavalli:2015:CST

[MP15] N. Murugavalli and A. Pushpalatha. τ∗–gλ-closed sets in topological spaces. International Journal of Computer Applications, 125(??):28–32, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number14/22501-2015906266/.

Rajashri:2017:PEA

[MP17] Chaudhari Rajashri M and Patil Manesh P. Performance evaluation of attack detection algorithms in delay tolerant networks. International Journal of Computer Applications, 171(??):1–7, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume171/number4/28166-2017915008/.

Mota:2018:LES

[MPJ18] Josué Batista Mota, Renata Imaculada Soares Pereira, and Sandro César Silveira Jucá. Linux embedded system for vehicle license plates recognition. International Journal of Computer Applications, 182(??):43–46, August 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume182/number9/29851-2018917687/.

K:2014:CCM

[MPR14] Prabusankarlal K M, Thirumoorthy P, and Manavalan R. Combining clustering, morphology and metaheuristic optimization technique for segmentation of breast ultrasound images to detect tumors. International Journal of Computer Applications, 86(??):28–34, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume86/number14/15055-3465/.

Mazouni:2011:CVP

[MR11] Romaissaa Mazouni and Abdellatif Rahmoun. On comparing verification performances of multimodal biometrics fusion techniques. International Journal of Computer Applications, 33(??):24–29, November 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume33/number7/4033-5774/.

Mahapatra:2012:MJM

[MR12a] G. S. Mahapatra and P. Roy. Modified Jelinski–Moranda software reliability model with imperfect debugging phenomenon. International Journal of Computer Applications, 48(??):38–46, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number18/7451-0534/.

Misra:2012:SDS

[MR12b] Santanu Kr. Misra and Amitava Ray. Software developer selection: a holistic approach for an eclectic decision. International Journal of Computer Applications, 47 (??):12–18, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number1/7151-9855/.

Mal:2013:AWP

[MR13a] Sandip Mal and Kumar Rajnish. Applicability of Weyuker’s Property 9 to inheritance metric. International Journal of Computer Applications, 66(??):21–26, March 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume66/number12/11135-6214/.

Morampudi:2013:ESW

[MR13b] Naga Sri Morampudi and Gaurav Raj. Evaluating strengths and weaknesses of agile scrum framework using knowledge management. International Journal of Computer Applications, 65(??):1–6, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume65/number23/11221-6058/.

Mayreddy:2015:SSF

[MR15] Megha Mayreddy and B. Tarakeswara Rao. Secure SOA framework for multi-cloud storage and computing. International Journal of Computer Applications, 114(??): 10–16, March 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume114/number8/19997-1746/.

Mahmood:2014:DED

[MREd14] R. A. Mahmood, M. Z. Rashad, and M. A. El-dosuky. A dashboard of an education data portal using big data solutions. International Journal of Computer Applications, 90(??):1–5, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume90/
number19/15825-3633/.

Mustafa:2019:ICS

[MRG19] Hassan M. H. Mustafa, Hany S. Ramzy, and Satish Gajawada. On interdisciplinary comparative study of analogical feedback/assessment models applied in blended learning versus computer aided learning using artificial neural networks. International Journal of Computer Applications, 182(??):31–43, March 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number43/30440-2019918550/.

Mahmoud:2015:SAR

[MRH15] M. A. W. Mahmoud, A. M. Rashad, and Z. M. Hussien. Stochastic analysis of a repairable cold standby system attacked by Poisson shocks considering inspection and post repair. International Journal of Computer Applications, 132(??):33–40, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number11/23640-2015907578/.

Moghadam:2014:INT

[MRJN14] Mehdi Bahari Moghadam, Alireza Rostami, Edris Joonaki, and Payam Mohammad Ali Nejad. Investigation of a novel technique for decline curve analysis in comparison with the conventional models. International Journal of Computer Applications, 99(??):1–11, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume99/number18/17470-8316/.

Mohamed:2013:FRU

[MRM13] Benouis Mohamed, Tlmesani Redwan, and Senouci Mohamed. Face recognition using two-dimensional subspace analysis and PNN. International Journal of Computer Applications, 72(??):1–8, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume72/number6/12495-6019/.

Martins:2015:TME

[MRP15] Andre L. Martins, Evandro L. L. Rodrigues, and Maria Stela V. Paiva. True-motion estimation and compensation with multitemporal block-matching search. International Journal of Computer Applications, 111(??): 16–21, February 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume111/number1/19501-0667/.

Martins:2016:FED

[MRP16] André L. Martins, Evandro L. L. Rodrigues, and Maria Stela V. Paiva. A feathering effect detector based on local average and effective values. International Journal of Computer Applications, 146(??):28–35, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number1/25364-2016910609/.

Mia:2021:GAF

[MRR21] Shisir Mia, Md. Mahfuz Reza, and Mohammad Motiur Rahman. Genetic algorithm and Fisher discriminant analysis based wavelet thresholding for speckle noise filtering in ultrasound images. International Journal of Computer Applications, 174(??):13–18, April 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number30/31869-2021921231/.

Chandrajit:2016:CSS

[MRTB16] Chandrajit M, Girisha R, Vasudev T, and Ashok C B. Cast and self shadow segmentation in video sequences using interval based eigen value representation. International Journal of Computer Applications, 142(??):27–32, May 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume142/number4/24886-2016909752/.

Mahajan:2010:SCI

[MS10a] Manish Mahajan and Akashdeep Sharma. Steganography in colored images using information reflector with 2k correction. International Journal of Computer Applications, 1(??):48–53, February 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume1/number1/22-130/.

Maurya:2010:SBP

[MS10b] Ashish K. Maurya and Dinesh Singh. Simulation based performance comparison of AODV, FSR and ZRP routing protocols in MANET. International Journal of Computer Applications, 12(??):23–28, December 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume12/number2/1651-2220/.

Mamatha:2011:TPA

[MS11] G. S. Mamatha and S. C. Sharma. TODV: Performance analysis of a time on demand distance vector protocol for MANETS. International Journal of Computer Applications, 28(??):32–36, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume28/
number2/3359-4636/.

S:2012:OSL

[MS12] Anzar S M and Sathidevi P S. Optimal score level fusion using modalities reliability and separability measures. International Journal of Computer Applications, 51(??): 1–8, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number16/8123-1953/.

Mafteiu-scai:2013:NDM

[Ms13] Liviu Octavian Mafteiu-scai. A new dissimilarity measure between feature-vectors. International Journal of Computer Applications, 64(??):39–44, February 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume64/number17/10730-5734/.

Maadi:2014:EPM

[MS14a] Mansoureh Maadi and Marzieh Soltanolkottabi. Extension of PROMETHEE method for solving multi-objective optimization problems. International Journal of Computer Applications, 89(??):23–29, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume89/number11/15677-4430/.

Modarres:2014:PLH

[MS14b] Amir Farid Aminian Modarres and Mohsen Soryani. A pyramidal layered HMM for multiview human behavior recognition in asynchronous video streams. International Journal of Computer Applications, 96(??):34–40, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume96/number7/16808-6539/.

B:2016:SEI

[MS16] Chethana Kumara B M and Nagendraswamy H S. Sign energy images for recognition of sign language at sentence level. International Journal of Computer Applications, 139(??):44–51, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number2/24466-2016909118/.

Malankar:2017:QAW

[MS17] Nupur Rajan Malankar and Ruchi Shah. QoS analysis over WiMAX network with varying modulation schemes and efficiency modes. International Journal of Computer Applications, 162(??):9–16, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume162/number8/27262-2017913390/.

Mohamed:2016:OCI

[MSA16] M. A. Mohamed, A. S. Samrah, and M. I. Fath Allah. Optical color images encryption based on double random phase encoding. International Journal of Computer Applications, 149(??):6–15, September 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume149/number5/25991-2016911397/.

Mohamed:2017:NCB

[MSA17] M. A. Mohamed, A. S. Samrah, and M. I. Fath Allah. A novel chaotic based optical color image encryption technique. International Journal of Computer Applications, 157(??):13–21, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume157/number7/26842-2017912738/.

Nivaashini:2018:SGM

[MSA18] Nivaashini M, Soundariya R S, and Dinesh Kumar A. Silkworm growth monitoring smart sericulture system based on Internet of Things (IOT) and image processing. International Journal of Computer Applications, 180(??):29–33, February 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume180/number18/29035-
2018916514/.

Minavathi:2012:CMB

[MSD12] Minavathi, Murali S, and M. S. Dinesh. Classification of mass in breast ultrasound images using image processing techniques. International Journal of Computer Applications, 42(??):29–36, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number10/5731-7801/.

Mahmoud:2013:MCM

[MSEEs13] M. A. W. Mahmoud, A. A. Soliman, A. H. Abd Ellah, and R. M. El-sagheer. Markov chain Monte Carlo to study the estimation of the coefficient of variation. International Journal of Computer Applications, 77(??):31– 37, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number4/13384-1000/.

Mansour:2015:VRU

[MSM15] Abdelmajid H. Mansour, Gafar Zen Alabdeen Salh, and Khalid A. Mohammed. Voice recognition using dynamic time warping and mel-frequency cepstral coefficients algorithms. International Journal of Computer Applications, 116(??):34–41, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number2/20312-2362/.

Malwatkar:2017:HAR

[MSM17] Shweta Malwatkar, Rekha Sugandhi, and Anjali R. Mahajan. Human affect recognition system based on survey of recent approaches. International Journal of Computer Applications, 158(??):10–17, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume158/number6/26911-2017912801/.

Mehrotra:2012:NAI

[MSN12] Akansha Mehrotra, Krishna Kant Singh, and M. J. Nigam. A novel algorithm for impulse noise removal and edge detection. International Journal of Computer Applications, 38 (??):30–34, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume38/
number7/4701-6855/.

Mishra:2012:PAV

[MSP12] B. K. Mishra, S. K. Singh, and Kalpana Patel. Performance analysis of various codecs using RSVP on VoIP quality of service over variable bandwidth. International Journal of Computer Applications, 41(??):28–36, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume41/number7/5555-7630/.

Murthy:2016:PAM

[MSR16] K. Sreerama Murthy, S. Pallam Setty, and G. S. V. P. Raju. Performance analysis of mail clients with RSA and ElGamal using SNORT. International Journal of Computer Applications, 151(??):26–33, October 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume151/number11/26279-2016911945/.

Mohanapriya:2013:PAG

[MSS13] T. K. Mohanapriya, A. Sivanantharaja, and D. Shanmuga Sundar. Performance analysis of 200-gb/s low complexity transmission in second window. International Journal of Computer Applications, 68(??):5–8, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume68/number22/11709-7282/.

Mishra:2014:CER

[MSS14] Ruchika Mishra, Utkarsh Sharma, and Manish Shrivastava. Contrast enhancement of remote sensing images using DWT with kernel filter and DTCWT. International Journal of Computer Applications, 87(??):43– 49, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number17/15304-4075/.

Mishra:2013:TSD

[MT13a] Kamal Sheel Mishra and Anil Kumar Tripathi. Task scheduling of a distributed computing software in the presence of faults. International Journal of Computer Applications, 72(??):1–9, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number13/12551-8925/.

Misra:2013:NQB

[MT13b] P. Misra and A. Tripathy. A novel quad-band printed antenna design using a multi-slotted patch for cellular communication. International Journal of Computer Applications, 63(??):1–5, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume63/number2/10435-5113/.

Masri:2014:SUP

[MTD14] Ahmed Masri, Saed Tarapiah, and Yousef Dama. Secondary user power saving in overlay cognitive radio networks. International Journal of Computer Applications, 86 (??):1–5, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number7/14994-2788/.

Makwana:2013:CVE

[MTJ13] Hemant Makwana, Sanjay Tanwani, and Suresh Jain. Circular visualization enhancement through complementary interaction. International Journal of Computer Applications, 82 (??):1–5, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume82/
number3/14093-2101/.

Mourad:2015:SIC

[MTKP15] Bousaaid Mourad, Ayaou Tarik, Afdel Karim, and Estraillier Pascal. System interactive cyber presence for E learning to break down learner isolation. International Journal of Computer Applications, 111(??):35–40, February 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume111/number16/19626-1544/.

MousaviRad:2012:AIC

[MTM12] S. J. MousaviRad, F. Akhlaghian Tab, and K. Mollazade. Application of imperialist competitive algorithm for feature selection: a case study on bulk rice classification. International Journal of Computer Applications, 40(??):41– 48, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume40/
number16/5068-7485/.

Mugo:2024:OIC

[Mug24] Bob Chege Mugo. Optimization of industrial conveyor based SCADA system using machine learning techniques. International Journal of Computer Applications, 186(??):15–27, July 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number29/
optimization-of-industrial-conveyor-based-scada-system-
using-machine-learning-techniques/.

Mukherjee:2013:COM

[Muk13] J. K. Mukherjee. A computation optimization method in computer mediated teleoperation. International Journal of Computer Applications, 66(??):46–53, March 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume66/number24/11268-6562/.

Munoz:2014:LBO

[Mun14] Roberto Alejandro Espi Munoz. Linux based operating system proposal for the acquisition and processing of data in embedded devices. International Journal of Computer Applications, 93(??):1–4, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume93/number17/16424-5828/.

Mustafa:2013:FGT

[Mus13] Heba I. Mustafa. On L-fuzzy generalized topology. International Journal of Computer Applications, 69(??): 4–7, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number28/12245-8513/.

Mathi:2012:MIR

[MV12a] Senthil Kumar Mathi and M. L. Valarmathi. Mobile IP registration protocols: a survey. International Journal of Computer Applications, 51(??):24–34, August 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume51/number17/8135-1853/.

Mohan:2012:WKC

[MV12b] S. Mohan and K. Venkatachalapathy. Wood knot classification using bagging. International Journal of Computer Applications, 51(??):50–53, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume51/number18/8146-1937/.

Muthumari:2010:CIS

[MVC10] S. Muthumari, K. Vijayarekha, and Amrit Chatterjee. Comparative investigation of split spectrum processing over classical approach of filtering for non-linear & non-stationary signals. International Journal of Computer Applications, 10 (??):35–38, November 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume10/
number3/1459-1973/.

Ma:2020:LPR

[MWS20] Tingyu Ma, Tao Wang, and Jinchao Shi. License plate recognition system based on improved BP neural network. International Journal of Computer Applications, 176(??):27–34, May 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume176/number21/
31326-2020920204/.

Mostafa:2013:PSP

[MY13] Almetwally M. Mostafa and Ahmed E. Youssef. A primary shift protocol for improving availability in replication systems. International Journal of Computer Applications, 72 (??):37–44, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number4/12485-8905/.

Mousavi:2018:HMM

[MZ18] Alireza Shafii Mousavi and Houman Zarrabi. A hidden Markov model (HMM) scheme for lip based identification utilizing vertical grooves angles. International Journal of Computer Applications, 179(??):45–50, March 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number29/29165-2018916678/.

Mohsen:2016:DPD

[MZAE16] Saeed Mohsen, Abdelhalim Zekry, Mohamed Abouela, and Ahmed Elshazly. Development of a portable DAQ-based electroencephalogram system. International Journal of Computer Applications, 142(??):33–38, May 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume142/number5/24896-2016909810/.

Mahmoud:2013:ECC

[MZEE13] Eman Mohammed Mahmoud, Abdelhalim Zekry, Ahmed Abd El Hafez, and Talaat A. Elgarf. Enhancing channel coding using AES block cipher. International Journal of Computer Applications, 61(??):28–33, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume61/number6/9933-4568/.

Nasereddin:2016:EOS

[NA16] Hebah H. O. Nasereddin and Yaman F. Abdullah. Enhancing open space method in data hiding technique. International Journal of Computer Applications, 155(??): 5–17, December 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume155/number7/26615-2016912103/.

Nayeem:2018:WAW

[NA18] Mazharul Islam Nayeem and Mahfida Amjad. Water automation for water pump controller using Android application — review. International Journal of Computer Applications, 182(??):34–38, November 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number29/30167-2018918165/.

N:2020:ESV

[NA20] Ononiwu R N and Okengwu U A. Efficient steganography on video file using discrete cosine transform method (DCTM). International Journal of Computer Applications, 176(??):22–28, April 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number11/31246-2020920051/.

Nadafi:2014:ACE

[Nad14] Reza Nadafi. Adaptive control of an electrostatic microactuator with unbounded dynamic uncertainties based on Lyapunov stability criterion. International Journal of Computer Applications, 103(??):34–38, October 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume103/number1/18041-8924/.

Nada:2017:TUF

[Nad17] Yasser A. Nada. Taif university female golf transportation cars application: Analysis, design, and implementation. International Journal of Computer Applications, 157(??):1–9, January 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume157/number4/26816-2017912658/.

NobreMorais:2019:ISA

[NAD19] Elton Nobre Morais, Pedro Hericson M. Araújo, and Leonardo Da Silva Costa. IoT system with access control integrating RFID and non-invasive air conditioning. International Journal of Computer Applications, 178(??): 5–11, September 2019. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume178/number52/30905-2019919378/.

Noordin:2011:SEU

[NAH11] Syazwan Noordin, Wan Fatimah Wan Ahmad, and Yew Kwang Hooi. Study of effectiveness and usability of multimedia courseware integrated with 3-dimensional model as a teaching aid. International Journal of Computer Applications, 16 (??):20–27, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume16/
number4/2001-2697/.

Nahhas:2013:DPR

[Nah13] Ahmed M. Nahhas. Differential precision rectifier using single CMOS DVCC. International Journal of Computer Applications, 64(??):1–6, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume64/number7/10643-5400/.

Nageh:2015:PSP

[NAH15] Mohamed Nageh, Wagdy R. Anis, and Ismail M. Hafez. A proposed solution for partial shadowing. International Journal of Computer Applications, 121(??):28–38, July 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume121/number18/21642-
4974/.

Nagy:2013:PAD

[NAhA13] Amr A. Nagy, Gamal M. Abdel-hamid, and Ahmed E. Abdalla. Performance analysis of DSTBC over VHF airto-ground communication channels. International Journal of Computer Applications, 61(??):34–38, January 2013. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume61/number9/9958-
4604/.

Najafabadi:2012:EFP

[Naj12] Mahdi Imanian Najafabadi. The estimated formation pressure during drilling operations, in drilling wells, and an exploration of steps towards efficient drilling. International Journal of Computer Applications, 44(??):24–28, April 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume44/number7/6276-

Nassar:2014:ICV

[NAK^{+}14] S. S. Nassar, N. M. Ayad, H. M. Kollash, O. S. Faragallah, and M. A. M. El-bendary. Image contents verification algorithm using transform domain techniques. International Journal of Computer Applications, 97(??): 1–8, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number9/17032-7333/.

Naserasadi:2011:ASE

[Nas11] Ali Naserasadi. Authorís security in electronic learning systems. International Journal of Computer Applications, 21 (??):25–29, May 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume21/
number10/2617-3458/.

Nader:2017:ACI

[NAZ17] Jihad Nader, Ziad A. A. Alqadi, and Bilal Zahran. Analysis of color image filtering methods. International Journal of Computer Applications, 174(??):12–17, September 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume174/number8/28426-2017915449/.

Nzeugaing:2011:OFC

[NB11] Gutembert Nganpet Nzeugaing and Elmarie Biermann. Optimized filtering and compression system for an earth observation satellite. International Journal of Computer Applications, 36(??):32–36, December 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume36/number10/4527-6446/.

Nassef:2014:ABR

[NBF14] Mohammad Nassef, Amr Badr, and Ibrahim Farag. An algorithm for browsing the referentially-compressed genomes. International Journal of Computer Applications, 86(??):1– 10, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number8/15002-3134/.

Neji:2013:FIC

[NBKM13] Nihel Neji, Anis Boudabous, Wajdi Kharrat, and Nouri Masmoudi. FPGA implementation of the CORDIC algorithm for fingerprints recognition systems. International Journal of Computer Applications, 63(??):39–45,

February 2013.  CODEN ????  ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number6/10473-5204/.

Nayak:2015:IBH

[NBL15] Padmalaya Nayak, V. Bhavani, and B. Lavanya. Impact of black hole and sink hole attacks on routing protocols for WSN. International Journal of Computer Applications, 116(??):42–46, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number4/20327-2495/.

Nagaraj:2010:FME

[NBR10] Srinivasan Nagaraj, Kishore Bhamidipati, and M. Ramachandra. Formal method of encryption using 9’s complement. International Journal of Computer Applications, 8(??):23–25, October 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume8/number5/1206-1728/.

Narwade:2014:RAS

[NC14] Ashish A. Narwade and Vrishali A. Chakkarwar. Result analysis of smoke detection in video for early warnings using static and dynamic features. International Journal of Computer Applications, 95(??):12–18, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number24/16741-6950/.

Namwongse:2011:SAE

[NCL11] Paradee Namwongse, Thitivadee Chaiyawat, and Yachai Limpiyakorn. Scenario analysis in expressway management via modeling Bayesian networks. International Journal of Computer Applications, 30(??):14–20, September 2011. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume30/number11/3695-5187/

.

Nithya:2011:IGB

[ND11] N. S. Nithya and K. Duraiswamy. An information gain based fuzzy classifier for predictive analysis in colon cancer data. International Journal of Computer Applications, 31 (??):45–48, October 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume31/
number6/3833-5329/.

Nettleton:2015:SGM

[ND15] David F. Nettleton and Anton Dries. A sub-graph matching method based on calibration of characteristics of topological footprint. International Journal of Computer Applications, 130(??):29–38, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number10/23248-2015907098/.

Neama:2017:HDR

[ND17] Azhar Hussein Neama and Ismael Shanan Desher. High data rate optical logic OR, and NOT gates at optimum injection current based on SOA-MZI. International Journal of Computer Applications, 177(??):24–32, November 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number3/28607-2017915692/.

Nayak:2010:CBO

[NDC10] G. Subramanya Nayak, Ottolina Davide, and Puttamadappa C. Classification of bio optical signals using K-means clustering for detection of skin pathology. International Journal of Computer Applications, 1(??):92–96, February 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume1/number2/32-141/.

Najari:2017:AFT

[NEA17] Hajer Najari, Rafika El Harabi, and Mohamed Naceur Abdelkrim. An active fault tolerant control strategy based on bond graph adaptive observers. International Journal of Computer Applications, 176(??):1–7, October 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number6/28553-2017915631/.

Nebti:2013:BIA

[Neb13] Salima Nebti. Bio-inspired algorithms for color image segmentation. International Journal of Computer Applications, 73(??):11–16, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number18/12840-9810/.

Negi:2013:IRR

[Neg13] Sandeep Negi. An improved round robin approach using dynamic time quantum for improving average waiting time. International Journal of Computer Applications, 69 (??):12–16, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number14/11909-8007/.

Nofal:2015:DWB

[NF15] Mostafa A. Nofal and Khaled M. Fouad. Developing Web-based semantic and fuzzy expert systems using proposed tool. International Journal of Computer Applications, 112(??):38–45, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number7/19682-1414/.

Nair:2017:MBA

[NF17] Ranjini S. Nair and Shaniba Fasal. Mobile banking and its adopting challenges. International Journal of Computer Applications, 160(??):24–30, February 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume160/number4/27062-2017913036/.

Nair:2010:LBY

[NG10] Mydhili K. Nair and V. Gopalakrishna. Look before you leap: a survey of Web service discovery. International Journal of Computer Applications, 7(??):22–30, September 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume7/number5/1159-1508/

.

Negi:2014:RNP

[NGA14] Ashish Negi, Ankit Garg, and Akshat Agrawal. A review on natural phenomenon of fractal geometry. International Journal of Computer Applications, 86(??):1– 7, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number4/14970-3157/.

Nguyen:2016:CSO

[Ngu16] Duc Hoang Nguyen. A comparison of DE and SFLA optimization algorithms in tuning parameters of fuzzy logic controller. International Journal of Computer Applications, 156(??):17–22, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume156/number11/26753-2016912557/.

Nguyen:2018:DPI

[Ngu18] Duc Hoang Nguyen. Dynamic parameter identification of UP6 robot manipulator using SFLA. International Journal of Computer Applications, 182(??):34–39, November 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number27/30150-
2018918115/.

Nguyen:2021:IKS

[Ngu21] Duc Hoang Nguyen. Inverse kinematics of SCARA manipulator based on SFLA. International Journal of Computer Applications, 183(??):20–25, November 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume183/number38/32179-2021921780/.

Nguyen:2022:SEM

[Ngu22] Duc Hoang Nguyen. Simulation and evaluation of the mattress manufacturing process design model. International Journal of Computer Applications, 184(??):16–21, November 2022. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume184/number36/
32548-2022922454/.

Nguyen:2023:MBF

[Ngu23] Duc Hoang Nguyen. Modified DE-based fuzzy PD controller for UP6 manipulator. International Journal of Computer Applications, 185(??):36–40, November 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number38/32941-2023923190/.

Nguyen:2024:ISF

[Ngu24a] Duc Hoang Nguyen. Improved shuffled frog leaping algorithm with self-adaptive shuffling for fuzzy logic PD+G controller optimization in robotic manipulators. International Journal of Computer Applications, 186(??):21– 26, November 2024. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume186/number47/improved-shuffled-frog-leaping-
algorithm-with-self-adaptive-shuffling-for-fuzzy-

logic-pdg-controller-optimization-in-robotic-manipulators/

.

Nguyen:2024:SBL

[Ngu24b] Duc Hoang Nguyen. SFLA-based line balancing and task optimization in master-slave controlled hanger transportation systems for garment production. International Journal of Computer Applications, 186(??):57– 63, December 2024. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume186/number53/sfla-based-line-balancing-and-
task-optimization-in-master-slave-controlled-hanger-
transportation-systems-for-garment-production/.

Neto:2013:DSR

[NH13] Leoncio Claro Barros Neto and Andre Riyuiti Hirakawa. Data structures in robot navigation optimized by adaptive straightness. International Journal of Computer Applications, 62(??):1–9, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume62/number11/10121-4906/.

Nabavi:2010:SWM

[NHH10] Seyed Mohammad Hosseini Nabavi, Nazanin Alsadat Hosseinipoor, and Somayeh Hajforoosh. Social welfare maximization by optimal locating and sizing of TCSC for congestion management in deregulated power markets. International Journal of Computer Applications, 6(??):16–20, September 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume6/number6/1082-1202/

.

Noori:2013:VLK

[NHMR13] Sheak Rashed Haider Noori, Khandaker Tabin Hasan, Md. Saddam Hossain Mukta, and Md. Saifur Rahman. Visualization of linguistic and knowledge using DAG of concept/ontology. International Journal of Computer Applications, 65(??):12–19, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume65/number13/10983-6137/.

N:2011:ADV

[NHS11] Manjunath T N, Ravindra S. Hegadi, and Mohan H S. Automated data validation for data migration security. International Journal of Computer Applications, 30(??):41– 46, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume30/
number6/3642-5088/.

Naciri:2014:TPA

[NI14] Samia Naciri and M. A. Janati Idrissi. Third-party application maintenance management. International Journal of Computer Applications, 103(??):1–7, October 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume103/number6/18075-8766/.

Nuc:2015:RDS

[NI15] Andrei Nuc and Cosmina Ivan. REVEDERE — distributed support system for code review process. International Journal of Computer Applications, 115(??):1–6, April 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume115/number14/20216-
2488/.

Nima:2015:NIT

[Nim15] Mohammed A. Nima. Numerical investigation of transient phase change in horizontal porous channel with localized heating using two-equation model. International Journal of Computer Applications, 131(??):5–16, December 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume131/number5/23443-2015907334/

.

Nagori:2016:BTD

[NJ16] Meghana Nagori and Madhuri S. Joshi. Brain tumour disease pattern identification from metabolites in magnetic resonance spectroscopy graph using data mining techniques. International Journal of Computer Applications, 138(??):19–22, March 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume138/number13/
24440-2016909027/.

Nandan:2015:CSD

[NJAK15] Harshshikha Nandan, Manisha Jindal, Arsh, and Debanjan Konar. A comparative study on the development of binary object extraction system using different self organizing neural network. International Journal of Computer Applications, 117(??):24–29, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number22/20687-3554/.

Nikoui:2016:PCB

[NJB16] Tina Samizadeh Nikoui, Sam Jabbehdari, and Alireza Bagheri. Providing a cloud broker-based approach to improve the energy consumption and achieve a green cloud computing. International Journal of Computer Applications, 138(??):42–49, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume138/number1/24347-2016908715/.

Nosrat:2012:HDM

[NK12] Arash Nosrat and Yousef S. Kavian. Hardware description of multi-directional fast Sobel edge detection processor by VHDL for implementing on FPGA. International Journal of Computer Applications, 47(??):1–7, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume47/number25/7533-9872/.

Nguyen:2014:UMS

[NK14a] Thanh Trung Nguyen and Takashi Komeda. Using motor speed profile and genetic algorithm to optimize the fuzzy logic controller for controlling DC servomotor. International Journal of Computer Applications, 94 (??):1–8, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume94/
number14/16408-6076/.

Noureddin:2014:SWC

[NK14b] Amaigarou Noureddin and Mohamed Khaldi. S@J-Wiki: a collaborative semantic wiki to manage juridical knowledge. International Journal of Computer Applications, 104(??):44–49, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number6/18209-9371/.

Naick:2015:AAT

[NK15] B. Suribabu Naick and P. Rajesh Kumar. An algorithmic approach towards construction of long binary sequences using modified Jacobi sequences. International Journal of Computer Applications, 120(??):8–15, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume120/number1/21189-3836/.

Nakai:2019:DWV

[NK19] Yuichi Nakai and Kuya Kohara. Digital watermarking for vector drawings based on discrete Haar wavelet transform. International Journal of Computer Applications, 182(??):1–6, April 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number49/30525-2019918752/.

Nara:2014:ESP

[NKD14] Sanjeev Nara, Manvinder Kaur, and Saurav Dutta. Effect of self preferred romantic music after physical exertion in males. International Journal of Computer Applications, 88(??):27–32, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume88/number17/15447-4063/.

N:2012:PSI

[NKG12] Asha N, M Varun Kumar, and Vaidhyanathan G. Preventing SQL injection attacks. International Journal of Computer Applications, 52(??):28–32, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume52/number13/8264-1809/.

Nafi:2012:ADL

[NKH12] Kawser Wazed Nafi, Tonny Shekha Kar, and Sayed Anisul Hoque. An advanced door lock security system using palmtop recognition system. International Journal of Computer Applications, 56(??):18–26, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume56/number17/8983-3008/.

Nabavi:2011:CMU

[NKM11] S. M. H. Nabavi, A. Kazemi, and M. A. S. Masoum. Congestion management using genetic algorithm in deregulated power environments. International Journal of Computer Applications, 18(??):19–23, March 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume18/number2/2257-2894/.

Njuguna:2012:WAG

[NKM12] Peter N. Njuguna, Andrew M. Kahonge, and Evans K. Miriti. Web application and GPS integration in motor vehicle accident detection — a case of Nairobi, Kenya.

International Journal of Computer Applications, 59(??):6– 11, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number8/9566-4036/.

Nadh:2014:ATD

[NKR14] K. Lakshmi Nadh, Y. K. Sundara Krishna, and K. Nageswara Rao. Acknowledgment time delay approach to optimize TCP performance in hybrid networked systems. International Journal of Computer Applications, 102(??): 5–11, September 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume102/number8/17833-8695/.

N:2015:IHS

[NL15] Shamala N and C. Lakshminarayana. An insight to harmonic suppression techniques with power filters in power electronics. International Journal of Computer Applications, 111(??):26–34, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number9/19566-1334/.

Negi:2012:CIC

[NLC12] Ashish Negi, Shashank Lingwal, and Yashwant Singh Chauhan. Complex and inverse complex dynamics of fractals using Ishikawa iteration. International Journal of Computer Applications, 40(??):1–9, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume40/number12/5013-7335/.

N:2011:SIP

[NM11] Velvizhi N and D. Manjula. Securing information in peer-to-peer environment. International Journal of Computer Applications, 21(??):30–34, May 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume21/number8/2531-3448/.

Naushad:2013:CDA

[NM13] Alamgir Naushad and Tufail Muhammad. Condition driven adaptive music generation for computer games. International Journal of Computer Applications, 64(??):6– 10, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number8/10652-5416/.

Nasr:2014:CPI

[NM14] Ayman H. Nasr and Mohamed R. Metwalli. Comparative performance of the integration of ETM-8 and ERS-1 data for geological application. International Journal of Computer Applications, 94(??):1–8, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume94/number7/16352-5721/.

Noorikhameneh:2016:PLW

[NM16] Mostafa Noorikhameneh and Morteza Moosavi. Prolong the lifetime of wireless sensor network using fuzzy logic to control the movement of the mobile sink. International Journal of Computer Applications, 147(??):5–14, August 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume147/number7/25663-2016911166/.

Negm:2019:SOS

[NM19] Noha Negm and Hany Mahgoub. Survey on online social networks analysis concepts and knowledge discovery techniques. International Journal of Computer Applications, 178(??):12–21, May 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number14/30597-2019918904/.

Nayak:2020:SAP

[NM20] Shriniwas Nayak and Aditya Mahaddalkar. Statistical approach for predicting the most accurate classification algorithm for a data set in analysis. International Journal of Computer Applications, 176(??):1–7, June 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number28/31373-2020920306/.

Narvekar:2024:CPE

[NM24] Aditya Narvekar and Shubh Mehta. Cat2Vec with position encoding: a new approach for handling ordinal features using learned embeddings with positional encoding. International Journal of Computer Applications, 186(??):9–15, October 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number44/
cat2vec-with-position-encoding-a-new-approach-for-
handling-ordinal-features-using-learned-embeddings-
with-positional-encoding/.

Nunoo-Mensah:2014:RON

[NMAB14] Henry Nunoo-Mensah, Emmanuel Kofi Akowuah, and Kwame Osei Boateng. A review of opensource network access control (NAC) tools for enterprise educational networks. International Journal of Computer Applications, 106(??):28–33, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number6/18526-9721/.

Nemati:2013:BBH

[NMB13] Mostafa Nemati, Hossein Momeni, and Navid Bazrkar. Binary black holes algorithm. International Journal of Computer Applications, 79(??):36–42, October 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume79/number6/13748-1561/.

Naseriparsa:2013:HFS

[NmBV13] Mehdi Naseriparsa, Amir masoud Bidgoli, and Touraj Varaee. A hybrid feature selection method to improve performance of a group of classification algorithms. International Journal of Computer Applications, 69(??): 28–35, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number17/12065-8172/.

Nazari:2013:SIT

[NmEM13] Sara Nazari, Amir masoud Eftekhari, and Mohammad Shahram Moin. Secure information transmission using steganography and morphological associative memory. International Journal of Computer Applications, 61(??):23–29, January 2013. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume61/number7/9940-
4572/.

Nombo:2014:GLF

[NMK14] Josiah Nombo, Alfred Mwambela, and Michael Kisangiri. A grey level fitting mechanism based on Gompertz function for two phase flow imaging using electrical capacitance tomography measurement systems. International Journal of Computer Applications, 101(??):7–12, September 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume101/number8/17705-8704/.

Nombo:2015:PAG

[NMK15] Josiah Nombo, Alfred Mwambela, and Michael Kisangiri. Performance analysis of grey level fitting mechanism based Gompertz function for image reconstruction algorithms in electrical capacitance tomography measurement system. International Journal of Computer Applications, 109(??):9–14, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume109/number15/19263-0960/.

Niknazar:2015:ESP

[NMN15] Hamid Niknazar, Keivan Maghooli, and Ali Motie Nasrabadi. Epileptic seizure prediction using statistical behavior of local extrema and fuzzy logic system. International Journal of Computer Applications, 113(??):24–30, March 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume113/number2/19799-1578/.

Nair:2010:DTR

[NMS10] Binoy B. Nair, V. P. Mohandas, and N. R. Sakthivel. A decision tree- rough set hybrid system for stock market trend prediction. International Journal of Computer Applications, 6(??):1–6, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume6/number9/1106-1449/.

Nazir:2012:IER

[NMSA12] Mian Hammad Nazir, Shahid Mehmood, Nathirulla Sheriff, and Ahsan Adeel. Improved efficient RFID tag estimation scheme. International Journal of Computer Applications, 47(??):16–19, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume47/number17/7280-0380/.

Nguyen:2016:WMN

[NNR16] Duy-Huy Nguyen, Hang Nguyen, and Éric Renault. WE-MQS: a new LTE downlink scheduling scheme for voice services based on user perception. International Journal of Computer Applications, 142(??):28–36, May 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume142/number10/24934-2016909948/.

Nwiabu:2018:PFO

[NO18] Nuka D. Nwiabu and Okpu E. Okpomo. Predicting fire outbreak caused by electrical faults using artificial bee colony algorithm. International Journal of Computer Applications, 179(??):9–16, January 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number14/28866-2018916197/.

Nonum:2016:CWN

[NOO16] E. O. Nonum, P. O. Otasowie, and K. C. Okafor. Campus wireless network classification for enterprise adoption: Perspectives and dimensions for large scale computing. International Journal of Computer Applications, 142(??):19–31, May 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume142/number12/
24948-2016909953/.

Norton:2013:PPT

[Nor13] Andre Norton. Proactive policing in the Trinidad and Tobago Police Service (TTPS) using “Big Data”. International Journal of Computer Applications, 72(??): 43–46, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number18/12646-9453/.

Narasimhulu:2011:RWT

[NP11] C. Venkata Narasimhulu and K. Satya Prasad. A robust watermarking technique based on nonsubsampled contourlet transform and SVD. International Journal of Computer Applications, 16(??):27–36, February 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume16/number8/2032-2642/.

Neelamegam:2012:ESQ

[NP12] C. Neelamegam and M. Punithvalli. Enhanced software quality metrics for fault prediction in object oriented components using SVM classifier. International Journal of Computer Applications, 46(??):30–36, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume46/number22/7099-9834/.

N:2013:PES

[NP13] Thomas Rincy N and Yogadhar Pandey. Performance evaluation on state of the art sequential pattern mining algorithms. International Journal of Computer Applications, 65 (??):8–15, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number14/10990-6157/.

Nema:2015:CKD

[NP15] Rahul Nema and Anjana Pandey. Community kernels detection in OSN using SVM clustering and classification. International Journal of Computer Applications, 113(??):9–13, March 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume113/number11/
19869-1854/.

N:2019:PPU

[NP19] Ramya Shree A N and Kiran P. Privacy preserving unstructured data publishing (PPUDP) approach for big data. International Journal of Computer Applications, 178(??):4–9, June 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number28/
30711-2019919091/.

Nishat:2024:SLP

[NPA^{+}24] Nazma Hossen Nishat, Pranta Paul, Farzina Akther, Tahmina Akter, and Muhammad Anwarul Azim. Skin lesion prediction from dermoscopic images using deep learning. International Journal of Computer Applications, 186(??):17– 29, April 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number18/
skin-lesion-prediction-from-dermoscopic-images-using-
deep-learning/.

Ng:2012:ECD

[NR12] Kin Wei Ng and Ahmad Rohanin. The effects of control domain size on optimal control problem of monodomain model. International Journal of Computer Applications, 47 (??):6–11, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number10/7222-0046/.

Nagalakshmi:2016:ADE

[NR16] G. Nagalakshmi and M. Babu Rao. Automatic detection of enrolled distorted fingerprints. International Journal of Computer Applications, 147(??):19–24, August 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume147/number3/25634-2016911053/.

Naik:2017:USS

[NRL17] Milind Naik, Govind Ranade, and Rajesh B. Lohani. Underwater sonar systems: a methodology to generate bathymetry and sub-strata data. International Journal of Computer Applications, 170(??):1–5, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume170/number3/28047-2017914724/.

Narayana:2012:AIS

[NRP12] Ch. V. Narayana, E. Sreenivasa Reddy, and M. Seetharama Prasad. Automatic image segmentation using ultra fuzziness. International Journal of Computer Applications, 49 (??):6–13, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number12/7677-0977/.

Nagaraj:2010:DTM

[NRR^{+}10] Srinivasan Nagaraj, K. Koteswara Rao, G Appa Rao, G. Anuradha, and GSVP Raju. Designing a thread migration facility for Linux operating system. International Journal of Computer Applications, 1(??):89–94, February 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume1/number27/493-806/.

Nimbhore:2012:PED

[NRR12] S. S. Nimbhore, G. D. Ramteke, and R. J. Ramteke. Pitch estimation of devnagari vowels using cepstral and AutocorrelationTechniques for original speech signal. International Journal of Computer Applications, 55(??):38– 43, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume55/
number17/8849-3118/.

Narendar:2012:DHP

[NRRM12] V Narendar, Wanjul Dattatray R, Sanjeev Rai, and R. A. Mishra. Design of high-performance digital logic circuits based on FinFET technology. International Journal of Computer Applications, 41(??):40–44, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume41/number20/5812-8104/.

Nesakumari:2012:INC

[NRV12] G. Rosline Nesakumari, S. Rajendran, and V. Vijayakumar. Integrated normalized content system for efficient watermarking. International Journal of Computer Applications, 53(??):33–39, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume53/number15/8500-2448/.

Nagarajan:2013:SHG

[NS13a] S. Nagarajan and T. S. Subashini. Static hand gesture recognition for sign language alphabets using edge oriented histogram and multi class SVM. International Journal of Computer Applications, 82(??):28–35, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume82/number4/14106-2145/.

Nandanwar:2013:AEA

[NS13b] Jashweeni Nandanwar and Urmila Shrawankar. Aggregation of EDF and ACO for enhancing real time system performance. International Journal of Computer Applications, 73(??):24–33, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number17/12834-9974/.

Nogorabi:2013:ICE

[NS13c] Elham Masoumi Nogorabi and Hedieh Sajedi. Increasing classifier ensemble efficiency using KSBC algorithm. International Journal of Computer Applications, 70(??): 43–50, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number13/12025-8088/.

Nyamawe:2014:RIR

[NS14] Ally S. Nyamawe and Hassan Seif. The role of ICT in reducing maternal and neonatal mortality rate in Tanzania. International Journal of Computer Applications, 95 (??):39–42, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number13/16658-6639/.

N:2015:ERZ

[NS15a] Anand Kumar N and A. Grace Selvarani. Efficient routing in Zigbee wireless network using shortcut tree routing. International Journal of Computer Applications, 117 (??):23–28, May 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume117/number2/20527-2864/.

Naji:2015:MMD

[NS15b] Ahmed Mohammed Naji and N. D. Soner. The minimum monopoly distance energy of a graph. International Journal of Computer Applications, 128(??):1–6, October 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume128/number3/22850-2015906457/

.

Naresh:2016:EIW

[NS16] Shalki Naresh and Navjot Singh. Evaluating impact of wormhole on distance error and connectivity of network. International Journal of Computer Applications, 148(??):22–28, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume148/number7/25770-2016911213/.

Nooraldaim:2023:ETB

[NS23] Almonzer Salah Nooraldaim and Adil Ali Saed. An electronic tool based on bloom taxonomy and blueprint for producing lesson plan. International Journal of Computer Applications, 185(??):23–27, May 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number7/32714-2023922723/.

Noviana:2024:PCR

[NS24] Reski Noviana and Enny Itje Sela. Performance comparison random forest and logistic regression in predicting time deposit customers with feature selection. International Journal of Computer Applications, 186(??):33–38, April 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number16/
performance-comparison-random-forest-and-logistic-
regression-in-predicting-time-deposit-customers-with-
feature-selection/.

Naghdi:2014:FDI

[NSA14] Maryam Naghdi, Mohamad Ali Sadrnia, and Javad Askari. Fault detection and isolation for nonlinear system via ESO. International Journal of Computer Applications, 88(??):8– 13, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume88/
number16/15434-3663/.

Nayak:2013:CAB

[NSM13] Deepak Ranjan Nayak, Sumit Kumar Sahu, and Jahangir Mohammed. A cellular automata based optimal edge detection technique using twenty-five neighborhood model. International Journal of Computer Applications, 84(??):27– 33, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number10/14614-2869/.

Niknam:2014:IAP

[NSM14] Behzad Niknam, Kourosh Shahriar, and Eng Hassan Madani. An intelligence approach to predict fire flame length under tunnel ceiling. International Journal of Computer Applications, 106(??):39–43, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number18/18709-9859/.

Nouaiti:2017:ISP

[NSMK17] Ayoub Nouaiti, Abdallah Saad, Abdelouahed Mesbahi, and Mohamed Khafallah. Implementation of a single phase switched-capacitor nine-level inverter for PV system applications with selective harmonic elimination. International Journal of Computer Applications, 168(??):9–15, June 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume168/number7/27885-2017914416/

.

Natarajan:2015:LEB

[NSS15] Muralidaran Natarajan, Nandlal L. Sarda, and Sharad C. Srivastava. Lean and efficient business tier for performance scaling. International Journal of Computer Applications, 118(??):31–40, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number7/20760-3182/.

Nayyar:2017:ERD

[NSS17] Tushar Nayyar, Sarvpreet Singh, and Karamdeep Singh. Extraction of random data from .png/.tif/.jpeg image using Prewitt operator. International Journal of Computer Applications, 160(??):7–12, February 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume160/number2/27043-2017912966/.

Ntaganda:2014:QPD

[Nta14] James Ntaganda. A quick pre-deployment fibre optic link design methodology based on Q-factor in a digital eye. International Journal of Computer Applications, 90 (??):8–16, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume90/
number13/15778-4472/.

Nower:2018:ESD

[NTL18] Naushin Nower, Yasuo Tan, and Yuto Lim. Efficient spatial data recovery scheme with error refinement for cyberphysical systems. International Journal of Computer Applications, 179(??):34–40, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number26/29100-2018916558/.

Nagm:2019:NWA

[NTY19] Ahmad M. Nagm, Mohamed Torky, and Khaled Y. Youssef. A novel watermarking approach for protecting image integrity based on a hybrid security technique. International Journal of Computer Applications, 178(??):14–22, July 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume178/number30/30727-
2019919136/.

Nanthini:2012:BIB

[NU12] V. P. Nanthini and V. Rhymend Uthariaraj. Behavioral identification of budding professionals through “Neo personality inventory traits”. International Journal of Computer Applications, 56(??):41–46, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume56/number12/8946-3112/.

Nwiabu:2018:TLC

[NU18] Nuka D. Nwiabu and Emem E. Udoudom. Traffic light control system using genetic algorithm. International Journal of Computer Applications, 182(??):37–43, October 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number22/30069-
2018918040/.

Nair:2016:MLB

[NUK16] Anu B. Nair, R. Umamaheswari, and P. Kuppusamy. A machine learning-based state-of-the-art approach to identifying the person behind an e-mail ID. International Journal of Computer Applications, 136(??):1–4, February 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume136/number5/24146-2016908431/.

Nurmaini:2011:ILC

[Nur11] Siti Nurmaini. Intelligent low cost mobile robot and environmental classification. International Journal of Computer Applications, 35(??):1–7, December 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume35/number12/4545-6280/.

Narasimhan:2012:QAR

[NV12] B. Narasimhan and R. Vadivel. QoS aware reliable cluster based routing protocol (QoS-RCBRP) for heterogeneous mobile ad-hoc networks. International Journal of Computer Applications, 59(??):22–27, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume59/number12/9601-4226/.

Nair:2012:DFL

[NVR12] Vishnu G Nair, Dileep M V, and Prahalad K R. Design of fuzzy logic controller for lateral dynamics control of aircraft by considering the cross-coupling effect of yaw and roll on each other. International Journal of Computer Applications, 47(??):44–48, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume47/number13/7252-0368/.

N:2016:MET

[NVV16] Mohammed Muddasir N, Ravi Kumar V, and Prajwal V. Methods to enhance transformation in near real time ETL. International Journal of Computer Applications, 137(??):20–24, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume137/number5/24272-2016908733/.

Nataliani:2013:IBV

[NW13] Yessica Nataliani and Theophilus Wellem. Implementation of bit-vector algorithm for approximate string matching on rhodopsin protein sequence. International Journal of Computer Applications, 72(??):34–38, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume72/number14/12565-9214/.

Noviada:2015:AIS

[NWS15] Kadek Bondan Noviada, I Made Agus Wirawan, and I Made Gede Sunarya. AMI information system of Ganesha University of Education with decision support system. International Journal of Computer Applications, 131(??):32–37, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume131/number15/23529-2015907469/.

Nirmala:2015:SMR

[NYS15] G. Nirmala, A. Yesubabu, and P. Seetharamaiah. In silico multivariate regressio n analysis and validation studies on selective MMP-13 inhibitors. International Journal of Computer Applications, 130(??):24–37, November 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume130/number6/23213-2015907019/.

Nguyen:2014:RSR

[NYU14] Do Van Nguyen, Koichi Yamada, and Muneyuki Unehara. Rough sets and rule induction in imperfect information systems. International Journal of Computer Applications, 89 (??):1–8, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number5/15495-4286/.

Omri:2012:UNL

[OA12] Bessem Omri and Kaicar Ammous. Using non linear average model of NPC inverter in DTC control of synchronous machines. International Journal of Computer Applications, 59 (??):11–16, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number4/9534-3968/.

Ojeyinka:2015:PAD

[OA15] Taiwo O. Ojeyinka and Olusola Olajide Ajayi. Performance analysis of dual core, Core 2 Duo and Core i3 Intel processor. International Journal of Computer Applications, 120(??):1–7, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number10/21260-8736/.

Omri:2013:NMB

[OAA13] Bessem Omri, Kaicar Ammous, and Anis Ammous. New method for balancing capacitors voltages in NPC inverter without DC-link voltages sensors. International Journal of Computer Applications, 65(??):17–24, March 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume65/number22/11216-6370/.

O:2014:TMR

[OAA14] Akinwunmi A O, Olajubu E A, and Aderounmu G A. A trustworhty model for reliable cloud service discovery. International Journal of Computer Applications, 87(??):23– 30, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number16/15293-3962/.

Obayya:2016:LCI

[OAA16] Marwa I. M. Obayya, Nihal F. F. Areed, and Abdulhadi Omar Abdulhadi. Liver cancer identification using adaptive neuro-fuzzy inference system. International Journal of Computer Applications, 140(??):1–7, April 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume140/number8/24611-2016909402/.

Okai:2024:AFA

[OAA24] George Essah Yaw Okai, William Akotam Agangiba, and Millicent Agangiba. Assessment of farmers’ acceptance of intelligent agriculture system using technology acceptance model. International Journal of Computer Applications, 186(??):16–22, September 2024. CODEN ????

ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume186/number38/assessment-of-farmers-

acceptance-of-intelligent-agriculture-system-using-
technology-acceptance-model/.

Omotosho:2016:RLT

[OAE16] Adebayo Omotosho, Omotanwa Adegbola, and Michael Edobor. Reanalyzing li and tao. (2014): Investigating algorithm recognition on dark irises. International Journal of Computer Applications, 155(??):44–52, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume155/number10/26645-2016912460/.

Okai:2017:RCN

[OAG17] Felix Attuquaye Okai, Ing. Komi Agbesi, and Victor Gbedawo. Resilient campus network to support research and educational network in Ghana [case study of Accra Technical University]. International Journal of Computer Applications, 179(??):19–23, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number4/28724-2017915918/.

Olagunju:2018:SAM

[OAO18] M. Olagunju, A. E. Adeniyi, and T. O. Oladele. Staff attendance monitoring system using fingerprint biometrics. International Journal of Computer Applications, 179(??):8–15, February 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number21/28990-2018916370/.

O:2016:MSM

[OAR16] Awoyelu I O, Awosan O A, and Adagunodo E R. Modeling and simulation of K-means clustering learning object adaptability model for selecting materials in e-learning. International Journal of Computer Applications, 141(??):10–18, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume141/number1/24747-2016908822/.

Obeidat:2019:IRE

[OAS19] Mohammad A. Obeidat and Rasha M. Al-Sarayrah. The impact of renewable energy integration on stability of the Jordanian National Grid. International Journal of Computer Applications, 182(??):26–35, January 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number36/30300-2019918367/.

O:2012:EAS

[OAW12] Akinwunmi A O, Aderounmu G A, and Onifade O F W. Enhanced accounting scheme for grid computing architecture. International Journal of Computer Applications, 37 (??):21–28, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume37/
number9/4637-6687/.

Obeidavi:2023:AIE

[Obe23] Sahereh Obeidavi. Automatic identification of extremely tiny brain hemorrhages in susceptibility weighted images using convolutional neural network. International Journal of Computer Applications, 185(??):43–48, November 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number43/32979-2023923244/.

Orike:2013:CEG

[OC13] Sunny Orike and David W. Corne. Constrained elitist genetic algorithm for economic load dispatch: Case study on Nigerian power system. International Journal of Computer Applications, 76(??):27–33, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume76/number5/13245-0706/.

Odunola:2020:IEE

[OC20] Akinyemi Bodunde Odunola and Egunlayi Olutayo Cyril. An improved energy-efficient prediction-based model for animal tracking in wireless sensor networks. International Journal of Computer Applications, 177(??):15–24, February 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume177/number37/31146-
2020919860/.

Ozturk:2013:VCW

[OCZ13] Muhammed Maruf Ozturk, Unal Cavuhohlu, and Ahmet Zengin. A voice controlled Web based FTP tool. International Journal of Computer Applications, 64(??):22– 27, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number10/10671-5456/.

Olabisi:2013:DSP

[OD13] Samuel Babafemi Olabisi and Aderibigbe Adesoji David. Digital smart pen: a portable media with endless benefits. International Journal of Computer Applications, 74 (??):28–30, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number14/12954-0041/.

Odeh:2014:NER

[Ode14] Jehad Q. Odeh. New and efficient recursive-based string matching algorithm (RSMA-FLFC). International Journal of Computer Applications, 86(??):1–6, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume86/number15/15058-3190/.

Oladokun:2014:UMF

[OE14] V. O. Oladokun and C. G. Emmanuel. Urban market fire disasters management in Nigeria: a damage minimization based fuzzy logic model approach. International Journal of Computer Applications, 106(??):1–6, November 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume106/number17/18609-9525/.

Okon:2018:IOB

[OEA18] E. Uko Okon, B. O. Eke, and P. O. Asagba. An improved online book recommender system using collaborative filtering algorithm. International Journal of Computer Applications, 179(??):41–48, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number46/29486-2018917193/.

Osman:2016:FTS

[OESKF16] Mohamed S. Osman, Mahmoud M. El-Sherbiny, Hamdeen A. Khalifa, and Hanan H. Farag. A fuzzy technique for solving rough interval multiobjective transportation problem. International Journal of Computer Applications, 147(??):49–57, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume147/number10/25693-2016911204/.

Olowookere:2015:UDL

[OFE15] Toluwase A. Olowookere, Ayodeji I. Fasiku, and Ifeanyi C. Emeto. UPH digital library miner: a topic modelling-based software application for mining document collections of a digital library. International Journal of Computer Applications, 132(??):1–8, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number13/23651-2015907559/.

Ogunlere:2017:ACA

[OFI17] S. O. Ogunlere, S. A. Fagbohun, and O. K. Iwala. Animal care automated system. International Journal of Computer Applications, 177(??):23–30, November 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number2/28600-2017915674/.

Osunade:2016:ELS

[OG16] O. Osunade and I. A. Ganiyu. Enhancing the least significant bit (LSB) algorithm for steganography. International Journal of Computer Applications, 149(??): 1–8, September 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume149/number3/25974-2016911363/.

Ouchitachen:2015:MOW

[OHI15] Hicham Ouchitachen, Abdellatif Hair, and Najlae Idrissi. Multi-objective weighted clustering algorithm minimizing jointly the costs of mission and communication in wireless sensor network. International Journal of Computer Applications, 127(??):24–31, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number15/22807-2015906673/.

Opoku:2013:RSA

[OHT13] Daniel Opoku, Abdollah Homaifar, and Edward Tunstel. The A-r-Star (Ar) Pathfinder. International Journal of Computer Applications, 67(??):32–44, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume67/number8/11417-6753/.

Otman:2011:CSA

[OJ11] Abdoun Otman and Abouchabaka Jaafar. A comparative study of adaptive crossover operators for genetic algorithms to resolve the traveling salesman problem. International Journal of Computer Applications, 31(??):49– 57, October 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume31/
number11/3945-5587/.

Ojha:2011:NSO

[Ojh11] Muneendra Ojha. A novel service oriented model for query identification and solution development using Semantic Web and multi agent system. International Journal of Computer Applications, 35(??):12–18, December 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume35/number9/4428-6165/.

O:2011:DRM

[OJO11] Alamu F O, Emuoyibofarhe O J, and Oke A O. Developing a robust multimedia picture archiving and communication system (PACS). International Journal of Computer Applications, 34(??):12–25, November 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume34/number4/4085-5568/.

Olama:2011:DIN

[OJR11] Ehsan Khadem Olama and Hooshang Jazayeri-Rad. Design and implementation of a noise tolerant polynomial nonlinear ARX model using the averaging wavelet method. International Journal of Computer Applications, 35(??):1– 5, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume35/
number1/4362-6012/.

Otieno:2014:AMP

[OK14] Erick O. Otieno and Andrew M. Kahonge. Adoption of mobile payments in Kenyan businesses: a case study of small and medium enterprises (SME) in Kenya. International Journal of Computer Applications, 107(??): 5–12, December 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume107/number7/18761-0041/.

Okpor:2017:EEV

[Okp17] Margaret Dumebi Okpor. An enhanced electronic voting system (EnEVoS) with hybrid authentication technique. International Journal of Computer Applications, 169(??):1–5, July 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume169/number1/27946-2017912060/.

Omur:2014:NPG

[OKS14] Nese Omur, Sibel Koparal, and Cemile Duygu Sener. A new perspective to the generalization of sequences of torder. International Journal of Computer Applications, 86 (??):29–33, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume86/
number6/14991-2633/.

Otasowie:2018:TDD

[OKS18] Owolafe Otasowie, Alese Bonifacem K, and Adewale Olumide S. Towards detecting deception using K-nearest neighbour model. International Journal of Computer Applications, 181(??):14–18, October 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number20/30000-2018917846/.

Outair:2014:TAE

[OLT14] Anas Outair, Abdelouahid Lyhyaoui, and Mariam Tanana. Towards an automatic evaluation of UML class diagrams by graph transformation. International Journal of Computer Applications, 95(??):36–41, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number21/16721-7063/.

Oluwaseun:2014:LPS

[Olu14] Longe Idowu Oluwaseun. Logic of perfect secrecy in aroko computational cryptography. International Journal of Computer Applications, 95(??):32–42, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume95/number22/16729-6987/.

Oujaoura:2014:RIP

[OMF^{+}14] M. Oujaoura, B. Minaoui, M. Fakir, R. El Ayachi, and O. Bencharef. Recognition of isolated printed tifinagh characters. International Journal of Computer Applications, 85 (??):1–13, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number1/14802-3005/.

Obaida:2011:RED

[OMH11] Mohammad Abu Obaida, Md. Sanaullah Miah, and Md. Abu Horaira. Random early discard (RED-AQM) performance analysis in terms of TCP variants and network parameters: Instability in high-bandwidth-delay network. International Journal of Computer Applications, 27(??):40– 44, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume27/
number8/3317-4556/.

Ochieng:2014:SSE

[OMM14] Peter Ochieng, Waweru Mwangi, and Solomon Mwanjele Mwgha. Software size estimation in incremental software development based on improved pairwise comparison matrices. International Journal of Computer Applications, 93 (??):29–39, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number5/16213-5519/.

Omar:2013:AID

[ONJ13] Salima Omar, Asri Ngadi, and Hamid H. Jebur. An adaptive intrusion detection model based on machine learning techniques. International Journal of Computer Applications, 70(??):1–5, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number7/11971-6640/.

Ocharo:2024:APM

[ONK24] Dominic Ocharo, H. Chege Nganga, and Stephen Kiambi. Automatic PPE monitoring system for construction workers using YOLO algorithm based on deep reinforcement learning. International Journal of Computer Applications, 186(??):10–15, November 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number48/automatic-ppe-monitoring-system-
for-construction-workers-using-yolo-algorithm-based-
on-deep-reinforcement-learning/.

Omoanatse:2014:DIM

[OO14] Alenoghena Caroline Omoanatse and Emagbetere Joy Okumo. Design and implementation of a map building model for base stations placement. International Journal of Computer Applications, 90(??):31–35, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume90/number9/15605-4149/.

Ojekudo:2018:MDU

[OO18] Nathaniel A. Ojekudo and Eva N. Obinna. Malaria diagnosis using swarm intelligence. International Journal of Computer Applications, 181(??):24–28, August 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number9/29800-2018917602/.

Oladipo:2021:TSS

[OO21] Francisca O. Oladipo and Abdulaziz Baba-Ali Ohiani. Text summarization system: an extractive approach using hierarchical text clustering. International Journal of Computer Applications, 174(??):15–19, March 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number23/31813-2021921015/.

Otieng:2018:ACD

[OOC18] Okiyai Otieng, George Okeyo, and Wilson Cheruiyot. Approaches to curbing data breaches in Internet banking based on cloud computing. International Journal of Computer Applications, 181(??):11–16, July 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number2/29687-2018917411/.

Ogini:2018:IBM

[OOE18] Nicholas Oluwole Ogini, Noah Oghenefego Ogwara, and Obeten Obi Ekabua. Intelligent based multi-agent approach for university timetable scheduling system. International Journal of Computer Applications, 182(??):10–21, July 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number1/29724-2018917432/

.

Olanrewaju:2017:FLC

[OOJ17] O. M. Olanrewaju, A. A. Obiniyi, and S. B. Junaidu. Fuzzy logic concept for safety driven vehicle-pedestrian traffic interaction. International Journal of Computer Applications, 167(??):24–28, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number1/27736-2017914058/.

Oluwadare:2017:TST

[OON17] Samuel A. Oluwadare, Bosede A. Ogunsanmi, and John C. Nwaiwu. Towards solving travelling salesperson problem using hybrid of genetic algorithm and Lin–Kernighan algorithm: a comparative evaluation with neural network model. International Journal of Computer Applications, 179(??):11–19, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number7/28747-2017915953/.

Ojelabi:2019:CDC

[OOO19] Ayobami I. Ojelabi, Oluwabusayo I. Omotosho, and Olajide A. Oladejo. Classification and detection of citrus disease using feature extraction and support vector machine (SVM). International Journal of Computer Applications, 177(??):17–25, November 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number17/30990-2019919582/.

Ojha:2018:OAR

[OP18] Devyani Ojha and Pragya Pandey. Optimizing association rule using genetic algorithm and data sampling approach. International Journal of Computer Applications, 179(??):15–19, January 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number11/28843-2018916083/.

Orike:2014:NAE

[Ori14] Sunny Orike. Novel adaptive evolutionary computation approaches to the dynamic economic load dispatch problems with non-smooth fuel cost function. International Journal of Computer Applications, 103(??):11–18, October 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume103/number9/18101-9166/

.

Omran:2014:DIM

[OS14] Safaa S. Omran and Maryam Abdulmunem Salih. Design and implementation of multi-model biometric identification system. International Journal of Computer Applications, 99(??):14–21, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume99/number15/17448-8255/.

Oleksiy:2016:SVD

[OS16] Marygin Oleksiy and Xing Shen. Synthesis of vibration damping control systems based on the identified model of a mechanical construction at broadband disturbance influence. International Journal of Computer Applications, 142(??):18–23, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number2/24868-2016909692/.

Osamy:2014:SSR

[OSA14] Walid Osamy, Ahmed Salim, and Ahmed Aziz. Sparse signals reconstruction via adaptive iterative greedy algorithm. International Journal of Computer Applications, 90 (??):5–11, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume90/
number17/15810-4715/.

Omar:2014:IPI

[OSS14] Azizah Che Omar, Norshuhada Shiratuddin, and Siti Mahfuzah Sarif. Impulse purchase in iTV advertising: a conceptual model of gap analysis. International Journal of Computer Applications, 91(??):20–26, April 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume91/number11/15925-5178/.

Otair:2013:EEC

[Ota13] Mohammed Otair. Effectively and efficiency consideration for spatial database. International Journal of Computer Applications, 74(??):32–37, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume74/number18/12987-0232/.

Ouali:2012:PES

[Oua12] Mohammed Ouali. Performance evaluation of stereo matching algorithms in the lack of visual features. International Journal of Computer Applications, 53(??):7–11, September 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume53/number5/8415-
0636/.

Ouda:2013:FCS

[Oud13] Osama Ouda. Fuzzy commitment scheme for masked iris codes. International Journal of Computer Applications, 79 (??):1–5, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number11/13782-1915/.

Omidvar:2013:PME

[OV13] Mohammad Nabi Omidvar and Reza Vaziri. Provide a method for evaluation of software architecture using ontology. International Journal of Computer Applications, 64 (??):1–7, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number16/10715-1307/.

Priya:2012:SNN

[PA12] R. Priya and P. Aruna. SVM and neural network based diagnosis of diabetic retinopathy. International Journal of Computer Applications, 41(??):6–12, March 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume41/number1/5503-7503/.

P:2013:AEC

[PA13] Brindha P and Senthilkumar A. Area efficient counting bloom filter (A-CBF) design for NIDS. International Journal of Computer Applications, 70(??):17–21, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume70/number4/11949-7767/.

P:2014:EMS

[PA14] Binu V P and Sreekumar A. Efficient multi secret sharing with generalized access structures. International Journal of Computer Applications, 90(??):1–6, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume90/number12/15769-4446/.

Parveen:2017:BCA

[PA17] Rihana Parveen and Mohd Abdulla. Broadband compact antenna for S-band and C-band application. International Journal of Computer Applications, 160(??): 5–9, February 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume160/number8/27091-2017912979/.

Prasad:2019:ATD

[PA19] M. V. V. Siva Prasad and P. S. Avadhani. Attack tree design and analysis of offshore oil and gas process complex SCADA system. International Journal of Computer Applications, 181(??):12–18, February 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number41/30334-2019918440/.

Pratama:2023:OMV

[PA23a] Alfansyah Nori Pratama and Farida Ardiani. Optimization of model-view-ViewModel (MVVM) architecture pattern and RESTfull API on Android-based e-learning application. International Journal of Computer Applications, 185(??):4–11, November 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number45/32990-2023923261/.

Putera:2023:IRB

[PA23b] Cornellius Agova Madu Putera and Joko Aryanto. Implementation of rule base-fuzzy logic on Android based plant watering system with Internet of Things technology. International Journal of Computer Applications, 185(??):31–35, November 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number39/32948-2023923198/.

Patel:2024:EPP

[PA24] Twinkle Patel and Kiran Amin. Enhancing privacy preservation: Multi-attribute protection with P -sensitive Kanonymity. International Journal of Computer Applications, 186(??):1–8, March 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number14/enhancing-privacy-preservation-
multi-attribute-protection-with-p-sensitive-k-anonymity/

.

Proloy:2017:CAV

[PAH17] Md. Razaul Hoq Proloy, Syed Mushfiq Ahmed, and Md. Asif Hossain. Comparative analysis of various wireless digital modulation techniques with different channel coding schemes under AWGN channel. International Journal of Computer Applications, 161(??):30–34, March 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume161/number3/27130-2017913139/.

Paliwal:2016:ARB

[PAK16] Shefali Paliwal, Darpan Anand, and Sartaj Khan. Application of rule based fuzzy inference system in prediction of Internet phishing. International Journal of Computer Applications, 148(??):30–37, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume148/number14/25843-2016911334/.

Paramasiven:2011:RWA

[Par11] Appavoo Paramasiven. Rendering wormhole attacks trivial using the scalability features of a geocasting protocol. International Journal of Computer Applications, 19 (??):24–28, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume19/
number5/2357-3083/.

Parmar:2014:LSM

[Par14a] K. P. Singh Parmar. LFC system of multi-area interconnected power systems using TVAC-PSO based controller. International Journal of Computer Applications, 88(??):13– 19, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume88/
number8/15372-3923/.

Parmar:2014:PBP

[Par14b] K. P. Singh Parmar. PSO based PI controller for the LFC system of an interconnected power system. International Journal of Computer Applications, 88(??):20– 25, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume88/
number7/15364-3856/.

Patel:2014:MOJ

[Pat14] Pritibahen Sumanbhai Patel. Multi-objective job scheduler using genetic algorithm in grid computing. International Journal of Computer Applications, 92(??):34– 43, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume92/
number14/16079-5312/.

Priya:2012:RRI

[PB12] G. Nirmala Priya and R. S. D. Wahida Banu. A robust rotation invariant multiview face detection in erratic illumination condition. International Journal of Computer Applications, 57(??):46–51, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume57/number20/9234-3802/.

Pardeshi:2013:HBA

[PB13] Hrishikesh Pardeshi and Chandranath Bhattacharyya. Heuristic-based approach to detect global touch gestures performed on touch devices. International Journal of Computer Applications, 82(??):37–43, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume82/number8/14139-2278/.

Pokle:2014:NMI

[PB14] Pravin B. Pokle and Narendra G. Bawane. Novel method of image compression using angular domain concept to achieve high compression rate. International Journal of Computer Applications, 108(??):23–27, December 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume108/number6/18915-0234/.

Phulpagar:2015:SIC

[PB15] B. D. Phulpagar and R. S. Bichkar. Segmentation of images containing multiple intensity levels using genetic algorithms. International Journal of Computer Applications, 126(??):29–39, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume126/number9/22581-2015906186/.

Pathak:2016:CAA

[PB16] Neha Pathak and Abhinav Bhargava. A comparative analysis of asymmetrical U-slot and substrate integrated waveguide fed microstrip patch antenna. International Journal of Computer Applications, 138(??):33–36, March 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume138/number11/24426-2016909047/.

Pavlic:2019:FBC

[PB19] Jani Pavlic and Jernej Burkeljca. FFmpeg based coding efficiency comparison of H.264/AVC, H.265/HEVC and VP9 video coding standards for video hosting websites. International Journal of Computer Applications, 182(??):1–8, January 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume182/number37/
30303-2019918381/.

Papakyriakou:2023:HPL

[PB23] Dimitrios Papakyriakou and Ioannis S. Barbounakis. High Performance Linpack (HPL) benchmark on Raspberry Pi 4B (8GB) Beowulf cluster. International Journal of Computer Applications, 185(??):11–19, July 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number25/32847-2023923005/.

Pande:2012:SEM

[PBG12] Shaifali Pande, Vivek Singh Bhadouria, and Dibyendu Ghoshal. A study on edge marking scheme of various standard edge detectors. International Journal of Computer Applications, 44(??):33–37, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume44/number9/6294-8491/.

Pramanik:2015:CCM

[PBG15] Surapati Pramanik, Durga Banerjee, and B. C. Giri. Chance constrained multi-level linear programming problem. International Journal of Computer Applications, 120(??):1–6, June 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume120/number18/
21324-4275/.

Patre:2014:DMC

[PBK14] Samiksha Patre, Nilesh Bodane, and Vinay Keswani. Design of 0.13μ m CMOS two stage low noise amplifier. International Journal of Computer Applications, 100(??):17–20, August 2014. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume100/number15/
17600-8190/.

PereiraDaSilva:2022:CTD

[PBVD22] João Vagner Pereira Da Silva, Sandra Kalil Bussadori, Maria Lucia Z. Varellis, and Alessandro M. Deana. Classification of tooth decay by computer-aided diagnostics using laser speckle image analysis. International Journal of Computer Applications, 184(??):12–17, April 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number10/32361-2022922008/.

Paramasivam:2011:DLF

[PC11a] B. Paramasivam and I. A. Chidambaram. Design of loadfrequency controller using artificial bee colony algorithm for an interconnected power system coordinated with UPFC and RFB. International Journal of Computer Applications, 36 (??):25–36, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume36/
number5/4488-6318/.

Perwej:2011:MRH

[PC11b] Yusuf Perwej and Ashish Chaturvedi. Machine recognition of hand written characters using neural networks. International Journal of Computer Applications, 14(??): 6–9, January 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume14/
number2/1819-2380/.

Perera:2015:NEB

[PC15a] T. A. S. Achala Perera and John Collins. A novel eigenface based species recognition system. International Journal of Computer Applications, 115(??):19–23, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume115/number20/20268-2675/.

Pongle:2015:RTI

[PC15b] Pavan Pongle and Gurunath Chavan. Real time intrusion and wormhole attack detection in Internet of Things. International Journal of Computer Applications, 121(??):1–9, July 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume121/number9/21565-4589/.

Patra:2012:TAS

[PCC12] Rusha Patra, Chandan Chakraborty, and Jyotirmoy Chatterjee. Textural analysis of spinous layer for grading oral submucous fibrosis. International Journal of Computer Applications, 48(??):33–37, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume48/number22/7513-0563/.

Pant:2014:CDJ

[PCD14] Suman Pant, Yashwant S Chauhan, and Priti Dimri. Complex dynamics of Jungck Ishikawa iterates for hyperbolic cosine function. International Journal of Computer Applications, 101(??):14–23, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number4/17674-8506/.

Papoutsidakis:2017:BGC

[PCKD17] M. Papoutsidakis, A. Chatzopoulos, K. Kalovrektis, and C. Drosos. A brief guide for the continuously evolving controller Raspberry PI Mod.B. International Journal of Computer Applications, 176(??):30–33, October 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number8/28577-2017915651/.

Padmanabhan:2012:EPA

[PCR12] S. Padmanabhan, M. Chandrasekaran, and V. Srinivasa Raman. Evaluation of the performance of ant colony optimization over particle swarm optimization. International Journal of Computer Applications, 38(??):12–18, January 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume38/number6/4691-
6826/.

Papoutsidakis:2018:AUP

[PCS18] M. Papoutsidakis, A. Chatzopoulos, and E. Symeonaki. Automation upgrade via PLC applications for big harbor delivery facilities. International Journal of Computer Applications, 179(??):39–44, March 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number30/29173-2018916690/.

Patil:2010:FAC

[PCW10] N. J. Patil, R. H. Chile, and L. M. Waghmare. Fuzzy adaptive controllers for speed control of PMSM drive. International Journal of Computer Applications, 1(??):94–101, February 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume1/number11/233-387/.

Pandharipande:2014:CSP

[PD14] Shekhar L. Pandharipande and Aasheesh Kumar Dixit. Comparative study of performance of chemical reaction optimization with genetic algorithm (GA). International Journal of Computer Applications, 107(??):

1–8, December 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume107/number8/18768-0069/.

PedrozaFerreira:2016:AGQ

[PDG16] Matheus Pedroza Ferreira, Glêndara A. De S.Martins, and Warley Gramacho Da Silva. Analysis of guava quality by image processing. International Journal of Computer Applications, 156(??):30–36, December 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume156/number3/26691-2016912404/.

Praveen:2012:ITS

[PDK12] Bagadi K. Praveen, Susmita Das, and Sridhar K. Image transmission over space time coded MIMO-OFDM system with punctured turbo codes. International Journal of Computer Applications, 51(??):1–6, August 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume51/number15/8115-1734/.

Pandharipande:2014:GAC

[PDK14] Shekhar L. Pandharipande, Aarti R. Deshmukh, and Rohit P. Kalnake. Genetic algorithm for constrained optimization with stepwise approach in search interval selection of variables. International Journal of Computer Applications, 87(??):43–52, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume87/number11/15256-4017/.

Prasad:2011:IET

[PDR11] M Seetharama Prasad, T Divakar, and L S S Reddy. Improved entropic threshold based on GLSC histogram with varying similarity measure. International Journal of Computer Applications, 23(??):25–32, June 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume23/number1/2852-3661/.

Petkovic:2022:IPF

[Pet22] Dušan Petković. Identifying possible financial frauds using SQL row pattern recognition. International Journal of Computer Applications, 184(??):31–34, November 2022. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number35/32542-2022922446/.

Petkovic:2023:CTS

[Pet23] Dušan Petković. Comparison of techniques for storing JSON data in relational form. International Journal of Computer Applications, 185(??):9–12, July 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number22/32822-2023922962/.

PereiraDaSilva:2020:MOP

[PF20] Tcharles Pereira Da Silva and Kleinner Farias. Mind overflow: a process proposal for decomposing monolithic applications in microservices. International Journal of Computer Applications, 176(??):1–9, April 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number18/31298-2020920125/.

Paul:2012:PHA

[PFS12] Michael Paul, Andrew Finch, and Eiichiro Sumita. Predicting human assessment of machine translation quality by combining automatic evaluation metrics using binary classifiers. International Journal of Computer Applications, 59 (??):1–7, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number10/9581-4062/.

Patil:2011:CSM

[PG11] Netra Patil and Arpita Gopal. Comparative study of mechanisms for Web service discovery based on centralized approach focusing on UDDI. International Journal of Computer Applications, 14(??):28–31, January 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume14/number1/1809-2328/.

Pal:2013:RAS

[PG13a] Dipankar Pal and Nabin Ghoshal. A robust audio steganographic scheme in time domain (RASSTD). International Journal of Computer Applications, 80(??):1– 8, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume80/
number15/13934-1803/.

Pradhan:2013:ETA

[PG13b] Mohan P Pradhan and M K Ghose. Efficient technique for automatic extraction and identification of text from an image using friend pattern chain and Euclidean distance. International Journal of Computer Applications, 71 (??):30–41, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number10/12396-8774/.

Punitha:2013:TBE

[PG13c] A. Punitha and M. Kalaiselvi Geetha. Texture based emotion recognition from facial expressions using support vector machine. International Journal of Computer Applications, 80(??):1–5, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume80/number5/13854-1715/.

Pagare:2015:CRF

[PG15a] Sanjay Pagare and Surendra Gupta. Comparison of RBFNN, FBNN and KNN algorithms for face recognition using PCA and rectangular feature. International Journal of Computer Applications, 115(??):42–48, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number9/20183-2403/.

Prabakar:2015:FJS

[PG15b] D. Prabakar and C. Gowthami. FHQA JDRA schemes for reliability and energy efficiency — soft-DF-based ORassisted SC-FDMA. International Journal of Computer Applications, 118(??):21–26, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number20/20862-3573/.

Pathirana:2023:CEI

[PGPM23] H. P. A. I. Pathirana, V. B. Godagama, H. K. S. H. Premadas, and A. B. Mailewa. Cost-effective investment to monitor the network infrastructure of small and mediumscale enterprises. International Journal of Computer Applications, 185(??):33–38, July 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number23/32834-2023922983/.

Pure:2013:NIF

[PGS13] Anjali A. Pure, Neelesh Gupta, and Meha Shrivastava. A new image fusion method based on integration of wavelet and fast discrete curvelet transform. International Journal of Computer Applications, 69(??):31–35, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume69/number18/12073-8217/.

Perera:2023:SGM

[PH23] Waat Perera and B. Hettige. Subtitle generating media player using Mozilla DeepSpeech model. International Journal of Computer Applications, 185(??):34–42, August 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume185/number28/32871-
2023923033/.

Palaniappan:2015:HAS

[PHK^{+}15] Satish Palaniappan, Naveen Hariharan, Naren T Kesh, Vidhyalakshimi S, and Angel Deborah S. Home automation systems — a study. International Journal of Computer Applications, 116(??):11–18, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number11/20379-2601/.

Paul:2011:EAA

[PIB11] Bijan Paul, Md. Ibrahim, and Md. Abu Naser Bikas. Experimental analysis of AODV & DSR over TCP & CBR connections with varying speed and node density in VANET. International Journal of Computer Applications, 24(??): 30–37, June 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume24/
number4/2937-3897/.

Pilipczuk:2013:IVC

[Pil13] Olga Pilipczuk. Integration of the visible color spectrum and the rough sets theory in the process of expert knowledge extraction. International Journal of Computer Applications, 68(??):24–30, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number19/11689-7387/.

Prakash:2012:NIP

[PJ12] Aaditya Prakash and Rajan K Jha. New interface protocol to connect multiple bank networks from a single outlet. International Journal of Computer Applications, 55 (??):1–9, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume55/
number12/8804-3034/.

Pande:2012:LHR

[PK12a] Sharon Pande and Preeti Khanna. Leveraging human resource information systems: Alignment of business with technology. International Journal of Computer Applications, 56 (??):22–29, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number3/8872-2851/.

Patnaik:2012:NDB

[PK12b] Ansuman Patnaik and Diptimayee Konhar. A novel dual band circular microstrip patch antenna for wireless applications. International Journal of Computer Applications, 40 (??):1–2, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume40/
number11/5005-7322/.

Paralkar:2015:CSP

[PK15a] Suryakant R. Paralkar and U. V. Kulkarni. A comprehensive survey on pattern classifier’s security evaluation under attack. International Journal of Computer Applications, 115(??):23–28, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number21/20276-2701/.

Patil:2015:BMC

[PK15b] Ruchika Patil and Amreen Khan. Bisecting K-means for clustering Web log data. International Journal of Computer Applications, 116(??):36–41, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number19/20448-2799/.

Pradeebha:2016:CLD

[PK16] R. Pradeebha and S. Karpagavalli. Classification of lung disease using local and global descriptors. International Journal of Computer Applications, 135(??): 19–23, February 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume135/number6/24055-2016908368/.

Patil:2019:AUC

[PK19] Pooja Patil and Subhash S. Kulkarni. Analytics of unconstrained convex problems in sparse and redundant representations: a case study on image processing applications. International Journal of Computer Applications, 178(??):35–42, July 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number34/
30762-2019919223/.

Panigrahi:2013:TSF

[PKB13] Sibarama Panigrahi, Yasobanta Karali, and H. S. Behera. Time series forecasting using evolutionary neural network. International Journal of Computer Applications, 75(??): 13–17, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number10/13146-0553/.

Papoutsidakis:2017:IDA

[PKDS17] M. Papoutsidakis, K. Kalovrektis, C. Drosos, and G. Stamoulis. Intelligent design and algorithms to control a stereoscopic camera on a robotic workspace. International Journal of Computer Applications, 167(??):32–35, June 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume167/number12/27824-
2017914495/.

Prakash:2014:SCC

[PKJ14] A. Praveen Prakash, M. P. Kannan, and J. Esther Jerlin. A study on the causes for child trafficking using indeterministic trapezoidal fuzzy cognitive maps. International Journal of Computer Applications, 104(??): 36–39, October 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume104/number8/18226-9393/.

Prakash:2015:CSC

[PKJ15] Khairnar Vinayak Prakash, Abhijeet Kumar, and Prerana Jain. Circumventing short channel effects in FETs: Review. International Journal of Computer Applications, 117(??):24– 30, May 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume117/number17/
20648-3407/.

Papakyriakou:2018:BRP

[PKK18] Dimitrios Papakyriakou, Dimitra Kottou, and Ioannis Kostouros. Benchmarking Raspberry Pi 2 Beowulf cluster. International Journal of Computer Applications, 179(??):21–27, April 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume179/number32/
29203-2018916728/.

P:2013:SBS

[PKP13] Satyanarayana Murty P, Sekhar Dileep Kumar, and Rajesh Kumar P. A semi blind self reference image watermarking in discrete cosine transform using singular value decomposition. International Journal of Computer Applications, 62(??):28–36, January 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume62/number13/10141-4948/.

Priyanka:2013:PEW

[PM13] Priyanka and Jyoteesh Malhotra. Performance evaluation of WiMAX networks through large scale fading channel conditions. International Journal of Computer Applications, 70 (??):22–25, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number2/11935-7722/.

Pivezhandi:2015:SBN

[PM15] Mohammad Pivezhandi and Baback Mazloom-Nezhad Maybodi. Statistical based neural network in human activity recognition. International Journal of Computer Applications, 124(??):1–5, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number12/22153-2015905678/.

Pereira:2021:ADS

[PM21] Silas Santiago L. Pereira and Jose E. B. Maia. Anomaly detection in surveillance video of natural environment. International Journal of Computer Applications, 183(??):1–7, May 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number1/31889-2021921288/

.

Poyil:2013:SCC

[PMA13] Azeemsha Thacham Poyil, Nasimudeen K M, and Sultan Aljahdali. Significance of Cohen’s class for time frequency analysis of signals. International Journal of Computer Applications, 72(??):1–8, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume72/number12/12543-8960/.

Podder:2020:RST

[PMBP20] Prajoy Podder, M. Rubaiyat Hossain Mondal, Subrato Bharati, and Pinto Kumar Paul. Review on the security threats of Internet of Things. International Journal of Computer Applications, 176(??):37–45, July 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number41/31476-2020920548/.

Pudaruth:2013:HRP

[PMD13] Sameerchand Pudaruth, Nicolas Medard, and Zaynah Bibi Dookhun. Horse racing prediction at the Champ De Mars using a weighted probabilistic approach. International Journal of Computer Applications, 72(??):37–42, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume72/number5/12493-9048/.

Prakash:2010:ACM

[PMH10] Pattan Prakash, V. D. Mytri, and P. S. Hiremath. Active contour multigrid model for segmentation and automatic quantification of material phases of cast iron. International Journal of Computer Applications, 9(??):32–37, November 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume9/number4/1372-1849/

.

Parmar:2012:LIP

[PMK12] K. P. Singh Parmar, S. Majhi, and D. P. Kothari. LFC of an interconnected power system with thyristor controlled phase shifter in the tie line. International Journal of Computer Applications, 41(??):27–30, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume41/number9/5571-7665/.

Pramanik:2018:TSB

[PMM18] Surapati Pramanik, Indrani Maiti, and Tarni Mandal. A Taylor series based fuzzy mathematical approach for multi objective linear fractional programming problem with fuzzy parameters. International Journal of Computer Applications, 180(??):22–29, May 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number45/29447-2018917154/.

Patel:2012:FRL

[PMP12] Rahul Patel, Hiren Mewada, and Suprava Patnaik. Fast and regularization less active contour. International Journal of Computer Applications, 47(??):26–31, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume47/number6/7194-9958/.

Purwanto:2014:ESM

[PMP14] Purwanto, Danny Manongga, and M. A. Ineke Pakereng. Efficiency of small- and medium-sized tofu enterprises (SME) in Salatiga using Data Envelopment Analysis (DEA). International Journal of Computer Applications, 91(??): 44–50, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume91/
number12/15936-5252/.

Poornima:2017:MRQ

[PMU17] Aradhyamath Poornima, Naghabhushana N M, and Rohitha Ujjinimatad. Matrix representation of quantum gates. International Journal of Computer Applications, 159(??):1–6, February 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume159/number8/27018-2017913011/.

Prakash:2010:SST

[PN10] J. Thomas Joseph Prakash and L. Ruby Nirmala. Synthesis, spectral and thermal properties of bis thiourea zinc acetate (BTZA) nonlinear optical single crystal. International Journal of Computer Applications, 8(??):7–11, October 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume8/number2/1188-
1660/.

P:2011:EUG

[PN11] Leela Rani P and Revathi N. Enhancing the utility of generalization for privacy preserving re-publication of dynamic datasets. International Journal of Computer Applications, 13(??):42–49, January 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume13/number6/1781-2456/.

Padmashree:2012:CAQ

[PN12] S. Padmashree and Rohini Nagapadma. Comparative analysis of quad-tree partitioned iterated function systems for medical images. International Journal of Computer Applications, 59(??):25–27, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume59/number14/9619-4261/.

Pandey:2014:DMS

[PN14] Chanchal Pandey and Sandeep Nemad. Distinctive methods for speech enhancement using Kalman filtering. International Journal of Computer Applications, 105(??): 1–5, November 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume105/number5/18370-9511/.

Phimthong-Ngam:2016:ENO

[PNIS16] Yutthana Phimthong-Ngam, Kannakorn Intharakham, and Kesorn Suwanprasert. Evaluation of nitric oxide in lacunar stroke and young healthy during cerebrovascular reactivity by support vector machine. International Journal of Computer Applications, 146(??):28–35, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number9/25428-2016910891/.

Poonia:2022:PAM

[Poo22] Praveen Kumar Poonia. Performance assessment of a multistate standby series system using copula distribution and catastrophic failure. International Journal of Computer Applications, 184(??):1–7, December 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number39/32568-2022922478/.

Pourvali:2013:DNN

[Pou13] Ghahreman Pourvali. On design of a novel nano metric parity preserving reversible random access memory. International Journal of Computer Applications, 75(??): 1–7, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number4/13096-0384/.

Parvathi:2010:NHS

[PP10] R. Parvathi and S. Palaniammal. A novel hybrid spatial association rule mining algorithm for neuro imaging. International Journal of Computer Applications, 8(??):32–37, October 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume8/number9/1233-
1616/.

Pandey:2012:MMC

[PP12a] Anjana Pandey and K. R. Pardasani. A model for mining course information using vague association rule. International Journal of Computer Applications, 58(??):1– 5, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number20/9395-7825/.

Preetha:2012:WBR

[PP12b] S. Preetha and M. Punithavalli. Web based reachability testing for hetergeneous application using Nash equilibrium. International Journal of Computer Applications, 45 (??):25–30, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume45/
number2/6752-9004/.

P:2013:MSB

[PP13a] Mani P and Kamalakkannan P. Mitigating selfish behavior in mobile ad hoc networks: a survey. International Journal of Computer Applications, 73(??):1–7, July 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume73/number22/13031-8846/.

P:2013:MWD

[PP13b] Satyanarayana Murty P and Rajesh Kumar P. A method for watermarking in digital videos by using hybrid transforms and edge detection. International Journal of Computer Applications, 71(??):16–23, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number13/12418-8918/.

Patil:2013:SPM

[PP13c] Pooja K. Patil and P. M. Pawar. Security of PHR model on public cloud using MultiAuthority and key policy attribute based encryption. International Journal of Computer Applications, 84(??):46–52, December 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume84/number12/14631-2999/.

Patil:2013:SRV

[PP13d] Sandip S. Patil and Harshal S. Patil. Study and review of various image texture classification methods. International Journal of Computer Applications, 75(??):33– 38, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number16/13197-0897/.

Patel:2015:NII

[PP15] Pritam Kumar Patel and M Premchand. Numerical investigation of the influence of water depth on ship resistance. International Journal of Computer Applications, 116(??):10– 17, April 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume116/number17/
20427-2750/.

Prasad:2016:DES

[PP16] V. R. Prasad and G. M. Phade. Design of electronic stethoscope and heart rate monitor for remote area application. International Journal of Computer Applications, 137(??):33–36, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume137/number10/24311-2016908881/.

Pandey:2017:PLU

[PP17a] Mrinal Pandey and Rajiv Pandey. Provenance linking using bundles in OWL ontology. International Journal of Computer Applications, 164(??):5–9, April 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume164/number11/27525-2017913721/.

Patil:2017:AOD

[PP17b] Divya C. Patil and Pravin S. Patil. Abandoned object detection using temporal consistency modeling. International Journal of Computer Applications, 164(??):15–21, April 2017. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume164/number10/
27519-2017913707/.

Patil:2015:EED

[PPB15] Hemlata Patil, A. J. Patil, and S. G. Bhirud. Energy efficient design for external cooperative sensing in cognitive radio. International Journal of Computer Applications, 119(??):26–29, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number20/21185-4260/.

Patel:2013:GSP

[PPK13] Manoj Kumar Patel, B. M. Pandeya, and V. Kumar. Generalization of semi-projective modules. International Journal of Computer Applications, 83(??):1–6, December 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume83/number8/14465-2757/.

Parameswaran:2012:TMB

[PPM12] T. Parameswaran, C. Palanisamy, and K. Madheswari. Topology management based energy balancing model for IPS in MANET using MEC clustering algorithm. International Journal of Computer Applications, 48(??): 37–43, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number15/7428-0456/.

Prabhakar:2011:FIB

[PPP11] Telagarapu Prabhakar, K. Satya Prasad, and P. M. K. Prasad. Free inter-band aliasing sub-band adaptive filtering with critical sampling filter bank analysis. International Journal of Computer Applications, 12(??):6– 10, January 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume12/
number11/1731-2341/.

Patel:2016:DLB

[PPP16a] Srushti Patel, Hiren Patel, and Nimisha Patel. Dynamic load balancing techniques for improving performance in cloud computing. International Journal of Computer Applications, 138(??):1–5, March 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume138/number3/24356-2016908717/.

Patil:2016:IRU

[PPP16b] Amol M. Patil, Dilip S. Patil, and Pravin S. Patil. Iris recognition using gray level co-occurrence matrix and Hausdorff dimension. International Journal of Computer Applications, 133(??):29–34, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume133/number8/23809-2016907953/.

Prakasan:2022:LCC

[PPP22] Vismaya Prakasan, Romita Pawar, and Aditee Pachpande. Land cover classification of satellite imagery using deep learning. International Journal of Computer Applications, 184(??):1–7, September 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number28/32490-2022922357/.

Pranav:2012:CSH

[PPR^{+}12] P K Pranav, T Patel, M Rathore, A J Sonowal, Tage Tapang, and Subu Laji. Computer simulation for haulage performance of power tiller. International Journal of Computer Applications, 58(??):19–25, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume58/number5/9279-3467/.

Parassa:2017:ATI

[PPR17] Yonatan Parassa, Mex U. Pesik, and Yohanis S. Rompon. Analysis of tourist information system model in North Sulawesi Province. International Journal of Computer Applications, 175(??):11–16, October 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number2/28459-2017915479/.

Patel:2011:ETA

[PPU11] Rajesh Patel, Sunil Pariyani, and Vijay Ukani. Energy and throughput analysis of hierarchical routing protocol (LEACH) for wireless sensor network. International Journal of Computer Applications, 20(??):32–36, April 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume20/number4/2420-3247/.

P:2010:NMA

[PR10a] Vani P and Jagannathan R. A novel mathematical algorithm to represent and reconstruct some bio-metric patterns. International Journal of Computer Applications, 11 (??):26–33, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number7/1594-2141/.

Purohit:2010:EIA

[PR10b] G. N. Purohit and Asmita Singh Rawat. Efficient implementation of arithmetic operations in ECC over binary fields. International Journal of Computer Applications, 6(??):5–9, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume6/number2/1056-1376/.

P:2012:ERR

[PR12] Meena Priya P and Nirmala P. Ratchagar. Effect of reaction rate on dispersion of atmospheric aerosols in the presence of electric field. International Journal of Computer Applications, 43(??):47–53, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number5/6103-8306/.

P:2015:DCE

[PR15] Vishnuprakash P and Rajesh R. Design and concept evaluation of tricycle for aged and orthopedic differentially abled persons. International Journal of Computer Applications, 111(??):18–25, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number15/19614-1505/.

Patel:2016:HAI

[PR16] Paresh M. Patel and Hitesh A. Ravani. Hybrid approach of image stitching using normalized gradient correlation and Harris corner detector. International Journal of Computer Applications, 155(??):30–35, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume155/number8/26627-2016912396/.

Pratiwi:2014:PEA

[Pra14] Dyah Pratiwi. Psychological effects of attachment on the user application SPSS with technology acceptance model. International Journal of Computer Applications, 96(??): 25–29, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number5/16790-6378/.

Priyadharshini:2013:AEF

[Pri13] P. Priyadharshini. ATC enhancement with facts devices using quantum inspired PSO. International Journal of Computer Applications, 71(??):39–45, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number22/12622-9360/.

Phaneendra:2017:TWL

[PRK17] A. Naga Phaneendra, V. Diwakar Reddy, and G. Krishnaiah. Taming of whip-lash effect in forecasting for automotive spare parts industry using Minitab and Excel spread sheet. International Journal of Computer Applications, 161(??):12–17, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number9/27175-2017913210/.

Patel:2015:IFB

[PRP15] Rajvi Patel, Manali Rajput, and Pramit Parekh. Image fusion based on local pixel information with rough-fuzzy Cmeans approach. International Journal of Computer Applications, 120(??):12–17, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number12/21278-4152/.

Padmalatha:2015:MCD

[PRR15] E. Padmalatha, C. R. K. Reddy, and Padmaja Rani. Mining concept drift from data streams by unsupervised learning. International Journal of Computer Applications, 117(??):27–34, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number15/20632-3255/.

Prajapati:2016:PAS

[PRT^{+}16] G. L. Prajapati, Abhijeet Singh Rathore, Bhavana Tanwar, Surbhi Bhadviy, and Tushar Jain. Performance analysis of selected string matching algorithms based on good suffix and bad character rule. International Journal of Computer Applications, 140(??):28–37, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume140/number9/24623-2016909445/.

Paul:2010:IMT

[PS10] Sanchita Paul and Abhimanyu Kumar Singh. Identification of metastatic tumors by using DNA nanorobot: a fuzzy logic approach. International Journal of Computer Applications, 1(??):5–10, February 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume1/number24/565-747/.

Pathan:2012:DPC

[PS12a] A. T. Pathan and A. M. Shaikh. Dielectric properties of co-substituted Li–Ni–Zn nanostructured ferrites prepared through chemical route. International Journal of Computer Applications, 45(??):24–28, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number21/7075-9745/.

Purkayastha:2012:DPL

[PS12b] Basab B Purkayastha and Kandarpa Kumar Sarma. A digital phase locked loop based system for Nakagami-m fading channel model. International Journal of Computer Applications, 42(??):1–8, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number9/5718-7778/.

Patra:2013:NNA

[PS13] Anuradha Patra and Divakar Singh. Neural network approach for text classification using relevance factor as term weighing method. International Journal of Computer Applications, 68(??):37–41, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number17/11674-7301/.

P:2014:GGM

[PS14a] Radhakrishnan P and Kanmani S. GSOFT: a generic model for teaching and learning large-scale software programs. International Journal of Computer Applications, 92 (??):38–43, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume92/
number1/15976-4869/.

Ps:2014:EPM

[Ps14b] Sheeba Ps. Estimation of parameters for model matching using genetic algorithms. International Journal of Computer Applications, 108(??):1–6, December 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume108/number10/18944-9954/.

P:2015:TTU

[PS15a] Jyothi A P and Usha Sakthivel. Trends and technologies used for mitigating energy efficiency issues in wireless sensor network. International Journal of Computer Applications, 111(??):32–40, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number3/19521-1150/.

Paramasivan:2015:NDS

[PS15b] P. Paramasivan and R. K. Santhi. Non-dominated sorting flower pollination algorithm for dynamic economic emission dispatch. International Journal of Computer Applications, 130(??):19–26, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number9/23237-2015907113/.

Phway:2015:RPS

[PS15c] Linn Linn Phway and Myint Myint Sein. Radix priority search tree in circular range searching for GIS. International Journal of Computer Applications, 109(??): 25–29, January 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume109/number9/19218-0980/.

Pradhan:2015:DTF

[PS15d] Shrijana Pradhan and Kalpana Sharma. Development of technique for formation of equi- sized clusters in wireless sensor networks for efficient enery management. International Journal of Computer Applications, 112(??):1–8, February 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume112/number14/
19731-1519/.

Prayudi:2015:DCC

[PS15e] Yudi Prayudi and Azhari Sn. Digital chain of custody: State of the art. International Journal of Computer Applications, 114(??):1–9, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume114/number5/19971-1856/.

Pawar:2016:CWB

[PS16a] Shailendra Singh Pawar and Madhu Shandilya. A compact wide band printed LPDA antenna for WLAN applications in 5 GHz band. International Journal of Computer Applications, 136(??):18–22, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume136/number10/24189-2016908587/.

Pei:2016:SPL

[PS16b] Junxiang Pei and Hongtao Shan. Software phase-locked loop based on virtual three-phase for power grid. International Journal of Computer Applications, 136(??): 34–38, February 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume136/number3/24136-2016908424/.

P:2021:LTM

[PS21] Paramesh S P and Shreedhara K S. Leveraging the text mining to automate the customer helpdesk systems. International Journal of Computer Applications, 183(??):35–41, July 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume183/number17/
32021-2021921519/.

Parveen:2018:FFD

[PSA18] Eyasmina Parveen, Koj Sambyo, and Kausar Ahmed. Foreign fiber detection in cotton using HSI approach for industrial automation. International Journal of Computer Applications, 179(??):39–42, May 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number44/29431-2018917098/.

Patwardhan:2013:PSD

[PSD13] A. M. Patwardhan, S. P. Shirsat, and S. D. Dawande. Pressure swing distillation simulation and optimization of azeotropic IBA + IBAc mixture. International Journal of Computer Applications, 66(??):9–13, March 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume66/number10/11118-6079/.

Piarsa:2012:WBG

[PSG12] I. Nyoman Piarsa, A. A. Kompyang Oka Sudana, and Gde Wahyu M. Gunadi. Web-based GIS by using spatial decision support system (SDSS) concept for searching commercial marketplace — using Google MAP API. International Journal of Computer Applications, 50(??):1–5, July 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume50/number7/7780-
0867/.

Prasad:2015:NMI

[PSM15] R. Satya Prasad, Marri Suneetha, and R. Mahesh. A novel methods of investigation on bio-medical cancer tissues using advanced clustering techniques. International Journal of Computer Applications, 129(??):34–39, November 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume129/number12/
23128-2015907059/.

Patil:2013:IEU

[PSR13] Umesh Patil, Kirtee Shevade, and Rajneesh Raikwar. Image enhancement using edge sharpening and bidirectional smooth histogram stretching. International Journal of Computer Applications, 71(??):31–39, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number2/12334-8599/.

Poonia:2011:DRP

[PSS11] Rakesh Poonia, Amit Kumar Sanghi, and Dharm Singh. DSR routing protocol in wireless ad-hoc networks: Drop analysis. International Journal of Computer Applications, 14(??): 18–21, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume14/
number7/1896-2527/.

Pai:2012:EAP

[PSS12] Smitha N. Pai, K. C. Shet, and Mruthyunjaya H S. Energy aware path establishment and maintenance for sensor network with parameters as used in agricultural field. International Journal of Computer Applications, 57(??):30– 37, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number10/9153-3406/.

Pandey:2013:CCR

[PT13] Prashant Pandey and Rajeev Tripathi. Computational complexity reduction of OFDM signals by PTS with alternate optimised grouping phase weighting method. International Journal of Computer Applications, 78(??):1– 7, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume78/
number1/13450-1027/.

Pandey:2015:NSP

[PT15] Mahesh Shankar Pandey and Bramha Swaroop Tripathi. A novel strategy for performance improvement of MISO wireless system using Alamouti STBC with 4th order FIR filter. International Journal of Computer Applications, 110(??):9–12, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number16/19399-0789/.

Papoutsidakis:2019:SIS

[PTPS19] M. Papoutsidakis, E. Theocharis, D. Piromalis, and A. Sort. SCADA implementations to supervise the water networks infrastructures in the City of Athens. International Journal of Computer Applications, 178(??):14–19, June 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume178/number25/30691-2019919046/.

P:2013:CSA

[PUW13] Constatina A. Widi P, Wiranto Herry Utomo, and Agustinus Fritz Wijaya. Customer satisfaction analysis to health service by Servqual 5 dimension method and customer satisfaction index. International Journal of Computer Applications, 70(??):17–21, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume70/number12/12014-8039/.

Panwar:2017:DSN

[PV17] Uday Panwar and Himanshu Vishnoi. Diode switch: a novel technique for mitigation of leakage power in DSM technologies. International Journal of Computer Applications, 167(??):13–19, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number14/27936-2017914209/.

Parthasarathy:2013:EUS

[PVA13] C. Parthasarathy, A. Vinodkumar, and M. Mallika Arjunan. Existence, uniqueness and stability of neutral stochastic functional integro-differential evolution equations with infinite delay. International Journal of Computer Applications, 65(??):1–7, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume65/number15/10997-6174/.

Prakash:2010:GCO

[PVL10] J. Thomas Joseph Prakash, J. Felicita Vimala, and M. Lawrence. Growth and characterization of organic NLO material. International Journal of Computer Applications, 8(??):18–21, October 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume8/number6/1214-1743/.

Patvardhan:2011:RWP

[PVL11] C. Patvardhan, A. K. Verma, and C. Vasantha Lakshmi. A robust wavelet packet based blind digital image watermarking using HVS characteristics. International Journal of Computer Applications, 36(??):6–12, December 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume36/number9/4517-6403/.

Pandey:2020:DAF

[PVR20] Shraddha Pandey, Pankaj Vyas, and Paresh Rawat. Design and analysis of frequency reconfigurable dual-band microstrip patch antenna for wireless communication. International Journal of Computer Applications, 177(??):43–47, January 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume177/number31/
31100-2020919777/.

Prihatini:2023:ILU

[PW23] Retno Prihatini and Tri Widodo. Implementation of lean UX for coworking space booking room application design. International Journal of Computer Applications, 185(??):54–60, June 2023. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume185/number17/
32792-2023922894/.

Pancardo:2012:MWM

[PWA12] Pablo Pancardo, Miguel A. Wister, and Francisco Acosta. MAIZ: a web Mashup Application for Information services to citiZens during disaster situations. International Journal of Computer Applications, 55(??):1–8, October 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume55/number3/8732-2590/.

Prudviraj:2015:CBI

[PWG15] Jeripothula Prudviraj, Rajesh Wadhvani, and Manasi Gyanchandani. Content-based image retrieval using conflation of wavelet transformation and CIECAM02 color histogram. International Journal of Computer Applications, 126(??):15–20, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume126/number5/22548-2015906052/.

Prajapati:2019:PTG

[PY19] Manisha Prajapati and Archit Yajnik. POS tagging of Gujarati text using VITERBI and SVM. International Journal of Computer Applications, 181(??):32–35, March 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume181/number43/30405-
2019918514/.

Patel:2010:PFE

[PZ10] Narendra M. Patel and Mukesh Zaveri. Parametric facial expression synthesis and animation. International Journal of Computer Applications, 3(??):34–40, June 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume3/number4/719-1011/.

Prastya:2021:IMC

[PZN21] Septyan Eka Prastya, Muhammad Zulfadhilah, and Nurhaeni. Implementation of K-modes clustering in determining traffic accident patterns. International Journal of Computer Applications, 183(??):32–37, October 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number30/32122-2021921689/.

Qasim:2014:FCT

[QA14] Sara Raad Qasim and Zainab T. Alisa. A fuzzy controller for TCP improvement over mobile networks. International Journal of Computer Applications, 100(??):42–47, August 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume100/number11/17573-

Qazdar:2015:AMA

[QCERM15] Aimad Qazdar, Chihab Cherkaoui, Brahim Er-Raha, and Driss Mammass. AeLF: Mixing adaptive learning system with learning management system. International Journal of Computer Applications, 119(??):1–8, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume119/number15/21140-4171/.

Qureshi:2013:TBM

[QJ13] Asiya Abdus Salam Qureshi and Syed Muhammad Khalid Jamal. Taxonomy based metadata classifier. International Journal of Computer Applications, 73(??):46– 52, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number19/12998-0289/.

Qiegen:2011:AID

[QJY11] Liu Qiegen, Luo Jianhua, and Zhu Yuemin. Adaptive image decomposition by improved bilateral filter. International Journal of Computer Applications, 23(??):16–22, June 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume23/number7/2900-
3798/.

Quaiyum:2011:ANN

[QKRB11] Salman Quaiyum, Yousuf Ibrahim Khan, Saidur Rahman, and Parijat Barman. Artificial neural network based short term load forecasting of power system. International Journal of Computer Applications, 30(??):1– 7, September 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume30/
number4/3633-5073/.

Qasem:2015:CNQ

[QT15] Safwan W. Qasem and Ameur A. Touir. Cardinal neighbor quadtree: a new quadtree-based structure for constant-time neighbor finding. International Journal of Computer Applications, 132(??):22–30, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number8/23615-2015907501/.

Qteishat:2014:IIS

[Qte14] Mazen Kamal Qteishat. The impact of information system success factors, human resource staff satisfaction, and e-human resource use on organizational benefit. International Journal of Computer Applications, 105(??): 1–7, November 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume105/number2/18346-9453/.

Rajkumar:2010:ORM

[RA10a] Rajkumar and Lalit Kishore Arora. Optimising up*/down* routing by minimal paths. International Journal of Computer Applications, 5(??):17–20, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume5/number1/884-1256/.

RenuMadhavi:2010:QHR

[RA10b] CH. RenuMadhavi and A. G. Ananth. Quantification of heart rate variability (HRV) data using symbolic entropy to distinguish between healthy and disease subjects. International Journal of Computer Applications, 8(??):10–13, October 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume8/number12/1258-
1770/.

Rodriguez:2012:AAL

[RA12a] Veronica Rodriguez and Gerardo Ayala. Adaptivity and adaptability of learning object’s interface. International Journal of Computer Applications, 37(??):6–13, December 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume37/number1/4570-
6535/.

Rokbani:2012:IPP

[RA12b] Nizar Rokbani and Adel M Alimi. IK-PSO, PSO inverse kinematics solver with application to biped gait generation. International Journal of Computer Applications, 58 (??):33–39, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number22/9432-3844/.

Ramadan:2018:API

[RA18] Dina A. Ramadan and Walaa Magdy A. On the alphapower inverse Weibull distribution. International Journal of Computer Applications, 181(??):6–12, August 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number11/29814-2018917657/.

Roqip:2023:IDG

[RA23] Imam Roqip and Farida Ardiani. Inventory and demand of goods Android-based: Rest API implementation on information system. International Journal of Computer Applications, 185(??):45–52, November 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number41/32964-2023923221/.

Rajinikanth:2014:GLH

[RAA14] V. Rajinikanth, J. P. Aashiha, and A. Atchaya. Gray-level histogram based multilevel threshold selection with bat algorithm. International Journal of Computer Applications, 93(??):1–8, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number16/16296-6035/.

Ranjarajan:2015:AG

[RAA15] R. Ranjarajan, Akram Alqesmah, and Anwar Alwardi. On V_{n}-arithmetic graph. International Journal of Computer Applications, 125(??):1–7, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number9/22457-2015905563/.

Rezk:2012:DSP

[RAB12] Amira Rezk, H. A. Ali, and S. I. Barakat. Database security protection based on a new mechanism. International Journal of Computer Applications, 49(??):32–38, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume49/number19/7879-1188/.

Rabie:2014:RCL

[Rab14] Tamer Rabie. Rapid 2D-3D conversion for low-cost 3D television. International Journal of Computer Applications, 102(??):1–7, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume102/number6/17816-8751/.

Rashed:2012:NMC

[RAC12] Md. Golam Rashed, Raquib Ahsan, and Sharmin Reza Chowdhury. Numerical modelling of concrete tensile strength test by wrapping scripting language with compiled library. International Journal of Computer Applications, 40(??):34– 38, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume40/
number14/5050-7468/.

Roy:2024:DCM

[RACI24] Joy Roy, Muhammad Anwarul Azim, Abu Nowshed Chy, and Mohammad Khairul Islam. Detection and classification of mahogany tree species via satellite imagery, Google Earth engine, and deep learning. International Journal of Computer Applications, 186(??):18–26, July 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number31/
detection-and-classification-of-mahogany-tree-species-
via-satellite-imagery-google-earth-engine-and-deep-
learning/.

Rad:2013:IAL

[Rad13] Javad Mohammadi Rad. Intensity adaptive LSB method applying a revised matching. International Journal of Computer Applications, 70(??):20–26, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume70/number27/12240-8468/.

Ramadan:2015:HCS

[RAE15] Noha Ramadan, Hossam Eldin H. Ahmed, and Said E. Elkhamy. Hybrid ciphering system of images based on fractional Fourier transform and two chaotic maps. International Journal of Computer Applications, 119(??):12–17, June 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume119/number11/
21110-3892/.

Rafiuzzaman:2014:FCS

[Raf14] Mohammad Rafiuzzaman. Forecasting chaotic stock market data using time series data mining. International Journal of Computer Applications, 101(??):27–34, September 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume101/number10/17725-
8169/.

Rafiei:2015:FSN

[Raf15a] Mehdi Rafiei. Full scaled numerical simulation of the critical velocity in sloping tunnels during the high load fire incidents. International Journal of Computer Applications, 126(??):38–42, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume126/number10/22592-2015906211/.

Rafiei:2015:NSF

[Raf15b] Mehdi Rafiei. Numerical simulation of a full scaled fire test of the tunnel with natural ventilation. International Journal of Computer Applications, 115(??):9–16, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume115/number1/20113-2158/.

Ragab:2015:OSE

[Rag15] Mohammad Ehab Ragab. Overcoming singularity of Euler angles in robot pose estimation using axis rotation. International Journal of Computer Applications, 128(??):6–12, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number5/22867-2015906429/.

Rahman:2012:DAP

[Rah12] Sonia Rahman. Development of an accident prediction model for intersections of Dhaka City, Bangladesh. International Journal of Computer Applications, 47(??): 10–16, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number16/7271-0389/.

Rahman:2013:SEG

[Rah13] Mahmudur Rahman. Search engines going beyond keyword search: a survey. International Journal of Computer Applications, 75(??):1–8, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume75/number17/13200-0357/.

Rahman:2014:CEF

[Rah14] Md. Riazur Rahman. Cost efficient fault tolerant decoder in reversible logic synthesis. International Journal of Computer Applications, 108(??):7–12, December 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume108/number2/18881-0160/.

Rayan:2016:PSE

[RAHH16] Ahmed M. Rayan, Ahmed A. Abdel-Hafez, and Ismail Mohamed Hafez. Provably secure encryption algorithm based on Feistel structure. International Journal of Computer Applications, 139(??):1–8, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number1/24451-2016908849/.

Rahman:2012:ENL

[RAJ^{+}12] Muhammad Mizanur Rahman, Faisal Ahmed, Mohammad Imrul Jubair, Syed Ashfaqueuddin Priom, and Imtiaz Masud Ziko. An enhanced non-linear adaptive filtering technique for removing high density salt-and-pepper noise. International Journal of Computer Applications, 38 (??):7–12, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume38/
number11/4744-6932/.

Rajendraprasad:2019:LLG

[Raj19] Hari Chandana Karnam Rajendraprasad. Lion/lioness gender recognition. International Journal of Computer Applications, 177(??):9–12, October 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number12/30947-2019919493/.

Rani:2013:NRA

[Ran13] S. Geetha Rani. A new ranking algorithm for ranking search results of search engine based on personalized user profile. International Journal of Computer Applications, 74(??):7–11, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number15/12959-9961/.

Raslan:2024:ISC

[RAR24] Khaled SH. Raslan, Almohammady S. Alsharkawy, and K. R. Raslan. iHHO-SMOTe: a cleansed approach for handling outliers and reducing noise to improve imbalanced data classification. International Journal of Computer Applications, 186(??):1–10, August 2024. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume186/number32/ihho-smote-a-cleansed-
approach-for-handling-outliers-and-reducing-noise-
to-improve-imbalanced-data-classification/.

Rasham:2016:AWS

[Ras16] Ali M. Rasham. Analysis of wind speed data and annual energy potential at three locations in Iraq. International Journal of Computer Applications, 137(??):5–16, March 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume137/number11/
24317-2016908862/.

Rasham:2017:AAW

[Ras17] Ali M. Rasham. Assessment of annual wind energy potential at three sites in Iraq for wind energy applications. International Journal of Computer Applications, 168(??):35–43, June 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume168/number5/27874-2017914417/.

Rane:2011:IMS

[RB11] Kantilal P. Rane and S. G. Bhirud. Image mosaicing with strip search algorithm based on a novel similarity measure. International Journal of Computer Applications, 27 (??):42–47, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume27/
number2/3272-4439/.

R:2014:SRR

[RB14a] Ruban R and S. Santhosh Baboo. A secure and robust reversible watermarking algorithm using fuzzy matching — quad tree segmentation (F-QTS) technique for digital images. International Journal of Computer Applications, 108(??):4–10, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number9/18937-7868/.

Rao:2014:NFB

[RB14b] Ch. Srinivasa Rao and B. Raveendra Babu. A new fuzzy based evolutionary optimization for job scheduling with TLBO. International Journal of Computer Applications, 105(??):6–11, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number3/18355-9462/.

Raval:2015:IEE

[RB15] Gaurang Raval and Madhuri Bhavsar. Improving energy estimation based clustering with energy threshold for wireless sensor networks. International Journal of Computer Applications, 113(??):41–47, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume113/number19/19938-2104/.

Refaat:2016:HAF

[RB16] Hosam E. Refaat and Usama Badawi. High available fault tolerant technique in distributed shared memory. International Journal of Computer Applications, 140(??):1–9, April 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume140/number12/24643-
2016908243/.

Ren:2011:TTF

[RBB11] Qun Ren, Marek Balazinski, and Luc Baron. Type-2 TSK fuzzy logic system and its Type-1 counterpart. International Journal of Computer Applications, 20(??): 8–13, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume20/
number6/2440-3292/.

Reges:2016:IPS

[RBDD16] J. P. Reges, E. J. Braga, L. C. Dos S.Mazza, and A. R. De Alexandria. Inserting photovoltaic solar energy to an automated irrigation system. International Journal of Computer Applications, 134(??):1–7, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number1/23875-2016907751/.

Radwan:2012:TAD

[RBF12] Muhammad Radwan, Amr Badr, and Ibrahim Farag. Task allocation in distributed artificial intelligence using boids model. International Journal of Computer Applications, 53 (??):40–46, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number2/8397-2028/.

Rattal:2019:TMM

[RBM19] Salma Rattal, Abdelmajid Badri, and Mohammed Moughit. Trajectory management of mobile VoIP over WiMAX. International Journal of Computer Applications, 178(??):38–44, July 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number29/
30723-2019919181/.

Rahman:2016:OEP

[RBN^{+}16] Md. Hafizur Rahman, Faisal Bin Al Abid, M. Naderuzzaman, Md. Arifur Rahman, and Md. Masud Reza. Optimizing and enhancing performance of MVC architecture based on data clustering technique. International Journal of Computer Applications, 134(??):42–46, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number12/23970-2016908099/.

Rodrigues:2019:EWM

[RBO19] Thiago M. B. Rodrigues, Glenda M. Botelho, and Ary H. M. Oliveira. ePerícia: Web and mobile computing environment for public forensic data management and collect. International Journal of Computer Applications, 177(??):59–69, October 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number13/30960-2019919535/.

Reddy:2010:BPI

[RBT10] PVGD Prasad Reddy, Geeta R B, and Shashikumar G. Totad. Batch processing for incremental FP-tree construction. International Journal of Computer Applications, 5(??):28–32, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume5/number5/910-1288/.

Roy:2012:BDB

[RC12] Debdutta Barman Roy and Rituparna Chaki. Baids: Detection of blackhole attack in Manet by specialized mobile agent. International Journal of Computer Applications, 40 (??):1–6, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume40/
number13/5037-7355/.

Raj:2014:SLD

[RC14] Tej Raj and Brijlata Chauhan. Single layer dual band microstrip patch antenna using probe feed. International Journal of Computer Applications, 92(??):1–5, April 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume92/number12/16058-5143/

.

Rajak:2015:ITE

[RC15] Jeetesh Kumar Rajak and Achint Chugh. Improved texture enhanced image denoising. International Journal of Computer Applications, 121(??):13–18, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume121/number2/21511-4472/.

Rafiuzzaman:2016:FLB

[RÇ16] Mohammad Rafiuzzaman and Ibrahim Çil. A fuzzy logic based agricultural decision support system for assessment of crop yield potential using shallow ground water table. International Journal of Computer Applications, 149 (??):20–31, September 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume149/number9/26025-2016911526/.

R:2013:PAB

[RD13] Sathe S R and Shrimankar D D. Parallelizing and analyzing the behavior of sequence alignment algorithm on a cluster of workstations for large datasets. International Journal of Computer Applications, 74(??):18–30, July 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume74/number21/13042-0091/.

Rashid:2017:PAV

[RD17a] Md. Mamunur Rashid and Prithwiraj Datta. Performance analysis of vehicular ad hoc network (VANET) considering different scenarios of a city. International Journal of Computer Applications, 162(??):1–7, March 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume162/number10/27276-2017913329/.

Rubasinghe:2017:TVM

[RD17b] Iresha Dilhani Rubasinghe and T. N. K. De Zoysa. Transaction verification model over double spending for peerto-peer digital currency transactions based on blockchain architecture. International Journal of Computer Applications, 163(??):24–31, April 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume163/number5/27393-2017913531/.

Rudrapal:2012:VRA

[RDD^{+}12] Dwijen Rudrapal, Smita Das, S. Debbarma, N. Kar, and N. Debbarma. Voice recognition and authentication as a proficient biometric tool and its application in online exam for P. H. people. International Journal of Computer Applications, 39(??):6–12, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number12/4870-7297/.

Riad:2015:PSR

[RDE^{+}15] R. Riad, H. Douzi, M. El Hajji, R. Harba, and F. Ros. A print-scan resilient watermarking based on Fourier transform and image restoration. International Journal of Computer Applications, 128(??):13–17, October 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume128/number15/22948-2015906192/.

Ranjani:2012:CDC

[REA12] R. Ranjani, S. Anitha Elavarasi, and J. Akilandeswari. Categorical data clustering using cosine based similarity for enhancing the accuracy of squeezer algorithm. International Journal of Computer Applications, 45(??): 41–45, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume45/
number20/7036-9705/.

Reddaiah:2019:SGA

[Red19] B. Reddaiah. A study on genetic algorithms for cryptography. International Journal of Computer Applications, 177(??):1–4, December 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number28/31071-2019919509/.

Rao:2018:IPM

[REUHTG18] Tariq Aziz Rao, Ehsan-Ul-Haq, Sadia Tariq, and Humera Batool Gill. Influence of project management in requirement engineering process for global software development. International Journal of Computer Applications, 181(??):28–35, November 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume181/number28/
30118-2018918132/.

Rezaeian:2012:ACC

[Rez12] Mehdi Rezaeian. Automatic classification of collapsed buildings using stereo aerial images. International Journal of Computer Applications, 46(??):35–42, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume46/number21/7069-9818/.

Rahman:2014:PSS

[RFAH14] Md. Mustafizur Rahman, Wang Fang, Mursheda Akter, and Shahina Haque. Performance study of several methods and selected wavelets for image compression. International Journal of Computer Applications, 103(??):21–27, October 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume103/number13/18135-
9288/.

Rehal:2021:MIF

[RG21] Munish Rehal and Akhil Goyal. Multimodal image fusion based on hybrid of Hilbert transform and intensity hue saturation using fuzzy system. International Journal of Computer Applications, 183(??):40–45, May 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number4/31920-2021921336/.

Rosas-gonzalez:2013:EAS

[RGCGCMFJ13] Armando Rosas-Gonzalez, Dulce-Maria Clemente-Guerrero, Santiago-Omar Caballero-Morales, and Jorge-Carmen Flores-Juan. An evolutionary approach for solving the N-jobs Mmachines permutation flow-shop scheduling problem with break-down times. International Journal of Computer Applications, 83(??):1–6, December 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume83/number1/14409-2488/.

Roger:2011:MLI

[RGJ11] Atsa Etoundi Roger, Abessolo Alo’o Ghislain, and Simo Bonaventure Joel. Migration of legacy information system based on business process theory. International Journal of Computer Applications, 33(??):27–34, November 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume33/number2/3994-5656/.

Reddy:2017:AWA

[RGK17] Ch. Ram Mohan Reddy, D. Evangelin Geetha, and T. V. Suresh Kumar. An appraisal of Web applications vs. Web services with respect to performance engineering using software performance engineering approach. International Journal of Computer Applications, 158(??): 20–31, January 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume158/number4/26896-2017912779/.

Roger:2013:FFI

[RGNA13] Atsa Etoundi Roger, Nkoulou Onanena Georges, Nkondock Mi Bahanag Nicolas, and Mboupda Moyo Achille. A formal framework for intrusion detection within an information system based on workflow audit. International Journal of Computer Applications, 81(??):1–10, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume81/number1/13973-1964/.

Rangaswamy:2015:IPR

[RGR15] B. Rangaswamy, N. Geethanjali, and T. Ragunathan. Improving the performance of read operations in distributed file system using global cache mechanism and anticipated parallel processing. International Journal of Computer Applications, 116(??):1–5, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number8/20353-2545/.

Rawat:2015:TSS

[RGS15] Indira Rawat, M. K. Gupta, and Virendra Singh. Test scheduling of stacked 3D SoCs with thermal aware considerations. International Journal of Computer Applications, 120(??):28–35, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number7/21241-4021/.

Rehioui:2014:TMA

[RH14] Fadoua Rehioui and Abdellatif Hair. Towards a modeling approach based on software components. International Journal of Computer Applications, 99(??):10– 17, August 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume99/
number1/17336-7178/.

Ranveer:2015:CAF

[RH15] Smita Ranveer and Swapnaja Hiray. Comparative analysis of feature extraction methods of malware detection. International Journal of Computer Applications, 120(??):1–7, June 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume120/number5/21220-3960/.

Raj:2012:ESM

[RHD12] T. Mohan Raj, S. Shahul Hammed, and Amala Deepan. Enhancing security measures by tunnelling protocol in distributed grid network. International Journal of Computer Applications, 38(??):32–36, January 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume38/number1/4654-6736/.

Rehman:2015:ASA

[RHK15] Rizwan Rehman, Gopal Chandra Hazarika, and D. Kardong. Auditory scale analysis and evaluation of phonemes in MIS-ING language. International Journal of Computer Applications, 113(??):1–5, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume113/number15/19899-2001/.

Rani:2014:IEP

[RHM14] Dimple Rani, Javad Hatami, and Diba Meysamiazad. Investigation the effect of particle swarm optimization in performance of mixture of experts. International Journal of Computer Applications, 95(??):28–32, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume95/number2/16569-6246/.

Rasham:2015:EIS

[RHM15] Ali M. Rasham, Emad Talib Hashim, and Shaymaa A. Mahdi. Experimental investigation for small horizontal portable wind turbine of different blades profiles under laboratory conditions. International Journal of Computer Applications, 127(??):25–34, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number14/22798-2015906563/.

Rahman:2018:PDT

[RHS18] Md. Abdur Rahman, Md. Abu Hasan, and Md. Saeed Siddik. Prioritizing dissimilar test cases in regression testing using historical failure data. International Journal of Computer Applications, 180(??):1–8, January 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number14/28927-2018916258/.

Rosalina:2019:ECR

[RHT19] Vidila Rosalina, Hamdan, and Agung Triayudi. Electronic customer relationship management (E-CRM) application as efforts to increase customer retention of micro small and medium enterprises (MSMEs) in Banten Indonesia. International Journal of Computer Applications, 181(??):1–8, January 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number39/30319-2019918310/.

Rahman:2012:DLL

[RHU12] Md. Mahfuzur Rahman, Mamun Hasan, and Saeed Mahmud Ullah. Design of a low-loss Y-splitter for optical telecommunication using a 2D photonics crystal. International Journal of Computer Applications, 60(??):34– 39, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number14/9763-3613/.

Rosy:2023:PAS

[RJA23] Muhamad Andika Rosy, Asep Juarna, and Dina Anggraini. Performance analysis of six chaos cryptographic algorithms for image encryption. International Journal of Computer Applications, 184(??):44–51, January 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number43/32600-2023922545/.

Raut:2015:DPS

[RJC15] Amol M. Raut, K. D. Joshi, and V. K. Chandrakar. Damping of power system oscillation using STATCOM integrated with SMES. International Journal of Computer Applications, 119(??):11–14, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number8/21086-3775/.

Roushan:2017:PAS

[RJC17] Priyesh Roushan, Rupesh Kumar Jha, and Piyush Chawla. Performance analysis of spectrum sharing based on Bayesian spectrum sensing in cognitive radio networks. International Journal of Computer Applications, 166(??):22–27, May 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume166/number6/27675-2017914045/

.

Rajput:2013:NHP

[RJS^{+}13] Navya Rajput, Ankit Jindal, Sahil Saroha, Ritesh Kumar, and Geetanjali Sharma. A novel and high performance implementation of 8 × 8 multiplier based on Vedic mathematics using 90nm hybrid PTL/CMOS logic. International Journal of Computer Applications, 69(??):27–33, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume69/number27/12146-8461/.

Radha:2016:DMD

[RJSJ16] D. Radha, R. Jayaparvathy, M. Shanmughi, and V. Jothilakshmi. Data mining in DHCP security. International Journal of Computer Applications, 142(??):6–17, May 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume142/number11/24938-2016909566/.

R:2013:FSM

[RK13] Vincent O R and Bamiro K. Fluctuations in stock market prices: What went wrong, its implications to Nigerian economyh. International Journal of Computer Applications, 63(??):13–20, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume63/number19/10573-5612/.

Rastogi:2014:DMA

[RK14] Mayuri Rastogi and Pooja Khanna. Development of morphological analyzer for Hindi. International Journal of Computer Applications, 95(??):1–5, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number17/16683-6799/.

Repe:2015:FPM

[RK15] Vishal T. Repe and Veeresh G. Kasabegoudar. Four ports multiple input multiple output diversity antenna for UWB applications. International Journal of Computer Applications, 128(??):40–48, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number13/22937-2015906702/.

Rana:2016:OCC

[RK16] Keshav Singh Rana and Rupinderdeep Kaur. Online customer care: an Android application for mobile customers using speech synthesis. International Journal of Computer Applications, 145(??):18–25, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume145/number2/25250-2016910578/.

Rathore:2017:OEE

[RK17] Preeti Rathore and Vivek Kumar. Optimization and energy efficient analysis of shortest path algorithm in WSN for node failure. International Journal of Computer Applications, 167(??):36–42, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number10/27811-2017914396/.

Razaq:2019:OPB

[RK19] Amir Razaq and Shahzada Alamgir Khan. Optical power budgeting in colorless WDM PON. International Journal of Computer Applications, 177(??):13–20, October 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume177/number8/30916-2019919455/

.

Rugendo:2021:ACM

[RK21] Titus Murithi Rugendo and Andrew Mwaura Kahonge. Access control model for container based virtual environments. International Journal of Computer Applications, 174(??):21–29, February 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number20/31792-2021921091/.

Rafiq:2013:CAE

[RKA13] Mohammad Rafiq, Manoj Kumar, and Shabeer Ahmad. Comparative analysis of energy efficient routing protocols in MANET. International Journal of Computer Applications, 82(??):1–6, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume82/number5/14109-8589/.

Rahman:2019:LSS

[RKB19] Anisur Rahman, Shafayet Khan, and Khalid Bin Salahuddin. Localization of a single sensor with respect to a single beacon using received signal strength (RSS) in terrestrial environment. International Journal of Computer Applications, 181(??):45–52, January 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number38/30286-2019918389/.

Rghioui:2014:TSC

[RKBB14] Anass Rghioui, Anass Khannous, Said Bouchkaren, and Mohammed Bouhorma. 6lo technology for smart cities development: Security case study. International Journal of Computer Applications, 92(??):54–59, April 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume92/number15/16089-5402/.

Ramschie:2023:IIM

[RKE23] Ali Ramschie, Ronny Katuuk, and Sulastri Eksan. IoT implementation for monitoring and controlling solar power plant systems. International Journal of Computer Applications, 185(??):23–31, October 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number37/32931-2023923177/.

Rashad:2014:GCO

[RKEdK14] M. Z. Rashad, A. E. Keshk, M. A. El-dosuky, and M. M. Kamal. Genetic cuckoo optimization algorithm (GCOA). International Journal of Computer Applications, 90(??): 7–12, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume90/
number3/15552-4187/.

Revankar:2010:OMS

[RKG10] P. S. Revankar, Avinash S. Kapse, and W. Z. Gandhare. Overview of a MAC scheduling algorithm for IEEE 802.16 wireless networks. International Journal of Computer Applications, 2(??):5–8, June 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume2/number8/689-968/.

Reddy:2015:ESM

[RKGK15] Ch. Ram Mohan Reddy, K. Sailaja Kumar, D. Evangelin Geetha, and T. V. Suresh Kumar. Experimental study on mail services: a comparison with activity-based performance prediction. International Journal of Computer Applications, 131(??):15–24, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume131/number9/23477-2015907382/.

Reddy:2016:CWS

[RKGK16] Ch. Ram Mohan Reddy, K. Sailaja Kumar, D. Evangelin Geetha, and T. V. Suresh Kumar. Composite Web services: a game theory application for assessing the workload sharing using Shapley value. International Journal of Computer Applications, 147(??):37–44, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume147/number1/25620-2016910980/.

Rao:2010:UUB

[RKR10] V. V. R. Maheswara Rao, V. Valli Kumari, and KVSVN Raju. Understanding user behavior using Web usage mining. International Journal of Computer Applications, 1(??):55–61, February 2010. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume1/number7/162-
286/.

Rao:2011:NAE

[RKR11] N. Gnaneswara Rao, V. Vijaya Kumar, and P. S. V. Srinivasa Rao. Novel approaches of evaluating texture based similarity features for efficient medical image retrieval system. International Journal of Computer Applications, 20 (??):20–26, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume20/
number7/2446-3303/.

Rafii:2015:DPR

[RKR15] Fadoua Rafii, M’Hamed Aït Kbir, and Badr Dine Rossi. Data preprocessing and reducing for microarray data exploration and analysis. International Journal of Computer Applications, 132(??):20–26, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number16/23679-2015907691/.

R:2017:ISL

[RKR17] Pushpalatha K R, Asha Gowda Karegowda, and D. Ramesh. Identification of similar looking bulk split grams using GLCM and CGLCM texture features. International Journal of Computer Applications, 167(??):30–36, June 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume167/number6/27778-2017914328/.

Rizvi:2016:SRP

[RKS16] Syed Wajahat A. Rizvi, Raees Ahmad Khan, and Vivek Kumar Singh. Software reliability prediction using fuzzy inference system: Early stage perspective. International Journal of Computer Applications, 145(??):16–23, July 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume145/number10/25314-
2016910774/.

Rizvi:2017:ESS

[RKS17] Syed Wajahat A. Rizvi, Raees Ahmad Khan, and Vivek Kumar Singh. Early stage software reliability modeling using requirements and object-oriented design metrics: Fuzzy logic perspective. International Journal of Computer Applications, 162(??):44–59, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume162/number2/27219-2017913290/.

Rodd:2012:MDM

[RKY12] S. F. Rodd, U. P. Kulkrani, and A. R. Yardi. Modeling of DBMS memory for performance tuning. International Journal of Computer Applications, 42(??):35– 39, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number5/5692-7738/.

Rostami:2014:NAO

[RL14] Belal Rostami and Shahriar Lotfi. A new approach to organize the results of searching the web, using a combination of ranking and genetic structure-based clustering. International Journal of Computer Applications, 89 (??):34–40, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number6/15510-4347/.

Ray:2011:DIM

[RM11] Mitrabinda Ray and Durga Prasad Mohapatra. Designing influence metric at the architectural level for improving the reliability of a system. International Journal of Computer Applications, 29(??):16–23, September 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume29/number10/3600-5001/.

Rahman:2012:PAS

[RM12] Muhammad Mokhlesur Rahman and Shalima Binta Manir. Performance analysis of sc-FDMA and OFDMA in LTE frame structure. International Journal of Computer Applications, 45(??):31–38, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number23/7090-9795/.

Rahman:2015:PAM

[RM15] Md. Habibur Rahman and M. A. Matin. On the prediction of average monsoon rainfall in Bangladesh with artificial neural network. International Journal of Computer Applications, 127(??):45–52, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number5/22729-2015906392/.

Rajarapollu:2016:DDS

[RM16] Prachi Rohit Rajarapollu and Vijay R. Mankar. Diagonal direct sub pixel based down sampling filter for antialiasing the image. International Journal of Computer Applications, 139(??):34–38, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume139/number5/24489-2016909163/.

Rahman:2019:WIO

[RMAM19] Md. Aminur Rahman, Sarker Miraz Mahfuz, and S. M. Abdullah Al-Mamun. Writer-independent offline handwritten signature verification using novel feature extraction techniques. International Journal of Computer Applications, 177(??):21–27, October 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number14/30966-2019919537/.

Ray:2010:SSA

[RMD10] Dipankar Ray, D. Dutta Majumder, and Amit Das. Synergistic study of Alzheimer diseased brain MRI with PET and SPECT images using shape based registration and fuzzy-Dempster Shafer evidence accumulation model. International Journal of Computer Applications, 12(??):18– 25, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume12/
number7/1691-2120/.

Rathore:2013:PAC

[RMG13] Suman Rathore, Priyanka Mehta, and Kapil Gupta. Performance analysis of cooperative communication network with relay selection over Rayleigh fading channel. International Journal of Computer Applications, 68(??):48– 51, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume68/
number25/11755-7151/.

Rosalina:2020:ECR

[RMH20] Vidila Rosalina, Tb Ai Munandar, and Achmad Nizar Hidayanto. Electronic citizen relationship management (e-CiRM) modeling towards serang as a smart city. International Journal of Computer Applications, 175(??):27–32, October 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume175/number25/
31608-2020920795/.

Roger:2011:FAI

[RML11] Atsa Etoundi Roger, Fouda Ndjodo Marcel, and Atouba Christian Lopez. A formal approach for the inclusion of key performance indicators in a business process modeling. International Journal of Computer Applications, 12(??):24– 31, January 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume12/
number12/1736-2357/.

R:2011:ELB

[RMP11] Arockia Xavier Annie R, Vanitha Sheba M, and Yogesh P. Enhanced link based congestion control (ELCC) in peerto-peer (P2P) based video on demand (VoD) system. International Journal of Computer Applications, 36(??):18– 26, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume36/
number1/4455-6238/.

Ramschie:2017:MFL

[RMP17] Ali A. S. Ramschie, Johan F. Makal, and Veny V. Ponggawa. Method of freon leak detection and dirty air filter in air conditioning for electrical savings. International Journal of Computer Applications, 172(??):35–40, August 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume172/number1/28218-2017915061/.

Rahman:2014:RVM

[RMR14] Md. Habibur Rahman, Md. Monzur Morshed, and Meftah Ur Rahman. Realistic vehicular mobility impact of FTM, IDM, IDM-IM and IDM-LC on VANETs. International Journal of Computer Applications, 90(??):5–12, March 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume90/number11/15762-4440/.

Raghuwanshi:2012:CBI

[RMS12] Ghanshyam Raghuwanshi, Nishchol Mishra, and Sanjeev Sharma. Content based image retrieval using implicit and explicit feedback with interactive genetic algorithm. International Journal of Computer Applications, 43(??): 8–14, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume43/
number16/6186-8665/.

Rasmi:2014:ECE

[RMS14a] Mohammad Rasmi, Hani Mimi, and Mohammad Sharif. Evaluating composite EC operations and their applicability to the on-the-fly and non-window multiplication methods. International Journal of Computer Applications, 105(??):19–26, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number6/18382-9621/.

Rizzi:2014:SGP

[RMS14b] Francesco Rizzi, Gianmario Motta, and Daniele Sacco. A system for green personal integrated mobility: Compensation engine. International Journal of Computer Applications, 105(??):5–11, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number1/18339-9447/.

Ridha:2013:FIM

[RMSS13] Oday Abdul Lateef Abdul Ridha, Bassim Sayed Mohammed, Thamir Rashed Saeed, and Jafar Wadi Abdul Sadah. FPGA implementation of multistage knapsack public key cryptosystem. International Journal of Computer Applications, 71 (??):15–21, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number19/12594-9247/.

R:2014:BEE

[RMVP14] Tanuja R, S. H. Manjula, K. R. Venugopal, and L. M. Patnaik. BHnFDIA: Energy efficient elimination of black hole and false data injection attacks in wireless sensor networks. International Journal of Computer Applications, 96 (??):1–8, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number23/16931-6957/.

Randeniya:2024:TIA

[RN24] Samitha P. Randeniya and Ruwan D. Nawarathna. Traffic image analysis using deep learning for safe vehicle navigation in roads controlled by police. International Journal of Computer Applications, 186(??):8–18, August 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number35/
traffic-image-analysis-using-deep-learning-for-safe-
vehicle-navigation-in-roads-controlled-by-police/.

Razmjooy:2014:ARC

[RNM14] Navid Razmjooy, S. Shahram Naghibzadeh, and B. Somayeh Mousavi. Automatic redeye correction in digital photos. International Journal of Computer Applications, 95 (??):9–17, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number9/16621-6473/.

Riadh:2011:WCA

[RO11] May H. Riadh and Akram M. Othman. Web content adaptation system. International Journal of Computer Applications, 23(??):32–39, June 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume23/number9/2978-3817/.

RochaDaSilva:2017:UGD

[Roc17] Fernanda Aparecida Rocha Da Silva. Using goal diagram and business process modeling to facilitate the identification of domain specific services. International Journal of Computer Applications, 175(??):1–9, October 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number7/28497-2017915597/.

Rokaya:2020:SAS

[Rok20] Mahmoud Rokaya. Sentiment analysis of social media micro blogs using power links and genetic algorithms. International Journal of Computer Applications, 177(??):28–35, March 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume177/number44/
31202-2020919967/.

Rajan:2010:OSA

[RP10] E. G. Rajan and Manas Ranjan Pradhan. Orbital structure analysis in molecular electronics. International Journal of Computer Applications, 4(??):35–43, August 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume4/number8/846-1181/.

Rajalakshmi:2014:EOL

[RP14] M. C. Rajalakshmi and A. P. Gnana Prakash. Energy optimization for large scale wireless sensor network using realtime dynamics. International Journal of Computer Applications, 108(??):40–46, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number7/18927-0293/.

Radhakrishnan:2015:EAM

[RP15] Sridevi Radhakrishnan and D. Shanmuga Priyaa. An ensemble approach on missing value handling in hepatitis disease dataset. International Journal of Computer Applications, 130(??):23–27, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number17/23302-2015907197/.

Reddy:2024:ESD

[RP24] Sareddy Shiva Reddy and Suresh Pabbojur. Enhancing software defect prediction with ensemble models based on defect relations rule learning. International Journal of Computer Applications, 186(??):7–14, September 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number42/
enhancing-software-defect-prediction-with-ensemble-
models-based-on-defect-relations-rule-learning/.

G:2014:SEM

[rPGM14] Ravi (ravinder) Prakash G and Kiran M. Shuffling expressions with MapReduce arrangements and the role of binary path symmetry. International Journal of Computer Applications, 102(??):23–30, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume102/number16/17900-8895/.

Regis:2024:SPD

[RPMH24] Adriano Regis, Daniel J. Pagano, Francisco R. M. Mota, and Marduck M. Henao. A semi-parallel data acquisition method in electrical impedance tomography using undersampling technique. International Journal of Computer Applications, 186(??):1–7, May 2024. CODEN ????

ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume186/number23/a-semi-parallel-data-
acquisition-method-in-electrical-impedance-tomography-
using-undersampling-technique/.

Rajeev:2014:SDT

[RPN14] Prajapati Rajeev, Adhikari Prabhat, and Lama Norsang. Sphere detection technique: an optimum detection scheme for MIMO system. International Journal of Computer Applications, 100(??):25–29, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number2/17498-8036/.

Roy:2012:MMC

[RPR12] Anandarup Roy, Swapan K. Parui, and Utpal Roy. A mixture model of circular-linear distributions for color image segmentation. International Journal of Computer Applications, 58(??):6–11, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume58/number9/9308-3539/.

Roshan:2012:RSB

[RPS12] Rakesh Roshan, Rabins Porwal, and Chandra Mani Sharma. Review of search based techniques in software testing. International Journal of Computer Applications, 51(??):42– 45, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number6/8050-1387/.

Rao:2010:CPS

[RPV10] S. N. Tirumala Rao, E. V. Prasad, and N. B. Venkateswarlu. A critical performance study of memory mapping on multicore processors: an experiment with k-means algorithm with large data mining data sets. International Journal of Computer Applications, 1(??):1–8, February 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume1/number9/211-358/.

Radhika:2011:NSH

[RR11] K. Radhika and A. Venugopal Reddy. Network selection in heterogeneous wireless networks based on fuzzy multiple criteria decision making. International Journal of Computer Applications, 22(??):7–10, May 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume22/number1/2550-3495/.

Reddy:2012:OWS

[RR12] B. Brahma Reddy and K. Kishan Rao. Optimization of wireless sensor network lifetime by deploying relay sensors. International Journal of Computer Applications, 49 (??):20–26, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number6/7632-0712/.

R:2016:PCP

[RR16] Umesh D R and B. Ramachandra. Parallel computing to predict breast cancer recurrence on SEER dataset using mapreduce approach. International Journal of Computer Applications, 149(??):31–35, September 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume149/number12/26052-2016911669/.

Rashid:2018:DIS

[RR18] Abdulmuttalib T. Rashid and Osama T. Rashid. Design and implementation a smart energy saving system using an Arduino and RF module. International Journal of Computer Applications, 182(??):36–44, November 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume182/number26/30143-2018918166/.

Robbani:2022:RAI

[RR22] Nizar Robbani and Imam Riadi. Risk assessment of integrated library system using COBIT 5 framework. International Journal of Computer Applications, 184(??): 25–36, September 2022. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume184/number27/32485-2022922337/.

Rahim:2017:ECT

[RRA17] Mufassir Abdur Rahim, Ghulam Rasool, and Nasir Ahmad. EMG-controlled transradial prostheses — an investigation into machine learning techniques. International Journal of Computer Applications, 174(??):1– 8, September 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume174/number3/28384-2017915354/.

Rashid:2019:DBV

[RRAFAM19] Osama T. Rashid, Abdulmuttalib T. Rashid, Israa Sabri A. Al-Forati, and Alaa I. Al-Mayoof. Designing and building an over voltage protection system for a smart home. International Journal of Computer Applications, 178(??):25–30, August 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number39/
30794-2019919276/.

Reddy:2012:ASC

[RRB12] P. Ammi Reddy, E. V. Krishna Rao, and I. Ramesh Babu. Automatic segmentation of connected thin peripheral vessels in chest CT scan images. International Journal of Computer Applications, 54(??):17–24, September 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume54/number9/8594-2353/.

Rogers:2016:TAT

[RRB16] Mwavu Rogers, Mugonza Robert, and John Businge. Tools, architectures and techniques for monitoring energy efficiency in computer networks: State of the art survey. International Journal of Computer Applications, 146(??):16–23, July 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume146/number14/25465-
2016910939/.

Rosmiati:2016:MLF

[RRP16] Rosmiati, Imam Riadi, and Yudi Prayudi. A maturity level framework for measurement of information security performance. International Journal of Computer Applications, 141(??):1–6, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume141/number8/24801-2016907930/.

Rao:2011:TAT

[RRR11] Kuda Nageswara Rao, K. Srinivas Rao, and P. Srinivasa Rao. Transient analysis of a tandem communication network with dynamic bandwidth allocation having two stage direct bulk arrivals. International Journal of Computer Applications, 13(??):14–22, January 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume13/number7/1794-2485/.

Rao:2013:IRS

[RRR13] B. Vara Prasad Rao, K. Gangadhara Rao, and B. Srinivasa Rao. Inverse Rayleigh software reliability growth model. International Journal of Computer Applications, 75 (??):1–5, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number6/13112-0470/.

Row:2017:TCN

[RRR17] S. Viziananda Row, K. Srinivasa Rao, and P. Srinivasa Rao. Truncated compound normal with gamma mixture model for mixture density estimation. International Journal of Computer Applications, 157(??):6–12, January 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume157/number3/26809-2017912643/.

Rashid:2012:PPP

[RRudS12] Asif Rashid, Zainab Riaz, and Muiz ud-din Shami. Predicting, prioritizing & planning enterprise system’s CSFs for smart factory: a balanced scorecard FPG-data-mining scenario approach. International Journal of Computer Applications, 48(??):33–39, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume48/number24/7531-0615/.

R:2012:DBO

[RRVP12] Prashanth C R, K. B. Raja, K. R. Venugopal, and L. M. Patnaik. DWT based offline signature verification using angular features. International Journal of Computer Applications, 52(??):40–48, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume52/number15/8280-1929/.

Rao:2011:CSA

[RS11] O. Srinivasa Rao and S. Pallam Setty. Comparative study of arithmetic and Huffman compression techniques for enhancing security and effective bandwidth utilization in the context of ECC for text. International Journal of Computer Applications, 29(??):44–60, September 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume29/number6/3566-4905/.

Ramachandran:2012:KCB

[RS12] G. Ramachandran and K. Selvakumar. Knowledge and cache based adaptive query searching in unstructured P2P networks. International Journal of Computer Applications, 48 (??):4–12, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number11/7390-0247/.

R:2013:PAN

[RS13a] Surender R and P. Samundiswary. Performance analysis of node mobility in beacon and non-beacon enabled IEEE 802.15.4 based wireless sensor network. International Journal of Computer Applications, 76(??):32– 35, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume76/
number12/13302-0838/.

Rabiee:2013:JSG

[RS13b] Maryam Rabiee and Hedieh Sajedi. Job scheduling in grid computing with cuckoo optimization algorithm. International Journal of Computer Applications, 62(??):38– 44, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number16/10168-5076/.

Rajput:2013:PFS

[RS13c] U S Rajput and Bal Govind Shukla. P9-factorization of symmetric complete bipartite digraph. International Journal of Computer Applications, 66(??):14–21, March 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume66/number17/11175-6199/.

Reshma:2013:FCI

[RS13d] R. Reshma and S. K. Srivatsa. A fast construction of intrusion relieved communication path based on trust level and heuristic search. International Journal of Computer Applications, 63(??):1–6, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume63/number22/10763-5168/.

Raslan:2014:TMH

[RS14a] Kamal Raslan and Zain F. Abu Shaeer. The tanh methods for the Hirota equations. International Journal of Computer Applications, 107(??):5–9, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number11/18793-0134/.

Reddy:2014:ECB

[RS14b] V. Lokeswara Reddy and B. Sailendar. Enhanced chaos based image steganography using edge adaptive and cat mapping techniques. International Journal of Computer Applications, 104(??):18–23, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number11/18247-9310/.

Rochimah:2016:MEL

[RS16] Siti Rochimah and Alhaji Sheku Sankoh. Migration of existing or legacy software systems into Web service-based architectures (reengineering process): a systematic literature review. International Journal of Computer Applications, 133(??):43–54, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume133/number3/23770-2016907801/.

Rashnodi:2011:UBA

[RSA11] Omid Rashnodi, Hedieh Sajedi, and Mohammad Saniee Abadeh. Using box approach in Persian handwritten digits recognition. International Journal of Computer Applications, 32(??):1–8, October 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume32/number3/3882-5428/.

Rajagopal:2016:FIP

[RSA16] A. Rajagopal, K. L. Sudha, and Dundi Ajay. FPGA implementation of pseudo noise sequences based on quadratic residue theory. International Journal of Computer Applications, 134(??):10–14, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number9/23941-2016907989/.

Rachid:2012:WCA

[RSB12] Maouedj Rachid, Diaf Said, and Benyoucef Boumediene. Wind characteristics analysis for selected site in Algeria. International Journal of Computer Applications, 56(??): 39–46, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number5/8890-2896/.

Rathore:2016:GFP

[RSC16] G. P. S. Rathore, Omprakash Sikhwal, and Ritu Choudhary. Generalized Fibonacci polynomials and some identities. International Journal of Computer Applications, 153(??):4–8, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume153/number12/26539-2016911990/.

Rahman:2012:PEC

[RSHI12] Md. Mahfuzur Rahman, Sanjit Kumar Saha, Md. Zakir Hossain, and Md. Babul Islam. Performance evaluation of CMN for mel-LPC based speech recognition in different noisy environments. International Journal of Computer Applications, 58(??):6–10, November 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume58/number10/9316-3548/.

R:2014:IIM

[RSJT14] Meenakshisundaram R, Albert Rabara S, Amutha J, and Daisy Premila Bai T. An integrated IP-MPLS architecture for next generation networks. International Journal of Computer Applications, 104(??):25–29, October 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume104/number3/18183-9078/.

Rani:2015:INB

[RSK15a] Pallvi Rani, Gurmohan Singh, and Manjit Kaur. Impact of negative bias temperature instability on 6T CMOS SRAM cell performance. International Journal of Computer Applications, 128(??):1–6, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number12/22922-2015906678/.

Reddy:2015:OBS

[RSK15b] V. Madhu Sudana Reddy, B. Subramanyam, and M. Surya Kalavathi. Optimal bidding strategy in south region dayahead market model using new aggregated demand model and hybrid technique. International Journal of Computer Applications, 131(??):24–33, December 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume131/number7/23462-2015907383/.

Raghuwanshi:2015:EIS

[RSM15] Kapil Raghuwanshi, Amit Saxena, and Manish Manoria. An enhanced integrated solution for identification and elimination of wormhole attack in MANET. International Journal of Computer Applications, 110(??): 16–21, January 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume110/number7/19329-0849/.

Rajalakshmi:2011:CMD

[RSP11] L. Rajalakshmi, M. V. Suganyadevi, and S. Parameswari. Congestion management in deregulated power system by locating series FACTS devices. International Journal of Computer Applications, 13(??):19–22, January 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume13/number8/1801-2404/.

Raza:2020:END

[RSRT20] Naeem Raza, Aized Amin Soofi, Habib Ur Rehman, and Mubashir Tariq. Effect of node density over the performance of DSR, TORA, and OLSR routing protocols of MANET. International Journal of Computer Applications, 177(??):34–41, February 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number39/31166-2020919888/.

Rastogi:2011:CTF

[RSS^{+}11] Ravi Rastogi, Amit Singh, Nikhil Singhal, Nitin, and Durg Singh Chauhan. Case tool: Fast interconnections with new 3-Disjoint paths MIN simulation module. International Journal of Computer Applications, 19(??):24–29, April 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume19/number6/2365-
3110/.

Raghavendran:2014:TAC

[RSSV14] Ch. V. Raghavendran, G. Naga Satish, M. V. Rama Sundari, and P. Suresh Varma. Transient analysis of communication network model with homogeneous Poisson arrivals and dynamic bandwidth allocation. International Journal of Computer Applications, 98(??):33–39, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume98/number3/17165-7226/.

Rajavat:2011:SDF

[RT11] Anand Rajavat and Vrinda Tokekar. SysRisk ña decisional framework to measure system dimensions of legacy application for rejuvenation through reengineering. International Journal of Computer Applications, 16(??):16– 19, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume16/
number2/1985-2674/.

R:2013:SAI

[RT13] Sethunadh R and Tessamma Thomas. Spatially adaptive image denoising using undecimated directionlet transform. International Journal of Computer Applications, 84(??):43– 49, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number11/14624-2968/.

Rexiramdhan:2020:DAC

[RT20] Jovy Rexiramdhan and Farid Thalib. Design and analysis of circular polarized receiver antenna as video stabilizer in UHF television. International Journal of Computer Applications, 175(??):12–18, November 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number30/31640-2020920836/.

Rana:2010:PAE

[RTD10] J. L. Rana, Deepak Singh Tomar, and Kaveesh Dashora. A practical approach for evidence gathering in windows environment. International Journal of Computer Applications, 5(??):21–27, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume5/number10/948-1326/.

R:2015:CDI

[RTK15] Dhafer R, Mohammed Ali Tawfeeq, and Ihab Ibrahim Kadhim. Correction of deformed images based on deformationmap. International Journal of Computer Applications, 122(??):40–45, July 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume122/number18/21804-5114/.

Ray:2012:PHI

[RTM12] B. N. B. Ray, Alok Ranjan Tripathy, and S. P. Mohanty. Parallel Hermite interpolation on extended Fibonacci cubes. International Journal of Computer Applications, 54(??):36– 41, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume54/
number17/8661-2534/.

Radhakrishnan:2011:JSP

[RUB11] K. Radhakrishnan, A. Unnikrishnan, and K. G. Balakrishnan. Joint state and parameter estimation of squirrel cage induction motor — a system identification approach using EM based extended Kalman filter. International Journal of Computer Applications, 27(??):24–29, August 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume27/number1/3267-4426/.

Rao:2013:PPC

[RURS13] T. C. Srinivasa Rao, Fatima Unnisa, D. Shobha Rani, and N. Srinivas. Photovoltaic power converter as an input source for SEPIC converter. International Journal of Computer Applications, 78(??):6–13, September 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume78/number13/13581-1111/.

Ruslia:2015:EPV

[Rus15] Muhammad Ruslia. The effect of presentation variety of interactive multimedia learning to the learning result. International Journal of Computer Applications, 122(??):7–12, July 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume122/number10/
21734-4921/.

Raj:2012:UMI

[RV12a] V Naga Prudhvi Raj and T Venkateswarlu. Ultrasound medical image denoising using hybrid bilateral filtering. International Journal of Computer Applications, 56(??):44– 51, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number14/8963-3171/.

Ranjan:2012:OOD

[RV12b] Rajeev Ranjan and Shirshu Varma. Object-oriented design for wireless sensor network assisted global patient care monitoring system. International Journal of Computer Applications, 45(??):8–15, May 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume45/number3/6758-9018/.

Rao:2013:CBR

[RV13] B. Bhaskara Rao and Valli Kumari Vatsavayi. Concept based ranking of results using an ontology and fuzzy network for a personalized Web search engine. International Journal of Computer Applications, 81(??):17– 24, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume81/
number13/14073-2350/.

Rani:2016:BTD

[RV16] Neha Rani and Sharda Vashisth. Brain tumor detection and classification with feed forward back-prop neural network. International Journal of Computer Applications, 146(??):1–6, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number12/25447-2016910738/.

Rahalkar:2020:EEL

[RVGP20] Chaitanya Rahalkar, Anushka Virgaonkar, Dhaval Gujar, and Sumedh Patkar. End-to-end lung cancer diagnosis on computed tomography scans using 3D CNN and explainable AI. International Journal of Computer Applications, 176(??):1–6, April 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number15/31274-2020920111/.

R:2012:INE

[RVKB12] Srinivasan R, Ramamoorthi P V, Vijayarekha K, and Venkatraman B. Image de-noising and enhancement — a comparative study on ultrasound TOFD scans. International Journal of Computer Applications, 49(??):1–5, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume49/number4/7612-0656/.

Rao:2020:DNU

[RW20] B. Bamleshwar Rao and Akhilesh A. Waoo. Design a novel user based authentication system that identify by user and device. International Journal of Computer Applications, 175(??):30–33, November 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number33/31665-2020920858/.

S:2013:ARA

[S13] Manasa S. ANLRED: a robust AQM mechanism for congestion avoidance. International Journal of Computer Applications, 81(??):1–9, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume81/number15/14196-2259/.

S:2014:GDG

[S14] Rajasree R S. Generation of dynamic group digital signature. International Journal of Computer Applications, 98(??):3–5, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number9/17209-7430/.

S:2022:SCI

[S22] Mahadevaswamy B S. Stability of conditional invariant sets of control systems. International Journal of Computer Applications, 184(??):1–13, December 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number40/32576-2022922482/.

S:2023:SGD

[S23] Mahadevaswamy B S. Stability of general dynamical systems. International Journal of Computer Applications, 184(??):1–17, January 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number44/32602-2023922552/.

S:2024:ESS

[S24] Irene Getzi S. An efficient scheme for secure similaritybased medical image retrieval using searchable symmetric encryption. International Journal of Computer Applications, 186(??):19–24, May 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number22/an-efficient-scheme-for-secure-
similarity-based-medical-image-retrieval-using-searchable-
symmetric-encryption/.

Sharma:2010:FRE

[SA10] Manju Sharma and Lalit Awasthi. Frequency reliance on EEHC-based routed wireless sensor networks. International Journal of Computer Applications, 3(??):20–24, June 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume3/number2/709-996/.

Seerat:2012:OMI

[SA12a] Bakhtawar Seerat and Farouque Azam. Opinion mining: Issues and challenges (a survey). International Journal of Computer Applications, 49(??):42–51, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume49/number9/7658-0762/.

Sehgal:2012:HNA

[SA12b] Manpreet Singh Sehgal and Anuradha. HWPDE: Novel approach for data extraction from structured Web pages. International Journal of Computer Applications, 50(??): 22–27, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume50/
number8/7791-0897/.

Shahnaz:2012:EFV

[SA12c] Andleeb Shahnaz and Farooque Azam. Enhancement in a formal V&V approach for real-time databases. International Journal of Computer Applications, 46(??): 25–32, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number17/7010-9582/.

Soni:2012:NIE

[SA12d] Aradhana Soni and Anuja Kumar Acharya. A novel image encryption approach using an index based chaos and DNA encoding and its performance analysis. International Journal of Computer Applications, 47(??):1–6, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume47/number23/7493-9944/.

Sonmez:2012:FBB

[SA12e] Mehmet Sonmez and Ayhan Akbal. FPGA-based BASK and BPSK modulators using VHDL: Design, applications and performance comparison for different modulator algorithms. International Journal of Computer Applications, 42 (??):34–40, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number13/5756-7995/.

Saleh:2013:ECM

[SA13a]  Ahmed A. Saleh and M. A. Abdou.  Evolutionary

computation methods in image watermarking. International Journal of Computer Applications, 63(??):1– 6, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume63/
number10/10499-5263/.

Singh:2013:KOP

[SA13b] Aarti Singh and Basim Alhadidi. Knowledge oriented personalized search engine: a step towards wisdom web. International Journal of Computer Applications, 76(??): 1–9, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume76/
number8/13264-0744/.

Suthendran:2014:PCB

[SA14] K Suthendran and T Arivoli. Performance comparison of blind equalization algorithms for wireless communication. International Journal of Computer Applications, 85 (??):1–6, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number13/14898-3388/.

Sa:2015:VMB

[SA15a] Inkyu Sa and Ho Seok Ahn. Visual 3D model-based tracking toward autonomous live sports broadcasting using a VTOL unmanned aerial vehicle in GPS-impaired environments. International Journal of Computer Applications, 122(??):1–7, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number7/21709-4825/.

Salam:2015:VSC

[SA15b] Asiya Abdus Salam and Ruba Mahmoud Al Salah. Vertically scrambled Caesar cipher method. International Journal of Computer Applications, 118(??):9–13, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume118/number21/20868-3269/.

Shete:2015:PIS

[SA15c] Prasanna J. Shete and R. N. Awale. Performance investigation of signal strength based gossip AODV. International Journal of Computer Applications, 115(??):6–10, April 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume115/number13/20209-
2468/.

Singh:2015:CBE

[SA15d] Harjeevan Singh and Mohit Arora. Comparison of bit error rate performance of different modulation techniques over turbulent FSO link. International Journal of Computer Applications, 109(??):20–24, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume109/number12/19241-0973/.

Saraswati:2016:IPM

[SA16] Snehi Saraswati and Neelesh Agrawal. Improvement of the performance of microstrip patch antenna. International Journal of Computer Applications, 148(??):25–34, August 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume148/number10/25794-
2016911275/.

Samarasuriya:2024:RNN

[SAA24] SR Samarasuriya, DVDS Abeysinghe, and KGK Abeywardhane. Recurrent neural network for stock market forecasting using long short-term memory and an analysis of how social media affects share prices. International Journal of Computer Applications, 186(??):9–14, February 2024. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.
org/archives/volume186/number10/recurrent-neural-
network-for-stock-market-forecasting-using-long-short-
term-memory-and-an-analysis-of-how-social-media-affects-
share-prices/.

Satu:2017:PAL

[SAAM17] Md. Shahriare Satu, Tania Akter, Md. Sadrul Arifen, and Md. Raza Mia. Predicting accidental locations of Dhaka– Aricha Highway in Bangladesh using different data mining techniques. International Journal of Computer Applications, 165(??):1–6, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number12/27622-2017914096/.

Sabharwal:2016:ACO

[Sab16] Munish Sabharwal. The assessment of concerns, opinions and perceptions of bank customers to find the significant metrics for deployment of biometrics in ebanking. International Journal of Computer Applications, 140(??):28–41, April 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume140/number5/24592-2016909301/.

Sheta:2012:DOS

[SABA12] Alaa Sheta, Mouhammd Alkasassbeh, Malik Braik, and Hafsa Abu Ayyash. Detection of oil spills in SAR images using threshold segmentation algorithms. International Journal of Computer Applications, 57(??):10–15, November 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume57/number7/9125-
3291/.

Shakiba:2013:FEF

[SaDN13] Nafiseh Mohamadi Shakiba, Mohamad ali Doostari, and Shahin Norouzi. I-FOO: an enhancement to FOO protocol. International Journal of Computer Applications, 70 (??):27–36, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number3/11944-7752/.

Sarlak:2017:NNA

[ŞAG17] Gül Gözde Biçer Şarlak, Ayşe Anapali, and Mustafa Gülsu. A new numerical approach for solving fractional Bagley– Torvik equation. International Journal of Computer Applications, 178(??):55–59, November 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number2/28650-2017915771/.

Siraj:2012:MIT

[SAH12] Fadzilah Siraj, Hashim Asman, and Md. Rajib Hasan. Modeling information technology competency using neural networks. International Journal of Computer Applications, 58 (??):1–6, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number11/9323-3623/.

Shini:2016:HTB

[SAI16] Ahlam Hanoon Shini, Zainab Ibrahim Abood, and Tariq Ziad Ismaeel. Hybrid techniques based speech recognition. International Journal of Computer Applications, 139(??):12–18, April 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume139/number10/
24525-2016909340/.

Sharma:2013:CSG

[SAK13a] Dinkar Sharma, Sangeeta Arora, and Sheo Kumar. A comparative study of Galerkin finite element and Bspline methods for two point boundary value problems. International Journal of Computer Applications, 67(??): 1–6, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number23/11532-5386/.

Siddeq:2013:AMM

[SAk13b] Mohammed Mustafa Siddeq and Ghadah Al-khafaji. Applied minimized matrix size algorithm on the transformed images by DCT and DWT used for image compression. International Journal of Computer Applications, 70(??): 33–40, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number15/12040-8000/.

Sakai:2015:MRB

[Sak15] Motoki Sakai. Modeling the relationship between heart rate and features of vocal frequency. International Journal of Computer Applications, 120(??):32–37, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume120/number6/21233-3986/.

Shehab:2018:SSA

[SAK18] Sara Shehab, Sameh Abdulah, and Arabi E. Keshk. A survey of the state-of-the-art parallel multiple sequence alignment algorithms on multicore systems. International Journal of Computer Applications, 182(??):1–9, August 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume182/number12/29869-2018917658/.

Shrestha:2024:ORP

[SAKS24] Santosh Shrestha, Pranith Varma Appani, Praveen Reddy Kota, and Alaa Sheta. Optimizing robot path planning in 2D static environments using GA, PSO and ACO search algorithms. International Journal of Computer Applications, 186(??):1–10, February 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number7/optimizing-robot-path-planning-
in-2d-static-environments-using-ga-pso-and-aco-search-
algorithms/.

Salabun:2014:AFM

[Sal14] Wojciech Salabun. Application of the fuzzy multi-criteria decision-making method to identify nonlinear decision model. International Journal of Computer Applications, 89(??): 1–6, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number15/15704-4635/.

Saleh:2015:DAF

[Sal15] Manal H. Saleh. 3-D air flow in a toroidal natural convection loop. International Journal of Computer Applications, 121(??):24–29, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number8/21561-4594/.

Salim:2017:RCS

[Sal17] Ahmed Salim. Randomized clustering scheme for heterogeneous wireless sensor networks. International Journal of Computer Applications, 173(??):1–6, September 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume173/number2/28304-2017915239/

.

Saleh:2018:SAR

[Sal18] Ma’En Saleh. Security aware routing protocol for intelligent transportation distributed multi-agent system. International Journal of Computer Applications, 180(??):5–13, January 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number10/
28896-2018916153/.

Salem:2019:OAD

[Sal19] Sameh A. Salem. Optimal assistive drive system using mobile cloud computing. International Journal of Computer Applications, 182(??):45–51, March 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number46/30465-2019918624/.

Samak:2013:NNA

[Sam13] Ahmed H. Samak. A new nonlinear anisotropic Wiener method for speckle noise reduction in optical coherence tomography. International Journal of Computer Applications, 65(??):10–15, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number12/10975-6116/.

Siham:2014:UGA

[SAM14a] Rekiek Siham, Azmani Abdellah, and El Harzli Mohamed. Using genetic algorithms for the classification of employees in the context of a transformation to the eadministration. International Journal of Computer Applications, 106(??):11–18, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number4/18507-9576/.

Subbiah:2014:CFT

[SAM14b] Dharmalingam Subbiah, Balaji Arulmozhi, and Hariharasudhan Maruthamuthu. Constraint free testing using service virtualization. International Journal of Computer Applications, 105(??):23–26, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number17/18470-9882/.

Sudana:2016:OAT

[SAP16] A. A. K. Oka Sudana, I. Gd. Km. Puja Aditya, and I. Pt. Agus Eka Pratama. Optimization algorithm in traditional card game rummy 21. International Journal of Computer Applications, 146(??):41–47, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number8/25422-2016910841/.

Sargolzaei:2013:LPE

[Sar13] M. H. Sargolzaei. A low power and error tolerant structure for motion estimation core in H.264/AVC standard. International Journal of Computer Applications, 67(??): 19–23, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number3/11375-6643/.

Sarhan:2014:WST

[Sar14] Ahmad M. Sarhan. A WPD scanning technique for iris recognition. International Journal of Computer Applications, 85(??):6–12, January 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume85/number14/14907-3446/.

Saraswat:2016:COS

[Sar16a] Manish Saraswat. Cost optimization of SRGM using genetic algorithm. International Journal of Computer Applications, 144(??):13–20, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number5/25174-2016910266/.

Singhai:2016:CLP

[SAR16b] Vikas Singhai, Saima Ayyub, and Paresh Rawat. Comparison of leakage power reduction techniques in 65nm technologies. International Journal of Computer Applications, 134(??):28–32, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number8/23936-2016907962/.

Shafiuzzaman:2018:PSA

[SAR18] Md. Shafiuzzaman, Amit Seal Ami, and Rayhanur Rahman. A proactive self-adaptive framework for context-aware mobile applications. International Journal of Computer Applications, 180(??):1–8, February 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number20/29045-2018916477/.

Sharma:2013:CTM

[SAS13] Seema Sharma, Jitendra Agrawal, and Sanjeev Sharma. Classification through machine learning technique: C4.5 algorithm based on various entropies. International Journal of Computer Applications, 82(??):28–32, November 2013. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume82/number16/14249-2444/

.

Sayed:2021:GPU

[SAS21] M. Sayed, A. Arafa, and H. I. Saleh. Graphical processing unit based implementations of crystal identification algorithms. International Journal of Computer Applications, 183(??):12–16, December 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number43/32218-2021921833/.

Sani:2024:IHM

[SAS24] Halliru Sani, Usman Nuruddeen Abdulhamid, and Musbahu Salisu. Impacts of hedonic motivation and social influence in self-broadcasting using social media platforms. International Journal of Computer Applications, 186(??):1–11, July 2024. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume186/number28/impacts-of-hedonic-motivation-
and-social-influence-in-self-broadcasting-using-social-
media-platforms/.

Soni:2011:PDM

[SASS11] Jyoti Soni, Ujma Ansari, Dipesh Sharma, and Sunita Soni. Predictive data mining for medical diagnosis: an overview of heart disease prediction. International Journal of Computer Applications, 17(??):43–48, March 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume17/number8/2237-2860/.

Sajjad:2013:VMS

[SATS13] Asma Sajjad, Sadia Arshad, Afsheen Tariq, and Ayesha Shoaib. Virtual makeover software. International Journal of Computer Applications, 80(??):9–12, October 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume80/number14/13927-1734/.

Sankoh:2016:EPS

[SAW16] Alhaji Sheku Sankoh, Agus Zainal Arifin, and Arya Yudhi Wijaya. Extracted pixels similarity features (EPSF) using interactive image segmentation techniques. International Journal of Computer Applications, 136(??): 5–12, February 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume136/number2/24123-2016908236/.

Saeed:2016:VTE

[SAY16] Saad Zaghlul Saeed, Muhamad Azhar Abdilatef, and Zead Mohammed Yosif. Visual tracking enhancement of object on circular path based on tuned Kalman filter by particle swarm optimization. International Journal of Computer Applications, 146(??):43–50, July 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume146/number4/25390-2016910685/.

Sayed:2022:EPD

[Say22] M. Sayed. An efficient program to design linear accelerator structural shielding. International Journal of Computer Applications, 183(??):15–21, January 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number47/32246-2022921870/.

Shashi:2010:IBG

[SB10] Shashi and Seema Bawa. Interoperability: Bridging the gap. International Journal of Computer Applications, 7 (??):15–21, October 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume7/
number12/1300-1603/.

Shivakumar:2011:AFM

[SB11a] B. L. Shivakumar and S. Santhosh Baboo. Automated forensic method for copy-move forgery detection based on Harris interest points and SIFT descriptors. International Journal of Computer Applications, 27(??):9–17, August 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume27/number3/3283-4472/.

SIALA:2011:SEM

[SB11b] Med Aymen Siala and Slim Ben Saoud. A survey on existing MPSOCs architectures. International Journal of Computer Applications, 19(??):28–41, April 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume19/number3/2341-3053/.

Sur:2011:CAV

[SB11c] Samarendra Nath Sur and Rabindranath Bera. Capacity analysis of the VBLAST MIMO system in fading condition with M-QAM modulation. International Journal of Computer Applications, 25(??):38–41, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume25/number3/3009-4059/.

Subha:2012:SEW

[SB12a] S. Thai Subha and Vidhyacharan Bhaskar. Spectrum efficiency of WiMAX networks in the presence of interference with diversity combining. International Journal of Computer Applications, 40(??):5–11, February 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume40/number5/4950-7190/.

Sudha:2012:GBG

[SB12b] L. R Sudha and R Bhavani. Gait based gender identification using statistical pattern classifiers. International Journal of Computer Applications, 40(??):30–35, February 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume40/number8/4985-
7248/.

Sutar:2012:UTB

[SB12c] Sanjay R. Sutar and Rajan S. Bichkar. University timetabling based on hard constraints using genetic algorithm. International Journal of Computer Applications, 42 (??):1–7, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number15/5766-7964/.

Saelee:2013:MSA

[SB13a] Jarunee Saelee and Veera Boonjing. A metadata search approach to keyword query in relational databases. International Journal of Computer Applications, 69(??): 1–7, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number3/11819-7500/.

Sidram:2013:NSS

[SB13b] M. H. Sidram and Nagappa U. Bhajantri. A novel shape signature of geometric mean of segmented centroid distance function to track the object through Fourier descriptors. International Journal of Computer Applications, 83 (??):1–6, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number14/14513-2896/.

Sharma:2015:SSM

[SB15] Ruchi Sharma and S. Balaji. SMEH: Stochastic method of energy harvesting for powering up mobile phones. International Journal of Computer Applications, 109(??): 38–45, January 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume109/number2/19164-0617/.

Soyemi:2018:WBD

[SB18a] Jumoke Soyemi and Adesi Adesola Bolaji. A Web-based decision support system with SMS-based technology for agricultural information and weather forecasting. International Journal of Computer Applications, 180(??):1–6, February 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume180/number16/29013-
2018916338/.

Suhasini:2018:TSA

[SB18b] Matla Suhasini and Srinivasu Badugu. Two step approach for emotion detection on Twitter data. International Journal of Computer Applications, 179(??):12–19, June 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number53/29539-2018917350/.

Sukoon:2018:WTF

[SB18c] Madhup Sukoon and Haider Banka. Water-tank fish algorithm: a new metaheuristic for optimization. International Journal of Computer Applications, 182(??):1–5, September 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number16/29943-
2018917835/.

Sabale:2021:RPE

[SB21] Priyanka B. Sabale and Prabhavati D. Bahirgonde. REpeating pattern extraction technique (REPET), a novel and simple approach for separating the repeating “Background” from the non-repeating “Foreground” in a mixture. International Journal of Computer Applications, 183(??):24–28, August 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number21/32049-2021921578/.

Souahlia:2012:AIT

[SBC12] Seifeddine Souahlia, Khmais Bacha, and Abdelkader Chaari. Artificial intelligence tools aided-decision for power transformer fault diagnosis. International Journal of Computer Applications, 38(??):1–8, January 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume38/number3/4665-6768/.

Shahdib:2013:ODO

[SBHM13] Fayaz Shahdib, Md. Wali Ullah Bhuiyan, Md. Kamrul Hasan, and Hasan Mahmud. Obstacle detection and object size measurement for autonomous mobile robot using sensor. International Journal of Computer Applications, 66 (??):28–33, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number9/11114-6074/.

Saini:2016:PCT

[SBKS16] Anu Saini, Ankita Bahl, Supriya Kumari, and Mitali Singh. Plagiarism checker: Text mining. International Journal of Computer Applications, 134(??):8–11, January 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume134/number3/23892-2016907833/

.

Sing:2015:LOF

[SBM^{+}15] Tan Yong Sing, Emran Bin Siraj Syahrel, Pratap Nair Marimuthu, Raman Raguraman, and K. Nithiyananthan. Local outlier factor based data mining model for three phase transmission lines faults identification. International Journal of Computer Applications, 130(??): 17–23, November 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume130/number2/23181-2015906878/.

Shabbir:2018:REH

[SBR18] Aisha Shabbir, Kamalrulnizam Abu Bakar, and Raja Zahilah Raja Mohd. Radzi. Replication effect over Hadoop MapReduce performance using regression analysis. International Journal of Computer Applications, 181(??):33–38, October 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume181/number24/30037-
2018918034/.

Shehab:2013:GFE

[SBS13] M. Elemam Shehab, K. Badran, and Gouda I. Salama. A generic feature extraction model using learnable evolution models (LEM+ID3). International Journal of Computer Applications, 64(??):27–32, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume64/number11/10680-5562/.

Singh:2012:ERA

[SC12] Sukhpal Singh and Inderveer Chana. Enabling reusability in agile software development. International Journal of Computer Applications, 50(??):33–40, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume50/number13/7834-1132/.

Sindhura:2013:ANN

[SC13] Lalam S. Sindhura and Kalpana Chaudary. Artificial neural network implementation for maximum power point tracking of optimized solar panel. International Journal of Computer Applications, 78(??):1–6, September 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume78/number10/13522-1176/.

Shanmugapriya:2014:TIS

[SC14a] N. Shanmugapriya and E. Chandra. A thorough investigation on speech enhancement techniques for hearing aids. International Journal of Computer Applications, 99(??): 9–12, August 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume99/
number13/17432-7957/.

Singh:2014:ETM

[SC14b] Ranjan Singh and Meenu Chawla. An efficient trust management technique for delay tolerant network. International Journal of Computer Applications, 98(??): 8–12, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number21/17305-7736/.

Saxena:2015:LBC

[SC15a] Richa Saxena and Sushil Kumar Chaturvedi. Learning based clustering for the automatic annotations from Web databases. International Journal of Computer Applications, 113(??):18–23, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume113/number7/19838-1692/.

Sharma:2015:ESB

[SC15b] Anubhuti Sharma and Shaila Chugh. An efficient shearlet Bayesian network based approach for image denoising. International Journal of Computer Applications, 128(??):15–20, October 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume128/number10/
22908-2015906646/.

Singh:2015:NNB

[SC15c] Puneet Kumar Singh and D. K. Chaturvedi. Neural network based modeling and simulation for estimation of maximum transformer inrush current. International Journal of Computer Applications, 123(??):35–40, August 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume123/number5/21958-2015905331/.

Song:2016:EEM

[SC16] Kyu-Hyun Song and Tae-Ho Cho. An energy-efficient method for preventing internal sinkhole attacks on IN-SENS based WSNs using interactive authentications. International Journal of Computer Applications, 153(??): 38–44, November 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume153/number2/26379-2016911978/.

Shirwal:2019:EIC

[SC19a] Vijaykumar S. Shirwal and Mahesh S. Chavan. Efficient implementation of complex matrix inversion for LMMSE decoder. International Journal of Computer Applications, 178(??):12–17, June 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number20/30649-2019918991/.

Sudharsan:2019:MAV

[SC19b] Bharath Sudharsan and Manigandan Chockalingam. A microphone array and voice algorithm based smart hearing aid. International Journal of Computer Applications, 178(??):1–6, August 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number41/30806-2019919295/.

Sangeeta:2010:GBS

[SCD10] Sangeeta, Vinay Chopra, and H. P. S. Dhami. Generating Boolean SAT based test pattern generation using multiobjective genetic algorithm. International Journal of Computer Applications, 6(??):1–4, September 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume6/number8/1099-1438/.

Sadouki:2014:EMO

[SCD14] Bouamama Reda Sadouki, Hichem Chaker, and Mouhamed Djebbouri. The effect of multipath on the OFDM system. International Journal of Computer Applications, 89 (??):6–10, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number13/15689-3933/.

Sakthivel:2014:DML

[SCE14] V. Sakthivel, Rathlavath Chandru, and Elizabeth Elias. Design of multiplier-less MDFT filter banks with perfect reconstruction using ABC algorithm. International Journal of Computer Applications, 96(??):13–22, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume96/number1/16758-6312/.

Sanya:2021:MLN

[SCH21] Waheed Sanya, Hashim Chande, and Haji Ali Haji. Maize lethal necrosis disease detection for maize crop real-time prediction yield modeling through colour pixel feature. International Journal of Computer Applications, 183(??):1–9, July 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume183/number18/32023-
2021921522/.

Suryawanshi:2017:PEV

[SCK17] U. V. Suryawanshi, S. S. Chowhan, and U. V. Kulkarni. Performance evaluation of various segmentation techniques on MRI of brain tissue. International Journal of Computer Applications, 157(??):12–20, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume157/number8/26850-2017912784/.

Sharifahmadian:2014:MDC

[SCL14] Ershad Sharifahmadian, Yoonsuk Choi, and Shahram Latifi. Multichannel data compression using wavelet subbands arranging technique. International Journal of Computer Applications, 91(??):17–22, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume91/number4/15869-4814/.

Simon:2013:EID

[SCV13] Heloisa Simon, Eros Comunello, and Aldo Von Wangenheim. Enrichment of interactive digital TV using second screen. International Journal of Computer Applications, 64 (??):58–64, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number22/10782-5764/.

Shaikh:2016:ADI

[SCW16] Md. Shahnawaz Shaikh, Ankita Choudhry, and Rakhi Wadhwani. Analysis of digital image filters in frequency domain. International Journal of Computer Applications, 140(??):12–19, April 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume140/number6/24597-2016909330/.

Sharma:2012:EAC

[SD12] Meenakshi Sharma and Sonia Dora. Efficient approach for compression in data warehouse. International Journal of Computer Applications, 53(??):1–3, September 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume53/number9/8446-
2233/.

Singh:2015:PSM

[SD15] Ankita Singh and Shishir Dixit. Placement of SVC for multiobjective function using RCGA. International Journal of Computer Applications, 120(??):34–41, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume120/number11/21275-4268/.

Saluja:2016:SSR

[SD16] Rachit Saluja and Susmita Deb. Speech signal reconstruction using two-step iterative shrinkage thresholding algorithm. International Journal of Computer Applications, 153(??):1–4, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume153/number11/26444-2016912212/.

Senarathne:2017:IIL

[SD17] A. N. Senarathne and Kasun De Zoysa. ILSB: Indexing with least significant bit algorithm for effective data hiding. International Journal of Computer Applications, 161(??):28–42, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number5/27147-2017913201/.

S:2018:NCB

[SD18] Venkatesh Prasad K S and B. P. Divakar. A new charge balancing and equalization mechanism for batteries. International Journal of Computer Applications, 180(??):16–26, January 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number11/
28906-2018916226/.

Soares:2021:ELM

[SD21] Wedson L. Soares and Roberta A. De A.Fagundes. Extreme learning machine models for predicting student performance. International Journal of Computer Applications, 174(??):1–7, February 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number22/31801-2021921122/.

S:2015:CPA

[SDB15a] Tekanyi A M S, Dajab D D, and Muazu M B. Comparative performance analysis of single-server and multiple-server Markovian models. International Journal of Computer Applications, 131(??):16–24, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume131/number14/23517-2015907507/.

Si:2015:GSB

[SDB15b] Tapas Si, Arunava De, and Anup Kumar Bhattacharjee. Grammatical swarm based segmentation methodology for lesion segmentation in brain MRI. International Journal of Computer Applications, 121(??):1–8, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume121/number4/21525-4481/.

Shaikh:2014:EBG

[SDD14] Saleha A. Shaikh, Ami A. Desai, and Amit A. Deshmukh. Electromagnetic band gap structures in MSA. International Journal of Computer Applications, 106(??): 6–14, November 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume106/number9/18546-9781/.

Schreiber:2018:RWC

[SDG18] Fábio Victor Schreiber, Roberto Alexandre Dias, and Victoria Zanetti Marçal Gallo. Remote water consumption metering using MQTT over sub-gigahertz RF environment. International Journal of Computer Applications, 180(??):10–17, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number51/29584-2018917343/.

Sarkar:2013:HEU

[SDJ13] B. C. Sarkar, S. Das, and R. N. Jana. Hall effects on unsteady MHD free convective flow past an accelerated moving vertical plate with viscous and Joule dissipations. International Journal of Computer Applications, 70 (??):19–28, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number24/12214-8351/.

Singh:2015:QAS

[SDS15] Jai Singh, Maitreyee Dutta, and Abhishek Swaroop. A QoS aware self adaptive general scheme to solve GME problem. International Journal of Computer Applications, 109(??):25–29, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume109/number7/19201-0840/.

Sabhadiya:2020:CCM

[SDS20] Parth Sabhadiya, Vaikunth Desai, and Nayankumar Sorathiya. COVID-CAM: a method of detection COVID using active map classification, CNN and deep learning. International Journal of Computer Applications, 176(??):7–13, July 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume176/number38/31449-
2020920445/.

Sendiang:2021:KSM

[SDW21] Maksy Sendiang, Fanny Doringin, and Stephy Walukow. Kinematics simulation with mobile app approachment. International Journal of Computer Applications, 174(??):11–16, January 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume174/number10/
31713-2021920966/.

Saberi:2015:UMR

[SDZ15] Farhad Fani Saberi, Sadegh Ahmadi Dastgerdi, and Mehdi Zandieh. Unified model reference adaptive attitude control of a satellite in presence of uncertain parameters: Design and implementation. International Journal of Computer Applications, 121(??):25–32, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number12/21593-4687/.

Sivagami:2010:YAI

[SE10] V. M. Sivagami and L. Anita Elizabeth. Yet another improvement in RSFT algorithm in mobile grid. International Journal of Computer Applications, 5(??):7–14, August 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume5/number4/907-1285/.

Smiti:2011:OMC

[SE11] Abir Smiti and Zied Elouedi. Overview of maintenance for case based reasoning systems. International Journal of Computer Applications, 32(??):49–56, October 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume32/number2/3881-5423/.

S:2012:DML

[SE12a] Bindiya T S and Elizabeth Elias. Design of multiplierless reconfigurable non-uniform channel filters using metaheuristic algorithms. International Journal of Computer Applications, 59(??):1–11, December 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume59/number11/9589-4213/.

Sabri:2012:GRL

[SE12b] Yassine Sabri and Najib El Kamoun. Geographic routing with logical levels forwarding for wireless sensor network. International Journal of Computer Applications, 51 (??):1–8, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number11/8083-1620/.

S:2014:DML

[SE14] Bindiya T S and Elizabeth Elias. Design of multiplier-less sharp transition width MDFT filter banks using modified metaheuristic algorithms. International Journal of Computer Applications, 88(??):1–14, February 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume88/number2/15321-3642/.

Seckiner:2015:EHS

[SE15] Serap Ulusam Seçkiner and Yunus Eroğlu. Evaluation of a harmony search algorithm for manual lifting tasks optimization. International Journal of Computer Applications, 131(??):10–17, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume131/number6/23452-2015907442/.

Sabri:2016:GRW

[SE16] Yassine Sabri and Najib El Kamoun. Geographic routing in wireless sensor networks based on a partitioned architecture. International Journal of Computer Applications, 153(??):1–8, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume153/number5/26396-2016912059/.

Sabri:2017:BDI

[SE17] Yassine Sabri and Najib El Kamoun. BIG data: Implementation a Scala approach for large scale classification. International Journal of Computer Applications, 172(??):1–6, August 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume172/number3/28228-2017915123/.

Sidi:2016:SSA

[SEA16] Emany Sidi, Mohamed El Merouani, and El Amin A. Abdelouarit. Star schema advantages on data warehouse: Using bitmap index and partitioned fact tables. International Journal of Computer Applications, 134(??):11–13, January 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume134/number13/23973-
2016908108/.

S:2013:NAM

[SEb13] Ahmed Abou Elfetouh S and Hazem M. El-bakry. A novel adaptive mobile e-learning model. International Journal of Computer Applications, 63(??):12–25, February 2013.

CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume63/number14/10533-5517/

.

Sami:2013:MMB

[SEE13] Y. Sami, R. M. Gad Elrab, and Y. Elshaer. Modeling of mechanical behavior of notched composite laminates. International Journal of Computer Applications, 71 (??):15–20, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number8/12378-8725/.

Salama:2017:NIR

[SEF17] A. A. Salama, Mohamed Eisa, and A. E. Fawzy. A neutrosophic image retrieval classifier. International Journal of Computer Applications, 170(??):1–6, July 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume170/number9/28095-2017914798/.

Selwyn:2022:DKE

[Sel22] Justus Selwyn. Design of knowledgebase for EES using object — relational data modeling. International Journal of Computer Applications, 184(??):8–14, July 2022. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number21/32439-2022922234/.

Sultan:2012:LSI

[SErEs^{+}12] Eman Sultan, Sayed El-rabaie, Fathi E. Abd El-samie, Nawal El-fishawy, and Said E. El-khamy. Least-squares interpolation of fused MR and CT images in the wavelet domain. International Journal of Computer Applications, 55 (??):1–6, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume55/
number13/8812-0622/.

Soliman:2015:MSM

[SET15] Ahmed M. Soliman, Mohamed A. Eldosoky, and Taha E. Taha. Modelling and simulation of microparticles separation using standing surface acoustic waves (SSAWs) microfluidic devices for biomedical applications. International Journal of Computer Applications, 129(??):30– 38, November 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume129/number9/23104-2015906978/.

Sbeity:2014:MMN

[SF14] Hassan Sbeity and Ahmad Fadlallah. Moodle mobile notifier. International Journal of Computer Applications, 93 (??):1–7, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number7/16224-5715/.

Salah:2015:FID

[SF15a] Ghodhbani Salah and Jemili Farah. Filtering intrusion detection alarms using ant clustering approach. International Journal of Computer Applications, 111(??):1–5, February 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume111/number17/19754-
1048/.

Siddiqui:2015:AND

[SF15b] Tariq Siddiqui and Tanveer Farooqui. Anomalous node detection and removal in ad-hoc network using enhanced prime product scheme. International Journal of Computer Applications, 122(??):28–33, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number16/21787-5081/.

Saberi:2014:DIS

[SFA14] Farhad Fani Saberi, Alireza Fazlyab, and Abbas Ajorkar. Design and implementation of a sliding mode attitude controller of a satellite in software in the loop test bed. International Journal of Computer Applications, 98(??): 28–34, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number16/17270-7653/.

Saputra:2019:IAS

[SFAS19] Tri Ichsan Saputra, Fauziah, Andez Apriansyah, and Gatot Soepriyono. Implemented analyzer and syntesis in compiler process. International Journal of Computer Applications, 178(??):35–39, August 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number43/30828-2019919324/.

Sarhan:2017:OHI

[SFH17] Amany Sarhan, Marwa Fayez, and Noha Hussen. Optimized hardware implementation of enhanced TRIPLE-DES using cluster LUT and pipelining on SPARTEN FPGA. International Journal of Computer Applications, 164(??):5–14, April 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume164/number4/27469-2017913608/.

Shaikh:2014:SSL

[SFVA14] Arshad Shaikh, Safiullah Faizullah, Danish Vasan, and Zohaib Ahmed. SLIM — simple lightweight and intuitive multicast protocol for MANETs. International Journal of Computer Applications, 94(??):45–47, May 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume94/number3/16327-5593/.

Sidhu:2010:PPA

[SG10] Jagroop Singh Sidhu and Anudeep Gandam. A postprocessing algorithm for detection & removal of corner outlier. International Journal of Computer Applications, 4(??): 9–16, July 2010. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume4/number2/804-
1144/.

Sharma:2012:CBP

[SG12a] Ritika Sharma and Kamlesh Gupta. Comparison based performance analysis of UDP/CBR and TCP/FTP traffic under AODV routing protocol in MANET. International Journal of Computer Applications, 56(??):28– 35, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number15/8969-3194/.

Singh:2012:EIF

[SG12b] Narender Singh and Nasib Singh Gill. The early identification of functional and non-functional crosscutting concerns. International Journal of Computer Applications, 46 (??):25–32, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number13/6971-9507/.

Sarhadi:2013:NAM

[SG13a] Reza Manouchehri Sarhadi and Vahid Ghafori. New approach to mitigate XML-DOS and HTTP-DOS attacks for cloud computing. International Journal of Computer Applications, 72(??):27–31, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume72/number16/12579-9201/.

Sathya:2013:VBT

[SG13b] R. Sathya and M. Kalaiselvi Geetha. Vision based traffic police hand signal recognition in surveillance video — a survey. International Journal of Computer Applications, 81 (??):1–10, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume81/
number9/14037-2192/.

Shaikh:2014:ACR

[SG14] Md. Shahnawaz Shaikh and Kamlesh Gupta. Analysis of cognitive radio spectrum sensing techniques. International Journal of Computer Applications, 102(??):1–7, September 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume102/number12/17864-
8805/.

Singh:2015:IWS

[SG15] Harsimran Singh and Vishal Gupta. An insight into word sense disambiguation techniques. International Journal of Computer Applications, 118(??):32–39, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume118/number23/20888-3666/.

Sharma:2016:CAB

[SG16a] Mukta Sharma and R. B. Garg. Comparative analysis of block key encryption algorithms. International Journal of Computer Applications, 145(??):26–35, July 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume145/number7/25292-2016910656/.

Siddiqi:2016:CAA

[SG16b] Murtaza A. Siddiqi and Naveed Ghani. Critical analysis on advanced persistent threats. International Journal of Computer Applications, 141(??):46–50, May 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume141/number13/24847-2016909784/.

Shivanka:2013:CEA

[SGA13] Shivanka, Ashu Grover, and Nikhil Arora. Controlling electrical appliances through PC and GSM technology. International Journal of Computer Applications, 76(??):44– 50, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume76/
number2/13222-0635/.

Sharma:2014:FFR

[SGB14] Nisha Sharma, Anjali Goyal, and Y. S. Brar. Fusion framework for robust and secured watermarking. International Journal of Computer Applications, 101(??):10–18, September 2014. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume101/number15/
17762-8873/.

Shabeer:2016:SHP

[SGB16] Saduf Shabeer, Ashish Grover, and Abdul Hamid Bhat. Shunt hybrid power filter for harmonic reduction to improve power quality using PI and hysteresis controller. International Journal of Computer Applications, 147(??): 7–12, August 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume147/number4/25639-2016910968/.

Singh:2013:PCD

[SGD13] Vinay Kumar Singh, Shilpi Gupta, and Upena D. Dalal. Performance comparison of Discrete Hartley Transform (DHT) and Fast Fourier Transform (FFT) OFDM system in AWGN channel. International Journal of Computer Applications, 70(??):1–4, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number9/11987-7865/.

Soares:2020:UIR

[SGD^{+}20] Carlos Soares, Eliza Gomes, Fabiano Dahlke, Carlos De Rolt, Patricia Plentz, Mario Dantas, and Vildes Scussel. Use of IoT to real-time monitoring of storage silo and ozone gas fungal decontamination strategy. International Journal of Computer Applications, 175(??):1–7, September 2020. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume175/number16/31533-2020920663/.

Sudha:2012:EAP

[SGJ12] A. Sudha, P. Gayathri, and N. Jaisankar. Effective analysis and predictive model of stroke disease using classification methods. International Journal of Computer Applications, 43(??):26–31, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number14/6172-8599/.

Soman:2012:ECT

[SGKS12] K. P. Soman, Manu Unni V G, Praveen Krishnan, and V. Sowmya. Enhancing computational thinking with spreadsheet and fractal geometry: Part 3: Mandelbrot and Julia set. International Journal of Computer Applications, 55 (??):16–23, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume55/
number14/8822-2743/.

Sepehrband:2011:EAL

[SGM^{+}11] Farshid Sepehrband, Pedram Ghamisi, Ali Mohammadzadeh, Mohammad Reza Sahebi, and Jeiran Choupan. Efficient adaptive lossless compression of hyperspectral data using enhanced DPCM. International Journal of Computer Applications, 35(??):6–11, December 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume35/number4/4386-6078/.

Singh:2013:PEE

[SGS13] Kuwar Pratap Singh, P. K. Gupta, and G. Singh. Performance evaluation of enhanced interior gateway routing protocol in IPv6 network. International Journal of Computer Applications, 70(??):42–47, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume70/number5/11962-7802/.

Shrivastava:2015:SMB

[SGS15] Ashay Shrivastava, Manish Gupta, and Shashank Swami. SPV and mutation based artificial bee colony algorithm for travelling salesman problem. International Journal of Computer Applications, 116(??):19–22, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume116/number14/20405-2753/.

Sigar:2018:IEM

[SGS18] Toar Romario Sigar, Vincencius Gunawan, and Jatmiko Endro Suseno. Implementation of the ELECTRE method for determining the location of evacuation of Web-based tsunami disaster. International Journal of Computer Applications, 180(??):25–31, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number52/29594-2018917362/.

Singh:2011:PID

[SGT^{+}11] Brijmohan Singh, Nitin Gupta, Rashi Tyagi, Ankush Mittal, and Debashish Ghosh. Parallel implementation of Devanagari text line and word segmentation approach on GPU. International Journal of Computer Applications, 24 (??):7–14, June 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume24/
number9/2987-3988/.

Singh:2012:IRB

[SH12] S. Mangijao Singh and K. Hemachandran. Image retrieval based on the combination of color histogram and color moment. International Journal of Computer Applications, 58 (??):27–34, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number3/9263-3441/.

Sayari:2015:EMT

[SH15] Zeinab Sayari and Ali Harounabadi. Evaluation and making a tradeoff between load balancing and reliability in grid services using formal models. International Journal of Computer Applications, 127(??):5–10, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number9/22755-2015906458/.

Sahu:2017:EVV

[SH17] Smita Sahu and G. C. Hazarika. The effects of variable viscosity and thermal conductivity of a MHD micropolar fluid past a continuously moving plate with Soret and Dufour effects. International Journal of Computer Applications, 180(??):19–27, December 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number1/28764-2017915894/.

Safira:2023:ICA

[SH23] Alfaeni Syafa Safira and Arief Hermawan. Implementation of C4.5 algorithm for classification of nutritional status of toddlers. International Journal of Computer Applications, 185(??):1–4, December 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number49/33019-2023923317/.

Shahbazi:2013:FTM

[Sha13] Mozhdeh Shahbazi. 2.5D feature tracking and 3D motion modeling. International Journal of Computer Applications, 64(??):43–50, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume64/number5/10634-5375/.

Shao:2014:MID

[Sha14] Deo Shao. MoMath: an innovative design of a mobile based system for supporting primary school mathematics in Tanzania. International Journal of Computer Applications, 95 (??):22–27, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number15/16671-6671/.

Shah:2016:PMB

[Sha16] Ashish Shah. Prediction of malignant and benign tumor using machine learning. International Journal of Computer Applications, 135(??):19–23, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number5/24046-2016908385/.

Sharaby:2018:IIS

[Sha18] Wafaa Hanna Zaki Sharaby. Improving image steganography using a proposed mutated Levy-flight firefly algorithm. International Journal of Computer Applications, 182(??):23–29, October 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number20/30050-2018917979/.

Shaker:2013:MAP

[SHD13] Ghada Mounir Shaker, Alaa A. Hefnawy, and Moawed I. Dessouky. A modified adaptive PCA learning based method for image denoising. International Journal of Computer Applications, 74(??):10–18, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume74/number20/13025-0103/.

Shedeed:2011:NIM

[She11] Howida Abdel Fattah Shedeed. A new intelligent methodology for computer based assessment of short answer question based on a new enhanced soundex phonetic algorithm for Arabic language. International Journal of Computer Applications, 34(??):40–47, November 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume34/number10/4136-6054/.

Sheikh:2012:CCW

[She12] Hayyan R. Sheikh. Comparing CORBA and web-services in view of a service oriented architecture. International Journal of Computer Applications, 39(??):47–55, February 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume39/number6/4828-
7083/.

Saadi:2013:COP

[SHHL13] Mostafa Saadi, M. Lahcen Hasnaoui, Abderrahim Beni Hssane, and Mohamed Laghdir. Computation of the optimal probability of becoming a cluster head in hierarchical clustered WSNs. International Journal of Computer Applications, 75(??):1–7, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume75/number1/13072-0212/.

Sekar:2011:NAA

[SHR11] P. Ananda Sekar, S. Hariharan, and P. Raguraman. A novel approach for analysis of CLR & JVM by performance metrics ñ a survey. International Journal of Computer Applications, 19(??):8–11, April 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume19/number1/2328-3025/.

Setayesh:2014:PEV

[SHR14] Sara Setayesh, Ali Harounabadi, and Amir Masoud Rahmani. Presentation of an extended version of the PageRank algorithm to rank Web pages inspired by ant colony algorithm. International Journal of Computer Applications, 85 (??):7–10, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number17/14931-3357/.

Shalaby:2023:MSC

[SHS23] Mohamed Shalaby, Doaa Helmy, and Mona Shokair. Modeling and simulation of cooperative communication based on IRSs in tera Hertz bands. International Journal of Computer Applications, 184(??):17–22, February 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number46/32614-2023922566/.

Saafan:2012:ANN

[SHSA12] Mahmoud M. Saafan, Amira Y. Haikal, Sabry F. Saraya, and Fayez F. G. Areed. Artificial neural network control of permanent magnet synchronous motor. International Journal of Computer Applications, 37(??):9–18, January 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume37/number5/4602-6560/.

Shankar:2013:HSZ

[SI13] R. Shankar and E. Ilavarasan. A highly scalable zone based multicast protocol for redundant overhead minimization. International Journal of Computer Applications, 79 (??):41–47, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number14/13813-1924/.

Salem:2014:MFB

[SIAI14] Fatty M. Salem, Maged H. Ibrahim, Ihab A. Ali, and I. I. Ibrahim. Matched-filter-based spectrum sensing for secure cognitive radio network communications. International Journal of Computer Applications, 87(??):41– 46, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number18/15312-4025/.

Singh:2013:WSP

[Sin13] Gajendra Pratap Singh. A wheel 1-safe Petri net generating all the {0; 1}^{n} sequences. International Journal of Computer Applications, 84(??):1–7, December 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume84/number16/14657-2946/.

Singh:2014:TSP

[Sin14] Rachhpal Singh. Task scheduling in parallel systems using genetic algorithm. International Journal of Computer Applications, 108(??):34–40, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number16/18999-0470/.

Singhal:2016:EBN

[Sin16] Shefali K. Singhal. An evolutionary Bayesian network learning algorithm using feature subset selection for Bayesian network classifiers. International Journal of Computer Applications, 135(??):1–8, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number13/24106-2016908057/.

Santra:2012:BCW

[SJ12] A. K. Santra and S. Jayasudha. Bayesian classifier Web object filters for user navigation with validation cross point. International Journal of Computer Applications, 52(??): 42–48, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number16/8290-1887/.

Singla:2014:PCR

[SJ14] Supriya Singla and Sushma Jain. Performance comparison of routing protocols of MANET in real world scenario using NS3. International Journal of Computer Applications, 99(??):17–23, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume99/number14/17441-8223/.

Supuwiningsih:2019:OLD

[SJ19] Ni Nyoman Supuwiningsih and I. Wayan Jepriana. Online learning development of interactive multimedia based geographic information systems. International Journal of Computer Applications, 181(??):16–20, January 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume181/number36/30266-2019918327/.

Sabouri:2013:CSE

[SJA13] Neda Sabouri, Hamid Haj Seyyed Javadi, and Toktam Zhiani Asoudeh. A comparative study on the effect of used crossover operator on performance of GA as a Web page classifier. International Journal of Computer Applications, 71(??):32–37, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume71/
number23/12629-9443/.

Salman:2021:ESM

[SJK21] Yasin K. Salman, Jalal M. Jalil, and Ahmed F. Khudheyer. Experimental study of mixed convection heat transfer to thermally developing air flow in a horizontal rectangular duct. International Journal of Computer Applications, 174(??):44–53, March 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number25/31833-2021921172/.

Shukla:2013:ERQ

[SJV13] Diwakar Shukla, Anjali Jain, and Kapil Verma. Estimation of ready queue processing time using transformed factortype (t-f-t) estimator in multiprocessor environment. International Journal of Computer Applications, 79(??):40– 48, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number16/13948-1938/.

Sunilkumar:2015:RBC

[SJV^{+}15] G Sunilkumar, Thriveni J, K R Venugopal, Manjunatha C, and L M Patnaik. Reinforcement based cognitive algorithms to detect malicious node in wireless networks. International Journal of Computer Applications, 109(??):29–34, January 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume109/number16/19273-
0990/.

Sanampudi:2010:TRN

[SK10a] Suresh Kumar Sanampudi and G. Vijaya Kumari. Temporal reasoning in natural language processing: a survey. International Journal of Computer Applications, 1(??):53–57, February 2010. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume1/number4/100-
209/.

Sharma:2010:HEE

[SK10b] Ajay K. Sharma and Harneet Kour. Hybrid energy efficient distributed protocol for heterogeneous wireless sensor network. International Journal of Computer Applications, 4(??):37–41, July 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume4/number6/828-1173/.

Singh:2010:RMC

[SK10c] P. K. Singh and Rajendra Kumar. Role of multiblocks in control flow prediction using parallel register sharing architecture. International Journal of Computer Applications, 4(??):29–32, July 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume4/number4/815-1156/.

Salhi:2012:STC

[SK12a] Houda Salhi and Samira Kamoun. Self-tuning control of a nonlinear stochastic systems described by a Hammerstein mathematical model. International Journal of Computer Applications, 39(??):15–22, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number8/4839-7101/.

Singh:2012:NWB

[SK12b] Arun Vikas Singh and Srikanta Murthy K. Neuro-wavelet based efficient image compression using vector quantization. International Journal of Computer Applications, 49 (??):32–37, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number3/7610-0653/.

Srinivasan:2012:PTI

[SK12c] M. Srinivasan and A. Krishnan. Prediction of transformer insulation life with an effect of environmental variables. International Journal of Computer Applications, 55(??):43– 48, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume55/
number5/8755-2658/.

Sandhya:2013:EAC

[SK13a] Sandhya and Vijay Katiyar. An enhanced ant colony system for solving vehicle routing problem with time window. International Journal of Computer Applications, 73 (??):27–31, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number12/12794-0080/.

Sihag:2013:GBT

[SK13b] Vikas Kumar Sihag and Subhash Kumar. Graph based text document clustering by detecting initial centroids for k-means. International Journal of Computer Applications, 62(??):1–4, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number19/10185-5005/.

S:2014:CRP

[SK14a] Muthukumar S and Jawahar P K. Cache replacement policies for improving LLC performance in multi-core processors. International Journal of Computer Applications, 105(??):12–17, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number8/18396-9655/.

Sonawane:2014:GBR

[SK14b] S. S. Sonawane and P. A. Kulkarni. Graph based representation and analysis of text document: a survey of techniques. International Journal of Computer Applications, 96(??):1–8, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number19/16899-6972/.

Subbarao:2014:HIM

[SK14c] A. V. Subbarao and P. Anil Kumar. Hardware implementation of modified weighted median filtering on FPGA. International Journal of Computer Applications, 105(??):1–8, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number4/18362-9495/.

Saiteja:2015:LFC

[SK15a] K. Saiteja and M. S. Krishnarayalu. Load frequency control of two-area smart grid. International Journal of Computer Applications, 117(??):1–9, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number14/20619-3323/.

Sharma:2015:MFH

[SK15b] B. R. Sharma and Hemanta Konwar. MHD flow, heat and mass transfer due to axially moving cylinder in presence of thermal diffusion, radiation and chemical reactions in a binary fluid mixture. International Journal of Computer Applications, 110(??):52–59, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number15/19396-1074/.

Sruthi:2015:CSS

[SK15c] G. Sruthi and G. Naveen Kumar. Comparison of series and shunt FACTS controllers for voltage stability in wind power network. International Journal of Computer Applications, 125(??):7–15, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number11/22474-2015906104/.

Suryavanshi:2015:IGC

[SK15d] Rajani Mohan Suryavanshi and Rohan Koul. Integration of GPS and cellular technologies for development of smart public transport network. International Journal of Computer Applications, 130(??):1–6, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number13/23266-2015907162/.

Sethi:2016:SSI

[SK16] Pratiksha Sethi and V. Kapoor. A secured system for information hiding in image steganography using genetic algorithm and cryptography. International Journal of Computer Applications, 144(??):5–11, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number9/25205-2016910425/.

Saroj:2018:NPE

[SK18] Preeti Saroj and Sanjay Kumar. A novel power efficient clustering technique with a fault tolerance mechanism based on Type-2 fuzzy logic, in a multi-hop WSN. International Journal of Computer Applications, 181(??):5–9, July 2018. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume181/number5/29710-2018917514/.

Singh:2013:EUM

[SKA13] Gajendra Pratap Singh, Sangita Kansal, and Mukti Acharya. Existence and uniqueness of a minimum crisp Boolean Petri net. International Journal of Computer Applications, 73 (??):35–41, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number20/13013-0316/.

Shinde:2015:ECR

[SKC15] S. V. Shinde, U. V. Kulkarni, and A. N. Chaudhary. Extracting the classification rules from general fuzzy min-max neural network. International Journal of Computer Applications, 121(??):1–7, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number23/21837-4095/.

Sharma:2014:FLM

[SKG14] Bharti Sharma, Vinod Kumar Katiyar, and Arvind Kumar Gupta. Fuzzy logic model for the prediction of traffic volume in week days. International Journal of Computer Applications, 107(??):1–6, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number17/18840-0026/.

Sharma:2017:SIW

[SKG17] Renuka Sharma, Pankaj Kumar, and Devottam Gaurav. A systematic investigation of WiMAX and its standards, development, technology and security. International Journal of Computer Applications, 160(??): 30–38, February 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume160/number3/27055-2017913012/.

Suthar:2012:ACS

[SKGP12] Krunal Suthar, Parmalik Kumar, Hitesh Gupta, and Hiren Patel. Analytical comparison of symmetric encryption and encoding techniques for cloud environment. International Journal of Computer Applications, 60(??):16– 19, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number19/9807-3054/.

Sugandhi:2017:LDM

[SKGV17] Rekha Sugandhi, Tejaswini Kasture, Yash Gupta, and Ashish Varghese. Learning disability monitor. International Journal of Computer Applications, 160(??): 37–42, February 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume160/number6/27081-2017913069/.

Singh:2015:MSC

[SKJ15] Vivek Kumar Singh, Mayur Kumar, and A. K. Jaiswal. Modeling and simulation of compact dual band rectangular micro strip patch antenna for WLAN/WiMAX applications. International Journal of Computer Applications, 121(??):30–33, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number15/21617-4896/.

Saberi:2016:SMA

[SKM16] Farhad Fani Saberi, Soheil Khalajian, and Mahmoud Mani. Sliding mode active vibration control and three axis attitude control using four reaction wheels. International Journal of Computer Applications, 152(??): 29–36, October 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume152/number1/26283-2016911754/.

Sharma:2010:IUW

[SKS10] Dhirendra Sharma, Vikram Kumar, and Vikram Singh. ICT in universities of the Western Himalayan Region of India: Study of networks, performance and efficiency. International Journal of Computer Applications, 6(??):10–15, September 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume6/number10/1111-
1455/.

Sachdeva:2013:BBO

[SKS13] Gagan Sachdeva, Ruchi Kansal, and Ashwani Singla. Biogeography based optimization for gain maximization of nine-element Yagi-Uda antenna. International Journal of Computer Applications, 68(??):1–4, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume68/number15/11652-7157/.

Singh:2016:ELS

[SKS16] Nikhil Kumar Singh, Ankit Kasana, and Vibhav Kumar Sachan. Enhancement in lifetime of sensor node using data reduction technique in wireless sensor network. International Journal of Computer Applications, 145(??):1–5, July 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume145/number11/25319-
2016910802/.

Shukla:2014:MFB

[SKT14] H S Shukla, Narendra Kumar, and R P Tripathi. Median filter based wavelet transform for multilevel noise. International Journal of Computer Applications, 107(??):11–14, December 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume107/number14/18818-
0225/.

Sylvester:2022:ODA

[SKZ22] Akpah Sylvester, Alese Boniface Kayode, and Yao Yevenyo Ziggah. Object detection in artisanal small-scale gold mining environments using a modified YOLOv4 algorithm. International Journal of Computer Applications, 183(??):36–49, February 2022. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume183/number50/
32267-2022921906/.

Sinha:2011:DAP

[SL11] Adwitiya Sinha and D. K. Lobiyal. A dynamic aggregation protocol for energy efficient data fusion in wireless sensor network. International Journal of Computer Applications, 16(??):32–38, February 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume16/number3/1991-2683/.

Sharifahmadian:2014:SCT

[SL14] Ershad Sharifahmadian and Shahram Latifi. Smart compression for telemedicine. International Journal of Computer Applications, 98(??):1–6, July 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume98/number6/17184-7363/.

Steen:2015:RST

[SL15] Tanya Steen and Ray Lindsay. RecB: Set theory based technique for large scale pattern mining in Web logs. International Journal of Computer Applications, 124(??):1–9, August 2015. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume124/number8/22128-
2015905584/.

Suri:2013:DHP

[SLK^{+}13] Lakshay Suri, Devesh Lamba, Kunwar Kritarth, Bhavna Ghai, and Geetanjali Sharma. Design of high performance and power efficient 16-bit square root carry select adder using hybrid PTL/CMOS logic. International Journal of Computer Applications, 69(??):32–35, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume69/number10/11881-7696/.

Shen:2014:MSA

[SLU14] Chi Shen, Wen Li, and Mike F. Unuakhalu. Multistep sparse approximation technology in information retrieval. International Journal of Computer Applications, 99 (??):1–8, August 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume99/
number10/17406-7991/.

Seyezhai:2010:HMI

[SM10] R. Seyezhai and B. L. Mathur. Hybrid multilevel inverter using ISPWM technique for fuel cell applications. International Journal of Computer Applications, 9(??):41–47, November 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume9/number1/1347-1817/

.

Shabrandi:2011:PPP

[SM11] A. Shabrandi and M. Mohammadi. Production planning for perishable products with partial postponement strategy. International Journal of Computer Applications, 31 (??):28–37, October 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume31/
number5/3821-5295/.

Shiroie:2012:OAC

[SM12a] Masoud Shiroie and Karim Mohammadi. Optimized asynchronous circuit design based on evolutionary algorithm. International Journal of Computer Applications, 40(??):1– 6, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume40/
number4/5029-7177/.

Singh:2012:SEE

[SM12b] Brajesh Kumar Singh and A. K. Misra. Software effort estimation by genetic algorithm tuned parameters of modified constructive cost model for NASA software projects. International Journal of Computer Applications, 59(??):22– 26, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number9/9577-4053/.

Sarkar:2013:MGE

[SM13a] S. Sarkar and R. Mahapatra. 3D microelectrode geometry effects the multilayer dense osteo intra-organelle membrane potential characterization. International Journal of Computer Applications, 80(??):1–9, October 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume80/number17/13966-1065/.

Shamsoddin-motlagh:2013:ATC

[Sm13b] Ebrahim Shamsoddin-motlagh. Automatic test case generation for orchestration languages at service oriented architecture. International Journal of Computer Applications, 80 (??):35–44, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume80/
number7/13876-1756/.

Singh:2013:ARR

[SM13c] Abhishek Singh and Rajesh Mehra. Adaptive RLSreceived signal strength algorithm in wireless network area for multi-mobile nodes location estimation system. International Journal of Computer Applications, 64(??):12– 15, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number15/10709-5671/.

Singh:2013:OLL

[SM13d] Satvir Singh and Meenaxi. An optimized low loss energyaware routing protocol for wireless sensor networks. International Journal of Computer Applications, 65(??):1– 4, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number21/11205-6468/.

Saifee:2014:DIA

[SM14a] Mufaddal A. Saifee and Usha S. Mehta. Design and implementation of 2-axis circular interpolation controller in field programmable gate array (FPGA) for computer numerical control (CNC) machines and robotics. International Journal of Computer Applications, 106(??):1–7, November 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume106/number13/18577-
9844/.

Singh:2014:PAF

[SM14b] Akanksha Singh and Pramod Kumar Mishra. Performance analysis of Floyd Warshall algorithm vs rectangular algorithm. International Journal of Computer Applications, 107(??):23–27, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number16/18837-0372/.

Sheikh:2015:MSI

[SM15a] Md. Abdul Alim Sheikh and Sumitra Mukhopadhyay. Multistage sub-image filtering technique with improved de-noising feature for road detection in disaster management. International Journal of Computer Applications, 121(??):1–8, July 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume121/number22/21829-
5086/.

Shirazi:2015:FED

[SM15b] Mohammad Shokrolah Shirazi and Brendan Tran Morris. Fast edge detection architecture using different levels of parallelism on a FPGA. International Journal of Computer Applications, 113(??):1–8, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume113/number13/19883-1914/.

Saher:2017:PBM

[SM17a] Arshed A. Saher and Jawad R. Mahmood. PLC based multi three-phase induction motors motion controller. International Journal of Computer Applications, 175(??): 37–44, October 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume175/number1/28456-2017915477/.

Salman:2017:LNC

[SM17b] Yasin K. Salman and Humam Salih Mahdi. Laminar natural convection heat transfer to air from a vertically arranged array of horizontal cylinders. International Journal of Computer Applications, 160(??):1–10, February 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume160/number5/27066-2017913047/.

Saxena:2018:IID

[SM18] Apoorv Saxena and Carsten Mueller. Intelligent intrusion detection in computer networks using swarm intelligence. International Journal of Computer Applications, 179(??):1–9, January 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume179/number16/
28880-2018916224/.

Sojib:2019:EHR

[SM19] Noushad Sojib and Sheikh Nabil Mohammad. Experiment with humanoid robot hand to reach object by measuring objects 3D coordinates using binocular stereo vision. International Journal of Computer Applications, 178(??):1–4, August 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume178/number44/
30830-2019919332/.

Soyusiawaty:2023:SMT

[SM23] Dewi Soyusiawaty and Bella Okta Sari Miranda. Statistical machine translation from Indonesian to regional languages in Indonesia. International Journal of Computer Applications, 184(??):18–23, March 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number49/32634-2023922603/.

Saleh:2015:SLF

[SMAM15] M. H. Saleh, D. Sh. Mohamed, M. H. Ahmed, and M. K. Marjan. System of linear fractional integro–differential equations by using Adomian decomposition method. International Journal of Computer Applications, 121(??):9–19, July 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume121/number24/21872-
5117/.

Srinath:2017:MIS

[SMC17] V. Srinath, Man Mohan, and D. K. Chaturvedi. A multilevel inverter system for an induction motor with openended windings. International Journal of Computer Applications, 163(??):6–13, April 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume163/number10/27429-2017913528/.

Saha:2017:BDT

[SMGS17] Payal Saha, Mohit Mittal, Shreya Gupta, and Marwa Sharawi. Big data trends and analytics: a survey. International Journal of Computer Applications, 180(??): 9–20, December 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume180/number8/28819-2017915926/.

Senthil:2010:RCD

[SMHS10] K. Senthil, K. Manikandan, Ignatius A. Herman, and L. Saravanan. Reduction of carbon dioxide emission in thermal power plants by using particle swarm optimization technique. International Journal of Computer Applications, 3(??):1–5, June 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume3/number1/704-988/.

Selvalingam:2012:RAW

[SMJ12] T. Pon Selvalingam, Siva Mahalakshmi, and S. Kurshid Jinna. Re-authentication in wireless sensor network. International Journal of Computer Applications, 55(??):32– 41, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume55/
number8/8777-2720/.

Singh:2014:IIS

[SMK14] Yashveer Singh, A K Malik, and Satish Kumar. An inflation induced stock-dependent demand inventory model with permissible delay in payment. International Journal of Computer Applications, 96(??):14–18, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume96/number25/16949-6917/.

Shuma:2017:MBC

[SMK17] Shivachi Newton Shuma, Elijah Mwangi, and Patrick Karimi. A microcontroller based carbon monoxide monitoring and mapping system using GPS technology. International Journal of Computer Applications, 163(??):1–8, April 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume163/number11/27436-2017913749/.

Singh:2011:AEA

[SMM11] Mayank Singh, Shailendra Mishra, and Rajib Mall. Accessing and evaluating AspectJ based mutation testing tools. International Journal of Computer Applications, 31(??): 33–38, October 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume31/
number1/3791-5220/.

S:2022:LST

[SMN^{+}22] Sunil Kumar D S, Mamatha Mallesh, Bharath K N, Susmitha B C, and Kiran. Long short-term memory (LSTM) based epileptic seizure recognition. International Journal of Computer Applications, 184(??):23–29, July 2022. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume184/number22/32449-2022922260/.

Susetianingtias:2016:AED

[SMR16] Diana Tri Susetianingtias, Sarifuddin Madenda, and Rodiah. An approach to exudates detection using color reference segmentation in retinal fundus image. International Journal of Computer Applications, 146(??):25–29, July 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume146/number2/25372-2016910692/.

Sarathe:2016:MMA

[SMS16] Reena Sarathe, Amit Mishra, and Shiv Kumar Sahu. Maxmin ant system based approach for intelligent VM migration and consolidation for green cloud computing. International Journal of Computer Applications, 136(??):15–18, February 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume136/number13/24213-
2016908629/.

Sahu:2014:TUA

[SMSC14] Pradip Kumar Sahu, Kanchan Manna, Tapan Shah, and Santanu Chattopadhyay. Thermal uniformity-aware application mapping for network-on-chip design. International Journal of Computer Applications, 99(??):8–22, August 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume99/number3/17351-7838/.

Selvam:2010:AQF

[SMSM10] P. M. Sithar Selvam, M. S. Muthuraman, M. Sridharan, and R. Muthuraj. Anti Q-fuzzy BG —ideals in BG — algebra. International Journal of Computer Applications, 4(??):27–31, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume4/number11/867-1217/.

Srinivasan:2017:ABD

[SMT17] Satish M. Srinivasan, Michael Martin, and Abhishek Tripathi. ANN based data mining analysis of the Parkinson’s disease. International Journal of Computer Applications, 168(??):1–7, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume168/number1/27836-2017914254/.

Safarian:2019:GBE

[SMT19] Mohammad Reza Safarian, Ghazal Moradi, and Saber Izadpanah Toos. Gain and bandwidth enhancement in CMOS low-voltage low-power operational amplifiers. International Journal of Computer Applications, 177(??):8–14, December 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume177/number23/31035-
2019919671/.

El-kenawy:2012:DGR

[sMTEkEdAr12] El sayed M. T. El-kenawy, Ali Ibraheem El-desoky, and Mohamed F. Al-rahamawy. Distributing graphic rendering using grid computing with load balancing. International Journal of Computer Applications, 47(??):1–6, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume47/number9/7213-0001/.

Sakthivel:2011:OLS

[SMVK11] S. Sakthivel, D. Mary, R. Vetrivel, and V. Senthamarai Kannan. Optimal location of SVC for voltage stability enhancement under contingency condition through PSO algorithm. International Journal of Computer Applications, 20 (??):30–36, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume20/
number1/2396-3183/.

Santhi:2010:MNA

[SN10a] V. Santhi and A. M. Natarajan. MNEWQUE: a new approach to TCP/AQM with ECN. International Journal of Computer Applications, 3(??):28–35, June 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume3/number3/713-1005/.

Sivakumar:2010:CSS

[SN10b] R. Sivakumar and D. Nedumaran. Comparative study of speckle noise reduction of ultrasound B-scan images in matrix laboratory environment. International Journal of Computer Applications, 10(??):46–50, November 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume10/number9/1506-2024/.

Sharma:2014:RCA

[SN14a] Manoj Sharma and Arti Noor. Reconfigurable CPL adiabatic gated logic — RCPLAG based universal NAND/NOR gate. International Journal of Computer Applications, 95 (??):27–32, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number26/16961-7078/.

Sharma:2014:PEV

[SN14b] Shivani Sharma and Ravindra Pratap Narwaria. Performance evaluation of various window techniques for noise cancellation from ECG signal. International Journal of Computer Applications, 93(??):1–5, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume93/number19/16464-5826/.

Saha:2015:DCB

[SN15a] Sanjib Saha and Debashis Nandi. Data classification based on decision tree, rule generation, Bayes and statistical methods: an empirical comparison. International Journal of Computer Applications, 129(??):36–41, November 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume129/number7/23089-2015906891/.

Samreen:2015:DNT

[SN15b] Shirina Samreen and G. Narsimha. Design of a novel trust model and its application in trust based routing to defend against dishonest recommenders. International Journal of Computer Applications, 122(??):16–23, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume122/number2/21672-4754/.

Samreen:2015:TBD

[SN15c] Shirina Samreen and G. Narsimha. Trust based data plane security mechanism for a mobile ad hoc network through acknowledgement reports. International Journal of Computer Applications, 129(??):6–13, November 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume129/number6/23075-2015906926/.

Sundaram:2016:HTB

[SN16] P. Kalyana Sundaram and R. Neela. Hilbert transform based fuzzy expert system for diagnosing and classifying power quality disturbances. International Journal of Computer Applications, 142(??):48–55, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number3/24880-2016909729/.

S:2022:IPL

[SN22] Pemasiri S S B P S and Vidanagama V G T N. Identification of paddy leaf diseases using machine learning techniques. International Journal of Computer Applications, 183(??):1–5, January 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number49/32260-2022921898/.

Soglo:2013:ABN

[SNA13] Giovani Soglo, Michael Osei Ntow, and Amevi Acakpovi. Automated business name reservation and registration system — a case study of Registrar General’s Department, Ghana. International Journal of Computer Applications, 77 (??):39–45, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number16/13571-1421/.

Saleh:2015:LWH

[SNA15] M. H. Saleh, A. S. Nagdy, and M. E. M. Alngar. Legendre wavelet and He’s homotopy perturbation methods for linear fractional integro-differential equations. International Journal of Computer Applications, 110(??):25–31, January 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume110/number10/19354-
1063/.

Shivamurthy:2018:CEM

[SNB18] P. M. Shivamurthy, T. N. Nagabhushan, and Vijaya Basavaraj. A combined expectation maximization and marker controlled watershed driven distance regularized level sets for nuclear segmentation in histopathological images. International Journal of Computer Applications, 180(??):1–8, April 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume180/number34/
29263-2018916884/.

Sahu:2013:IVC

[SND13] Sanjaya Kumar Sahu, D. D. Neema, and T. V. Dixit. Indirect vector control of induction motor using ANN estimator and ANFIS controller. International Journal of Computer Applications, 66(??):1–6, March 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume66/number14/11149-6246/.

Ismail:2013:MAC

[sNIALY13] Nor syahidatul N. Ismail, Sharifah H. S. Ariffin, N. M. Abdul Latiff, and Farizah Yunus. Medium access control with token approach in wireless sensor network. International Journal of Computer Applications, 65(??):51–57, March 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume65/number17/11020-6395/.

Singh:2016:BPA

[SNJ16] Thakur Pushpendra Singh, Navendu Nitin, and A. K. Jaiswal. Ber performance analytical approach for conventional DFT and DWT for different modulation techniques using SUI channel in OFDM. International Journal of Computer Applications, 149(??):15–20, September 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume149/number10/26033-2016911602/.

Sathyabama:2015:EAD

[SNP15] A. R. Sathyabama, C. M. Nalayini, and S. Priyadharshini. Earnest access of divulging and aversion of DDOS attack. International Journal of Computer Applications, 123(??):18–21, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number11/22003-2015905588/.

Sakai:2013:ESM

[SO13] Motoki Sakai and Yuichi Okuyama. Estimation of skin moisture and elasticity from facial image by using kernel ridge regression. International Journal of Computer Applications, 72(??):12–18, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume72/number3/12473-8866/.

Salama:2014:BPA

[SO14a] Ahmed S. Salama and Amany A. Omar. A back propagation artificial neural network based model for detecting and predicting fraudulent financial reporting. International Journal of Computer Applications, 106(??): 1–8, November 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume106/number2/18489-9521/.

Saleh:2014:SCB

[SO14b] Ameer L. Saleh and Adel A. Obed. Speed control of brushless DC motor based on fractional order PID controller. International Journal of Computer Applications, 95(??):1–6, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number4/16579-6269/.

Soriyan:2014:TMD

[SO14c] Abimbola Soriyan and Theresa Omodunbi. Trends in multi-document summarization system methods. International Journal of Computer Applications, 97(??):46– 52, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number16/17095-7804/.

S:2016:DGD

[SOME16] Falohun A S, Oke A O, Abolaji B M, and Oladejo O E. Dangerous gas detection using an integrated circuit and MQ-9. International Journal of Computer Applications, 135(??):30–34, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number7/24064-2016908473/.

Segun:2015:SMA

[SOO15] Adebayo Segun, Akinwunmi A O, and Ogunti E O. A survey of medium access control protocols in wireless sensor network. International Journal of Computer Applications, 116(??):1–8, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number22/20465-9980/.

Samuel:2013:ECB

[SOOA13] O. W. Samuel, M. O. Omisore, B. A. Ojokoh, and E. J. Atajeromavwo. Enhanced cloud based model for healthcare delivery organizations in developing countries. International Journal of Computer Applications, 74(??): 34–39, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number2/12861-9834/.

Saturday:2018:DAG

[SOT18] E. G. Saturday, J. C. Ofodu, and M. S. Torbira. Development and application of gas turbine performance analysis software: Part II — modified cycles. International Journal of Computer Applications, 179(??):6–14, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number51/29522-2018917250/.

Soultanian:2015:RHI

[Sou15] Victor Setrag George Soultanian. The role of hiding information in information systems on organization continuation and strength. International Journal of Computer Applications, 119(??):34–40, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number2/21043-3661/.

Stuwart:2010:RFB

[SP10] S. Lenty Stuwart and N. R. Nantha Priya. Robust feature based image watermarking process. International Journal of Computer Applications, 4(??):13–16, July 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume4/number5/825-1168/.

Sathappan:2011:EVQ

[SP11a] S. Sathappan and S. Pannirselvam. An enhanced vector quantization method for image compression with modified fuzzy possibilistic C-means using repulsion. International Journal of Computer Applications, 21(??):27–34, May 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume21/number5/2505-
3387/.

Sibanda:2011:NAM

[SP11b] Wilbert Sibanda and Philip Pretorius. Novel application of multi-layer perceptrons (MLP) neural networks to model HIV in South Africa using seroprevalence data from antenatal clinics. International Journal of Computer Applications, 35(??):26–31, December 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume35/number5/4398-6106/.

Shanmugapriya:2012:MPM

[SP12] B. Shanmugapriya and M. Punithavalli. A modified projected K-means clustering algorithm with effective distance measure. International Journal of Computer Applications, 44(??):32–36, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number8/6285-8468/.

Sahu:2013:CCO

[SP13] Yatendra Sahu and R. K. Pateriya. Cloud computing overview with load balancing techniques. International Journal of Computer Applications, 65(??):40– 44, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number24/11236-6559/.

S:2015:ETT

[SP15a] Ratnaparkhi Punam S and S. E. Pawar. Extended term tXSchema uses SAX parsing partially. International Journal of Computer Applications, 113(??):31–37, March 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume113/number16/19914-
2043/.

Saini:2015:NIR

[SP15b] Dinesh Kumar Saini and Lakshmi Sunil Prakash. Nature inspired recommender algorithms for collaborative Web based learning environments. International Journal of Computer Applications, 114(??):16–22, March 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume114/number14/20046-2048/.

Satish:2015:SOL

[SP15c] A. Rama Satish and P. Bala Krishna Prasad. Spotting outliers in large distributed datasets using cell density based approach. International Journal of Computer Applications, 122(??):1–7, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number8/21717-4858/.

Shah:2015:OFP

[SP15d] Arpan H. Shah and Pratik A. Patel. Optimum frequent pattern approach for efficient incremental mining on large databases using map reduce. International Journal of Computer Applications, 120(??):25–29, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume120/number4/21217-3933/.

Singh:2015:EET

[SP15e] Jai Vardhan Singh and Girijesh Prasad. Enhancing an eyetracker based human-computer interface with multi-modal accessibility applied for text entry. International Journal of Computer Applications, 130(??):16–22, November 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume130/number16/23293-
2015907194/.

S:2016:IEC

[SP16] Jadhav Sonali S and B. M. Patil. Integrating encrypted cloud database services using query processing. International Journal of Computer Applications, 148(??):25–30, August 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume148/number12/25810-
2016911323/.

S:2018:CHD

[SP18a] Shyam Mohan J S and Shanmugapriya P. Clustering of huge datasets using machine intelligence techniques. International Journal of Computer Applications, 181(??):8–14, September 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume181/number18/29961-
2018917856/.

Sendiang:2018:IMT

[SP18b] Maksy Sendiang and Jerry H. Purnama. Implementation of microservice technology on mobile learning development. International Journal of Computer Applications, 182(??):1–5, October 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number23/30070-2018918013/.

Singhal:2018:RTT

[SP18c] Varun Kumar Singhal and Shaik Raheem Pasha. Real time traffic sign detection recognition using adaptive neuro fuzzy inference system. International Journal of Computer Applications, 179(??):30–36, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number49/29510-2018917284/.

Siccardi:2012:ECA

[SPB12] Stefano Siccardi, Rita Pizzi, and Giuliano Benenti. Entanglement computation in atoms and molecules. International Journal of Computer Applications, 60(??):43– 48, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number1/9660-4081/.

Sarawagi:2017:RTS

[SPBP17] Ankit Sarawagi, Rajeev Pandey, Raju Barskar, and S. P. Pandey. A real time stream data processing and analysis model and catchments over Twitter stream data. International Journal of Computer Applications, 179(??): 22–33, December 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume179/number1/28700-2017915663/.

Sarode:2013:CDA

[SPC13] Vidya Sarode, Sneha Patkar, and Alice N Cheeran. Comparison of 2-D algorithms in ElT based image reconstruction. International Journal of Computer Applications, 69 (??):6–11, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume69/
number8/11860-7642/.

Satsangi:2016:ASR

[SPK16] Swanti Satsangi, C. Patvardhan, and P. K. Kalra. Automatic synthesis of reversible circuits. International Journal of Computer Applications, 140(??):31–36, April 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume140/number11/24641-2016909518/.

Shaikh:2015:PDT

[SPL15] Mustafa M. Shaikh, Ashwini A. Pawar, and Vibha B. Lahane. Pattern discovery text mining for document classification. International Journal of Computer Applications, 117(??):6–12, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume117/number1/20516-2101/.

Sethi:2011:EEE

[SPP11] Hemant Sethi, Devendra Prasad, and R. B. Patel. EIRDA: an energy efficient interest based reliable data aggregation protocol for wireless sensor networks. International Journal of Computer Applications, 22(??):20–25, May 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume22/number7/2595-3605/.

S:2013:BMR

[SPP13] Matilda S, B. Palaniappan, and Thambidurai P. Bufferbloat mitigation for real-time video streaming using adaptive controlled delay mechanism. International Journal of Computer Applications, 63(??):1–6, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume63/number20/10579-5747/.

Solanky:2016:RPT

[SPP16] Jigna Solanky, Kajal Patil, and Gayatri Patel. Resemblance of PhoneGap and titanium for mobile application development. International Journal of Computer Applications, 144(??):10–19, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number10/25214-2016910445/.

Sreedhar:2012:ERM

[SPSR12] B. R. Sreedhar, K. Pushpanjali, G. Y. Sagar, and Y. R. Reddy. Estimation of the reliability measures of a threecomponent system with human errors and common cause failures. International Journal of Computer Applications, 49(??):22–28, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number18/7728-1155/.

Sinha:2015:EOE

[SPW15] Rashmi Sinha, Shweta Parashar, and Pawan Whig. Effect of output error on fuzzy interface for VDRC of second order systems. International Journal of Computer Applications, 125(??):11–16, September 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume125/number13/22491-2015906161/.

Smyrnaiou:2017:CSS

[SPXT17] G. P. Smyrnaiou, M. Papoutsidakis, A. Xatzopoulos, and D. Tseles. Control of SIMO systems in simulation: the challenge of the multiple axes actuating pneumatic arm. International Journal of Computer Applications, 168(??):1–7, June 2017. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume168/number10/
27908-2017914499/.

Shorman:2018:PID

[SQ18] Amaal Shorman and Mohammad Qatawneh. Performance improvement of Double Data Encryption Standard algorithm using parallel computation. International Journal of Computer Applications, 179(??):1–6, March 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number25/29085-2018916527/.

Sallal:2013:GVP

[SR13a] Kawther Abbas Sallal and Abdul Monem Saleh Rahma. Generation of video panorama system. International Journal of Computer Applications, 73(??):20–26, July 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume73/number5/12737-9614/.

Sharma:2013:MRA

[SR13b] G. V. K. Sharma and K. Raja Rajeswari. MIMO radar ambiguity optimization using phase coded pulse waveforms. International Journal of Computer Applications, 61(??): 18–23, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number10/9964-4615/.

Sultana:2014:RAC

[SR14] Sayeeda Sultana and Katarzyna Radecka. Reversible architecture of computer arithmetic. International Journal of Computer Applications, 93(??):6–14, May 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume93/number14/16281-5852/.

Saleh:2015:REA

[SR15a] Muna Hadi Saleh and Noor Abdul Razzaq Raheem. Reducing the effect of the atmospheric disturbance on longitudinal fight control system usage PID controller. International Journal of Computer Applications, 127(??):27–31, October 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume127/number13/22791-
2015906586/.

Soni:2015:RHD

[SR15b] Trapti Soni and Narendra Rathor. Removal of high density impulse noise using efficient median filter for digital image. International Journal of Computer Applications, 115(??):25–31, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number5/20148-2280/.

Srinivas:2015:NMV

[SR15c] Rayudu Srinivas and K. V. Ramana. A novel method VESTAL to label lumber vertebrae and intervertebral discs. International Journal of Computer Applications, 113(??):15– 21, March 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume113/number18/
19925-2034/.

Surajudeen:2017:BDS

[SR17] Abdulrasaq Surajudeen and Zwiggelaar Reyer. Breast density segmentation based on fusion of super pixels and watershed transform. International Journal of Computer Applications, 161(??):1–7, March 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume161/number12/27197-2017913208/.

Sembiring:2023:WIF

[SR23] Richa Retno Rahmadhani Sembiring and Imam Riadi. Wireless implementation on fake Wifi using network forensic development life cycle method. International Journal of Computer Applications, 185(??):1–8, August 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number26/32849-2023922042/.

Skandar:2015:EDR

[SRA15] Arfa Skandar, Mariam Rehman, and Maria Anjum. An efficient duplication record detection algorithm for data cleansing. International Journal of Computer Applications, 127(??):28–37, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number6/22736-2015906401/.

Singh:2012:FTA

[SRC12] Pratik Singh, Aman Raj, and Debdutta Chatterjee. Flood tolerant AODV protocol (FT-AODV). International Journal of Computer Applications, 53(??):18–27, September 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume53/number6/8424-
2180/.

Samad:2012:TRS

[SRF12] Abdus Samad, M. Qasim Rafiq, and Omar Farooq. Two round scheduling (TRS) scheme for linearly extensible multiprocessor systems. International Journal of Computer Applications, 38(??):34–40, January 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume38/number10/4726-6910/.

Srivastava:2012:IJS

[Sri12] Riktesh Srivastava. Implementation of job scheduling algorithms in a cloud computing environment for B2C electronic commerce architecture using GI/G1 queuing model. International Journal of Computer Applications, 42(??): 1–8, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number1/5654-7541/.

Singh:2010:IEA

[SRK10] Yogesh Singh, B. V. R. Reddy, and R. Rama Kishore. Incremental error analysis of 3D polygonal model through MAYA API. International Journal of Computer Applications, 8(??):22–27, October 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume8/number1/1178-1617/.

Soulemane:2012:CHW

[SRM12] Moumie Soulemane, Mohammad Rafiuzzaman, and Hasan Mahmud. Crawling the hidden web: an approach to dynamic Web indexing. International Journal of Computer Applications, 55(??):7–15, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume55/number1/8717-7290/.

Singh:2013:DGS

[SRM13] Gurpreet Singh, Rajni, and Anupma Marwaha. Design of G-shaped defected ground structure for bandwidth enhancement. International Journal of Computer Applications, 75 (??):7–11, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number9/13137-0516/.

Srinivasan:2012:CIS

[SRS12] E. M. Srinivasan, K. Ramar, and A. Suruliandi. Color image segmentation using fuzzy local texture patterns. International Journal of Computer Applications, 41(??):16– 23, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume41/
number18/5641-8021/.

Shukla:2014:FFO

[SRS14a] Annpurna Shukla, Ravimohan, and Atulika Shukla. Fine frequency offset based section synchronization in cognitive radio systems. International Journal of Computer Applications, 108(??):38–43, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number17/19007-0560/.

Singh:2014:OCP

[SRS14b] Baldeep Singh, Arvind Rajput, and Sabhyata. Optimized conditional privacy preservation protocol for NFC applications using genetic algorithm. International Journal of Computer Applications, 102(??):7–11, September 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume102/number9/17841-8733/.

Satish:2014:PET

[SRSV14] G. Naga Satish, Ch. V. Raghavendran, M. V. Rama Sundari, and P. Suresh Varma. Performance evaluation of threenode tandem communication network model with feedback for first two nodes having non homogeneous Poisson arrivals. International Journal of Computer Applications, 98 (??):26–36, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number12/17236-7564/.

Saoud:2011:MPN

[SRTB11] Lyes Saad Saoud, Faycal Rahmoune, Victor Tourtchine, and Kamel Baddari. Modeling pH neutralization process using fuzzy dynamic neural units approaches. International Journal of Computer Applications, 28(??):22– 29, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume28/
number4/3375-4666/.

Saini:2010:DDM

[SS10a] Shiwani Saini and C. R. Suri. Design and development of microcontroller based auto-flow assembly for biosensor application. International Journal of Computer Applications, 6(??):21–27, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume6/number1/1047-1359/.

Saleena:2010:NAD

[SS10b] B. Saleena and S. K. Srivatsa. A novel approach to develop a self-organized domain specific search mechanism for knowledge acquisition using ontology. International Journal of Computer Applications, 12(??):1–4, December 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume12/number4/1670-2251/.

Saroha:2010:VLS

[SS10c] Kriti Saroha and Pradeep Kumar Singh. A variant of LSB steganography for hiding images in audio. International Journal of Computer Applications, 11(??):12– 16, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number6/1588-2128/.

Shaik:2010:PCA

[SS10d] Syed Basha Shaik and S. P. Setty. Performance comparison of AODV, DSR and ANODR for grid placement model. International Journal of Computer Applications, 11 (??):6–9, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume11/
number12/1637-2201/.

Sharma:2010:PAS

[SS10e] Ashutosh Sharma and Poonam Sinha. Performance analysis of space time trellis codes over Nakagami fading & Rayleigh fading channels: with comparative analysis. International Journal of Computer Applications, 7(??):19–26, October 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume7/number7/1264-
1684/.

Stuwart:2010:IPU

[SS10f] S. Lenty Stuwart and T. C. Sabareeswari. Identification of a person using multimodal biometric system. International Journal of Computer Applications, 3(??):12–16, July 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume3/number9/769-1077/.

S:2011:ESF

[SS11] Arun K S and Sarath K S. Evaluation of SUSAN filter for the identification of micro calcification. International Journal of Computer Applications, 15(??):41– 45, February 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume15/
number3/1924-2569/.

S:2012:OIS

[SS12a] Abirami S and Murugappan S. Online information search from Tamil document images in World Wide Web. International Journal of Computer Applications, 51(??):31– 39, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume51/
number5/8039-1350/.

Sasikumar:2012:GBA

[SS12b] S. Sasikumar and S. Singaravelu. GeneticAlgorithm based analysis of six-phase self-excited induction generators for wind energy conversion. International Journal of Computer Applications, 41(??):34–41, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume41/number12/5596-7842/.

Saudi:2012:PPM

[SS12c] Azali Saudi and Jumat Sulaiman. Path planning for mobile robot using 4EGSOR via nine-point Laplacian (4EGSOR9L) iterative method. International Journal of Computer Applications, 53(??):38–42, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume53/number16/8509-2568/.

Shanmugasundar:2012:SDI

[SS12d] G. Shanmugasundar and R. Sivaramakrishnan. Software development for an inverse kinematics of seven- degrees of freedom newly designed articulated inspection robot. International Journal of Computer Applications, 58(??):29– 36, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume58/
number18/9384-3858/.

Singh:2012:FFD

[SS12e] Ashutosh K. Singh and Harshit Singh. Forest fire detection through wireless sensor network using Type-2 fuzzy system. International Journal of Computer Applications, 52 (??):19–23, August 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume52/
number9/8230-1315/.

Som:2012:DSW

[SS12f] Sukalyan Som and Moumita Som. DNA secret writing with Laplace transform. International Journal of Computer Applications, 50(??):44–50, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number5/7771-0853/.

Sridevi:2012:STL

[SS12g] N. Sridevi and P. Subashini. Segmentation of text lines and characters in ancient Tamil script documents using computational intelligence techniques. International Journal of Computer Applications, 52(??):7–12, August 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume52/number14/8268-1826/.

Sachdeva:2013:MEB

[SS13a] Gagan Sachdeva and Dilpal Singh. Migration effects on BBO evolution in optimizing fifteen element Yagi–Uda antenna design. International Journal of Computer Applications, 68(??):1–5, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number18/11676-7346/.

Saparkhojayev:2013:IMH

[SS13b] Nurbek Saparkhojayev and Olzhas Shaiken. An implementation of MD5 hash algorithm for RFID tags. International Journal of Computer Applications, 72(??): 28–34, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number22/12674-9352/.

Sehgal:2013:ECI

[SS13c] Parul Sehgal and Vijay Kumar Sharma. Eliminating cover image requirement in discrete wavelet transform based digital image steganography. International Journal of Computer Applications, 68(??):37–42, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number3/11562-6848/.

Sharma:2013:OPF

[SS13d] Nidhi Sharma and Atul Srivastava. Ontological paradigm for focused crawling based on lexical analysis. International Journal of Computer Applications, 62(??):32– 36, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume62/
number6/10086-4715/.

Singh:2013:WNL

[SS13e] Reena Singh and V. K. Srivastava. Weighted non-linear diffusion filtering with wavelet thresholding in image denoising. International Journal of Computer Applications, 78 (??):1–6, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume78/
number14/13588-1318/.

Swaroop:2013:HAS

[SS13f] Abhishek Swaroop and Awadhesh Kumar Singh. A hybrid algorithm to solve group mutual exclusion problem in message passing distributed systems. International Journal of Computer Applications, 67(??):51–59, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume67/number9/11426-6777/.

Sagar:2014:EEM

[SS14a] Harish Kumar Sagar and Varsha Sharma. Error evaluation on K- means and hierarchical clustering with effect of distance functions for iris dataset. International Journal of Computer Applications, 86(??):1–5, January 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume86/number16/15066-3429/.

Sahu:2014:CIY

[SS14b] Hemlal Sahu and B. K. Sharma. Cryptanalysis and improvement of Yanlin and Xiaoping’s signature scheme based on ECDLP and factoring. International Journal of Computer Applications, 85(??):39–41, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume85/number3/14825-3065/.

Salamy:2014:ERP

[SS14c] Hassan Salamy and Olalekan Sopeju. An effective resource partitioning heuristic for embedded applications on an MP-SoC. International Journal of Computer Applications, 97 (??):1–8, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number10/17040-7350/.

Shao:2014:EOS

[SS14d] Deo Shao and Hassan Seif. Exploitation of online social networks (OSNs) among university students: a case study of the university of dodoma. International Journal of Computer Applications, 94(??):10–14, May 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume94/number12/16394-6012/.

Sharma:2014:IMD

[SS14e] Anil Sharma and Manu Sood. Incorporating MDA to design business intelligence services by using SaaS model of cloud computing. International Journal of Computer Applications, 106(??):1–8, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number16/18601-9876/.

Sejpal:2015:OPQ

[SS15a] Shveti Sejpal and Nikesh Shah. Optimized perceptual quality, robust and highly secured randomly scrambled image watermarking using 3-level DWT domain. International Journal of Computer Applications, 132(??): 30–39, December 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume132/number6/23600-2015907515/.

Singh:2015:FDM

[SS15b] Parvinder Singh and Mandeep Singh. Fraud detection by monitoring customer behavior and activities. International Journal of Computer Applications, 111(??):23–32, February 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume111/number11/19584-
1340/.

Singh:2015:MCE

[SS15c] Ram Govind Singh and Akhil Pratap Singh. Multiple complex extreme learning machine using holomorphic mapping for prediction of wind power generation system. International Journal of Computer Applications, 123(??):24–33, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume123/number18/22060-2015905735/.

Sahu:2016:ESC

[SS16a] Rani Sahu and Shailendra Kumar Shrivastava. An efficient source code auditing using fuzzy decision tree. International Journal of Computer Applications, 156(??):19–22, December 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume156/number14/26786-
2016912549/.

Sharma:2016:IEE

[SS16b] Shivangi Sharma and Gaurav Sharma. Impact of energyefficient and eco-friendly green computing. International Journal of Computer Applications, 143(??):20–28, June 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume143/number7/25089-2016910250/

.

Singh:2016:HHD

[SS16c] Tejkanwar Singh and Mandeep Singh. Hand held device for detection of pesticides using NDVI. International Journal of Computer Applications, 154(??):29– 35, November 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume154/number9/26521-2016912226/.

Suman:2016:MDA

[SS16d] Manish Kumar Suman and Sini Shibu. MAODV: a defense against Sybil attacks in wireless networks. International Journal of Computer Applications, 135(??):8–16, February 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume135/number12/24099-
2016908488/.

Sagheer:2017:HIS

[SS17a] Dania Sagheer and Fadel Sukkar. A hybrid intelligent system for abstractive summarization. International Journal of Computer Applications, 168(??):37–44, June 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume168/number9/27904-2017914505/.

Shapoorifard:2017:IDU

[SS17b] Hossein Shapoorifard and Pirooz Shamsinejad. Intrusion detection using a novel hybrid method incorporating an improved KNN. International Journal of Computer Applications, 173(??):5–9, September 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume173/number1/28297-2017914340/.

Suri:2017:ROC

[SS17c] Ashish Suri and Jaswinder Singh. Receiver operating characteristic for variable threshold and sample values using energy detection for secondary user in cognitive environment. International Journal of Computer Applications, 159(??):25–35, February 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume159/number5/26998-2017912935/.

SchrankoDeOliveira:2021:BMD

[SS21] Angelo Schranko De Oliveira and Renato Jose Sassi. Behavioral malware detection using deep graph convolutional neural networks. International Journal of Computer Applications, 174(??):1–8, April 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number29/31858-2021921218/.

Sprankle:2023:DAT

[SS23a] John-Tyler Sprankle and Selvanayaki Kolandapalayam Shanmugam. A design of assistive technology application for visually disabled elderly citizens. International Journal of Computer Applications, 185(??):18–21, May 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number8/32721-2023922734/.

Swapna:2023:CSD

[SS23b] S. L. Swapna and V. Saravanan. Chi-square detective ensembled cardinal gradient bootstrap aggregating classifier for secured big data communication. International Journal of Computer Applications, 185(??):1–8, August 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number32/32897-2023923078/.

Salem:2011:MCM

[SSA11] Dina Ahmed Salem, Rania Ahmed Abul Seoud, and Hesham Arafat Ali. MGS-CM: a multiple scoring gene selection technique for cancer classification using microarrays. International Journal of Computer Applications, 36(??):30– 37, December 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume36/
number6/4498-6349/.

Sharma:2014:GGB

[SSA14] Mradula Sharma, Kunal Sain, and Suneeta Agarwal. GES: a group based encoding of shares for visual cryptography without pixel expansion. International Journal of Computer Applications, 91(??):1–7, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume91/number3/15858-4801/.

Singh:2016:RTP

[SSA16] Kajal Singh, Divya Sharma, and Shipra Aggarwal. A real time patient monitoring system based on artificial neural fuzzy inference system (ANFIS). International Journal of Computer Applications, 146(??):22–28, July 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume146/number15/25475-2016910959/.

Sayson:2023:SWD

[SSA^{+}23] Butz M. Sayson, Kenneth N. Samson, Jerome S. Austria, Jackylyn I. Ferrer, and Maynard Gel F. Carse. Smart wearable device for physical distance and location monitoring. International Journal of Computer Applications, 185(??):24–29, May 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number10/32737-2023922770/.

Sagvekar:2010:MCL

[SSAK10] Kalpana Sagvekar, Tanuja K. Sarode, Archana Athawale, and H. B. Kekre. Mixing codebooks of LBG, KPE and KFCG algorithms to increase capacity of information hiding. International Journal of Computer Applications, 5(??):15–20, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume5/number3/898-1273/.

Shaker:2013:TCM

[SSAZ13] Ahmed Shaker, Gihan T. Sayah, Mohamed Abouelatta, and Abdelhalim Zekry. Thyristor compact modeling based on Gummel–Poon model including parameter extraction procedure. International Journal of Computer Applications, 61 (??):12–20, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number16/10010-4904/.

Singhal:2011:NAF

[SSC11] Manoj Singhal, Girish Sharma, and R. K. Chauhan. A new approach for finding the various optimal variable ordering to generate the binary decision diagrams (BDD) of a computer communication network. International Journal of Computer Applications, 31(??):1–8, October 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume31/number3/3801-2502/.

Shukla:2015:IRA

[SSC15a] Rajesh Shukla, Sanjay Silakari, and P K Chande. An improved recommendation approach based on user’s standard and interests. International Journal of Computer Applications, 119(??):6–14, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number1/21029-2840/.

Syal:2015:RCS

[SSC15b] Neeraj Syal, Niti Syal, and Urvashi Chaudhary. A review on CT scan screening for early lung cancer detection in India. International Journal of Computer Applications, 114(??):17–20, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume114/number13/20038-1748/.

Sethi:2013:NHP

[SSD^{+}13] Manan Sethi, Karna Sharma, Paanshul Dobriyal, Navya Rajput, and Geetanjali Sharma. A novel high performance dual threshold voltage domino logic employing stacked transistors. International Journal of Computer Applications, 77 (??):30–35, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume77/
number5/13393-1035/.

Singh:2011:BBC

[SSG11] Manwinder Singh, Maninder Singh, and Anudeep Goraya. Block based channel estimation algorithms for OFDM-IEEE 802.16e (mobile WiMAX) system. International Journal of Computer Applications, 13(??):33–36, January 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume13/number3/1759-
2405/.

S:2013:NRF

[SSG13] Karthik S, Snehanshu Saha, and Chaithra G. A new relevance feedback based approach for efficient image retrieval. International Journal of Computer Applications, 61 (??):1–6, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number14/9993-4846/.

Singh:2014:DAH

[SSG14] Jatinder Pal Singh, Amandeep Singh, and Nancy Gupta. Design analysis of a hexaband slot loaded microstrip patch antenna using ANN. International Journal of Computer Applications, 101(??):7–12, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number14/17753-8772/.

Sunny:2012:DWT

[SSJ12] Sonia Sunny, David Peter S, and K Poulose Jacob. Discrete wavelet transforms and artificial neural networks for recognition of isolated spoken words. International Journal of Computer Applications, 38(??):9–13, January 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume38/number9/4714-6871/.

Santhi:2010:NPP

[SSK^{+}10] M. Santhi, Arun Kumar S, G. S. Praveen Kalish, Siddharth Sarangan, and G. Lakshminarayanan. A novel pseudo 4 phase dual rail asynchronous protocol with self reset logic & multiple reset. International Journal of Computer Applications, 1(??):17–21, February 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume1/number28/518-837/.

Shapii:2011:DPP

[SSK11] Azrulhizam Shapi’i, Riza Sulaiman, and Abd Yazid Mohd Kassim. Digital preoperative planning for total hip replacement using two dimensional X-ray imaging. International Journal of Computer Applications, 17(??):20– 27, March 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume17/
number2/2193-2784/.

Sharma:2013:AEB

[SSK13a] Anjali Sharma, Richa Singh, and Pankaj Kajla. Area efficient 1-bit comparator design by using hybridized full adder module based on PTL and GDI logic. International Journal of Computer Applications, 82(??):5–13, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume82/
number10/14150-2316/.

Sivia:2013:DSC

[SSK13b] Jagtar Singh Sivia, Amarpartap Singh, and Tara Singh Kamal. Design of Sierpinski carpet fractal antenna using artificial neural networks. International Journal of Computer Applications, 68(??):5–10, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume68/number8/11597-6955/.

Sahlin:2012:OQD

[SSM12] John P. Sahlin, Shahram Sarkani, and Thomas Mazzuchi. Optimizing QoSin distributed systems/cloud computing architectures. International Journal of Computer Applications, 42(??):14–20, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number18/5792-8097/.

Sharma:2013:AEC

[SSMP13] Gaurav Sharma, Vaishali Sahu, Prashant Kumar Maurya, and Rajeev Paulus. Analyzing the effect of constant and lognormal shadowing model on ad-hoc routing protocols. International Journal of Computer Applications, 66 (??):16–19, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number8/11104-5512/.

Salim:2018:HMF

[SSN18] D. Jayus Nor Salim, Jatmiko Endro Suseno, and Oky Dwi Nurhayati. Hybrid method of first come first served and priority queue for queue system in hospital. International Journal of Computer Applications, 182(??):15–22, August 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number10/29855-
2018917713/.

S:2014:AFV

[SSO14] Adewale Olumide S, Ogundele Oloruntoba S, and Iyare Otasowie. An analytical framework for vision testing in driving license allocation in Nigeria. International Journal of Computer Applications, 97(??):34–38, July 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume97/number4/16998-7156/.

Sarangi:2013:HDE

[SSP13a] Partha Pratim Sarangi, Abhimanyu Sahu, and Madhumita Panda. A hybrid differential evolution and backpropagation algorithm for feedforward neural network training. International Journal of Computer Applications, 84 (??):1–9, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number14/14641-2943/.

Singh:2013:WUM

[SSP13b] Arun Kumar Singh, Dheeraj Sharma, and Avinav Pathak. Web usage mining: a concise survey on tools and applications. International Journal of Computer Applications, 74(??):1–7, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume74/
number1/12846-9076/.

Shah:2015:DBP

[SSP15a] Hardik A. Shah, Satish K. Shah, and Viranchi C. Pandya. DSP based pulse generation for induction motor speed control. International Journal of Computer Applications, 122(??):1–5, July 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume122/number14/21765-5014/.

Singh:2015:ESR

[SSP15b] Nidhi Singh, Hari Mohan Singh, and Shiv Pratap Pal. An enhanced service ring protocol for service discovery in MANET. International Journal of Computer Applications, 120(??):31–37, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number24/21411-4448/.

Shrivastava:2014:RIN

[SSR14] Priyanka Shrivastava, Uday Pratap Singh, and Vineet Richhariya. Removal of impulse noise using first order neighborhood mean filter. International Journal of Computer Applications, 87(??):34–39, February 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume87/number4/15199-3585/.

Subhas:2012:CAP

[SSRD12] S. Subhas, V. F. Saji, S. Ramakrishna, and H. N. Das. CFD analysis of a propeller flow and cavitation. International Journal of Computer Applications, 55(??):26– 33, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume55/
number16/8841-3125/.

Singh:2012:SBA

[SSS12] Gurjeevan Singh, Ashwani Kr. Singla, and K. S. Sandha. Superiority of Blowfish algorithm in wireless networks.

International Journal of Computer Applications, 44(??): 23–26, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume44/
number11/6308-8632/.

Sarumathi:2013:CAD

[SSS13a] S. Sarumathi, N. Shanthi, and M. Sharmila. A comparative analysis of different categorical data clustering ensemble methods in data mining. International Journal of Computer Applications, 81(??):46–55, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume81/number4/14004-2050/.

Sharma:2013:NZK

[SSS13b] B. K. Sharma, Hemlal Sahu, and Neetu Sharma. A new zero knowledge identification scheme based on Weil pairing. International Journal of Computer Applications, 83(??): 34–37, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number2/14423-2556/.

Singh:2013:PED

[SSS13c] Richa Singh, Anjali Sharma, and Rohit Singh. Power efficient design of multiplexer based compressor using adiabatic logic. International Journal of Computer Applications, 81(??):45–50, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume81/number10/14052-2217/.

Saxena:2015:MAR

[SSS15a] Sanjay Saxena, Neeraj Sharma, and Shiru Sharma. Multithreaded approach for registration of medical images using mutual information in multicore environment and its applications in medical imaging. International Journal of Computer Applications, 113(??):23–32, March 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume113/number3/19807-1598/.

Shah:2015:CMG

[SSS15b] Dharin Shah, Chirag Sachdev, and Bhavik Shah. Classification of music genre using neural networks with cross-entropy optimization and soft-max output. International Journal of Computer Applications, 119(??):33–38, June 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume119/number12/21123-4013/.

Singh:2016:CHA

[SSS16] Rupinder Singh, Jatinder Singh, and Ravinder Singh. Cluster head authentication technique against hello flood attack in wireless sensor networks. International Journal of Computer Applications, 156(??):43–49, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume156/number4/26701-2016912423/.

Shirazi:2012:HSS

[SSSS12] Syed Fahad Shirazi, Syed Hamad Shirazi, Syed Muslim Shah, and Muhammad Khalil Shahid. Hybrid spectrum sensing algorithm for cognitive radio network. International Journal of Computer Applications, 45(??):25–30, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume45/number17/7003-9557/.

S:2012:ADM

[SsTM12] Hagar Fady S, Taha El sayed T, and Mervat Mahmoud M. Applying data mining technique for the optimal usage of neonatal incubator. International Journal of Computer Applications, 52(??):11–20, August 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume52/number3/8181-1508/.

Sood:2010:BPO

[SSU10] Neetu Sood, Ajay K. Sharma, and Moin Uddin. BER performance of OFDM-BPSK and -QPSK over generalized gamma fading channel. International Journal of Computer Applications, 3(??):13–16, June 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume3/number6/736-1035/.

Sharma:2011:AAM

[SSV11] Pranjali Sharma, Amit Sharma, and R. K. Vats. Analysis of adjacency matrix and neighborhood associated with zero divisor graph of finite commutative rings. International Journal of Computer Applications, 14(??):38–42, January 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume14/number3/1822-

Singh:2014:NAD

[SSV14] Tripty Singh, Sanju S, and Bichu Vijay. A new algorithm designing for detection of moving objects in video. International Journal of Computer Applications, 96(??): 4–11, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume96/
number2/16764-6324/.

Syaiin:2011:NAN

[SSY11] Mat Syai’in, Adi Soeprijanto, and Eko Mulyanto Yuniarno. New algorithm for neural network optimal power flow (NN-OPF) including generator capability curve constraint and statistic-fuzzy load clustering. International Journal of Computer Applications, 36(??):1–8, December 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume36/number7/4500-6351/.

Supli:2016:CAA

[SSZ16] Ahmad Affandi Supli, Norshuhada Shiratuddin, and Syamsul Bahrin Zaibon. Critical analysis on algorithm visualization study. International Journal of Computer Applications, 150(??):18–22, September 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume150/number11/26137-2016911633/.

Shrivastava:2011:DTA

[ST11] Shailendra K. Shrivastava and Manisha Tantuway. A decision tree algorithm based on rough set theory after dimensionality reduction. International Journal of Computer Applications, 17(??):29–34, March 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume17/number7/2231-2846/.

Samuel:2012:RPC

[ST12] Ibiyemi Tunji Samuel and Ajiboye Aye Taiwo. Robust PID control of mobile satellite dish network within Nigeria. International Journal of Computer Applications, 41 (??):37–42, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume41/
number21/5828-8154/.

Soman:2014:NAM

[ST14] Remya Soman and Jency Thomas. A novel approach for mixed noise removal using ‘ROR’ statistics combined WITH ACWMF and DPVM. International Journal of Computer Applications, 86(??):11–17, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume86/number17/15076-3442/.

Saleh:2016:NMG

[ST16] Muna Hadi Saleh and Hadeel Mohammed Tuama. Novel multi-gen multi parameter genetic algorithm representation for attributes selection and porosity prediction. International Journal of Computer Applications, 141(??):34–39, May 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume141/number4/24775-2016909614/

.

Stamoulis:2022:HDT

[Sta22] Dimitrios S. Stamoulis. How to digitally transform the communication of the weather forecast science using an intelligent systems approach. International Journal of Computer Applications, 184(??):31–35, June 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number17/32412-2022922177/.

Shrivastava:2014:DDT

[STD14] Sonam Shrivastava, Ravi Tiwari, and Susmita Das. Dynamicdouble-threshold energy detection scheme under noise varying environment in cognitive radio system. International Journal of Computer Applications, 87(??):23– 27, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number14/15277-3987/.

Shahsavari:2014:SFB

[STE14] H. R. Shahsavari, M. R. Taheri, and R. Esmaelzadeh. Satellite flatness based fault tolerant control. International Journal of Computer Applications, 99(??):49– 53, August 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume99/
number4/17364-7883/.

Shah:2012:ARH

[STG12] Komal Shah, Amit Thakkar, and Amit Ganatra. Association rule hiding by heuristic approach to reduce side effects and hide multiple R.H.S. items. International Journal of Computer Applications, 45(??):1–7, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume45/number1/6741-7813/.

Sharma:2013:ISM

[STG13] Pallavi Sharma, G. Thakar, and R. C. Gupta. Interpretive structural modeling of functional objectives (criteria’s) of assembly line balancing problem. International Journal of Computer Applications, 83(??):14–22, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number13/14508-2883/.

Sadananda:2013:MAF

[STM13] Priyanka Sadananda, Wassim Trojet, and Joseph Mouzna. Multicast authentication framework for hierarchical networks using Chinese Remainder Theorem. International Journal of Computer Applications, 82(??):1–7, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume82/number11/14157-2233/.

Sari:2019:IDM

[STS19] Yuli Nur Indah Sari, Agung Triayudi, and Ira Diana Sholihati. Implementation of data mining to predict food sales rate method using apriori. International Journal of Computer Applications, 178(??):22–28, July 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number35/30768-2019919228/.

Sharmila:2012:DMS

[SU12] S. Sharmila and G. Umamaheswari. Defensive mechanisms of selective forward attack in wireless sensor networks. International Journal of Computer Applications, 39(??):46– 52, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume39/
number4/4812-7048/.

Sridevi:2014:NHO

[SU14] K. Sridevi and R. Umarani. A novel and hybrid ontology ranking framework using semantic closeness measure. International Journal of Computer Applications, 87(??):44– 48, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number5/15208-3612/.

Sanghamitra:2018:IDF

[SU18] Nath Sanghamitra and Sharma Utpal. Incorporating dialectal features in synthesized speech using voice conversion techniques. International Journal of Computer Applications, 180(??):1–8, February 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number19/29037-2018916443/.

Soyusiawaty:2020:IJC

[SU20] Dewi Soyusiawaty and Indriyani Putri Utami. Implementation of Jaccard coefficient method for searching report findings of internal quality audit in Ahmad Dahlan University. International Journal of Computer Applications, 175(??):28–35, December 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number38/31702-2020920954/.

Subha:2010:OAN

[Sub10] S. Subha. An optimized algorithm for network on chip. International Journal of Computer Applications, 3(??):23–25, July 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume3/number12/789-
1118/.

Subhadra:2020:ADS

[Sub20] Akella Subhadra. An analysis of data security and privacy in cloud computing. International Journal of Computer Applications, 176(??):31–40, July 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number39/31461-2020920508/.

Subramaniam:2024:LIV

[Sub24] Nantha Kumar Subramaniam. Leveraging interactive videocentric courseware for enhancing “Assessment as learning” in programming. International Journal of Computer Applications, 186(??):43–48, August 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number34/leveraging-interactive-video-centric-
courseware-for-enhancing-assessment-as-learning-in-
programming/.

Sudha:2010:SCB

[Sud10] G. Sudha. A simple classification of binary document into vector image or scalar image using feed forward neural networks with back propagation training. International Journal of Computer Applications, 9(??):24–27, November 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume9/number3/1365-1840/

.

Sujanarko:2014:SDC

[Suj14] Bambang Sujanarko. Simulation development of carrier based pulse width modulation for cascaded multilevel inverter. International Journal of Computer Applications, 104(??):36–41, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number5/18202-9124/.

Sulaiman:2017:OZA

[Sul17] Alyaa Ghanim Sulaiman. Overview of ZUC algorithm and its contributions on the security success and vulnerabilities of 4G mobile communication. International Journal of Computer Applications, 168(??):34–38, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume168/number8/27898-2017914485/.

Sun:2019:RBS

[Sun19] Junyuan Sun. 3D-rebuild based SFM algorism improvement. International Journal of Computer Applications, 177(??):5–9, November 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume177/number19/31005-2019919545/.

Sundararaja:2010:ONP

[SV10] S. Sundararaja and V. E. Vairavan. Optimization of network packet scheduling — using multi objective genetic algorithm with virtual deadline as a schedule parameter. International Journal of Computer Applications, 1(??):61–65, February 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume1/number19/402-598/.

S:2012:TRM

[SV12a] Allirani S and Jagannathan V. Torque ripples minimization in DTC based induction motor drive using fuzzy logic technique. International Journal of Computer Applications, 40 (??):25–31, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume40/
number1/4921-7144/.

Sudhamathy:2012:EHF

[SV12b] G. Sudhamathy and C. Jothi Venkateswaran. An efficient hierarchical frequent pattern analysis approach for Web usage mining. International Journal of Computer Applications, 43(??):1–7, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume43/
number15/6176-8603/.

S:2014:IAM

[SV14a] Premananda B S and Uma B V. Incorporating auditory masking properties for speech enhancement in presence of near-end noise. International Journal of Computer Applications, 106(??):1–6, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number15/18593-9835/.

Sivakumar:2014:EML

[SV14b] S. Sivakumar and R. Venkatesan. Error minimization in localization of wireless sensor networks using modified cuckoo search with mobile anchor positioning (MCS-Map) algorithm. International Journal of Computer Applications, 95 (??):1–8, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number6/16595-6398/.

S:2018:EMN

[SV18a] Suhas S and C. R. Venugopal. An efficient MRI noise removal technique using linear and nonlinear filters. International Journal of Computer Applications, 179(??):17–20, January 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume179/number15/28875-
2018915777/.

Sirisha:2018:NME

[SV18b] D. Sirisha and G. Vijayakumari. A new makespan estimation model for scientific workflows on heterogeneous processing systems. International Journal of Computer Applications, 179(??):18–26, February 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number17/28959-2018916026/.

Sudha:2013:FDT

[SVB13] G. Sudha, K. R. Valluvan, and T. Basavaraju. Fault diagnosis of transmission lines with rogowski coils as current sensors. International Journal of Computer Applications, 70(??):19–25, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number25/12223-8447/.

Sikhwal:2017:GMC

[SVB17] Omprakash Sikhwal, Yashwant Vyas, and Shikha Bhatnagar. Generalized multiplicative coupled Fibonacci sequence and its properties. International Journal of Computer Applications, 158(??):14–17, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume158/number10/26941-2017912838/.

Sylvester:2023:TEA

[SVB23] Akpah Sylvester, Selasie Vondee, and Desmond Boahen. Tracking of excavators in the artisanal small-scale mining sector in Ghana using TX 140-4G tracker. International Journal of Computer Applications, 185(??):18–24, June 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume185/number16/32778-
2023922855/.

Sumbaly:2014:DBC

[SVJ14] Ronak Sumbaly, N. Vishnusri, and S. Jeyalatha. Diagnosis of breast cancer using decision tree data mining technique. International Journal of Computer Applications, 98 (??):16–24, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number10/17219-7456/.

Soni:2015:BCB

[SVJ15] Ashish Soni, Gagan Vishwakarma, and Yogendra Kumar Jain. A bee colony based multi-objective load balancing technique for cloud computing environment. International Journal of Computer Applications, 114(??): 19–25, March 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume114/number4/19967-1825/.

Syed:2013:SDE

[SVMLb13] Tahir Q. Syed, V. Vigneron, C. Montagne, and S. Lelandaisbonad. Suitability of digital elevation models for watershed segmenting images with directional illumination. International Journal of Computer Applications, 78(??):1– 7, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume78/
number12/13572-0946/.

Singh:2015:ITS

[SVRB15] Vikas Pratap Singh, Vivek Vijay, B. Ravindra, and M. Siddhartha Bhatt. Impact of trend and seasonality in forecasting of 5 MW PV plant generation using single exponential smoothing method. International Journal of Computer Applications, 130(??):5–9, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume130/number1/23171-2015906840/.

Sulistianingsih:2017:PAM

[SW17] Neny Sulistianingsih and Edi Winarko. Performance analysis of molecular complex detection in social network datasets. International Journal of Computer Applications, 175(??):10–15, October 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number4/28475-2017915509/.

Shum:2016:DBR

[SWH16] Andrew Shum, Yunyi Wang, and Sheng-Jen Hsieh. Design, build and remote control of a miniature automated robotic sorting system. International Journal of Computer Applications, 141(??):12–17, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume141/number3/24763-2016909576/.

Saha:2019:CSO

[SWP19] Debjyoti Saha, Pravin Wararkar, and Shashikant Patil. Comprehensive study and overview of vehicular ad-HOC networks (VANETs) in current scenario with respect to realistic vehicular environment. International Journal of Computer Applications, 178(??):26–40, May 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number15/30606-2019918910/.

Saini:2012:SCT

[SY12] Dinesh Kumar Saini and Jabar H Yousif. Soft computing techniques for mishaps prediction. International Journal of Computer Applications, 47(??):1–7, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume47/number8/7205-9989/.

Shukla:2015:VBE

[SY15] Sanyam Shukla and R. N. Yadav. Voting based extreme learning machine with accuracy based ensemble pruning. International Journal of Computer Applications, 115(??):14–18, April 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume115/number22/
20282-2837/.

Saleh:2017:PSE

[SY17] Huda M. Saleh and Hameed A. Younis. Private searching on encrypted data in cloud. International Journal of Computer Applications, 165(??):20–26, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number7/27585-2017913924/.

Saleh:2011:DDB

[SYAS11] H. Saleh, A. Yahya, M. Ashour, and M. Sayed. DST and DCT-based depth of interaction (DOI) determining techniques for LSO and LuYAP scintillation detectors in PET. International Journal of Computer Applications, 28 (??):41–45, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume28/
number7/3397-4725/.

Srivatsa:2010:RMF

[SYP10] Kopparapu Srivatsa, Madamshetti Yashwanth, and A. Parvathy. RFID & mobile fusion for authenticated ATM transaction. International Journal of Computer Applications, 3(??):5–10, June 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume3/number5/731-1025/.

Singh:2013:MTD

[SYR13] Archana Singh, Avantika Yadav, and Ajay Rana. K-means with three different distance metrics. International Journal of Computer Applications, 67(??):13–17, April 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume67/number10/11430-6785/.

Shaheryar:2017:RFE

[SYR17] Ahmad Shaheryar, Xu-Cheng Yin, and Waheed Yousuf Ramay. Robust feature extraction on vibration data under deeplearning framework: an application for fault identification in rotary machines. International Journal of Computer Applications, 167(??):37–45, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number4/27763-2017914249/.

Saberi:2015:DAG

[SZ15] Farhad Fani Saberi and Mehdi Zandieh. Design and analysis of gimbal thruster configurations for 3-axis satellite attitude control. International Journal of Computer Applications, 112(??):29–38, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number6/19673-1234/.

Sheta:2016:PME

[SZEM16] Mohamed Adel Sheta, Mohamed Zaki, Kamel Abd El Salam El Hadad, and H. Aboelseoud M. Performance and maintainability evaluation of anti-spyware system. International Journal of Computer Applications, 150(??): 31–39, September 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume150/number4/26085-2016911493/.

Saad:2018:DFH

[SZM18] Sabah Saad, Bassam A. Zafar, and Ahmed Mueen. Developing a framework for e-healthcare applications using the Semantic Internet of Things. International Journal of Computer Applications, 182(??):25–33, December 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume182/number34/30252-2018918279/.

Salkanovic:2024:DAS

[SZOK24] Emina Salkanović, Nejla ZukorliĆ, Lamija Oković, and Dino Kečo. A data analysis of steam’s game catalog and diverse recommendation strategies. International Journal of Computer Applications, 186(??):39–49, December 2024. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume186/number54/a-
data-analysis-of-steams-game-catalog-and-diverse-
recommendation-strategies/.

Sultan:2015:TPM

[SZS15] Kiran Sultan, Bassam A. Zafar, and Babar Sultan. Transmit power minimization using fuzzy rule based system in relay assisted cognitive radio networks. International Journal of Computer Applications, 130(??): 29–34, November 2015. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume130/number3/23190-2015906953/.

Soheli:2019:CPA

[SZUSNK19] Sultana Jahan Soheli, Zayed-Us-Salehin, Farjana Afroz Nila, and Md. Ashikur Rahman Khan. Comparative performance analysis of DSDV, DSR and OLSR routing protocols to determine the best suited routing approach through simulation in mobile Wimax network. International Journal of Computer Applications, 177(??):19–28, December 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number22/31029-2019919656/.

Sarhadi:2011:PTO

[SZZR11] Ali Sarhadi, Hosein Zohrevand, Ebad Zohrevandi, and Rasoul Rostaei. Proffering task oriented grid resource discovery based on learning automata. International Journal of Computer Applications, 21(??):41–46, May 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume21/number2/2480-3339/.

Tripathi:2010:PCW

[TA10] Sandesh Tripathi and S. Q. Abbas. Performance comparison of Web services under simulated and actual hosted environments. International Journal of Computer Applications, 11(??):20–23, December 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume11/number5/1578-2110/.

Tajani:2012:AKA

[TA12] Chakir Tajani and Jaafar Abouchabaka. An alternating KMF algorithm to solve the Cauchy problem for Laplace’s equation. International Journal of Computer Applications, 38 (??):30–36, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume38/
number8/4709-6876/.

Tuaimah:2013:ACO

[TAH13] Firas M. Tuaimah, Yaser Nadhum Abd, and Fahad A. Hameed. Ant colony optimization based optimal power flow analysis for the Iraqi super high voltage grid. International Journal of Computer Applications, 67(??):13– 18, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number11/11438-6700/.

Taha:2014:DRP

[Tah14] Saleem M. R. Taha. Design of reversible programmable gate array based on new reversible logic modules. International Journal of Computer Applications, 93(??): 26–31, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number10/16252-5857/.

Taher:2017:IAB

[Tah17] Zahraa Tariq Mohammed Taher. An improved artificial bee colony algorithm for linear least squares and constrained optimization problems. International Journal of Computer Applications, 171(??):22–29, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume171/number5/28178-2017915036/.

Talib:2011:IGK

[Tal11] Mohammad Talib. ICT to go kaput due to impending solar flare: Some counteractive measures. International Journal of Computer Applications, 23(??):6–11, June 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume23/number8/2909-3823/.

Toihria:2013:DEC

[TAM13] Intissar Toihria, Rim Ayadi, and Mohamed Masmoudi. Design of an effective charge pump-phase locked loops architecture for RF applications. International Journal of Computer Applications, 74(??):38–44, July 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume74/number3/12867-9740/.

Taheri:2014:AAC

[TAM14a] Behnaz Seyed Taheri, Mostafa Ghobaei Arani, and Mehrdad Maeen. ACCFLA: Access control in cloud federation using learning automata. International Journal of Computer Applications, 107(??):30–40, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number6/18758-0028/.

Tarapiah:2014:APN

[TAM14b] Saed Tarapiah, Shadi Atalla, and Ahmed Masri. Analysis the performance of network coding for ad hoc networks in realistic simulation scenarios. International Journal of Computer Applications, 85(??):13–20, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume85/number10/14875-3253/.

Tariquzzaman:2019:SFS

[TAM19] Md. Tariquzzaman, Md. Tanvir Ahammed, and Md. Moznuzzaman. Switching frequency selection technique for model predictive control based multilevel inverter. International Journal of Computer Applications, 177(??):46–50, November 2019. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume177/number18/31003-
2019919647/.

Talha:2013:BAN

[TAMA13] Umar Talha, Mir Asif, Sheraz Mohani, and Jahanzeb Ahmad. Body area networks (BANs) — an overview with smart sensors based telemedical monitoring system. International Journal of Computer Applications, 84(??):20– 23, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number8/14596-2836/.

Thakker:2015:GWM

[TAP15] Shruti Thakker, Jhummarwala Abdul, and M. B. Potdar. GeoProcessing workflow models for distributed processing frameworks. International Journal of Computer Applications, 113(??):33–38, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume113/number1/19794-1574/.

Tuaimah:2014:OCD

[TArM14] Firas M. Tuaimah, Nihad M. Al-rawi, and Waleed A. Mahmoud. Optimal controller design for a turbogenerator automatic voltage regulator and governor using two degree of freedom linear quadratic Gaussian (2DOFLQG). International Journal of Computer Applications, 85(??): 1–7, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number12/14890-3289/.

Taouil:2016:HII

[TAS^{+}16] Y. Taouil, E. B. Ameur, A. Souhar, A. El Harraj, and M. T. Belghiti. High imperceptibility image steganography methods based on HAAR DWT. International Journal of Computer Applications, 138(??):38–43, March 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume138/number10/24419-2016908984/.

Topal:2015:MBA

[TAY15] Ali Osman Topal, Oguz Altun, and Yunus Emre Yildiz. Micro bat algorithm for high dimensional optimization problems. International Journal of Computer Applications, 122(??):1–10, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume122/number12/21749-5015/.

Taherian:2012:SOI

[TB12a] Taher Taherian and Samira Bairamzadeh. Scheduling oriented improvement of Kanban card resupplying in logistic systems at automotive industry. International Journal of Computer Applications, 49(??):13–21, July 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume49/number14/7694-1022/.

Taherian:2012:SSO

[TB12b] Taher Taherian and Samira Bairamzadeh. Sequencing and scheduling oriented customer driven production planning in CONWIP systems. International Journal of Computer Applications, 41(??):18–24, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume41/number14/5608-7876/.

Thamilselvan:2012:IGA

[TB12c] R. Thamilselvan and P. Balasubramanie. Integrating genetic algorithm, tabu search and simulated annealing for job shop scheduling proble. International Journal of Computer Applications, 48(??):42–54, June 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume48/number5/7348-0283/.

Tair:2013:DEP

[TB13] Mohammed M. Abu Tair and Rebhi S. Baraka. Design and evaluation of a parallel classifier for large-scale Arabic text. International Journal of Computer Applications, 75 (??):13–20, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume75/
number3/13090-0370/.

Taruna:2014:PBC

[TB14] S. Taruna and Nidhi Bhartiya. Parameter based cluster head election in wireless sensor network: a fuzzy approach. International Journal of Computer Applications, 106(??):6–10, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number7/18530-9733/.

Trojet:2015:SRU

[TB15] Wassim Trojet and Tahar Berradia. System reliability using simulation models and formal methods. International Journal of Computer Applications, 132(??):1–8, December 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume132/number17/23683-
2015907684/.

Thabet:2012:RTD

[TBA12] A. Thabet, M. Boutayeb, and M. N. Abdelkrim. Real time dynamic state estimation for power system. International Journal of Computer Applications, 38(??):11– 18, January 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume38/
number2/4659-6755/.

Tarik:2013:IRS

[TBKA13] Ayaou Tarik, Mourad Boussaid, Afdel Karim, and Amghar Abdelah. Improving road signs detection performance by combining the features of Hough transform and texture. International Journal of Computer Applications, 73 (??):5–9, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number9/12767-8795/.

T:2013:CBR

[TBKG13] Pavan Kumar T, Ramesh Babu B, Rajasekhar Rao K, and Dinesh Gopalni. Cluster based routing protocol for cognitive radio wireless mesh networks. International Journal of Computer Applications, 80(??):27–31, October 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume80/number12/13914-1812/.

Tammi:2015:ANN

[TBN^{+}15] Wasima Matin Tammi, Noor Ahmed Biswas, Ziad Nasim, Khadizatul Zannat Shorna, and Faisal Muhammad Shah. Artificial neural network based system for intrusion detection using clustering on different feature selection. International Journal of Computer Applications, 126(??):21–28, September 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume126/number12/22604-
2015906244/.

T:2015:RAR

[TBR15] Shiva Prakash T, Raja K B, and Venugopal K R. Reliable adaptive replication routing for wireless sensor networks. International Journal of Computer Applications, 114(??):1–7, March 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume114/number17/
20067-2023/.

Thakur:2010:IAR

[TC10] Uma K. Thakur and S. A. Chhabria. Implementation analysis of RETT-gen routing protocol. International Journal of Computer Applications, 1(??):58–59, February 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume1/number17/361-548/.

Tiwari:2016:RCR

[TC16] Abhavya Tiwari and Richa Chouhan. A robust chaotic randomization for LSB image watermarking. International Journal of Computer Applications, 137(??):15–20, March 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume137/number4/24263-2016908682/

.

Tensubam:2014:CAF

[TCS14] Bidyalaxmi Devi Tensubam, Nongmaithem Lalleima Chanu, and Sonika Singh. Comparative analysis of FBMC and OFDM multicarrier techniques for wireless communication networks. International Journal of Computer Applications, 100(??):27–31, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number19/17636-8382/.

Tenruh:2019:ENT

[TCS19] Mahmut Tenruh, Periklis Charchalakis, and Elias Stipidis. Evaluating a new time-triggered CAN FD protocol with periodic messages. International Journal of Computer Applications, 178(??):1–9, September 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number48/30872-2019919383/.

Thiruvengatanadhan:2016:IRM

[TD16a] R. Thiruvengatanadhan and P. Dhanalakshmi. Indexing and retrieval of music using Gaussian mixture model techniques. International Journal of Computer Applications, 148(??):35–41, August 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume148/number3/25740-2016911095/.

Thomas:2016:VAP

[TD16b] Winnie Thomas and Rohin Daruwala. Variability analysis in the power appetite of GPGPU applications. International Journal of Computer Applications, 147(??):28–34, August 2016. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume147/number12/25708-
2016911281/.

Tawhid:2018:GRS

[TD18] Md. Nurul Ahad Tawhid and Emon Kumar Dey. A gender recognition system from facial image. International Journal of Computer Applications, 180(??):5–14, February 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume180/number23/29070-
2018915852/.

Tiwari:2021:TBM

[TD21] Pradeep Tiwari and A. D. Darji. Therapy bot: a multimodal stress/emotion recognition and alleviation system. International Journal of Computer Applications, 183(??):1–8, October 2021. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume183/number33/
32142-2021921719/.

Thenmozhi:2016:GCR

[TE16] M. Thenmozhi and T. P. Ezhilarasi. Group coupon recommendation system for mobile users. International Journal of Computer Applications, 143(??):31–36, June 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume143/number10/25115-2016910379/.

Tatane:2015:AMA

[TERCM15] Khalid Tatane, Brahim Er-Raha, Chihab-Eddine Cherkaoui, and Sanaa Mouhim. Alignment methodological approach of evolving domain sub-ontologies using terminological and structural matchers applied to tourism domain. International Journal of Computer Applications, 123(??):6–13, August 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume123/number15/
22033-2015905710/.

Tesha:2015:ITF

[Tes15] Thomas Tesha. The impact of transformed features in automating the Swahili document classification. International Journal of Computer Applications, 127(??):37–42, October 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume127/number16/22817-
2015906707/.

Taha:2011:POW

[TFT11] Mohammed Abu Taha, Mousa Farajalla, and Radwan Tahboub. A practical one way hash algorithm based on matrix multiplication. International Journal of Computer Applications, 23(??):33–37, June 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume23/number2/2859-3677/.

Thabet:2018:ODC

[TG18] Assem Thabet and Noussaiba Gasmi. Observer design for a class of nonlinear discrete time systems: Real time application to the one-link flexible joint robot. International Journal of Computer Applications, 179(??):1–6, April 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number37/29279-2018916875/.

Tanner:2014:FDI

[TGcO14] Kesler Tanner, Christophe Giraud-carrier, and Dan R. Olsen, Jr. Formatting by demonstration: an interactive machine learning approach. International Journal of Computer Applications, 86(??):41–47, January 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume86/number18/15090-3080/.

Tabar:2013:UVP

[TGL13] Mahdi Taheri Tabar, Imran Ghani, and Mohd Shafie Abd Latiff. User’s verification to prevent malicious access in Webbased forums. International Journal of Computer Applications, 63(??):11–16, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume63/number9/10493-5250/.

Tauro:2012:SBO

[TGR12] Clarence J M Tauro, N Ganesan, and Anjan Kumar Rs. A study of benefits in object based storage systems. International Journal of Computer Applications, 42 (??):28–30, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number8/5715-7773/.

Thaher:2014:SVD

[TH14] Raad H. Thaher and Zaid K. Hussein. Stereo vision distance estimation employing SAD with Canny edge detector. International Journal of Computer Applications, 107(??):38–43, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number3/18735-9977/.

Thakur:2015:EVV

[TH15] Parash Moni Thakur and G. C. Hazarika. Effects of variable viscosity and thermal conductivity on unsteady free convective heat and mass transfer MHD flow of micropolar fluid with constant heat flux through a porous medium. International Journal of Computer Applications, 110(??):20–30, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number8/19338-0900/.

Thayananthan:2011:AQC

[Tha11] Vijey Thayananthan. Analysis of quantum computing and trellis coding based on PUM codes. International Journal of Computer Applications, 30(??):8–12, September 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume30/number5/3640-
5084/.

Thakur:2013:NSB

[Tha13] Shivani Thakur. Node scheduling based on grid optimization for enhancing packet delivery in WSN. International Journal of Computer Applications, 83(??):7– 10, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume83/
number6/14450-2708/.

Thaher:2014:HSE

[Tha14] Raad H. Thaher. A high spectral efficient non-binary TTCM-assisted G2 STBC-OFDM for 4G systems. International Journal of Computer Applications, 93(??): 17–22, May 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume93/
number20/16544-5599/.

Twum:2019:PNF

[THAP19] Frimpong Twum, J. B. Hayfron-Acquah, and J. K. Panford. A proposed new framework for securing cloud data on multiple infrastructures using erasure coding, dispersal technique and encryption. International Journal of Computer Applications, 181(??):38–49, April 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number50/30502-2019918681/.

Talaat:2011:RTT

[THEE11] Ahmad H. Talaat, Nael A. Hussein, Hussein A. Elsayed, and Hadia Elhennawy. Real-time traffic performance for WiFi handovers over MIPv4 versus MIPv6. International Journal of Computer Applications, 26(??):10–15, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume26/number8/3127-4303/.

Thant:2011:FTM

[THT11] Win Win Thant, Tin Myat Htwe, and Ni Lar Thein. Function tagging for Myanmar language. International Journal of Computer Applications, 26(??):34–41, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume26/number2/3074-4204/.

Tabassum:2015:IFU

[TIA15] Fahima Tabassum, Md. Imdadul Islam, and M. R. Amin. Identification of fingerprint using discrete wavelet packet transform. International Journal of Computer Applications, 128(??):38–44, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume128/number7/22889-2015906613/.

Tithi:2014:OSB

[TIT14] Fahmida Hossain Tithi, M. S. Islam, and Md. Tawhidul Anwar Tanna. Overview of stimulated Brillouin scattering effect and various types of method to eliminate this effect. International Journal of Computer Applications, 92 (??):23–26, April 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume92/
number7/16022-5037/.

Titus:2016:PPU

[Tit16] Carina Titus. A proposed prototype on using online social networks as learning platforms. International Journal of Computer Applications, 145(??):19–26, July 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume145/number9/25307-2016910760/.

Thilagam:2012:AEH

[TJ12] K. Thilagam and K. Jayanthi. Analytical evaluation of a hybrid chaotic shift keying [CS-QCSK] in a CDMA system. International Journal of Computer Applications, 54(??): 38–43, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume54/
number15/8645-2544/.

Taheri:2014:EEA

[TK14] M. Taheri and Yousef S. Kavian. Energy efficient algorithm for wireless sensor networks using fuzzy logic. International Journal of Computer Applications, 89(??): 1–5, March 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume89/
number14/15696-4055/.

Thepade:2015:VCU

[TK15] Sudeep D. Thepade and Madhura M. Kalbhor. Video classification using sine, cosine, and Walsh transform with Bayes, function, lazy, rule and tree data mining classifier. International Journal of Computer Applications, 110(??):18–23, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number3/19296-0732/.

Thorat:2017:RBK

[TK^{+}17] S. S. Thorat, K. V. Kale, , S. C. Mehrotra, Y. D. Rajendra, and V. B. Waghmare. Rule base knowledge and fuzzy approach for classification of specific crop and acreage estimation. International Journal of Computer Applications, 165(??):38–47, May 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume165/number6/27581-2017913900/.

Tanvee:2013:MMA

[TKC^{+}13] Moin Mahmud Tanvee, Shaikh Jeeshan Kabeer, Tareque Mohmud Chowdhury, Asif Ahmed Sarja, and Md. Tayeb Hasan Shuvo. Mining maximal adjacent frequent patterns from DNA sequences using location information. International Journal of Computer Applications, 76(??):26–32, August 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume76/number15/13322-0819/.

Trivedi:2015:PER

[TKR15] Neha Trivedi, G. Kumar, and Teena Raikwar. Performance and evolution of routing protocol DSR, AODV and AOMDV in MANET. International Journal of Computer Applications, 109(??):1–8, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume109/number8/19205-0876/.

Tanbakuchi:2016:VTU

[TL16] Mahdi Tanbakuchi and Mojtaba Lotfizad. Visual tracking using corner based centrist descriptor with a robust localization algorithm. International Journal of Computer Applications, 153(??):1–11, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume153/number6/26404-2016912072/.

Tripathi:2013:TPI

[TM13] Sudhakar Tripathi and R. B. Mishra. Two phase integrated rule based model (TPC-IRBM) for clustering of gene expression data of CA1 region of rat hippocampus. International Journal of Computer Applications, 84(??):23– 29, December 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume84/
number6/14580-2803/.

Tripathi:2018:FOT

[TM18] Mukesh Kumar Tripathi and Dhananjay D. Maktedar. A framework with OTSU’S thresholding method for fruits and vegetables image segmentation. International Journal of Computer Applications, 179(??):25–32, June 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume179/number52/29533-2018917336/.

Tarish:2017:WSN

[TMC17] Hiba A. Tarish, Mahmoud S. Mahdi, and Muayad Sadik Croock. Wireless sensor network structure for ground water well’s field in karbala city. International Journal of Computer Applications, 159(??):5–8, February 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume159/number9/27027-2017912926/.

Thepade:2010:IRS

[TMK^{+}10] Sudeep Thepade, Priyadarshini Mukherjee, Miti Kakaiya, Shobhit Wadhwa, and Satyajit Singh. Image retrieval with shape features extracted using morphological operators with BTC. International Journal of Computer Applications, 12 (??):1–5, December 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume12/
number3/1662-2238/.

Thethi:2011:ESI

[TMP11] H. Pal Thethi, Babita Majhi, and G. Panda. Efficient system identification using a low complexity nonlinear network with differential evolution and its variant based training schemes. International Journal of Computer Applications, 31(??):38–46, October 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume31/number8/3848-5350/.

Thakur:2015:CDA

[TMS15] Yashika Thakur, Rajesh Mehra, and Anjali Sharma. CMOS design of area and power efficient multiplexer using tree topology. International Journal of Computer Applications, 112(??):32–36, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number11/19714-1493/.

Taha:2014:NQR

[TN14a] Saleem M. R. Taha and Abbas K. Nawar. A new quantum radial wavelet neural network model applied to analysis and classification of EEG signals. International Journal of Computer Applications, 85(??):7–11, January 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume85/number7/14851-3216/.

Toloei:2014:SET

[TN14b] Alireza Toloei and Saeid Niazi. State estimation for target tracking problems with nonlinear Kalman filter algorithms. International Journal of Computer Applications, 98 (??):30–36, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number17/17277-7708/.

Thampatty:2010:RAB

[TNC10] K. C. Sindhu Thampatty, M. P. Nandakumar, and Elizabeth P. Cheriyan. RTRL algorithm based adaptive controller for non-linear multivariable systems. International Journal of Computer Applications, 1(??):94–101, February 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume1/number5/115-230/.

Timilsina:2015:GES

[TNKS15] Sandesh Timilsina, Rohit Negi, Yashika Khurana, and Jyotsna Seth. Genetically evolved solution to timetable scheduling problem. International Journal of Computer Applications, 114(??):12–17, March 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume114/number18/20077-2100/.

Tigani:2013:MAS

[TO13] Smail Tigani and Mouhamed Ouzzif. Multi-agent system design for synchronizing software components communication: Application on orchestration in complex SOA architecture. International Journal of Computer Applications, 78 (??):17–20, September 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume78/
number4/13477-1170/.

Tchao:2013:PCS

[TOD13] E. T. Tchao, W. K. Ofosu, and K. Diawuo. A performance case study of electronic governance implementation in Ghana with WiMAX technology. International Journal of Computer Applications, 71(??):1–5, June 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume71/number14/12423-8972/.

Tigani:2014:LCA

[TOHS14] Smail Tigani, Mouhamed Ouzzif, Abderrahim Hasbi, and Rachid Saadane. Low complexity algorithm for probability density estimation applied in big data analysis. International Journal of Computer Applications, 101(??): 1–5, September 2014. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume101/number7/17696-8650/.

Tapia-o:2013:APC

[TOMGBSS13] Ruben Tapia-O., Omar Aguilar M, Abel Garcia-B., and Omar J. Santos-S. Adaptive PI controllers for doubly fed induction generator using B-spline artificial neural networks. International Journal of Computer Applications, 80 (??):37–42, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume80/
number16/13956-1949/.

Tuyaa:2016:EWT

[TOO16] Koriata P. Tuyaa and W. Okelo-Odongo. Enhanced weighted trust evaluation scheme for detection of malicious nodes in wireless sensor networks. International Journal of Computer Applications, 155(??):32–38, December 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume155/number4/26595-2016912298/.

Teymouri:2012:CDA

[TR12] Sanaz Teymouri and Amir Masoud Rahmani. A continues double auction method for resource allocation in economic grids. International Journal of Computer Applications, 43 (??):7–11, April 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume43/
number17/6193-8703/.

Tyagi:2013:GCO

[TR13] B. K. Tyagi and Rachna. On generalized closure operators in generalized topological spaces. International Journal of Computer Applications, 82(??):1–5, November 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume82/number15/14236-2128/.

Thalatam:2014:CSP

[TR14a] M. N. Vamsi Thalatam and Allam Appa Rao. Computational study and performance evaluation of different genomic signal processing methods for identification of protein coding regions (exon regions) of DNA sequence. International Journal of Computer Applications, 85(??):11– 15, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number11/14884-3318/.

Thomas:2014:SMS

[TR14b] Liya Thomas and Syama R. Survey on MapReduce scheduling algorithms. International Journal of Computer Applications, 95(??):9–13, June 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume95/number23/16733-6903/.

T:2015:STB

[TR15] Ramaswamy T and K. Chennakesava Reddy. Space-time block coding (STBC) for multiple transmit antennas over time selective fading channels. International Journal of Computer Applications, 130(??):38–45, November 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume130/number12/23265-2015907143/.

Thakur:2024:WRI

[TRDB24] Rajeev Kumar Thakur, Rajesh Kumar Rai, Mohan Dholvan, and Ashish Bhakoriya. Wideband radiation inverted Uslot microstrip patch antenna design for multiband applications. International Journal of Computer Applications, 186(??):1–5, January 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number3/33050-2024923354/.

Tripathi:2010:BAT

[Tri10] Rajeev R. Kumar Tripathi. Balancing of AVL tree using virtual node. International Journal of Computer Applications, 7(??):12–15, October 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume7/number14/1331-1695/.

Tiwari:2023:AEI

[TRM23] Amitosh Tiwari, Amit Ranjan, and Subhransu Mania. Activation energy impacts on hydromagnetic convective heat transfer flow of nanofluid past a surface of vertical wavy with variable properties. International Journal of Computer Applications, 184(??):35–42, March 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number50/32647-2023922632/.

Tagore:2016:AOD

[TRR16] M. R. N. Tagore, E. V. Krishna Rao, and B. Prabhakar Rao. An accurate optic disc segmentation from peripapillary atrophy incident retinal images. International Journal of Computer Applications, 147(??):14–20, August 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume147/number11/25697-2016911224/.

Tiwari:2015:RMI

[TRSK15] Anoop Kumar Tiwari, Ajay Rajpoot, K. K. Shukla, and S. Karthikeyan. A robust method for image steganography based on chaos theory. International Journal of Computer Applications, 113(??):35–41, March 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume113/number4/19817-1637/.

T:2010:KKO

[TS10] Hamsapriya T and Poomagal S. KOMOS — a keyword occurrence method for ordering search results. International Journal of Computer Applications, 2(??):1–6, June 2010. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume2/number7/683-961/.

Tiwari:2011:NFA

[TS11] Rajdev Tiwari and Manu Pratap Singh. A neuro- fuzzy approach for formulating survey and managing recorded information in data warehouses. International Journal of Computer Applications, 26(??):1–9, February 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume26/number5/3103-2199/.

Tamilpavai:2012:MRP

[TS12] G. Tamilpavai and S. Tamil Selvi. Multiple representations of perceptual features for texture classification and retrieval. International Journal of Computer Applications, 48(??):5–11, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number20/7462-0479/.

Tayel:2014:OTD

[TS14] Mazhar Tayel and Hamed Shawky. An optimized tradeoff decomposition steganography algorithm. International Journal of Computer Applications, 86(??):19–23, January 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume86/number5/14981-3187/

.

Thakur:2015:MBS

[TS15] Tejeshwari Thakur and Birendra Kumar Sharma. Mobile banking system based on certificateless chameleon hash function. International Journal of Computer Applications, 121(??):1–5, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number7/21549-4543/.

Tuncay:2016:SUR

[TS16] Mesut Tunçay and Aslihan Sezgin. Soft union ring and its applications to ring theory. International Journal of Computer Applications, 151(??):7–13, October 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume151/number9/26259-2016911867/.

Tayel:2017:NSC

[TS17] Mazhar Tayel and Mahmoud Sabry. A new scattering clusters equalization algorithm for airtarget acoustic passive detection and classification. International Journal of Computer Applications, 170(??):29–34, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume170/number1/28036-2017914665/.

Tara:2012:SLR

[TSA12] Rayi Yanu Tara, Paulus Insap Santosa, and Teguh Bharata Adji. Sign language recognition in robot teleoperation using centroid distance Fourier descriptors. International Journal of Computer Applications, 48(??):8–12, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume48/number2/7318-0100/.

Tabaa:2015:NIU

[TSCD15] Mohamed Tabaa, Safa Saadaoui, Mouhammad Chehaitly, and Abbas Dandache. NLOS identification for UWB body communications. International Journal of Computer Applications, 124(??):12–17, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number6/22107-2015905496/.

Taher:2017:NIC

[TSE17] Nesma Taher, Shaimaa Salama, and Doaa ElZanfaly. New insight into customer value analysis using data mining techniques. International Journal of Computer Applications, 176(??):27–38, October 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume176/number3/28533-2017915560/.

T:2010:IDP

[TSRM10] Mohamed Mubarak T, Syed Abdul Sattar, Appa Rao, and Sajitha M. Intrusion detection: a probability model for 3D heterogeneous WSN. International Journal of Computer Applications, 6(??):15–20, September 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume6/number12/1125-1475/.

Talebzadeh:2012:ACS

[TSS12] Saeed Talebzadeh, Bahram Hashemi Shahraki, and Ahmad Shariati. Application of CFD in studying and modification of gas burners of catalytic reforming unit of Abadan Refinery in Iran. International Journal of Computer Applications, 55(??):46–51, October 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume55/number9/8787-2760/.

Thakur:2018:BAD

[TSY18] Tulsi Thakur, Kavita Singh, and Arun Yadav. Blind approach for digital image forgery detection. International Journal of Computer Applications, 179(??):34–42, January 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume179/number10/28839-
2018916108/.

Tavassoli:2011:EPF

[TT11] Sude Tavassoli and Samane Tavassoli. An effective pinglish filter for morphological analyzer of Persian adjectives. International Journal of Computer Applications, 28(??):41– 44, August 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume28/
number1/3349-4618/.

Taha:2014:ACE

[TT14] Saleem M. R. Taha and Zahraa K. Taha. Analysis and classification of EEG signals based on a new quantum inspired wavelet neural network model. International Journal of Computer Applications, 92(??):23–30, April 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume92/number5/16005-5008/.

Tanomsakyut:2010:CSC

[TTA10] Teerapon Tanomsakyut, Natcha Thawesaengskulthai, and Pongpun Anuntavoranich. A case study of concept development of service quality measurement system for mobile telecom service encounter in Thailand. International Journal of Computer Applications, 9(??):33–37, November 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume9/number11/1427-
1923/.

Tuaimah:2017:WES

[Tua17] Firas M. Tuaimah. Wind energy system MVA profile enhancement based on feedback quadratic function controller. International Journal of Computer Applications, 178(??):21–26, November 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume178/number6/28679-2017915833/.

Tarigan:2020:IMM

[TUS20] Rehulina Tarigan, Andi Usri, and Ahmad Surahmat. Implementation of MVC (model-view-controller) concept in develop management information system for transportation fees (uang jalan operasional) and revenue in trucking companies. International Journal of Computer Applications, 175(??):1–8, November 2020. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume175/number29/31630-2020920818/.

Tripathi:2010:DSS

[TV10] Pawan Tripathi and Gyanendra K. Verma. A digital security system with door lock system using RFID technology. International Journal of Computer Applications, 5(??):6–8, August 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume5/number11/957-
1334/.

Thankam:2013:ASM

[TV13] T. Akhila Thankam and K. S. Angel Viji. Abnormality segmentation of MRI brain images using fuzzy nearest neighbour approach. International Journal of Computer Applications, 67(??):6–12, April 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume67/number19/11501-7215/.

Thukral:2018:SLC

[TVGT18] Tarun Thukral, Shradha Verma, Swati Gaur, and Nalini Tyagi. Segmentation of lung cancer using mark region growing and median filter. International Journal of Computer Applications, 181(??):15–21, July 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number6/29720-2018917492/.

Tiwari:2019:WES

[TVS19] Varun Tiwari, Vikas Rao Vadi, and Mukta Sharma. WOT enabled smart cities integrated big data analytics. International Journal of Computer Applications, 177(??):39–44, October 2019. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume177/number11/
30945-2019919540/.

Thota:2017:PAE

[TVY17] Jayanth Thota, Priyanka Vangali, and Xiaokun Yang. Prototyping an autonomous eye-controlled system (AECS) using Raspberry-Pi on wheelchairs. International Journal of Computer Applications, 158(??):1–7, January 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume158/number8/26925-2017912853/.

Taheri:2020:GLB

[TX20] Hamid Taheri and Zhao Chun Xia. Geometric landmarkbased pose tracking for holonomic mobile robots. International Journal of Computer Applications, 175(??):11–17, August 2020. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume175/number11/
31496-2020920587/.

Tatar:2013:ALC

[TY13a] Yetkin Tatar and Gungor Yildirim. An application of a low-cost monitoring and guidance system in a WSN. International Journal of Computer Applications, 73(??): 31–38, July 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume73/
number16/12827-0225/.

Tojima:2013:QCS

[TY13b] Tatsuya Tojima and Takashi Yukawa. A query classification system based on snippet similarity for a one-click search. International Journal of Computer Applications, 82 (??):1–8, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume82/
number1/14077-2146/.

Thanh:2015:RSA

[TY15a] Han Trong Thanh and V. U. Van Yem. Robust system architecture for DOA estimation based on total forward backward matrix pencil algorithm. International Journal of Computer Applications, 126(??):32–36, September 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume126/number4/22541-2015906033/.

Tiwari:2015:DBH

[TY15b] Nidhi Tiwari and Raghav Yadav. Detection of black hole attack using control packets in AODV protocol for MANET. International Journal of Computer Applications, 118(??):23–29, May 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume118/number24/20957-3381/.

Tuaimah:2016:VSE

[TY16] Firas M. Tuaimah and Massara Gala Yahya. Voltage stability enhancement for Iraqi (400 kV) super high voltage grid system based on cat swarm optimization algorithm. International Journal of Computer Applications, 134(??):7–11, January 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume134/number16/23996-2016907484/.

Taygi:2015:IMP

[TYSS15] Babita Taygi, Ajay Singh Yadav, Sanjai Sharma, and Anupam Swami. An inventory model with partial backordering, Weibull distribution deterioration under two level of storage. International Journal of Computer Applications, 129(??):15–22, November 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume129/number16/23157-2015907008/.

UlKabir:2014:BAS

[UA14] Md. Wasi Ul Kabir and Mohammad Shafiul Alam. Bat algorithm with self-adaptive mutation: a comparative study on numerical optimization problems. International Journal of Computer Applications, 100(??):7–13, August 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume100/number10/17559-8172/.

Udo:2016:HMB

[UAA16] Ifiok J. Udo, Babajide S. Afolabi, and Bernard I. Akhigbe. Hybrid minutiae-based architecture for automated fingerprint verification system. International Journal of Computer Applications, 133(??):6–12, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number11/23828-2016907908/.

Ujakpa:2017:EAP

[UAMN17] Martin Mabeifam Ujakpa, Rajesh Arora, Aussie K. N. Mutalya, and Laina Namulo. Employee adoption of eprocurement and its implication on supply chain in developing countries. International Journal of Computer Applications, 171(??):11–15, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume171/number9/28208-2017914850/.

Umadevi:2015:RMA

[UC15] V. Umadevi and C. Chandrasekar. Removal of malicious attacks using hybrid symmetric cryptograhic technique. International Journal of Computer Applications, 132(??):1–8, December 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume132/number7/23603-2015907445/.

U:2013:PEH

[UCNC13] Onyishi D U, Nwalozie G C, Aniedu A N, and Ejiofor A C. Performance evaluation of handover channel exchange scheme in GSM network. International Journal of Computer Applications, 70(??):28–33, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume70/number10/12000-7887/.

Uwaechia:2013:PMA

[UDA13] A. N. Uwaechia, D. D. Dajab, and B. A. Adegboye. Performance modeling of ahmadu bello university data network using a storage area network (SAN) design concept. International Journal of Computer Applications, 65 (??):6–13, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number10/10958-5924/.

Udoh:2016:IDD

[UI16] Linus Udoh and Ibrahim Inuwa. Implementation of a data driven transaction processing system for the AUN’s Restaurant, Yola — Nigeria. International Journal of Computer Applications, 149(??):9–20, September 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume149/number6/26000-2016911419/.

Umbarkar:2013:DPG

[UJ13] A. J. Umbarkar and M. S. Joshi. Dual population genetic algorithm (GA) versus OpenMP GA for multimodal function optimization. International Journal of Computer Applications, 64(??):29–36, February 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume64/number19/10744-5516/.

Utsai:2014:ARU

[UJ14] Sonali Utsai and Ram B. Joshi. DOS attack reduction by using Web service filter. International Journal of Computer Applications, 105(??):4–9, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume105/number14/18443-9813/.

Umoh:2013:PFF

[UN13] Uduak A. Umoh and Mfon M. Ntekop. A proposed fuzzy framework for cholera diagnosis and monitoring. International Journal of Computer Applications, 82(??):1– 10, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume82/
number17/14252-1626/.

Ulle:2018:AAT

[UNM18a] Anand Raj Ulle, T. N. Nagabhushan, and Nandini Manoli. An automated approach towards detection of mitosis in histopathological images. International Journal of Computer Applications, 180(??):1–7, April 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number35/29287-2018916886/.

Ulle:2018:CHI

[UNM18b] Anand Raj Ulle, T. N. Nagabhushan, and Nandini Manoli. Classification of histopathological images based on modified clump splitting approach. International Journal of Computer Applications, 182(??):1–8, September 2018. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume182/number18/29959-2018917888/.

Uwho:2023:DII

[UO23] Sophie Uwho and Bourdillon O. Omijeh. Design and implementation of an intelligent based remote health monitoring system using wireless sensor network. International Journal of Computer Applications, 185(??):1–7, July 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number20/32805-2023922817/.

Upadhye:2023:ARC

[Upa23] Akshata Upadhye. Automating resume classification: Leveraging NLP and AI for efficient candidate screening. International Journal of Computer Applications, 185(??):46–50, November 2023. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume185/number40/
32956-2023923208/.

Umamaheswari:2012:HMI

[UR12] J. Umamaheswari and G. Radhamani. Hybrid medical image segmentation based on fuzzy global minimization by active contour model. International Journal of Computer Applications, 39(??):1–6, February 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume39/number13/4877-7309/.

Atta-ur-rahman:2013:TAP

[ur13] Atta ur rahman. Teacher assessment and profiling using fuzzy rule based system and apriori algorithm. International Journal of Computer Applications, 65(??):22– 28, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number5/10921-5859/.

Umardand:2016:MSB

[UR16] Dheeraj D. Umardand and Chirchi V R. Mobile sink-based self adaptive and energy efficient proactive data collection protocol for wireless sensor network. International Journal of Computer Applications, 144(??):17–24, June 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume144/number7/25191-2016910438/.

Pasha:2013:BTS

[urPP13] Muhammad Anwar ur-rehman Pasha and Shaheen Pasha. Bloom’s taxonomy for standardizing BPM education of IT under-graduates students. International Journal of Computer Applications, 78(??):6–13, September 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume78/number6/13491-1204/.

U:2012:NCM

[US12] Shajeesh K U and K. P. Soman. Noise cancellation method for robust speech recognition. International Journal of Computer Applications, 45(??):38–44, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume45/number11/6827-9438/.

Umamageswari:2014:NCD

[US14] A. Umamageswari and G. R. Suresh. A new cryptographic digital signature for secure medical image communication in telemedicine. International Journal of Computer Applications, 86(??):4–9, January 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume86/number11/15027-3343/.

Upadhyay:2015:VBA

[US15] Rashmi Upadhyay and Ritesh Shah. A VPRS based approach for enhancement in the classification efficiency of machinery datasets. International Journal of Computer Applications, 112(??):11–15, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number15/19741-0870/.

Utpatadevi:2012:IMM

[USC12] Ni L. P. Pravina Utpatadevi, A. A. K. Oka Sudana, and A. A. Kt. Agung Cahyawan. Implementation of MVC (model-view-controller) architectural to academic management information system with Android platform base. International Journal of Computer Applications, 57(??):1– 6, November 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume57/
number8/9131-3313/.

Upadhyay:2017:SMN

[USK17] Navin Mani Upadhyay, Kumari Soni, and Arvind Kumar. Security management in network visualization environment by using modified root management algorithm for creation of virtual node and virtual link. International Journal of Computer Applications, 162(??):37–43, March 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume162/number6/27250-2017913367/.

U:2017:RRD

[USR17] Vasanthakumar G U, P. Deepa Shenoy, and Venugopal K R. RDUP3: Relative distance based user profiling from profile picture in multi-social networking. International Journal of Computer Applications, 157(??):7–15, January 2017. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume157/number10/26865-
2017912834/.

Uzoma:2013:CBU

[Uzo13] Ojemeni Uzoma. Correlation between uplink noise, uplink load and call drop rate in a WCDMA network. International Journal of Computer Applications, 66(??):39– 41, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume66/
number5/11084-5881/.

Vadgaonkar:2020:NAD

[Vad20] Priti H. Vadgaonkar. Network anomaly detection and user behavior analysis using machine learning. International Journal of Computer Applications, 175(??):47–53, August 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume175/number13/31517-
2020920635/.

Vaisla:2010:SAI

[VB10] Kunwar Singh Vaisla and Manoj Kumar Bisht. SWOT analysis of e-initiative in Uttarakhand. International Journal of Computer Applications, 12(??):5–14, December 2010. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume12/number5/1677-
2263/.

Voola:2013:FRP

[VB13] Persis Voola and A Vinaya Babu. 4A frameworks for requirements prioritization. International Journal of Computer Applications, 76(??):38–44, August 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume76/number1/13215-0598/.

Vijayaraj:2010:DIC

[VD10] A. Vijayaraj and P. DineshKumar. Design and implementation of census data collection system using PDA. International Journal of Computer Applications, 9(??):28– 32, November 2010. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume9/
number9/1412-1908/.

Verma:2015:MOE

[VD15] Ramakant Verma and Maitreyee Dutta. Moving object extraction through a real-world variable-bandwidth network using KFDA-based RBF. International Journal of Computer Applications, 116(??):15–22, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume116/number20/20452-2806/.

Vyza:2022:SPB

[VDM22] Yashwanth Vyza, Suman Dabbu, and Malini Mudigonda. A spectral power based index using wavelet analysis of EEG for drowsiness detection. International Journal of Computer Applications, 183(??):1–8, January 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number46/32236-2022921731/.

Sundar:2012:DDC

[VDS12] Bala Sundar V, T Devi, and N Saravanan. Development of a data clustering algorithm for predicting heart. International Journal of Computer Applications, 48(??): 8–13, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume48/
number7/7358-0095/.

Venkataraman:2014:ISE

[Ven14a] Ganesh Venkataraman. Intelligent search engine ranking algorithm inspired by recommendation engines. International Journal of Computer Applications, 85(??):16– 18, January 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume85/
number8/14860-3231/.

Vennila:2014:SET

[Ven14b] S. Mary Vennila. Security enhanced trust based G-LeaSel for grid environment. International Journal of Computer Applications, 106(??):29–34, November 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume106/number8/18542-9770/.

Verma:2016:AAC

[Ver16] Surbhi Verma. Android app controlled Bluetooth robot. International Journal of Computer Applications, 152(??): 35–40, October 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume152/number9/26351-2016911912/.

Verma:2017:CBR

[Ver17] Mayuri Verma. Cluster based ranking index for enhancing recruitment process using text mining and machine learning. International Journal of Computer Applications, 157(??):23–30, January 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume157/number9/26860-2017912812/.

Vashist:2012:RSA

[VG12] Renu Vashist and M. L. Garg. A rough set approach for generation and validation of rules for missing attribute values of a data set. International Journal of Computer Applications, 42(??):31–35, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number14/5762-7938/.

Virak:2016:EIM

[VG16] Sajjan Virak and Amit Grover. The effect of interference of mobile ADHOC networks on WIMAX. International Journal of Computer Applications, 149(??):

16–22, September 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume149/number7/26009-2016911459/.

Varghese:2012:SNS

[VGJ12] Jibin M. Varghese, Anish Gopinath, and P. V. John. A single neuron speed controller for BLDC motor driven by pseudo sinusoid commutation. International Journal of Computer Applications, 43(??):26–33, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number2/6078-8201/.

Vahabi:2012:FNN

[VGo12] Hamid Reza Vahabi and Hasan Ghasabi-oskoei. A feedback neural network for solving nonlinear programming problems with hybrid constraints. International Journal of Computer Applications, 54(??):41–46, September 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume54/number5/8565-2164/.

Vankawala:2015:SDI

[VGP15] Fagun Vankawala, Amit Ganatra, and Amit Patel. A survey on different image deblurring techniques. International Journal of Computer Applications, 116(??):15–18, April 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume116/number13/20396-
2697/.

Rajkumar:2011:EBR

[VGS11] Franklin Rajkumar V, Manekandan GRS, and V. Santhi. Entropy based robust watermarking scheme using Hadamard transformation technique. International Journal of Computer Applications, 12(??):14–21, January 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume12/number9/1712-2293/.

Varghese:2024:HHM

[VJ24] Samuel Benny Varghese and Eldho K J. Harvest horizon: machine learning advancements in crop projection. International Journal of Computer Applications, 186(??):1–6, April 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number17/
harvest-horizon-machine-learning-advancements-in-
crop-projection/.

Vijayarani:2022:EKS

[VJD22] S. Vijayarani, R. Janani, and S. Dinesh. An efficient keyword searching algorithm for information retrieval from desktop. International Journal of Computer Applications, 184(??):40–44, April 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number8/32350-2022922053/.

Vaneshani:2011:NCC

[VJR11] Saeed Vaneshani and Hooshang Jazayeri-Rad. Nonlinear control of a chemical plant employing a combination of fuzzy logic and particle swarm optimization techniques. International Journal of Computer Applications, 33(??):58– 63, November 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume33/
number9/4047-5807/.

Vinayakray-jani:2012:PTM

[VjS12] Preetida Vinayakray-jani and Sugata Sanyal. Proactive TCP mechanism to improve handover performance in mobile satellite and terrestrial networks. International Journal of Computer Applications, 48(??):25–29, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume48/number25/7546-0643/.

Varade:2011:PAM

[VK11] Suchita Varade and Kishore Kulat. Performance analysis of MIMO-OFDM system using space-time turbo codes and adaptive beamforming. International Journal of Computer Applications, 22(??):27–32, May 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume22/number3/2563-3524/.

Vanjulavalli:2013:OBI

[VK13] N. Vanjulavalli and A. Kovalan. Ontology based information retrieval for semi structure data using bagging. International Journal of Computer Applications, 67(??): 6–11, April 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume67/
number15/11469-7077/.

Vahidi:2014:EGB

[VKB14] Majid Reza Vahidi, Mohammad Mansour Riahi Kashani, and Alireza Bagheri. An efficient gradient based algorithm for improving performance of image edge detection. International Journal of Computer Applications, 103(??):7–14, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume103/number4/18060-8991/.

Visave:2015:VFG

[VKJ^{+}15] Ashwini Visave, Pramod Kachare, Amutha Jeyakumar, Alice Cheeran, and Gurmit Bachher. Vocal features for glottal pathology detection using BPNN. International Journal of Computer Applications, 118(??):1–6, May 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume118/number17/20834-3571/.

Velusamy:2012:PAU

[VM12a] K. Velusamy and R. Manavalan. Performance analysis of unsupervised classification based on optimization. International Journal of Computer Applications, 42(??):22– 27, March 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume42/
number19/5802-8090/.

Verma:2012:EFB

[VM12b] H. K. Verma and Prachi Mafidar. Environment friendly BB-BC optimized economic dispatch with real and reactive power constraints. International Journal of Computer Applications, 43(??):10–15, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number4/6090-8266/.

Verma:2014:CCW

[VM14a] Anu Verma and Nitin Mittal. Congestion controlled WSN using genetic algorithm with different source and sink mobility scenarios. International Journal of Computer Applications, 101(??):8–15, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number13/17746-8819/.

Victor:2014:DMB

[VM14b] Crallet Victor and Nixon Mtonyole. Design of multicarrier broadband power line communication system for Tanzania homes. International Journal of Computer Applications, 100(??):25–32, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number14/17595-8380/.

Vangala:2023:MST

[VMA23] Sreenivasa Rao Vangala, Ravi Kiran Mallidi, and V. L. Prasuna Appili. Micro-services transactions resilience across bounded domains: an architecture perspective. International Journal of Computer Applications, 184(??):30–35, February 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume184/number45/32608-
2023922557/.

Veeramuthu:2014:NAA

[VMM14] A. Veeramuthu, S. Meenakshi, and P. S. Manjusha. A new approach for Alzheimer’s disease diagnosis by using association rule over PET images. International Journal of Computer Applications, 91(??):9–14, April 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume91/number9/15908-5009/.

Viswanathan:2013:CAC

[VMT13] C. Viswanathan, K. M. Mehata, and Latha Tamilselvan. Call admission control based on bandwidth allocation in ultra wide band (UWB) ad hoc networks. International Journal of Computer Applications, 69(??):14–20, May 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume69/number1/11805-7453/.

Vidyabanu:2010:MBF

[VN10] R. Vidyabanu and N. Nagaveni. A model based framework for privacy preserving clustering using SOM. International Journal of Computer Applications, 1(??):17–21, February 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume1/number13/288-450/.

I:2015:DMS

[VNM15] Accamma I V, Suma H N, and Dakshayini M. Decoding multiple subject fMRI data using manifold based representation of cognitive state neural signatures. International Journal of Computer Applications, 115(??):1–7, April 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume115/number15/20224-2512/.

Venkataramani:2010:ZZP

[VP10] Y. Venkataramani and M. Padmaa. ZIG-ZAG PVD — a nontraditional approach. International Journal of Computer Applications, 5(??):5–10, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume5/number7/928-1305/.

Volpe:2010:OVP

[VPGF10] Yary Volpe, Matteo Palai, Lapo Governi, and Rocco Furferi. From 2D orthographic views to 3D pseudo-wireframe: an automatic procedure. International Journal of Computer Applications, 5(??):18–24, August 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume5/number6/918-1296/.

Vijayalakshmi:2011:WWA

[VR11] S. Vijayalakshmi and S. Albert Rabara. Weeding wormhole attack in MANET multicast routing using two novel techniques — LP3 and NAWA2. International Journal of Computer Applications, 16(??):26–33, February 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume16/number7/2024-2729/.

Viswanadham:2012:QBS

[VR12] K. N. S. Kasi Viswanadham and Y. Showri Raju. Quartic B-spline collocation method for fifth order boundary value problems. International Journal of Computer Applications, 43(??):1–6, April 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume43/number13/6160-8566/.

Vaisi:2014:PIP

[VR14] Bahareh Vaisi and Sadigh Raissi. Productivity improvement in the pride’s spare parts manufacturing using computer simulation and data envelopment analysis. International Journal of Computer Applications, 95(??):12–18, June 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume95/number7/16605-6431/.

Veni:2015:GDB

[VR15a] S. Krishna Veni and G. S. N. Raju. Generation of dualbeam patterns using particle swarm optimization. International Journal of Computer Applications, 113(??):51–56, March 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume113/number10/
19866-1849/.

Viswanadham:2015:NSS

[VR15b] K. N. S. Kasi Viswanadham and S. M. Reddy. Numerical solution of seventh order boundary value problems by Petrov– Galerkin method with quintic B-splines as basis functions and septic B-splines as weight functions. International Journal of Computer Applications, 122(??):41–47, July 2015. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume122/number5/21700-4811/.

Viswanadham:2010:ULI

[VRS10] Y. K. Viswanadham, Ch. D. V. Subba Rao, and T. V. Subrahmanyam. Ultra long integer multiplication on GDPS. International Journal of Computer Applications, 7(??):22–25, October 2010. CODEN ???? ISSN 0975-8887. URL https:/

/www.ijcaonline.org/archives/volume7/number8/1271-
1751/.

Vaithianathan:2012:GAI

[VRS12] V. Vaithianathan, J. Raja, and R. Srinivasan. A 1.7 ñ 2.5 GHz active inductor based low power low noise amplifier for multi standard applications. International Journal of Computer Applications, 39(??):16–24, February 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume39/number5/4816-7054/.

Venkateswaran:2010:IST

[VS10] R. Venkateswaran and V. Sundaram. Information security: Text encryption and decryption with poly substitution method and combining the features of cryptography. International Journal of Computer Applications, 3(??):28–31, June 2010. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume3/number7/741-1047/.

Vivekananthamoorthy:2012:EEL

[VS12] N. Vivekananthamoorthy and K. Sarukesi. Enhancing effectiveness of e-learning framework using UML modeling and self regulation: a case study. International Journal of Computer Applications, 48(??):49–56, June 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume48/number6/7354-0171/.

Vittapu:2015:SCO

[VSA15] Mani Sarma Vittapu, Venkateswarlu Sunkari, and Atoyoseph Abate. Synchronous CR-OLAP tool for efficient parallel computing. International Journal of Computer Applications, 115(??):1–13, April 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume115/number17/20240-2593/.

Venkataramanan:2013:CAC

[VSN13] K. Venkataramanan, B. Shanthi, and S. P. Natarajan. Comparative analysis on carrier overlapping PWM strategies for seven level symmetrical inverter. International Journal of Computer Applications, 79(??):35– 40, October 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume79/
number12/13795-1872/.

Vidyasagar:2016:TST

[VSS16] K. Vidyasagar, K. Swathi, and K. Sudharsanarao. Three stage toll gate alarming mechanism on road highways. International Journal of Computer Applications, 143(??):24–29, June 2016. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume143/number12/
25130-2016910136/.

Vishwakarma:2017:IDS

[VST17] Satyendra Vishwakarma, Vivek Sharma, and Ankita Tiwari. An intrusion detection system using KNN-ACO algorithm. International Journal of Computer Applications, 171(??):18–23, August 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume171/number10/28291-2017914079/.

Varghese:2015:CEJ

[VT15] Arathy Varghese and Sudha T. Capacity enhancement with joint subchannel and power allocation scheme for OFDMA femtocell networks. International Journal of Computer Applications, 124(??):1–6, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number14/22170-2015905757/.

Vishesh:2012:FVA

[VV12] Kinchit Vishesh and Amandeep Verma. Formal verification of authenticated AODV protocol using AVISPA. International Journal of Computer Applications, 50(??): 38–43, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume50/
number19/7914-1179/.

Valmy:2014:BIC

[VV14] Larissa Valmy and Jean Vaillant. Bayesian inference on a Cox process associated with a Dirichlet process. International Journal of Computer Applications, 95(??): 1–7, June 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume95/
number18/16691-6825/.

Ventayen:2018:ELP

[VV18] Lemuel Magno Ventayen and Randy Joy Magno Ventayen. Ethical and legal perspective of software piracy and business effects on the illegal use of peer-to-peer systems architecture. International Journal of Computer Applications, 179(??):30–34, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume179/number47/29494-2018917278/.

Vyavahare:2014:CBD

[Vya14] Arati J. Vyavahare. Canny based DRLSE algorithm for segmentation. International Journal of Computer Applications, 102(??):1–5, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume102/number7/17824-8483/.

Waziri:2011:PAD

[WAG11] M. Y. Waziri, H. Aisha, and A. I. Gambo. On performance analysis of diagonal variants of Newton’s method for large-scale systems of nonlinear equations. International Journal of Computer Applications, 35(??):18–22, December 2011. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume35/number3/4381-
6064/.

Waris:2012:SIM

[WAM12] Madiha Waris, Farooque Azam, and Abdul Wahab Muzaffar. A survey of issues in multimedia databases. International Journal of Computer Applications, 46(??):1–9, May 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume46/number7/6917-
8972/.

Wahed:2014:EMM

[WAmMS14] Mohamed Wahed, Hassan Al-mahdi, Tarek M Mahmoud, and Hassan Shaban. The effect of mobility models and traffic patterns on the performance of routing protocols in MANETs. International Journal of Computer Applications, 101(??):52–58, September 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume101/number9/17719-8778/.

Wilfred:2022:CUE

[WBK22] Ajeba S. Wilfred, Adebanji Bankole, and Owoeye F. Kemisola. Catalyzing uninterrupted energy availability in Nigeria through decentralized off-grid hybrid energy exploitation. International Journal of Computer Applications, 184(??):9–20, May 2022. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number13/32382-2022922114/.

Wickramasinghe:2023:DAS

[WBL^{+}23] Lakshika Medhavini Wickramasinghe, Pasindu Bhasura, Ishanka Pramidu Lakshan, Charith Widanapathirana, and Nelum Amarasena. DriveCare — automated and sustainable points manipulation system for drivers and traffic police officers. International Journal of Computer Applications, 185(??):11–16, April 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number5/32700-2023922687/.

Wakif:2016:NST

[WBS16] Abderrahim Wakif, Zoubair Boulahia, and Rachid Sehaqui. Numerical study of a thermal convection induced by a purely internal heating in a rotating medium saturated by a radiating nanofluid. International Journal of Computer Applications, 135(??):33–42, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number10/24088-2016908529/.

Wazirali:2015:EBH

[WC15] Raniyah Abdullah Wazirali and Zenon Chaczko. EA based heuristic segmentation for efficient data hiding. International Journal of Computer Applications, 118(??):1–7, May 2015. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume118/number5/20738-3126/

.

Waziri:2015:WCB

[WDAJ15] Victor Onomza Waziri, Hakimi Danladi, John K. Alhassan, and Faiza Babakano Jada. Web-cloud-based security services based-on elliptic curves cryptosystem. International Journal of Computer Applications, 115(??):12–18, April 2015. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume115/number23/
20290-2146/.

Wang:2016:ACD

[WDD16] YiMing Wang, Chen Deng, and GaoXu Deng. Adaptive controller design for the control systems with deadzone. International Journal of Computer Applications, 139(??):29–33, April 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume139/number6/24495-2016909151/.

Wathsara:2020:EWM

[WDN^{+}20] H. P. H. Wathsara, M. L. P. D. Dhananjana, D. A. Nanayakkara, B. B. K. R. Dharmarathne, L. H. Sanjeevi Chandrasiri, and J. K. Joseph. Employee work motivation detection using image and audio processing. International Journal of Computer Applications, 175(??):34–40, November 2020. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume175/number31/31651-
2020920862/.

Wilson:2013:GAB

[WDS13] Ian David Wilson, Ross Davies, and Nigel Stanton. A genetic algorithm based solution to the teaching assignment problem. International Journal of Computer Applications, 81(??):1–6, November 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume81/
number19/14268-0056/.

Wirdiani:2017:AAB

[WES17] Ni Kadek Ayu Wirdiani, Ni Nyoman Triana Anggra Emi, and A. A. K. Oka Sudana. Application of Android-based ear biometrics identification. International Journal of Computer Applications, 172(??):11–17, August 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume172/number10/28286-2017915181/.

Worch:2014:CSF

[WG14] Jan-Hendrik Worch and Bjoern Gottfried. Choosing shape features by means of genetic algorithms for gylph-clustering of historical documents. International Journal of Computer Applications, 102(??):1–6, September 2014. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume102/number3/17792-8585/.

Wijendra:2021:ACC

[WH21] Dinuka R. Wijendra and K. P. Hewagamage. Analysis of cognitive complexity with cyclomatic complexity metric of software. International Journal of Computer Applications, 174(??):14–19, February 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume174/number19/31784-2021921066/.

Wisaeng:2014:ADO

[WHP14] Kittipol Wisaeng, Nualsawat Hiransakolwong, and Ekkarat Pothiruk. Automatic detection of optic disc in digital retinal images. International Journal of Computer Applications, 90(??):15–20, March 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume90/number5/15569-4105/.

Huang:2012:TPH

[wHsYwT12] Ko wei Huang, Chu sing Yang, and Chun wei Tsai. A two-phase hybrid particle swarm optimization algorithm for solving permutation flow-shop scheduling problem. International Journal of Computer Applications, 48(??):11–18, June 2012. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume48/number1/7311-
9883/.

Win:2014:CXS

[Win14] Thidar Win. Conversion of XML schema to data warehouse schema using automatic approach. International Journal of Computer Applications, 108(??):12–18, December 2014. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume108/number15/
18986-0424/.

Wisaeng:2013:ECD

[Wis13] Kittipol Wisaeng. An empirical comparison of data mining techniques in medical databases. International Journal of Computer Applications, 77(??):23–27, September 2013. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume77/number7/13408-1061/

.

Wanjala:2016:SMF

[WK16] George Wafula Wanjala and Andrew M. Kahonge. Social media forensics for hate speech opinion mining. International Journal of Computer Applications, 155(??): 39–47, December 2016. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume155/number1/26573-2016912258/.

Wynn:2024:SIP

[WKAL24] Scott Wynn, Alec Kyritsis, Stephora Alberi, and Enyue Lu. Selection improvements on the parallel iterative improvement algorithm for stable matching. International Journal of Computer Applications, 186(??):49–56, September 2024. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume186/number43/
selection-improvements-on-the-parallel-iterative-
improvement-algorithm-for-stable-matching/.

Waghmare:2011:PAM

[WKC11] Vivek N. Waghmare, Sandip V. Kendre, and Sanket G. Chordiya. Performance analysis of matrix–vector multiplication in hybrid (MPI + OpenMP). International Journal of Computer Applications, 22(??):22–25, May 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume22/number5/2579-3561/.

Wakchaure:2018:AAO

[WKG^{+}18] M. A. Wakchaure, Nishant L. Kale, Pradeep G. Gawade, Lakhan V. Gangule, and Laxman D. Chate. Automation in agriculture to optimize utilization of water and crop monitoring using IOT. International Journal of Computer Applications, 180(??):1–6, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number50/29574-2018917235/.

Wisana:2021:DOS

[WNE21] I. Dewa Gede Hari Wisana, Priyambada Cahya Nugraha, and Dwiana Estiwidani. Design oxygen saturation monitoring in telemedicine smartphone system (TmSS). International Journal of Computer Applications, 174(??):42–45, February 2021. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume174/number17/31773-
2021921062/.

Wornyo:2018:CKE

[WOBK18] Dickson Keddy Wornyo, Elias Nii Noi Ocquaye, and Bright Bediako-Kyeremeh. Coupled kernel ensemble regression. International Journal of Computer Applications, 181(??):1–8, December 2018. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume181/number34/
30207-2018918278/.

Williams:2023:CMS

[WRJ23] Kemp Williams, Gregory Roberts, and Jason James. Combining multiple sentiment analysis dimensions into a comprehensive sentiment metric. International Journal of Computer Applications, 185(??):14–19, August 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number27/32859-2023923018/.

Waly:2013:SRF

[WSW^{+}13] M. Waly, Amr Sharawy, Khaeld Wahba, Ayman Salah, and Islam Ibrahem. Sensitivity of the risk factors for the progression of ocular hypertension to primary open angle glaucoma. International Journal of Computer Applications, 70 (??):34–42, May 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume70/
number22/12202-8486/.

Wagh:2013:WSP

[WT13] Kishor S. Wagh and R. C. Thool. Web service provisioning on Android mobile host. International Journal of Computer Applications, 81(??):5–11, November 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume81/number14/14182-2338/.

Wu:2021:FAR

[Wu21] Shinq-Jen Wu. Fuzzy approach to regulate S-type biological systems. International Journal of Computer Applications, 183(??):50–55, June 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number10/31967-2021921409/.

Wu:2023:CVT

[Wu23a] Shinq-Jen Wu. Construct and visualize three fuzzy controllers for biological S-systems. International Journal of Computer Applications, 185(??):53–60, April 2023. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume185/number3/32688-2023922697/.

Wu:2023:FRA

[Wu23b] Shinq-Jen Wu. Frequency response analysis for T -S fuzzy systems. International Journal of Computer Applications, 184(??):17–22, February 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume184/number47/32622-2023922580/.

Wisana:2012:IPV

[WWSS12] I. Dewa Gede Hari Wisana, Thomas Sri Widodo, Mochammad Sja’bani, and Adhi Susanto. Identification of premature ventricular contraction ECG signal using wavelet detection. International Journal of Computer Applications, 46(??):11–15, May 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume46/
number16/6992-9424/.

Xu:2011:RCP

[XCY11] Yunlong Xu, Yuping Cao, and Songlin Yang. Research on care of postoperative patient based on rough sets theory. International Journal of Computer Applications, 31 (??):8–12, October 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume31/
number10/3859-5383/.

Xezonakis:2021:AHE

[XL21] Ioannis S. Xezonakis and Svoronos Leivadaros. N-ary Huffman encoding using high-degree trees — a performance comparison. International Journal of Computer Applications, 183(??):33–39, May 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number5/31925-2021921315/.

Xiao:2014:MVF

[XRL14] Jingjing Xiao, Hongyu Ren, and Wusheng Luo. Modification of view field in wireless multimedia sensor networks. International Journal of Computer Applications, 98 (??):7–14, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume98/
number5/17177-7272/.

Yakubu:2014:RFV

[YA14] Osman Yakubu and Osei Adjei. Reliability of fingerprint verification in Ghana. International Journal of Computer Applications, 107(??):9–13, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume107/number10/18785-0116/.

Younes:2018:ETS

[YA18] Yousuf Mohammad Younes and Abdul Ghani A. A. Earthing transformer to support the normal operation of the power system at failure of a protection system. International Journal of Computer Applications, 182(??):1–6, December 2018. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume182/number31/30223-
2018918222/.

Younes:2014:FAS

[YAA14] Nesma Younes, Said M. Abdallah, and M. Abdel Alim. Fuzzy approach for the synthesis of mass exchange network. International Journal of Computer Applications, 104(??):5–11, October 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume104/number2/18172-9058/.

Yacine:2015:TAW

[Yac15] El Younoussi Yacine. Towards an Arabic Web-based information retrieval system (ARABIRS): Stemming to indexing. International Journal of Computer Applications, 109(??):28–33, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume109/number14/19257-1009/.

Yehia:2011:AIR

[YAE11] Mohamad A. Yehia, Mohammed S. Aziz, and Hussein A. Elsayed. Analysis of IGP routing protocols for real time applications: a comparative study. International Journal of Computer Applications, 26(??):11–17, July 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume26/number3/3085-4223/.

Yakubu:2018:IMV

[Yak18] Kamaluddin Yahaya Yakubu. Implementation of mobile voting application in Infrastructure University Kuala Lumpur, Malaysia. International Journal of Computer Applications, 180(??):25–31, June 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number47/29555-2018917288/.

Yahyaoui:2015:AOS

[YAT15a] Imene Yahyaoui, Mahmoud Ammous, and Fernando Tadeo. Algorithm for optimum sizing of a photovoltaic water pumping system. International Journal of Computer Applications, 111(??):21–28, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume111/number6/19544-1409/.

Yildiz:2015:CCM

[YAT15b] Yunus Emre Yildiz, Oguz Altun, and Ali Osman Topal. Computational chemotaxis in micro bacterial foraging optimization for high dimensional problems: a comparative study on numerical benchmark. International Journal of Computer Applications, 124(??):1–8, August 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume124/number4/22096-2015905406/.

Yusuf:2012:SMP

[YB12] Ibrahim Yusuf and Saminu I. Bala. Stochastic modeling and performance measures of redundant system operating in different conditions. International Journal of Computer Applications, 50(??):23–29, July 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume50/number22/7936-1264/.

Yadav:2015:SIS

[YB15] Shailaja S Yadav and D. G. Bhalke. Speaker identification system using wavelet transform and VQ modeling technique. International Journal of Computer Applications, 112(??):19–23, February 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume112/number9/19694-1453/.

Yassine:2012:SHR

[YBA12] Himeur Yassine, Boudraa Bachir, and Khelalef Aziz. A secure and high robust audio watermarking system for copyright protection. International Journal of Computer Applications, 53(??):33–39, September 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume53/number17/8515-2540/.

Yeshambel:2016:IOS

[YBM16] Tilahun Yeshambel, Mengistu Belete, and Yelkal Mulualem. Impact of online social networking on employees productivity at work place in university of gondar — a case study. International Journal of Computer Applications, 135(??):18–24, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number3/24030-2015907844/.

Younes:2014:SAQ

[YEE14] Ali Younes, Mohamed Essaaidi, and Ahmed El Moussaoui. SFL algorithm for QoS-based cloud service composition. International Journal of Computer Applications, 97 (??):42–49, July 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume97/
number17/17103-7700/.

Gicquel:2013:SCP

[yGL13] Pierre yves Gicquel and Dominique Lenne. Semantic and contextual proximities for informal learning: the case study of museum visits. International Journal of Computer Applications, 69(??):10–17, May 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume69/number26/12134-8410/.

Yazdi:2011:ENM

[YJ11] Mohammad Rezaee Yazdi and Hooshang Jazayerirad. The effect of nonlinearity measure on model order selection in identification of chemical processes. International Journal of Computer Applications, 32(??):1–5, October 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume32/number9/3929-5134/.

Younis:2014:DVE

[YJ14] Mohammed Younis and A. I. A. Jabbar. Design and validation of Ethernet network interface card using Simevents technique. International Journal of Computer Applications, 87(??):48–56, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume87/
number10/15247-3799/.

Yarava:2012:RNN

[YK12] Anitha Yarava and S. Murali Krishna. Reachable nearest neighbour query processing of R-trees for moving query object. International Journal of Computer Applications, 49(??):1–3, July 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume49/
number17/7716-0954/.

Yan:2013:MBE

[YK13] Mu Yan and Younggoo Kwon. Multi-beacons energy-efficient geographical routing for low rate wireless personal area networks. International Journal of Computer Applications, 72 (??):28–33, June 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume72/
number7/12507-9100/.

Yadav:2016:WFM

[YK16] Rohit Kumar Yadav and Ravi Khatri. A weather forecasting model using the data mining technique. International Journal of Computer Applications, 139(??):4–12, April 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume139/number14/24665-2016908900/.

Yadav:2012:DBI

[YM12] Ravi Kant Yadav and Bireshwar Dass Mazumdar. Detection of bold and italic character in Devanagari script. International Journal of Computer Applications, 39(??):19– 22, February 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume39/
number2/4792-7037/.

Yousaf:2013:VLT

[YMA13] Kanwal Yousaf, Muhammad Nadeem Majeed, and Adeel Akram. Virtual Learner 2.0: a technique to evaluate the importance of Web 2.0 based e-learning application. International Journal of Computer Applications, 64(??):7– 14, February 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume64/
number4/10620-4248/.

Yalagi:2016:LBC

[YMA16] Pratibha S. Yalagi, Mitali M. Mohire, and Sulabha S. Apte. Load balancing in cluster system using entropy value. International Journal of Computer Applications, 135(??):26–32, February 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume135/number4/24039-2016908319/.

Yassien:2017:GWO

[YMAS17] Eman Yassien, Raja Masadeh, Abdullah Alzaqebah, and Ameen Shaheen. Grey wolf optimization applied to the 0/1 knapsack problem. International Journal of Computer Applications, 169(??):11–15, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume169/number5/27980-2017914734/.

Yousfi:2012:CBH

[YMR^{+}12] Najah Yousfi, Pierre Melchior, Chokri Rekik, Nabil Derbel, and Alain Oustaloup. Comparison between H8 and CRONE control combined with QFT approach to control multivariable systems in path tracking design. International Journal of Computer Applications, 45(??):1–9, May 2012. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume45/number12/6829-9192/.

Yousif:2013:NLP

[You13] Jabar H. Yousif. Natural language processing based soft computing techniques. International Journal of Computer Applications, 77(??):43–49, September 2013. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume77/number8/13418-1089/.

Yousif:2018:CIP

[You18] Abbas Mohammed Younus Yousif. Controlling for IoT projects systems based on smart phone using firebase notification technique. International Journal of Computer Applications, 181(??):21–24, December 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number33/30203-2018918218/.

Yunianto:2019:BDD

[YPS19] Eko Yunianto, Yudi Prayudi, and Bambang Sugiantoro. B-DEC: Digital evidence cabinet based on blockchain for evidence management. International Journal of Computer Applications, 181(??):22–29, March 2019. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number45/30422-2019918580/.

Yadav:2012:ICD

[YR12] Vinay Kumar Yadav and Ashwani K. Rana. Impact of channel doping on DG-MOSFET parameters in nano regime-TCAD simulation. International Journal of Computer Applications, 37(??):36–41, January 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume37/number11/4732-6926/.

Yoga:2023:HML

[YRA23] Castro A. Yoga, Anthony J. Rodrigues, and Silvance O. Abeka. Hybrid machine learning approach for attack classification and clustering in network security. International Journal of Computer Applications, 185(??):45–51, August 2023. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume185/number31/32896-
2023923076/.

Yazdanpanah:2015:RTH

[YRMB15] Ali Pour Yazdanpanah, Emma E. Regentova, Venkatesan Muthukumar, and George Bebis. Real-time horizon line detection based on fusion of classification and clustering. International Journal of Computer Applications, 121(??):5–11, July 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume121/number10/21574-4625/.

Yadav:2014:UCB

[YRV14] Nalini Yadav, K. Rajeswari, and V. V. Vaithiyanathan. Usage of classification based association for removal of noisy attributes. International Journal of Computer Applications, 108(??):1–5, December 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume108/number13/18968-5716/.

Yadav:2017:DBM

[YRV17] K. Rajesh Yadav, T. Srinivasa Rao, and P. Suresh Varma. Dynamic bandwidth management for wireless ad hoc networks for two zones under homogeneous conditions. International Journal of Computer Applications, 159(??): 25–32, February 2017. CODEN ???? ISSN 0975-

8887.  URL https://www.ijcaonline.org/archives/
volume159/number4/26991-2017912906/.

Yadav:2011:CAS

[YS11] Vedpal Yadav and Alka Sharma. CARVER+Shock (agriculture) — software for agriculturists to fight against bioterrorism. International Journal of Computer Applications, 24(??):5–10, June 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume24/
number8/2973-2776/.

Yousef:2012:OAV

[YSH12] Marwa T. Yousef, Elsayed M. Saad, and Shahira M. Habashy. Obstacle avoidance with virtual sensor in mobile robot’s motion using the advanced potential field controller. International Journal of Computer Applications, 53(??):9– 14, September 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume53/
number4/8407-2032/.

Yadav:2015:CPC

[YT15] Nalini Yadav and Sudeep D. Thepade. Comprehensive performance comparison of Fourier, Walsh, Haar, sine and cosine transforms for video retrieval with partial coefficients of transformed video. International Journal of Computer Applications, 120(??):38–43, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume120/number19/21339-4345/.

Yadav:2012:FCD

[YWT12] Anamika Yadav, Prarthana Walayani, and A. S. Thoke. Fault classification, distance location and faulty section identification in teed transmission circuits using artificial neural network. International Journal of Computer Applications, 47 (??):19–25, June 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume47/
number15/7264-0234/.

Yang:2023:PBG

[YYS^{+}23a] Jianjun Yang, Mingyuan Yan, Abi Salimi, Bryson Payne, Jason Porter, and Janet Morales. Parabola based geographic forwarding in wireless sensor networks. International Journal of Computer Applications, 185(??):1–8, August 2023. CODEN ???? ISSN 0975-8887. URL https:

//www.ijcaonline.org/archives/volume185/number30/
32881-2023923054/.

Yang:2023:MVT

[YYS^{+}23b] Jianjun Yang, Mingyuan Yan, Abi Salimi, Jason Porter, Ying Luo, and Ju Shen. A mechanism on video tracking and recognition with fast transmission in wireless communications. International Journal of Computer Applications, 185(??):1–8, June 2023. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume185/number14/32761-2023922816/.

Ziada:2013:DDS

[ZAA13] Maha E. M. Ziada, Ali Mohammed Abderahman Ajoub, and Elsedige A. Aljaak. Development of a decision support system for irrigation management at Gezira Scheme in Sudan. International Journal of Computer Applications, 65 (??):32–36, March 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume65/
number4/10916-5845/.

Zayed:2018:DPM

[ZAAS18] M. Mokhtar Zayed, Wagdy R. Anis, Mohamed Abouelatta, and A. I. Salem. Design of PV modules including a layer between solar cells and glass cover to increase PV module lifetime. International Journal of Computer Applications, 182(??):18–23, October 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume182/number19/30041-2018917950/.

Zaeri:2013:BLS

[Zae13] Naser Zaeri. Blended learning system performance evaluation. International Journal of Computer Applications, 76 (??):33–39, August 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume76/
number4/13237-0673/.

Zaytar:2018:AAQ

[ZAE18] Mohamed Akram Zaytar, Mohamed Amrani, and Chaker El Amrani. Air2Day: an air quality monitoring adviser in Morocco. International Journal of Computer Applications, 181(??):1–6, September 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume181/number17/29911-2018917851/.

Zhar:2012:SDD

[ZAER12] Noureddine Zhar, Mohamed Ait Ali, Mohsine Eleuldj, and Amine Raji. A specific-domain design tool for FPGA-based image and video processing system. International Journal of Computer Applications, 56(??):16– 21, October 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume56/
number11/8935-3070/.

Zafar:2011:PSC

[Zaf11] Saima Zafar. Post scan correction of step, linear and spiral motion effects in CT scans. International Journal of Computer Applications, 35(??):13–19, December 2011. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume35/number10/4436-6182/.

Zarad:2016:SEA

[ZAHH16] Mohamed M. Zarad, Ahmed A. Abdel-Hafez, and Ahmed H. Hassanein. Secure and efficient authentication scheme for cloud computing. International Journal of Computer Applications, 142(??):19–24, May 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume142/number8/24916-2016909883/.

Zarri:2024:UGW

[Zar24] Gian Piero Zarri. Using the generalized world entities (GWEs) paradigm in a Semantic Web of Things (SWoT) context. International Journal of Computer Applications, 186(??):32–40, January 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number4/33063-2024923392/.

Zhao:2014:JAM

[ZBI14] Xiao Zhao, Elyes Bdira, and Mohamed Ibnkahla. Joint adaptive modulation and adaptive MAC protocols for wireless sensor networks. International Journal of Computer Applications, 85(??):32–39, January 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume85/number19/15098-3492/.

Zhu:2024:ERM

[ZC24] Ao Zhu and Peng Cao. An efficient restoration method for the faint dot matrix images. International Journal of Computer Applications, 186(??):1–10, October 2024. CODEN ???? ISSN 0975-8887. URL https://

www.ijcaonline.org/archives/volume186/number45/an-
efficient-restoration-method-for-the-faint-dot-matrix-
images/.

Zoby:2013:HNA

[ZCDR13] L. T. M. Zoby, M. P. Carvalho, J. C. Da Costa, and L. B. Ruiz. HACBA: New approach to hierarchical routing for wireless sensor network. International Journal of Computer Applications, 64(??):38–42, February 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume64/number20/10753-5758/.

Zhou:2024:DMS

[ZCZ24] Xingyu Zhou, Peng Cao, and Xia Zhang. Design of multi-spectral anti-counterfeit image acquisition and detection system. International Journal of Computer Applications, 186(??):25–32, January 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number2/33045-2024923348/.

Zol:2015:CSM

[ZD15] Smita Haribhau Zol and R. R. Deshmukh. A comparative study of MRI image segmentation based on fast kernel clustering analysis. International Journal of Computer Applications, 110(??):26–29, January 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume110/number6/19321-0890/.

Zaghloul:2017:EPE

[ZEB17] Zaghloul Saad Zaghloul, Nelly Elsayed, and Magdy Bayoumi. Early prediction of epilepsy seizures system based on artificial immune BCI system. International Journal of Computer Applications, 169(??):35–43, July 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume169/number9/28016-2017914881/.

Zaied:2015:EHR

[ZEE15] Abdel Nasser H. Zaied, Mohammed Elmogy, and Seham Abd Elkader. Electronic health records: Applications, techniques and challenges. International Journal of Computer Applications, 119(??):38–49, June 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume119/number14/21139-4153/.

Zeineldin:2017:FFR

[ZEF17] Ramy Ashraf Zeineldin and Nawal Ahmed El-Fishawy. FRANSAC: Fast RANdom sample consensus for 3D plane segmentation. International Journal of Computer Applications, 167(??):30–36, June 2017. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume167/number13/27832-2017914558/.

Zaid:2013:PES

[ZEfNF13] Osama M. Abu Zaid, Nawal A. El-fishawy, E. M. Nigm, and Osama S. Faragallah. A proposed encryption scheme based on Henon chaotic system (PESH) for image security. International Journal of Computer Applications, 61 (??):29–39, January 2013. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume61/
number5/9925-4549/.

Zeid:2011:TIP

[ZGAS^{+}11] Amir Zeid, Hani Ghalayini, Mohammad Al-Sulaiman, Ghanim Al-Ghanim, Abdulla Al-Jassim, and Abdulrahman Al-Failakawi. Towards IDEs for physically challenged programmers. International Journal of Computer Applications, 18(??):4–9, March 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume18/
number7/2299-1944/.

Zhu:2021:PBP

[Zhu21] Yongwen Zhu. The plum-blossom product method of large digit multiplication and its application to computer science. International Journal of Computer Applications, 183(??):17–23, December 2021. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume183/number41/32202-2021921805/.

Zainudin:2016:LEG

[ZHUR16] Dyana Zainudin, Thaier Hamid, and Atta Ur-Rahman. Leadership by example in e-government security management system. International Journal of Computer Applications, 144(??):10–17, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number4/25166-2016910239/.

Zia:2019:UCM

[ZJR^{+}19] Muhammad Bilal Zia, Zhao Juan Juan, Zia Ur Rehman, Kamran Javed, Saad Abdul Rauf, and Arooj Khan. The utilization of consignable multi-model in detection and classification of pulmonary nodules. International Journal of Computer Applications, 177(??):23–28, December 2019. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume177/number27/31070-2019919741/.

Zina:2012:DOR

[ZK12] Nakhla Zina and Nouira Kaouther. Development of ontology for the representation of adverse drug events of diabetes disease. International Journal of Computer Applications, 42(??):10–16, March 2012. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume42/number20/5815-8123/.

Zolfy:2010:OTB

[ZKN10] Mina Zolfy, Ziaddin Daie Koozehkanani, and Zainalabedin Navabi. An OSCI TLM-2.0 based platform for grid computing simulation. International Journal of Computer Applications, 3(??):7–13, June 2010. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume3/number8/752-1063/.

Zahid:2014:NSF

[ZMB14] M. Zahid, A. Mezrioui, and A. Belmekki. A new scalable framework for emulating huge networks. International Journal of Computer Applications, 88(??):11– 17, February 2014. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume88/
number6/15355-3815/.

Zongur:2017:HAS

[Zon17] Aykut Zongur. High-accuracy serial pattern detection system with optimized annealing. International Journal of Computer Applications, 176(??):26–32, October 2017. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume176/number2/28527-2017915543/.

Zyout:2011:CMC

[ZPAQP11] Imad Zyout, PhD, Ikhlas Abdel-Qader, and PE PhD. Classification of microcalcification clusters via PSO-KNN heuristic parameter selection and GLCM features. International Journal of Computer Applications, 31(??):34– 39, October 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume31/
number2/3798-5235/.

Zakary:2015:EOA

[ZRE15] Omar Zakary, Mostafa Rachik, and Ilias Elmouki. On effectiveness of an optimal antiviral bitherapy in HBV-HDV coinfection model. International Journal of Computer Applications, 127(??):1–10, October 2015. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume127/number12/22778-2015906554/.

Zakary:2016:IAP

[ZRE16] Omar Zakary, Mostafa Rachik, and Ilias Elmouki. On the impact of awareness programs in HIV/AIDS prevention: an SIR model with optimal control. International Journal of Computer Applications, 133(??):1–6, January 2016. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume133/number9/23811-2016908030/.

Zulfadhilah:2016:LCU

[ZRP16] Muhammad Zulfadhilah, Imam Riadi, and Yudi Prayudi. Log classification using K-means clustering for identify Internet user behaviors. International Journal of Computer Applications, 154(??):34–39, November 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume154/number3/26475-2016912076/.

Zanganeh:2011:AAN

[ZRZ11] Tayebeh Zanganeh, Meysam Rabiee, and Masoud Zarei. Applying adaptive neuro-fuzzy model for bankruptcy prediction. International Journal of Computer Applications, 20 (??):15–21, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume20/
number3/2415-3229/.

Zouhri:2014:STG

[ZSR14] S. Zouhri, S. Saadi, and M. Rachik. Simulating the tumor growth with cellular automata models. International Journal of Computer Applications, 108(??):5–11, December 2014. CODEN ???? ISSN 0975-8887. URL https://www.

ijcaonline.org/archives/volume108/number14/18977-
0392/.

Zougagh:2014:MOP

[ZTL^{+}14] H. Zougagh, A. Toumanari, R. Latif, Y. Elmourabit, and Noureddine Idboufker. Modified OLSR protocol for detection and prevention of packet dropping attack in MANET. International Journal of Computer Applications, 100(??):32–38, August 2014. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume100/number17/17617-8254/.

Zghal:2011:IIW

[ZTM11] Nadia Smaoui Zghal, Khaled Taouil, and Dorra Sellami Masmoudi. Implementation of an improved watershed algorithm in a virtex 5 platform. International Journal of Computer Applications, 21(??):40–45, May 2011. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume21/number1/2472-3326/.

Zahoor:2016:CPS

[ZUDS16] Zarka Zahoor, Moin Ud-Din, and Karuna Sunami. Challenges in privacy and security in banking sector and related countermeasures. International Journal of Computer Applications, 144(??):24–35, June 2016. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume144/number3/25161-2016910173/.

Zali-vargahan:2012:CTI

[ZvAS12] Behrooz Zali-vargahan, Mehdi Chehel Amirani, and Hadi Seyedarabi. Contourlet transform for iris image segmentation. International Journal of Computer Applications, 60 (??):41–44, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume60/
number10/9732-4209/.

Zhu:2021:MID

[ZYL^{+}21] Xiaoming Zhu, Lijun Yao, Fan Luo, Kejun Wang, Zhou Che, Jing Yan, Min Zhou, Yongchang Cai, Lingling Wang, Zelong Cao, Lan Peng, Fengqing Bai, Zifang You, Hongqiu Xiao, and Haocheng Qi. Motion image deblurring using AS-cycle generative adversarial network. International Journal of Computer Applications, 183(??):32–37, November 2021. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume183/number39/32190-2021921768/.

Zin:2013:DRC

[ZYM13] Zalhan Mohd Zin, Rubiyah Yusof, and Ehsan Mesbahi. Dimension reduction and clustering of high dimensional data using auto-associative neural networks. International Journal of Computer Applications, 72(??):31–37, June 2013. CODEN

???? ISSN 0975-8887. URL https://www.ijcaonline.org/
archives/volume72/number11/12540-9090/.

Zyout:2012:CCM

[Zyo12] Imad Zyout. Classification of clustered microcalcifications in mammograms using particle swarm optimization and least-squares support vector machine. International Journal of Computer Applications, 59(??):23– 28, December 2012. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume59/
number17/9640-4349/.

Zhang:2018:HSA

[ZYW^{+}18] Yunxiang Zhang, Xiaokun Yang, Lei Wu, Archit Gajjar, and Han He. Hierarchical synthesis of approximate multiplier design for field-programmable gate arrays (FPGA)- CSRmesh system. International Journal of Computer Applications, 180(??):1–7, February 2018. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume180/number17/29021-2018916380/.

Zou:2011:ANG

[ZZ11] Xianlin Zou and Qingsheng Zhu. Adaptive neighborhood graph for LTSA learning algorithm without free-parameter. International Journal of Computer Applications, 19(??): 28–33, April 2011. CODEN ???? ISSN 0975-8887.

URL https://www.ijcaonline.org/archives/volume19/
number4/2348-3070/.

Zhang:2024:GAB

[ZZW24] Wei Zhang, Qingtian Zhang, and Huaqiang Wu. Genetic algorithm-based training method for memristor neural networks. International Journal of Computer Applications, 186(??):7–13, November 2024. CODEN ???? ISSN

0975-8887. URL https://www.ijcaonline.org/archives/
volume186/number49/genetic-algorithm-based-training-
method-for-memristor-neural-networks/.